A Bibliography of Publications about the Computation of Elementary Functions in Computer Programming Languages

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

18 October 2024 Version 5.299

Title word cross-reference

#3506 [Fet81b, Köl97]. #6725 [Fet81a].

(0, 2) [GRS87]. (1 + 2x) exp(x^{2})(x) [SL81]. (a, b) ← (\frac{a+3b}{4}, \frac{\sqrt{ab}+b}{2}) [BB89]. (η, π) [BDG^{+}84]. (η, ρ) [BDG^{+}72]. (M, p, k) [BEM12, WMBE14]. (n, 1) [Kra85b]. (n, 2) [Kra85a]. (ψ(x) − ψ(y >))^{2} [De 91]. −1/2 [CT67a, CT67b]. −∞ < N < +∞ [Kog57, Kog58b]. 0 [Col80, Col84]. 0 < N < 1 [Kog58a]. 0 ≤ x < ∞ [SL81]. 0 ≤ x ≤ 1 [Ano48]. 1 [AB93, Her61c, NDT69, Pas92]. 1/2 [CT67a, CT67b, Fuk15d]. 1/9 [Mag94]. 1/n [Boy09]. 1/\sqrt{x} [TBDS11]. 10 [Col80, Col84, NDT69]. 1271 [Knu62a]. $140.00 [Lip21]. 2 [Bak73, Bak75a, Bas01, BJK^{+}11, CL94a, CP99, Her61c, LO95, MIT^{+}95, MC93, MC95, PZ95]. 29, 360, 000 [Bai88]. 2^{p} [Mer94]. 3 [Lin79a, NKY08]. 3/2 [CT67a, CT67b]. 38 [Gus78a]. 3j [SG76]. 4 [BH01, EL90, LA01, LA03, NL99]. $46.36 [Lip21]. 6j [SG76]. 8 [HF95, PZ95]. 9 [Wei98]. [−1, 1] [HR88]. [0, ∞) [CMV69]. ^{2} [Sel93]. 0F_{3} [Lar69]. _{1}F_{1}(a, b, z) [AG08]. _{2}F_{1}(a, b; c; x) [TC01]. _{2}F_{1}(a, b; c; z) [BS00]. 2F_{2}(1, α; ρ_{1}, ρ_{2}; z) [Kim72]. _{3}F_{2} [Büh87]. _{n} [LDP93]. _{p}F_{p−1}(a; b; z) [BS06]. pF_{p−1}(z) [Sko04]. A [IP81, AG08]. a, b [DJ67]. Ai(x) [MPG92a, MPG92b]. B [Wen09, AG08, NPA16]. c [FLS06]. χ^{2} [Fet79]. cos [Tur69]. cos(iπ/n) [Oli75]. cos(x)/x [Gau66a, GK70b]. cos^{−1} [MMM93]. cosh z = az [Fet76]. cos N [Kog59]. E [Cod66, FC64, FC64, Cod65a, Yar67]. E(m) [GMMIDM13]. e^{−x} [Bar72, CMV69]. e^{−z} [Sie76]. e^{N} [Kog57]. e^{x} [Bar72, ECSD20, Mar86]. e^{x}/x [Gau66a, GK70b]. e^{z}E_{1}(z) [Has02]. E_{1}(x) [CT68]. E_{1}(x), x > 0 [SZ77]. E_{n}(x) [SZ74]. E_{ν}(x) [CLM88, NP18a]. ϵ [Bre78a, Kal04, KWY07a, KWY07b, Mat92]. η [BFSG74, BFSG84]. exp(Bt) [PN85]. exp(x) [AJMT81]. F [Dor68]. F (a; b; c; z) [FLS06]. f(∞) [BB94]. F (n/2, 1/2, x) [Uni49]. f(x) [BB94]. F_{0}(x) [SC71]. F_{1} [EG97]. F_{1/2}(x) [WR63]. F_{2} [ABF^{+}23, OSS09]. F_{3}^{m} [AH10]. F_{a}(α; β, β′; γ; x, y) [TS76]. F_{c} [BK16]. F_{j}(x) [Let01]. F_{λ}(η, x) [Bar82a]. F_{m}(x) [CP98]. \frac{2}{π} ∫_{0}^{∞} (\frac{sin t}{t})^{n}dt

[HKW60]. \frac{∂}{∂θ} P_{μ}^{1}(cos θ) [Wil68]. G [Lev91a, BS13, SS10]. g_{3/2} [Mac97]. g_{5/2} [Mac97]. G_{λ}(η, x) [Bar82a]. Γ [BB82, BF92, Sha76]. Γ_{3} [CS11]. H [KS04]. H_{1} [AJ03]. H_{n}^{(1)}(z) [AM79, AM84a]. H_{n}^{(2)}(z) [AM79, AM84a]. Ei(x) [CT69]. Ki_{n}(x) [Amo83a]. I [GS81, Soo73a]. I_{0} [Hil77]. I_{0}(x) [Bla74]. I_{0}(ξ) [Las82]. I_{1}(x) [Bla74]. I_{n}(b) = 2π^{−1} ∫_{0}^{∞} (\frac{sinx}{x})^{n} cos(bx)dx [Tho66].

I_{n}(b) = \frac{2}{π} ∫_{0}^{∞} (\frac{sin x}{x})^{n} cos bx dx [Fet67]. I_{n}(b) = \frac{2}{π} ∫_{0}^{∞} (\frac{sin x}{x})^{n} cos(bx)dx

[MR65]. I_{n}(X) [Leh44]. I_{n}(x + jy) [Sca71]. I_{ν}(x) [ADW77a, ADW77b, Amo78]. I nu(z) [Tho04, Cam81, Cam84a, TB87]. I_{x}(a, b) [YA24]. ∫_{0}^{∞} ^{\frac{t^{2α−1}J_{ν}(x\sqrt{1+t^{2}})(1+t^{2})^{α+β−1}}{d}} t [Sch78a]. ∫_{0}^{∞} t^{2α−1}(1 + t^{2})^{1−α−β}J_{ν}(x\sqrt{1 + t^{2}})dt [Gla79]. ∫_{x}^{∞} e^{−t^{2}/2}t^{p}dt, x > 0, p [DiD78]. ∫_{x}^{∞} J_{0}(t)t^{−n}dt [Lon59]. ∫_{x}^{∞} J_{1}(t)t^{−n}dt [Lon59]. ∫_{0}^{π/2} J_{0}^{2}(γ sin x)dx [SF87]. ∫_{0}^{π/2} J_{ν}^{2}(λ cos θ)dθ [Won88]. ∫_{0}^{h} e^{x^{2}} dx [Daw97]. ∫_{0}^{∞} exp(−ct) dt/\sqrt{t(1 + t^{2})} [Act74]. ∫_{0}^{∞} t^{n} exp(−t^{2} − x/t) dt [CP79]. ∫_{0}^{z} − log |1 − y|/y dy [Mit49]. ∫_{1}^{∞} e^{−xt}t^{−n} dt [Gau59]. ∫ e^{b(x+a cosx)} dx [Rut51, Rut51]. J [Gun65b, Gun65c, Soo73a, JF16, Wei98]. j = −3/2 [Let01]. J′_{l}(x) [Bea58]. J_{−ν}(x) [CK79]. J_{0} [AM77, AM84c]. J_{0}(x) [GM88, Let96, MG85, Sch64, WBR82]. J_{1} [AM77, AM84c]. J_{1}(x) [GM88, WBR82]. j_{l}(r) [Del79, Del84]. J_{l}(x) [Bea58]. j_{n} [AM78b, AM84b]. J_{n}(x) [Col80, Col84]. J_{n}(x + jy) [Sca71]. J_{n}(z) [AM79, AM84a, dT93, CM83]. J_{ν}(x) [ADW77a, ADW77b, Amo78, BGV81, Cam79, Cam84b, MG89]. J_{ν}(z) [ES99]. J_{v}(x) [LW96]. Je_{0}(a, z) [DK90]. K [Cod66, FC64, Lau73, FC64, BST13, Cod65a, GS81, Bas01, Lem88, LC87]. K_{0}(x) [Ano48]. K_{0}(z) [Bur74]. K_{1}(x) [Ano48]. K_{1}(z) [Bur74]. K_{iζ}(x) [BO75]. K_{n}(Z) [Par72, DC81, LG79]. K_{ν}(t)/I_{ν}(t) [LL72]. K_{ν}(z) [Cam81, Cam84a, TB87, Tho04]. K_{p}(x) [Aha69]. Ki_{n}(x) [BEJ78]. Ki_{s}(x)

[GP64]. L [ZSA86, BB15, Cra12a]. L^{∞} [BF23]. l_{2} [GLT^{+}15]. L_{p} [Hei88, Ubh89, JMS98]. λ [Bar84b]. ln(x) [MTP82]. log_{e}^{2}(1 + x) [MM90a]. log_{2} x [KP75]. log n [KR79]. log Z [Luk57]. M [Kog59, SK11]. M(a, b, x) [Mul01]. m^{2} + y^{2}n^{2} [AB93]. M_{k,1/2} [Lau73]. GF(2^{m}) [RHMLL08]. μ [Kun81]. N [CH89a, Kog59, BH99, CH89b, Hig93, CL00, Gau59, GLT^{+}15, Har14, Lev91b, Rei69]. ν [CH70a, CK79]. O(N^{2}/ log^{3/2N}) [SR81]. O(n^{3}L) [Gon89]. arc tan Z [Luk57]. bei [Tha62a]. ber [Tha62a]. cos x [Bar72, Bar72]. Ei(x) [Ng70, Pac70, Red70, TM68]. inverfcθ [Phi60]. modp [Sch85]. Shi(x) [TM68]. sin x [Bar72, Bar72]. p [ADN03, CL11, Tor02, Tur94]. P_{ν}^{1}(cos θ) [Wil68]. p_{1/2+iτ}^{m} (x) [GST12b]. P_{n}(x) [LW95]. ϕ [Pro88]. π [Bai88, BBB89, CDR71, Sal76a, Sha76, Yar67]. ψ [Gla66]. ψ(x) [DR98]. ψ(x) − ψ(y >) [De 91]. ψ(x) − θ(x) [Per85b, Per87]. ψ(x; 3, l) [McC84]. ψ(z) [McC81]. ψ^{(k)}(x) [Köl96]. Q [BTT16, CB09, CD02, CPG12, Dev12, DB13, DD08, Fre09, GST14a, IR08, Jan11a, Jan11b, KL07, LKF10, LBC11, MJ13, PUHM12, SK11, SF10, WLK11, Bor88, CIM94, GA08, Kra95, SJ12]. Q(x) [BS79]. QD [AN68]. R [JKD92]. R^{2} [Bos89]. ρ [BFSG74, BFSG84]. S_{n} [SM77, SM78]. sin [Tur69]. sin(x)/x [Gau66a, GK70b]. sin 1^{∘} [Aab54]. sin^{−1} [MMM93]. sin N [Kog59]. sin z = az, cos z = az [Fet76]. \sqrt{(}x) [And82]. \sqrt{(x/d)} [ALB98]. \sqrt{6} [AB93]. \sqrt{x} [Kah80, Mou67]. \sqrt{x^{2} + y^{2}} [BJK^{+}11]. ∑_{n=1}^{∞} n^{p−1}E_{n}^{∗}(f) < ∞ [HS83]. t [Hil70a, Hil70b, Hil73a, Hil81a, Hil81b, eL79]. Θ [Fdi97]. θ(x; 3, l) [McC84]. θ_{3}(q) [WV20]. U(a, b, z) [Tem83a]. U(a, b, z) [GST15a]. U(a, x) [GST06a, GST06b, SGA81]. V (a, x) [GST06a, GST06b, RL76, SGA81]. W [Gau11, Goe14, JHC96, Jef17, Joh20a, MS16, Miy19, Pak18, Veb12]. W (a, ±x) [SGA81]. W (a, x) [GST11a, GST11b]. W_{k,1/2} [Lau73]. W_{k,m}(z) [Won73a, WR71]. ℘ [Eck76, Sou57, T.57]. ℘(z|e_{1}, e_{2}, e_{3}) [Tod90]. X [HJ67b, De 91]. x = 0 [BB94, Mid75]. x = 1/4 [Köl96]. x = 3/4 [Köl96]. x^{n} [McC77, NL94]. x^{x−1} [Mor14]. xC′_{ν}(x) + γC_{ν}(x) = 0 [GS12]. x exp(iπ/4) [Wal84]. x ≥ 0, ν ≥ 0 [ADW77a, ADW77b, Amo78]. y [AB93, De 91]. Y_{0}(x) [Ano48, WBR82]. Y_{0}(z) [ZB95, Zha96a]. Y_{1}(x) [Ano48, WBR82]. Y_{1}(z) [Zha95, Zha96a]. y_{n} [AM78b, AM84b]. Y_{n}(z) [AM79, AM84a, ZB97, dT93]. Y_{ν}(x) [Cam79]. Z [Som82, NPA16]. z = 1 [BS06, Sko04]. Z^{2} [Lef05]. Zα [FR76]. Zα^{2} [FR76]. ζ [Lee97]. ζ(α, s) [BS06]. |x|^{α} [GS91].

-acceptability [IP81]. -algorithm [Bre78a, Mat92]. -analogues [GA08]. -Approximations [BF23]. -Bessel [BST13]. -binomial [Kra95]. -continued [Lev91a]. -D [BJK^{+}11]. -DIIS [Sel93]. -discrepancy [Pro88]. -Distribution [Hil70a, Fet79, Hil81a, eL79, Dor68]. -elementary [Bor88]. -expansion [Kal04, KWY07a, KWY07b]. -Exploring [Sho02]. -fractions [Lev91b]. -Friendly [WMBE14, BEM12]. -Function [DD08, Eck76, Gla66, Jan11a, BTT16, GST14a, IR08, Jan11b, KL07, LBC11, SK11, WLK11, BF92, CB09, CD02, CPG12, Dev12, DB13, Fre09, LKF10, MJ13, PUHM12]. -functions [BS13, BB82, Gau11, SF10]. -gamma [GA08]. -interpolation [GRS87]. -Laguerre [CIM94]. -norm [JMS98]. -Norms [GLT^{+}15]. -periodic [Lem88]. -plane [BDG^{+}84]. -Quantiles [Hil70b, Hil81b]. -saturation [Hei88]. -select [CL00]. -series [BB15, Cra12a]. -spaces [Som82]. -Spline [Lin79a]. -th [Har14, CH89b]. -transforms [KS04]. -type [Fdi97]. -Vectors [GLT^{+}15]. -Zeros [CH70a, CK79]. -Zeta [SJ12].

1 [Frö63a, Lov53]. 1006 [AM20]. 1014 [Bor21a]. 10th [KM91]. 11th [SIJ93]. 121 [Laz85]. 121-122 [Laz85]. 122 [Laz85]. 123 [HJ67b]. 12th [KM95]. 13 [Joh87b]. 135 [Ano84]. 13th [LMT97]. 14 [Bea60]. 144 [ACG^{+}87]. 149 [Mer62, Tha63b]. 14th [KK99]. 15th [BC01]. 16 [Wim72]. 1604 [Mae60]. 163 [Fet63]. 165 [Far69, Tha63a]. 167 [PZ95]. 16BST [Cat85]. 16th [VDR05]. 179 [BB74, Lud63, PH67]. 17th [IEE05]. 18-19 [IEE81a]. 180 [HJ67b]. 181 [HJ67b]. 18th [KM07]. 191 [Kop74, Rel63a, Rel63b]. 192 [Rel63c]. 1958 [Lan59]. 1975 [Ask75, Tra76]. 1976 [Cow77]. 1981 [Mul82]. 1983/84 [Laz85]. 1985 [Mar85, MC87]. 1988 [MC90]. 1989 [MMH89]. 1990 [BL91]. 1993 [BCEP94, Zah94]. 1996 [Boi97]. 1999 [CDS00, DIW00, SC01]. 19th [BCDH09].

2 [IM99a, MLS^{+}18]. 2001 [BC01]. 2002 [Bor02, KV03]. 2005 [HUY07, IL07]. 2006 [Ano06, Men06]. 2008 [IEE08, Sht19]. 2010 [ISO10]. 2013 [IEE13]. 2015 [MTV15]. 2016 [MSH^{+}16]. 2017 [BBdD17]. 2019 [TBL19]. 2020 [CLT20]. 2021 [IEE21]. 2022 [IEE22]. 2023 [IEE23]. 209 [HJ67b, Ibb63]. 20th [Bre00, SO11]. 213 [Gra63a, Gra64]. 21st [IEE13]. 221 [Gau64b]. 222 [Gau64c, Gau64d]. 225 [Bra64, CC64]. 226 [Cyv64, HJ67b]. 22nd [MTV15]. 236 [Gau64a, Gau65b]. 23nd [MSH^{+}16]. 244 [LG64]. 245 [Mac89]. 24th [BBdD17]. 25 [Hen77, Sha76]. 252 [Gun65a]. 259 [Gau65a]. 25th [TT18]. 260 [Gun65b]. 261 [Gun65c]. 26th [TBL19]. 272 [HJ67b, Mac65, Mac68]. 27th [CLT20]. 282 [Gau66a]. 28th [IEE21]. 291 [PH66]. 292 [Gau66b, Gau69b, Köl69a]. 299 [HP67, OW68]. 29th [IEE22].

30 [ACG^{+}87]. 30-ns [MIT^{+}95]. 300 [Gun67, Köl69b, Köl69c, Vos73]. 301 [BP67, Pit67]. 304 [Ber68, HJ67a, HJ67b]. 30th [IEE89b, IEE23]. 312 [Fri67]. 32 [ASA06]. 32-bit [LN21]. 322 [Dor68]. 327 [Köl68]. 332 [Sko75b, Wit68]. 33rd [Spr00]. 34 [Fra81]. 349 [TS69]. 352 [Cle69]. 360 [Fik66, Kuk66]. 3600 [Cod64]. 363 [Gau69a, Köl72a]. 370 [ACG^{+}86]. 38 [Bin68]. 3800 [WOG95]. 385 [Pac70]. 392 [Fri72, SM70]. 395 [Hil70a]. 396 [Hil70b]. 3DNow [IM99a].

40 [Fet81a]. 404 [LT71]. 421 [Kuk72a]. 43 [Fet81b, Köl97]. 435 [Ful72, Sch78b]. 44 [Per87, Woj61]. 442 [HD73]. 44th [Men06]. 45 [Lew87]. 46 [Her61b]. 465 [Hil73a]. 47 [Cob69, Tho04]. 471 [Gau73]. 484 [Bur74]. 487 [Pom74]. 488 [Bre74]. 49 [Her61a]. 490 [GZ75, Mor76]. 498 [Pri75]. 4th [HY14, IEE78].

50th [Gau94]. 511 [ADW77a, Amo78]. 518 [Hil77]. 54 [Her61c]. 542 [Gau79a]. 549 [Eck80]. 55 [Her61d, Tha63c]. 55-b [MIT^{+}95]. 56 [Her61e, Lar66]. 57 [Her61f, Tha62a]. 577 [CN81]. 585 [Bre82a]. 599pp [Köl90]. 5th [Ano03, Gre16, IEE81a, IEE81b, IL07].

6 [Gun65b]. 6- [Gun65b]. 60 [Tem79a]. 6000 [Mar90]. 602 [Gel51]. 602-A [Gel51]. 609 [Amo83a]. 610 [Amo83b]. 64 [HKST99, Mar00, ST99]. 644 [Amo86, Amo90, Amo95, Kod07]. 650 [Joh87a]. 654 [DM87]. 66 [Hil73b]. 680 [PW90a]. 69-0172 [Fet81a]. 6th [FBM04].

708 [BL94, DM92]. 714 [Cod93a]. 723 [And00, Sny93, Sny96]. 7289 [FR98]. 73 [Jef62, Mey63, vdR63]. 736 [DR94a]. 745 [Goa95, Goa97]. 754 [AHM^{+}98, AAHTH10, BT99, DR04, Hou81]. 754r [LSZ08]. 757 [Mac96]. 77 [GST02b, WNO94]. 78 [Has66]. 794 [Wie99]. 7th [Ano06, BCD^{+}20, Hwa85].

8008/8080 [Arn77]. 804 [Alh00a]. 814 [Smi01]. 819 [GST02b]. 81f [Fra81]. 822 [GST02c]. 831 [GST04a]. 838 [Fab04]. 850 [GST06a]. 858 [VC06a]. 86k [Per87]. 86m [Lew87]. 87 [Alz98]. 877 [Kod08]. 88 [Cun62a]. 885 [Lin09]. ’89 [IEE89c, ACM89, Cun62b]. 895 [BC09].

9 [Ano84, Gun65c]. 9- [Gun65c]. ’90 [Koo91, Cun62c, DG99, EC13, GST12b, SSG^{+}18, Smi01, Tho97a]. 903 [CZ10]. ’91 [IEE91, Koo91]. 910 [Kor11]. 911 [Smi11]. 912 [Kod11]. 914 [GST11a]. 916 [ZA11]. 92 [VRS^{+}99]. 926 [Tho13]. 934 [EC13]. 935 [RKZ^{+}14]. 939 [GST14a]. 954 [Flo15]. ’96 [Lak96, RDK98]. 969 [GRAST16]. 978-0-367-40820-6 [Lip21]. 9th [ES89].

A-Priori [Krä87]. A. [Leh83b]. abc} [CLZ04]. Ableitungen [Töl68]. Abramowitz [Dav59, GJL20]. Abründungsfestigkeit [Ost37]. Absolute [MM91b, Mat90a, Fri67]. abstract [Ost54]. Abstraction [KI16]. abstracts [Spr00]. Academy [Mil75a]. Accelerated [HR05, DGB82]. Accelerating [BMR04, Vep08]. Acceleration [Alb62, Boy09, Bra65, Bre82b, CRZ00, DGS65, FFS83, Gus66, HBS00, Hom96a, Hom96b, Hom98a, IBM05, KP03, LB90, Osa90, Ovt03, Pas03, Sel93, SF79, Tro84, UE05, Wil12, Wyn72, Bre77, Che88, Del81, AGJ86, Ano67, BD85, Bel90, Bre77, Bre78c, Bre78a, Bre78b, BDGB83, Bre83, Bre85, Bre88, Bre99, Bre00, BCG91, Che88, Cro92, DK82, DGB80, Del81, Fdi97, GB90, Gus78a, Gus78b, Gus84, Gus85, GS91, HP93, Har77, Hau88, Hom94, Hom98b, Hom99, Ise91, Ise94, JW88, Joh83, Kru99, Kza92, Kza99, Lem88, LP88, Lev91b, Lev91a, LP95, Lon81, Mat90a, Mat92, Mül93, Mül99, Now06, Ole96, Osa94, Pas92, Pas08, Pas10, Pas11, RdL01, Sch68, Sed90, SL83, SS98, Wen89, Wen01, DGB80]. Accelerator [MLS^{+}18]. accelerators [Sid06]. acceptability [IP81]. according [Töl68]. account [Mit49]. Accuracy [AMM21, BBM84, BHK09, Gal85, GB68, Ike76, LDP93, Sal51b, Sch96, SF59, WMC21, Zag19, Bar81, Bar84b, Bra64, CC64, Gal86, Jab20, Krä88, Loe01, Sch80b, TB87, WMC18, Zag24b, Bra87, Rot71, von80, Mül90]. Accurate [AM20, AFC10, ASA06, AM79, AM84a, Bee91, Bin68, CG24, DB13, FB91, Flo15, FR76, GB91, GST12a, GF88, HT13, Har09b, How22, JGB13, JF16, KB90, LBC11, LJ13, Mac97, MNP83, MT85, MT86, SS97b, SS97a, SK11, Sho05, SZ05, SS99b, Tan90a, Tan90c, Van89, VLCSFN^{+}12, Zag17, ZOHR01, von84, AM12, AIS^{+}17, BCK^{+}93, Che16d, DMMM95, Di 86, GST11b, Lut95, Mit92, MN16, NP18a, PKS00, PN85, SZ77, SS10, SS23, Tan91a]. Achieving [BTDS11, Gal85, Gal86]. Ackermann [Sun71]. Acknowledgment [Wis48]. ACL2 [Bor02]. ACL2-2002 [Bor02]. ACM [ACM89, Joh87b, Koo91, Bre74, Bur74, Ful72, Gau64a, Gau73, GZ75, Hil70a, Hil70b, HD73, Hil73a, Kuk72a, LT71, Mer62, Mor64, Pom74, Rel63a, Wit68]. ad [Yun14]. Ada [ACM87, DPC95, Dri89, Dri91d, Dri91a, Dri91b, Dri91c, Mat87, Squ91c, Squ91a, Squ91d, Squ91b, Tan90c, Tan91a]. Adaptive [BM04, ZCL15, Was89]. Add [BDL09]. addenda [Luk56b]. Addendum [Fra81]. adders [CP99]. Addition [SS97b, SS99b, MC91]. Additional [LM93]. Additions [LJ13, ITY96]. adds [Bic81]. Adelaide [KK99]. Ado [Kah87]. advance [BMR04]. Advanced [Luk99, Ask75, CDS00]. Advances [CD08, IEE99, VK95]. Aerospace [Fet81a]. afficionado [BCR^{+}07]. affiliated [Bor02]. Again [Hin77]. AGM [Bre17]. ähnliche [Völ83]. Ai [CM13b]. Aids [SR53]. Air [Fet81a]. Airy [BP67, CM13b, CH78, CJR91, CJR92a, CJR92b, FO99, Fab04, FLO04, FK19, GST02b, GST03a, GS14a, GST20, HB75, JL12, Kra14, LO93, Mac94a, MPG92a, MPG92b, MP79, Moo81, Moo84, Pit67, Pri75, RS81, oKSAG79, SC81, TV09, VS04, VS10]. Airy-type [GST03a, GST20]. AIZ [GST02b]. Al [Aab54]. Al-Kāshı [Aab54]. Alan [Mil04]. Albuquerque [IEE91]. Alexander [Lip21]. Algebra [Fat16, QN18, Arm82, Fab02, Joh20c, Joh20d, Joh88, Ost54, Tre80]. Algebraic [ACM89, Bro07, FS96, Lak96, Ris79, SE11, Her82]. algebraically [Boy09]. algebraically-converging [Boy09]. Algebras [Var95, Hai60]. ALGOL [Tem79a, Wyn62, Chr62, Chr65, Frö63b]. Algorithm [ALMN05, ABJ19, Bai17, Bai93, BV85, BO93, Bic81, Bor19, Bor21a, Bor95b, Bre74, BDK^{+}05, Bur74, Cod83, Cod93b, Col78, CG85b, Cus83, DB13, Erc78, FFS83, Ful72, Gau64a, Gau73, Gau16, Gil16, GZ75, GLR07, Gus78a, Hei96, Hil70a, Hil70b, HD73, Hil73a, HMO19, Joh87b, KDDH94, Kir88, Kuk72a, KNS95, Let01, LC87, LT71, Mac89, Mer62, MKY92, Mor64, MW83, MCT18, OE82, PSB76, PEB04, PB08, Pom74, Pre78, Ree90, Rel63a, Rol87, SG02a, oKSAG79, Sel93, SR14, Smi91, Sny21, Soo73d, Soo73c, Tak01, Tem79a, VB04, VVA09, Wit68, Zag16, Zag17, Zen04, eL76, AN68, Bai88, BKM94, Bar81, Bor95a, Bre18, Bre19, BJ15, Bre78a, Bre80b, Bro89b, Bun09, Car72, CL94a, CO80, DMMM95, DAV12a, Dev12, EH89, Fdi97, Fet74, Flo61, Fuk83]. algorithm [GEK04, GST12b, Gon89, Gre82, GR19, Har14, Has00a, Jab12, Jab13, Jab20, KB89, KMY89, KCYL06, LO95, LP88, LC96, LC97b, LMS73, Mat92, McC77, Mit92, Mos69, NP18a, OY91, PN85, Per85a, SGA81, SS10, SS23, SD85, Tan90b, Ter81b, Tom00, Vep08, WMC18, Wal71, WTM05, Woź10, YA24, ZA14, ZLC04, AM20, Alh00a, ADW77a, Amo78, Amo83a, Amo83b, Amo86, Amo90, Amo95, And00, BC09, Bea60, Ber68, BP67, Bor21a, BB74, Bra64, Bra70, Bre82a, BL94, CN81, CZ10, Cle69, Cob69, CP70, Cod93a, Cun62a, Cun62b, Cun62c, CC64, Cyv64, DM87, DM92, Dor68, DR94a, Eck80, EC13, Fab04, Far69, Fet63, Fil66, Flo15, Fri67, Gau64b, Gau64c, Gau64d, Gau65a, Gau65b, Gau66a, Gau66b, Gau69a, Gau69b, GK70b, Gau79a, GST02b, GST02c, GST04a, GST06a]. Algorithm [GST11a, GST14a, GRAST16, Goa95, Goa97, Gra63a, Gra64, Gun65b, Gun65c, Gun65a, Gun67, HH18, Her61f, Her61d, Her61e, Her61b, Her61c, Her61a, HP67, HJ67a, HJ67b, Hil73b, Hil77, Hil81a, Hil81b, HP85, Hol70, Ibb63, Jan77, Jef62, Joh87a, Kod07, Kod08, Kod11, Köl68, Köl69a, Köl69b, Köl69c, Köl72a, Kop74, Kor11, Lar66, Lew75, Lin09, LG64, Lud63, Mac65, Mac68, Mac89, Mac96, Mey63, Mor76, Ng70, OW68, Pac70, PH66, PH67, Pit67, Pom76, PW90a, Pri75, RT61, RKZ^{+}14, RS81, Red70, Rel63c, Rel63b, Sch78b, Sko75b, Sko75a, Sko78, SM70, Smi01, Smi11, Sny93, Sny96, TS69, Tha62a, Tha63a, Tha63c, Tha63b, Tho13, VC06a, Vos73, Wie99, Woj61, ZA11, eL79, vdR63, Zag19]. Algorithmen [Mül90, Völ83]. Algorithmes [Bre78c]. Algorithmic [Cra12b, Che81]. Algorithms [Alh00b, Ano68, AH16, Bak75b, Bre80a, Bre06, Bre78c, CPCD19, CN81, CMW63, CH98, CHI^{+}03, DeL70, Deu76, Eps75, EL94, FLR03, Fan89, Fat16, Fow93, GST12a, God20, Gra63b, Gus86, Has90, Hil73b, Int00, Int03a, Int03b, JSH^{+}11, Knu98b, LF92, dL70, Luk77, The21, Mac89, Maj85, Mar72, Mar04a, MC87, MC90, Mil65, Mül90, Mul97, Nes84, Ng75, OS72, Pen81, Rus98, Sg92, Sor94, ST99, Tan91b, TBDS11, Tem94b, Ter81a, Tro84, Völ83, WG91, WG94b, Abe16, AIS^{+}17, Bak73, Bak75a, Bol06, BST13, Bor21c, BJM22, BM80, Bre10, CL61, CO86, CHGM99, Cra12a, DAV12b, Egb77, EH16, EM03, GA08, Gor82, Har03, Kog89, Lit89, Luk99, Mat90b, MMH93, MC89, MM90b, Obe99, Olv80, Saw02, Sch93a, Sch93b, Sch95, Tan87, Yey92, Zag24b, Fri72]. Algorithms [IL07, MC87, MC90, Rix82, Tha62b]. Algorithmus [Fil66, Rix82]. algoritmo [AN68]. alike [BCR^{+}07]. all-order [KWY07b]. Alley [Hei96]. Allgemeine [Töl68]. allgemeiner [von80]. already [FKY86]. Alternating [Alb62, CRZ00]. Alternative [ES10]. alternierenden [Alb62]. AMD [IM99a, Obe99, Rus98, Rus99]. AMD-K7 [Rus98]. AMD-K7^{T M} [Obe99]. Amherst [TT18]. Among [JMMW79, Ye17]. amplitude [Sal89]. AMS [DM08, Ric92]. Amsterdam [GST14b]. analiz [PS93]. Analog [BH65, Bin68, Pre55, Rob55, HFS99]. Analogue [SR53]. analogues [GA08]. Analyse [CHG^{+}11, Bre77]. Analysis [Abe88, BSY82, Das94, FP68, IM99a, IP87, Mil58, Par99, Pri66, Pug04, SK11, SA00, Sle64, Sle78, Smi95, Tan91b, AGA^{+}80, BCD^{+}02, Bre77, Bre78b, CH78, CDJ^{+}01, GST20, Has02, KW21, LP61, LP62, LČJ63, Mil75a, PS93, Ris79, SP61, Sob88, Tod54, Tod55, vRdL88]. Analytic [BG84, BS98, DB13, Ric73, Tra76, TW80, Wal00, BS00, BL71, BR94, Cri82, Fer86, Kza92, LP95, Mül93, Sea84, Sko04, Ter79, Tra76]. Analytical [Ano98a, Dar70, Fuk14a, GM04, How22, Sho60, CK89a, MWH^{+}16, MWH^{+}18, Van89, Hen77, Lov43, Lov89]. Analytically [IR08, LBC11]. Analytische [Hen77]. Ancient [BB12, Yon70]. Anger [Olv67a]. Angular [TTW79, V^{+}75, TTW84]. ANL [Kuk66]. Ann [IEE81a, IEE81b]. Anniversary [Gau94]. announcement [SSG^{+}18]. Annual [Spr00, IEE89b, Koo91, Men06]. Anomalous [Gau77a]. Answer [Wal80]. Antonio [Koo91]. any [AG08]. APMathLib [ZOHR01]. Appearing [Squ70]. Appell [ABF^{+}23, EG97, OSS09, TS76]. Appendix [Lov53]. Appl [Alz98, Ano84]. Application [AEF92, Hou81, IEE08, SF96, TBDS11, T^{+}97, VDR05, Ask75, CB09, CD02, Cro92, EC79, Fab02, GS91, Hor17, LG79, LKF10, Luk56a, Luk56b, PB82, Sho86, ZXBZH01, Lef05]. Application-Specific [IEE08, T^{+}97, VDR05]. Applications [AKS01, Ban01, BS79, Bor02, Bow53b, Bow61, Cor61, Flo15, Gre92, Gre59, Hol90, IK05, JKD92, Koe16, Kor02, Leb65, LKK19, MS68, Ng77, Sam02, SK11, SC01, Sid03, VS04, VS10, Wim81, XD18, AAS93, BB15, BB16, Buc69, CZ94, CZ95, CDS00, CMV69, AEF92, KS04, Leb72, MS67, NU88, OSS09, PT80, RM07, Ric92, SO89, TV09, Veb12, XC14, YC15, Ars62]. Applied [Car77b, ECSD20, HDG^{+}15, MH08, Ano98b, GS91, TW80]. Applying [QN18]. Approach [BJG^{+}00, BM15, Gal85, Har97b, Ike76, SA00, Tal68, Bre88, BCG91, FK88, Gal86, GS14c, JS89, Kre89, Lew94, LKF10, MWH^{+}16, MWH^{+}18, VPK99]. approaches [Wan16]. Approximable [AMM21]. Approximants [Bar74, GSS12, AGJ86, Bor88, DF78, Sab08]. Approximate [Bee99, Lit93, Lit94, Mak80, Mar72, MLS^{+}18, Mul20, OJL22, RX07, Via89, DK77]. Approximating [GW91, Ham78, Haw82, Jar12, Joh89, Lin89, NL94, QN18, SS99a, ZLA13, Boo55, BMST97, Epp89, HR05, HR07, KK96, MST89, Özb06, SLD16, Yey92]. Approximation [AJ03, Ach56, Ach92, ABJ19, BGV81, BM15, BM04, BR94, CPCD19, CG89, Caw00, CST18, CJL09, DB13, DiD78, Eck76, EW76, GS82a, Gil16, God01, Har66, IL07, IR08, ITY95, KL07, KP75, KS88, Kra85a, Lan64, Lin90, LBC11, Mei67, MS87, Mus87, Nor89, Osi13, PB82, PA86, POMB05, Plo96, Plo01, PLT23, Pug04, Qua18, Sch99, SS97a, SF96, Sha85, SL81, SK11, Sim64, Sou57, Squ70, SS99b, VLCSFN^{+}12, WS05a, WEX14, Wej83, WR63, Wyn60, YBR11, Yun09, Zag17, ZSA86, AQ18, AL20, AGS99, Bad76, BKKC09, Bra84, Bry02, CB09, Che16d, CL61, CD02, Cho14, Cuy94, Dan77, DAV12a, DAV12b, Dev12, Dri95, DD08, DZKK77, Ebe89, Eck77, Ell83, Fik67a, Fuk83, Fuk15d, Fuk15b, Fuk15c, GRS87, Glo89, Gor82, GS91]. approximation [Hor19, ITY97, JMS98, Kra85b, KSVW07, Lan59, Let96, LG79, LN22, Lor89, LSM14, LSM16, Mag94, Man23, Mar86, MC87, MC90, MRS17, MRS18, McC90, Mon83, MV98, MS15, Ost38, Peĭ63, Phi79, Pro87, Rei86, Roc82, SF93a, Sei99, Sho86, SS87, SR16, Sti59, Sun88, Tan90b, Tes69, Tim87, Ubh89, VT11, Ved93, VM87, VB81, WR96, WdZ04, Wat90, Yun14, Zwi87, vRdL88, Mei64, PT80, T.57, Zah94]. approximation-theoretic [VB81]. Approximationen [Hol69]. Approximations [BTT16, BS79, BCE^{+}08, BF23, Bur63, CG55, CG94, Car75, CDR71, Cod65a, Cod66, CH67, CT69, Cod69, CH70b, CST73, Dar70, Der77, Dun94a, Dun94b, DD07, Eve63, FL67, FB91, Fie65, Fik66, Fik67b, FR98, Fra65b, Frö61, Frö63a, HCL^{+}68a, HCL^{+}68b, Has55, Hit57, HBS00, Hol69, How22, KPPS08, KZ90, Kra14, LB09, Luk68, Luk69a, Luk69b, Luk70, Luk75b, Mae60, Moo67, New84, Nin70, Pag77, Pri75, RS81, ST07, Sch84, SS97b, SW99a, Sho05, Spi61, SDP11, Wad58, WC61, Wil70, Yoh70, Bla74, BEJ76, BEJ78, BBB89, BPD81, Cer07, CDS03, Cle54, CT67a, CT67b, CT68, CMV69, CPT70, CHT71, Dan77, Dun87a, Dun87b, Dun88, EG97, FW13, Fer86, FH62, GP64, Gau02, GL83, GM88, IP81, Jam99, Kas80, Kee82, Kon76, LW95, Let97, Lit93]. approximations [Lit94, LSM15, LM15b, Luk56a, Luk56b, Luk76, MT72, MG85, MG89, MPG92a, MPG92b, Mul06, OG82, PUHM12, Pie84b, Pow88, RM07, Rev90, Saa91, SK08, Sha76, Sie76, SE14, Spe72, Van89, Wan16, WC90, Wen07, Wes05, WBR82, ZB95, Zha95, Zha96a, ZB97, Ben83]. approximators [HSW89]. April [Ask75, Bor02, DM08, IEE13, IP87, KK99, Lan59, Tra76]. Aptitude [Völ83]. Aquifer [Har97b, Har98, Har01, Har08, PHH08]. Arbeit [Rut51]. Arbitrarily [How22, Joh20c, Joh20d]. Arbitrary [BH65, Bra87, Fat16, JM18, Joh21, Krä87, Lew75, LDP93, Pot14, Rot71, TS69, Bai21, BS00, Che09, DSK83, DSK84, Joh20a, Joh12, LS95, KR86, Sca71, Shi93, VC06a]. Arbitrary-Precision [JM18, Joh21, Joh20a]. Arbor [IEE81a, IEE81b]. arc [BBC08a]. Architecture [AFC10, ASA06, ALMN05, CHI^{+}03, ECSD20, HKST99, Int00, Int03a, Int03b, LP16, PEB04, PBLM08, RJM^{+}20, CL94a, GIJL23, KP98, MMW91, MM91a, MS82, ZA14, MS94]. Architectures [IEE08, JSH^{+}11, LP17b, LF92, T^{+}97, VDR05, Luk99, MMH93]. Arcsin [Kog58a]. Arctan [Kog58b, SK71]. arctangent [JL94]. arctangents [Car72]. Area [NKY08, SL95, Tom00, Cho14, LKH93, SR81]. Area-Efficient [NKY08, Tom00]. Area/Performance [SL95]. Areas [Nor89, Sha85, McC90]. Argument [Amo86, Amo90, Amo95, AM77, Bar61, BDL09, BS98, Chr65, GRAST16, GZ75, Kod08, Kod11, Köl72c, Mat04, Mec66, Min70, Mor76, Mor79, Smi95, Soo73a, Soo73b, Soo73d, Soo73c, Sou57, AM78b, AM79, AM84a, AM84c, AM84b, BY07, Büh87, Büh92, Büh03, Cam79, Cam81, Cam84b, Cam84a, Di 86, DGS24, Fuk10, GST04a, Gra61, HTHR94, Köl84, KRVZ98, KB86, Lew85, Lew87, Sca71, SW50, TB87, Tur72, VC06b, YM97, Zak69, Zha96b, ZB97, de 77, dT93, T.57, Tho04]. Arguments [Arc43, Bur74, CP98, Fox54, Fox60, GB68, Gau65a, GT58, Jan77, Krä87, Raf70, RL80, Tho13, TM68, Wan78, Bar82b, Bar84a, Bat19, Bre19, Cai11, Cob69, CJR91, CJR92a, CJR92b, GST04b, GST15a, Joh12, Krä88, Mas83, Moo81, Moo84, SG98, TB86, VPK99, Wal84, ZA24, Zag24b, L.61]. arise [SL83]. arising [BG84, GST20, Har77, Ye17]. ARITH [BC01, BBdD17, CLT20, IEE05, IEE21, IEE22, IEE23, MSH^{+}16, MTV15, TBL19]. ARITH-15 [BC01]. ARITH-17 [IEE05]. ARITH-26 [TBL19]. Arithmetic [Ber84, BHK09, BB84, BB97, BB00, BB04, BCDH09, BC01, BBdD17, CLT20, DLM09, DHL^{+}04, ES89, God20, Hul78, Hwa85, IEE78, IEE81a, IEE81b, IEE05, IEE13, IEE21, IEE22, IEE23, JF16, KM95, KK99, KM91, KM07, LMT97, MSH^{+}16, MTV15, OJL22, PL19, RJM^{+}20, Sal76a, SO11, SOL81, Smi91, SIJ93, TBL19, Tan89b, Tan89a, Tan90e, Tan92, TT18, Ull90, Wyn62, XLW^{+}20, von80, BG84, Bic81, BCK^{+}93, CG24, Che71, CA00, Eps75, EMR82a, EMR82b, Flo61, GLM15, Hou81, HS20, HJZ23a, HJZ23b, HJZ^{+}24, JM17, Joh20a, Joh20c, Joh20d, KMB94, LM00, LMOT01, MT72, MH72, PH90, RN96, Sal89, Sch93a, Sch93b, Smi01, Swa90, TRHH94, VCV01, dDM06, Atk79, HUY07]. Arithmetic-Geometric [BB84, Sal76a, MH72, Sal89]. Arithmétique [Ber84]. Arizona [DM08]. ARL [Fet81a]. ARM [BH07]. Array [Dav81, LA85, MK71, Maj72, CL95, LC97b]. Arrays [LE95, EM86, LE93]. Arsenal [Wyn62]. Art [IP87, MRH19, PTVF07, Toc63]. articles [Ban01]. Artificial [AGL93]. ASAP [VDR05]. ASIC [WMDM92]. Asilomar [LMT97, Sin95]. Askey [IA20, KS20]. Aspect [Cod67]. Aspects [AKS01, Gau67, Fer07, Tem94a, Tem00, Tem07]. Assessment [Boi97]. assistant [BMZ06]. Associated [Ano60a, Bra73, BL45, Cle69, Fri72, Lew58, Lew81, SJ01, Wim62, AB93, Bre18, CB12, CM13a, CE15, Che15a, Che15b, Che16c, CV03, Fre05, GQ09, Sai89, SSG^{+}18, SS65, SJ12, Wim72, Cob69]. Asymptotic [AS03, Bor10, BS78, CG85a, CG94, CE15, CS11, Cun62a, Din58, Din73, FO99, Kas80, KRT02, Kim72, Lóp00, ND19, Olv91, Sch64, Sle65, SZW11, SF87, Tem92, TV09, Tem15, Tho65, Wen07, Won73a, Won88, XHT16, XC17, YC15, ZZ11, AS05, Air37, Ano84, BB94, Che13b, CP15, Che15a, Che16e, CL16, Che16a, Che16c, CT16, CC17, DT09, DiD84, DGS18, EW63, FLMR00, Fet79, FF54, FF57, FW85, FKY86, GST03a, GST20, JW88, Kar01, KC11, LFM14, LW14, LSM15, Luk75a, Mac72, MP79, McC83, Mor11, Nem13, Par02a, Par02b, Par03, Par16, Sal83, SLH06, Tem75b, Tem78, Tem79b, Tem83c, Tem85, Tem00, TC01, Wen96, Won73b, Dop42, Sch80a, Sch68, Tri50]. Asymptotical [WCS16]. Asymptotically [Wan82, Hom98b]. Asymptotics [CTV96, CIM94, CV03, Har20, Mag94, Olv74, Wal96, Boe91, BBC08b, Che16b, FLO04, GG08, PW92, Tem83b, Tem86, Tem96b, Wil12]. Asymptotische [Dop42, Sch80a, Sch68, Tri50]. Atan2 [dDI15]. Atkin [KCYL06]. Atlantic [AFI71]. Atlas [OMS09, SO87, Tho97b, Tho97a]. ATM [MRS17, SR16]. ATM-forward [MRS17, SR16]. atmospheric [RM07]. atomic [SC81]. attempt [Che81]. Augmented [BJK^{+}11]. August [BL91, Bro07, Gau94, HY14, Mar85]. Austin [IEE13]. Australia [KK99, Lav80]. auto [PIH14]. auto-vectorisable [PIH14]. autocorrection [Lit93, Lit94]. Automated [Rev16]. Automatic [AK93, Bal00, Cah54, Che72, DeL70, Gra02, LM15a, dL70, MS82, Pie82, Pie84a, SZ70, SZ74, SZ76, VM19, AOB64, AN68, HXX22, Lin57, Wen64]. automatically [MS03, dLSG18]. automatico [AN68]. Auxiliary [LDP93]. available [Bur82]. Average [LB09, CD02]. avoiding [Hom99]. Axis [Zag19].

B [Köl90, Bor19, MIT^{+}95, Sab08]. B-splines [Sab08]. Babbage [CK89a].

Babylonian [FR98, Mur88, Mur99]. background [Kog89]. Backward [OS72, Tha79]. Backwards [Fat16]. Bad [MT85, MT86]. Balanced [BS98]. Baltimore [IEE99]. band [LP62, RX07]. band-limit [RX07]. band-limited [LP62]. Bandlimited [Osi13]. Bandwidth [Bru22, Far81]. Bartky [Bul69a]. Bartky-transformation [Bul69a]. Base [CVV65, Fet81a, Has66, LC87]. Base- [LC87]. Based [BO93, BEM12, DB13, GZC22, HAK91, KZ90, NS09, Sho05, SL07, Tal68, WMBE14, WG94b, BK16, Che16b, CHGM99, Dev12, DR04, HFS99, Kru99, KI16, LF92, LFM14, LSM15, Mac94b, MC87, MC90, MRS17, MRS18, Mat90a, Nis94, OJL22, PUHM12, PLT23, RdC93, SdCR97, TVG00, WG91, Yun09]. Basel [Ull90]. Bases [Sau93]. Basic [Ars68, Gas81, Kuk71, CIM94, Joh88, Kra95, Sch80b, Shi93, Ver77, DF84, VCV01]. Basis [AH10, Yon70, Hig77, Mul85a, Pow88, KR86, Sun15]. Bateman [IA20, KS20]. Bath [KM95]. Bayesian [JKD92]. BCD [Sch73]. BCY [Che81]. Be [PW76, DGB82, AK09]. beach [Ehr89]. Behandlung [Mei64]. behaved [Vai89]. behavior [Büh87, MP79]. Belgium [IEE08]. beliebige [Krä87]. beliebigen [Bra87]. beliebiger [Rot71]. Bell [NR15]. Bemerkung [Bau89]. Bemerkungen [Rut51]. Benchmarks [WOG95]. Ber [Her61f]. Berechnung [Fil66, Rot70a, Rot71, Sti53, von84]. beregner [Ped80]. Berkeley [Liu87]. Berlin [Gre16, AGA^{+}80]. Bernoulli [BH13, Che88, Fil92, Har14, HG64, KB67]. Bernstein [Bor82, Cri82, Hei88, Tim87, Whi82]. Beruf [Bar14]. besonderer [Völ83]. Bessel [Ano84, Bra70, Sko75a, Tho04, VRS^{+}99, Wim72, AS97, Abr45, Air37, AJDG02, Amo74, ADW77a, ADW77b, Amo78, Amo86, Amo90, Amo95, Ano46, Ano48, Ano52, Ano60a, Ant64, AM77, AM78b, AM79, AM84a, AM84c, AM84b, AWH13a, Ars81, BC79, BDGP13, BDGP14, Bal00, BL96, BGV81, Bar10a, Bar10b, Bar81, Bar82b, Bar84a, BA44, Bea58, Ben98, BK78, Ben83, BH65, Bin68, Bla74, Boe91, BST13, BO75, BBC08a, Bow58, BPD81, Bre19, Bur74, Cai11, Cam79, Cam81, Cam84b, Cam84a, Cha82, CIM94, CY18, CP66, CMF77, Cod83, Col80, CM83, Col84, CU59, CV03, Cur64, DCL^{+}92, DTLM98, Del79, Del84, DC81, ES99, EL01, Elb01, EM94, Fab02, Fat16, FS92, FMC82, Fox54, FO93, FKY86, Gab79, GKK00, GP64, GS81, Gau64a, Gau65b, GS78, GF97]. Bessel [GST02d, GST03a, GST04a, GST04b, GS14a, GF88, GL83, Gla74, Gla81, GT58, GT59, GCP67, Gra61, GM88, Har00, Har08, HF09, Har09b, Hil77, Hin77, Hit57, Hor17, Hum85, ISK87, IS88, IS90, IKK91, IS92, Ike76, IKF91, Ism77, Jab94, JL12, KW21, Kas80, KT00, KRT02, KKI67, Kor02, Kra14, KRVZ98, KB86, Laf86, Laf91, LM93, LW96, Lar69, Las82, Leh44, Len76, Len90, LG79, Lew69, LW82, Lin72, LK73, Lir71, Lóp18, LU96, LM08, LO93, LS95, Luk59, Luk64, Luk71a, Luk71b, Luk72, Luk75a, Mac94b, Mak65, Mar95, MG85, MG89, Mas83, MS68, Max91, McL34, McL55, Mec66, Mec68, MV98, Mul95, Nag01, Nas74, NR15, Nes84, NF63, New84, O’B80, Olv64, Par09, Par72, PSS03, PS77, PS79, PB82, Pie82, PB83, Pie84a]. Bessel [Pie84d, Pie84b, Pie84c, PA86, PT91a, PT91b, Puo88, Rap94, Rap01, Rap14, RdC93, RKZ^{+}14, Saf10, Sai89, Sal83, Sca71, Sch64, SdCR97, SG99, SS10, SS23, Sol69, Soo73a, Soo73b, Soo73d, Soo73c, SA57, Tak22, Tal83, Tal84, Tal09, Tem75a, Tem76, Tem79a, Tem81, Tem86, Ter81a, Ter81b, Tha79, TB86, TB87, VC06a, Vel11, VRS^{+}95, Wal84, Wat66, Wat95, WC90, Wil71, WBR82, Wim62, Woj61, XD18, YM97, YM03, Zai93, ZB95, Zha95, Zha96a, Zha96b, ZB97, Zhu10, de 77, dT93]. Bessel-function [Boe91]. Best [Hol69, Nin70, SS87, Zwi87, Kee82, Mag94, OG82, LW96, LU96]. bester [Hol69]. Bestimmung [Hol69]. Beta [BB74, BL94, DJ67, DM88a, DM92, OM68, PSB76, Sau92, Wis48, Gau64c, Gau64d, Lud63, Tem75b, TW80, YA24, PH67]. Beta-Function [DJ67]. Betrachtungen [Rut63]. Betriebssystems [Rot70b]. Better [BBD03, Wes05, BD02, LM00]. Between [BJG^{+}99, Hin77, Lar69, Lef05, Car69, QL19]. beyond [Ars81]. Bibliography [Ein79, Ful80]. Bickley [Amo83a, BEJ78]. bicomplex [Tre80]. biharmonic [Mon83]. bilinear [Töl68, Töl68]. Bilinear-Entwicklungen [Töl68]. billion [BBB89]. Binary [Che71, CVV65, DeL70, Har09a, Has66, LA85, dL70, Maj72, Mar72, Pat18, Por20, Rot70a, Bak76, CKL89, FHL^{+}07, Hig93, RN96]. binary64 [HJZ23a, HJZ23b, HJZ^{+}24]. binomial [Kra95]. Bipartite [SS97a, SS99a]. Birmingham [IP87]. BIT [Frö63a, Kah87, RHMLL08, Ziv91, LN21, LMS73, Per85a]. Bit-Parallel [RHMLL08]. Bits [Pat18]. Bivariate [Cad51, Div79, Dre78, JKMR11, LW89, Ruy89]. BIZ [GST02b]. BKM [BKM94]. BLAST [Has00a, Has00b, ZLC04]. Blinn [Bli97]. Block [Cas83]. Boca [Lip21]. body [CZ10]. Bombelli [RS98]. Book [Ars62, Köl87, Köl90, L.61, Lip21, Wig67, BBB97, BBB00, BBB04]. Boolean [Bel89, CG89]. Boost [BD99]. bootstrap [LW89]. borné [Del81]. Borwein [Bre17]. Borweins [Bai88]. Bose [Din58, Mac97, NDT69]. Boston [ACM87]. both [RX07, Sca71]. Boulder [Mul82]. Bound [AMS76, Jan11a, Jan11b, BJ15, BMST97, Cha82, CPG12, Ziv99]. Boundary [AO08, LL09, Cro92, JS89]. boundary-value [JS89]. bounded [Del81, GK89b, HR88, Zie89]. Bounding [Düm10, HKA93, Iac21]. Bounds [Alb62, Amo73, BST13, CCM11, FK11, IS88, Laf91, MDM16, Olv67a, Olv88, Rot70a, Tro84, WLK11, Yoh70, Yoh73, AM89, Bel89, BW89, Cer07, CD02, CDS03, Fre09, ISK87, Kra98, LW96, LKF10, Lit89, LU96, MJ13, Mor14, Nem13, Par02a, SS98, YCZ15]. Bourbaki [Laz85]. Bowman [Ars62]. box [DR89]. Boys [WO15]. Bozeman [BL91]. Brahmagupta [KB89]. BRAM [PLT23]. Branch [Har97a, Jef17, Kah82, Kah87, Sht19]. branches [JHC96]. Braunschweig [BCD^{+}20]. Brent [BJ15]. Briggs [AM12]. Brighton [IEE89a]. British [Gau94]. Broadening [WKG66]. Brook [Cow77]. Build [Sal92]. Built [Cow64]. Built-In [Cow64]. bulk [AM89]. Bureau [Low49, Mul82]. Burnside [Che16e, LFM14, LSM15, Qi14]. Byrd [Fet81b, Köl97].

C [Fet81a, Köl87, Sha76, Bak92b, ISO10, Joh12, Kor11, Lau95, MHSK16, Sch88b, Sel93, SP20, Tho97b, USE88]. C249 [Mar86]. C9X [McC02]. CA [IEE78, SE11]. Cairo [Wah04]. cake [Bel89]. Calcolo [RS98, AN68]. calcul [Che88, Mul85b, Mul86]. calculate [Ped80, RdC93]. calculateurs [Ber84]. Calculating [Bag95, Gel51, God20, Kal34, Kod11, LaF54, Raf70, Sal51a, TS76, Abe16, AOB64, DSK83, DSK84, Jab12, Jab13, Jab20, MKS83, Tal83, Tal84, Tal09]. Calculation [AS54, Chr62, DJ67, Div79, DP92, Eve63, Fet65, Fet67, Fik66, FW84, Gau99, Gel51, GLR07, GT59, Har97b, Hon71, Kin21, Kin24, Kin07, Lin57, LP83, Mec66, MKY92, MM91b, Mor80, Mor78, Mor99, Pot14, KR86, Saa91, SH72, SR53, SF59, Spi71, Str68, TM68, War60a, ZA24, Zak69, vdLT84, Air37, AN68, Bel58, Bul65a, Bul65b, Bul69b, Che88, DM91, FMC82, HvdR63, Hor17, JL12, LR74, Len90, Mid75, MH72, MWH^{+}16, MWH^{+}18, Mou67, Mul85b, Pan59a, RL76, RS98, Sal89, Saw02, Sti53, Tak22, Tur43, VRS^{+}95, VRS^{+}99, Wei98, Zag24b, Köl87]. Calculations [DLM09, DPC95, GLT^{+}15, Hen77, KKI67, Kor11, Len76, AJ03, Sal76b, Tri66, Tur53, Yar67]. Calculator [Hen77, MW83, Nor89, OMS09, Pag77, Egb77, Lin89]. Calculators [BR73, Cus83, Der77, Kir88]. Calculus [FS96, SO89]. California [ES89, LMT97, MSH^{+}16, Sin95]. Cambridge [IEE89c]. Canada [Bro07]. Cancellation [CM13b]. cannot [DGB82]. Canonical [Rap01]. capability [Bic81, LMT^{+}92]. Cape [IEE05]. Capital [MMH89]. Carathéodory [Mag94, VT11]. Carathéodory-Fejér [VT11]. cardinality [Was89]. Carefully [HH18, Wil96]. Carlson [Gra02, JKD92]. Carnegie [Tra76]. Carnegie-Mellon [Tra76]. Carolina [BCEP94, IEE89b]. Cascaded [KNS95]. Case [Cas83, EH16, OLHA95, Sle78, Sou57, ABFP24, Eck77, Ein89, Gau11, Har20, Kra98, KI16, LW96, SK08, Sko04, T.57]. Cases [Car89, LM03, LSZ08, SLZ05]. Caslin [Fet81a]. Cause [Zag19]. CD [Spr00]. CD-ROM [Spr00]. CDC [ADW77a, ADW77b, Amo78, Cod64]. CDC-3600 [Cod64]. CDF [Sho05, YBR11]. CELEFUNT [Cod91, Cod93a]. cell [JL94, AK09]. cells [MO18]. Cellular [MK71, Maj72, CL95]. censoring [Ruy89]. Centenary [BCEP94]. centennial [AGA^{+}80]. Center [Ask75, IEE85, AGH89, Saf10, Vai89]. center-unstable [AGH89]. central [GST15b]. Century [Lid01, Lid02, Bre00, Gau94, Hay99, Plo96]. Certain [Ano46, DeL70, Deu76, Erc73, Lan60, dL70, Pan59a, Sch64, Sri70, Ver77, Abr45, Akr85, BC79, BL71, CCV11, De 91, EW63, GK70a, Kra85a, Las82, Leh71, Lew94, LP95, MM95, MM98, RX07, SS93a, Sie76, SZW11, Win70, ZSA86, ZZ11]. Certification [Ber68, Bra64, Bra70, BL94, Cob69, Far69, Gau64d, Gau65b, Gra64, Köl69a, Köl69b, Köl72a, Kop74, Lar66, Lew75, Mey63, Ng70, OW68, RT61, Soo73d, Soo73c, TBDS11, Tha62a, Tha62b, Tha63c, Tha63b, vdR63]. Certified [CJL09]. Certifying [dDLM11]. Cesaro [MS87]. chain [Paw11, Wil96]. challenging [Sid06]. Chandrasekhar [Jab12, Jab13, Jab20]. Channels [LB09, SA00, SA05, CDS03]. chaos [PH90]. CHAPLIN [BD14]. CHAPLIN-Complex [BD14]. Chapman [Lip21]. character [BB16]. Characteristic [Fri72, AB88, Bos89, Fuk09b, McN73, Cle69].

Characterizing [Kee82]. charge [MO18]. Charles [CK89a]. Chebyschev [BH63]. Chebyshev [CT67a, Frö63a, Hol69, Per87, Bar61, Bla74, BEJ76, BEJ78, Boy89, Boy01, Boy14, Cle55, CP66, Cod65a, Cod65b, Cod66, CH67, CT67b, CT68, CT69, CMV69, Cod69, CH70b, CPT70, CHT71, CST73, CM83, Dar70, Del79, Del84, EW76, FP68, Frö61, GP64, GCP67, Hum64, KP03, LLK14, Luk76, MT72, MH03, Ném65, NDT69, Per85b, Pie84b, Pri75, RS81, Sch78c, Sch80b, She74, She76, SL81, Tan90b, Wat90, WC61, WBR82, Woo67, ZB97]. Checked [Rus98, Rus99]. Chemie [Sne63]. Chemistry [Sne56, Sne61, Sne80, Sne63]. chemists [Wen09]. Chernoff [CCM11, CPG12]. Chernoff-Type [CCM11, CPG12]. Chester [IL07]. Chi [HP85, Rob69, Tho65, eL76, DH77, GST15b, SLD16, HP67, OW68]. Chi-Square [Rob69, DH77, GST15b, SLD16]. Chi-Squared [HP85, eL76, OW68, HP67]. Chinese [Che81, Liu93, XLX^{+}83, XL94, Yon70]. Chip [MLS^{+}18]. choice [Ge89]. choices [SL97]. Cholesky [Sch84]. Choong [Sha76]. Choosing [SL97]. CIMPA [HUY07]. Circles [Kin21]. Circuit [IEE89a, MIT^{+}95, MCT18, Bel89, CCC96, CP99, HFS99, MI97, SL63]. Circuits [DM88b, Maj85, OTY91, Bel89, KK98, Liu95]. Circular [Ano74, DJ62b, DJ62a, Ken14, Ken21, WD61, Zuc64, ZS08]. City [AFI71]. Civic [IEE85]. Clara [MSH^{+}16]. Clarification [ACG^{+}87]. Class [DeL70, Dub83, Gla81, Lir71, dL70, CR08, Cro92, GQ09, Gus78b, Gus84, Mak80, MM95, Mos69, Som82]. Classes [OG82, Cri82, Kaz65, Liu93, Lu14, SZW11, Ubh89, ZZ11]. Classical [Loz00, PS05, Tem96a, BG84, Håv81, Ye17]. Clebsch [Buj75, Gun65a, RB75, SM77, SM78]. Clebsch-Gordan [SM78]. Clenshaw [PB83]. Clever [Kah04]. Close [Har66]. Closed [GW71, Jan11a, Jan11b, RM07, XC15, CTV96, Dav88]. Closed-Form [Jan11b, Jan11a, RM07]. cm [L.61]. CMOS [HFS99, TRHH94]. co [DMS95, USE88]. co-processor [DMS95]. Coaxial [Kin21]. Coconut [AK09]. Cod [IEE05]. Code [BdDKL15, KL14, RdC93, SdCR97]. Codes [CJL10, FLR03, Gus78a, SG99]. coding [Flo61]. Coefficients [Bra65, Buj75, Der77, RB75, SG76, Str56, Boy09, FNC75, FW80, Fra81, Gun65a, Luk56a, Luk56b, Pan59a, Pan59b, Sch76a, Sch77a, SM77, SM78, SMCH41, WV20, Ye17]. Coerror [Amo73, Gau77b, Gau16]. colleagues [Bir54]. Collection [WNO94, Bat19, Sne63]. Collective [Sti88]. College [MC87, MC90]. Colorado [BC01, Mul82, Luk99]. Coloring [PP12]. Columbia [Gau94]. Combination [LA03, Lit83, Sho86]. Combinatorics [GI01]. Combined [LM95, Nan11, ALB98, LM99, NL99]. combining [EL89a, MI97]. Comment [T.57]. Comments [Fat16, Kuk66, Tra60, DD08]. Common [BK78, Ben83, DMS95, LM07, PSS03]. Commun [Tho04, VRS^{+}99]. Communication [SA00, SA05]. Communications [BS79]. Comp [Fra81, Lew87, Per87, Sha76, Wim72]. Compact [Caw00, Sch99, CK88, Whi82]. Companion [HDG^{+}15]. compare [QW06]. Compared [JJ65]. Comparison [And82, Ng75, SL95, LMS73].

Comparisons [FLR03]. Compatible [Tay81]. compensated [Bor21c, GIJL23]. compile [Mul06]. Compiler [DG99]. Compilers [BD99, LKK19]. complement [CL94a]. Complementary [Bai17, Cod90, FCC73b, VM19, Chr62, EL08, FCC73a, HR72, Phi79, RM07, Zag24a, Zak69, HJ67b]. Complete [AVV90, AVV92, Cod65a, Cod65b, Cod66, FC64, Goa95, Mer62, OM68, QGG10, QV96, Sko78, War60a, BZ92, Boy15, Fet81a, Fuk09a, Fuk11, GM07, Ism77, MH72, Mor78, PDK96, Qi14, RT89, Far69, Her61d, Her61e, Lar66, Tha63a, Tha63c, Tha63b]. Completely [QL19, BP01, CE15, CP15, Che15b, GQ09]. Completeness [Hig77]. Complex [Ahr96, Amo86, Amo90, Amo95, Ano74, AM77, Bak92a, Bra87, Bur74, CP98, Chr65, Cod91, Cod93a, Cor61, Dor66, EW76, EM04a, EM04b, EM07, FK19, FCC73a, Fet76, FK75, Fri67, Gel51, God01, Her61b, HFT94, Ise91, Kah82, Kah87, Ken14, Ken21, Kod08, Kod11, Kol70, Köl72c, Krä87, Kuk72a, Kuk72b, Lan60, LT71, Mat04, Mec66, Neh07, Ng75, Par72, PS11, PP12, PW90a, PW90b, Rap94, Sal51a, Sal55, oKSAG79, Sht19, SMG94, Soo73a, Soo73c, TTW79, WEX14, Wyn62, Zag19, AQ18, AM79, AM84a, AM84c, BKM94, BCD^{+}02, Cai11, CG24, Cam81, Cam84a, CL11, Car95, CA00, CM83, CJR91, CJR92a, CJR92b, CDJ^{+}01, DGS24, EC13, FLO04, Fet81a, GST02b, GST02c, GS14a, GJL20, Gon52a, Gra61, HHR00, JM17, Joh20a, Joh12, JT85, Jür66, Köl84, Krä88]. complex [KRVZ98, KB86, LG79, Mas83, MY91, Moo81, Moo84, MH72, Mor78, Mor99, PS86, Rap01, Sca71, SK08, Squ91a, Squ91d, SW50, Str59, TTW84, TR84, Tan90b, Tem94a, TB86, TB87, Tho04, Töl67b, Tur72, Wat90, YM97, ZA24, Zag24b, Zak69, Zha96b, ZB97, dT93, Bea60, BD14, Gau69a, Köl72a, RT61]. Complex-Valued [Mat04, PS11, Wat90]. Complexities [NS09]. Complexity [AH10, BB88, Mil75b, RHMLL08, Vav89, Bel89, Bre75b, Bre76b, MST89, Per85a, Tra76]. Compliance [Rus98]. complying [AHM^{+}98]. Component [vRdL88]. compounding [Leh71]. Comprehensive [AWH13c]. Comput [Alz98, Ano84, Tho04, VRS^{+}99]. Computation [ACM89, AS95, ARH14, Aha69, AH10, Alh00b, Amo74, ALB98, BB16, Bal00, Bee91, Bee17, Boe62, BB84, BB97, BB00, BB04, Bre06, BH13, BL94, Bro07, CS19, Che70, Che72, CJL09, CMF77, CVV65, Cri89, DH76, DH77, DM86, DM88a, DiD90, DM92, Dre78, Dub83, DGS24, Ehr61, Ekl01, FLO04, Fil66, Fle68, Frö63b, Fuk14b, Fuk15a, Ful99, GI01, GSS12, Gau61, GST03a, GST12a, GST14a, GRAST16, Goa95, God01, HT13, HCK09, HKST99, Har09b, Has66, Hul78, JSH^{+}11, JKD92, Joh14b, JM18, Joh21, KT00, KB67, Kog57, Kog58b, Kog58a, Kog59, Köl72d, Lak96, Leh44, Leh96, Let01, LDP93, Loe01, LJ13, Luk57, Luk77, Mac94a, MMV81, Mar90, MR71, Mec68, Mül04, Nad15, NKY08, NPA16, ND70, Pea09, PB02, PEB04, PW90b, Rap94, Rob69]. Computation [RHMLL08, Rot70a, Rot71, Rus13, Sal76a, Sar59, Sch85, SE11, Spo94, Swe63, TAY91, Tem79a, Tra60, VVA09, Wen96, Woo92, XD18, YA24, ZJ96, AQ18, AGA^{+}80, Alh00a, Bai88, BBP97, BB15, BL96, Bat19, BY19, BM80, Cai11, Car95, CZ10, CH89a, CH89b, CA00, CD02, CDS03, CCV11, Cra06, DMMM95, DKK80, DHM89, DK90, EH16, Fil92, FH62, FI94, Fuk09a, Fuk09b, Fuk10, Fuk11, Fuk13a, Fuk13b, Fuk14a, Fuk15d, Fuk15b, Fuk15c, Gau63, GK70a, GS78, Gau94, Gau11, GST02b, GST02c, GST11b, GST15b, GST17, GS99, HHv^{+}73, Hit74, JJW86, Joh15b, JT85, Kat78, Kon76, KB86, Las82, LKK19, LK73, Mar81, MNP83, Mer94, Miy19, Mul85a, Mul86, NP18b, Ng84, Pan59b, POP17, PB82, Pie82, PB83, Pie84a, Pie84d, Rap01, Sar60]. computation [SS10, SS23, Tem83a, Tem86, Tem87, Wim74, XL94, Zag24a, Zah94, Zha96b, BL91, LČJ63, Low49, Mül90, SR53, von84, Köl90]. Computational [BBC98, BBC00, BW00, Gau67, Gau75, Gau79b, Mil75b, Ng75, Qua18, Sun71, Tem94a, Wim82, Gau94, Jam89, SW88, Tra76]. Computations [Ahr96, BH18, Erc75, Erc77, Krä93, MT85, Ng77, WG91, WG94b, Bel89, Dan77, Joh83, KO94, Kra98, MT86, Vav89]. compute [BBB89, Boy15, EC13, Joh12, Mac96, Per85a, Sca71, SSG^{+}18, Win70]. Computed [Woj61, Yoh70, Che09]. Computer [AFI69, AFI71, Atk79, BH65, Bin68, BCDH09, BC01, BBdD17, CET62, CLT20, DeL70, Erc75, Erc77, ES89, Fat16, Fik68, HCL^{+}68a, HCL^{+}68b, Hwa85, IEE78, IEE81a, IEE81b, IEE89b, IEE89c, IEE05, IEE13, IEE21, IEE22, IEE23, JJ65, KB90, KM95, Kog57, Kog58b, Kog58a, Kog59, KK99, KM91, KM07, LMT97, dL70, Mar88, MSH^{+}16, MTV15, PH90, PW76, Pre55, Pri66, QN18, SO11, Swa90, SIJ93, TBL19, Tan89b, TT18, Tra76, Ull90, Wah04, WWG51, WWG57, WWG82, von80, AGA^{+}80, DSK83, DSK84, Fab02, GP84, Gus78a, Knu62b, Knu63, McC90, Pie84d]. computer-oriented [AGA^{+}80]. Computers [Ano03, Ano06, CU59, Cus83, FBM04, Has55, Köl90, Mol83, Sin95, SA57, Sug58, Ber84, RW77, Tak00, Tri66]. Computing [AK93, API^{+}19, Bow53a, Bur82, Cah54, Car79, DR98, DH59, DJ62b, DJ62a, DM87, DR94b, For97, Fra65b, Gal85, Gal86, GST02a, GST03b, GKS04, GST04b, GST06b, GS12, GST15a, HH18, Har97a, Ike76, IKF91, JKMR11, Joh16, Joh19, Joh20a, Köl72c, Kuz15, LBdDM16, LW63a, LW63b, Lin09, Lin72, LCY65, Mar05, MMM93, MY91, Mul01, Mul20, PL19, PTVF07, Rob55, ST07, SKL93, SLD16, SZ70, SZ74, SZ76, Tak01, Tho97b, Tho97a, TKK17, Van69, Vol59a, Vol59b, YM03, ZA11, Zag16, AIS^{+}17, BDM81, BPT76, BCJV18, Car72, Dom03, Fet74, Gab79, GST12b, GLM15, Har14, HHR00, mH12, JM17, Kan96, Köl84, KCYL06, KRVZ98, Let96, McC81, PKP74, PNV01, Pól49, Pro83, Puo88, Saf10, Sko05, Vep08, XC15, Zen04, VC06a]. Concept [von80, Gon52b]. Concerning [Wre68, Arm82, CC17, SW88, Wre73]. Concise [PC11]. condition [HS83]. conditional [Fuk13a]. Conditioning [Ric65]. conditions [JW88]. conduction [CMV69]. conductive [RdL01]. conductive-convective [RdL01]. Conference [AFI69, AFI71, Ano03, Ano06, Ban01, BCD^{+}20, Boi97, BCEP94, FBM04, Gre16, HY14, IEE89a, IEE85, IEE89c, IEE08, IP87, IL07, MMH89, MC87, MC90, Men06, Mil75a, MS94, Mul82, SKL93, Sin95, Spr00, T^{+}97, USE88, VDR05, Wah04, ACM87, AGA^{+}80, BL91, AEF92, IEE21, IEE22, Mar85, Zah94]. Configuration [Yos97]. Confluence [Nag04].

Confluent [AS95, GN89, Luk59, Mil57, NPA16, Olv91, Sla60, Sla64, Tha62b, AS97, AG08, Buc69, BJLP19, CQSP04, DT09, DGS18, Gau02, KW21, Lew44, Mul01, NP18b, POP17, Tem78, Tem81, Tem83a, Uni49, Rel63c]. Congress [FvdHJT10, Lav80]. conical [GST12b]. conjecture [CLZ04, EL01]. conjugate [MS87]. connect [CK88]. connected [Ost54]. Connection [GSS12, Car69]. considerations [Wen01, Rut63]. consistent [Hom98b]. Constant [Chl11, Knu62a, Swe63, TAY91, AL20, Bau89, BY19, Blo88, BM80, CM13a, Che16a, Che16b, Fuk09b, Lu14, LSY15, Luk56a, Luk56b, MWH^{+}16, MWH^{+}18, XC17]. constant-time [BY19]. Constants [HMO19, Str56, BBP97, CC17, CB87, GS99, KC11, Sti53, SMCH41, Yar67]. Constrained [Let97, MT72, SDP11]. Constraints [Kuk71]. constructing [Kai89]. Construction [Ske86a, Ske86b, Dan77, Lit93, Lit94, OWS^{+}14, Pas95]. contained [CC69]. Containing [NF63, Sch64, De 91, Las82, Lóp98, Mak65, Tem86]. contemporary [Ano98b]. Contents [Ful81b]. Context [FR98]. contiguous [Seg08]. Continuation [BS98, BS00, Mül93, Sko04]. Continued [Bar74, Bar82a, BS83, Che13a, CPV^{+}08, Len76, Len90, Mor64, Pas03, She54, SB71, Smi95, Spi61, Wal00, BC09, BHJ05, BS78, Cha80, Dij77, Gau77a, Hau88, JW88, Kra89, Lee92, Lem88, Lev91a, LSY15, LM15b, LSM16, MM90a, McC74, McC83, Now06, Pas92, RdC93, Sar60, SdCR97, Sha76]. Continued-Fraction [Smi95, Bar82a]. Continuous [PS93, Ebe89, FH62, GS82a, Hor19, MS87, PS78, Som82]. Contour [ST07]. contoured [Was89]. Contribution [Chr62, Frö63b]. Contributions [Ng84, Ull90]. Control [Kuk72a, Kuk72b, Bel90, BL91, Bre83, Roc82, Mae60]. controlled [Bra64, CC64]. convective [RdL01]. Convergence [Alb62, Ano67, Bre78a, Bre82b, BDGB83, Bre85, Bre00, Chl11, CRZ00, DGS65, DK82, GB90, Gus66, Gus78b, GS91, HP93, Har77, Hau88, HBS00, Hom96a, Hom96b, Hom98a, Hom99, Ise94, JW88, Kru99, Kza92, Kza99, KP03, Lem88, LP88, LB90, Lev91b, Mat92, MH72, Mül99, Osa90, Osa94, Ovt03, Pas92, Pas03, Pas08, Sed90, Sel93, SF79, UE05, Wyn72, AL20, ABFP24, AGJ86, BD85, Bel90, BHJ05, Bre77, Bre78c, Bre78b, Bre83, Bre88, Bre99, BR94, BCG91, Cro92, DGB80, Del81, DF78, Fdi97, FS92, Gau77a, GK89b, Gus78a, Gus84, Gus85, HTHR94, HR05, HR07, Hom94, Hom98b, Ise91, Joh83, KI04, Lev73, Lev91a, LP95, Lon81, Lu14, Mat90a, Mül93, Now06, Ole96, Ost37, Pas10, Pas11, PR04, PKP74]. convergence [RdL01, Sau93, Sch68, SL83, SS98, Sid06, Vep08, VPK99, Wen89, Wen01, Bre77, Del81]. Convergent [Bar74, BJLP19, God01, Lóp18, Osa90, UE05, Bai88, DGB82, FGG82, FF54, FF57, Hom99, Lew94, McC81, WN09, Woź10]. Converging [Din58, ITY95, Air37, Boy09, Jam89, Kat78, Sch95]. Convex [Lit83, EH89, Fuk83, Zwi87]. convexity [GR92]. conveyors [KK98, Liu95]. convolution [Mül99]. Coprocessor [KZ90, TVG00]. Coq [BMZ06, MDM16]. CORDIC [Mül90, Ekl01, HTHR94, Hit74, SR14, TAY91, VVA09, Vol59a, Vol59b, Völ83, ZCL15, MMM93]. CORDIC-Algorithmen [Mül90]. Cornelius [BCEP94]. Corner [Bli97]. Corners [Tho93]. Correct [LM03, SF16, FHL^{+}07, Mul24]. Correction [BHK09, Yun14]. Corrections [DD08, Jan11a, WMC21, PC93]. Correctly [Bor21b, BJK^{+}11, BHMZ24, God20, Kah96, LA01, LM00, Ziv91, Ano14, BJM22, BMR04, CLZ04, HJZ23a, HJZ23b, HJZ^{+}24, LMT98, LN21, LN22, SS93b, dDLM04]. Correctly-rounded [BHMZ24, BJM22, HJZ23a, HJZ23b, HJZ^{+}24]. Correctness [CH98, CHGM99, Rus99]. Correspondence [Hon71, vRdL88]. corresponding [McC83]. Corrigenda [Ano94, Cod66]. Corrigendum [Alz98, CT67a, Joh87b, Lew87, Per87, Sny21, Wim72, Fra81]. Cosine [BH63, Cun62a, Cun62c, SZ76, TAY91, Tan90d, Tha61, Bak76, Har00, JL94, PKS00, Saf10]. Cosines [Pat18]. Cost [CL00, Dan82, XLW^{+}20, BHMZ24, FLMR00]. Cost/performance [CL00]. CoStLy [Neh07]. Cotes [Fre09]. COULFG [Bar82b, Bar84a]. COULN [NT84]. Coulomb [Abr64a, BDG^{+}72, BDG^{+}84, BFSG74, Bar76a, Bar76b, Bar81, Bar82b, Bar82a, Bar82c, Bar84a, BFSG84, Bar84b, Bar84c, BS84, Boe69, CH70b, CP70, DGA18, Frö55, Gau66b, Gau69b, Gun67, Hum85, Ike75, Ike76, Köl69a, Köl69b, Köl69c, Köl72d, LVV05, MKS83, Nes84, NT84, PGA18, Sea84, SG72, TTW79, TTW84, TR84, TB86, Vos73, Wil71]. count [MCT18]. coupling [Gun65a]. Course [BW00, Tka03]. Coverage [DJ62b, DJ62a, DiD90]. CP [LVV05]. CptnHook [PI16]. CR [Ano14]. CR-Libm [Ano14]. Crandall [BB15]. Cray [???90, KB90]. CRC [Lip21, Zwi03]. Create [ES10]. Criteria [Ike76]. criterion [Fik67a]. Critical [BH18]. crossover [PUHM12]. crunching [Atk79]. cryptography [Per85a]. cryptoprotocols [Per85a]. crystalline [Har77]. Cube [AH10, Ded26, Joh11, Lan42, Mur88, PBLM08, Yoh70, Yoh73, Bur82, Pen81, RS98]. Cubic [Car89, Flo15, WdZ04]. Cubiche [RS98]. Cubing [AH10]. Cumulants [FC60]. Cumulative [Gil16, Ham78, Pag77, Pom74, Pom76, BKKC09, GST15b, MRS18, Rev90, SE14, Wes05, Yun09]. current [BIS01, KK98, Liu95]. Curtis [PB83]. curvature [Ron86]. Curve [Che70, HJ67a, HJ67b, Hol70, Nor89, Sha85, Squ70, Ber68, Cho14, McC90]. Curves [Sch85, Nis94]. Cuts [Kah82, Kah87, Sht19]. Cutting [Tho93]. Cylinder [Dor66, Fox60, L.61, Mil64, WR71, DGS24, GST06a, GST06b, GST11a, GST11b, GS12, GR92, Jon06, Kre89, LR74, MMV81, RL76, SGA81, SG98, SMCH41, Tem00]. Cylindrical [Kod08, Kod11, Mas83, KB86].

D [Fet81b, Köl97, Sha76, BJK^{+}11, NKY08]. d’accélération [Bre78c]. Dagstuhl [FBM04]. Dallas [Koo91]. d’après [Laz85]. Data [Joh88, Krä87, LK21, Mae60, MC87, Ran77, BCG91, EMR82a, FK88, Wen01]. datapath [CCCP99]. Date [Ano22]. Datenformate [Krä87]. Davenport [Laz85]. Davidson [Ovt03]. Dawson [AQ18, CPT70, Dij77, Hum64, Let97, McC74, MM88, Ryb89, Zag24a, Zag24b]. Daykin [Sha76]. DC [MMH89]. DDMF [BCD^{+}10]. Debye [ND70]. decade [Bre85]. December [ACM87, BCEP94, MMH89, Zah94]. Decimal [Har09a, VVA09, WS05b, Bai88]. decimal64 [LSZ08]. decision [Ye17]. decomposition [CZ95]. Decreasing [BFHT85]. Dedicated [IK05].

Defektberechnung [Mül90]. Defined [Wyn60, Tre80]. definite [Gla74, Won73b]. definition [BL71]. Degenerate [PS79]. degree [Bre18, Tes69]. degrees [QL19]. Dekker [Bol06]. demonstration [CCC96]. Denominator [BM04]. density [CKL89, McN73]. Denver [Luk99, USE88]. Department [Tra76]. dependent [Kre89]. Depth [BFHT85, OTY91]. derivation [Din73]. derivative [CS82, CS83, DC81, GST11a, IKK91, LLK14, War75, Wen64]. Derivatives [Amo83b, AG08, Dan82, GK70b, Ike75, Ike76, Lau73, ST74, SB71, Tem79a, XD18, Bar82b, Bar82a, Bar84a, Bus74, Che81, Dom03, Fuk13b, GK70a, GS14a, GOST23, Joh15b, MP79, McC81, Sai89, Sun88, Tak22, TV09, Töl68, XC15, ZZ96, Gau66a]. describe [MO18]. Design [ASA06, BT99, BH01, BH07, CCC96, HXX22, IEE89a, IEE89c, LKH93, LMS99, MCT18, OLHA95, Rev16, SS94, TBDS11, ZG87, SL97, ZA14]. designed [CTA22]. designing [WN03]. designs [CL94a, Sg92, SS93b]. Desirable [Hul78]. Desk [BR73]. Details [Bai17]. Detection [QXG18]. DeTemple’s [Che16b]. Determinant [CT88]. Determination [BB94, Hol69, Liu93, de 77, Aab54, Ter79, VRS^{+}95, VRS^{+}99]. Deuflhard [YA24]. Deuxièmes [AEF92]. Develop [ACG^{+}87]. Development [CJL10, Lev73, PN85, Ypm95]. developments [Dop42, GST14b, Sch80a, Töl68]. Deviate [HD73]. diagonal [Kra89]. dicembre [Ano98b]. Dictionary [BCD^{+}10]. Difference [FK19, LA85, Olv67a, CH78, Luk56a, Luk56b, OWS^{+}14, Olv67b, QL19]. differences [MNP83]. different [Her71, KW21]. Differentiable [PS93, Pro87, Zie89]. DIfferential [BK16, AO08, Cas83, Cle69, Fri72, HBF09, IS90, Arm82, Dev12, EM94, FLO04, GST04b, IS92, Luk56a, Luk56b, KR86]. differently [DM91]. Differentsiruemye [PS93]. differintegrations [Nis84]. Difficult [Alt79]. Difficulties [Lon81]. Difficulty [Tur69]. diffuse [MO18]. Digamma [Köl72c, Lin79a, Moo67, Pai19, Spo94, Fer07, Köl84, QM15]. Digit [ALMN05, BL94, BTDS11, DM88a, DiD90, DM92, EL94, EM03, HSE^{+}24, Kor05, LA01, PBLM08, Rus13, TBDS11, CA00, LKKH99, LE93, LP83, MC93, MC95, PKP74]. Digit-by-Digit [PBLM08, LP83]. Digit-by-Rounding [BTDS11, TBDS11]. Digit-Recurrence [ALMN05, EL94, EM03, LE93]. Digital [Ano22, Bow53a, CET62, CU59, Erc75, Erc77, Has55, IEE99, LMS99, Loz03, Maj85, MKY92, SA00, SA05, Sug58, WWG51, WWG57, WWG82, Lin57, RW77]. digits [Bai88, BBB89, ZA24]. DIIS [Sel93]. Dilemma [Kah18]. Dilogarithm [GZ75, Mor76, Mor79, O’S16, Köl68]. Dilogarithms [Lew58]. dimension [LP62, vRdL88]. Dimensional [BTT16, WEX14, PS93, SF10, VB81]. Dimensions [Sle64, Pól49]. Dirac [CT67a, MN16, BDM81, CT67b, Din58, Fer86, Fuk14a, Fuk14b, Fuk15d, Fuk15b, Fuk15c, GOST23, Goa95, Let01]. Direct [GBS19, HAK91, Kin24, Kin07, PDK96, LKK19, Ng68, Wei98]. direction [DR89]. directions [BIS01]. Dirichlet [BB15, BM15]. discrepancy [Pro88]. Discrete [JKK92, JMS98, Mul85a, Roc82, Sle78, Per85a, Rei86, Sch96, Tur72].

Discussion [BJK^{+}11]. Discussions [Cah54, Ded26, Esc37, Kal34]. dispersion [CH78]. displacement [YH89]. Distance [Lef05]. distributed [Tak00]. distributed-memory [Tak00]. Distribution [Dor68, Gil16, Gup60, Ham78, HJ67b, Hil70a, Hil77, Lin09, Mac65, Mar04b, Mor80, PR82, PR96, Pom74, Pom76, Rob69, Sho02, Sho05, Tim87, VLCSFN^{+}12, YBR11, Bad76, BKKC09, Cyv64, DH77, DC81, Fet79, FW84, Fuk14b, GST15b, HB75, KPPS08, Leh70, Mac68, MZM94, MRS18, Rev90, Sho86, Ved93, Wri84, Yun09, Hil81a, eL79]. Distributions [FC60, JKK92, SZW11, Wan82, ZZ11]. divergent [Wen89, Wen96, Wen10]. Divide [BH07, CH98, D^{+}89, GBKK09, Int00, Int03a, PZ95, SL95, Wil70, ZG87, AGS99, BH01, CHGM99, Mat90b, MM92, Saw02, WMDM92]. Divide/Square [SL95]. divided [MNP83]. Divider [XLW^{+}20, LMT^{+}92]. Divine [Bat07]. Division [BR73, BC23, Bru22, CHI^{+}03, CL94b, EL94, EIM^{+}99, EIM^{+}00, HSE^{+}24, Int03b, IM99b, ITY95, JJ65, Kah99, KM97, Kor05, LM95, LA03, LF92, LD89, Mar04a, MIT^{+}95, Nan11, PB02, PB08, RJM^{+}20, Rus98, Sar59, SL97, Tra60, WG92, AHM^{+}98, ALB98, BMR04, CL94a, CL95, EL89b, EL89a, EM03, Fan89, GREL96, HTHR94, HF95, ITY97, KM93, KMB94, KP98, LM99, LKH93, LKKH99, MI97, MMW91, MM91a, MMH93, MC91, MC93, MC95, NL99, Obe99, OJ04, SF93b, Sch95, Ste89, Tak00, Tay81, TE06, WN03, ZXBZH01, Zim00]. division-free [HTHR94]. division/square [ALB98, LM99]. division/square-root [ALB98]. Divisionless [Sar60]. divisor [BMR04]. DL [Mor81]. DLMF [Loz00]. dlya [PT80]. do [Ped80, Wen09]. Docev [Ano84, Sal83]. Document [MHSK16]. Documentation [Rot70b]. Dokumentation [Rot70b]. Domain [PP12, Cro92, HTHR94, JT85]. domains [Ric92, Sau93, Shi93]. dont [Del81]. Doppler [WKG66]. Double [Cod64, DAV12a, DAV12b, God20, Kin65, LL09, LM03, PB02, SH72, Sca71, She76, Spe72, Swa65, ABF^{+}23, Ano14, GKK00, HHv^{+}73, Hon71, MS87, Paw11, Ped03, Tem86, TE06, WG94a, dDLM04]. double-index [GKK00]. Double-Length [SH72, HHv^{+}73, Hon71]. Double-Precision [Cod64, God20, Kin65, Swa65, Ano14, dDLM04]. Downward [GM04]. Driven [ASA06, Tan89a, Tan90e, Tan92]. Drivers [Cod93b]. drobno [Saa91]. drobno-ratsional’nykh [Saa91]. DSP [BJG^{+}99, BJG^{+}00, LP16, MKY92, Ste89, WMDM92, dDP10]. DSP-Enabled [LP16, dDP10]. DSP-Oriented [MKY92]. dual [KMY89, MI97]. dual-rail [MI97]. Dualmaschinen [Rot70a]. Dumka [Köl90]. duplication [Car79, Fuk13a]. durch [Sti53]. during [Bre00]. Dutch [Dop42]. Dynamic [BCD^{+}10, Mül90, Dem89, Krä88, MI97]. Dynamical [Mey01, Ise94]. dynamics [AGH89, Ise91]. dynamischer [Mül90]. Dyson [Hom96a]. Dzjadik [Kon76].

early [Dut84]. eBook [Lip21]. ECCTD [IEE89a]. Eclectic [Sri07]. Economic [SK71]. Economical [Cod67]. edge [Mit92]. edition [Fet81b, Köl97]. Editor [WTA^{+}90, Cat86, Cod67, Fik67a, Kin65, Knu63, McC90, Rev90, Swa65].

EDSAC [WWG82]. Education [LMS99]. effect [Lit93, Lit94]. Effective [BBC08b, HMO19, Kra85b, AM89, Jab12, KC11]. èffektivnom [DZKK77]. efficiencies [Jam89]. Efficiency [Zag16]. Efficient [Ahm89, AIS^{+}17, Bor95b, BCE^{+}08, BF23, CL94a, CVW06, DJ67, DZKK77, Flo15, Fow93, FR76, GST12a, GST17, GLT^{+}15, ITY95, ITY97, Joh14a, Joh15a, Joh87a, Joh20c, Joh20d, Kir88, KB86, LL09, LJ13, NKY08, PW90b, PLT23, Ric73, SW99a, SK11, Smi89, Spi61, Tan90a, WS05a, Woź10, Zag17, Zag24a, Zag24b, AQ18, Bak75a, Bor95a, CGKZ05a, CGKZ05b, Dan77, Ell83, Flo61, Has00a, HF95, Joh87b, LKH93, Mac97, MC89, dLSG18, SS23, Sun15, Tom00, Vep08, Wal84]. efficiently [NL94]. Effusions [Bat07]. Egypt [Wah04]. Eigenschaften [Tri50]. Eigenvalue [Ovt03]. Eigenvalues [SS65, Sti53]. Eigenwerten [Sti53]. Eignung [Völ83]. Einbettung [von80]. einfacher [von84]. einiger [Fil66]. Einschließung [Sch68]. Einstein [Din58, Mac97, NDT69]. Electrical [Wah04]. electrochemical [MO18]. Electromagnetic [LKL02]. Electronic [BR73, Kog57, Kog58b, Kog58a, Kog59, Köl90, Rob55, Wah04, WWG51, WWG57, WWG82, Har77, Tri66, WO15]. electronic-structure [Har77]. Electrostatics [Mul95]. element [LG79, Mon83]. Elementaarsed [Jür66]. élémentaires [Ber84, Laz85, Mul85b, Mul86]. Elementare [Töl43, Töl50]. elementaren [Rot71, Völ83]. Elementarer [Emd40, Emd45, Emd48, Emd59, Fil66]. èlementarnykh [Bar72, DZKK77, LČJ63, LP83, PS93, Saa91]. Elementary [ACG^{+}86, Ahm89, AFC10, Bak75b, Bak92a, BSY82, BBM84, BB84, BB97, BB00, BB04, Bra87, Bre75a, Bre76a, Bre80a, Bre06, Cod71, CW80, Cod91, Cod93a, CB87, CVW06, Dam94, Dav81, DeL70, Dri89, Ekl01, Emd40, Emd45, Emd48, Emd59, Erc72, Erc73, ELMT00, ECSD20, Far81, Fil66, Fra78, Frö61, Frö63a, Gal85, GB91, GL83, Gon52a, Gon52b, Gus86, HKA93, Hul78, HFT94, JGB13, Joh15a, Jür66, Kah82, Kah87, KDDH92, KDDH94, KB90, Kog66, KZ90, Krä87, LM03, Leh96, Liu87, LJ13, dL70, LCY65, Mar72, MDM16, Mat87, MLS^{+}18, MCM90, Mor64, Mül90, Mul97, Mul20, NS09, NKY08, Nis94, OTY91, PW76, Ric68, Rit25, Rot70b, Rot71, dLSG18, Sal92, SS94, SS99a, Spi61, SDP11, Tan89b, Tan90c, Tan91b, Tho87, Tur72, Völ83, Wan74, WG91, WG94b]. Elementary [Ziv91, Zuc64, dDLM11, von80, von84, ACG^{+}87, AM78a, Ano14, Arm82, BKM94, Bak73, Bak75a, Bar72, BBD03, Ber84, BL71, Bor88, BCD^{+}02, BD02, Bre75b, Bre76b, Bre10, Bro89b, Bro89a, Bro90, BJLP19, Car69, Car90, Che81, CA00, Cle54, Cod82, CDJ^{+}01, Dan77, Das94, DMMM95, Dri91d, Dri91a, Dri91c, DM88b, DK77, DF78, DZKK77, Eps75, EC79, FLMR00, Fer86, Fog51, Gal86, GM84, Hai60, HHv^{+}73, Her69, Her71, Hit75, Jam81, Joh14a, Kai89, Kaz65, Kog60, Kon76, Kra85a, Kra85b, Kui52, Laz85, LM07, LN22, Lin57, Liu93, LČJ63, Loe01, Lóp18, vdD88, Lut95, LP83, MS67, Mak65, Mak80, MT72, Mar81, Mar90, Mar00, Mar03, McC77, Mer94, MY91, MV98, Mul85a, Mul85b, Mul86, Nag68, Nis84, OGWY94, Olv80, Peĭ63, PKP74]. elementary [PS93, Ric88, Ris79, KR86, Saa91, Sai89, Sal76b, Sg92, SS93a, SS93b, Shi93, Sho60, Sim64, Smi89, Spe72, Squ91a, Squ91d, Squ91b, Tan91a, Tes69, Töl43, Töl50, Tre80, Wal71, War60b, WG94a, WG95, XLX^{+}83, XL94, ZSA86, dDM06, PBM86a, Töl43, Töl50]. Elementary-Function [BBM84]. elements [Mus87]. ELF [SG99]. Elfun18 [Bat19]. elimination [EH16]. Ellipse [Adl12]. Ellipsoidal [Hob31, Hob55]. Ellipsoids [DR94a, DR94b]. Elliptic [AVV90, AVV92, Ars62, Boe61, Bow53b, Bow61, BF71, Car77a, CN81, Car87, Car88, Car89, Car92, CG94, Car99, Cod65a, Cod65b, Cod66, DH59, Eck80, Ehr61, FL67, Fet65, FC64, Fra65a, Fuk15a, GW71, Gla71, Gre92, Gre59, Hol90, Kin24, Kin07, Lan60, LW63a, LW63b, Lóp00, Luk68, Luk70, Mas75, Mer62, Mey01, MT64a, MT64b, NC66, PS05, PT90, QV96, QVV98, Sch88a, Sch85, Sch15, Sid95, Sko78, Sou64, SMCH41, Tka03, Van69, War60a, Wri73, ZC70, Ali11, Bat19, BZ92, Boy15, Bul65a, Bul65b, Bul69a, Bul69b, Car79, CG85a, Car91, Car95, CF00, Car05, Cri89, DCL^{+}92, Eck77, Far69, FGG82, Fet81b, Fet81a, FI94, Fuk09a, Fuk09b, Fuk10, Fuk11, Fuk12, Fuk13a, Fuk13b, GMMIDM13, Gra02, Hay56, Her61d, Her61e, HvdR63]. elliptic [Jef62, Köl97, Lan80, Lar66, Mey63, Mid75, MH72, Mor78, Mor99, MO18, Nis94, PDK96, Paw11, Sch76a, Sch77a, Tha63a, Tha63c, Tha63b, Töl67b, Töl67a, Töl69, Töl70, VM87, VPK99, Wac00, vdR63]. Elliptical [DiD90, Töl67a, Töl67b]. elliptically [Was89]. Elliptische [Töl67b, Töl67a]. elliptischen [Töl69, Töl70]. Eloquent [Adl12]. embedded [Ste08]. Embedding [von80]. Emden [Rut51]. Emphasis [BS98, Buc69]. empirical [Ruy89]. Enabled [LP16, LP17b, dDP10]. enclosure [Sch68]. encoding [ZXBZH01]. Encyclopedia [IA20, KS20, MS03]. end [Fet74]. energies [BS84, TR84]. energy [Bar84b, LVV05, NT84, Sea84, Sti88]. engine [CK89a, Lov43, Lov89]. Engineering [PC11, RW04, Wah04]. Engineers [Bel04, BF71, Lau95, Lau04, McL34, McL55, And98, Bar14, Bel68, Fet81b, Köl97, Sen67]. England [KM95]. English [AEF92, Hom96b]. Enhanced [BHK09, ZCL15, BR94]. Enhancement [Boi97]. ensembles [AEF92]. Entire [BDG^{+}72, BDG^{+}84, Zal89]. Entwicklungen [Sch80a, Töl68]. Environment [CJL10]. environments [ISO10]. Episode [Ful99]. epsilon [GS14c, HL13]. Epstein [AB93]. Equal [KP75]. Equation [Bra65, Cle69, Flo15, Fri72, Hom96a, Mil57, FLO04, GST04b, GS12, Mon83]. Equations [AO08, Cas83, Cow64, FK19, HBF09, HBS00, Lir71, Olv67a, Ars68, BBB89, Boy14, EM94, Hor19, Luk56a, Luk56b, NS13, Olv67b, PS77, PS78, PS79]. Equator [OMS09]. Equianharmonic [Sou57, T.57]. equivalence [Her71]. ERF [HJ67b, Che09]. ERFC [Che09]. erfinv [Jar12]. ergodic [Kra89]. ERL [GEK04]. Erlang [Whi63]. Errata [Ano60b, Ano84, Coh12, Luk56b, Fet81b, Fet81a, Köl97, Sha76]. Erratum [Amo78, Frö63a, LW63a, Szm13, Tho04, VRS^{+}99, Wre73]. erreurs [Del81]. Error [Alz10, ABJ19, Bai17, BS79, Bre83, Bre99, Car63, Car75, CCM11, CR68, Chr65, Cod69, Cod90, DJ62a, DD07, FCC73b, Fra78, Gau61, Gau64e, Gau69a, Har66, Has02, How22, Iac21, Jan11a, Jan11b, KP69, Köl87, KK20, Krä87, Kuk72a, Kuk72b, Lip21, LB09, MR71, Mil89, MM91b, Mül90, NSR10, Nem13, Olv88, Par02a, Plo01, PW90a, PW90b, Sal51a, Sho02, Str68, Tan91b, VM19, WD61, Woo67, Zag19, AQ18, Akr85, Bel90, BEJ76, BJ15, CST18, CD02, CDS03, Del81, EL08, FCC73a, Fet74, GK70a, GLM15, HJ67b, HR72, JM17, Joh12, KW21, Köl72a, Kra98, Kra14, KT59, LS88, Lit93, Lit94, Mag94, Mat90a, McC74, Phi79, QXG18, RM07, Sal55, Sch78c, SK08, Van89, Wil96, Yun14, Zag24a, Zag24b, Zak69, Ziv99, vdLT84, Alb62, HJ67b]. Error [SZ70]. Error-Like [Köl87, vdLT84]. Error-Optimized [ABJ19]. Errors [Bra65, CJL09, KP75, Bor10, Wen07]. ESF [MS03]. essays [War08]. essentially [LP62]. est [Del81]. estimate [McC74]. Estimates [Fra78, Krä87, McC84, Ovt03, Per85b, Per87, QV96, BD85, Bre99, CG89, Che13a, Mar85]. Estimation [CKL89, Dem89, GK89a, Mat90a]. Estimations [Ele15]. estimators [LW89, Wan82, vR89]. Estonian [Jür66]. ETAPS [Bor02]. étude [Bre78c]. Euclidean [HH18]. Eugene [Tal68]. Euler [AL20, Bau89, Bor10, BM80, CM13a, Che16a, Che16b, Chl11, Dav59, Fre05, HG64, Kar01, Knu62a, KB67, Lin79a, Lu14, LSY15, Rza̧12, Swe63, XC17]. Eulersche [Bau89]. European [IEE89a]. Evaluate [WMBE14, McC77, SdCR97, SG99]. Evaluating [BCE^{+}08, Dan82, Fik67b, KZ90, Lin79a, LS95, Mar04b, Wac00, Whi63, FLMR00, FGG82, NT84, Oli75, Mul82]. Evaluation [AM20, AMT78, AJMT81, ABJ19, Bar61, BSY82, Bre75a, Bre76a, BEM12, Bur74, CET62, CG02, CM13b, CR68, Cod90, CP79, DeL70, DM88b, Eis63, Erc72, Erc73, Erc75, Erc77, ECSD20, Far81, Fik67b, Fik68, Fra65a, Fra78, FR76, Ful80, GA08, Gau77b, Gau16, GST02d, GOST23, GJL20, Gla66, GW71, GN89, GBS19, GM04, Hay56, HS20, Jab94, JKMR11, JGB13, JF16, Knu62b, Knu98a, Kuk66, KR79, Lan60, LL72, Lin78, Lin79b, LW82, LB09, dL70, Mar72, Mar88, MTP82, MR65, MM88, Mot55a, Mot55b, NC66, OJL22, PS11, PW90a, Ric73, dLSGDR17, Sch78a, SG76, oKSAG79, ST74, SC71, SF87, SG72, Tak66, Tem75a, Tho66, Wan78, WOG95, Zag17, Zag19, Ziv91, ABF^{+}23, BC79, BDGP13, BBC^{+}14, Bar82a, Bar82c, BS06, BO75, BZ92, BBC08a, Bre18]. evaluation [Bre19, Bre75b, Bre76b, BHMZ24, Bun09, CY18, CGKZ05a, CGKZ05b, CLM88, CJR91, CJR92a, CJR92b, DF84, Di 86, Dor62, Dun90, DM88c, Ein79, EW63, Eve64, Fis73, FW85, GS81, GEK04, GST20, Gra63b, GR19, GM07, HF09, Has02, HR72, KW21, LVV05, LGML05, LO94a, LO94b, Mac97, MT72, MZM94, Max91, Mil89, MN16, NSU20, OGWY94, PKS00, Par09, Pas95, RW72, SGA81, SZ77, Smi89, Tem76, Wal84, WO15, Wen64, Wil71, WG94a, WG95, Zai93, Cun62a, Cun62b, Cun62c, Knu63, LDP93]. Evaluations [GS82b, Lav86]. evaluator [PBN93]. Even [ARH14]. evolution [Dev12, War08]. Evolving [PUHM12]. Exact [Lef05, Pom74, Pom76, dLSGDR17, Sau92, SS93a, BWKM91, CZ10]. Exactly [SS94]. examine [Wil96]. Example [Bol06, Plo96]. Excellence [Lid01, Lid02]. Exception [HFT94]. Exchange [ABJ19, SR81]. Execution [Bai93]. Exercise [BB12]. Existence [Som82, Wan74]. existing [FKY86]. exp [BBC08a]. Expansion [BH63, FO99, Kim72, Luk59, LC61, NDT69, PL19, Pie84c, Sch64, Tho65, Van69, Won73a, Won88, Wyn60, AS97, AS03, Ali11, CG85a, Dij77, Ehr89, Fet79, GST03a, GS14c, HL13, Kal04, KWY07a, KWY07b, LW14, MMV81, McC74, Nag01, NP18b, Nem13, Par02a, Par02b, Par03, Par16, PHH08, Sch76a, Sch77a, Tem79b, Tem81, TC01, Won73b]. Expansions [Boe69, Cod65b, Cur64, Din58, FW61, GHT03, Hom98a, Hum64, Lóp00, Olv91, SR53, She74, She76, Sle65, Tha61, Wim62, WR71, Woo67, AS05, BF92, Bor10, BJLP19, CE15, CP15, Che15a, Che16e, CL16, Che16a, Che16c, CT16, CC17, CM83, CB87, DiD84, Din73, Dom03, DGS18, Fuk12, Fuk14a, Har00, Hum85, Kat78, Lan80, LT13, Lóp18, Luk75a, Luk76, Mac72, Nag04, ND19, Ném65, Rza̧12, Sch78c, Sch80b, Tem75b, Tem78, Tem83c, Tem85, TV09, Wim72, XHT16, XC17]. experiment [PIH14]. Explanatory [Ful81b]. Explicit [Buj75, McC84, RB75, GF97, Kra14, OG82]. explicitly [SE14]. Exploring [Sho02, SL97]. Expm1 [Tan92]. Exponent [BHK09]. Exponential [AAHTH10, AS10, Bee99, CET62, Caw00, CLOT86, CT69, Cor61, Fri72, FK11, Gau59, GC64, Gau73, God20, HJ96, HA86, Kar84, Köl87, LP17b, LSZ08, Mil85, Ng70, OJL22, PS11, PEB04, Red70, Sch99, SMG94, Smi11, SZ74, SZ76, Sto41, Tan88, Tan89a, Tan90a, WLK11, Zuc64, BSW95, Bea60, Bor10, Bra84, BT99, Che71, CD02, CDS03, CLM88, CO80, CT68, DR04, Dri95, DR89, Fog51, GHT03, GEK04, Har00, Has02, Hit74, IP81, JMS98, Kan96, KSVW07, Luk56a, Luk56b, Mag94, MNP83, NP18a, Nem13, Ng84, Olv94, PBE03, RT61, Ran77, Sab08, SZ77, Sti53, WdZ04, XC14, XC15, dDP10, vdLT84, Her61b, Pac70]. Exponential-Type [FK11]. Exponentially [Olv91, PW92]. Exponentially-improved [PW92]. Exponentials [Che72, ZLA13]. Exponentialsummen [Sti53]. exponentiation [CC69]. Expression [Cun62a]. Expressions [BFHT85, MDM16, Sho60, SF10]. Extended [CQSP04, SOL81, KPPS08, KA71]. extended-precision [KA71]. Extended-Range [SOL81]. extends [GP84]. Extension [ARH14, BKdLSGL23, GZC22, Hea65, Rus88, Bul69a, Dav35, Ge89, Kai89]. Extensions [Ein89, ISO10, Sle64, Ker83]. Extracting [Esc37, Pen81]. Extraction [BR73, Lan42, MK71, Maj72, Mur88, Mur99, Sug58, BDGB83, GREL96, Hay99, Joh11]. Extractor [HAK91]. extracts [Fre81, Lov53]. Extrapolation [Bre82a, BR91, GSS12, Sid03, Wal96, Bre80b, Håv79, Håv81, Saf10]. extreme [YM03]. Extremely [TKK17].

F [Ars62, Coh12, Fet81b, Köl97, Leh83b, Szm13]. Faber [Ell83]. Factor [TAY91, Air37, Car91]. Factorial [BC18, Bor10, Fer07, Wen10]. Factorization [JF16]. Factors [Car92, Din58, Sch84, CF00]. Faddeeva [Joh12]. Faddeyeva [Zag16, ZA11, Zag17]. Fading [LB09, SK11, SA00, SA05, CDS03]. Failure [KI04, Wan82]. Faithfully [GLT^{+}15]. Fall [AFI69]. Familiar [BB88]. Families [AS10, FW13]. family [NS13]. Fast [AM20, Ahr96, Ano12, Bee99, BY19, BS06, Bor21c, BB84, BZ92, BB97, BB00, BB04, Bre75a, Bre76a, Bre06, BH13, Caw00, CH89a, CH89b, CJL09, Fuk09a, Fuk09b, Fuk10, Fuk15a, GST11b, GLR07, Gus86, HT13, Har09b, Hei96, HJ96, HMO19, ITY95, JGB13, Joh14b, JF16, JM18, KB90, KP98, Mar05, MWH^{+}16, MWH^{+}18, NP18a, RW72, Rol87, Sch99, Sor94, Tak22, Tro84, WMC21, WG91, WG94a, WG94b, WG95, Ziv91, Bak76, Bun09, Bur82, Cra06, DMMM95, Fuk11, Fuk13a, Fuk15d, Fuk15b, Fuk15c, GEK04, Has00b, HJZ23a, HJZ23b, HJZ^{+}24, Lyo91, Mar03, Mit92, OGWY94, PKS00, Phi79, PIH14, Sch96, Tan90b, Tom00, WMC18, Zim00, ZA14]. Fast-Start [Mar05]. Faster [Bow53a, Fil92, LKK19, DM91]. fatigue [Sob88]. Favard [GS82a]. Feasibility [Dun90, Via89]. February [Koo91]. feedforward [HSW89]. Fehlerabschätzungen [Krä87]. fehlerhaften [Bra65]. Fehlerschranken [Alb62]. Fejér [Mag94, VT11]. Fermi [CT67a, MN16, BDM81, CT67b, Din58, Fer86, Frö63b, Fuk14a, Fuk14b, Fuk15d, Fuk15b, Fuk15c, GOST23, Goa95, Let01, WR63]. festschrift [Zah94]. Fibonacci [RS98]. Field [LE95, Ngu98, Ehr89, LE93, OSS09, Töl69, Töl70]. Fields [ARH14, HCK09, Mül04, Sch85, KCYL06]. Fifteenth [Plo96]. Fifteenth-Century [Plo96]. Fifth [Ded26, SC01]. Filter [MKY92]. find [Che81, LG79, NS13]. Finding [AMM21, CH70a, Bre75b, Bre76b, Jam89, KI04, Lyo91, Vai89]. Finite [BM15, Bus74, HCK09, Lew69, Mül04, Sch85, Fog51, Hai60, KCYL06, LG79, LU96, Mon83, OWS^{+}14]. First [Ban01, Bar10a, Bea58, Bra70, Car77a, Cod83, DH59, ES99, FL67, FK11, Gau64a, Jab94, LW63a, LW63b, Pie84c, Ree90, Sko75a, Van69, Vel11, BL96, BZ92, Boy15, CG85a, Cha82, Cob69, CV03, DAV12a, DAV12b, Fet81a, Fuk10, Gau65b, Her61d, HvdR63, ISK87, IS88, LW96, LU96, Mor99, Sca71, SG99, Tha63c, Wal84, YM97, Zai93]. First-Order [FK11]. Fisher [Ano60b]. fits [Ran77]. Fitting [Sto41, BCG91, Phi79]. Five [MW83]. Five-Function [MW83]. Fixed [HHR00, IP17, MLS^{+}18, dDIS13, dDI15, LBdDM16, Vav89]. Fixed-Point [IP17, MLS^{+}18, dDI15, dDIS13, LBdDM16]. FL [Lip21]. Fleshing [MS16]. Flexible [IP17]. Floating [AAHTH10, Bli97, Bra87, CH98, D^{+}89, DHL^{+}04, GB91, Has90, Hul78, JKMR11, JMMW79, Kah80, KMB94, LP16, LP17a, LP17b, LM15a, LD89, MMNP91, Obe99, Rev16, Rus98, Sch88b, Smi91, Ste08, Tan89a, Tan90e, Tan92, WS05b, dDP10, dDLM11, BBC^{+}91, BWKM91, BCK^{+}93, Bol06, BMR04, CG24, CHGM99, FHL^{+}07, GLM15, Hou81, HS20, JM17, Kra98, Kul07, LBdDM16, LKH93, LC97a, LMOT01, LN21, MT72, Mat90b, OY91, OJ04, PKS00, Rus99, SF93b, Smi01, SL97, Ste89, TE06, TRHH94, VCV01, WN03, dDM06]. Floating-Point [AAHTH10, Bli97, Bra87, CH98, DHL^{+}04, Has90, Hul78, JKMR11, JMMW79, Kah80, LP16, LP17b, LM15a, LD89, MMNP91, Rev16, Smi91, Tan89a, Tan90e, Tan92, WS05b, dDLM11, dDP10, BBC^{+}91, BCK^{+}93, Bol06, BMR04, CG24, CHGM99, FHL^{+}07, GLM15, Hou81, HS20, JM17, Kra98, LBdDM16, LKH93, LMOT01, MT72, OY91, OJ04, PKS00, SF93b, Smi01, SL97, Ste89, TRHH94, VCV01, WN03, dDM06]. Florida [Men06, Ric92]. flous [AEF92]. flow [Joh83]. flows [Fei78]. fly [EL89b]. FNLIB [Ful81b, Ful81a]. following [Laz85]. fonctions [Ber84, Laz85, Mul85b, Mul86]. Force [Fet81a]. Forensic [BB12]. Form [AMS76, GW71, Jan11b, CTV96, Dav88, FI94, Jan11a, Mar95, RM07, Tem85, Wal84]. Formal [Har03, MRH19, Rus13, Saw02, dLMF23, Bol06]. Formally [BDL09, Wyn60]. Format [LSZ08, MMNP91]. Formats [Krä87]. Formeln [Bar14]. Formelsammlung [Sne63]. Forms [Ric65]. Formula [Adl12, Chi20, Sau92, Spi71, Sun71, CL16, FNC75, FO93, GF97, Har01, KC11, Kza99, LFM14, LSM14, LSM15, LM15b, LSM16, Mak80, Mei83, Qi14, Sch77a, Sne63, Wei98]. Formulae [Kin21, KP03, Wim68, Zwi03, Akr85, Ben98, Jam89, RX07, Ver77]. Formulas [AS64, Boe61, Buj75, Car70a, Coh12, DH59, Gas81, Jef95, Jef00, Jef04, JD08, LW63b, MO49, MOS66, RB75, Sal51a, Ske86a, Ske86b, Bar14, Che13b, Che15a, Che16e, CS11, FKY86, Kza92, Let96, Mor11, OSS09, Sim64, WCS16, XC15, YC15, LW63a, Wig67, Szm13]. Forth [Koo91]. FORTRAN [WB78, Wan80, DG99, HH18, Amo83a, Amo83b, Bai93, Cod93b, DM87, HKA93, KDDH92, Kuk66, KA71, Sca71, Smi91, WB78, BD14, CZ10, Col80, Col84, DG99, EC13, GST02b, GST02c, GST12b, GN89, KDDH94, Mil04, NSU20, SSG^{+}18, Sht19, Smi01, Tho97a, WNO94, Wie99, Zag16]. Fortran-77 [KDDH94]. Forum [IEE99]. Forward [LDP93, MRS17, SR16]. foundational [Fer07]. Foundations [IEE89b]. Four [Boy15, Cus83, DMS95, HFS99]. Four-Function [Cus83]. four-quadrant [HFS99]. Fourier [AO08, Boy89, Boy01, Boy09, CZ95, HBF09, Hom98b, LP61, LP62, Lan80, MS87, Ole96, Sch96, SP61, Sle64, Sle78]. Fourth [Atk79]. Fox [L.61, MM95]. FPGA [DR04, TVG00]. FPGA-based [DR04]. FPGAs [IP17, LP16, LP17a, LP17b, LC97a, PLT23, WN03, dDP10, dDIS13]. FPU [AHM^{+}98]. fractals [PH90]. Fraction [Bar74, She54, Smi95, Spi61, Bar82a, Cha80, Che13a, Dij77, Gau77a, Kra89, Lee92, Len90, LSY15, LM15b, LSM16, MM90a, McC74, McC83, SdCR97]. Fractional [Ano46, Ant64, CMF77, GCP67, MG85, PT91a, PT91b, Sai89, VM87, ZZ96, Abr45, CP66, GM88, KT00, KRT02, MG89, MPG92a, MPG92b, Nis84, SO89, TV09, Tim87]. Fractions [CPV^{+}08, Len76, Mor64, Pas03, SB71, Wal00, BC09, BHJ05, BZ92, BS78, BS83, Hau88, JW88, Lan80, Lem88, Lev91b, Lev91a, Now06, Pas92, RdC93, Sha76, Spe72]. framework [Per85a]. France [Ano03, Ano06, Bor02, KM91, MTV15, KM07]. Francis [Lip21]. Franklin [NP18b]. FRB [CZ10]. Free [ES10, LF92, CZ10, HTHR94, Joh12, LLK14]. free/open [Joh12]. free/open-source [Joh12]. French [Ber84, Bre77, Bre78c, Che88, DGB80, Del81, Laz85, Mul85b, Mul86]. Frequency [Hil73a]. Fresnel [Sny21, Act74, And00, Chr62, Cun62a, Cun62b, Cun62c, Fle68, Gau64e, Gra63a, Gra63b, Gra64, LG64, Ném65, Sny93, Sny96, ZA24]. FRG [MT85, MT86]. Friedman [Fet81b, Köl97, NP18b]. Friendly [BEM12, WMBE14]. friends [Bir54]. Fritz [Wig67]. Fukuoka [Ano91]. full [AHM^{+}98, Bol06, LKKH99, Spr00]. Function [AS95, AM20, Ahm89, AAHTH10, Amo83b, AK09, AWH13b, Bai17, Bak92b, BTT16, Bar61, BY07, BJG^{+}99, Bee17, BM15, BBM84, BH18, BP67, BS79, Bor95b, BBC98, BBC00, BCE^{+}08, BEM12, BI86, CET62, Car63, Car75, Caw00, CCM11, CM13b, CR68, Chi20, CH67, Cod69, CST73, Cod75, Cod93b, Col78, Cus83, Dav59, Dav64, DHL^{+}04, DJ62b, DJ62a, DJ67, DM86, DM87, DM88a, DiD90, DM92, DD08, Eck76, Ekl01, ES99, ECSD20, FLR03, FCC73b, Fik67b, FK11, Ful72, Ful77, GB68, GZC22, Gau61, Gau64e, Gau69a, Gau77b, Gau16, GST02d, GST12a, GST14a, GRAST16, Gil16, GZ75, Gla66, God20, GBS19, Gus86, HP62, Har97b, Har98, Har09b, Har66, HG64, Hea65, Her61f, Her61a, HJ96, Hil77, How22, HA86, Hum64, Iac21, IP17, IR08, Jan11a, Jan11b, Jar12]. Function [JKD92, Joh21, KL07, KDDH92, KDDH94, Kim72, Kol70, Köl72c, Kol72b, Köl87, Kor11, KK20, Kuk66, Kuk71, Kuk72a, Kuk72b, Laf84, Lan64, LSZ08, Leh44, Let01, LDP93, Lin79a, Lin72, Liu87, LBC11, LT71, dL70, Luk59, Mac89, Mar65, Mat04, MR71, MLS^{+}18, Mil85, Min70, Moo67, Mor80, Mor76, Mor79, MW83, NSR10, NF63, Ng75, OMS09, Olv91, OM68, Pag77, PW76, PH66, POMB05, PW90a, PW90b, PLT23, Raf70, Rei69, Sal51a, Sau92, ST07, Sch64, Sch16, Sch78b, Sch76b, Sch77b, Sch99, SS97b, SS97a, SG02b, SB71, SK11, Sou57, Spi71, Sri07, SLZ05, SS99b, Sto41, Str68, Sun71, Tak66, Tan88, Tan89a, Tan90c, Tan90e, Tan92, Tem75a, Tho87, TKK17, Vel11, VM19, WS05a, WEX14, WMBE14, WC61, Wis48, WG91, WG94b, Woo67, Wre68, WLK11, Zag17]. Function [Zag19, dDLM11, AJ03, AS97, AS03, AQ18, Ali11, Alz93, Alz97a, Alz97c, Alz98, AR99, Alz03, Alz08, Alz10, ABF^{+}23, AG08, AM78a, AB93, BC79, Bad76, BB15, BSW95, BGV81, BS00, BP01, BEJ76, Boe91, BS06, BF92, BST13, BO75, BZ92, Bor95a, Bos89, BS78, Bra84, BCJV18, Bra64, Bre75b, Bre76b, BHMZ24, Buc69, Büh87, BT99, BK16, CS19, Cau22, CST18, Cha80, CTV96, CB09, Che13a, Che13b, CP15, Che15a, Che15b, Che16e, CL16, Che16a, Che16d, Che16c, CT16, CD02, CS11, CO80, CHT71, Cod82, Col80, CM83, Col84, CPG12, CB87, CS82, CS83, CC64, Cyv64, DAV12a, Del79, Del84, Dev12, DB13, DH76, DiD84, DC81, Dri95, Dun90, Dut84, Eck77, EG97, EL08, Ele15, Eps75, EMR82a, FW13, FNC75, FLS06, Fet63, FCC73a, Fet74]. function [Fis73, FK75, For97, FW80, Fra81, FW84, Fre09, Frö63b, Fuk15d, GA08, GP64, Gau64d, GS78, GS88, GK89a, GST03a, GST11a, GST11b, GST12b, GST15a, GLP98, GL01, Goe14, Gon52b, GR19, GP84, GQ09, HXX22, Har20, Har01, HF09, HHR00, Her61c, HJ67b, HJZ23a, HJZ23b, HJZ^{+}24, Hum85, HR72, IS88, IP81, Jab12, Jab13, Jab20, Jam81, Jam99, JHC96, JS89, Joh15b, Joh20a, Jon70, JMS98, Kai89, KW21, Kar01, Kat78, Kaz65, Ker83, KPPS08, KRT02, Köl72a, Köl84, Köl96, KT59, KSVW07, Kuz15, LW96, Las82, Lee97, LGML05, Leh70, Leh56, LG79, Lew44, LKF10, LN22, LW82, LT13, LFM14, LSM14, LW14, LSM15, LM15b, LSM16, Luk56a, Luk56b, Mac94b, Mag94, MS16, Mar03, Mar86, MZM94, MG85, MG89, MPG92a, MPG92b, MJ13, MRS18, McC83, MNP83]. function [Mer05, MM88, Mit49, Miy19, Moo81, Moo84, Mor11, Mor14, MS15, Mul01, Mul06, Nag01, Nag04, Nem13, ND19, OY91, OSS09, Pak18, Pal98, PDK96, PW92, Par02a, Par02b, Par03, Par72, Ped03, Per85b, Per87, Phi60, Phi79, PUHM12, Pie82, PB83, Pie84a, PHH08, Pit67, Pow88, Pré10, Puo88, Qi14, QM15, RL76, Ron86, Rza̧12, Sab08, Sal89, Sal55, Sch88a, Sch80a, SK08, SLH06, Sho86, Sko04, Sko05, SE14, SW50, SC81, T.57, Tak22, Tan91a, Tem75b, Tem76, Tem83a, Tem86, Ter81b, Tha62a, TC01, Tri50, Tur43, Tur53, Uni49, Vai89, VC06b, Van89, Veb12, Ved93, WV20, WdZ04, Wan16, WO15, WC90, Wen96, Wim74, Wri84, XHT16, XC17, YC15, YCZ15, YA24, YM97, Yun14, Zag24a, Zak69, ZB95, Zha95, ZB97, ZLA13, Ziv99, vdLT84, Hol69]. Function [Mar85, Wre73, Lip21]. Functional [AVV90, AVV92, Deh89, Kog89, Liu93, Töl43, Töl50, Töl66, Töl67b, Töl67a, Töl68, Töl69, Töl70, vRdL88]. functionals [Deh89]. Functions [Abr64a, AS64, Abr64b, ACG^{+}86, AO08, AMM21, Alh00b, AFC10, AGL93, Amo73, ADW77a, ADW77b, Amo78, Amo83a, Amo86, Amo90, Amo95, AVV92, ASA06, AMT78, AJMT81, AAR99, Ano46, Ano48, Ano52, Ano60a, Ano74, Ano22, Ant64, AH16, Arc43, AM77, AWH13a, Arn77, Ars62, AKS01, Bak75b, Bak92a, Bal00, BDG^{+}72, Bar10a, Bar10b, BA44, BD99, BJG^{+}00, Bat07, BKdLSGL23, BSY82, Bea58, Bee99, Bel04, BCD^{+}10, BK78, Ben83, BH65, Bin68, BC62, Bla64, Boe62, Boe69, BL02, BB84, BB88, BB97, BB00, BB04, Bow53b, Bow58, Bow61, Bra87, Bra70, Bre75a, Bre76a, Bre80a, Bre06, BL45, BLP24, BF23, BdDKL15, BGVHN99, Bur63, Bur64, Bur74, Bus76, Cam84b, Car70a, CG55, Car70b, Car77b, CP98, CD08, CHG^{+}11, CMW63, CLOT86, CH70a, CP70, Cod71]. Functions [CMF77, CW80, Cod83, Cod90, Cod91, Cod93a, CU59, Cur64, CVW06, CPV^{+}08, CBBV10, D^{+}89, Dav64, Dav81, DeL70, DGA18, Deu76, Din58, Div79, Dop42, Dor66, Dri89, DM88b, DD07, Eck80, Emd40, Emd45, Emd48, Emd59, Erc72, Erc73, Erc75, Erc77, ELMT00, FO99, Fab04, Far81, Fat16, FK19, FB91, FW61, Fie65, Fik68, Fil66, Fla57, Fle68, Fow93, Fox54, Fox60, Fra78, Fra65b, Frö55, Frö61, Frö63a, Fuk15a, Ful80, Gal85, GI01, Gau64a, GC64, Gau65a, Gau66b, Gau75, Gau79a, Gau79b, Gau99, Gel51, GST03b, GKS04, GST07a, GLR07, GT58, GT59, GCP67, GN89, Gre92, Gre59, Gun67, GM04, HKST99, HBF09, HJ67b, Hin77, Hit57, Hoc71, Hoc86, Hol90, HKA93, Hul78, HFT94, Ike75, Ike76, IK05, IA20, Jab94, Jan77, JGB13, Joh14b, Joh15a]. Functions [Joh19, Joh20b, Kah82, Kah87, Ken14, Ken21, KB90, KKI67, Kin24, Kin07, Kod08, Kod11, Kog66, Köl87, Köl90, KS20, KZ90, Kor02, Krä87, KL14, L.61, LaF54, LK21, Lar69, Lau73, LM15a, Leb65, LM03, Len76, Lew58, Lew81, Lew69, Lew75, LDP93, LKL02, Lin78, Lin79b, Lir71, LJ13, Low64, LO93, Loz94, Loz96a, Loz96b, Loz97a, Loz97b, Loz00, Luk59, LC61, Luk64, Luk69a, Luk69b, Luk71a, Luk71b, Luk72, Luk75b, Luk77, LASC95, LCY65, Mac65, Mar72, MDM16, MC87, Mas75, MH08, Mat87, MR71, MS68, McC02, MTP82, McL34, McL55, Mec66, Mec68, Mey01, Mil64, Mil65, MT64b, Mor64, Mos72, Mos89, Mül90, Mul97, Mul20, Nad15, NS09, NKY08, Nas74, NPA16, Nav83, NDT69, ND70, Obe64, OJL22, OTY91, OMS09, Olv64, Olv67a, Olv74, OLBC10]. Functions [Osi13, Pai19, PL19, PS05, Pie84c, PP12, Pot14, PT91a, PT91b, Pri75, PGA18, QN18, Qua18, Rai60, Rap94, Rap14, RS81, RL80, Rev16, Ric73, Ric68, Rot70b, Rot71, Rum01, dLSGDR17, Sal92, Sal51b, Sal53, SS94, SS99a, oKSAG79, Sch15, She76, SF59, SP20, Sht19, SC01, Sid95, Sko75a, Sla64, Sla66, SL07, Sle64, Sle65, Sle78, Smi11, Sne56, Sne61, Sne80, Sol69, Soo73a, Soo73b, Soo73d, Soo73c, Sou64, SS70, SO87, Spi61, Spo94, Sri70, SJ01, SA57, Ste64a, Ste64b, SZ70, SZ74, SZ76, ST99, Str56, SG72, SDP11, Szm13, TS69, TTW79, Tal68, Tan90a, Tan90d, Tan91b, Tem79a, Tem92, Tem94b, Tem96a, Tha79, Tho93, Tho97b, Tho97a, Tho13, TS76, VS04, VS10, V^{+}75, VB04, Var95, VVA09, Vel11, Ver67, VK95]. Functions [Vla02, Völ83, Vos73, Wan74, Wan78, WG89, Wat66, Wat95, Whi63, Wig67, Wil96, Wim62, Wim68, Woj61, WR71, Wyn60, XD18, Yos97, Zag16, ZS64, ZJ96, Ziv91, Zuc64, von80, von84, AS97, AS05, Abr45, Ada98, ACG^{+}87, Air37, AJDG02, Alh00a, AIS^{+}17, API^{+}19, Alz97b, Amo74, ABFP24, And98, Ano93, Ano98a, Ano14, AM78b, AM79, AM84a, AM84c, AM84b, Arm82, Ars81, Ars68, AN68, Ask75, BDGP13, BDGP14, BC09, BBC^{+}14, BKM94, Bak73, Bak75a, BL96, Ban01, BDM81, BDG^{+}84, Bar72, BFSG74, Bar81, Bar82b, Bar82a, Bar84a, BFSG84, Bar84b, Bat19, BW10, BS13, BBD03, BB82, Bel68, BS84, Ben98, Ber84, BG84, Ber01, BPT76, Bla74, BEJ78, BL71, BHJ05, Boo55, Bor88, BBC08a, BS83, BCD^{+}02, BD02, BPD81, Bre18, Bre19, Bre10, Bro89b]. functions [Bro89a, Bro90, BR94, Büh92, Büh03, BJLP19, Bul65a, Bul65b, Bul69b, Bun09, Bus74, BIS01, BK16, Cai11, Cam79, Cam81, Cam84a, CL11, Car69, Car05, Car90, Car82, Cer07, CKT07, Cha82, CZ94, CZ95, CQSP04, Che71, Che81, CKL89, CIM94, Che11, CB12, CE15, CP15, Che15b, CA00, Che09, CH78, Cle54, CP66, Cob69, Coc65, CDS00, Coh12, CCV11, CJR91, CJR92a, CJR92b, CDJ^{+}01, CV03, Cri89, Dam94, Dan77, Das94, DMS95, DCL^{+}92, DTLM98, DMMM95, Dav35, DAV12b, DT09, DF84, Di 86, Die80, DKK80, Dom03, DM08, Dri91d, Dri91a, Dri91b, Dri91c, DIW00, DGS18, DGS24, DHM89, DK77, DF78, DZKK77, El 06, EL01, Elb01, EC79, EC13, EM94, Fab02, FLO04, FLMR00, Fei78, FS92, FGG82, FMC82, Fer86, Fer07, Fog51]. functions [FO93, FH62, Fre05, Fre09, FKY86, Fuk09a, Fuk09b, Fuk13a, Fuk13b, Gab79, Gal86, GKK00, GAZK03, GS81, Gau63, Gau64b, Gau64c, Gau65b, Gau69b, Gau77a, Gau98, Gau02, GHT03, GG08, Gau11, GS82a, GEK04, GRS87, GF97, GST02b, GST02c, GST02a, GST04a, GST06a, GST06b, GS12, GS14a, GST14b, GST15b, GF88, GJL20, GL83, GR92, GM84, Gla74, Gon52a, Gon52b, Gra61, GS14c, GM88, GK89b, GQ09, GM07, HTHR94, Hai60, Har08, HB75, HHv^{+}73, Her69, Her71, Hig77, Hit74, Hit75, Hoc61, HvdR63, HUY07, Hor17, HR88, HL13, Hum85, ISO10, ISK87, IS90, IKK91, IS92, IKF91, Ism77, JJW86, Jam16, JL12, Joh14a, Joh16, Joh12, JT85, Jon06, Jür66, Kal04, KWY07a, KWY07b, Kan96, Kas80, KT00, KV03, Koe16, Kog60, Köl69a]. functions [Köl69b, Köl69c, Kon76, KS88, Krä88, Kra14, KRVZ98, Kra85a, Kra85b, Kui52, Kul07, KB86, KI16, Kyr74, Kza92, Laf86, LS88, Laf91, LM93, LP61, LP62, Lan80, LR74, Laz85, Leb72, LVV05, LM00, LM07, Len90, Lew85, Lew87, Lin57, LK73, Liu93, LČJ63, Loe01, Lóp98, Lóp18, LU96, LM08, vdD88, LO94a, LO94b, LS95, Luk75a, Lut95, LP83, Mac68, Mac94a, Mac97, MS67, Mac96, MO49, MOS66, Mah30, MMV81, Mak65, Mak80, MT72, Mar81, Mar90, Mar00, Mar88, Mas83, Mat93, MP79, Max91, MMM93, McN73, Mei64, Mei67, Mer94, MS03, MY91, Mil89, Mil68, MM95, MM98, Mit49, MKS83, MC04, MCM90, MS87, MV98, Mos69, Mot55b, Mul95, Mul85a, Mul85b, Mul86, Mul24, MO18, Nag68, Nag01, NR15, NP18b, Neh07]. functions [Nes84, New84, Ng68, NU88, Nis84, Nis94, NT84, Nør50, Nør55, O’B80, OWS^{+}14, OGWY94, OG82, Olv80, ÖE16, PDK96, Par09, Par16, Paw11, Pea09, POP17, Ped03, PSS03, Peĭ63, PS77, PS78, PS79, PB82, Pie84d, Pie84b, PA86, PIH14, PI16, PKP74, PNV01, PT80, PS93, Pro87, Pro88, PBM86a, PBM86b, PBM92, QGG10, QXG18, QL19, Ran77, Rap01, RdC93, RKZ^{+}14, RL76, RM07, Ric92, Ric88, Ris79, Rit25, RX07, Rus88, RDK98, KR86, Saa91, Saf10, Sai89, dLSG18, Sal76b, Sau93, Sca71, Sch76a, Sch77a, SSG^{+}18, Sch78c, Sch80b, SS93a, SS93b, SGA81, Sea84, SdCR97, SG98, SG99, Seg08, Sen67, Shi93, Sho60, Sim64, Sla60, SP61, SS65, SS10, SS23, SS87, Smi89, Smi01, SF10, Spe72, Squ91a, Squ91d, Squ91b, SO89]. functions [SZW11, SJ12, Sti53, SMCH41, Sun88, Sun15, TTW84, TR84, Tem75b, Tem78, Tem79b, Tem81, Tem83b, Tem87, Tem94a, Tem96b, Tem00, Tem07, TV09, Ter79, Ter81a, Tes69, TB86, TB87, Tho04, Tim87, Tka03, Töl43, Töl50, Töl66, Töl67b, Töl67a, Töl68, Töl69, Töl70, Tre80, TW80, Tri66, Tur72, Ubh89, VC06a, Vep08, Vil88, VM87, VRS^{+}95, VPK99, Wac00, Wal84, Wal71, WCS16, War60b, Wat90, WS84, Wen07, Wen09, Wes05, Wil12, Wil71, WBR82, Wim72, Wim82, Win70, WG94a, WG95, XLX^{+}83, XL94, XC14, XC15, Ye17, YM03, Yun09, ZA11, Zag24b, ZZ96, Zai93, Zal89, ZSA86, Zha96a, Zha96b, ZS08, Zhu10, Zie89, Zwi87, de 77, dDM06, dDP10, dDIS13, dT93, vdLT84, Her82, Loz03, Tha62b, Sne63, Uni49, Köl87, VRS^{+}99, Wim72]. Fundamentals [AGA^{+}80]. FunGrim [Joh20b]. Funktionen [Fil66, Mei64, Rot71, Sne63, Sti53, Töl43, Töl66, Töl67b, Töl67a, Töl68, Töl69, Töl70, Völ83, Emd40, Emd45, Emd48, Emd59, Töl50, Töl66, Töl67a]. Funktionenlehre [Töl43, Töl50, Töl66, Töl67b, Töl67a, Töl68, Töl69, Töl70]. funktsii [PT80, PS93, PS93, LČJ63]. funktsioonid [Jür66]. funktsioonide [Jür66]. funktsiĭ [LP83, Bar72, PS93]. funktsiĭ. [DZKK77]. funktsiĭs [Saa91]. FUNPACK [Cod75]. Funtionen [Töl69, Töl70]. funzioni [AN68]. Further [GAZK03, GR92, JJW86, Laf84, Luk70]. Fused [BDL09, RJM^{+}20]. future [BIS01]. Fuzzy [AEF92, DP92].

G [Köl87, Köl90]. G5 [Bre74]. Gal [SZ05]. Galatasaray [HUY07]. Gamma [AM20, Alz08, BB15, Bar61, BY07, BP01, BC18, BI86, CP98, Chi20, CH67, Dav59, Dav64, DM86, DM87, Din58, Dop42, Ful72, GB68, Gau79a, Gau79b, Gau99, GST12a, GRAST16, God01, Gup60, GM04, Joh21, Kol70, Köl72c, Kol72b, Köl87, Kuk72a, Kuk72b, Laf84, Lan64, LDP93, LT71, Mac89, Mat04, Ng75, PH66, Pug04, Raf70, Rza̧12, ST07, Sch78b, SG02b, SF59, SLH06, Spi71, Spo94, Sri07, Tak66, Tem92, Tem94b, Vel11, WC61, Whi63, Wis48, Wre68, Wre73, AS05, Alz93, Alz97a, Alz97b, Alz97c, Alz98, Alz03, BZ92, BS78, BCJV18, CS19, Cau22, Cha80, CZ94, CZ95, CTV96, Che11, CB12, Che13b, CE15, CP15, Che15a, Che15b, Che16e, CL16, Che16d, Che16c, CT16, DF84, DH76, DiD84, FW13, Fis73, FW80, Fra81, FW84, GA08, Gau98, GHT03]. gamma [GST15b, GLP98, GL01, GR19, GQ09, GM07, Har08, JJW86, Jam16, JT85, Kar01, Ker83, Köl84, LS88, Lóp98, LFM14, LSM14, LW14, LSM15, LM15b, LSM16, Mah30, Mar77, Mar86, McC83, Mer05, MM98, Mor11, Mor14, MS15, Nem13, ND19, ÖE16, Pal98, PW92, Par02a, Par02b, Par03, Par16, Ped03, Phi79, Qi14, Sch88a, Sch80a, Smi01, SW50, Tem75b, Tem79b, Tem87, Tem94a, Tem96b, Ter79, TW80, Tri50, VC06b, Wan16, Wri84, XHT16, YC15, vdLT84, AWH13b, Bra64, CS11, CC64, Gau64b, Har97a, Her61c, Tho13]. gamma-functions [Mah30]. gamma-ray [Phi79]. GammaCHI [GST15b]. Gammafuncties [Dop42]. Gammafunktion [Sch80a, Tri50]. Gammafunktionen [Mah30]. Gappa [dDLM11]. Gate [LE95, LE93]. Gatteschi [GG08]. Gauss [Dav35, Fer07, FLS06, Fet65, Gau02, Gau16, HT13, Har20, HJ67b, Ibb63, JM18, KWY07a, Kza99, KP03, LW95, LT13, Nad15, POP17, Yun14, ZZ11]. Gaussian [DD08, Jan11a, Akr85, BTT16, BS00, Bre74, CCM11, CB09, CPG12, Dev12, DB13, Düm10, Fre09, Hay56, IR08, Jan11b, KL07, Kza92, LBC11, MJ13, PUHM12, SK11, Sun15, Ter81a, WCS16, Wen01, WLK11]. Gautschi [Zah94, Alz03, GLP98, Ker83]. gcd [BY19, Por20, Sor94, WTM05]. Gebiet [Töl69, Töl70]. gems [mH12]. genauer [von84]. Genauigkeit [Rot71, von80]. General [Bre82a, Erc75, Erc77, GW91, Kog89, Mil57, Mos72, von80, BD85, Bre80b, Fdi97, Fuk83, FI94, Fuk11, Fuk14b, Gon52b, Gus78b, HR07, LFM14, Mid75, SZ77, Sun15, Töl68, Wen01]. Generalised [Bar74, Ver67, NSU20, Par03]. generalization [Dij77, EM94, Pal98, Paw11, Rac82]. Generalizations [Dor62]. Generalized [AM20, AJDG02, AS54, Bar10a, Bar10b, Büh92, BS98, Bur64, Car70a, CZ94, Che70, CLOT86, DTLM98, DJ62a, Fie65, Gom89, Håv79, HP62, LC87, Mil85, Min70, Ovt03, Sch16, Sla66, Sle64, Sri70, Vel11, Vla02, WD61, Ali11, Ano84, BS11, BK16, CZ95, CTV96, CLM88, CH78, DCL^{+}92, Fuk14a, Fuk15c, GKK00, GAZK03, GS82a, GOST23, GS14c, GM07, Håv81, Har08, HB75, Jon70, KWY07b, KCYL06, KI04, Lav86, LKF10, MS16, Mat93, MM98, Mus87, Nag04, NP18a, Olv94, PBN93, Sal83, Sko05, WCS16, Wil12, MN16]. Generated [AK09, LSM14, MS03, dLSG18]. Generating [DCL^{+}92, Kyr74, Len76, LC87, Mul06, Sch16, Wil70, FW84, Mar77, Olv80, Wri84]. Generation [Ahm89, BH65, Bin68, CU59, IP17, JJ65, Kog60, Kog66, LA85, SM78, SA57, Wal80, AN68, GF88, KI16, Mat90b, OY91, XLX^{+}83]. Generator [KL14, AM78a, Bak76, Bre74, HXX22, War75]. Generators [BdDKL15, AHM^{+}98]. generazione [AN68]. Generic [Dri89, Tan88, Tan90c, Dri91a, Dri91b, Dri91c, Squ91a, Squ91d, Squ91b, Tan91a]. Genetic [LK21, PUHM12, Tan87]. gentle [BS13]. Geometric [Bar10b, BB84, BB97, BB00, BB04, Sal76a, AL20, Her82, LKF10, MH72, Sal89]. Geometrical [Yon70]. geometry [HUY07, Sti88]. German [Alb62, Bar14, Bau89, Bra65, Emd40, Emd45, Emd48, Emd59, Hen77, Hol69, Mah30, Mei64, Ost37, Rut51, Rut63, Sch80a, Sch68, Sne63, Sti53, Töl43, Töl50, Töl66, Töl67b, Töl67a, Töl68, Töl69, Töl70, Tri50]. Germany [BCD^{+}20, Gre16, FBM04, SO11]. GEV [Gom89]. Ginkgo [CTA22]. Givens [Bor21c]. GIZ [GST02c]. Gleichungssystemen [Bra65]. Gleitpunktrastern [Bra87]. global [HR07]. Glow [Che70, Squ70]. GMP [BMZ02, BMZ06, Zim00]. GNOME [SG99]. Golden [Hin77]. Goldschmidt [EIM^{+}99, EIM^{+}00, Mar04a]. Gompertz [Sto41]. Good [Gal85, Gal86, Mat90a]. Gopengauz [CG89]. Gopengauz-type [CG89]. Gordan [Buj75, Gun65a, RB75, SM77, SM78]. Gordon [Paw11]. Gosper [Che16e]. GPU [mH12]. GPUs [API^{+}19]. gradient [Mit92]. graduate [BW10]. Graph [NS09, DC81]. Graph-Based [NS09]. Graphics [NKY08]. Graphs [AS64, JMMW79]. Greece [SC01, VDR05]. Green [JS89, MKS83]. Grenoble [Bor02, KM91]. grid [Joh83]. Grosswald [Ano84, Sal83]. Group [AKS01, ???90, Koo91, Lip21, Tal68, WS84, Vil88]. Grouping [LA85]. Groups [Var95, VK95, CK88, Töl67b, Die80]. Grove [Sin95]. Growth [QVV98, Bos89]. guaranteed [Loe01, PR04]. Guide [AWH13c, BA44, LFB60, Tho97a]. Guidebook [Tho97b]. Gumbel [Gom89]. Gupta [Ano60b]. Gurland [Mer05].

H [Ano84, T.57]. Hadamard [Par09]. Half [DJ67, Gau16, Kah04, Lir71, Sau92, Di 86, Fuk10, Fuk15b, Gau94, GJL20, KWY07a, NR15, SG98]. half-century [Gau94]. Half-Integer [DJ67, Sau92, Di 86, SG98]. half-integral [NR15]. Half-Range [Gau16]. half/integer [KWY07a]. Hall [Lip21]. Hall/CRC [Lip21]. Halley [LLK14, NS13]. Halley-like [NS13]. Hamiltonian [Fei78]. Hand [Der77]. Handbook [AS64, Bak92b, Bee17, BL02, BF71, CPV^{+}08, Fet81b, Jef95, Jef00, Jef04, JD08, Loz97b, LCY65, OLBC10, PR82, PR96, PC11, RW04, Mat93, Köl90, Köl97]. Handheld [NKY08, GP84]. Handling [HFT94]. handyG [NSU20]. Hankel [Coc65, CH70a, CS82, CS83, DK90, Fet63, FW85, Glo89, JL12, Mac72, PB82, Wie99]. Hankel-norm [Glo89]. Hard [LP16, LP17b]. hardback [Lip21]. Hardware [Erc77, Fow93, Mar72, SS94, SF93a, SF96, SMG94, SDP11, Tak01, WG91, WG94b, XLW^{+}20, ZA14, dDI15, BKM94, Joh20c, Joh20d, Kan96, LGML05, Mul86, OY91, Sg92, SS93b, Tay81]. Hardware-Based [WG94b, WG91]. Hardware-Oriented [Erc77, OY91]. Hardwired [DM88c, LC87]. Harmonic [BD14, Alz97a, Alz98, Alz03, Ele15, MS67]. Harmonics [Bra73, Hob31, Hob55, XD18, MS67]. Hausdorff [Gre82]. Having [Bra65, FC60, Bos89, DGS24]. Hawaii [MS94, Spr00, Spr00]. hazard [Ruy89]. heat [CMV69]. Held [Mul82, AGA^{+}80, HUY07, IP87, Mar85, MC87, MC90, Ric92, SKL93, Tra76, Ull90]. Help [Car74, Wen96]. HEP [PIH14]. Her [L.61]. hereditary [Roc82]. Hermite [Bun09, Dri95, Fat16, Gau16, KSVW07, Nem13, Sab08, WdZ04]. Hidden [Car70b]. hierarchies [Her71]. hierarchy [Her69, Hom98b]. hierarchy-consistent [Hom98b]. High [Abe16, Ahr96, AFC10, Bar82c, BY07, BJM22, Bre06, Bru22, BTDS11, CCCP99, CG85b, CP99, CL94b, DPC95, Far81, Fer86, FW80, Fra81, Gal85, GB68, JSH^{+}11, Joh14b, KK96, KM93, KM97, LM95, LN21, LMT^{+}92, Maj85, MMW91, NP18b, Par09, PEB02, PB02, POMB05, RGK72, Sal51b, SW99b, Sch93a, Sch93b, Sei99, SF59, SC71, SA57, SG72, VLCSFN^{+}12, WMDM92, XLX^{+}83, ZG87, ALB98, Bar84b, BM80, CCC96, Fan89, Gab79, Gal86, Gau11, Jab20, Joh15b, Joh20c, Joh20d, KO94, LM99, MM92, PBE03, RdC93, Sch80b, SF93b, Bra87, Mül90]. High-Accuracy [Bra87, Mül90]. High-Bandwidth [Bru22]. High-level [BJM22]. high-order [Sch93a, Sch93b]. High-Performance [BTDS11, MM92]. High-Precision [Ahr96, BY07, Bre06, KM97, Bar82c, Fer86, FW80, Fra81, NP18b, Par09, XLX^{+}83, BM80, Gau11, Joh15b]. High-Radix [CL94b, PEB02, Sch93a, Sch93b, PBE03, SF93b]. High-Speed [JSH^{+}11, Maj85, PB02, POMB05, RGK72, SW99b, ZG87, Sei99, Fan89]. High-Throughput [AFC10]. Higher [CM90, GM88, LM92, Pot14, Ars81, BL71, Car69, EM94, MC89, Ng68, XC15]. higher-order [EM94, XC15]. Highly [DB13, Lut95, Mat90b, CY18]. Hilfe [Fil66]. Hilton [MMH89]. Historical [Dav59, Ypm95]. History [Ful99, Dut84, War08, Wen09]. Hitachi [WOG95]. HIZ [GST02c]. hoc [Yun14]. Hochgenaue [Bra87, Mül90]. HOL [AAHTH10]. Holdred [Ful99]. Holiday [IEE85]. Hong [DIW00]. honneur [AEF92]. honor [AEF92, Zah94]. Horner [CG24, CG02, Dor62, Ful99, GIJL23, Ost54]. Hotel [MMH89]. Howland [LW82]. HP [Hen77, LMOT01]. HP-25 [Hen77]. HP-UX [LMOT01]. hundred [DC81]. HURRY [FFS83]. Hurwitz [BB15, BS06, Joh15b, Sch16, Vep08]. Hvordan [Ped80]. Hvpergeometric [Sla64]. Hyatt [IEE89c]. Hybrid [SR14, CP99]. Hydra [Rot70b, Rot70b]. HYP [Kra95]. Hyperbolic [Ano74, AH16, Fri72, Ken14, Ken21, dLSGDR17, Sal53, SM70, Tha61, Zuc64, HTHR94, Hit74, Yun09, ZS08]. hypercube [Tan87]. HYPERDIRE [BK16]. Hyperelliptic [DR94a, DR94b]. Hypergeometric [AS95, AVV92, AS54, BS98, Bus76, Cah54, Car70a, FW61, Fie65, Gas81, Gel51, GS82b, GKS04, GN89, JKD92, Joh19, Kim72, Kop74, Luk59, LC61, Mil57, Min70, Nad15, NPA16, Ng77, Nør50, Nør55, Obe64, Olv91, Rel63a, Sla66, Sri70, Tha62b, Var95, Ver67, Wim68, Yos97, AS97, AS03, ABF^{+}23, ABFP24, AG08, BS00, BS06, Buc69, Büh87, Büh92, Büh03, BJLP19, BK16, CR08, CQSP04, CGKZ05a, CGKZ05b, CCV11, DT09, DGS18, Dut84, Fer07, FLS06, For97, GAZK03, Gau02, GST07b, GS14c, Har20, HUY07, HL13, Joh16, Jon70, Kal04, KWY07a, KWY07b, KW21, KT00, KRT02, Kra95, Kyr74, Lav86, Lew85, Lew87, LP95, Lew44, LT13, Mul01, Nag01, Nag04, NP18b, OSS09, Pas95, Pea09, POP17, PBN93, Rel63c, Seg08, Sko04]. hypergeometric [Sko05, Sla60, SZW11, Tem78, Tem81, Tem83a, TC01, Uni49, Ver77, WCS16, Wen01, Wil12, Zei91, ZZ11, BK16, Rel63b, Ric92]. hypergeometric-Bessel [KRT02]. Hypersingular [Sam02]. hypersurfaces [Sti88]. Hypervirial [FMC82]. hypot [Bor19, Bor21a]. hypotenuse [Bor21c, Ziv99]. Hypothesis [BCR^{+}07]. HYPQ [Kra95]. i860 [Sin96]. IA [ASA06, HKST99, Mar00, ST99]. IA-32 [ASA06]. IA-64 [HKST99, Mar00, ST99]. IBESS [ADW77a, ADW77b, Amo78]. ibid [Sha76]. IBM [ACG^{+}87, ACG^{+}86, ACG^{+}87, Fik66, Gel51, Mar90, WB78, Wan80, ZOHR01]. ICCD [IEE89c]. ICEEC’04 [Wah04]. ICMS [FvdHJT10, Gre16, BCD^{+}20, HY14]. ideas [Ano98b]. identification [Roc82]. identities [XC14, Zei91]. Identity [Gla71, Wri73, EC79, GMMIDM13, Mim88]. IEC [ISO10]. IEEE [BCDH09, BC01, BBdD17, CLT20, IEE05, IEE13, IEE21, IEE22, IEE23, KM07, MSH^{+}16, MTV15, SO11, TBL19, VDR05, AHM^{+}98, AAHTH10, BT99, CH98, DR04, GB91, Hou81, IEE85, IEE89c, IEE99, Int03b, Kah80, KK99, KM91, LMT97, LSZ08, LD89, Rus98, Tan89a, Tan90e, Tan92, T^{+}97]. IEEE-754 [AHM^{+}98, AAHTH10, BT99, DR04]. IFIP [Boi97, Lav80]. ihre [Mei64]. II [BL91, MC90, WB78, Ano52, Ben83, Bul65b, CG89, CKT07, Cuy94, Dun87b, Eck77, EMR82b, KSVW07, LP61, Luk71b, Mil65, Ovt03, Pro87, RL76, Sch68, Sne63, SZ74, VB81]. iii [L.61, Ano60a, Bul69b, Din58, Dun88, LP62, Lov53, Luk72, SZ76]. IIPBF [RKZ^{+}14]. Illinois [Cow77, Hwa85]. Illustrated [Tho97b, Tho97a]. im [Töl67b]. IMA [MC87, IP87]. IMA/SIAM [IP87]. image [JL94]. Imaginary [Bar61, GST02d, NPA16, BF92, BST13, Cob69, GST03a, GST04a, GST04b, VC06b]. immiscible [YH89]. impedance [AJ03]. Implementation [AAHTH10, AFC10, Bai17, BV85, BLP24, CVW06, DMS95, DHL^{+}04, EL89a, GBKK09, HH18, IM99a, Jef17, Joh15a, Joh87a, KNS95, LA01, LP17a, LM15a, LC97a, LE95, Mul97, Nav83, PP12, Sht19, SMG94, SL97, SDP11, Tak00, Tan88, Tan89a, Tan90e, Tan92, VM19, Zai93, dDLM11, von80, DR04, Has00b, Joh14a, Joh87b, LKKH99, LO95, LE93, MC89, Neh07, Obe99, Tom00, VT11, ZXBZH01]. Implementation-Oriented [PP12]. Implementations [EL94, PB08, SL95, Tan90d, XLW^{+}20, dDI15, CL00, Kul07, LC96, LC97b, Luk99, dLSG18, SK08, SL97, Zim00]. Implementierung [von80]. Implementing [HFT94, Tho87, Che71]. Implied [Sho02, MRS17, SR16]. Improved [Bor19, Bor21a, CD02, HCK09, Jab13, Jab20, KL07, KCYL06, Mor11, Nin70, Olv91, ST99, Di 86, DD08, GST12b, Lu14, Mid75, PW92, PC93, ZLC04]. Improvement [LK21, Mec68, OGWY94, Wan80, LLK14, Nem13]. Improvements [Tru11, Zag16, PS86]. Improving [Chl11, EIM^{+}99, EIM^{+}00, WMC18, WMC21, Lev73]. incident [Ehr89]. Including [She54, Str56, HTHR94, SMCH41]. Inclusion [Cri82, PC93]. Incomplete [AM20, Bar61, BL94, CN81, CP98, DH59, DJ67, DM86, DM87, DM88a, DM92, Dop42, Ehr61, FL67, Ful72, Gau79a, Gau79b, Gau99, GST12a, GRAST16, Goa95, GM04, Har08, Hil77, Kol70, Kol72b, LW63a, LW63b, LDP93, Mat04, OM68, PSB76, Sau92, Sch78b, Tak66, Tem92, Tem94b, Tho13, Van69, Vel11, Whi63, Wis48, Alz97c, BB15, Bul69a, CZ94, CZ95, CTV96, DH76, DiD84, DK90, Fis73, FI94, Fuk09b, Fuk10, Fuk13b, Gau98, GHT03, GR19, GM07, HF09, Hay56, JJW86, Jam16, JT85, Mah30, Mar86, MM98, ND19, ÖE16, Par02a, Par02b, Par03, Par16, Sch80a, SS10, SS23, Tem75b, Tem79b, Tem87, Tem94a, Tem96b, Ter79, Ter81b, TW80, Tri50, VC06b, YA24, BB74, Gau64c, Gau64d, Jef62, Lud63, Mey63, PH67, vdR63]. Incorporated [PW76]. increasing [Wan82]. incrementation [DD76]. Indefinite [CK89b]. Independent [Fra65b, Bre18, VCV01]. Index [Ano68, GKK00]. India [Ban01]. Indian [BB12, Hay99, Plo01]. indices [GF88]. Induction [Kin21]. Industry [BBM84]. Inequalities [AVV90, AVV92, Bus76, BI86, CD08, CP15, Che15a, Che15b, CL16, Che16a, Che16b, GL01, Laf84, Laf86, Nas74, QVV98, She54, Alz93, Alz97b, Alz97c, Alz08, Alz10, CB12, CM13a, Che16c, CT16, CT88, GLP98, IS90, Ker83, LS88, Liu93, LW14, Pal98, QM15, RT89, Whi82, ZS08, Zhu10]. Inequality [WLK11, Alz97a, Alz98, Alz03, IS92, Rac82, Zal89]. Infinite [Lir71, Sch88b, BC79, Cro92, LW96, LS95, RKZ^{+}14, VC06a]. Infinitely [IM99b]. Information [GS14b, ISO10, Lav80, CT88, Ste08, Töl68, War08, Was89]. Ingenieure [Bar14]. inhomogeneous [Mac94a]. Initial [Cas83, ITY95, EL89c, EL90, ITY97, SS98, Spi85]. Initiative [Loz00]. Inn [IEE85]. inner [AB88]. Input [Wan78, Wen01]. insertion [OJ04]. Instantiation [PLT23]. institutions [BL71]. Instruction [GZC22, PW76]. Instructions [GBKK09, IM99a]. Integer [Alh00b, BO93, Cat85, Cat86, Cre98, D^{+}89, Der77, DJ67, Fox54, Has90, Hei96, JF16, MRH19, ND70, OLHA95, Olv64, Rev16, Rol87, Sau92, Soo73a, Soo73b, Alh00a, AM78b, AM79, AM84a, AM84b, Di 86, FNC75, Fuk15b, GLM15, Hig93, KWY07a, KWY07b, Lyo91, Pro83, Sca71, SG98, WTM05, dT93]. integers [QW06]. Integrable [PS93]. Integral [Ano93, Cad51, Car75, Che70, Chr65, CT69, CP79, Cor61, Cun62b, Cun62c, Dav59, Dre78, Fet67, Fra65a, Fri72, GC64, Gau16, GW71, Gla71, Gla79, Goa95, HP67, HJ67a, HJ67b, Hil73b, Ism77, KC76, Mec68, MR65, Mer62, NF63, Ng70, Olv91, Pac70, PS11, Qi14, Raf70, Red70, Rei69, RL80, Ryb89, Sch78a, She54, Sko78, Smi11, Squ70, SZ74, VLCSFN^{+}12, War60a, WR63, Wri73, eL76, AQ18, Act74, Bea60, BB82, Ber68, Boy15, Bry02, CG85a, Car05, CS11, CT68, CPT70, Dij77, DK90, Fet81a, FKY86, Fuk10, Fuk11, Fuk14b, Fuk15d, Fuk15b, Fuk15c, Gau59, GHT03, GOST23, GMMIDM13, Gra61, Has02, Her61d, Her61e, Kuz15, Lar66, Lee92, Let97, Mar95, Max91, McC74, Mid75, NR15, NP18a, OW68, Olv94, Pak18, Pól49, RT61, RKZ^{+}14]. integral [Sun15, Tem86, Tha63c, Tha63b, Töl67b, Van89, WR96, Zag24a, Zag24b, HP85, Hol70]. Integrale [Töl68]. Integralgruppen [Töl67b]. Integrals [AVV90, AVV92, BH63, Boe61, BF71, Car77a, CN81, Car87, Car88, Car89, Car92, CG94, Car99, CR68, Cod65a, Cod65b, Cod66, Cun62a, Cur64, DH59, Din58, DR94a, DR94b, Ehr61, FL67, Fet65, FC64, Fle68, Fre05, Fuk15a, Gau61, Gau64e, Gau73, Gau77b, Gla66, GRJZ07, Gra63a, Jef95, Jef00, Jef04, JD08, KKI67, Kin24, Kin07, Köl87, KK20, Lan60, LW63a, LW63b, Lew69, LDP93, Lin72, Lóp00, LG64, Luk64, Luk68, Luk70, MT64a, NC66, PT90, PBM86a, PBM86b, PBM92, QV96, QVV98, Ric73, Sam02, ST07, Sch64, Sch15, Sid95, Sny93, Sny96, SZ76, Tem15, Tho65, Tho66, Töl67a, Van69, WKG66, Woo67, ZC70, Act74, And00, BC79, Bat19, BO75, BZ92, BBC08a, BS11, Bul65a, Bul65b, Bul69a, Bul69b, Car79, Car91, Car95, CLM88]. integrals [Chr62, CT67a, CT67b, Deh89, EW63, Far69, FGG82, FMC82, Fet81b, FF54, FF57, FI94, Fuk09a, Fuk09b, Fuk12, Fuk13b, Fuk14a, Gab79, GP64, Gau11, GST20, Gla74, Gra63b, Gra64, Gra02, Har00, Hay56, HvdR63, Jef62, KT00, KRT02, Köl97, Lan80, Las82, LW82, LK73, LS95, Mac72, Mak65, Mey63, Mim88, MH72, Mor78, Mor99, Ném65, PDK96, Pie82, PB83, Pie84a, Puo88, Saf10, Sai89, Sch88a, SZ77, SJ12, Tem83b, Tem83c, Tem85, Ter81a, Tha63a, Töl68, VC06a, Won73b, ZA24, vdR63, vdLT84, MN16, Sny21, Lip21]. integrands [CF00]. integrated [DM88b]. Integration [Bro90, Car99, Cas83, CK89b, Dav88, LO93, Pro88, BGV81, Bro89b, EM86, Hay56, Mos69, Pro87, Ter79]. Integriruemye [PS93]. Integro [Mil85]. Integro-Exponential [Mil85]. Intel [ASA06, CHI^{+}03]. Intelligence [AGL93]. interact [Wen09]. Interactions [BJG^{+}99]. Intercalation [Wyn72]. Interest [Koo91]. interfaces [ISO10]. interference [Dem89]. Interior [Meg89b, KMY89]. Interior-Point [Meg89b]. Interlacing [Seg08]. internals [SF93b]. International [ACM89, AGL93, Ano91, BCD^{+}20, Bor02, BCEP94, Bro07, FvdHJT10, Gre16, HY14, IEE89c, IEE08, IL07, Lak96, MC90, MS94, SE11, SC01, Spr00, TT18, T^{+}97, VDR05, Wah04, ACM87, DIW00, RDK98]. interpolate [IP81]. Interpolation [Bre82a, Fat16, Sti53, VB04, WEX14, BSW95, Blo88, Cau22, FK88, GRS87, Håv81, PDK96, RT89]. Interpolator [POMB05]. interpretation [Din73]. Interpretations [Yos97]. Interval [BSY82, CG02, DLM09, Krä87, PLT23, Rot70a, Yoh73, Ge89, Joh20a, Kai89, Krä88, Mar81, dDM06]. Intervallargumente [Krä87]. Intervalle [Bra87]. Intervals [Bra87, KP75, EMR82b, Shi93]. INTLIB [KDDH92, KDDH94]. Intrinsic [BJG^{+}99, BD99, BJG^{+}00, Sht19, Mil89]. Introduction [Bow53b, Bow58, Bow61, Dri91d, Hol90, PP12, SG02b, Tem96a, Tka03, BS13, MC04, NU88, Ars62]. Inverse [AH16, Car63, DM86, DM87, ELMT00, Gil16, Ham78, Hom96a, HMO19, Krä87, Krä88, KNS95, LaF54, Lin90, Mar05, MS68, Pag77, PB02, Sal53, SW99b, Sol69, Str68, Tha61, WMC21, ZA14, Ano12, BPT76, BEJ76, Boy15, DAV12a, Dom03, Fet74, Fuk15d, HTHR94, Kre89, Lin89, MZM94, MWH^{+}16, MWH^{+}18, PDK96, Rev90, Sho86, VPK99, WMC18, Krä87]. Inverse-Square-Root [KNS95]. Inverses [Sal51b, Sch84, Car05, Rit25, Wac00]. Inversion [GST12a, Por20, Tem92, BY19, GST15b]. invertible [SE14, Ved93]. investigate [PI16]. Investigation [Mak65]. inviscid [Joh83]. involved [Bel58]. Involving [BFHT85, Gla66, Lew69, Ric68, Sri70, UE05, WR71, BB82, DT09, FMC82, FO93, Gab79, GMMIDM13, LW82, LK73, LS95, Ng68, PS77, PS79, QGG10, XC14, XC15, Zhu10]. Inzell [zCFF05]. Irish [Mil75a]. Irregular [SG72, Wen01, Bar81, Bar82c, Nes84, SSG^{+}18]. ISBN [Lip21]. isn’t [Wil96]. ISO/IEC [ISO10]. ISSAC [ACM89, Lak96, SE11]. issue [T.57]. issues [BH01]. Istanbul [HUY07]. Italia [RS98]. Italian [AN68, RS98]. Italy [AGL93, CDS00, RS98]. Itanium [CHI^{+}03, Int00, Int03a, Int03b, LMOT01, Mar03]. Iterated [Amo73, Tha61, Deh89, GLM15]. Iteration [Alb62, BB89, Bra65, CL11, LC87, Nin70, WS05b, Aab54, Che71, DD76, Epp89, MC93, MC95, PS86, PR04, Rut63, Tuk48]. Iterations [Gar75, Gar79, Gar80, HHR00, SS98]. Iterationsbeschleunigung [Bra65]. Iterationsfolge [Alb62]. Iterationsfolgen [Bra65]. Iterative [CH98, Eve63, Hom96a, Joh89, KNS95, Plo96, RGK72, GA08, Har49, Hom98b, ITY96, Jam89, KK96, Kru99, KI04, ZXBZH01]. IV [WB78, Dun94a, Dun94b, Sle64, WB78, Wan80]. IX [Luk99]. izbytochnykh [LP83].

J [ACG^{+}87, Alz98, Fet81a]. J. [Ano84]. Jack [Ric92]. Jackson [CK88]. Jacobi [Ali11, DCL^{+}92, HT13, HvdR63, Mey01, PDK96, Paw11, Sko75b, Spe72, WV20, Wit68, Ye17]. Jacobi-elliptic [DCL^{+}92]. Jacobi-fractions [Spe72]. Jacobian [Bat19, Car05, Fuk09a, Fuk09b, Fuk13a, Hol90, Lan80, MT64b, MO18, PDK96, Sal89, Sch76a, Sch77a, Töl67b, Töl67a, VPK99]. Jacobische [Töl67b, Töl67a]. January [Mul82, Spr00]. Japan [Ano91, FvdHJT10, Lav80, TBL19]. Japanese [OGWY94]. Java [Lau04]. JBESS [ADW77a, ADW77b, Amo78]. Jensen [Fre09, WLK11]. Jersey [AFI71]. Jim [Bli97]. job [Bar14]. Jodhpur [Ban01]. Joint [AFI69, AFI71, AGL93, IP87]. Jonathan [Bre17]. Jonquiere [FK75]. Jordan [Zhu10]. Jose [SE11]. journées [AEF92]. July [Ano06, BCD^{+}20, Boi97, Bro07, BBdD17, Gre16, IEE08, IL07, KM95, Lak96, LMT97, Luk99, Mar85, MC87, MC90, MSH^{+}16, SO11, SIJ93, T^{+}97, VDR05]. June [AGA^{+}80, BCDH09, BC01, CLT20, Cow77, DIW00, HUY07, Hwa85, IEE05, IEE21, KM91, KM07, MTV15, SE11, SIJ93, TBL19, TT18].

K. [Rut51]. K5 [LASC95, Rus99]. K6 [IM99a]. K6-2 [IM99a]. K7 [Rus98]. K7^{T M} [Obe99]. Kāshı [Aab54]. Kōbe [FvdHJT10]. Kōriyama [SO89]. Kapteyn [DTLM98]. Karatsuba [Zim99]. Kaufman [AEF92]. Kaufmann [AEF92]. Kelvin [Bur63]. kernel [WO15]. kernels [Mus87]. Kershaw [GLP98]. Key [AFC10]. Kiev [Köl90]. Kind [Bar10a, Boe61, Car77a, Car87, Car88, CMF77, Cod83, DH59, Fet65, Fra65a, Gau64a, GST02d, Jab94, Mec66, PS05, Sko75a, Tem75a, Vel11, War60a, ZC70, BL96, BZ92, Boy15, Bul69a, Cob69, CV03, Fet81a, Fuk10, Fuk11, Gau65b, GST03a, Gla74, Her61d, Her61e, HvdR63, Lar66, Mid75, PDK96, Sca71, SG99, Tak22, Tem76, Tha63c, WC90, Wen96, YM97, YM03, Zai93, Bra70]. Kinds [FL67, LW63a, LW63b, Van69, Mor99]. Kingdom [Boi97, IEE89a]. kink [Paw11]. Klein [Bar84b]. Known [FC60, BMR04]. Koecher [Bau89].

Koechers [Bau89]. Koeffizienten [Bra65]. Kohlrausch [ZLA13]. Kolmogorov [Pom76, Pom74]. Komlos [Ein89]. Kompleksmuutuja [Jür66]. komplexe [Bra87, Krä87]. Komplexen [Töl67b]. Kong [DIW00]. Konstante [Bau89]. Konstanten [Sti53]. Konvergenz [Ost37]. konvergenzbeschleunigenden [Sch68]. Konvergenzbeschleunigung [Alb62]. Konvergenzverbesserung [Hom96b]. Korea [HY14]. Korotkov [Lip21]. Kryachko [Har01]. Kummer [Gau77a, GST15a]. Kuramoto [AGH89]. Kutta [Cas83]. kvadratroden [Ped80]. Kyoto [TBL19].

L. [L.61]. Laan [Köl87]. Lab [Fet81a]. Laboratory [L.61, Low49]. Lagrange [Blo88]. Laguerre [Air37, BBC08b, CIM94, DGA18, DGS18, GST17, GM84, Kza99, PGA18]. Lambert [Gau11, Goe14, JHC96, Jef17, Joh20a, MS16, Miy19, Pak18, Veb12]. Lanczos [BCEP94, God01, Pug04]. land [Ars81]. language [Che81]. languages [ISO10]. Laplace [Fab02, Gus85, Lee92, Tem83c, Tem85]. Large [Bin68, Cur64, Fox54, Fox60, GT58, HJ67b, L.61, NPA16, TKK17, Wan78, Won73a, FLS06, Gau59, GOST23, GF88, Har20, Jon06, Mas83, O’B80, RX07, Tem87, Tem94a, Zha96b]. Larger [Gau65a, Jan77, de 77]. Last [Ziv91]. Latency [ALMN05, Bru22, PB08, SR14]. Lattice [SLZ05, Har77]. Lauricella [BK16]. law [Deh89, Pak18]. layer [Bel89]. layout [SR81]. Leaky [Har97b, Har98, Har01, Har08, PHH08]. Learning [OJL22, BL71, Lit89]. Least [DB13, Sto41, Rei86]. Lebesgue [Blo88, Bos89]. Lecture [Mil58, HUY07]. Lectures [Tal68, Tri66, zCFF05]. Legendre [Jan77, Töl67a, Bra73, Bre18, BL45, Cob69, Gau65a, HT13, HS20, JM18, LW95, Lew69, SSG^{+}18, SOL81, Ste64a]. Legendresche [Töl67a]. lemniscatic [Eck77]. Length [SH72, HHv^{+}73, Hon71, von84]. Leonhard [Dav59]. Lerch [GS88]. Less [Bak92a]. Letter [Cat86, Cod67, Fik67a, Kin65, Knu63, McC90, Rev90, Swa65]. Letters [Rah01, WTA^{+}90]. Leuven [IEE08, KV03]. level [BJM22]. libcerf [Ano15]. Libcint [Sun15]. Libm [Ano14, LMOT01, dDEG05]. Libraries [IEE99, MMNP91, WB78, Wan80, dLMF23, Kuk71, LKK19, LN21]. Library [AK09, Ano22, Bee17, BLP24, DLM09, FHS77, GB91, Joh20b, KDDH92, KDDH94, Kuk66, Lau95, Lau04, LMS99, Loz03, The21, McC02, Mor75, Mor81, Mor90, Mor93, PW76, Sch76b, Sch77b, Sch88b, SP20, TGC89, Tho87, WWG82, ZOHR01, Ano14, CTA22, FHL^{+}07, KA71, LMOT01, Mar03, Mor84, Neh07, PIH14, Sun15, Wil96, DG99, ISO10]. Lie [Die80, CK88, Mil68, Var95, VK95]. Like [Köl87, Ngu98, NS13, vdLT84]. likelihood [Gom89, vR89]. Limit [Hin77, JW88, Lem88, LW89, RX07]. limited [EM03, LP62]. Limits [CM13a, Gon52b]. Line [BH18, Erc78, OE82, SMG94, Boe91, DHM89, Frö63a, PBE03]. Linear [DGS65, Jam81, KP75, LB90, Ng77, Olv67a, Olv88, SF79, YH89, Ano67, BDGB83, CG89, Die80, Ge89, Gon89, Joh20c, Joh20d, Joh88, KMY89, Lev73, Lev91a, Lit89, Luk56a, Luk56b, Meg89a, Olv67b, PS05, Tan90b, Via89, Was89, vR89]. linear-threshold [Lit89]. Linearized [HBS00, Ehr89]. lines [BF92]. link [Wil96]. Linux [IBM05]. Lipschitz [DK90, Som82]. Literature [Pri66]. LNS [DMS95]. local [Kee82]. locally [Vai89]. Locating [PNV01]. Log [MJ13, OTY91, BS11, Har97a]. Log-Depth [OTY91]. log-Gamma [Har97a]. Log-quadratic [MJ13]. log-sine [BS11]. Logarithm [AS10, CH67, CVV65, Cus83, Has66, Kah04, Kar84, Kir88, Köl72c, LP16, LP17b, Lin09, LA85, LC87, Mac89, MW83, PEB04, Rev16, SK71, SMG94, Tan90a, Tan90e, AM12, Che71, Deh89, Epp89, Köl84, Per85a, PEB02, dDLM04, PH66]. Logarithmic [CET62, CLOT86, KP69, SF79, XD18, CS83, GEK04, Hit74, Joh20c, Joh20d, Lit89, QL19, Sed90, Sko04, Fog51, Zuc64]. Logarithmically [Osa90, DGB82, GQ09, Hom99]. Logarithms [Che72, Mar72, Car72, LBdDM16]. Logic [Cow64, LA85, PL19, Sch73]. Logistic [Lin90, BKKC09]. Lommel [Boe62, FS92]. London [BBdD17, L.61]. Lookahead [KNS95]. Lookup [GBS19, dLSGDR17, SS97b, Tan91b, Par99]. loop [Dam94, Mim88]. LORIA [Ano06]. Loss [Zag19, DAV12a, DAV12b]. Lost [Car74]. Lösung [Die83, Rut51]. Love [Yos97]. Low [AH10, ALMN05, Bru22, Car75, HFS99, NL99, RHMLL08, SR14, KP98, MI97]. Low-Complexity [RHMLL08]. Low-Latency [Bru22]. Low-Order [Car75]. Low-power [NL99, KP98]. Lower [Tro84, WLK11, Bel89, CPG12, Fre09, Töl43]. lower-bounds-to-circuit-size [Bel89]. LSFBTR [Tal83, Tal84]. LSI [Atk79]. Lubbock [IEE85]. Luigi [GG08]. Lyon [Ano03, MTV15].

M [Fet81b, Köl87, Köl97]. MA [TT18]. Macdonald [FLR03]. Machine [BF23, Lan42, OJL22, Sch76b, AOB64, Bar81, Bel89, GB90]. Machine-Efficient [BF23]. Machinery [Cah54]. Machines [Bow53a, Gus86, Lin57, Rot70a]. Macrocell [BH07, AHM^{+}98]. MAD [BBC^{+}91]. Madelung [CB87]. Madison [Ask75, Lan59]. Magic [HMO19, MWH^{+}16, MWH^{+}18]. Magma [ES10]. magnitude [Zha96b]. Magnus [Coh12, Szm13, Wig67]. Majesty [L.61]. major [SL97, Ein89]. Make [Wil96]. man [Ped80]. manifold [AGH89]. manipulation [Kra95]. Manipulative [Sun71]. Mantissa [von84]. Mantissenlänge [von84]. Manual [CW80, Ful81b, Ful81a]. Many [Sal53, Sle64, KW21]. MAPLE [Qua18, ES10, Zei91]. mapping [ZA14]. March [Ask75, Ban01, IEE85, Koo91, Men06, MT85, MT86, Ric92, SKL93]. Marcum [FK11, GST14a, LKF10, SF10]. Markov [Rac82, Whi82]. Martingale [RT89]. Maryland [IEE99]. Mascheroni [CM13a, Che16a, Che16b, Chl11, XC17]. Massachusetts [ACM87, IEE89c, IEE05]. massively [BBC^{+}91]. massively-parallel [BBC^{+}91]. Matematicheskiıanaliz [LČJ63]. matematicheskiĭ [PS93]. Math [Alz98, ASA06, Ano84, BLP24, Fra81, Joh87b, Kuk66, Lew87, MMNP91, McC02, Per87, Sha76, Wil96, Wim72, CTA22, LN21]. MathCW [Bee17]. Mathematica [ABFP24, BK16, ES10, Kra95, Tho97b, Tho97a]. Mathematica-based [BK16]. Mathematical [AS64, And82, Ano22, Arf85, AW01, AW05, AWH13c, Arn77, Bak92b, Bee17, BCD^{+}10, BCD^{+}20, BL02, BF23, BdDKL15, CHG^{+}11, DHL^{+}04, ES10, Fik68, FHS77, FvdHJT10, Ful80, GB91, GS14b, HBF09, Hoc71, Hoc86, HY14, Jef95, Jef00, Jef04, JD08, JS56, JS99, Kuk71, KL14, L.61, LFB60, LČJ63, Loz97b, LMS99, Loz03, Luk75b, Luk77, Meg89b, Mol83, Mor75, Mos89, Mul82, OLBC10, Osi13, RW77, Rap14, Ric71, SP20, Sne56, Sne61, Sne80, SS70, Szm13, Tem96a, Tho97b, Tho97a, WB78, Wan80, WNO94, Wig67, Ziv91, ZOHR01, Zwi03, dLMF23, Ars68, BBC^{+}14, Bar14, CC17, Coh12, Her82, Hoc61, ISO10, KRVZ98, Kul07, KI16, MO49, MOS66, NU88, PIH14, PI16, PS93, QXG18, Sne63, FvdHJT10, Gre16, IBM05, Sne63]. Mathematical-Function [Bee17]. Mathematician [Bat07]. mathematicians [Hay99, Wen09]. Mathematics [Ask75, BB12, Bir54, BDL^{+}01, Car77b, DG99, FR98, HDG^{+}15, Mor81, Mor90, Mor93, Mur88, PC11, RW04, And98, Ano98b, BL71, Gau94, Joh11, Mor84, Mur99, Gau94]. Mathematische [Sne63]. mathematischen [Sne63]. mathematischer [Bar14]. Mathieu [Alh00b, Alh00a, BC62, Bla64, Cle69, EC13, Fri72, RL80]. Maths [LK21]. MATLAB [Bat19, ES10, API^{+}19, Zen04, RKZ^{+}14]. Matrices [Leh96, HHR00]. Matrix [AH16, Ike76, AIS^{+}17, API^{+}19, CS19, HHR00, Kru99, Miy19, Ng84]. Maui [Spr00]. maximal [von84]. Maximally [von84]. Maximum [KP75, vR89, GLM15, HF95, QXG18]. maximum-redundancy [HF95]. May [AFI71, CDS00, IEE81a, IEE81b, IEE99]. McDougall [Fuk14b]. McMillan [BJ15]. Mean [BB84, BB89, BB97, BB00, BB04, Sal76a, Alz97a, Alz98, Alz03, KPPS08, MH72, Sal89, Tod90]. Means [Fet65, Fil66, Gus66, Mül90, von80, BS06, CS11, Fer86, Kon76, Leh71, MMV81, MS87, Saa91, Tes69, Wen10]. Measure [DR94a, DR94b]. Measurement [BDL^{+}01]. Measurements [Lid01, Lid02]. measures [Was89]. Mechanical [SG02a, CK89a]. Mechanically [Rus98, Rus99]. Mechanics [Bir54, Sti88]. media [YH89]. medieval [Joh11]. Medium [Joh15a, Joh14a]. Medium-Precision [Joh15a, Joh14a]. Meijer [BS13]. Melbourne [Lav80, Men06]. Melfi [CDS00]. Mellon [Tra76]. Memoirs [Lov53]. Memoriam [Dav59]. Memory [LJ13, Kog89, Tak00]. Memory-Efficient [LJ13]. mesh [DR89]. Mesopotamian [Man23]. Metalibm [KL14]. Method [BC23, BCE^{+}08, BEM12, Car74, Cow64, Ded26, DJ62b, DJ62a, Erc75, Erc77, Esc37, FK19, Fle68, GZC22, GW91, Hol69, Joh89, Kal34, KP69, LaF54, Lan42, Lin79a, Lin72, LJ13, Mar05, Osa90, Ovt03, Plo96, Qua18, Sar59, SF96, SZ05, SS99b, Sto41, Tra60, TS76, Van69, WMBE14, WN09, Wyn72, Yon70, Ypm95, Aab54, AM12, Ali11, Bar82b, Bar84a, Bor82, BR94, DK82, Dom03, Fab02, Fuk14b, HTHR94, Hai60, HR05, Hit74, HF95, Hor19, Kon76, KI04, LG79, LP95, LP83, Lyo91, Mag94, Mar77, Mat90a, Now06, Ole96, Ost37, Ost38, PR04, PB83, Puo88, Rap01, RdC93, Rei86, Sch68, SdCR97, SZ77, Sko05, Tru11, Tur43, Vai89, VT11, Via89, VB81, Yun14, Zai93, ZB95, Zha95, Zha96a, Rap94]. méthodologie [Mul86]. Methodology [Sho02, Mul86]. Methods [AW05, AWH13c, BR73, BPT76, Bre82b, BR91, Cas83, Dub83, Fik67b, Fra65b, FR76, GSS12, Gau75, GST07a, HF09, HBF09, HR07, ITY95, JS56, JS99, Meg89b, Mos89, Ng75, RGK72, Sid03, SZ70, SZ74, SZ76, TAY91, Tem15, Ull90, Vla02, AGS99, Ano98a, Arf85, AW01, BDGP13, BDGP14, BL71, Boy89, Boy01, Bre75b, Bre76b, Bre78b, Bre85, Bre88, CG24, Cuy94, Epp89, FGG82, Gab79, Gus84, HP93, KK96, Kru99, LVV05, LLK14, McC77, Mei67, Mon83, Mül93, NS13, Oli75, Osa94, Özb06, Pas10, Pas11, POP17, PC93, PKP74, Pol88, RW77, Roc82, RDK98, Saf10, SS98, Spi85, Sti59, Wat90, Wim82, Die83]. metodom [LP83]. metric [Som82, ZSA86]. Mexico [IEE91]. MHz [PZ95]. Michigan [IEE81a, IEE81b]. Micro [DSK83, DSK84, Ber84, Bic81]. Micro/mini [DSK83, DSK84]. microcode [Rus99]. microcomputer [Mil89]. Microcomputers [AMT78, AJMT81, MTP82, Völ83]. Microprocessor [And82, Die83, HAK91, Obe99, Sin96, SL97, Kun81]. microprogramming [Hit75]. Mie [Len76]. Mikrorechner [Völ83]. Mikrorechnerprozeduren [Die83]. Military [MC87, MC90]. Miller [LP88, Mil04, Ter81b, Zai93]. Milton [Dav59]. mincing [Bel89]. mini [Ber84, DSK83, DSK84]. mini-micro [Ber84]. Miniaturized [Luk71a, Luk71b, Luk72]. Minimal [Met65]. Minimax [Fra65b, LW95, POMB05, Fuk15d, Fuk15b, Let97, Pol88, Wan82]. minimization [Dan77]. Minimum [XLW^{+}20]. Minkowski [Kra89]. Mirakyan [Sun88]. Miscellanea [Ham78, Lin90, Pag77]. Miscellaneous [Ste64b]. MISCFUN [Mac96]. Mises [Bir54, Deh89, Hil77]. Mistake [Lit89]. mittels [Mül90, von80]. mixed [IS88]. Mizan [IK05]. mnogochlenami [DZKK77]. Mnogoznachnye [Bar72]. Mod [WB78, Tur94]. Model [CHG^{+}11, Fra78, Gom89, Wen01]. Modeling [Sho02]. modelling [Sob88]. models [Ric88, SC81, vR89]. moderate [GOST23]. moderate/large [GOST23]. Modern [HH18, BL71]. modes [LN22]. Modification [Bar76a, Bar76b, WS06, Bar84c, ITY97]. Modifications [Ree90]. Modified [BHK09, Bra73, Bur74, Car74, Cod83, Ded26, Ful72, GST02d, Hin77, Lin79a, Mec66, Nas74, PB83, PT91b, Rap14, Sch78b, Tem75a, TB87, Vel11, XD18, AS97, Amo74, Bla74, BO75, CST18, CV03, DT09, DC81, Fab02, GS81, GS78, GST03a, GST04b, Hei88, KMB94, Kas80, Laf91, Las82, LG79, MV98, NR15, Par72, Sca71, SdCR97, Tak22, Tem86, Wal84, WC90, Zhu10, Fet63, GST04a, Tho04]. Modifying [WMC21]. Modular [DMMM95, Por20, Yos97, BY19, BBB89, WTM05]. Module [Kod11, LD89, Bas01, EL89a, GST12b, MM92]. Modulo [ADN03, Tor02, BH99, Per85a]. Modulus [CS83, AM89, Fet81a, MH72, Mor78, Mor99]. Molecular [LDP93, Saf10, WO15]. Møller [Hom96b, Hom96b]. moment [FW84, Gre82, Wri84]. Momenta [TTW79, TTW84]. Moments [LL72, FW84]. Monica [ES89, IEE78]. Monotone [Alb62, Dun94a, Dun94b, FB91, BP01, Dun87a, Dun87b, Dun88, SS87, Bra65]. monotonen [Alb62]. Monotonic [LM08, CE15, CP15, Che15b, GQ09, QL19]. Monotonicity [Che16c, LS88, CB12, GR92, GL01, Ism77, LM93, QGG10, Qi14]. Montana [BL91]. Montgomery [Ngu98]. Montgomery-Like [Ngu98]. Monthly [BC18]. Montpellier [KM07]. Mordell [BB16, Kuz15]. Mortici [CL16]. most [LM07]. motion [AB93]. motions [CZ10]. Motivation [Tod54, Tod55]. Moursund [Fik67a]. MP [KB90]. MPFR [FHL^{+}07, The21]. MPFUN2020 [Bai21]. MR [Fet81b, Fet81a, Fra81, Köl97, Lew87, Per87]. MTAC [T.57]. Much [Kah87, QW06]. Multi [Sch76b, Saf10, Zag24b]. multi-accuracy [Zag24b]. multi-center [Saf10]. Multi-Machine [Sch76b]. multi-precision [Zag24b]. Multidimensional [Dem89, Var95]. multigrid [BD85, RdL01]. Multilayer [HSW89]. Multiplace [Bar72]. Multiple [Bre75a, Bre75b, Bre76a, Bre76b, CM13b, EG97, Gar80, JKD92, Joh83, Kor11, Krä93, Smi91, Smi11, CV03, Deh89, FHL^{+}07, LN22, Smi89, Smi01, Tak00, Zag24a]. Multiple-grid [Joh83]. Multiple-Precision [Bre76a, CM13b, Kor11, Krä93, Smi91, Smi11, Bre75a, Bre75b, Bre76b, FHL^{+}07, Smi89, Tak00, Zag24a]. Multiple-term [EG97]. Multiplication [Alt79, BR73, GKK00, JJ65, LD89, RGK72, Rus98, WS06, CL95, EL89a, ITY97, McC77, MM91a]. Multiplication/Division/Square [LD89]. Multiplications [ST74, GLM15, Win70]. Multiplicative [Bak75b, SL95, Che71, ITY97]. multiplicities [Zen04]. multiplier [HFS99, SF93b, WG92]. Multipliers [Ahm89, ELMT00, WS05a, WG94b]. Multiply [BDL09, ZG87, ITY96, MM92, WMDM92]. Multiply-Add [BDL09]. multiply-additions [ITY96]. Multipolar [Hom98a]. Multiprecision [Bai93, BM15, CVW06]. Multistep [Ske86a, Ske86b]. Multivariable [KS20, ABFP24, DCL^{+}92]. multivariate [CKL89, Ein89, Gom89, Rac82, SLD16, Ter81a, vR89]. Multiword [JF16]. MultRoot [Zen04]. Murphy [Mei83]. must [Wil96]. Mutual [Kin21]. mutually [BPT76]. My [Yos97].

N [Köl87, Lip21, Kog58b, Kog58a]. Nīlakantha [Hay99]. nach [Töl68]. NAG [Sch77b]. Nakagami [SK11]. Nakagami- [SK11]. Nancy [Ano06]. narrow [MO18]. Narrowing [HJ96]. National [Ban01, L.61, Low49, Mul82, AEF92]. nationales [AEF92]. Natural [CH67, LP16, Bar14, Epp89, Nag68]. Naturwissenschaftler [Bar14]. Naukova [Köl90]. Navier [HBS00]. NBS [Loz97b]. Near [BS98, Fra65b, Zag19, Ehr89, Let97]. near-field [Ehr89]. Near-Minimax [Fra65b, Let97]. nebst [Rot70a]. necessary [Win70]. Need [BBM84]. needed [QW06]. Needs [Loz94, Loz96b]. négatifs [DGB80]. Negative [CP98, GRAST16, Tho13, Ada98, BS84, LVV05, NT84, Tem96b, DGB80]. neighborhood [Sko04]. neighbourhood [Mid75]. nel [AN68]. Nemes [Che15a, Che16e, LSM15, LM15b]. Nepreryvnye [PS93]. Nested [Dom03]. Nesting [BFHT85]. network [SR81]. networks [HSW89]. Neuenahr [MT85, MT86]. Neumann [Col78, GF88, Her61a]. Neuromorphic [MLS^{+}18]. Nevada [AFI69]. Neville [Håv79]. Newman [Bor82]. Newton [BC23, Bor21b, Car74, CHGM99, DK82, HR05, KMB94, KP69, KI04, MM91b, Mou67, Nin70, Ost37, Ost38, Vai89, WMC21, WS05b, Ypm95]. Newton-type [Vai89]. Newtonschen [Ost37]. Nikolai [Lip21]. Nimes [AEF92]. Ninth [Sin95]. NIST [Loz03, OLBC10]. No [Fet81a, Laz85, Wil70, ACG^{+}87, Chr62, Fra81, Frö63b, Joh87b, Lew87, Per87]. Nodes [HT13, JM18, Ben98]. nombres [Che88, Mul85b]. nome [FNC75]. Non [Arc43, Int03b, MCT18, OLHA95, PS05, Raf70, AB88, Lev73, LC96, LC97b]. non-characteristic [AB88]. Non-IEEE [Int03b]. Non-Integral [Raf70]. Non-linear [PS05, Lev73]. Non-Restoring [MCT18, OLHA95, LC96, LC97b]. Non-Sexagesimal [Arc43]. Noncentral [Rob69, GST15b]. nonconvex [Ubh89]. Nondiagonal [Dri95]. Nonlinear [BKdLSGL23, Cuy94, DB13, Hom94, MKY92, Wen89, Mus87, NS13, Wen96]. Nonnegative [Amo86, Amo90, Amo95, Kod08, SG98]. Nonrandom [Bat07]. Nonrestoring [MK71, Wal80]. Nonstandard [Ric88, Bel89]. Nonstiff [Cas83]. nonuniformly [FK88]. Norfolk [SKL93]. norm [Akr85, Glo89, JMS98, Rei86, HH18]. Normal [Cad51, Car75, Div79, Dre78, Ham78, Haw82, HJ67b, HD73, Hil73b, Hol70, KC76, Lin90, Lin09, Mac65, Mar04b, Mor80, Nor89, Pag77, PR82, PR96, Sha85, She54, Sho05, VLCSFN^{+}12, YBR11, Bad76, Bag95, BKKC09, Bry02, Cho14, DAV12a, DAV12b, KPPS08, Lee92, Lin89, Mac68, MZM94, MRS18, McC90, Rev90, SE14, Töl67a, Ved93, WR96, Wes05, Yun09, vR89, Ber68, Cyv64, HJ67a, HJ67b]. Normalintegrale [Töl67a]. Normalized [SOL81, Bun09]. normally [Ehr89]. Norms [BJK^{+}11, CJL09, GLT^{+}15, Mag94]. North [BCEP94, IEE89b]. notation [CK89a]. Note [Ano84, BH99, Cle55, Cra06, Gau99, Gla79, God01, Leh44, Mil57, NSR10, OS72, Sal83, Ver67, Boo55, Fis73, FW85, HR72, LK73, Pro83, Tuk48, Wri84, Zha96a]. Notes [Boe61, Boe62, Bur63, Bur64, Esc37, Fet67, Fra65a, GB68, Gla66, Har66, Har49, HKW60, Hea65, Hum64, JJ65, Jam99, Kal34, Lan60, Lon59, Lor89, Mec68, Mil58, Moo67, SF59, Tak66, Tho65, Tho66, TM68, Ver67, WC61, Wil68, WKG66, Woo67, HUY07, LM07, Mul24]. Nothing [Kah87]. Notices [Moo67]. Nov [FBM04]. Novel [Bas01, RJM^{+}20, SF10]. November [AFI69, Ano91, AEF92, IEE89b, IEE91, Sin95]. novembre [Ano98b, AEF92]. NP [RT89]. NP-complete [RT89]. ns [MIT^{+}95]. NS32000 [Cam86]. NSWC [DG99, Mor81, Mor84, Mor90, Mor93, TGC89]. NSWC/DL [Mor81]. Nullstellen [Mah30]. Number [BW00, DLM09, Fri67, GI01, Has66, Her61b, Hig18, Ngu98, OTY91, SH72, ST74, Tan89b, Bre74, Che88, GB90, Har14, Hon71, LP83, Per85a, KR86, Str59, VRS^{+}95, VRS^{+}99, Win70]. number-machine [GB90]. Numbers [Ano03, Ano06, BB88, BH13, CVV65, FBM04, GS99, Has90, KB67, Maj72, Sch16, CL11, DM91, Ele15, Fil92, Kul07, LKKH99, Mul85b, Nag68, SW88, Tre80, XC14]. Numerical [Abe88, BM15, Boi97, BO75, Bul65a, Bul65b, Bul69b, Car95, Cas83, CY18, CJL10, CH70a, CJR91, CJR92a, CJR92b, Cow77, Daw97, Dub83, DK90, Ehr61, FP68, Fra65a, Frö55, FI94, Gar79, GS81, GST07a, GST20, Gus66, HL13, Hul78, IP87, Jab94, JL12, KW21, Kin21, Kin24, Kin07, KZ90, Lau95, Lau04, Let01, LB90, LO94a, LO94b, Mat04, Mil58, Mil75b, Olv67b, POP17, PTVF07, Pri66, Pro87, Sti59, Tem75a, Tem00, Tem07, Ull90, Wat90, Wie99, AGA^{+}80, BGV81, Boy14, Bre18, Bre19, Bre77, Bre78c, Bre78b, Cuy94, Ein79, Ell83, EH16, Gor82, GP84, HvdR63, Hor19, Lan59, Mei64, Mei67, Mil75a, MN16, Mül93, NSU20, PBN93, RdL01, SGA81, SL83, Spi85, Tem76, Tem83a, Tod54, Tod55, VRS^{+}95, VRS^{+}99, Wil71, Zag24b, Zha96b]. Numerically [GST07b, Tal09]. Numerics [Nav83, AAS93, Squ91c, Squ91b]. numérique [Bre77, Bre78c]. numerische [Mei64]. NumExp [HL13]. NumSBT [Tal09]. Nuttall [Pak18].

Oak [Cow77]. Ob [DZKK77]. Oberhettinger [Coh12, Szm13, Wig67]. Objectives [Kuk71]. Obtaining [Sal53, SK71]. occasion [AGA^{+}80]. occurring [AJ03, Boe91]. Oct [Ull90]. October [IEE89b, IEE89c, Lav80, Sin95, USE88]. ODEs [LO93]. Odnomernyĭ [PS93]. Off [Kuk71]. Office [L.61]. Ohio [Fet81a]. Old [FR98]. Olsson.wl [ABFP24]. On-Line [Erc78, OE82, SMG94, PBE03, DHM89]. On-the-fly [EL89b]. One [BTT16, Gau65a, Hit57, LN22, Lir71, PS93, SLZ05, BBB89, Pól49, SF10, Spi85, Tes69, Car91, Jan77]. One-Dimensional [BTT16, PS93]. One-Half [Lir71]. one-step [Spi85]. One-Variable [SLZ05]. Ontario [Bro07, SIJ93]. Ontwikkelingen [Dop42]. Onvolledige [Dop42]. open-source [Joh12]. OpenCL [MHSK16]. Operand [EM04a, EM07, WS06, EM04b, ITY97]. Operands [SF16]. Operating [Rot70b, Sho86]. Operation [SF16, SF96, SF93a]. Operational [BC79, BDGP14]. operations [BWKM91, Flo61, Gon89, LBdDM16, Sch93a, Sch93b, TE06, VCV01]. Operator [FS96, von80, Sun88]. Operatorkonzepts [von80]. operators [AHM^{+}98, CG89, GS82a, GK89b, Hei88, Mit92, Whi82]. Optimal [AM89, Ge89, MM91b, Mou67, Wil70, von80, EH16, Fik67a, McC77, Meg89a, Roc82]. optimaler [von80]. Optimization [Dan77, DB13, Gly89, OLHA95, RdL01, Sin96, Pol88]. Optimizations [BJG^{+}99]. Optimized [AK09, ABJ19, BM04, MCT18, Por20, SK08, WO15]. Optimizing [LGML05]. option [SR16]. oracles [Boy14]. Order [Amo86, Amo90, Amo95, Ano46, Ant64, BH65, Car75, CMF77, Cur64, FK11, FR76, GCP67, Gup60, Hit57, Kod08, Kod11, Lau73, Let01, LB90, Lir71, Mec68, Olv64, Olv67a, Osi13, PT91a, PT91b, Rap94, RL80, SW99a, Soo73a, Soo73b, Abe16, Abr45, Ada98, AG08, AM78b, AM79, AM84a, AM84b, Bar81, BST13, Bre18, Cam79, Cam81, Cam84b, Cam84a, CP66, Coc65, Col80, Col84, CS82, DAV12a, DAV12b, EM94, Fuk15d, Fuk15b, GST04a, Gla74, Gra61, GM88, Hai60, HR07, Jon06, KK96, KWY07a, KWY07b, KRVZ98, Kra85a, Kra85b, Lev91a, LS95, MG89, NR15, NP18a, O’B80, Olv67b, Rap01, RdC93, RX07, KR86, Sca71, Sch93a, Sch93b, SS65, TB86, TB87, Tho04, XC15, de 77, dT93]. order-differential [KR86]. Orders [Alh00b, Fox54, GST02d, NDT69, ND70, Alh00a, Bre19, CT67a, CT67b, GST03a, GST04b, Hor19, IKF91, SG98, SG99, Wal84]. Ordinary [AO08, Bra73, Cas83, Tem79a, Sca71, Tem76]. Oregon [BCDH09, CLT20]. organization [Sal76b]. Oriented [Erc77, MKY92, PP12, AGA^{+}80, OY91]. Original [Chl11]. Orlicz [Mus87]. Orthogonal [BGVHN99, DGA18, DX01, DX14, Hoc64, Hom98a, IA20, KV03, PGA18, SC01, Sue99, VB04, CKL89, CV03, DM08, El 06, Elb01, Hom94, Pas08, zCFF05]. OS/360 [Kuk66]. Oscillating [UE05]. oscillatory [BO75, CY18, KW21, Vep08]. Other [Bal00, Fox60, L.61, LC61, MR71, Olv91, Air37, Boy14, Squ91b, SR53]. outer [Fuk83]. outline [CHGM99]. Overconvergence [Ehr89]. Overflow [KNS95, LM07]. overhead [MI97]. Overlapped [PZ95]. Oxford [Boi97].

P [Coh12, Fet81b, Köl97, Kun81, Szm13, Tal68, Wig67, Frö63a, L.61, T.57]. P.C. [WNO94]. P.D.E. [Fri72, SM70]. Pacific [Sin95]. Package [Amo86, Amo90, Amo95, Cod75, Cod91, Cod93a, Cod93b, Dri89, Kod08, Mat87, Sal92, Smi91, Tan90c, BC09, Bai21, Dri91a, Dri91b, Dri91c, GST15b, Joh12, KRVZ98, Mac96, Squ91a, Squ91d, Tan91a, VRS^{+}95, VRS^{+}99, Zen04]. Packages [DPC95, ABFP24, BK16, Kra95, Squ91b]. Padé [AGJ86, Bor88, Dri95, DF78, GSS12, KSVW07, Sab08, WdZ04]. pair [HFS99]. Paleo [BB12]. Paleo-Mathematics [BB12]. paper [Tra60]. Paperback [L.61]. papers [Spr00, Boe61, Boe62, Bur63, Bur64, Fet67, Fra65a, GB68, Gla66, Har66, HKW60, Hea65, Hum64, JJ65, Lan60, Lon59, Mec68, SF59, Tak66, Tho65, Tho66, TM68, Ver67, WC61, Wil68, WKG66, Woo67]. Parabolic [Fox60, Jon06, L.61, Mil64, SG98, WR71, DGS24, GST06a, GST06b, GST11b, LR74, MMV81, RL76, SGA81, Tem00, GST11a]. Parallel [Bak75b, CA00, Gar75, Gar80, LC97b, LO93, RHMLL08, SS93b, SKL93, Tan87, BBC^{+}91, GIJL23, Joh88, Mat90b, PNV01, Sg92, Tak00, Tak22]. Parallel-array [LC97b]. Parameter [Sau92, Töl68, AG08, DD76, EC13, Fuk09b, Fuk15c, PS78, Roc82, YM03]. Parameters [DJ67, NPA16, Won73a, BS00, Bus74, GS88, GOST23, Har20, KWY07a, KWY07b, MM95, Tem87, Tem94a, Tem96b, Töl68]. Parametrized [RJM^{+}20]. Park [IEE89b]. Part [Fet81a, NPA16, SZ70, Kru99, LU96, Ano52, Ano60a, Ben83, Cas83, SZ74, SZ76]. Partial [AO08, Büh03, HBF09, Sch16, Bus74, MKS83, War75]. Particular [Völ83]. partition [FNC75]. partitions [AGJ86]. parts [Tim87]. Pascal [von80, MKS83]. past [Bre85]. Pathways [Meg89a]. Patras [SC01]. pattern [Sid06]. patterns [PI16]. Patterson [Fet81a]. peaks [Phi79]. Pendulum [PS05]. Pennsylvania [Tra76]. percentage [DH77, Fet79]. Percentile [FC60]. perfect [Dun90, dDM06]. Performance [ASA06, BTDS11, Cod90, Gal85, SK11, SA00, SL95, CL00, CP99, Gal86, LN21, MMW91, MM92, WMDM92]. Performance-Driven [ASA06]. Perimeter [Adl12]. Periodic [DR89, Wan78, JW88, KS88, Lem88]. periods [Tod90]. Personal [Egb77]. perspective [BIS01]. Perturbation [Hom96b, Boy14]. perturbed [JS89]. Phase [CH70b, FKY86, CS83, Hor17].

Phase-integral [FKY86]. phenomena [Sti88]. Phi [Hin77]. Phys [Tho04, VRS^{+}99]. Physical [Flo15, L.61, Mat93]. Physicists [AW05, AWH13c, Arf85, AW01, Fet81b]. Physics [GI01, HBF09, Hoc71, Hoc86, JS56, JS99, PC11, Sne56, Sne61, Sne80, SS70, Tem96a, VS04, VS10, Ars68, Coh12, Her82, Hoc61, MO49, MOS66, NU88, Sne63, Veb12, Szm13, Wig67]. Physik [Sne63]. Pi [BBB97, BBB00, BBB04, Bre17]. Pick [Ped03]. Piecewise [KP75, Qua18, SDP11, Fuk15b]. Piecewise-Polynomial [SDP11]. pipeline [OJ04]. Pipelined [Bru22, CCCP99]. Pipelining [TE06]. Pitfalls [Bol06]. Pittsburgh [Tra76]. PLA [EL89c, EL90]. Places [Knu62a, Sal53]. Plane [BDG^{+}72, EW76, BDG^{+}84, CM83, GJL20, Töl67b]. platform [CTA22]. Plesset [Hom96b, Hom96b]. Plethysm [Car90]. Pocket [McC90, Nor89, Pag77, Lin89]. Pocket-Calculator [Nor89]. Pocketbook [Bar14]. Poincaré [Hau88, Lev91a, Pas92]. Point [AAHTH10, BM04, Bli97, Bra87, CH98, DGS65, D^{+}89, DHL^{+}04, GB91, Has90, Hul78, IP17, JKMR11, JMMW79, Kah80, LP16, LP17a, LP17b, LM15a, LD89, MMNP91, Meg89b, Mey01, MLS^{+}18, Rev16, Rus98, Sch88b, Smi91, Tan89a, Tan90e, Tan92, WS05b, dDLM11, dDI15, Ano67, BBC^{+}91, BS06, BWKM91, BCK^{+}93, Bol06, BMR04, CG24, CHGM99, Cro92, FHL^{+}07, GLM15, GM88, HHR00, Hou81, HS20, JM17, KMB94, KMY89, Krä88, Kra98, Kul07, LBdDM16, LKH93, LC97a, LMOT01, LN21, MT72, MG89, MPG92a, MPG92b, Mat90b, OY91, Obe99, OJ04, PKS00, Rus99, SF93b, Sko04, Smi01, SL97, Ste89, Ste08, TE06, TRHH94, Vav89, VCV01, WN03, dDM06, dDP10, dDIS13]. Point/Integer [D^{+}89]. Points [Bra87, BEM12, FCC73b, FC60, WMBE14, Bos89, DK82, DH77, Fet79, Oli75, PA86]. pointwise [Som82]. pointwise-Lipschitz-continuous [Som82]. Poisson [Gil16]. Polarization [FR76]. Polya [EH89, MRS17, MRS18, Sta17]. Pólya-based [MRS17, MRS18]. polycrystals [AM89]. Polygamma [Din58, Lew75, BHJ05, BS83, CE15, Dav35, Di 86, Köl96, Lóp98, ÖE16, QGG10, Ada98, TS69]. polylogarithm [Cra06, Cra12a, FK75, Vep08]. polylogarithmic [BBP97, Fre05]. Polylogarithms [Bar74, Lee97, Lew81, Lew91, Woo92, BB16, NSU20, BD14]. Polynomial [ADN03, AH10, BCE^{+}08, CPCD19, Cle54, Cod65b, Eis63, ECSD20, Fik67b, Fra65b, Hit57, JKMR11, LB09, NDT69, Qua18, Ric65, SR53, ST74, SDP11, Wim62, Bos89, Boy14, CB09, Dor62, DM88c, Fuk15c, Kon76, LN22, PS86, PC93, Rei86, Zen04, Wim72]. Polynomials [ABJ19, Bar61, Bra73, Cod67, Dan82, DGA18, DX01, DX14, DM88b, FP68, HG64, Hoc64, Hum64, IA20, KR79, MH03, PGA18, SC01, Sko75b, SOL81, Sue99, Tho65, Wit68, Ano84, Bel58, Boo55, Bor10, CIM94, CV03, DM08, DGS18, DK77, DZKK77, El 06, Elb01, EMR82b, Eve64, FS92, GST17, HS20, Knu62b, Knu63, Knu98a, KV03, Kra85b, KI04, Mot55a, NR15, Pan59a, Pan59b, Pow88, RW72, Rac82, Rap01, Ric92, Sal83, SL83, SZW11, Tes69, Tim87, ZZ11, zCFF05]. polytype [Vai89]. pomoshch’yu [Saa91]. Popov [Köl90]. porous [YH89]. PORT [FHS77]. portability [CTA22, Cow77]. Portable [Amo83a, Amo83b, Amo86, Amo90, Amo95, Bee17, Cod91, Cod93a, Cod93b, Ful77, KDDH92, KDDH94, Kor11, Rum01, SP20, Tan88, Tan90c, ZOHR01, AHM^{+}98, Tan91a, VRS^{+}95, VRS^{+}99]. Portland [BCDH09, CLT20, IEE23]. Posit [OJL22, XLW^{+}20, RJM^{+}20]. Positive [CPCD19, Pie84c, TM68, Bar84b, CG89, GST04a, GST04b, HR88, IS88, IS90, IKK91, IS92, LW96, Let96, LU96, NP18a]. positivity [Ric92]. possible [LW96, LM07, LU96]. Post [dDEG05]. Post-ultimate [dDEG05]. posteriori [Kai89]. Potential [MS68, Sti88, Vai89, Wil96]. Potentials [FR76]. pow [LL09]. Power [NKY08, Rei69, SF16, Sta17, Wyn60, BBD03, BF92, Gus78b, Gus84, HJZ23a, HJZ23b, HJZ^{+}24, Kat78, KP98, MI97, NL99, Wen10]. Power3^{T M} [AGS99]. powerful [SW88]. Powering [PEB04]. powers [Boy09, GLM15]. pp [Lip21]. Practical [Ike76, Sid03, Luk56a, Luk56b, Töl43, Töl50, Töl66, Töl67b, Töl67a, Töl68, Töl69, Töl70]. Practice [BR91]. Practitioners [Tho97b, Tho97a]. Praktische [Töl43, Töl50, Töl66, Töl67b, Töl67a, Töl68, Töl69, Töl70]. Precise [Abe88, Fuk11, Fuk13a, Fuk15d, Fuk15b, Fuk15a, Fuk15c, HJ96, IM99b, LA85, Wil12]. Precision [Ahr96, BY07, Bre76a, Bre06, BJK^{+}11, CHG^{+}11, CM13b, Cod64, DPC95, Fat16, God20, HA85, HA86, Joh14b, Joh15a, JM18, Joh21, KM97, Kin65, Kor11, Krä93, Lan64, LP16, LP17b, LL09, LM03, Let01, Lew75, LE95, PB02, Pot14, SH72, Sch88b, Smi91, Smi11, SG72, Swa65, TS69, Yar67, Ano14, Bai21, Bar82c, Bre75a, Bre75b, Bre76b, BM80, Che09, DAV12a, DAV12b, DSK83, DSK84, EM03, Fer86, FHL^{+}07, FW80, Fra81, Gab79, Gau11, HXX22, Hon71, Joh14a, Joh15b, Joh20a, Joh20c, Joh20d, KM93, KO94, KA71, LC97a, Mar00, Mar03, NP18b, Par09, QW06, Sca71, Smi89, Smi01, Spe72, Tak00, TE06, VCV01, WG94a, WG95, XLX^{+}83, XL94, Zag24a, Zag24b, dDLM04]. precision- [VCV01]. Predicting [AMM21]. Predictive [Bak73]. prefixed [HR07]. Preparation [WWG51, WWG57, WWG82, RW72]. Prescaling [EM04a, EM07, LM95, EM04b, LM92, LM99]. prescribed [Hay99]. Presentations [Bor02]. Presented [Bir54]. Press [Lip21]. Priblizhenie [PT80]. priblizhenii [DZKK77]. priblizheniĭ [Saa91]. prices [SR16]. prilozhenii [PT80]. primal [KMY89]. primal-dual [KMY89]. Prime [JF16, Per85a]. primitive [Dri91b]. Primitives [Laz85, EM03]. Princeton [HDG^{+}15]. Principal [Har97a]. Principled [OLHA95]. priori [Krä87, Kra98]. priors [GK89a]. Probabilistic [Chi20]. Probabilities [GSS12, GW91, Whi63, Bag95, Düm10, SLD16]. Probability [Car75, Div79, Lin90, LB09, SZ70, ZS64, CD02, CDS03, Fei78, Lin89, Mar88, McN73, Pól49]. problem [Boe91, EC79, Gre82, JS89, JMS98, Mag94, Sko04, SW88, Via89]. Problems [AO08, Bre82a, Cas83, Ovt03, Ric68, Rob55, Bel58, CMV69, Cro92, OSS09, Ost54, Pol88, RdL01, RT89, Spi85, Was89, Wen01]. Procedure [Gau79b, Di 86, MH72, Ye17, HJ67b, Mac65, Mac68]. Procedures [Wyn62, Cod82]. Proceedings [AGA^{+}80, AGL93, Ano91, ???90, ES89, Hwa85, IEE78, IEE81a, IEE81b, IEE89c, IEE91, IEE99, IEE23, IP87, IL07, KM95, KM91, Lak96, Lav80, Luk99, MT85, Mul82, SE11, SIJ93, TT18, Tra76, Ull90, Ask75, Ban01, BL91, BC01, CDS00, CLT20, Koo91, KK99, LMT97, MMH89, Mar85, MC87, MC90, Mil75a, MT86, Ric92, T^{+}97, USE88, Wah04, Zah94, ACM89, Ano06, BCD^{+}20, Boi97, BCEP94, Bro07, BCDH09, DIW00, HY14, IEE05, IEE13, KM07, Lan59, Men06, MS94, SO11, SC01, SKL93, Spr00, FvdHJT10, Gre16, IEE21, IEE22, RDK98]. Process [Whi63, Bel90]. Processes [SK71, Bre83, Ebe89, Har49, Ruy89, Wen01]. Processing [Lav80, SKL93, JL94, Luk99]. Processor [D^{+}89, Dav81, Nav83, Rus98, AGS99, DMS95, Mar90, Mat90b, PKS00]. Processors [IEE89c, IEE08, Rev16, T^{+}97, VDR05, BBC^{+}91]. produce [LN22]. product [AB88, Gla74, LW89, Max91, MS15, VC06a]. Production [Sch76b, Sch77b]. Products [GRJZ07, KKI67, MS68, Sol69, XD18, LK73, RKZ^{+}14]. Professeur [AEF92]. Professor [AEF92]. Profile [Dav59]. Profiler [PI16]. Program [Bar76a, Bar76b, Tem79a, BDM81, Bar84c, DSK83, DSK84, DF84, LMS73, MKS83, NT84, SSG^{+}18, Wen64, Wie99, Zei91]. Programmable [LA85, LE95, LE93]. Programming [Bee17, Chr62, DGS65, Frö63b, ISO10, Meg89b, Mul82, Ano67, Chr65, Gon89, Joh88, KMY89, Meg89a, PUHM12]. Programs [Bai93, BBM84, Cah54, Cod90, Köl72c, Köl84, Mos89, TGC89, Tho97b, Tho97a, WNO94, WWG51, WWG57, WWG82, BW89, Dam94, Fuk83, Gon89]. Progress [FLR03, Meg89b]. project [GST14b, ES10, IA20, KS20, Loz00]. projection [Som82]. projections [Via89]. projective [Via89]. Prolate [LP61, LP62, MC04, Osi13, SP61, Sle64, Sle65, Sle78, RX07, SS65]. Prony [LB09]. Proof [BMZ02, Chi20, Rus98, BMZ06, Bol06, Leh83a, Leh83b, Mim88, Rus99, Zim00]. Proofs [The21, dLMF23, CHGM99]. propagation [GK70a]. Properly [HA85]. Properties [Bar10b, Kui52, Lew91, RGK72, Sch76a, Sho02, Sie76, AB88, BG84, Che11, CB12, Che16c, GR92, GL01, Her82, Hig77, Kra89, Mak80, Mat92, Mit49, MC04, MO18, Ris79, Ruy89, Tre80, Tri50]. property [AR99]. Proposal [BCK^{+}93, DHL^{+}04, OGWY94]. Proposed [AOB64, Dri89, Dri91a, Dri91b, Kah80, Loz96a, Loz97a, Sal92, Squ91a, Dri91d, Dri91c, Hou81, Squ91d]. Provably [Dun87a, Dun87b, Dun88, Dun94a, Dun94b]. proven [dDLM04]. Prover [Bor02]. provider [Ste08]. Proving [CH98, MDM16, Bel89, Zei91]. proxy [Boy14]. pseudo [Bre74]. pseudo-random [Bre74]. Pseudodivision [SK71]. Psi [Amo83b, CST73, Din58, Moo67, SB71, Tho65, Alz97b, AR99, Che11, CB12, Che13a, CE15, Che16a, Ele15, Kat78, MM95, QL19, XC17, YCZ15, Wim74]. Punch [Gel51]. Punkte [Bra87]. pupils [Bir54]. Purdue [Zah94]. PVM [PNV01]. Pythagorean [Dub83, MM83]. PZ [CDS00].

Q [FK11]. Q-Function [FK11]. Qrd [LF92]. Qrd-based [LF92]. quad [Mar03]. quadrant [HFS99]. Quadrate [RS98]. Quadratic [Car92, POMB05, Car91, CF00, Dri95, MJ13]. quadratically [Sch95]. Quadrature [Ben98, Gau16, GST03b, Gus66, HT13, JM18, KP03, Akr85, FO93, Gau02, GF97, GST02a, Kza92, Kza99, LW95, SL83]. Quadratwurzel [Rot70a, Kun81, Rix82]. Quadratwurzel-Algorithmus [Rix82]. Quadratwurzelfunktion [Hol69]. Quadratwurzeliteration [Rut63]. Quadratwurzeln [Die83]. Quality [Boi97, Sht19, Atk79]. quantifier [EH16]. quantile [SE14]. Quantiles [Hil70b, eL79, Hil81b]. Quantisation [PS05]. quantities [DP92]. Quantum [MCT18, Var95]. Quartic [Flo15]. quartically [Bai88]. quasi [Cri82, GM88, MG89, MPG92a, MPG92b]. quasi-analytic [Cri82]. quasi-fractional [GM88, MG89, MPG92a, MPG92b]. quasiconfluent [AS03]. Quasipower [Pas95]. quaternary [CA00]. Qubit [MCT18]. question [PKP74]. Questions [BL71, Ded26, Esc37, Kal34, Arm82]. Quick [Tho93]. quicker [LM15b, LSM16]. Quotient [Rus13, Che71]. quotients [Ism77].

R [Coh12, Sha76, Szm13, Wig67]. R66 [Has66]. R66-78 [Has66]. R68 [Bin68]. R68-38 [Bin68]. Radial [BC62, Pow88]. radiation [OSS09]. Radici [RS98]. radius [KT59]. Radix [Bru22, BTDS11, CL94b, Erc72, Erc73, EL89c, EL90, HSE^{+}24, LM95, LA01, LA03, MIT^{+}95, Nan11, PBLM08, PZ95, Sal51b, Sal53, ALB98, Bak73, Bak75a, Bas01, BH01, CM90, CP99, Fan89, HF95, LM92, LM99, LO95, Mer94, MC89, MC91, MC93, MC95, NL99, PEB02, PBE03, Sch93a, Sch93b, SF93b]. Radix- [EL90, LA01, LA03, PZ95, Bak75a, Bas01, BH01, CP99, HF95, LO95, MC95, NL99]. Radix-16 [Erc73, Nan11]. Radix-2 [PBLM08]. Radix-4 [EL89c, HSE^{+}24]. Radix-64 [Bru22]. Radix-8 [BTDS11]. Rahman [IK05]. rail [MI97]. Raleigh [BCEP94]. Ramanujan [BBB89, Che15a, Kar01, LSM15, Mar86, Mim88]. RAND [War08]. random [BCJV18, Bre74, Ruy89, Sho86]. Range [BHK09, BDK^{+}05, Gau16, HKA93, JSH^{+}11, Joh15a, SOL81, DMMM95, GF88, Her61c, Joh14a, KPPS08, Mas83, MV98, VC06a, VCV01]. range-independent [VCV01]. Range-Reduction [BDK^{+}05]. Ranges [Bin68]. Raphson [CHGM99, KMB94, MM91b, Mou67, WMC21, WS05b, Ypm95]. Rapid [Esc37, JJ65, MZM94, NSU20, BBP97, Bre19]. Rapidly [FGG82, Jam89, Kat78, McC81]. Rapidly-convergent [FGG82]. rapport [Del81]. Rate [Chl11, AL20, DF78, FS92, GK89b, Wan82]. rates [BD85, SS98]. Rathbone [Sha76]. Ratio [BB74, DJ67, Hin77, OM68, PSB76, AS05, BS78, Gom89, Lud63, McC83, Mer05, PH67]. Rational [AL20, BSW95, BM04, Bla74, BEJ76, BEJ78, BPD81, BCE^{+}08, BF23, BGVHN99, CG55, CDR71, CT67b, CT68, Cod69, Eck76, EW76, FL67, Fie65, Frö61, Frö63a, GB68, GP64, GSS12, KZ90, Let96, Nin70, Ric65, ST07, Sha76, Squ70, VB04, WC90, WBR82, Wyn60, AQ18, Ali11, BZ92, Bra84, CL61, CMV69, Cuy94, Dan77, DAV12a, DAV12b, Eck77, Ell83, Fik67a, FH62, Fuk15d, Fuk15b, IP81, Jam99, Kra85a, Kra85b, Lan80, Let97, Lit93, Lit94, Luk56a, Luk56b, Luk76, Mag94, Mar86, MM95, Mot55b, RW72, Saa91, SK08, Sie76, VT11, CT67a]. Rationale [Dri89, Dri91c, Squ91d]. rationals [DSK83, DSK84]. Ratios [Amo73, AVV90, Che72, DM86, DM87, DM88a, DM92, GST12a, XD18, Amo74, Che15b, Del81, DH76, DiD84, Gau64c, Gau64d, Gau77a, GS78, YA24]. Raton [Lip21]. ratsional’nykh [Saa91]. ray [Phi79]. RCWFF [Bar76a, Bar76b, Bar84c]. RCWFN [Bar76a, Bar76b, Bar84c]. Real [Ano03, Ano06, Bar76a, Bar76b, Bra87, CP98, CMF77, DLM09, DiD78, EH16, FBM04, GKS04, GZ75, Krä87, Mor76, Mor79, Mül90, RL80, Ric68, Soo73b, Soo73d, Sou57, Wan74, Zag19, AM78b, AM84b, BFSG74, Bar81, Bar82b, Bar84a, BFSG84, Bar84b, Bar84c, BY07, Bat19, BF92, Bre19, Bro89a, CG24, Cam79, Cam81, Cam84b, Cam84a, Car95, Cob69, CS82, Fab02, GS88, GST06b, GS12, Gon52b, Krä88, KRVZ98, Mul85b, NP18a, Pan59a, Pan59b, PS78, Ric88, SG99, T.57, TB87, Tho04, Tur72, ZA24, GST06a, HJ67b]. real-imaginary [BF92]. Realistic [Fra78]. Reasoning [BCD^{+}02, CDJ^{+}01]. Rechenverfahren [Hen77]. Rechnerarithmetik [von80]. Recipes [PTVF07]. Reciprocal [ALMN05, Bor21b, BTDS11, EIM^{+}99, EIM^{+}00, Int03b, IM99a, IM99b, LA01, LA03, PB02, Tak01, WS06, Bor21c, BJM22, CLZ04, FW84, GEK04, JL94, Sei99, SW50, Wri84]. Reciprocals [BTDS11, SW99a, CO86, Sin96]. Reciprocation [ELMT00]. Recoding [ZCL15]. recognition [MS82]. record [Rah01, Sin95]. recourse [BW89]. rectangles [Man23]. Rectangular [WG94b]. Recurrence [Act74, ALMN05, Fat16, Gau67, GT59, HSE^{+}24, LA01, LB90, Lew85, Lew87, Mec68, OS72, Olv88, Pré10, Sch77a, EL94, EM03, Fre05, HS20, HT73, Lev91a, LE93, VC06b]. Recurrences [Tha79]. Recursion [GM04, LDP93, OSS09, Wim68, Wil71]. recursions [GST07b]. Recursive [Chl11, Ekl01, Fuk13b, Gau61, Gau63, GK70a, Gau99, SG76, XD18, GF88, SS10, SS23]. Recursively [Woj61]. Redheffer [ZS08]. Redheffer-type [ZS08]. Reduced [CM13b, CS83, MKS83]. Reducibility [JMMW79]. Reducing [MC93, MC95]. Reduction [BDL09, BDK^{+}05, Car70a, CF00, JSH^{+}11, MM95, NC66, Shi93, Smi95, SLZ05, BK16, DMMM95, Gra02, HTHR94, VB81, vRdL88, BK16]. redundancy [HF95]. Redundant [TAY91, LKKH99, LP83, MC93, MC95, Sei99, ZXBZH01]. reelle [Bra87, Krä87, Mül90]. réels [Mul85b]. Reference [Car74, HH18, Mor75, WWG82]. refinement [KI16]. Reflections [Cra12b]. Regency [IEE89c]. regimes [YM03]. Region [IEE85, Fet74, MO18]. Regional [Men06]. regions [ABFP24, ND19]. Regular [AS95, CP70, Ike75, Ike76, Sha76, Bar81, Bar82c, Nes84, SSG^{+}18, Gau66b, Gau69b, Köl69a]. Reihen [Bau89]. Related [Abr64b, CR68, Dav64, Fle68, GC64, Har66, Hit57, KK20, Lip21, MR71, Meg89b, Pri66, Sch16, Smi11, Sou64, SJ01, SZ70, SZ76, BP01, Che11, CP15, Che16e, CL16, DGS18, FW80, Fra81, Fre09, LM08, LSM14, LM15b, LSM16, Mac72, Mit49, Ped03, Qi14, Sch78c, Smi01, Uni49]. Relating [BC62, Boe61]. Relation [DGA18, PGA18, FKY86, HT73]. Relations [Gau67, GM04, HP62, Lar69, LB90, Olv88, Act74, Cri82, Fre05, Gra02, Lev91a, Lew85, Lew87, Wil71, Ye17]. Relationship [Hin77]. relative [GLM15, JM17]. Relatively [Zag17]. Reliable [CVW06, Mac89].

Remainder [CH98, CHI^{+}03, Int00, Int03a, LKK19, Qi14, VCV01, Wil12]. remainders [MC93, MC95]. Remark [Amo90, Amo95, And00, BB74, CP70, Col78, Fri72, Gau69b, GK70b, Goa97, Gra63b, HH18, Hil81a, Hil81b, HP85, Hol70, Jan77, Jef62, Kod07, Köl69c, Kop74, Mac68, Mor76, PH67, PSB76, Pit67, Pom76, RS81, Red70, Sch78b, Sko75b, Sko75a, Sko78, Sny96, Vos73, Zag16, Zag19, eL76, eL79, Bau89, Ber68, Köl69a, MC95, Sny21]. Remarks [Fri72, Håv81, HJ67b, Köl72d, ÖE16, Pól49, Hit75, Luk75a, Rut51]. Remedy [Zag19]. Remez [Ree90]. Rep [Fet81a]. Repeated [Gau61, Gau77b, Gau16, Sug58]. Replacing [MM83]. Representation [DGA18, GRS87, Kaz65, Mul85b, PGA18, Var95, VK95, Mul85b, CS11, FF54, FF57, Kar01, Mar95]. Representations [ADN03, LW89, Mat04, NS09, ABFP24, Bus74, Die80, Dut86, Ism77, KRT02, LKF10, LN21, LN22, Qi14, Vil88, WS84, Wen07]. representing [Hai60]. requirements [Wil96]. Res [ACG^{+}87, Fet81a]. Research [Ask75, IEE89b, IEE99]. residual [Kaz65]. Residue [OTY91]. resource [BCR^{+}07]. Respect [Lau73, AG08, Bus74]. respectively [Kra95]. Response [Sho02]. Restoring [MCT18, OLHA95, LC96, LC97b]. restricted [vdD88]. restrictions [Spi85]. Result [AK93, Krä93, Che16b, MC93, MC95]. Résultats [DGB80]. resulting [Che15b]. Results [Lef05, SF16, Sri70, DGB80, DHM89, Ein89, Elb01, Fdi97, GAZK03, JJW86, LS88, LM93, LN22, Per85a, SW88, VB81]. Reversion [FO99]. Review [Ars62, Köl87, Köl90, L.61, Lip21, Wig67]. Revised [Dun94b, Loz97b]. Revision [MHSK16]. Revisited [CG02, Eis63, SZ05, BF92, Par16]. RFSFNS [VRS^{+}95, VRS^{+}99]. Ricatti [Bra70]. Richard [Bir54]. Richardson [Fil66, Fil66]. Richardson-Algorithmus [Fil66]. Riemann [BM15, BH18, Bor95b, Bor95a, BBC98, BBC00, BCR^{+}07, CHT71, HP62, HG64, Kol70, Lee97, Leh70, Leh56, Mar65, Rza̧12, Sko05, TKK17, Tur53]. Riesz [TV09]. Right [SL97, Bry02, GJL20]. rigid [CZ10]. Rigorous [Joh14b, Joh15b, JM18, Rum01, Yoh70, WO15]. Rigorously [Joh19, Joh16]. RISC [Mar90, BKdLSGL23, GZC22]. RISC-V [BKdLSGL23, GZC22]. Risch [Laz85, Mos69]. RMM [Sho02, Sho05]. RMM-Based [Sho05]. RNC [Ano06]. RNC5 [Ano03]. RNC’6 [FBM04]. Robust [Mac89, DR04, VT11]. ROC2.wl [ABFP24]. ROM [Spr00, GP84]. Roma [Ano98b]. Romanian [Hai60]. Romberg [EM86]. Root [ARH14, AH10, Ale77, AMT78, ALMN05, AMS76, BV85, BO93, Bee91, BMZ02, BR73, Bor21b, Bru22, BH07, Cat86, Cod64, CH98, CHI^{+}03, CL94b, Cow64, D^{+}89, EL94, EIM^{+}99, EIM^{+}00, ELMT00, EM04a, EM07, Esc37, Fik66, FR98, Fri67, Ful99, Gar75, Gar80, GBKK09, HAK91, HSE^{+}24, Hei96, Hol69, HMO19, HA85, Int00, Int03a, Int03b, IM99a, IM99b, ITY95, Joh87a, Kah99, KM97, Kin65, KP69, Kog59, Kor05, KNS95, Lan42, LM95, LA01, LF92, LE95, LD89, Mar04a, MIT^{+}95, MRH19, MKY92, MM91b, MCT18, Nan11, Ngu98, Nin70, OLHA95, OE82, Par06, PB02, PBLM08, PB08, PZ95, RJM^{+}20, Rix82, RHMLL08, Rol87, Rot70a, Rus98, Rus13, SH72, SK71, SG02a, SF96, SL95, SL97, Swa65, Tak01, Tha61, Wad58, WMC21]. Root [Wal80, WS05b, Wil70, WS06, XLW^{+}20, Yon70, Zim99, ZG87, AHM^{+}98, AGS99, Ano12, ALB98, Bas01, BMZ06, Bic81, Bor21c, Bra84, BMST97, BH01, Bur82, CF00, Cat85, Che71, CH89a, CH89b, CL94a, CL95, CL00, CHGM99, DD76, DSK83, DSK84, EH16, EL89b, EL89a, EL89c, EL90, EM03, EM04b, Fan89, GREL96, Har03, Has00a, Has00b, Hay99, Hig93, Hit74, HF95, Hon71, HT73, HFS99, ITY97, JL94, Joh11, Joh87b, KM93, KMB94, KP98, KI04, LM92, LM99, LA03, LKH93, LKKH99, LO95, LC96, LC97a, LC97b, LE93, LMT^{+}92, Lyo91, Man23, MST89, MI97, Mat90b, MMW91, MM91a, MM92, MMH93, Mit92, MC89, MC91, MC93, MC95, MWH^{+}16, MWH^{+}18, NL99, Obe99, OJ04, Ped80, PS86, PR04, Pro83, Rus99, Rut63, Saw02, Sch73]. root [SF93a, SF93b, Sch95, Ste89, Ste08, SL63, Str59, Tak00, Tay81, TVG00, TE06, Tom00, Tuk48, VCV01, WMC18, WN03, WMDM92, ZXBZH01, ZLC04, Zim00, ZA14]. root-extraction [Hay99]. rootfinders [Boy14]. rootfinding [HP93]. Rooting [Alt79, Erc78, Gar79, Has90, Maj85, Met65, RGK72, CCC96, CCCP99, CM90, CP99, ITY96, KB89, KK98, Liu95, MM90b, Par99, WG92]. Roots [ADN03, Ahr96, BB12, BFHT85, BJK^{+}11, BTDS11, Cam86, Car74, Che72, CG85b, Cre98, Ded26, Die83, Eve63, Fet76, Gar80, GLR07, HCK09, JJ65, JKMR11, Joh89, Kal34, Kun81, Lir71, MK71, Maj72, MM83, Mül04, Mur88, Pat18, PS77, PS78, Sar59, Sch84, Sch85, SW99a, SW99b, Sug58, Tho93, Tor02, Tra60, Wej83, Yoh70, Yoh73, Abe16, BH99, BJM22, Bur82, CL11, CO86, Cre91, CLZ04, Dut86, Egb77, Fre81, GS12, HHR00, HR05, HR07, Jam89, Jam99, JM17, KK96, KCYL06, Leh56, Mur99, NS13, Özb06, Pen81, Per85a, PC93, PS79, PNV01, QW06, RS98, Sar60, Sin96, SD85, Tur94, VRS^{+}95, VRS^{+}99, Yey92, Zen04]. Rotation [ZCL15]. rotations [Bor21c]. Rough [Gor82]. Rounded [Bor21b, God20, GLT^{+}15, HA85, Kah96, LA01, SS94, Ziv91, Ano14, BJM22, BMR04, BHMZ24, CLZ04, HJZ23a, HJZ23b, HJZ^{+}24, LMT98, LM00, LN21, LN22, SS93b, dDLM04]. Rounding [BJK^{+}11, BTDS11, IM99b, LM95, LL09, LM03, Lef05, PB08, RN96, Sch95, TBDS11, ALB98, EL89b, FHL^{+}07, LM99, LN22, Mul24, Ost37, PEB02, PBE03, SS93a, Ziv99]. Routine [Ale77, Kin65, Swa65, Tur69, Cat85, Fre81]. Routines [Cod93b, Fik67b, Ful77, Sch76b, WB78, CZ10, GST02b, GST02c]. Royal [MC87, MC90, Mil75a]. Rule [CG02, Ric73, Dor62, Ost54]. Rules [GST03b, BDGP13, GST02a, Hay99]. Runge [Cas83]. runs [Bre18]. Russian [Köl90, Bar72, BSY82, BL71, Dan77, DK77, DF78, DZKK77, Fog51, Kai89, Kaz65, Kon76, Kra85a, Kra85b, Lin57, LP83, Mak65, Mak80, Peĭ63, PT80, PS93, KR86, Saa91, Sal76b, Tes69, Tre80, Tur72, ZSA86]. Rutishauser [AN68].

S [Fra81, Köl90, Wan80, WOG95]. S-3800 [WOG95]. S/360 [Wan80]. S13 [Gau73]. S14 [PSB76, Pom76, eL79, Dor68, Gau79a, HD73, Hil73a, Hil77, Kuk72a, Kuk72b, LT71, Sch78b, TS69]. S15 [eL76, Ber68, Gau69a, HJ67b, Mac65, OW68]. S16 [Cob69, Gau65a]. S17 [Sko75a, Bur74, Gau64a, Gau65b]. S18 [ADW77a, Amo78]. S20 [LG64]. S21 [Sko78, CN81, Eck80, Far69]. S22 [Cle69, CP70, Gau69b, GZ75, Köl68, Köl69a, Köl69b, Köl69c, Mor76]. Śaṅkara [Hay99]. Saddle [FCC73b]. safe [Bai21]. Sage [ES10]. SAIN [Kee82]. Salzer [Ano84]. Samos [VDR05]. Samples [Pom74, Pom76]. Sampling [Ryb89]. San [Koo91, SE11]. Sanskrit [Plo96]. Santa [ES89, IEE78, MSH^{+}16]. Santander [GST14b]. satisfactory [GST07b]. satisfying [EM94]. saturation [Hei88]. Scalable [OJ04]. Scalar [ACG^{+}86, FFS83, ACG^{+}87]. Scale [TAY91]. scaled [Zag24a]. scaling [ALB98]. SCAN [Ull90]. Scattering [Len76, MS68, Car82, Kre89]. SCELBAL [Arn77]. Scheme [Chl11, GIJL23, Kai89]. Schemes [Pan59b, CH78, Håv79, Pan59a]. schisleniya [LP83]. School [HUY07]. Schranken [Rot70a]. Schwarz [Sti53]. Schwarzschen [Sti53]. Science [IEE89b, LMS99, MC87, MC90, RW04, Tra76, RM07]. Sciences [MS94, PC11, Spr00, Bar14, Mat93]. Scientific [AK93, Lov53, MT85, PTVF07, Pri66, SKL93, AEF92, MT86]. scientifique [AEF92]. Scientists [Bel04, BF71, Lau95, Lau04, MH08, Bel68, Sen67, Köl97]. Scorer [GST02c]. search [Hig93]. Searching [SLZ05]. Secant [BH13, Plo96]. Second [Bra70, Car87, CMF77, DH59, AEF92, FL67, Köl97, LW63a, LW63b, LB90, MC90, Mec66, Olv67a, SW99a, Van69, BL96, BL91, DAV12b, Fet81b, Fuk11, Gla74, Her61e, HvdR63, IKK91, Jam81, Koo91, Lar66, LLK14, Mid75, Mor99, Olv67b, PDK96, RDK98, Tak22, Tem76, VC06b, WC90, Wen96, YM03]. Second-Order [LB90, Olv67a, Olv67b]. sections [Mül99, Rus88]. Segment [Lef05]. select [CL00]. Selected [Cra12b, Pot14, Ban01, TB87]. Selection [HSE^{+}24, Kor05, LM95, Rus13, ALB98, PEB02, PBE03]. selections [Som82]. Self [CC69, GREL96, LKKH99, MIT^{+}95, Ull90, CCCP99, MI97]. Self-contained [CC69]. Self-Timed [MIT^{+}95, LKKH99, CCCP99, MI97]. Self-Validating [Ull90]. semantic [PUHM12]. Semantics [BWKM91]. Semi [LM15a, VM19, Cro92]. Semi-Automatic [LM15a, VM19]. semi-infinite [Cro92]. Semicirculant [HBS00]. semigroups [Fei78]. seminar [Ask75, Laz85]. Seminumerical [Knu98b]. Sense [Dar70]. Seoul [HY14]. Separable [FLMR00]. Separation [Str56, SMCH41]. September [Ano03, ES89, FvdHJT10, IEE89a, IEE22, IEE23, SC01, Wah04]. Sequence [Alb62, Cod83, Tro84, Wim81, Hom98b, Man23, Mus87, Wen89, Wen96]. Sequences [Bra65, Ekl01, FFS83, Osa90, Bre78a, Del81, DGB82, GB90, Gon52b, Lev73, LM08, LSY15, Sed90, Yey92]. Sequential [Kar84]. Series [AGS99, AO08, AS54, BM15, BS98, Cah54, Cle55, CRZ00, Fat16, FFS83, Fuk12, Gas81, GS82b, GCP67, GRJZ07, HBF09, Hom96b, Hum64, Kal04, KC76, Luk59, LC61, PL19, Pie84c, Plo01, Pri75, RS81, Sch16, SJ01, Van69, WR71, Wre68, Wre73, Wyn60, Air37, AJDG02, ABFP24, Ano84, BB15, Bau89, BBD03, BB94, Boe91, BF92, Bor10, BS78, Boy09, Boy14, BCG91, CR08, CGKZ05a, CGKZ05b, CP66, Cra12a, DTLM98, De 91, DT09, Del79, Del84, Dom03, Gus78b, Gus84, Has02, Hom94, Hom98b, Hom99, Kat78, Kra95, Lav86, Lew94, LP95, Lóp98, LT13, McC83, McC81, MC04, MS87, Ng68, Ole96, Par09, Pas95, Pas08, PBN93, Pie84b, PBM86a, PBM86b, PBM92, Sal83, Saw02, Sch76a, Sch77a, SLH06, Sid06, SJ12, Vep08, Ver77]. series [Wen89, Wen96, Wen01, Wen07, Wen10, WN09, Woź10, ZB97]. Service [GS14b, Loz96a, Loz97a]. Session [DM08, Ric92]. Set [DPC95, GZC22, Tem94b, DGB82, GP84, Hay99, Meg89a]. Sets [AEF92, PW76, EH89, Hig77]. Setting [Rah01]. Seventh [MS94]. Several [DX01, DX14, Lóp98, Che13b, CC17, KS88, Liu93]. Sexagesimal [Arc43]. Shanks [AGJ86]. Shared [MIT^{+}95, Fan89, KP98, LKKH99]. Sharp [BD85, CT16, Mor14, QV96, YCZ15, Ziv99, MRS18, SR16]. sharpened [Kai89]. Shift [CH70b]. Shifts [BM04]. Short [ADN03, Boe61, Boe62, Bur63, Bur64, Fet67, Fox54, Fra65a, GB68, Gla66, Har66, HKW60, Hea65, Hum64, JJ65, Lan60, Lon59, Mec68, Moo67, SF59, Tak66, Tho65, Tho66, TM68, Ver67, WC61, Wil68, WKG66, Woo67]. Should [PW76]. Shrivenham [MC87, MC90]. shuffle [SR81]. shuffle-exchange [SR81]. SIAM [SKL93, IP87]. side [JW88]. Sided [Sch84]. Sieve [Ngu98]. SIGAda [ACM87]. SIGForth [Koo91]. sigma [Töl67a, Töl67a]. Sigma-Funktionen [Töl67a]. sigmoid [WR96]. Sign [Kah87, HHR00, Sid06]. signal [JL94, Luk99]. Signals [Sin95, LP62]. signed [CA00, LKKH99, RN96]. signed-digit [CA00]. Significant [BL94, DM88a, DiD90, DM92, ZA24]. SIGSAM [ACM89]. SIMD [GIJL23]. Similar [Tho66, Völ83, Act74]. Simple [BS79, HMO19, Jan11a, Jan11b, MC91, VLCSFN^{+}12, Zag17, von84, CB09, Cho14, Cre91, Dom03, KRVZ98, Leh83b, Lew94, NS13, PNV01, Vai89, Wen64]. Simpler [Lin90, Sha85]. simplex [Blo88]. Simplification [Bro89a, BBD03, BD02]. Simplified [TM68, Wal80]. simply [SE14]. Simulation [MMH89, Toc63, Gly89]. Simultaneous [Gar79, BL96, Bel90, KI04, Sun88]. Sinc [GS91]. since [Gau98]. Sine [BH63, Cun62a, Cun62b, Plo01, SZ76, TAY91, Tan90d, Bak76, BS11, Har00, JL94, PKS00, Saf10, Paw11]. sine/cosine [Bak76, JL94, PKS00]. Sines [Pat18]. Single [Fra65b, HHv^{+}73, LP16, LP17b, SH72, Hon71, LC97a, MI97, WG95]. Single- [HHv^{+}73]. single-rail [MI97]. singular [BS06, DK82, Mus87, Whi82]. Singularities [SL07, Tim87]. singularly [JS89]. sistemakh [LP83]. Sivashinsky [AGH89]. Six [AFC10, WOG95, ZS08]. sixteenth [Hay99]. sixteenth-century [Hay99]. Sixth [SKL93]. size [Bel89, Par99]. Sketch [Lov43, Lov89]. skew [Kru99]. skew-symmetric [Kru99]. skipping [MC91]. SLEEF [SP20]. Slepian [MC04]. Slimmer [dLMF23]. sloping [Ehr89]. Slow [LB09, Wil70]. Slow-Fading [LB09]. Slowly [UE05, Lew94, WN09, Woź10]. Small [Der77, ELMT00, Pom74, Pom76, BZ92, GST15a, Pie84d, Shi93]. Smirnov [Pom76, Pom74]. Smooth [Pol88, KS88, Pro88]. smoothing [BD85]. Smoothness [GK89a]. SMT [KI16]. SMT-based [KI16]. Sobolev [Zie89]. Society [Ban01]. Software [And82, Bee17, BCD^{+}20, Boi97, Cod71, CW80, Cow77, CBBV10, ES10, GS14b, Gre16, HY14, Joh87b, Joh12, Kah80, Loz94, Loz96a, Loz96b, Loz97a, Mar04a, Mil04, Mol83, Rap14, Ric71, Sch77b, Tan90d, WNO94, dDM06, Cod82, Ein79, FvdHJT10, GST14b, ISO10, KRVZ98, LKK19, Mac94b, Mac96, PIH14, Smi01]. solids [Har77]. Sollya [CJL10]. Solution [Die83, Hom96a, LB90, Mil57, Pre55, Rob55, DK77, Lev91a, Luk56a, Luk56b, Olv67b, Rut51, Spi85]. Solutions [Cle69, Fri72, Olv67a, GST04b, GST07b, RdL01]. Solver [Flo15]. solves [SD85]. Solving [Boy14, Fuk83, Gon89, Hor19]. Somastuvan [Hay99]. Some [AB88, Alz93, Bak75b, Bel58, BM80, Bre82b, Car70a, Che11, CB12, DHM89, Elb01, Erc72, ELMT00, Fdi97, FB91, Fil66, GB68, Gla66, Gla74, HP62, Her82, Hit75, HKA93, Ker83, Kin21, Lav86, LM07, Lu14, LSM15, LM15b, Luk75a, McN73, Mon83, Mul24, Osa90, Pas03, PC93, QM15, QVV98, RGK72, Ric68, Ruy89, ST74, Sle65, SW88, Tan90d, Tur53, Wan80, Wim62, XC14, Alz97b, Alz97c, Arm82, Bak73, Bak75a, Bre78a, CC17, De 91, DF78, FW80, Fra81, FW85, Gau11, GM84, GK89b, HP93, Hei88, Hol90, Kui52, Kza92, LW14, MT72, MY91, Mit49, MC91, Now06, OSS09, Pal98, PS86, QGG10, Rza̧12, Sed90, SL83, Sim64, Ter81a, Tre80, Wen01, Wim72, WN09, Woź10, VB81]. Sommerfeld [Fuk14a]. Soni [Coh12, Wig67, Szm13]. Sound [HJ96]. source [BBB97, BBB00, BBB04, Joh12]. sources [Wil96]. South [HY14]. Southard [T.57]. Southeast [Men06]. Space [CGKZ05a, CGKZ05b, LP62, Mus87]. Space-efficient [CGKZ05a, CGKZ05b]. spaced [FK88]. Spaces [Yos97, AB88, Som82, Zie89]. Spanish [MCM90]. SPECFUN [Cod93b]. Special [AO08, AGL93, AK09, And98, AAR99, AKS01, Bal00, Ban01, Bat07, BW10, Bel68, Bel04, BS11, Bre80a, BIS01, Car70b, Car77b, Cer07, CD08, CMW63, Cod75, Cod93b, CVW06, CPV^{+}08, CBBV10, Deu76, DM08, El 06, Fei78, Ful77, GI01, Gau75, GST03b, GST07a, GLR07, HBF09, Hoc61, IK05, IA20, Joh14b, Joh20b, Köl87, Koo91, KS20, Kor11, LM15a, Leb65, Leb72, Loz94, Loz96a, Loz96b, Loz97a, Loz00, Luk69a, Luk69b, MH08, Mil65, Mos72, NU88, Olv74, Rai60, Ric92, SF16, Sch76b, Sch77b, SC01, SL07, Sne56, Sne61, Sne63, Sne80, SZ70, SZ74, SZ76, Szm13, Tal68, Tem96a, Vil88, VK95, WG89, Wig67, WWG82, ZJ96, Ano93, Ano98a, Ars81, Ars68, AN68, Ask75, BC09, BBC^{+}14, Ber01, Bre10, Buc69, Bus74, Car82, CKT07]. special [CDS00, Coh12, DKK80, DIW00, Gau63, GG08, GST02a, GST14b, Hig77, ISO10, KV03, Koe16, LO94a, LO94b, Mac96, MO49, MOS66, Mar88, Mat93, MS03, Mil68, Mos69, PNV01, PBM92, RDK98, Sau93, Sen67, Sid06, Tem83b, Tem07, Töl66, Töl67a, Tri66, Ubh89, WS84, Wen01, Wen07, Wim82, vdLT84, Ban01, Die80, DM08, Her82, PBM86b]. Special-Function [Kor11]. speciali [AN68]. Specialness [Bat07]. Specific [GBKK09, HCK09, IEE08, T^{+}97, VDR05]. Specification [MHSK16]. spectral [Boy89, Boy01, Koe16]. Speculation [CL94b]. Speed [CG85b, JSH^{+}11, Maj85, PB02, POMB05, RGK72, SW99b, SC71, SA57, ZG87, BHJ05, CCCP99, Fan89, LMT^{+}92, Mar00, Sei99]. Speeding [PIH14]. spezielle [Töl66, Töl67a, Sne63]. Spherical [AM78b, AM84b, Bra73, Col78, CU59, GF97, Har00, Hob31, Hob55, Jab94, MS67, O’B80, PS05, BDGP14, Cai11, Del79, Del84, GF88, Len90, Max91, PS77, PS79, Tal83, Tal84, Tal09, Wil71, de 77, Her61a]. Spheroidal [Fla57, LKL02, Low64, Osi13, Sle64, Sle65, Sle78, Str56, V^{+}75, XD18, LP61, LP62, MC04, OWS^{+}14, RX07, SP61, SS65, SMCH41]. SPIE [Luk99]. SpiNNaker [MLS^{+}18]. SpiNNaker-2 [MLS^{+}18]. Spline [Lin79a]. splines [DR89, Sab08]. split [WG92]. Splitting [PLT23, Kre89]. sponsored [Ask75]. Spring [AFI71]. Springer [Fet81b, Köl97]. sqrt [Rah01, Kah96]. Square [ARH14, ADN03, Ahr96, Ale77, Alt79, AMT78, ALMN05, AMS76, BB12, BV85, BO93, Bee91, BMZ02, BR73, Bor21b, BFHT85, BJK^{+}11, Bru22, BH07, BTDS11, Cam86, Car74, Cat86, Che72, Cod64, CG85b, CH98, CHI^{+}03, CL94b, Cow64, Cre91, Cre98, D^{+}89, Die83, Egb77, ES99, EH16, Erc78, EL94, EIM^{+}99, EIM^{+}00, ELMT00, EM04a, EM07, Esc37, Eve63, Fik66, FR98, Fri67, Gar75, Gar79, Gar80, GBKK09, HAK91, HCK09, HSE^{+}24, Has90, Hei96, Hol69, HMO19, HA85, Int00, Int03a, Int03b, IM99a, IM99b, ITY95, ITY96, JL94, JJ65, JKMR11, Joh87a, Joh89, Kah99, Kal34, KM97, Kin65, KP69, KK98, Kor05, KNS95, LM95, LA01, LA03, LF92, Liu95, LE95, LD89, Maj85, MK71, Maj72, Mar04a, MIT^{+}95, MRH19, Met65, MKY92, MM83, MM91b]. Square [Mül04, MCT18, Nan11, NF63, Ngu98, Nin70, OLHA95, OE82, Par06, Pat18, PB02, PB08, PZ95, RGK72, RJM^{+}20, Rix82, Rob69, RHMLL08, Rol87, Rot70a, Rus98, Rus13, SH72, Sar59, SK71, SG02a, Sch84, Sch85, SW99a, SW99b, SF96, SL95, SL97, Swa65, Tak01, Tha61, Tho93, Tor02, Tra60, Tur94, Wad58, WMC21, Wal80, WS05b, Wej83, Wil70, WS06, XLW^{+}20, Yoh70, Yoh73, Yon70, ZA14, Zim99, ZG87, AHM^{+}98, AGS99, Ano12, BH99, Bas01, BMZ06, Bic81, Bor21c, BJM22, Bra84, BMST97, BH01, Bur82, CCC96, CCCP99, CF00, Cat85, Che71, CL94a, CL95, CL00, CM90, CO86, CHGM99, CP99, CLZ04, DD76, DSK83, DSK84, DH77, Dut86, EL89b, EL89a, EL89c, EL90, EM03, EM04b, Fan89, Fre81, GST15b, GREL96]. square [Har03, HHR00, Has00a, Has00b, HR05, HR07, Hig93, Hit74, HF95, Hon71, HT73, HFS99, ITY97, Jam89, Jam99, JM17, Joh87b, KB89, KK96, KM93, KMB94, KCYL06, KP98, LM92, LM99, LKH93, LKKH99, LO95, LC96, LC97a, LC97b, LE93, LMT^{+}92, Lyo91, Man23, MST89, MI97, Mat90b, MMW91, MM91a, MM92, MMH93, Mit92, MC89, MM90b, MC91, MC93, MC95, MWH^{+}16, MWH^{+}18, Mur99, NL99, Obe99, OJ04, Özb06, Par99, Ped80, Pen81, Per85a, PS86, PR04, Pro83, QW06, RS98, Rus99, Rut63, Sar60, Saw02, Sch73, SF93a, SF93b, Sch95, Sin96, SD85, SLD16, Ste89, Ste08, SL63, Str59, Tak00, Tay81, TVG00, TE06, Tom00, Tuk48, VCV01, WMC18, WN03, WG92, WMDM92, Yey92, ZXBZH01, ZLC04, Zim00, Kun81]. Square-Root [ALMN05, CL94b, Gar80, LA01, RJM^{+}20, Wad58, Yon70, JL94, LA03, ALB98, GREL96, Has00a, Has00b, HF95, HFS99, KP98, LKKH99, Mit92, PR04, TVG00, Tuk48, ZLC04]. Square-Rooting [Gar79, Maj85, RGK72, KK98, Liu95, Par99, WG92]. square-roots [BH99]. Squared [HP67, HP85, eL76, OW68]. Squarers [WS05a]. Squares [DB13, Maj72, Sto41, Rei86, SL63]. Squaring [Kar84, Sar60]. squeeze [Mar77]. SRT [MC95, BH01, HF95, Kor05, MMH93, MC93, Rus13].

Stability [Gar79, Mil75b, Gus84, Spi85, YH89]. Stable [Gus85, KR79, Oli75, Fet74, VC06b]. stacks [Dem89]. stage [Töl43]. Stages [PZ95]. Staggered [BHK09]. Standard [BBM84, Dri89, GB91, Lóp00, Mor75, Nor89, Rum01, Sal92, Sha85, SP20, YBR11, Zwi03, AHM^{+}98, Cho14, DAV12a, DAV12b, Dri91d, Dri91a, Dri91b, Dri91c, Düm10, Hou81, Krä88, McC90, Neh07, Squ91a, Squ91d, Squ91b, Tem85, WR96, Kah80]. Standardfunktionen [Bra87, Krä87, Mül90, Rot70b, von80, von84]. Standardization [CHG^{+}11, DHL^{+}04, Kul07, Squ91c]. Standards [Lid01, Lid02, Mul82, Low49]. Start [Mar05]. Starting [Eve63, Fik66, MM91b, Nin70, SF96, Wil70, Mou67, SF93a, Tem96b]. State [IP87, MRH19, Dem89, Roc82]. State-of-the-Art [MRH19]. States [IEE23]. static [MI97]. Stationery [L.61]. Stations [WNO94]. Statistic [Kra89, Pom74, Pom76]. Statistical [Hil73b, Mac89, TGC89, Mat93, Sti88]. Statistics [Gup60, Goe14, Gom89, Mar88]. STATLIB [TGC89]. status [Ste08]. Steed [Bar82b, Bar84a]. Step [Bor21b, Spi85]. steps [MC91]. Stepsize [Ske86a, Ske86b, Spi85]. Stieltjes [BB82, BHJ05, BS78, Cha80, KC11]. Stirling [Sch16, Spi71, XC14]. STO [LDP93]. Stochastic [PL19, Sob88, BW89, Ebe89, Gly89]. Stokes [HBS00]. Stoner [Fuk14b]. Störungsreihe [Hom96b]. straight [Rah01]. Strange [GS82b, Wen09]. Strategies [BBC98, BBC00]. strategy [WO15]. stratified [Kre89]. strength [Ost37]. stress [SC81]. strictly [Whi82]. Strong [Ebe89]. Strongly [Sal92]. Structural [Lew91]. Structure [CET62, Jef17, BB16, EC79, Har77, WO15]. Structures [FS96]. Struve [AJ03, Abr64b, BDGP14, New84, Olv67a]. Student [Hil81a, Hil81b, eL79, Hil70a, Hil70b, Hil73a]. Studies [Bir54, Har77]. Studium [Bar14]. Studv [WB78]. Study [OLHA95, Sun71, AB93, Bar14, Bre78c, EH16, GK70a, Gau11, Per85a, SK08]. Subject [Ano68]. Sublinear [BW89]. submultiplicative [AR99]. Subprograms [Kuk71]. subquadratic [Har14]. Subroutine [Amo83a, Amo83b, Bre82a, FHS77, Kod08, Col80, Col84, Tal83, Tal84, Tal09]. Subroutines [ADW77a, ADW77b, Amo78, Cod75, DM87, GN89, Mor75, Mor81, Mor90, Mor93, TGC89, Wan80, WWG82, Alh00a, EC13, Joh88, Mor84, Sca71, Die83]. subsequences [BDGB83]. subsets [Ubh89]. Subspace [Ovt03]. Subsystem [IBM05]. subtraction [TW80]. Subtractions [Sug58]. Subtractive [SL95, Ste89, LO95]. such [BO75]. Suggestion [Bak76]. suitable [MT72]. Suite [Liu87, Tan90c, Tan91a]. suites [Del81]. sum [BDGP13, Bus74, SL63, Vel11]. summability [Mül93]. Summation [Cle55, Deu76, Gas81, Sid06, Ver67, Wen10, Hom99, KK98, Lew94, Liu95, Ng68, Ver77, Wen89, Wen01, WN09, Woź10]. Summations [Gla81]. Summer [HUY07]. summing [Wen96]. Sums [Dub83, Lir71, MM83, Sch16, UE05, BB16, BG84, Büh03, CG89, Fre05, Har77, Kra95, QW06, Sti53, ZLA13]. super [Mag94]. Supercomputer [WOG95]. Supercomputing [Ano91, IEE91]. Superscalar [ECSD20].

Supplement [Arn77, Ske86b]. support [ISO10]. Supremum [CJL09]. sure [Wil96]. Surface [DR94a, DR94b]. surfaces [Lit83]. Survey [BTT16, Bre78b, Fra65b, Gau75, MM90b, Ars81, Lew44]. Sussex [IEE89a]. Switzerland [Lak96, T^{+}97]. swMATH [GS14b]. Symbolic [ACM89, BDGP13, Bro07, Car99, GI01, Joh20b, Lak96, SE11, Sko04, Zha96b, Ye17]. Symbols [JF16, Gun65b, Gun65c, MS82, Wei98]. Symmetric [CG94, JGB13, Lóp00, Ovt03, SS97b, SS97a, SS99a, SS99b, ZC70, Fuk12, Kru99]. Symmetrical [Sau92]. Symmetries [Car70b]. Symposium [ACM89, AGL93, Ano91, Atk79, Bow53a, Bro07, BCDH09, BC01, BBdD17, CLT20, ES89, Gau94, Hwa85, IEE78, IEE81a, IEE81b, IEE89b, IEE05, IEE13, IEE21, IEE22, IEE23, KM95, KK99, KM91, KM07, Lak96, LMT97, Lan59, MT85, MSH^{+}16, MTV15, SE11, SO11, SC01, SIJ93, TBL19, TT18, Tra76, MT86]. Synthesis [BLP24, BT99, EH16]. Synthesisable [BH07]. Synthesized [AAHTH10]. System [ACG^{+}87, Kor11, Kuk71, MS94, OTY91, Rot70b, Spr00, WB78, Fab02, Her82, ISO10, Ise94, Kaz65, KR86, ACG^{+}86, Fik66, Mar90]. system-theoretic [Her82]. System/360 [WB78, Fik66]. System/370 [ACG^{+}87, ACG^{+}86]. System/6000 [Mar90]. Systems [Bra65, Fri72, IEE08, Mey01, NKY08, QN18, Sin95, T^{+}97, VDR05, Dem89, Gly89, LP83, Roc82, Bra87, SM70]. systolic [EM86]. Szasz [Sun88]. Szego [Rus88].

T [MCT18]. T-count [MCT18]. T. [T.57]. tabellierten [Sti53]. Table [Bea58, BEM12, Car87, Car88, Car89, Car92, Coh12, Fet81b, Fet81a, Fox54, Ful81b, GRJZ07, Hea65, Köl97, Lon59, Mac94b, PLT23, Sal51b, Sal53, SS97b, Sha76, Sou57, SW50, SS99b, Tan89a, Tan90e, Tan91b, Tan92, TM68, WMBE14, WD61, Car91, DC81, KPPS08, Par99, T.57]. Table-Based [BEM12, WMBE14, Mac94b, PLT23]. Table-Driven [Tan89a, Tan90e, Tan92]. Table-Lookup [Tan91b]. Tables [AS64, AS10, Ano48, Ano74, Arc43, BA44, Ben83, BC62, BL45, DK77, Emd40, Emd45, Emd48, Emd59, FC64, Fox60, GBS19, Ken14, Ken21, L.61, LFB60, LJ13, Luk71a, Luk71b, Luk72, Mit49, Nis84, Pai19, Rus13, dLSGDR17, SS97a, SS99a, SZ05, Str56, Uni49, V^{+}75, War60b, Zwi03, Bar72, CH78, Lut95, Sal76b, SMCH41, L.61, Töl69, Töl70]. Tables-and-Additions [LJ13]. tablitsy [Bar72]. tabulated [Sti53]. Tafeln [Töl69, Töl70, Emd40, Emd45, Emd48, Emd59]. Tail [GSS12, GW91, Haw82, Lin90, Bry02, Düm10, Fet74, Lin89]. tail-end [Fet74]. tailored [CR08]. taking [BH99, Str59]. Tampa [Ric92]. Tangent [BH13, KB67, LP17a, Mar05, Pre78, Yun09]. Tangents [Pre55]. tanpi [Kah18]. Taschenbuch [Bar14]. Taschenrechner [Hen77]. Tau [ZB95, Zha95, Rap01, Zha96a, Rap94]. Tau-method [ZB95, Zha95, Zha96a, Rap94]. Taylor [Lip21, Fat16, Mül99, Plo01, SG02a, Sch76a, Sch77a, WV20, Ye17, Lov53]. TC2 [Boi97]. TC2/WG2.5 [Boi97]. Tchebycheff [Sti59, Tho65]. teaching [BL71]. Technical [AGA^{+}80, Moo67, PT80, Boe61, Boe62, Bur63, Bur64, Fet67, Fra65a, GB68, Gla66, Har66, HKW60, Hea65, Hum64, JJ65, Lan60, Lon59, Mec68, SF59, Tak66, Tho65, Tho66, TM68, Ver67, WC61, Wil68, WKG66, Woo67]. Technique [Rob55, Vol59a, Vol59b, Dan77, Har77, KMB94]. Techniques [CH70a, GT59, Mec68, Mul82, SH72, TBDS11, Cro92, Hon71, RN96, dDM06]. Technology [IEE99, Lid01, Lid02, LMS99, ISO10]. tekhnicheskikh [PT80]. Teljakowskii [Leh83a]. Temme [Köl87, DiD84]. teooria [Jür66]. Term [Gau67, BJ15, EG97, HS20, Kra14]. Terms [Tho65, UE05, AS97, BJLP19, Lóp18, MV98, Mor14, QM15, Sko05, Tem81]. Tesler [Köl90]. Test [CHG^{+}11, Cod91, Cod93a, Cod93b, Kah96, LL09, Liu87, Loz96a, Loz97a, Tan90c, Cod82, Gom89, KI16, Sid06, Tan91a]. test-case [KI16]. Testing [DG99, Sch77b, Tan89b, Tan90a, Wil96, Kul07, LMS73, Squ91c, VCV01]. Tests [Par06, Mac94b]. tetragamma [Che11]. Texas [IEE85, IEE13, Koo91, Koo91]. Text [Plo96, BW10]. th [CL11, CH89a, CH89b, Har14, Kog59, Rei69]. theatres [Sho86]. Their [AMM21, Amo73, AVV90, Dan82, DM86, Ike75, Ike76, Kor02, Leb65, Luk69a, Luk69b, Luk75b, Sam02, Tan91b, Wim81, Amo74, ABFP24, Ban01, Bar82b, Bar82a, Bar84a, Coc65, DM87, Din73, DC81, Dut86, FLO04, FKY86, GG08, Gau11, GS14a, ISO10, Jam89, Leb72, McN73, Mei64, MKS83, Neh07, PB82, RM07, Rit25, Sal51b, Sau93, SC01, SO89, Sun88, Tem79a, Wac00, Wei98]. Theorem [Bor02, Ryb89, SG02a, Bor82, Dav35, EC79, Leh83a, Leh83b, Rus88, CK88]. Theorems [Szm13, Wig67, CF00, Coh12, GKK00, MO49, MOS66]. Theoretic [Ike76, Tal68, Her82, VB81]. Theoretical [AKS01, Sun71, Wen09]. Theory [Ach56, Ach92, AH16, BCJV18, BW00, BR91, Che70, GI01, Hob31, Hob55, IEE89a, IK05, Jür66, LKL02, Mei67, Mos72, Sid03, SL07, Squ70, Tan89b, Var95, Vla02, Wal00, Wat66, Wat95, AAS93, BG84, Car82, CT88, Fog51, Håv81, KS04, Koe16, vdD88, Mil68, Per85a, Ric88, Töl43, Töl50, Töl66, Töl67b, Töl67a, Töl68, Töl69, Töl70, Vil88, WO15, Ask75]. Theta [MT64b, Töl66, Töl68, Töl69, Töl70, BG84, FGG82, FNC75, Kuz15, Ye17, Töl66]. Theta-Funktionen [Töl66, Töl68, Töl69, Töl70]. Thiele [Mak80]. Third [Boe61, Car88, Fet65, Fra65a, GST02d, Tem75a, War60a, ZC70, Bul69a, FvdHJT10, GST03a, Koo91, Lev91a, Bor02]. third-order [Lev91a]. those [GS14a, Tem79a]. Thought [Bow53a]. thread [Bai21]. thread-safe [Bai21]. Three [Gau67, Per85a, BK16, DR89, HS20, PS79]. Three-Term [Gau67, HS20]. threshold [Lit89]. Throughput [AFC10, CCC96]. Tight [MDM16, TBDS11, BMST97, Wil96]. Timan [Leh83b]. Time [KR79, BY19, Bre18, Ebe89, Kre89, LP62, MC93, MC95, Mul06]. time- [LP62]. Timed [MIT^{+}95, CCCP99, GREL96, LKKH99, MI97]. tiny [SG99]. together [Rot70a]. Tokyo [Lav80]. Too [Kah04]. tool [PI16, VCV01]. toolbox [RKZ^{+}14]. Tools [Osi13]. Topics [Mil75a]. Torino [Ano98b]. Tornheim [BB16]. Toronto [Hea65]. total [JMS98]. Totally [HR88]. Tour [Sri07]. Traces [Tem83b]. Tractable [IR08, LBC11]. Trade [Kuk71]. Trade-Off [Kuk71]. tradeoff [CL00]. Tradeoffs [WN03]. Trans [Joh87b]. Transcendence [Mas75]. Transcendental [FB91, Fil66, HKST99, Lir71, LASC95, Nav83, ST99, VVA09, Zuc64, Boy14, Car69, Eps75, GS88, HXX22, Ng68, PS77, PS78, PS79, Töl43, Töl50]. Transcendentals [Har09a, LMT98]. transfer [GK89a]. transform [AGJ86, Ell83, GM84, Sch96, Wie99]. Transformation [Fet65, Bul69a, Fuk10, Hom98b, SS10, Tem85, RDK98]. Transformations [Ng77, Sal89, Sch16, Wim81, ABFP24, Jam81, Lev73, KR86, Sko04, Wen89, Wen96]. Transforms [Olv91, Ano93, BB82, CR08, CZ95, CY18, FW85, Gus85, KW21, KS04, Mac72, PB82, Tal83, Tal84, Tal09]. transition [ND19]. Translation [Bai93, Zag16]. transmission [Boe91]. transmission-line [Boe91]. Transmutation [Car82]. Transportable [Cod82]. transzendente [Töl43, Töl50]. transzendenter [Fil66]. Trapezoidal [Ric73]. Treatise [Wat66, Wat95, MS67, Sen67]. Treatment [Frö55, Che13b, CC17, GLP98, GM07, Hit75, Mei64]. Triangle [IEE89b]. triangular [Kru99]. Tricks [Bli97]. Tricomi [Ano98b, Gau98, MMV81, Tem83b, Wim74]. Trig [Sal92, Tho93]. Trigamma [Pai19, Spo94, QM15, Ron86]. Trigonometric [AH16, Arc43, BEM12, Bur64, Fow93, GZC22, LaF54, Lew69, QN18, Qua18, Rob55, dLSGDR17, Sal51b, SR53, Smi95, Vol59b, WMBE14, AIS^{+}17, API^{+}19, Boo55, Kan96, LK73, Sch80b, dDIS13]. Trinomials [RHMLL08]. triple [CS11]. Truncated [WS05a, Fuk14a, Kuz15]. truncation [Bor10, McC74, Wen07]. Tschebyscheff [Hol69]. Tschebyscheff-Approximationen [Hol69]. tsifra [LP83]. tsifroĭ [LP83]. Tübingen [SO11]. Tucson [DM08]. Turin [AGL93]. Turing [Tru11]. Turkey [HUY07]. turning [PA86]. Tusnady [Ein89]. tutorial [Glo89]. Twenty [MS94, Sin95]. Twenty-Ninth [Sin95]. Twenty-Seventh [MS94]. Two [AS05, BK78, Ben83, Boe61, Boe62, Car92, CL61, CVV65, FW13, Has66, HMO19, MG89, MPG92a, MPG92b, MS68, Sch84, SG99, Sol69, Sor94, Sue99, Wad58, WEX14, Wre68, Wre73, ABFP24, CF00, Cro92, Das94, DC81, GST02b, GST02c, Gla74, GM88, Har20, LW96, Max91, Ost54, PS77, PS78, Pól49, QW06, RKZ^{+}14, Yey92, Gus78a]. Two-Dimensional [WEX14]. Two-point [MG89, MPG92a, MPG92b, Cro92, GM88]. Two-Sided [Sch84]. Type [BJG^{+}99, CCM11, FK11, Gel51, CG89, Chi20, CPG12, Fdi97, GST03a, GST20, GLP98, Håv79, Has02, Hau88, KT00, KRT02, Lev91a, LW82, ÖE16, Pas92, Tem85, Vai89, ZZ96, ZS08, Zhu10, KSVW07]. Typed [Sal92]. types [EMR82a].

UK [BBdD17]. ULP [Ziv99]. ultimate [dDEG05]. Ultra [EMR82a, EMR82b, KO94]. Ultra-arithmetic [EMR82a, EMR82b]. Ultra-high [KO94]. Uncertainty [Sle64, Sle78, LP61, LP62, SP61]. uncommon [Mac96]. Undecidability [Wan74]. Undecidable [Ric68]. under/overflow [LM07]. Unified [AFC10, AS10, AM78a, Che13b, CC17, Cra12a, GLP98, GM07, HSE^{+}24, NKY08, SA00, SL07, Wan16, CL94a, CL95, Håv81, HTHR94, Hit75, NU88, Wal71]. Uniform [Boe91, DGS18, MR71, Olv91, Tem75b, Tem78, Tem83c, Tem96b, Bry02, GST03a, GST20, Luk75a, OG82, Par02a, Par02b, Par16, Rei86, Tem85, Won73b]. Unifying [ES10, Par06]. Unit [Bru22, BS98, Min70, Nan11, ZG87, ALB98, BBC^{+}91, Büh87, Büh92, Büh03, BH01, Che71, DR04, KMB94, LM99, LKH93, Lew85, Lew87, MI97, MS82, TRHH94]. Unitary [Leh96]. United [Boi97, IEE89a, IEE23]. Units [WS06, OJ04]. univalence [KT59]. Univalent [SO89, Rus88]. Univariate [Div79, IA20, JKK92, MDM16]. universal [HSW89, Rix82]. Universeller [Rix82]. University [AGA^{+}80, Ask75, Bro07, HUY07, Hwa85, IEE89a, IEE81a, IEE81b, IP87, Tra76]. unlimited [HTHR94]. Unrestricted [Bre80a, Bre10, CO86, Fat16, Olv80, CO80]. unstable [AGH89]. Unterstufe [Töl43]. Untypical [Pas10, Pas11]. unvollstaendigen [Mah30]. unvollständige [Sch80a]. unvollständigen [Tri50]. Unwinding [Hig18, JHC96]. Upper [ISK87, Jan11a, Jan11b, SS98, BW89, Cha82, Fet79, Fre09, LW96, LU96]. Urbana [Hwa85]. USA [CLT20, IEE13, MSH^{+}16, BCDH09, ES89, IEE05, LMT97, SE11, SKL93, TT18]. Usage [Goe14, PI16]. Use [DB13, GBS19, Pag77, WWG82, Epp89, Lin89, Mim88, Mit92, MWH^{+}16, MWH^{+}18, Phi79, Tri66, Wil71]. useful [Vep08]. User [???90, Ful81b, Ful81a]. Users [Pri66]. Using [ACM87, Bee17, Der77, DG99, Ekl01, ELMT00, GST03b, GM04, HFT94, JF16, Kog57, Kog58b, Kog58a, Kog59, Len76, Mar04a, MIT^{+}95, MKY92, OTY91, PL19, POMB05, PZ95, PLT23, Pri75, RGK72, RS81, Sal76a, Sal92, Sar59, SG02a, ST07, SF93b, SLZ05, SDP11, Tra60, VVA09, WS05a, WS05b, WEX14, WG94b, WOG95, dDLM11, von84, AAHTH10, Bai88, Ben98, BMZ06, BZ92, Che81, CKL89, CA00, CP99, Dem89, DiD84, EM86, FLO04, GST02a, Gra02, Gus85, HBS00, Hon71, HFS99, KMB94, KK98, LKKH99, Liu95, MMM93, McC77, PNV01, SH72, Saw02, Spe72, WG92, Yun09]. UX [LMOT01].

v [LP83, BKdLSGL23, GZC22, Sch73, Sle78]. V. [Kon76, Rac82]. Vāriyar [Hay99]. VA [SKL93]. Vacuum [FR76]. Vacuum-Polarization [FR76]. Vail [BC01]. valid [RX07]. Validated [BBC^{+}14, CCV11, CBBV10]. Validating [Ull90]. Validation [AAS93, CK89b, Squ91b]. Value [AO08, Cas83, Daw97, Fri67, Cro92, JS89, Mat90a, Spi85]. Valued [Mat04, PS11, Fab02, Wat90]. Values [Ano60a, BY07, Cle69, DJ67, Fri72, GB68, GRAST16, HP62, HKW60, MM91b, SF16, TKK17, BS11, De 91, EC13, Fet81a, FW80, Fra81, Fuk09b, Gau59, GST15a, GOST23, Jab20, KWY07a, KWY07b, Kat78, Mou67, Pan59a, Pré10, RX07, Tem87, Tem96b]. Vancouver [Gau94]. Variable [CET62, Fra65b, HA85, HA86, Let01, LE95, PB08, Ric68, Sal51a, Ske86a, Ske86b, SLZ05, Won73a, XL94, Gon52a, Gon52b, HXX22, Jür66, MY91, Ric88].

Variable-Precision [LE95]. Variable-Stepsize [Ske86a, Ske86b]. Variables [Boe62, DX01, DX14, Sue99, ABFP24, BK16, KPPS08, KS88, MH72, Sho86]. variants [Cra12a]. variates [Mar77]. variation [GK89b, Zie89]. varies [AB93]. various [BBP97, Ism77, Joh12]. Varna [RDK98]. varying [Was89]. Vector [ACG^{+}86, BKdLSGL23, Mol83, NKY08, ACG^{+}87, BCK^{+}93, JS89, KK98, Liu95, Mat92, Rap01, TRHH94, Gun65a]. vector-arithmetic [TRHH94]. vectorisable [PIH14]. vectorizable [dLSG18]. Vectorized [BKdLSGL23, SP20]. Vectors [GLT^{+}15]. Vegas [AFI69]. Veltkamp [Bol06]. Veltkamp/Dekker [Bol06]. venue [IEE89a]. Verfahren [Hol69, Sch68, Die83, Ost37]. Verification [AK93, GBKK09, Krä93, LO95, MRH19, Rus13, SG02a, Har03, Saw02]. Verified [BDL09, DLM09, Miy19]. Verifying [AAHTH10]. Versatile [LBC11]. Version [MHSK16, SSG^{+}18]. versions [LW89]. versus [Ful99, GBS19]. Very [BHK09, LM95, LM99, Mit92, Sch80b, SE14, ALB98, MN16]. Very-High [LM95]. VFP11 [BH07]. Via [Gar79, Bor10, BBC08a, Cau22, CT88, GSS12, GRS87, Gly89, Har09a, Hig93, JKMR11, KW21, KT00, Kuz15, LKF10, NP18b, Par09, Sal89, WLK11]. Viable [ES10]. View [DGS65, Mey01, Ano67]. vignettes [War08]. Virtex [WN03]. virtual [IEE21, IEE22]. virtuoso [BCR^{+}07]. viscous [Joh83]. Vistas [CKT07]. VLSI [BV85, IEE89c, LC96, LD89, MMW91, MM91a, MM92, MMH93, Ste89]. Voigt [Zag16, Rei69, SK08, ZA11]. Vol [CK89a, Laz85, Lov53, MS94]. volatilities [SR16]. volatility [MRS17]. Vols [BC62]. voltage [HFS99]. Volume [IK05, IA20, KS20, L.61]. Vychislenie [LČJ63, LP83, Saa91].

W [Coh12, Szm13]. W. [Ker83]. Wallis [Chi20]. Wallis-type [Chi20]. Walter [Zah94]. Washington [MMH89]. Waterloo [Bro07]. Wave [Abr64a, Boe69, CP70, Fla57, Frö55, Gau66b, Gun67, Ike75, Ike76, LKL02, Low64, Osi13, She76, Sle64, Sle65, Sle78, Str56, SG72, V^{+}75, Vos73, BFSG74, BFSG84, Gau69b, Hum85, Köl69a, Köl69b, Köl69c, Kre89, LP61, LP62, MC04, OWS^{+}14, RX07, SP61, SS65, SMCH41]. Wavefunction [Bar76a, Bar76b, Bar84c]. Wavefunctions [Köl72d, She74, Bar82c]. waves [Ehr89, MKS83]. ways [Boy15]. weak [Wil96]. Weakly [Zie89]. Weber [L.61, Fox60, GST11b, Olv67a, SGA81]. Weierstrass [T.57, Eck76, Eck77, Eck80, KI04, Lin78, Lin79b, PC93, Sou57, Sou64, Tod90, Töl66, Töl67a, Töl68, ZZ96]. Weierstrass-type [ZZ96]. Weierstrasssche [Töl66, Töl67a, Töl68]. Weights [HT13, JM18]. well [Vai89]. well-behaved [Vai89]. WG2.5 [Boi97]. Where [Sau92]. Which [GBS19, Wyn60]. Whittaker [Lau73, LVV05, Wen96]. Wide [BHK09, MMNP91, MV98]. wide-range [MV98]. Wiener [Deh89]. Wigner [JF16, Tal68]. Wilhelm [Wig67]. Windschitl [LSM14, LSM16]. Windsor [SIJ93]. Winter [MMH89]. Wisconsin [Ask75]. Wisconsin-Madison [Ask75]. Without [BC23, Kah99, Sar59, Tra60, EL89c, EL90]. Witten [BB16]. WKB [FK19].

Work [WNO94, Fer07, GG08, Rut51, Tem83b]. Working [Boi97, Tod54, Tod55]. Works [Cra12b, CK89a]. Workshop [Bor02, Cow77, CDS00, DIW00, RDK98]. workshops [Koo91]. Worst [LM03, LSZ08, PS05, SLZ05, Kra98]. Wrigge [Fra81]. Wright [Fet81a, MM95]. Writing [DM91].

X [LDP93, Bor21a, HJ67b, LL09]. X8 [Rot70b]. XMeter [AMM21]. XXI [Hor19].

Y-MP [KB90]. YBC [FR98]. yield [Mei83]. York [Fet81b, Köl97].

Z [Gun65a]. Z80 [Kun81, Kun81]. za [LP83]. ZEBEC [KRVZ98]. Zero [BS98, ES99, Hit57, Osi13, Pie84c, BF92, Bre75b, Bre76b, Cha82, Gla74, IS88, MI97, MC93, MC95, SS65, Tes69]. Zero-Balanced [BS98]. zero-finding [Bre75b, Bre76b]. zero-overhead [MI97]. Zeros [Abr45, Ano46, Bal00, Bea58, BK78, Ben83, CH70a, CS82, Dor66, FO99, GKS04, Ike75, Ike76, Kol70, Kol72b, O’S16, Tem79a, Wan74, Wil68, AB93, Ben98, BPD81, Coc65, CK79, DC81, EL01, Elb01, EL08, FLO04, FS92, FCC73a, FK75, FO93, GG08, GS88, GS12, GS14a, GL83, GR92, HB75, ISK87, IS90, IKK91, IS92, IKF91, KRVZ98, LM93, LW96, Leh70, LW95, Let96, LG79, LU96, LM08, Mah30, Mul95, Par72, PS86, PSS03, Pie84b, Sal55, SG99, Seg08, SL83, SZW11, ZZ11, Ano60a]. zeroth [Wal84]. Zeta [BM15, BH18, Bor95b, BBC98, BBC00, HP62, Kol70, Lin78, Lin79b, Mar65, Rza̧12, Sch16, SJ01, SJ12, TKK17, Töl67a, AB93, BB15, BS06, Bor95a, CHT71, Cra12a, Joh15b, Leh70, Leh56, PDK96, Pré10, Sko05, Tur43, Tur53, Vep08, HG64]. Zeta- [Töl67a]. zeta-function [Leh70, Leh56, Tur43, Tur53]. zur [Die83, Hol69, Rut51, Rut63, Sti53, Hom96b, Sti53]. Zurich [Lak96, T^{+}97].

References

Cray:1990:PCU

[???90] ????, editor. Proceedings Cray User Group. ????, ????, Spring 1990. ISBN ???? LCCN ????

Aaboe:1954:AKI

[Aab54] Asger Aaboe. Al-Kāshǐ’s iteration method for the determination of sin 1^{∘}. Scripta Mathematica, 20(??):24–29, ???? 1954. ISSN 0036-9713.

Akbarpour:2010:VSI

[AAHTH10] Behzad Akbarpour, Amr T. Abdel-Hamid, Sofiène Tahar, and John Harrison. Verifying a synthesized implementation of IEEE-754 floating-point exponential function using HOL. The Computer Journal, 53(4):465–488, May 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/cgi/content/abstract/53/4/
465;   http://comjnl.oxfordjournals.org/cgi/reprint/
53/4/465.

Andrews:1999:SF

[AAR99] George E. Andrews, Richard Askey, and Ranjan Roy. Special Functions, volume 71 of Encyclopedia of mathematics and its applications. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-62321-9 (hardcover), 0-521-78988-5 (paperback), 1-107-32593-5 (e-book). xvi + 664 pp. LCCN QA351 .A74 1999. US$90.00 (hardcover), US$34.95 (paperback).

Albrecht:1993:VNT

[AAS93] R. Albrecht, G. Alefeld, and H. J. Stetter, editors. Validation numerics: theory and applications, volume 9 of Computing. Supplementum. Springer, Wien / New York, 1993. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Alonso:1988:SCN

[AB88] Javier Alonso and Carlos Benitez. Some characteristic and noncharacteristic properties of inner product spaces. Journal of Approximation Theory, 55(3):318–325, December 1988. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Arenstorf:1993:SMZ

[AB93] R. F. Arenstorf and L. L. Brewer. A study of the motion of zeros of the Epstein zeta function associated to m^{2} + y^{2}n^{2} as y varies from 1 to \sqrt{6}. Computers and Mathematics with Applications, 26(5):57–69, September 1993. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0898122193900746.

Aberth:1988:PNA

[Abe88] Oliver Aberth. Precise Numerical Analysis. William C. Brown Company Publishers, Dubuque, IA, USA, 1988. ISBN 0-697-06760-2. x + 225 pp. LCCN QA297 .A28 1988.

Abel:2016:HOA

[Abe16] Ulrich Abel. High order algorithms for calculating roots. The Mathematical Gazette, 100(549):420–428, November 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
2FACD442DB78364B04DA2E64BA06F269.

Ananthanarayan:2023:EAD

[ABF^{+}23] B. Ananthanarayan, Souvik Bera, S. Friot, O. Marichev, and Tanay Pathak. On the evaluation of the Appell F_{2} double hypergeometric function. Computer Physics Communications, 284(??):Article 108589, March 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465522003083.

Ananthanarayan:2024:OWR

[ABFP24] B. Ananthanarayan, Souvik Bera, S. Friot, and Tanay Pathak. Olsson.wl and ROC2.wl: Mathematica packages for transformations of multivariable hypergeometric functions and regions of convergence for their series representations in the two variables case. Computer Physics Communications, 300(??):??, July 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465524000857.

Arzelier:2019:EAE

[ABJ19] Denis Arzelier, Florent Bréhard, and Mioara Joldes. Exchange algorithm for evaluation and approximation error-optimized polynomials. In Takagi et al. [TBL19], pages 30–37. ISBN 1-72813-366-1. ISSN 1063-6889.

Abramowitz:1945:ZCB

[Abr45] Milton Abramowitz. Zeros of certain Bessel functions of fractional order. Mathematical Tables and Other Aids to Computation, 1(9):353–354, January 1945. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Abramowitz:1964:CWF

[Abr64a] Milton Abramowitz. Coulomb wave functions. In Abramowitz and Stegun [AS64], pages 537–554. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://

www.convertit.com/Go/ConvertIt/Reference/AMS55.ASP in

bitmap image format.

Abramowitz:1964:SFR

[Abr64b] Milton Abramowitz. Struve functions and related functions. In Abramowitz and Stegun [AS64], pages 495–502. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also avail-

able online at http://www.convertit.com/Go/ConvertIt/

Reference/AMS55.ASP in bitmap image format.

Agarwal:1986:NSV

[ACG^{+}86] Ramesh C. Agarwal, James W. Cooley, Fred G. Gustavson, James B. Shearer, Gordon Slishman, and Bryant Tuckerman. New scalar and vector elementary functions for the IBM System/ 370. IBM Journal of Research and Development, 30(2):126–144, March 1986. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). See clarification [ACG^{+}87].

Agarwal:1987:CNS

[ACG^{+}87] Ramesh C. Agarwal, James W. Cooley, Fred G. Gustavson, James B. Shearer, Gordon Slishman, and Bryant Tuckerman. Clarification: “New scalar and vector elementary functions for the IBM System/370” [IBM J. Res. Develop. 30 (1986), no. 2, 126–144]. IBM Journal of Research and Development, 31(2): 274, March 1987. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). See [ACG^{+}86].

Achieser:1956:TA

[Ach56] N. I. Achieser. Theory of Approximation. Frederick Ungar Publishing Company, New York, NY, USA, 1956. x + 307 pp. LCCN QA221 .A533 1956.

Achieser:1992:TA

[Ach92] N. I. Achieser. Theory of Approximation. Dover books on advanced mathematics. Dover, New York, NY, USA, 1992. ISBN 0-486-67129-1 (paperback). x + 307 pp. LCCN QA221 .A533

ACM:1987:UAA

[ACM87] ACM, editor. Using Ada: ACM SIGAda international conference, Boston, Massachusetts, December 8–11, 1987. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987.

ACM:1989:PAI

[ACM89] ACM, editor. Proceedings of the ACM–SIGSAM 1989 International Symposium on Symbolic and Algebraic Computation, IS-SAC ’89. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

Acton:1974:RRF

[Act74] Forman S. Acton. Recurrence relations for the Fresnel integral ∫_{0}^{∞} exp(−ct) dt/\sqrt{t(1 + t^{2})} and similar integrals. Communications of the ACM, 17(8):480–481, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Adamchik:1998:PFN

[Ada98] Victor S. Adamchik. Polygamma functions of negative order. Journal of Computational and Applied Mathematics, 100 (2):191–199, December 21, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042798001927.

Adlaj:2012:EFP

[Adl12] Semjon Adlaj. An eloquent formula for the perimeter of an ellipse. Notices of the American Mathematical Society, 59(8): 1094–1099, September 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.ams.

org/notices/201208/rtx120801094p.pdf.

Agou:2003:SPR

[ADN03] Simon Joseph Agou, Marc Deléglise, and Jean-Louis Nicolas. Short polynomial representations for square roots modulo p. Designs, Codes, and Cryptography, 28(1):33–44, January 2003. CODEN DCCREC. ISSN 0925-1022 (print), 1573-

7586 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/4630/45/2/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/4630/45/
2/fulltext.pdf.

Amos:1977:ACS

[ADW77a] D. E. Amos, S. L. Daniel, and M. K. Weston. Algorithm 511: CDC 6600 subroutines IBESS and JBESS for Bessel functions I_{ν}(x) and J_{ν}(x), x ≥ 0, ν ≥ 0 [S18]. ACM Transactions on Mathematical Software, 3(1):93–95, March 1977. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See erratum [Amo78].

Amos:1977:CSI

[ADW77b] D. E. Amos, S. L. Daniel, and M. K. Weston. CDC 6600 subroutines IBESS and JBESS for Bessel functions I_{ν}(x) and J_{ν}(x), x ≥ 0, ν ≥ 0. ACM Transactions on Mathematical Software, 3 (1):76–92, March 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

EC2:1992:DJN

[AEF92] Deuxièmes journées nationales: Les applications des ensembles flous, en l’honneur du Professeur A. Kaufmann, Nimes, 2–3 novembre 1992, conference scientifique (English: Second national conference: Application of Fuzzy Sets, in honor of Professor A. Kaufman, Nimes, 2–3 November 1992, scientific conference). EC2, Nanterre Cedex, France, 1992. ISBN 2-906899-78-X. LCCN ????

Alimohammad:2010:UAA

[AFC10] A. Alimohammad, S. F. Fard, and B. F. Cockburn. A unified architecture for the accurate and high-throughput implementation of six key elementary functions. IEEE Transactions on Computers, 59(4):449–456, ???? 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=5313801.

AFIPS:1969:ACPb

[AFI69] 1969 Fall Joint Computer Conference, November 18–20, 1969, Las Vegas, Nevada, volume 35 of AFIPS conference proceedings. AFIPS Press, Montvale, NJ, USA, 1969. LCCN TK7885.A1 J6 1969.

AFIPS:1971:ACP

[AFI71] 1971 Spring Joint Computer Conference, May 18–20, 1971, Atlantic City, New Jersey, volume 38 of AFIPS conference proceedings. AFIPS Press, Montvale, NJ, USA, 1971. LCCN ????

Ancarani:2008:DOC

[AG08] L. U. Ancarani and G. Gasaneo. Derivatives of any order of the confluent hypergeometric function _{1}F_{1}(a, b, z) with respect to the parameter a or b. Journal of Mathematical Physics, 49(6): 063508, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v49/i6/p063508_s1.

Alefeld:1980:PSE

[AGA^{+}80] G. Alefeld, R. D. Grigorieff, R. Albrecht, U. Kulisch, and F. Stummel, editors. Fundamentals of numerical computation (computer-oriented numerical analysis). Proceedings of a conference held June 5–8, 1979, on the occasion of the centennial of the Technical University of Berlin, volume 2 of Computing. Supplementum. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84.

Armbruster:1989:KSD

[AGH89] Dieter Armbruster, John Guckenheimer, and Philip Holmes. Kuramoto–Sivashinsky dynamics on the center-unstable manifold. SIAM Journal on Applied Mathematics, 49(3):676–691, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Andrews:1986:SCA

[AGJ86] George E. Andrews, Ian P. Goulden, and David M. Jackson. Shanks’ convergence acceleration transform, Padé approximants and partitions. Journal of Combinatorial Theory (Series A), 43 (1):70–84, 1986. CODEN JCBTA7. ISSN 0097-3165 (print), 1096-0899 (electronic).

Allasia:1993:PIJ

[AGL93] G. Allasia, Luighi Gatteshi, and Francesco Lerda, editors. Proceedings of the International Joint Symposium on Special Functions and Artificial Intelligence, (1993: Turin, Italy), volume 2(1/4) of Annals of numerical mathematics. Baltzer Science Publishers, Amsterdam, The Netherlands, 1993. ISSN 1021-2655. LCCN QA297 A614 v. 2, no. 1/4.

Agarwal:1999:SAM

[AGS99] R. C. Agarwal, F. G. Gustavson, and M. S. Schmookler. Series approximation methods for divide and square root in the Power3^{T M} processor. In Koren and Kornerup [KK99], pages 116–123. CODEN ???? ISBN 0-7803-5609-8, 0-7695-0116-8, 0-7695-0118-4. ISSN ???? LCCN

QA76.6 .S887 1999. URL http://euler.ecs.umass.edu/
paper/final/paper-144.pdf;   http://euler.ecs.umass.
edu/paper/final/paper-144.ps. IEEE Computer Society

Order Number PR00116. IEEE Order Plan Catalog Number 99CB36336.

Ahmadi:2010:LCC

[AH10] O. Ahmadi and F. R. Henríquez. Low complexity cubing and cube root computation over F_{3}^{m} in polynomial basis. IEEE Transactions on Computers, 59(10):1297–1308, October 2010. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5374372.

Aprahamian:2016:MIT

[AH16] Mary Aprahamian and Nicholas J. Higham. Matrix inverse trigonometric and inverse hyperbolic functions: Theory and algorithms. SIAM Journal on Matrix Analysis and Applications, 37(4):1453–1477, ???? 2016. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Aharoni:1969:C

[Aha69] Amikam Aharoni. Computation of K_{p}(x). Journal of Computational Physics, 4(2):270–271, August 1969. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0021999169900722.

Ahmed:1989:EEF

[Ahm89] H. M. Ahmed. Efficient elementary function generation with multipliers. In Ercegovac and Swartzlander, Jr. [ES89], pages 52–59. ISBN 0-8186-8963-3 (case), 0-8186-5963-7 (microfiche). LCCN QA 76.9 C62 S95 1989. IEEE catalog no. 89CH2757-3.

Aberbour:1998:PMF

[AHM^{+}98] M. Aberbour, A. Houelle, H. Mehrez, N. Vaucher, and A. Guyot. On portable macrocell FPU generators for division and square root operators complying to the full IEEE-754 standard. IEEE Transactions on Very Large Scale Integration (VLSI) Systems, 6(1):114–121, March 1998. CODEN IEVSE9. ISSN 1063-8210 (print), 1557-9999 (electronic).

Ahrendt:1996:FHC

[Ahr96] Timm Ahrendt. Fast high-precision computations of complex square roots. In Lakshman Y.N. [Lak96], pages 142–149. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996. URL http://www.

acm.org:80/pubs/citations/proceedings/issac/236869/p142-
ahrendt/.

Airey:1937:CFA

[Air37] J. R. Airey. The “converging factor” in asymptotic series and the calculation of Bessel, Laguerre and other functions. Philosophical Magazine, 24(162):521–552, ???? 1937. CODEN PHMAA4. ISSN 0031-8086.

Alonso:2017:EAA

[AIS^{+}17] Pedro Alonso, Javier Ibáñez, Jorge Sastre, Jesús Peinado, and Emilio Defez. Efficient and accurate algorithms for computing matrix trigonometric functions. Journal of Computational and Applied Mathematics, 309(??):325–332, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042716302321.

Aarts:2003:ASF

[AJ03] Ronald M. Aarts and Augustus J. E. M. Janssen. Approximation of the Struve function H_{1} occurring in impedance calculations. Journal of the Acoustical Society of America, 113(5):2635–2637, May 2003. CODEN JASMAN. ISSN 0001-4966.

Al-Jarrah:2002:GSB

[AJDG02] A. Al-Jarrah, K. M. Dempsey, and M. L. Glasser. Generalized series of Bessel functions. Journal of Computational and Applied Mathematics, 143(1):1–8, June 1, 2002. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042701005052.

Andrews:1981:EFM

[AJMT81] M. Andrews, D. Jaeger, S. F. McCormick, and G. D. Taylor. Evaluation of functions on microcomputers: exp(x). Computers and Mathematics with Applications, 7(6):503–508, ????

1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0898122181900341.

Adams:1993:ACA

[AK93] E. Adams and U. Kulisch, editors. Scientific Computing with Automatic Result Verification, volume 189 of Mathematics in science and engineering. Academic Press, New York, NY, USA, 1993. ISBN 0-12-044210-8. x + 612 pp. LCCN QA76.S368 1993.

Anand:2009:OCS

[AK09] C. K. Anand and W. Kahl. An optimized Cell BE special function library generated by Coconut. IEEE Transactions on Computers, 58(8):1126–1138, August 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4731241.

Akrivis:1985:ENC

[Akr85] G. Akrivis. The error norm of certain Gaussian quadrature formulae. Mathematics of Computation, 45(172):513–519, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Askey:2001:SFG

[AKS01] R. A. Askey, Tom H. Koornwinder, and Walter J. Schempp, editors. Special Functions: Group Theoretical Aspects and Applications. Mathematics and Its Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 90-277-1822-9. xxxiv + 311 pp. LCCN QA1 M428 v. 18 c.2.

Adell:2020:RAE

[AL20] José A. Adell and Alberto Lekuona. Rational approximation to Euler’s constant at a geometric rate of convergence. Mathematics of Computation, 89(325):2553–2561, January 2020. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL https://www.ams.org/AMSMathViewer;
https://www.ams.org/journals/mcom/2020-89-325/S0025-
5718-2020-03528-5;   https://www.ams.org/journals/mcom/
2020-89-325/S0025-5718-2020-03528-5/S0025-5718-2020-
03528-5.pdf;   https://www.ams.org/mathscinet/search/
authors.html?mrauthid=340766;   https://www.ams.org/
mathscinet/search/authors.html?mrauthid=663604.

Albrecht:1962:FKM

[Alb62] J. Albrecht. Fehlerschranken und Konvergenzbeschleunigung bei einer monotonen oder alternierenden Iterationsfolge. (German) [Error bounds and convergence acceleration with a monotone or alternating iteration sequence]. Numerische Mathematik, 4: 196–208, December 1962. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Antelo:1998:CVH

[ALB98] E. Antelo, T. Lang, and J. D. Bruguera. Computation of \sqrt{(x/d)} in a very high radix combined division/square-root unit with scaling and selection by rounding. IEEE Transactions on Computers, 47(2):152–161, February 1998. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=663761.

Alexander:1977:SRR

[Ale77] V. L. Alexander. Square root routine. IBM Technical Disclosure Bulletin, 20(3):1222, August 1977. CODEN IBMTAA. ISSN 0018-8689.

Alhargan:2000:ASC

[Alh00a] Fayez A. Alhargan. Algorithm 804: subroutines for the computation of Mathieu functions of integer orders. ACM Transactions on Mathematical Software, 26(3):408–414, September 2000. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Alhargan:2000:ACA

[Alh00b] Fayez A. Alhargan. Algorithms for the computation of all Mathieu functions of integer orders. ACM Transactions on Mathematical Software, 26(3):390–407, September 2000. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ali:2011:NGJ

[Ali11] Ahmad T. Ali. New generalized Jacobi elliptic function rational expansion method. Journal of Computational and Applied Mathematics, 235(14):4117–4127, May 15, 2011. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042711001257.

Antelo:2005:LLD

[ALMN05] Elisardo Antelo, Tomás Lang, Paolo Montuschi, and Alberto Nannarelli. Low latency digit-recurrence reciprocal and squareroot reciprocal algorithm and architecture. In IEEE [IEE05], page ?? ISBN ???? LCCN ???? URL http://arith17.

polito.it/final/paper-116.pdf.

Alt:1979:SRD

[Alt79] H. Alt. Square rooting is as difficult as multiplication. Computing: Archiv fur informatik und numerik, 21(3):221–232, September 1979. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Alzer:1993:SGF

[Alz93] Horst Alzer. Some gamma function inequalities. Mathematics of Computation, 60(201):337–346, January 1993. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Alzer:1997:HMI

[Alz97a] Horst Alzer. A harmonic mean inequality for the gamma function. Journal of Computational and Applied Mathematics, 87 (2):195–198, December 23, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042796001811.

See corrigendum [Alz98].

Alzer:1997:SIG

[Alz97b] Horst Alzer. On some inequalities for the gamma and psi functions. Mathematics of Computation, 66(217):373–389, January 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-97-00807-7&u=/mcom/
1997-66-217/.

Alzer:1997:SII

[Alz97c] Horst Alzer. On some inequalities for the incomplete gamma function. Mathematics of Computation, 66(218):771–778, April 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL http://www.ams.org/jourcgi/jour-
pbprocess?fn=110&arg1=S0025-5718-97-00814-4&u=/mcom/
1997-66-218/.

Alzer:1998:CHM

[Alz98] Horst Alzer. Corrigendum: A harmonic mean inequality for the gamma function [J. Comput. Appl. Math. 87 (1997) 195– 198]. Journal of Computational and Applied Mathematics, 90(2): 265, April 17, 1998. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042798000405. See [Alz97a].

Alzer:2003:GHM

[Alz03] Horst Alzer. On Gautschi’s harmonic mean inequality for the gamma function. Journal of Computational and Applied Mathematics, 157(1):243–249, August 1, 2003. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042703004564.

Alzer:2008:GFI

[Alz08] Horst Alzer. Gamma function inequalities. Numerical Algorithms, 49(1–4):53–84, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
1017-1398&volume=49&issue=1&spage=53.

Alzer:2010:EFI

[Alz10] Horst Alzer. Error function inequalities. Advances in Computational Mathematics, 33(3):349–379, October 2010. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.com/article/10.1007/s10444-
009-9139-2.

Ardill:1977:BFC

[AM77] R. W. B. Ardill and K. J. M. Moriarty. The Bessel functions J_{0} and J_{1} of complex argument. Computer Physics Communications, 13(1):17–24, May 1977. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0010465577900236.

Andrews:1978:UEF

[AM78a] M. Andrews and T. Mraz. Unified elementary function generator. Microprocessors and Microsystems, 2(5):270–273, October 1978. CODEN MIMID5. ISSN 0141-9331 (print), 1872-9436 (electronic).

Ardill:1978:SBF

[AM78b] R. W. B. Ardill and K. J. M. Moriarty. Spherical Bessel functions j_{n} and y_{n} of integer order and real argument. Computer Physics Communications, 14(3–4):261–265, May/June 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046557890019X.

Ardill:1979:ABF

[AM79] R. W. B. Ardill and K. J. M. Moriarty. Accurate Bessel functions J_{n}(z), Y_{n}(z), H_{n}^{(1)}(z) and H_{n}^{(2)}(z) of integer order and complex argument. Computer Physics Communications, 17(3):321–336, June 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Ardill:1984:ABF

[AM84a] R. W. B. Ardill and K. J. M. Moriarty. Accurate Bessel functions J_{n}(z), Y_{n}(z), H_{n}^{(1)}(z) and H_{n}^{(2)}(z) of integer order and complex argument. Computer Physics Communications, 35(1–3): C–559, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465584827344.

Ardill:1984:SBF

[AM84b] R. W. B. Ardill and K. J. M. Moriarty. Spherical Bessel functions j_{n} and y_{n} of integer order and real argument. Computer Physics Communications, 35:C–466, 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Ardill:1984:BFC

[AM84c] W. B. Ardill and K. J. M. Moriarty. The Bessel functions J_{0} and J_{1} of complex argument. Computer Physics Communications, 35:C–409, 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Avellaneda:1989:OBE

[AM89] Marco Avellaneda and Graeme W. Milton. Optimal bounds on the effective bulk modulus of polycrystals. SIAM Journal on Applied Mathematics, 49(3):824–837, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Al-Mohy:2012:MAB

[AM12] Awad H. Al-Mohy. A more accurate Briggs method for the logarithm. Numerical Algorithms, 59(3):393–402, March 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL http://www.springerlink.com/content/
4110609h521kg66m/;   http://www.springerlink.com/openurl.
asp?genre=article&issn=1017-1398&volume=59&issue=3&spage=
393.

Abergel:2020:AFA

[AM20] Rémy Abergel and Lionel Moisan. Algorithm 1006: Fast and accurate evaluation of a generalized incomplete gamma function. ACM Transactions on Mathematical Software, 46(1):10:1– 10:24, March 2020. CODEN ACMSCU. ISSN 0098-3500 (print),

1557-7295 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3365983.

Akram:2021:XFA

[AMM21] Riad Akram, Shantanu Mandal, and Abdullah Muzahid. XMeter: Finding approximable functions and predicting their accuracy. IEEE Transactions on Computers, 70(7):1081–1093, July 2021. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Amos:1973:BIC

[Amo73] D. E. Amos. Bounds on iterated coerror functions and their ratios. Mathematics of Computation, 27(122):413–427, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Amos:1974:CMB

[Amo74] D. E. Amos. Computation of modified Bessel functions and their ratios. Mathematics of Computation, 28(125):239–251, January 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Amos:1978:ECS

[Amo78] Donald E. Amos. Erratum: “Algorithm 511: CDC 6600 subroutines IBESS and JBESS for Bessel functions I_{ν}(x) and J_{ν}(x), x ≥ 0, ν ≥ 0 [S18]”. ACM Transactions on Mathematical Software, 4(4):411, December 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [ADW77a].

Amos:1983:APFa

[Amo83a] D. E. Amos. Algorithm 609: a portable FORTRAN subroutine for the Bickley functions Ki_{n}(x). ACM Transactions on Mathematical Software, 9(4):480–493, December 1983. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Amos:1983:APFb

[Amo83b] Donald E. Amos. Algorithm 610: a portable FORTRAN subroutine for derivatives of the psi function. ACM Transactions on Mathematical Software, 9(4):494–502, December 1983. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Amos:1986:APP

[Amo86] D. E. Amos. Algorithm 644: a portable package for Bessel functions of a complex argument and nonnegative order. ACM Transactions on Mathematical Software, 12(3):265–273, September 1986. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/

pubs/citations/journals/toms/1986-12-3/p265-amos/. See

also [Amo90, Amo95, Kod07].

Amos:1990:RPP

[Amo90] Donald E. Amos. Remark on “Algorithm 644: a portable package for Bessel functions of a complex argument and nonnegative order”. ACM Transactions on Mathematical Software, 16 (4):404, December 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/

pubs/citations/journals/toms/1990-16-4/p404-amos/. See

[Amo86, Amo95, Kod07].

Amos:1995:RAP

[Amo95] D. E. Amos. A remark on Algorithm 644: a portable package for Bessel functions of a complex argument and nonnegative order. ACM Transactions on Mathematical Software, 21(4): 388–393, December 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/

pubs/citations/journals/toms/1995-21-4/p388-amos/. See

[Amo86, Amo90, Kod07].

Assmus:1976:NFS

[AMS76] E. F. Assmus, Jr., H. F. Mattson, Jr., and Howard E. Sachar. A new form of the square root bound. SIAM Journal on Applied Mathematics, 30(2):352–354, March 1976. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Andrews:1978:EFM

[AMT78] M. Andrews, S. F. McCormick, and G. D. Taylor. Evaluation of functions on microcomputers: Square root. Computers and Mathematics with Applications, 4(4):359–367, 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Ascari:1968:LRG

[AN68] A. Ascari and P. G. Novario. L’algoritmo QD di Rutishauser e la generazione di funzioni speciali nel calcolo automatico. (Italian) [The QD algorithm of Rutishauser and the generation of special functions in automatic calculation]. Calcolo, 5(1):162–173, ???? 1968. CODEN CALOBK. ISSN 0008-0624 (print), 1126-

5434 (electronic). URL http://link.springer.com/article/
10.1007/BF02576063.

Andrews:1982:MMS

[And82] M. Andrews. Mathematical microprocessor software: a \sqrt{(}x) comparison. IEEE Micro, 2(3):63–79, May/June 1982. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Andrews:1998:SFM

[And98] Larry C. Andrews. Special functions of mathematics for engineers. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 1998. ISBN 0-19-856558-5 (Oxford hardcover), 0-8194-2616-4 (SPIE Press hardcover). xvii + 479 pp. LCCN QA351 .A75 1998.

Anderson:2000:RAF

[And00] Stuart Anderson. Remark on Algorithm 723: Fresnel integrals. ACM Transactions on Mathematical Software, 26(4):617, December 2000. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Anonymous:1946:MZC

[Ano46] Anonymous. More zeros of certain Bessel functions of fractional order. Mathematical Tables and Other Aids to Computation, 2 (15):118–119, July 1946. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Anonymous:1948:TBF

[Ano48] Anonymous. Tables of the Bessel Functions Y_{0}(x), Y_{1}(x), K_{0}(x), K_{1}(x), 0 ≤ x ≤ 1, volume 1 of Applied Mathematics Series / National Bureau of Standards. U.S. Government Printing Office, Washington, DC, USA, 1948. ix + 60 pp. LCCN QA3 .U5 no. 1.

Anonymous:1952:BFP

[Ano52] Anonymous. Bessel Functions, Part II, volume 10 of British Association Mathematical Tables. Cambridge University Press, Cambridge, UK, 1952. ???? pp.

Anonymous:1960:BFP

[Ano60a] Anonymous. Bessel Functions. Part III: Zeros and Associated Values, volume 7 of Royal Society Mathematical Tables. Cambridge University Press, Cambridge, UK, 1960. lx + 79 pp. Prepared under the direction of the Bessel Functions Panel of the Mathematical Tables Committee.

Anonymous:1960:EFG

[Ano60b] Anonymous. Errata to Fisher and Gupta. Technometrics, 2(4): 523–524, November 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1266462. See [FC60, Gup60].

Anonymous:1967:CAP

[Ano67] Anonymous. Convergence acceleration from the point of view of linear programming. BIT (Nordisk tidskrift for informationsbehandling), 7(3):256, September 1967. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=7&issue=3&spage=
256.

Anonymous:1968:ISA

[Ano68] Anonymous. Index by subject to algorithms, 1960–1968. Communications of the ACM, 11(12):827–830, December 1968. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1974:TCH

[Ano74] Anonymous, editor. Tables of Complex Hyperbolic and Circular Functions, volume 23 of Advanced series of mathematical and engineering tables. Corona Pub. Co., Tokyo, Japan, 1974. 621 pp. LCCN QA55 .T172.

Anonymous:1984:EJCb

[Ano84] Anonymous. Errata: J. Comput. Appl. Math. 9: H. E. Salzer, Note on the Dočev–Grosswald asymptotic series for generalized Bessel polynomials, (1983) 131–135. Journal of Computational and Applied Mathematics, 10(1):133, February 1984. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042784900773. See [Sal83].

Anonymous:1991:PIS

[Ano91] Anonymous, editor. Proceedings of the International Symposium on Supercomputing: Fukuoka, Japan, November 6–8, 1991. Kyushu University Press, Fukuoka, Japan, 1991. ISBN 4-87378-284-8. LCCN QA76.88.I 1991.

Anonymous:1993:ITS

[Ano93] Integral transforms and special functions, 1993. ISSN 1065-2469, 1476-8291. Gordon and Breach Science Publishers, Yverdon, Switzerland. Appears with variable frequency from 1993–2001, and six times yearly from 2002–date.

Anonymous:1994:C

[Ano94] Anonymous. Corrigenda. ACM Transactions on Mathematical Software, 20(4):553, December 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [HFT94].

Anonymous:1998:AMS

[Ano98a] Anonymous, editor. Analytical methods and special functions. Gordon and Breach Science Publishers, Amsterdam, The Netherlands, 1998. ISSN 1027-0264. ???? pp. LCCN A299.6 A533 v. 1 1998.

Anonymous:1998:TIC

[Ano98b] Anonymous, editor. Tricomi’s ideas and contemporary applied mathematics: (Roma, 28–29 novembre — Torino, 1–2 dicembre 1997), volume 147 of Atti dei convegni Lincei. Accademia Nazionale dei Lincei, Roma, Italy, 1998. ISSN 0391-805X. LCCN QA299.6.

Anonymous:2003:CRN

[Ano03] Anonymous, editor. 5th Conference on Real Numbers and Computers 2003 — RNC5, Lyon, France, September 2003. ????, ????, 2003. ISBN ???? LCCN ????

Anonymous:2006:PCR

[Ano06] Anonymous, editor. Proceedings of the 7th Conference on Real Numbers and Computers (RNC 7) LORIA, Nancy, France, July 10–12, 2006. ????, ????, 2006. ISBN ???? LCCN ????

Anonymous:2012:FIS

[Ano12] Anonymous. Fast inverse square root. Wikipedia article.,

March 20, 2012. URL http://en.wikipedia.org/wiki/Fast_

inverse_square_root. This article describes an algorithm for the inverse square root. The only novel feature is use of two IEEE-754 specific magic constants for 32-bit and 64-bit binary arithmetic that allow obtaining fast starting estimates for Newton–Raphson iterations by manipulating the floating-point representations as integers. The code fails to handle signed zero, Infinity, and NaN arguments, uses too few iterations, and does not adjust for rounding errors to obtain correctly-rounded results. See [Bli97].

Anonymous:2014:CLL

[Ano14] Anonymous. CR-Libm — a library of correctly rounded elementary functions in double-precision. Web site, 2014. URL

http://lipforge.ens-lyon.fr/www/crlibm/.

Anonymous:2015:L

[Ano15] Anonymous. libcerf. Web site, 2015. URL http://apps.jcns.

fz-juelich.de/doku/sc/libcerf.

Anonymous:2022:DLM

[Ano22] Anonymous. Digital library of mathematical functions: Date: 2010. NIST Web site, March 14, 2022. URL https:/

/dlmf.nist.gov/;   https://www.nist.gov/mathematics-
statistics/digital-library-mathematical-functions.

Antosiewicz:1964:BFF

[Ant64] H. A. Antosiewicz. Bessel functions of fractional order. In Abramowitz and Stegun [AS64], pages 435–478. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also avail-

able online at http://www.convertit.com/Go/ConvertIt/

Reference/AMS55.ASP in bitmap image format.

Agarwal:2008:OPD

[AO08] Ravi P. Agarwal and Donal O’Regan. Ordinary and Partial Differential Equations: with Special Functions, Fourier Series, and Boundary Value Problems. Universitext. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-79145-0 (paperback). xiv + 410 pp. LCCN ????

Aiken:1964:PAC

[AOB64] H. H. Aiken, A. G. Oettinger, and T. C. Bartee. Proposed automatic calculating machine. IEEE Spectrum, 1(8):62–69, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Alonso:2019:CMT

[API^{+}19] Pedro Alonso, Jesús Peinado, Javier Ibáñez, Jorge Sastre, and Emilio Defez. Computing matrix trigonometric functions with GPUs through Matlab. The Journal of Supercomputing, 75(3): 1227–1240, March 2019. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Abrarov:2018:RAD

[AQ18] Sanjar M. Abrarov and Brendan M. Quine. A rational approximation of the Dawson’s integral for efficient computation of the complex error function. Applied Mathematics and Computation, 321(??):526–543, March 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300317307312.

Alzer:1999:SPP

[AR99] Horst Alzer and O. G. Ruehr. A submultiplicative property of the psi function. Journal of Computational and Applied Mathematics, 101(1–2):53–60, January 15, 1999. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042798001903.

Archibald:1943:TTF

[Arc43] Raymond Clare Archibald. Tables of trigonometric functions in non-sexagesimal arguments. Mathematical Tables and Aids to Computation, 1(2):33–44, April 1943. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Arfken:1985:MMP

[Arf85] George B. (George Brown) Arfken. Mathematical methods for physicists. Academic Press, New York, NY, USA, third edition, 1985. ISBN 0-12-059820-5. xxii + 985 pp. LCCN QA37.2 .A74

1985. URL http://www.sciencedirect.com/science/book/
9780120598205.

Adj:2014:SRC

[ARH14] G. Adj and F. Rodriguez-Henriquez. Square root computation over even extension fields. IEEE Transactions on Computers, 63(11):2829–2841, November 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Armengou:1982:ASQ

[Arm82] Santiago Zarzuela Armengou. About some questions of differential algebra concerning to elementary functions. Publ. Sec. Mat. Univ. Autònoma Barcelona, 26(1):5–15, 1982.

Arnold:1977:SMF

[Arn77] Mark G. Arnold. SCELBAL Mathematical Functions Supplement (8008/8080). Scelbi Computer Consulting. Inc., 1322 Rear — Boston Post Road, Milford, CT 0646, USA, 1977.

31 pp. URL http://www.scelbi.com/files/docs/scelbal/
SCELBAL%20Mathematical%20Functions%20Supplement.pdf.

Arscott:1962:BRI

[Ars62] F. M. Arscott. Book review: Introduction to Elliptic Functions with Applications (F. Bowman). SIAM Review, 4(4):408, ???? 1962. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Arsenin:1968:BES

[Ars68] V. Ja. (Vasilii Jakovlevich) Arsenin. Basic equations and special functions of mathematical physics. Iliffe, London, UK, 1968. ISBN 0-592-05035-1. 7 + 361 pp. LCCN QC20 .A693. Translation by S. Chomet, Kings College, London of Matematicheskaëiìa fizika.

Arscott:1981:LBB

[Ars81] F. M. Arscott. The land beyond Bessel: a survey of higher special functions. Lecture Notes in Mathematics, 846:26–45, 1981. CODEN LNMAA2. ISBN 3-540-10569-7 (print), 3-540-38538-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0089822/.

Atta:1954:CGH

[AS54] Susie E. Atta and Ward C. Sangren. Calculation of generalized hypergeometric series. Journal of the Association for Computing Machinery, 1(4):170–172, October 1954. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Abramowitz:1964:HMF

[AS64] Milton Abramowitz and Irene A. Stegun, editors. Handbook of Mathematical Functions with Formulas, Graphs, and Mathematical Tables, volume 55 of Applied mathematics series. U. S. Department of Commerce, Washington, DC, USA, 1964. xiv + 1046 pp. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://www.convertit.com/

Go/ConvertIt/Reference/AMS55.ASP in bitmap image format.

Abad:1995:CRC

[AS95] Julio Abad and Javier Sesma. Computation of the regular confluent hypergeometric function. Mathematica Journal, 5(4): ??, Fall 1995. CODEN ???? ISSN 1047-5974 (print), 1097-

1610 (electronic). URL http://www.mathematica-journal.
com/issue/v5i4/article/abad/index.html.

Abad:1997:NEC

[AS97] Julio Abad and Javier Sesma. A new expansion of the confluent hypergeometric function in terms of modified Bessel functions. Journal of Computational and Applied Mathematics, 78(1):97–101, February 3, 1997. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042796001331.

Abad:2003:AEQ

[AS03] J. Abad and J. Sesma. Asymptotic expansion of the quasiconfluent hypergeometric function. Journal of Mathematical Physics, 44(4):1723–1729, April 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abad:2005:TNA

[AS05] Julio Abad and Javier Sesma. Two new asymptotic expansions of the ratio of two gamma functions. Journal of Computational and Applied Mathematics, 173(2):359–363, January 15, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042704001669.

Anand:2010:UTE

[AS10] Christopher Kumar Anand and Anuroop Sharma. Unified tables for exponential and logarithm families. ACM Transactions on Mathematical Software, 37(3):28:1–28:23, September 2010. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Anderson:2006:AMF

[ASA06] Cristina S. Anderson, Shane Story, and Nikita Astafiev. Accurate math functions on the Intel IA-32 architecture: a performance-driven design. In Anonymous [Ano06], page ?? ISBN ???? LCCN ????

Askey:1975:TAS

[Ask75] Richard Askey, editor. Theory and application of special functions: proceedings of an advanced seminar sponsored by the Mathematics Research Center, the University of Wisconsin-Madison, March 31–April 2, 1975, number 35 in Publication of the Mathematics Research Center, the University of Wisconsin. Academic Press, New York, NY, USA, 1975. ISBN 0-12-064850-4. LCCN QA3 .U45 no. 35 QA351.

Atkins:1979:FSC

[Atk79] D. E. Atkins. Fourth Symposium on Computer Arithmetic: crunching with quality and LSI. Computer, 12(4):94–97, April 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anderson:1990:FIC

[AVV90] G. D. Anderson, M. K. Vamanamurthy, and M. Vuorinen. Functional inequalities for complete elliptic integrals and their ratios. SIAM journal on mathematical analysis, 21(2):536–549, March 1990. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Anderson:1992:FIH

[AVV92] G. D. Anderson, M. K. Vamanamurthy, and M. Vuorinen. Functional inequalities for hypergeometric functions and complete elliptic integrals. SIAM journal on mathematical analysis, 23 (2):512–524, March 1992. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Arfken:2001:MMP

[AW01] George B. (George Brown) Arfken and Hans-Jurgen Weber. Mathematical methods for physicists. Harcourt/Academic Press, San Diego, CA, USA, fifth edition, 2001. ISBN 0-12-059825-6, 0-12-059826-4. xiv + 1112 pp. LCCN QA37.3 .A74

2001. URL http://catalog.hathitrust.org/api/volumes/
oclc/45705658.html.

Arfken:2005:MMP

[AW05] George B. Arfken and Hans-Jürgen Weber. Mathematical Methods for Physicists. Elsevier, Amsterdam, The Netherlands, sixth edition, 2005. ISBN 0-12-059876-0, 0-12-088584-0 (paperback). xii + 1182 pp. LCCN QA37.3 .A74 2005.

Arfken:2013:BF

[AWH13a] George B. (George Brown) Arfken, Hans-Jürgen Weber, and Frank E. Harris. Bessel functions. In Mathematical Methods for Physicists: a Comprehensive Guide [AWH13c], chapter 14, pages 643–713. ISBN 0-12-384654-4 (hardcover). LCCN QA37.3

.A74 2013. URL http://www.sciencedirect.com/science/
article/pii/B9780123846549000141.

Arfken:2013:GFb

[AWH13b] George B. (George Brown) Arfken, Hans-Jürgen Weber, and Frank E. Harris. Gamma function. In Mathematical Methods for Physicists: a Comprehensive Guide [AWH13c], chapter 13, pages 599–641. ISBN 0-12-384654-4 (hardcover). LCCN QA37.3

.A74 2013. URL http://www.sciencedirect.com/science/
article/pii/B978012384654900013X.

Arfken:2013:MMP

[AWH13c] George B. (George Brown) Arfken, Hans-Jürgen Weber, and Frank E. Harris. Mathematical Methods for Physicists: a Comprehensive Guide. Elsevier Academic Press, Amsterdam, The Netherlands, seventh edition, 2013. ISBN 0-12-384654-4 (hardcover). xiii + 1205 pp. LCCN QA37.3 .A74 2013.

Bateman:1944:GTB

[BA44] Harry Bateman and Raymond Clare Archibald. A guide to tables of Bessel functions. Mathematical Tables and Other Aids to Computation, 1(7):205–308, July 1944. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Badhe:1976:NAN

[Bad76] Sahadeo K. Badhe. New approximation of the normal distribution function. Communications in Statistics: Simulation and Computation, 5(4):173–176, 1976. CODEN CSSCDB. ISSN 0361-0918.

Bagby:1995:CNP

[Bag95] Richard J. Bagby. Calculating normal probabilities. American Mathematical Monthly, 102(1):46–48, January 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bailey:1988:CDD

[Bai88] David H. Bailey. The computation of π to 29, 360, 000 decimal digits using Borweins’ quartically convergent algorithm. Mathematics of Computation, 50(181):283–296, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bailey:1993:AMT

[Bai93] D. H. Bailey. Algorithm 719: Multiprecision translation and execution of FORTRAN programs. ACM Transactions on Mathematical Software, 19(3):288–319, September 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Baikov:2017:AID

[Bai17] Nikita Baikov. Algorithm and implementation details for complementary error function. IEEE Transactions on Computers, 66(7):1106–1118, July 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL https://www.computer.

org/csdl/trans/tc/2017/07/07792222-abs.html.

Bailey:2021:PMN

[Bai21] David H. Bailey. MPFUN2020: a new thread-safe arbitrary precision package. Web document, May 18, 2021. URL https://

www.davidhbailey.com/dhbpapers/mpfun2020.pdf.

Baker:1973:PAS

[Bak73] P. W. Baker. Predictive algorithms for some elementary functions in radix 2. Electronics Letters, 9:493–494, 1973. CODEN ELLEAK. ISBN 0013-5194. ISSN 0013-5194 (print), 1350-911X (electronic).

Baker:1975:MER

[Bak75a] P. W. Baker. More efficient radix-2 algorithms for some elementary functions. IEEE Transactions on Computers, C-24 (11):1049–1054, November 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1672725.

Baker:1975:PMA

[Bak75b] P. W. Baker. Parallel multiplicative algorithms for some elementary functions. IEEE Transactions on Computers, C-24(3):322– 325, March 1975. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1672808.

Baker:1976:SFB

[Bak76] P. W. Baker. Suggestion for a fast binary sine/cosine generator. IEEE Transactions on Computers, C-25(11):1134–1136, November 1976. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=1674566.

Baker:1992:LCE

[Bak92a] H. G. Baker. Less complex elementary functions. ACM SIG-PLAN Notices, 27(11):15–16, November 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1992:CMF

[Bak92b] Louis Baker. C Mathematical Function Handbook. McGraw-Hill programming tools for scientists and engineers. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-911158-0. xviii + 757 pp. LCCN QA351.B17 1991; QA351 .B17 1992.

Ball:2000:ACZ

[Bal00] James S. Ball. Automatic computation of zeros of Bessel functions and other special functions. SIAM Journal on Scientific Computing, 21(4):1458–1464, July 2000. CODEN SJOCE3.

ISSN 1064-8275 (print), 1095-7197 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/33907.

Banerji:2001:SFS

[Ban01] P. K. Banerji, editor. Special functions: selected articles: proceedings of the First National Conference of the Society for Special Functions and their Applications, March 3–4, 2000, Jodhpur, India. Published by Scientific Publishers (India) for the Society for Special Functions and their Applications, Jodhpur, India, 2001. ISBN 81-7233-267-X. LCCN QA351 .S665 2001.

Barakat:1961:EIG

[Bar61] Richard Barakat. Evaluation of the incomplete gamma function of imaginary argument by Chebyshev polynomials. Mathematics of Computation, 15(73):7–11, January 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bark:1972:MTF

[Bar72] L. S. Bark. Mnogoznachnye tablitsy èlementarnykh funktsiĭ (sinx, cosx, e^{x} i e^{−x}). (Russian) [Multiplace tables of the elementary functions (sin x, cos x, e^{x} and e^{−x} )]. Vyčisl. Centr Akad. Nauk SSSR, Moscow, USSR, second, unrevised edition, 1972. 134 pp. Tables processed and text translated from the English by L. S. Bark. Library of Mathematical Tables, No. 9. 1972.

Barlow:1974:CCF

[Bar74] R. H. Barlow. Convergent continued fraction approximants to generalised polylogarithms. BIT (Nordisk tidskrift for informationsbehandling), 14(1):112–116, March 1974. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=14&issue=1&spage=112.

Barnett:1976:MRC

[Bar76a] A. R. Barnett. RCWFF — modification of the real Coulomb wavefunction program RCWFN. Computer Physics Communications, 11(1):141–142, January/February 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Barnett:1976:RMR

[Bar76b] A. R. Barnett. RCWFF — modification of the real Coulomb wavefunction program RCWFN. Computer Physics Communications, 11(1):141–142, January/February 1976. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/001046557690045X.

Barnett:1981:ARI

[Bar81] A. R. Barnett. An algorithm for regular and irregular Coulomb and Bessel functions of real order to machine accuracy. Computer Physics Communications, 21(3):297–314, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Barnett:1982:CFE

[Bar82a] A. R. Barnett. Continued-fraction evaluation of Coulomb functions F_{λ}(η, x), G_{λ}(η, x) and their derivatives. Journal of Computational Physics, 46(2):171–188, May 1982. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0021999182900122.

Barnett:1982:CCB

[Bar82b] A. R. Barnett. COULFG: Coulomb and Bessel functions and their derivatives, for real arguments, by Steed’s method. Computer Physics Communications, 27(2):147–166, August 1982. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465582900704.

Barnett:1982:HPE

[Bar82c] A. R. Barnett. High-precision evaluation of the regular and irregular Coulomb wavefunctions. Journal of Computational and Applied Mathematics, 8(1):29–33, March 1982. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0771050X82900043.

Barnett:1984:CCB

[Bar84a] A. R. Barnett. Coulfg: Coulomb and Bessel functions and their derivatives, for real arguments, by Steed’s method. Computer Physics Communications, 35(1–3):C–812–C–813, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584829306.

Barnett:1984:KCF

[Bar84b] A. R. Barnett. Klein: Coulomb functions for real λ and positive energy to high accuracy. Computer Physics Communications, 35(1–3):C–753, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828842.

Barnett:1984:RMR

[Bar84c] A. R. Barnett. RCWFF — a modification of the real Coulomb wavefunction program RCWFN. Computer Physics Communications, 35(1–3):C–370, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584825850.

Baricz:2010:GBF

[Bar10a] Árpád Baricz. Generalized Bessel Functions of the First Kind, volume 1994 of Lecture Notes in Mathematics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNMAA2. ISBN 3-642-12229-9 (print), 3-642-12230-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 206 pp. LCCN QA3 .L28 no. 1994. URL http:/

/link.springer.com/book/10.1007/978-3-642-12230-9;
http://www.springerlink.com/content/978-3-642-12230-
9.

Baricz:2010:GPG

[Bar10b] Árpád Baricz. Geometric properties of generalized Bessel functions. Lecture Notes in Mathematics, 1994:23–69, 2010. CODEN LNMAA2. ISBN 3-642-12229-9 (print), 3-642-12230-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-12230-9_2.pdf.

Bartsch:2014:TMF

[Bar14] Hans-Jochen Bartsch. Taschenbuch mathematischer Formeln für Ingenieure und Naturwissenschaftler: [Für Studium und Beruf]. (German) [Pocketbook of mathematical formulas for engineers and natural sciences: [For study and job]]. Fachbuchverlag Leipzig im Hanser-Verlag, München, Germany, twenty-third edition, 2014. ISBN 3-446-43800-9. 832 pp. LCCN ???? URL

http://d-nb.info/1045240478/04.

Bashagha:2001:NRS

[Bas01] A. E. Bashagha. Novel radix-2 k square root module. Circuits, Devices and Systems, IEE Proceedings [see also IEE Proceedings G-Circuits, Devices and Systems], 148(4):190–196, August 2001. CODEN ???? ISSN ????

Batterman:2007:SSF

[Bat07] Robert W. Batterman. On the specialness of special functions (the nonrandom effusions of the divine mathematician). British Journal for the Philosophy of Science, 58(2):263–286, June 2007. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/58/2/263.full.pdf+html.

Batista:2019:ECM

[Bat19] Milan Batista. Elfun18 — a collection of MATLAB functions for the computation of elliptic integrals and Jacobian elliptic functions of real arguments. SoftwareX, 10(??): Article 100245, July/December 2019. CODEN ???? ISSN

2352-7110. URL http://www.sciencedirect.com/science/
article/pii/S2352711018302796.

Bauer:1989:BKR

[Bau89] Friedrich L. Bauer. Eine Bemerkung zu Koechers Reihen für die Eulersche Konstante. (German) [A remark on Koecher’s series for the Euler’s constant]. Bayer. Akad. Wiss. Math.-Natur. Kl. Sitzungsber., pages 27–33 (1990), 1989. ISSN 0340-7586.

Bosten:1974:RAI

[BB74] Nancy E. Bosten and E. L. Battiste. Remark on “Algorithm 179: Incomplete beta ratio”. Communications of the ACM, 17 (3):156–157, March 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Lud63, PSB76].

Belevitch:1982:SIT

[BB82] V. Belevitch and J. Boersma. On Stieltjes integral transforms involving Γ-functions. Mathematics of Computation, 38(157): 223–226, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Borwein:1984:AGM

[BB84] J. M. Borwein and P. B. Borwein. The arithmetic-geometric mean and fast computation of elementary functions. SIAM Review, 26(3):351–366, July 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1988:CFF

[BB88] J. M. Borwein and P. B. Borwein. On the complexity of familiar functions and numbers. SIAM Review, 30(4):589–601, December 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1989:MI

[BB89] J. M. Borwein and P. B. Borwein. On the mean iteration (a, b) ← (\frac{a+3b}{4}, \frac{\sqrt{ab}+b}{2}). Mathematics of Computation, 53(187):311–326, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL http://docserver.carma.newcastle.
edu.au/1586/.

Bender:1994:DAS

[BB94] Carl M. Bender and Stefan Boettcher. Determination of f(∞) from the asymptotic series for f(x) about x = 0. Journal of Mathematical Physics, 35(4):1914–1921, April 1994. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v35/i4/p1914_s1.

Borwein:1997:AGMa

[BB97] J. M. Borwein and P. B. Borwein. The arithmetic–geometric mean and fast computation of elementary functions. In Berggren et al. [BBB97], pages 537–552. ISBN 0-387-94924-0, 1-4757-2736-4 (e-book), 1-4757-2738-0 (print), 3-540-94924-0. LCCN

QA484 .P5 1997. URL http://link.springer.com/chapter/
10.1007/978-1-4757-2736-4_56. Reprint of [BB84].

Borwein:2000:AGM

[BB00] J. M. Borwein and P. B. Borwein. The arithmetic–geometric mean and fast computation of elementary functions. In Berggren et al. [BBB00], pages 537–552. ISBN 0-387-98946-3 (hardcover).

LCCN QA484 .P5 2000. URL http://link.springer.com/
chapter/10.1007/978-1-4757-3240-5_56. Reprint of [BB84].

Borwein:2004:AGMa

[BB04] J. M. Borwein and P. B. Borwein. The arithmetic–geometric mean and fast computation of elementary functions. In Berggren et al. [BBB04], pages 537–552. ISBN 0-387-20571-

3. URL http://link.springer.com/chapter/10.1007/978-

1-4757-4217-6_56. Reprint of [BB84].

Bailey:2012:AIS

[BB12] David H. Bailey and Jonathan M. Borwein. Ancient Indian square roots: An exercise in forensic paleo-mathematics. American Mathematical Monthly, 119(8):646–657, October 2012. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.08.646.pdf.

Bailey:2015:CCI

[BB15] D. H. Bailey and J. M. Borwein. Crandall’s computation of the incomplete gamma function and the Hurwitz zeta function, with applications to Dirichlet L-series. Applied Mathematics and Computation, 268(??):462–477, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300315008292.

Bailey:2016:CSC

[BB16] D. H. Bailey and J. M. Borwein. Computation and structure of character polylogarithms with applications to character Mordell–Tornheim–Witten sums. Mathematics of Computation, 85(297):295–324, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.ams.org/

journals/mcom/2016-85-297/S0025-5718-2015-02974-3;
http://www.ams.org/journals/mcom/2016-85-297/S0025-5718-
2015-02974-3/S0025-5718-2015-02974-3.pdf;   http://www.
ams.org/mathscinet/search/author.html?authorName=Borwein%
2C%20J.%20M; http://www.ams.org/mathscinet/search/author.
html?mrauthid=29355.

Borwein:1989:RME

[BBB89] J. M. Borwein, P. B. Borwein, and D. H. Bailey. Ramanujan, modular equations, and approximations to π or how to compute one billion digits of π. American Mathematical Monthly, 96(3): 201–219, March 1989. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Berggren:1997:PSB

[BBB97] Lennart Berggren, Jonathan M. Borwein, and Peter B. Borwein, editors. Pi, a source book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94924-0, 1-4757-2736-4 (e-book), 1-4757-2738-0 (print), 3-540-94924-0. xix + 716 pp. LCCN QA484 .P5 1997.

Berggren:2000:PSB

[BBB00] Lennart Berggren, Jonathan Borwein, and Peter Borwein, editors. Pi: a source book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 0-387-98946-3 (hardcover). xx + 736 pp. LCCN QA484 .P5 2000.

Berggren:2004:PSB

[BBB04] Lennart Berggren, Jonathan Borwein, and Peter Borwein, editors. Pi: a source book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2004. ISBN 0-387-20571-3. xx + 797 pp.

Bartoloni:1991:MFU

[BBC^{+}91] A. Bartoloni, C. Battista, S. Cabasino, N. Cabibbo, F. Del Prete, F. Marzano, P. S. Paolucci, R. Sarno, G. Salina, G. M. Todesco, M. Torelli, R. Tripiccione, W. Tross, and E. Zanetti. MAD, a floating-point unit for massively-parallel processors. Particle World, 2(3):65–73, ???? 1991. CODEN PARWEG. ISSN 1043-6790.

Borwein:1998:CSR

[BBC98] Jonathan M. Borwein, David M. Bradley, and Richard E. Crandall. Computational strategies for the Riemann zeta function. Report CECM-98-118, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, October 30, 1998. 68 pp. URL

http://docserver.carma.newcastle.edu.au/211;   http:
//people.reed.edu/~crandall/papers/attach01.pdf. Pub-

lished in [BBC00].

Borwein:2000:CSR

[BBC00] Jonathan M. Borwein, David M. Bradley, and Richard E. Crandall. Computational strategies for the Riemann zeta function. Journal of Computational and Applied Mathematics, 121(1–2): 247–296, September 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://people.reed.

edu/~crandall/papers/attach01.pdf.

Borwein:2008:EBF

[BBC08a] David Borwein, Jonathan M. Borwein, and O-Yeat Chan. The evaluation of Bessel functions via exp–arc integrals. Journal of Mathematical Analysis and Applications, 341(1):478–500, May 2008. CODEN JMANAK. ISSN 0022-247X (print), 1096-

0813 (electronic). URL http://adsabs.harvard.edu/abs/
2008JMAA..341..478B;   http://docserver.carma.newcastle.
edu.au/1231/;   http://www.sciencedirect.com/science/
article/pii/S0022247X07012346.

Borwein:2008:ELA

[BBC08b] David Borwein, Jonathan M. Borwein, and Richard E. Crandall. Effective Laguerre asymptotics. SIAM Journal on Numerical Analysis, 46(6):3285–3312, 2008. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://

docserver.carma.newcastle.edu.au/334/.

Backeljauw:2014:VES

[BBC^{+}14] Franky Backeljauw, Stefan Becuwe, Annie Cuyt, Joris Van Deun, and Daniel W. Lozier. Validated evaluation of special mathematical functions. Science of Computer Programming, 90 (part A)(??):2–20, September 15, 2014. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167642313001263.

Beaumont:2003:BSE

[BBD03] James Beaumont, Russell Bradford, and James H. Davenport. Better simplification of elementary functions through power series. In Proceedings of the 2003 International Symposium on Symbolic and Algebraic Computation, pages 30–36 (electronic). ACM Press, New York, NY 10036, USA, 2003.

Burgess:2017:ISC

[BBdD17] Neil Burgess, Javier Bruguera, and Florent de Dinechin, editors. 2017 IEEE 24th Symposium on Computer Arithmetic (ARITH 24), London, UK, 24–26 July 2017. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2017. ISBN 1-5386-1966-0 (print), 1-5386-1965-2, 1-5386-1964-4. ISSN 1063-6889. LCCN QA76.9.C62

S95 2017. URL http://ieeexplore.ieee.org/servlet/opac?
punumber=8019911.

Black:1984:NIS

[BBM84] Cheryl M. Black, Robert P. Burton, and Thomas H. Miller. The need for an industry standard of accuracy for elementaryfunction programs. ACM Transactions on Mathematical Software, 10(4):361–366, December 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bailey:1997:RCV

[BBP97] David Bailey, Peter Borwein, and Simon Plouffe. On the rapid computation of various polylogarithmic constants. Mathematics of Computation, 66(218):903–913, April 1997. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/journals/mcom/1997-66-
218/S0025-5718-97-00856-9/S0025-5718-97-00856-9.pdf.

Blanch:1962:TRR

[BC62] Gertrude Blanch and Donald S. Clemm. Tables relating to the radial Mathieu functions, vols. 1 & 2. Report, Aeronautical Research Labs. U. S. Government Printing Office, Washington, DC, USA, 1962.

Babb:1979:OEC

[BC79] Stanley E. Babb, Jr. and James W. Cafky. Operational evaluation of certain infinite Bessel function integrals. Mathematics of Computation, 33(147):1033–1039, July 1979. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burgess:2001:ISC

[BC01] N. Burgess and L. Ciminiera, editors. 15th IEEE Symposium on Computer Arithmetic: ARITH-15 2001: proceedings: Vail, Colorado, 11–13 June, 2001. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-1150-3; 0-7695-1152-X. ISSN 1063-6889. LCCN QA76.9.C62 S95 2001. US$145. IEEE order no. PR01150.

Backeljauw:2009:ACF

[BC09] Franky Backeljauw and Annie Cuyt. Algorithm 895: a continued fractions package for special functions. ACM Transactions on Mathematical Software, 36(3):15:1–15:20, July 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Borwein:2018:GFM

[BC18] Jonathan M. Borwein and Robert M. Corless. Gamma and factorial in the Monthly. American Mathematical Monthly, 125 (5):400–424, May 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Blanchard:2023:NMD

[BC23] Jeffrey D. Blanchard and Marc Chamberland. Newton’s method without division. American Mathematical Monthly, 130(7):606– 617, 2023. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bradford:2002:RAE

[BCD^{+}02] Russell Bradford, Robert M. Corless, James H. Davenport, David J. Jeffrey, and Stephen M. Watt. Reasoning about the elementary functions of complex analysis. Annals of Mathematics and Artificial Intelligence, 36(3):303–318, 2002. CODEN AMAIEC. ISSN 1012-2443 (print), 1573-7470 (electronic). Artificial intelligence and symbolic computation (Madrid, 2000).

Benoit:2010:DDM

[BCD^{+}10] Alexandre Benoit, Frédéric Chyzak, Alexis Darrasse, Stefan Gerhold, Marc Mezzarobba, and Bruno Salvy. The dynamic dictionary of mathematical functions (DDMF). In Fukuda et al. [FvdHJT10], pages 35–41. ISBN 3-642-15581-2 (paperback), 3-642-15582-0 (e-book). LCCN QA76.95 .I5654 2010. URL http:

//link.springer.com/book/10.1007/978-3-642-15582-6.

Bigatti:2020:MSI

[BCD^{+}20] Anna Maria Bigatti, Jacques Carette, James H. Davenport, Michael Joswig, and Timo de Wolff, editors. Mathematical Software — ICMS 2020: 7th International Conference, Braunschweig, Germany, July 13–16, 2020, Proceedings. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2020.

Bruguera:2009:PIS

[BCDH09] Javier D. Bruguera, Marius Cornea, Debjit DasSarma, and John Harrison, editors. Proceedings of the 19th IEEE Symposium on Computer Arithmetic, June 8–10, 2009, Portland, Oregon, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 0-7695-3670-0, 1-4244-4329-6. ISSN 1063-6889. LCCN QA76.6 .S887 2009.

URL http://www.ac.usc.es/arith19/.

Brisebarre:2008:EME

[BCE^{+}08] Nicolas Brisebarre, Sylvain Chevillard, Miloš D. Ercegovac, Jean-Michel Muller, and Serge Torres. An efficient method for evaluating polynomial and rational function approximations. In IEEE [IEE08], pages 233–238. ISBN 1-4244-1897-6 (paperback), 1-4244-1898-4. LCCN ???? URL http://ieeexplore.ieee.

org/servlet/opac?punumber=4569858;   http://www.gbv.
de/dms/tib-ub-hannover/631855815.pdf. IEEE catalog num-

ber CFP08063-PRT.

Brown:1994:PCL

[BCEP94] J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors. Proceedings of the Cornelius Lanczos International Centenary Conference, Raleigh, North Carolina, December 12–17, 1993, volume 73 of Proceedings in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Bunch:1991:DFA

[BCG91] K. J. Bunch, W. N. Cain, and R. W. Grow. A data fitting approach to series convergence acceleration. Applied Mathematics and Computation, 42(2 (part II)):189–195, ???? 1991. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Braumann:2018:RGF

[BCJV18] C. A. Braumann, J.-C. Cortés, L. Jódar, and L. Villafuerte. On the random gamma function: Theory and computing. Journal of Computational and Applied Mathematics, 335(??):142–155, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042717306064.

Bohlender:1993:PAF

[BCK^{+}93] G. Bohlender, D. Cordes, A. Knofel, U. Kulisch, R. Lohner, and W. V. Walter. Proposal for accurate floating-point vector arithmetic. In Adams and Kulisch [AK93], pages 87–102. ISBN 0-12-044210-8. LCCN QA76.S368 1993.

Borwein:2007:RHR

[BCR^{+}07] Peter Borwein, Stephen Choi, Brendan Rooney, Andrea Weirathmueller, et al., editors. The Riemann Hypothesis: a resource for the afficionado and virtuoso alike, volume 27 of CMS books in mathematics. Springer Science+Business Media, LLC, New York, NY, USA, 2007. ISBN 0-387-72125-8 (hardcover), 0-387-72126-6 (e-book). xiv + 533 pp. LCCN QA246 .R53 2008.

URL http://libanswers.liverpool.ac.uk/faq/182315.

Bank:1985:SEM

[BD85] Randolph E. Bank and Craig C. Douglas. Sharp estimates for multigrid rates of convergence with general smoothing and acceleration. SIAM Journal on Numerical Analysis, 22(4):617–633, August 1985. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Batten:1999:IFB

[BD99] D. Batten and P. D’arcy. Intrinsic functions boost compilers. Electrical Engineering Times, 1085:104, November 1999.

Bradford:2002:TBS

[BD02] Russell Bradford and James H. Davenport. Towards better simplification of elementary functions. In Proceedings of the 2002 International Symposium on Symbolic and Algebraic Computation, pages 16–22 (electronic). ACM Press, New York, NY 10036, USA, 2002.

Buehler:2014:CCH

[BD14] Stephan Buehler and Claude Duhr. CHAPLIN-Complex Harmonic Polylogarithms in Fortran. Computer Physics Communications, 185(10):2703–2713, October 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465514001969.

Brunie:2015:CGM

[BdDKL15] Nicolas Brunie, Florent de Dinechin, Olga Kupriianova, and Christoph Lauter. Code generators for mathematical functions. In Muller et al. [MTV15], pages 66–73. ISBN 1-4799-8665-8, 1-4799-8663-1. ISSN 1063-6889. LCCN QA76.9.C62

S95 2015. URL http://ieeexplore.ieee.org/servlet/opac?
punumber=7193754.

Bardin:1972:CFE

[BDG^{+}72] C. Bardin, Y. Dandeu, L. Gauthier, J. Guillermin. T. Lena. J.- M. Pernet, H. H. Wolter, and T. Tamura. Coulomb functions in entire (η, ρ) plane. Computer Physics Communications, 3(2): 73–87, March 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Bardin:1984:CFE

[BDG^{+}84] C. Bardin, Y. Dandeu, L. Gauthier, J. Guillermin, T. Lena, J.-M. Pernet, H. H. Wolter, and T. Tamura. Coulomb functions in entire (η, π)-plane. Computer Physics Communications, 35(1–3):C–125–C–126, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584823826.

Brezinski:1983:CAE

[BDGB83] C. Brezinski, J. P. Delahaye, and B. Germain-Bonne. Convergence acceleration by extraction of linear subsequences. SIAM Journal on Numerical Analysis, 20(6):1099–1105, December 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Babusci:2013:SME

[BDGP13] D. Babusci, G. Dattoli, K. Górska, and K. A. Penson. Symbolic methods for the evaluation of sum rules of Bessel functions. Journal of Mathematical Physics, 54(7):073501, July 2013. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Babusci:2014:SBS

[BDGP14] D. Babusci, G. Dattoli, K. Górska, and K. A. Penson. The spherical Bessel and Struve functions and operational methods. Applied Mathematics and Computation, 238(??):1–6, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300314005086.

Brisebarre:2005:NRR

[BDK^{+}05] Nicolas Brisebarre, David Defour, Peter Kornerup, Jean-Michel Muller, and Nathalie Revol. A new range-reduction algorithm. IEEE Transactions on Computers, 54(3):331–339, March 2005.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.computer.org/comp/trans/tc/
2005/03/t0331abs.htm;   http://csdl.computer.org/dl/
trans/tc/2005/03/t0331.htm;   http://csdl.computer.org/
dl/trans/tc/2005/03/t0331.pdf;   http://ieeexplore.ieee.
org/iel5/12/30205/01388197.pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388197.pdf?isnumber=30205&prod=
JNL&arnumber=1388197&arSt=+331&ared=+339&arAuthor=Brisebarre%
2C+N.%3B+Defour%2C+D.%3B+Kornerup%2C+P.%3B+Muller%2C+
J.-M.%3B+Revol%2C+N.;   http://ieeexplore.ieee.org/xpls/
abs_all.jsp?isnumber=30205&arnumber=1388197&count=13&
index=8;   http://ieeexplore.ieee.org/xpls/references.
jsp?arnumber=1388197.

Boisvert:2001:MM

[BDL^{+}01] Ronald F. Boisvert, M. J. Donahue, Daniel W. Lozier, R. McMichael, and B. W. Rust. Mathematics and measurement. NIST Journal of Research, 106(1):293–313, January/February 2001.

Boldo:2009:FVA

[BDL09] S. Boldo, M. Daumas, and Ren-Cang Li. Formally verified argument reduction with a fused multiply-add. IEEE Transactions on Computers, 58(8):1139–1145, August 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4711042.

Banuelos:1981:PCF

[BDM81] Alicia Bañuelos, Ricardo Angel Depine, and Roberto Claudio Mancini. A program for computing the Fermi–Dirac functions. Computer Physics Communications, 21(3):315–322, January 1981. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465581900126.

Beattie:1958:TFZ

[Bea58] Curtis L. Beattie. Table of first 700 zeros of Bessel functions — J_{l}(x) and J′_{l}(x). The Bell System Technical Journal, 37(3):689–697, May 1958. CODEN BSTJAN. ISSN

0005-8580. URL http://bstj.bell-labs.com/BSTJ/images/
Vol37/bstj37-3-689.pdf.

Beam:1960:ACE

[Bea60] A. Beam. Algorithm 14: Complex exponential integral. Communications of the ACM, 3(7):406, July 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Beebe:1991:ASR

[Bee91] Nelson H. F. Beebe. Accurate square root computation. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 4, 1991. 23 pp. Supplemental class notes prepared for Mathematics 118.

Beebe:1999:FAE

[Bee99] Nelson H. F. Beebe. Fast approximate exponential functions. Report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, December 7, 1999.

Beebe:2017:MFC

[Bee17] Nelson H. F. Beebe. The Mathematical-Function Computation Handbook: Programming Using the MathCW Portable Software Library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-64109-3 (hardcover), 3-319-64110-7 (e-book). xxxvi + 1114

pp. LCCN QA75.5-76.95. URL http://www.springer.com/
us/book/9783319641096.

Blair:1976:RCA

[BEJ76] J. M. Blair, C. A. Edwards, and J. H. Johnson. Rational Chebyshev approximations for the inverse of the error function. Mathematics of Computation, 30(136):827–830, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blair:1978:RCA

[BEJ78] J. M. Blair, C. A. Edwards, and J. H. Johnson. Rational Chebyshev approximations for the Bickley functions Ki_{n}(x). Mathematics of Computation, 32(143):876–886, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Belaga:1958:SPI

[Bel58] È. G. Belaga. Some problems involved in the calculation of polynomials. Doklady Akademii nauk SSSR, 123:775–777, 1958. CODEN DANKAS. ISSN 0002-3264.

Bell:1968:SFS

[Bel68] W. W. (William Wallace) Bell. Special functions for scientists and engineers. Van Nostrand, London, UK, 1968. xiv + 247 pp. LCCN QA351 .B4.

Belaga:1989:TMM

[Bel89] E. G. Belaga. Through the mincing machine with a Boolean layer cake: nonstandard computations over Boolean circuits in the lower-bounds-to-circuit-size complexity proving. Acta Informatica, 26(4):381–407, February 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bellalij:1990:SPC

[Bel90] M. Bellalij. A simultaneous process for convergence acceleration and error control. Journal of Computational and Applied Mathematics, 33(2):217–231, December 21, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Bell:2004:SFS

[Bel04] W. W. (William Wallace) Bell. Special Functions for Scientists and Engineers. Dover books on mathematics. Dover, New York, NY, USA, 2004. ISBN 0-486-43521-0. xiv + 247 pp. LCCN QA351 .B4 2004.

Brisebarre:2012:MPK

[BEM12] Nicolas Brisebarre, Milo D. Ercegovac, and Jean-Michel Muller. (M, p, k)-friendly points: a table-based method for trigonometric function evaluation. In IEEE, editor, 2012 IEEE 23rd International Conference on Application-Specific Systems, Architectures and Processors, 9–11 July 2012. Delft, The Netherlands, pages 46–52. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 0-7695-4768-0. ISSN 1063-6862.

Benton:1983:CZT

[Ben83] T. C. Benton. Common zeros of two Bessel functions. Part II. Approximations and tables. Mathematics of Computation, 41 (163):203–217, July 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

BenGhanem:1998:QFU

[Ben98] Riadh Ben Ghanem. Quadrature formulae using zeros of Bessel functions as nodes. Mathematics of Computation, 67(221):323– 336, January 1998. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL http://www.ams.org/jourcgi/
jour-pbprocess?fn=110&arg1=S0025-5718-98-00882-5&u=/
mcom/1998-67-221/.

Bergson:1968:ACR

[Ber68] A. Bergson. Certification of and remark on Algorithm 304 [S15]: Normal curve integral. Communications of the ACM, 11(4):271, April 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [HJ67a, HJ67b].

Berges:1984:AFE

[Ber84] J. C. Berges. Arithmétique et fonctions élémentaires sur minimicro calculateurs. (French) [arithmetic and elementary functions on mini-micro computers]. In Space mathematics, pages 193–229. Cépaduès, Toulouse, France, 1984.

Berry:2001:WSF

[Ber01] Michael Berry. Why are special functions special? Physics Today, 54(4):11–12, April 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v54/i4/p11_s1.

Byrd:1971:HEI

[BF71] Paul F. Byrd and Morris D. Friedman. Handbook of Elliptic Integrals for Engineers and Scientists, volume 67 of Die Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1971. ISBN 0-387-05318-2 (New York). xvi + 358 pp. LCCN QA343 .B95 1971.

Bohman:1992:FRP

[BF92] Jan Bohman and Carl-Erik Fröberg. The Γ-function revisited: power series expansions and real-imaginary zero lines. Mathematics of Computation, 58(197):315–322, January 1992.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brisebarre:2023:TME

[BF23] Nicolas Brisebarre and Silviu-Ioan Filip. Towards machineefficient rational L^{∞}-approximations of mathematical functions. In IEEE [IEE23], pages 119–126. LCCN ????

Borodin:1985:DND

[BFHT85] Allan Borodin, Ronald Fagin, John E. Hopcroft, and Martin Tompa. Decreasing the nesting depth of expressions involving square roots. Journal of Symbolic Computation, 1(2):169–188, June 1985. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Barnett:1974:CWF

[BFSG74] A. R. Barnett, D. H. Feng, J. W. Steed, and L. J. B. Goldfarb. Coulomb wave functions for all real η and ρ. Computer Physics Communications, 8(5):377–395, December 1974. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Barnett:1984:CWF

[BFSG84] A. R. Barnett, D. H. Feng, J. W. Steed, and L. J. B. Goldfarb. Coulomb wave functions for all real η and ρ. Computer Physics Communications, 35(1–3):C–285, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465584825151.

Berndt:1984:APA

[BG84] Bruce C. Berndt and Larry A. Goldberg. Analytic properties of arithmetic sums arising in the theory of the classical theta functions. SIAM journal on mathematical analysis, 15(1):143– 150, January 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baratella:1981:ABF

[BGV81] P. Baratella, M. Garetto, and G. Vinardi. Approximation of the Bessel function J_{ν}(x) by numerical integration. Journal of Computational and Applied Mathematics, 7(2):87–91, June 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0771050X81900401.

Bultheel:1999:ORF

[BGVHN99] Adhemar Bultheel, Pablo Gonzales-Vera, Erik Hendriksen, and Olav Njastad, editors. Orthogonal Rational Functions, volume 5 of Cambridge monographs on applied and computational mathematics. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-65006-2 (hardcover). xiv + 407 pp. LCCN QA404.5 .O75 1999. URL http://www.loc.

gov/catdir/description/cam029/98011646.html;   http:
//www.loc.gov/catdir/toc/cam024/98011646.html.

Barakat:1963:CES

[BH63] Richard Barakat and Agnes Houston. Chebyschev expansion of the sine and cosine integrals. Journal of Mathematics and Physics (MIT), 42(1–4):331–333, April 1963. CODEN JM-

PHA9. ISSN 0097-1421. URL https://onlinelibrary.wiley.
com/doi/epdf/10.1002/sapm1963421331.

Bingulac:1965:ACG

[BH65] S. P. Bingulac and E. A. Humo. Analog computer generation of Bessel functions of arbitrary order. IEEE Transactions on Electronic Computers, EC-14(6):886–889, December 1965. CODEN IEECA8. ISSN 0367-7508. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=4038609.

Bach:1999:NTS

[BH99] E. Bach and K. Huber. Note on taking square-roots modulo N. IEEE Transactions on Information Theory, 45(2):807–809, March 1999. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Burgess:2001:DIR

[BH01] N. Burgess and C. Hinds. Design issues in radix-4 SRT square root & divide unit. In Conference Record of the Thirty-Fifth Asilomar Conference on Signals, Systems and Computers, 2001, volume 2, pages 1646–1650. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. CODEN ???? ISSN ????

Burgess:2007:DAV

[BH07] Neil Burgess and Chris N. Hinds. Design of the ARM VFP11 divide and square root synthesisable macrocell. In Kornerup and Muller [KM07], pages 87–96. ISBN 0-7695-2854-6. ISSN 1063-6889. LCCN QA76.9.C62. URL http://www.lirmm.fr/

arith18/.

Brent:2013:FCB

[BH13] Richard P. Brent and David Harvey. Fast computation of Bernoulli, tangent and secant numbers. In D. H. Bailey, H. H. Bauschke, P. Borwein, F. Garvan, M. Théra, J. D. Vanderwerff, and H. Wolkowicz, editors, Computational and Analytical Mathematics, volume 50, pages 127–142. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-7620-8, 1-4614-7621-6 (e-book). ISSN 2194-1009.

Bober:2018:NCR

[BH18] Jonathan W. Bober and Ghaith A. Hiary. New computations of the Riemann zeta function on the critical line. Experimental mathematics, 27(2):125–137, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/10586458.2016.
1233083.

Bonan-Hamada:2005:SCF

[BHJ05] Catherine M. Bonan-Hamada and William B. Jones. Stieltjes continued fractions for polygamma functions; speed of convergence. Journal of Computational and Applied Mathematics, 179(1–2):47–55, July 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037704270400442X.

Blomquist:2009:MSC

[BHK09] Frithjof Blomquist, Werner Hofschuster, and Walter Krämer. A modified staggered correction arithmetic with enhanced accuracy and very wide exponent range. Lecture Notes in Computer Science, 5492:41–67, 2009. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

springerlink.com/content/k038294004403504/.

Brisebarre:2024:CRE

[BHMZ24] Nicolas Brisebarre, Guillaume Hanrot, Jean-Michel Muller, and Paul Zimmermann. Correctly-rounded evaluation of a function: why, how, and at what cost? Report hal-04474530, CNRS — Centre National de la Recherche Scientifique and others, Paris, France, February 23, 2024. 29 pp. URL https://hal.science/

hal-04474530.

Bustoz:1986:GFI

[BI86] Joaquin Bustoz and Mourad E. H. Ismail. On gamma function inequalities. Mathematics of Computation, 47(176):659–667, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bice:1981:AAS

[Bic81] P. K. Bice. Algorithm adds square root to micro’s arithmetic capability. Electronic Design, 29(11):146, May 1981. CODEN ELODAW. ISSN 0013-4872 (print), 1944-9550 (electronic).

Bingulac:1968:RAA

[Bin68] S. P. Bingulac. R68-38 accurate analog computer generation of Bessel functions for large ranges. IEEE Transactions on Computers, C-17(8):819, August 1968. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1687462.

Birkhoff:1954:SMM

[Bir54] Garrett Birkhoff, editor. Studies in Mathematics and Mechanics Presented to Richard von Mises by friends, colleagues, and pupils. Academic Press, New York, NY, USA, 1954. ISBN 1-4832-6356-8. ix + 353 pp. LCCN QA3 .S853.

Bustoz:2001:SFC

[BIS01] Joaquin Bustoz, Mourad Ismail, and S. K. (Sergei Konstantinovich) Suslov, editors. Special functions 2000: current perspective and future directions, volume 30 of NATO science series. Series II, Mathematics, physics, and chemistry. Kluwer Academic Publishers, Dordrecht, The Netherlands, 2001. ISBN 0-7923-7119-4, 0-7923-7120-8. xi + 520 pp. LCCN QA351 .S694 2001.

Brent:2015:BET

[BJ15] Richard P. Brent and Fredrik Johansson. A bound for the error term in the Brent–McMillan algorithm. Mathematics of Computation, 84(295):2351–2359, 2015. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/journals/mcom/2015-84-
295/S0025-5718-2015-02931-7;   http://www.ams.org/journals/
mcom/2015-84-295/S0025-5718-2015-02931-7/S0025-5718-
2015-02931-7.pdf.

Batten:1999:IBO

[BJG^{+}99] D. Batten, S. Jinturkar, J. Glossner, M. Schulte, R. Peri, and P. D’arcy. Interactions between optimizations and a new type of DSP intrinsic function. In ????, editor, Proceedings of the International Conference on Signal Processing Applications and Technologies, Orlando, Florida, November, 1999. ????, ????, 1999. ISBN ???? LCCN ???? URL http://mesa.ece.

wisc.edu/publications/cp_1999-09.pdf. Shortened version

in [BD99].

Batten:2000:NAD

[BJG^{+}00] D. Batten, S. Jinturkar, J. Glossner, M. Schulte, and P. D’arcy. A new approach to DSP intrinsic functions. In Sprague [Spr00], pages 2892–2901. ISBN 0-7695-0493-0, 0-7695-0494-9, 0-7695-0495-7. LCCN TA168 .H37 2000; TA168 .H37 2000xeb;

TA168. URL http://mesa.ece.wisc.edu/publications/cp_
2000-01.pdf.

Brisebarre:2011:APS

[BJK^{+}11] Nicolas Brisebarre, Mioara Joldes, Peter Kornerup, Érik Martin-Dorel, and Jean-Michel Muller. Augmented precision square roots and 2-D norms, and discussion on correctly rounding \sqrt{x^{2} + y^{2}}. In Schwarz and Oklobdzija [SO11], pages 23–

30. ISBN 0-7695-4318-9, 1-4244-9457-5. LCCN QA76.6.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=&
arnumber=5992105.

Bujanda:2019:CEC

[BJLP19] Blanca Bujanda, José, L. López, and Pedro J. Pagola. Convergent expansions of the confluent hypergeometric functions in terms of elementary functions. Mathematics of Computation, 88 (318):1773–1789, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.ams.

org/journals/mcom/2019-88-318/S0025-5718-2018-03389-
0;   https://www.ams.org/journals/mcom/2019-88-318/S0025-
5718-2018-03389-0/S0025-5718-2018-03389-0.pdf;   https:
//www.ams.org/mathscinet/search/authors.html?authorName=
Lopez%2C%20Jose%20L.; https://www.ams.org/mathscinet/
search/authors.html?mrauthid=636519;   https://www.ams.
org/mathscinet/search/authors.html?mrauthid=806866.

Borges:2022:HLA

[BJM22] Carlos F. Borges, Claude-Pierre Jeannerod, and Jean-Michel Muller. High-level algorithms for correctly-rounded reciprocal square roots. In IEEE [IEE22], pages 18–25. ISBN 1-66547-827-6, 1-66547-828-4. LCCN ????

Benton:1978:CZT

[BK78] T. C. Benton and H. D. Knoble. Common zeros of two Bessel functions. Mathematics of Computation, 32(142):533–535, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bytev:2016:HHF

[BK16] Vladimir V. Bytev and Bernd A. Kniehl. HYPERDIRE — HY-PERgeometric functions DIfferential REduction: Mathematicabased packages for the differential reduction of generalized hypergeometric functions: Lauricella function F_{c} of three variables. Computer Physics Communications, 206(??):78–83, September 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465516301059.

Bavier:2023:VNF

[BKdLSGL23] Eric Bavier, Nicholas Knight, Hugues de Lassus Saint-Geniès, and Eric Love. Vectorized nonlinear functions with the RISC-V vector extension. In IEEE [IEE23], pages 127–130. LCCN ????

Bowling:2009:LAC

[BKKC09] Shannon R. Bowling, Mohammad T. Khasawneh, Sittichai Kaewkuekool, and Byung Rae Cho. A logistic approximation to the cumulative normal distribution. Journal of Industrial Engineering and Management, 2(1):114–127, 2009. URL http:

//www.jiem.org/index.php/jiem/article/view/60.

Bajard:1994:BNH

[BKM94] Jean-Claude Bajard, Sylvanus Kla, and Jean-Michel Muller. BKM: a new hardware algorithm for complex elementary functions. IEEE Transactions on Computers, 43(8):955–963, 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Briggs:1945:TAL

[BL45] Lyman J. Briggs and Arnold N. Lowan. Tables of Associated Legendre Functions. Columbia University Press, New York, NY, USA, 1945. xlvi + 303 + 3 pp. LCCN QA406 .U5 1945.

Bohan:1971:ADC

[BL71] K. A. Bohan and K. V. Laščenov. The analytic definition of certain elementary functions. (Russian) questions of modern mathematics and methods of teaching it at institutions of higher learning. Leningrad. Gos. Ped. Inst. Učen. Zap., 404:59–78, 1971.

Bowers:1991:CCI

[BL91] K. L. (Kenneth L.) Bowers and J. (John) Lund, editors. Computation and control II: proceedings of the second Bozeman conference, Bozeman, Montana, August 1–7, 1990, volume 11 of Progress in systems and control theory. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1991. ISBN 0-8176-3611-0. LCCN TA329 .C644 1991. US$65.00.

Brown:1994:CAS

[BL94] Barry W. Brown and Lawrence Levy. Certification of Algorithm 708: Significant digit computation of the incomplete beta. ACM Transactions on Mathematical Software, 20(3):393–397, September 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://doi.

acm.org/10.1145/192115.192155;   http://www.acm.org/
pubs/citations/journals/toms/1994-20-3/p393-brown/. See

[DM92].

Balla:1996:SCB

[BL96] K. Balla and V. H. Linh. The simultaneous computation of Bessel functions of first and second kind. Computers and Mathematics with Applications, 31(4–5):87–97, February/March 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0898122195002200.

Boisvert:2002:HMF

[BL02] Ronald F. Boisvert and Daniel W. Lozier. Handbook of mathematical functions. In Lide [Lid02], pages 135–139. ISBN 0-8493-1247-7. LCCN QC100.U6 .C46 2002. URL https://nvlpubs.

nist.gov/nistpubs/sp958-lide/135-139.pdf;   https://

nvlpubs.nist.gov/nistpubs/sp958-lide/html/135-139.html.

Also printed as NIST Special Publication 958, Jan. 2001.

Blanch:1964:MF

[Bla64] Gertrude Blanch. Mathieu functions. In Abramowitz and Stegun [AS64], pages 721–750. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://www.convertit.com/

Go/ConvertIt/Reference/AMS55.ASP in bitmap image format.

Blair:1974:RCA

[Bla74] J. M. Blair. Rational Chebyshev approximations for the modified Bessel functions I_{0}(x) and I_{1}(x). Mathematics of Computation, 28(126):581–583, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Blinn:1997:JBC

[Bli97] James F. Blinn. Jim Blinn’s corner: Floating-point tricks. IEEE Computer Graphics and Applications, 17(4):80–84, July/August 1997. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic). Discusses use of IEEE 754 single-precision floatingpoint bit patterns as integers for implementations of fast, but low-accuracy, functions useful in computer graphics.

Bloom:1988:LCL

[Blo88] Thomas Bloom. The Lebesgue constant for Lagrange interpolation in the simplex. Journal of Approximation Theory, 54(3): 338–353, September 1988. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Briggs:2024:ISM

[BLP24] Ian Briggs, Yash Lad, and Pavel Panchekha. Implementation and synthesis of math library functions. Proceedings of the ACM on Programming Languages (PACMPL), 8(POPL):32:1–32:??, January 2024. CODEN ???? ISSN 2475-1421 (electronic). URL

https://dl.acm.org/doi/10.1145/3632874.

Brent:1980:SNA

[BM80] Richard P. Brent and Edwin M. McMillan. Some new algorithms for high-precision computation of Euler’s constant. Mathematics of Computation, 34(149):305–312, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Berrut:2004:APS

[BM04] Jean-Paul Berrut and Hans D. Mittelmann. Adaptive point shifts in rational approximation with optimized denominator. Journal of Computational and Applied Mathematics, 164–165 (??):81–92, March 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Proceedings of the 10th International Congress on Computational and Applied Mathematics University of Leuven, Belgium, 22–26 July 2002. Edited by M. J. Goovaerts, S. Vandewalle, and L. Wuytack.

Beliakov:2015:ARZ

[BM15] Gleb Beliakov and Yuri Matiyasevich. Approximation of Riemann’s zeta function by finite Dirichlet series: A multiprecision numerical approach. Experimental mathematics, 24(2):150–161, 2015. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Brisebarre:2004:ACR

[BMR04] N. Brisebarre, J.-M. Muller, and Saurabh Kumar Raina. Accelerating correctly rounded floating-point division when the divisor is known in advance. IEEE Transactions on Computers, 53 (8):1069–1072, August 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Bshouty:1997:TBA

[BMST97] Nader H. Bshouty, Yishay Mansour, Baruch Schieber, and Prasoon Tiwari. A tight bound for approximating the square root. Information Processing Letters, 63(4):211–213, September 10, 1997. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertot:2002:PGS

[BMZ02] Yves Bertot, Nicolas Magaud, and Paul Zimmermann. A proof of GMP square root. Journal of Automated Reasoning, 29(3–4): 225–252, September 2002. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (electronic). URL https://link.springer.

com/article/10.1023/A:1021987403425.

Bertot:2006:PGS

[BMZ06] Yves Bertot, Nicolas Magaud, and Paul Zimmermann. A proof of GMP square root using the Coq assistant. Research Report RR-4475, LORIA/INRIA Lorraine, Bâtiment A, Technopôle de Nancy-Brabois, 615 rue du jardin botanique, F-54602 Villers-lès-Nancy Cedex, France, 2006. 28 pp. URL ftp://ftp.inria.fr/

INRIA/publication/publi-pdf/RR/RR-4475.pdf;   ftp://
ftp.inria.fr/INRIA/publication/publi-ps-gz/RR/RR-4475.
ps.gz;   http://www.inria.fr/rrrt/rr-4475.html.

Boris:1975:NEO

[BO75] Jay P. Boris and Elaine S. Oran. Numerical evaluation of oscillatory integrals such as the modified Bessel function K_{iζ}(x). Journal of Computational Physics, 17(4):425–433, April 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999175900455.

Barrera:1993:IBS

[BO93] Tony Barrera and Pelle Olsson. An integer based square root algorithm. BIT (Nordisk tidskrift for informationsbehandling), 33(2):253–261, June 1993. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/contents/bit33.html;
http://www.springerlink.com/openurl.asp?genre=article&
issn=0006-3835&volume=33&issue=2&spage=253.

Boersma:1961:TFR

[Boe61] J. Boersma. Two formulas relating to elliptic integrals of the third kind (in Technical Notes and Short Papers). Mathematics of Computation, 15(75):296–298, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boersma:1962:CLF

[Boe62] J. Boersma. On the computation of Lommel’s functions of two variables (in Technical Notes and Short Papers). Mathematics of Computation, 16(78):232–238, April 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boersma:1969:ECW

[Boe69] J. Boersma. Expansions for Coulomb wave functions. Mathematics of Computation, 23(105):51–59, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Boersma:1991:UAB

[Boe91] J. Boersma. Uniform asymptotics of a Bessel-function series occurring in a transmission-line problem. Journal of Computational and Applied Mathematics, 37(1–3):143–159, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037704279190113X.

Boisvert:1997:QNS

[Boi97] Ronald F. Boisvert, editor. Quality of Numerical Software: Assessment and Enhancement. Proceedings of the IFIP TC2/WG2.5 Working Conference on the Quality of Numerical Software, Assessment and Enhancement, Oxford, United Kingdom, 8–12 July 1996. Chapman and Hall, Ltd., London, UK, 1997. ISBN 0-412-80530-8. LCCN QA297 .I35 1996.

Boldo:2006:PFF

[Bol06] Sylvie Boldo. Pitfalls of a full floating-point proof: Example on the formal proof of the Veltkamp/Dekker algorithms. Lecture Notes in Computer Science, pages 52–66, 2006. CO-DEN LNCSD9. ISBN 3-540-37187-7 (paperback), 3-540-37188-5. ISSN 0302-9743 (print), 1611-3349 (electronic).

Booth:1955:NAP

[Boo55] A. D. Booth. A note on approximating polynomials for trigonometric functions. Mathematical Tables and Other Aids to Computation, 9(49):21–23, January 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Borwein:1982:MNT

[Bor82] P. B. Borwein. On a method of Newman and a theorem of Bernstein. Journal of Approximation Theory, 34(1):37–41, January 1982. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Borwein:1988:PAE

[Bor88] Peter B. Borwein. Padé approximants for the q-elementary functions. Constructive Approximation, 4(4):391–402, 1988. ISSN 0176-4276 (print), 1432-0940 (electronic).

Borwein:1995:EARb

[Bor95a] P. Borwein. An efficient algorithm for the Riemann zeta function. In ????, editor, CMS Conference Proceedings, volume 27, pages 29–34. Canadian Mathematical Society, 616 Cooper Street, Ottawa, ON, K1R 5J2, Canada, January 1995.

URL https://web.archive.org/web/20140602151514/http:
//www.cecm.sfu.ca/personal/pborwein/PAPERS/P155.pdf.

Borwein:1995:EARa

[Bor95b] Peter Borwein. An efficient algorithm for the Riemann zeta function. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, January 20,

1995. 9 pp. URL http://docserver.carma.newcastle.edu.
au/107.

Borrione:2002:TIW

[Bor02] Dominique Borrione, editor. Third International Workshop on the ACL2 Theorem Prover and its Applications (ACL2-2002), April 8–9, 2002, in Grenoble, France. Presentations, affiliated with ETAPS 2002. ????, ????, 2002. ISBN ???? LCCN

???? URL http://www.cs.utexas.edu/users/moore/acl2/
workshop-2002/.

Borghi:2010:AFE

[Bor10] Riccardo Borghi. Asymptotic and factorial expansions of Euler series truncation errors via exponential polynomials. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1242– 1250, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Borges:2019:IAH

[Bor19] Carlos F. Borges. An improved algorithm for hypot(a,b). arXiv.org, ??(??):1–15, June 14, 2019. URL https://arxiv.

org/abs/1904.09481.

Borges:2021:AIA

[Bor21a] Carlos F. Borges. Algorithm 1014: an improved algorithm for hypot(x,y). ACM Transactions on Mathematical Software, 47 (1):9:1–9:12, January 2021. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://dl.acm.

org/doi/10.1145/3428446.

Borges:2021:CRN

[Bor21b] Carlos F. Borges. A correctly rounded Newton step for the reciprocal square root. arXiv.org, ??(??):1–8, December 28, 2021.

URL https://arxiv.org/abs/2112.14321.

Borges:2021:FCA

[Bor21c] Carlos F. Borges. Fast compensated algorithms for the reciprocal square root, the reciprocal hypotenuse, and Givens rotations. arXiv.org, ??(??):1–11, February 23, 2021. URL https:/

/arxiv.org/abs/2103.08694.

Bos:1989:CPR

[Bos89] L. Bos. A characteristic of points in R^{2} having Lebesgue function of polynomial growth. Journal of Approximation Theory, 56 (3):316–329, March 1989. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Bowden:1953:FTT

[Bow53a] Baron Bertram Vivian Bowden, editor. Faster Than Thought: a Symposium on Digital Computing Machines. Sir Isaac Pitman and Sons, Ltd., London, UK, 1953. LCCN QA76.5 .B66. URL

https://archive.org/details/FasterThanThought. With a

foreword by the Right Honourable the Earl of Halsbury.

Bowman:1953:IEF

[Bow53b] Frank Bowman. Introduction to Elliptic Functions with Applications. English Universities Press, London, UK, 1953. 115 pp. LCCN QA343 .B76 1953.

Bowman:1958:IBF

[Bow58] Frank Bowman. Introduction to Bessel Functions. Dover, New York, NY, USA, 1958. x + 135 pp. LCCN QA408 .B68i 1958.

Bowman:1961:IEF

[Bow61] Frank Bowman. Introduction to Elliptic Functions with Applications, volume 922. Dover, New York, NY, USA, 1961. 115 pp. LCCN QA343 .B76 1961.

Boyd:1989:CFS

[Boy89] John Philip Boyd. Chebyshev and Fourier spectral methods, volume 49 of Lecture notes in engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-51487-2, 3-540-51487-2. xvi + 798 pp. LCCN QA404.5 .B69 1989.

Boyd:2001:CFS

[Boy01] John Philip Boyd. Chebyshev and Fourier spectral methods. Dover Books on Mathematics. Dover, New York, NY, USA, second (revised). edition, 2001. ISBN 0-486-41183-4 (paperback), 0-486-14192-6 (e-book). 1375 pp. LCCN QA404.5 .B69

2001. URL http://www.freading.com/ebooks/details/r:
download/ZnJlYWQ6OTc4MDQ4NjE0MTkyMzpl.

Boyd:2009:AAC

[Boy09] John P. Boyd. Acceleration of algebraically-converging Fourier series when the coefficients have series in powers of 1/n. Journal of Computational Physics, 228(5):1404–1411, March 20, 2009. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Boyd:2014:STE

[Boy14] John P. (John Philip) Boyd. Solving transcendental equations: the Chebyshev polynomial proxy and other numerical rootfinders, perturbation series, and oracles. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2014. ISBN 1-61197-351-1 (paperback). xviii + 460 pp. LCCN

QA353.T7 B69 2014. URL http://www.loc.gov/catdir/
enhancements/fy1503/2014017078-b.html;   http://www.
loc.gov/catdir/enhancements/fy1503/2014017078-d.html;
 http://www.loc.gov/catdir/enhancements/fy1503/2014017078-
t.html.

Boyd:2015:FWC

[Boy15] John P. Boyd. Four ways to compute the inverse of the complete elliptic integral of the first kind. Computer Physics Communications, 196(??):13–18, November 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465515001733.

Bond:1967:AAF

[BP67] Gillian Bond and M. L. V. Pitteway. Algorithm 301: Airy function. Communications of the ACM, 10(5):291–292, May 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Berg:2001:CMF

[BP01] Christian Berg and Henrik L. Pedersen. A completely monotone function related to the gamma function. Journal of Computational and Applied Mathematics, 133(1–2):219–230, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042700006440.

Branrers:1981:RAZ

[BPD81] M. Branrers, R. Piessens, and M. De Meue. Rational approximations for zeros of Bessel functions. Journal of Computational Physics, 42(2):403–405, August 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0021999181902539.

Blagoveshchenskii:1976:MCM

[BPT76] Yu V. Blagoveshchenskii, B. A. Popov, and G. S. Tesler. Methods for computing mutually inverse functions. Cybernetics, 11(2):252–256, March 1976. CODEN CYBNAW. ISSN 0011-4235 (print), 2375-0189 (electronic). URL http://www.

springerlink.com/index/10.1007/BF01069867.

Besslich:1973:MDS

[BR73] P. W. Besslich and S. Raman. Multiplication, division and square root extraction methods for electronic desk calculators. Journal of the Institution of Telecommunication Engineers (India), 19(4), April 1973.

Brezinski:1991:EMT

[BR91] Claude Brezinski and Michela Redivo Zaglia. Extrapolation Methods: Theory and Practice, volume 2 of Studies in computational mathematics. North-Holland, Amsterdam, The Netherlands, 1991. ISBN 0-444-88814-4. ix + 464 pp. LCCN QA281 .B74 1991.

Bruno:1994:AAF

[BR94] Oscar P. Bruno and Fernando Reitich. Approximation of analytic functions: a method of enhanced convergence. Mathematics of Computation, 63(207):195–213, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bray:1964:CAGa

[Bra64] T. A. Bray. Certification of Algorithm 225: Gamma function with controlled accuracy. Communications of the ACM, 7(10): 586, October 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Braess:1965:MIG

[Bra65] Dietrich Braess. Monotone Iterationsfolgen bei Gleichungssystemen mit fehlerhaften Koeffizienten und Iterationsbeschleunigung. (German) [Monotone iteration sequences for equation systems with coefficients having errors, and iteration acceleration]. Numerische Mathematik, 7(1):32–41, February 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bray:1970:CAR

[Bra70] T. Bray. Certification of Algorithm 22, Ricatti–Bessel Functions of First and Second Kind. Communications of the ACM, 13(7): 448, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Braithwaite:1973:ALP

[Bra73] W. J. Braithwaite. Associated Legendre polynomials, ordinary and modified spherical harmonics. Computer Physics Communications, 5(5):390–394, May 1973. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Braess:1984:RAE

[Bra84] Dietrich Braess. On rational approximation of the exponential and the square root function. Lecture Notes in Mathematics, 1105:89–99, 1984. CODEN LNMAA2. ISBN 3-540-13899-4 (print), 3-540-39113-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0072401/.

Braune:1987:HSF

[Bra87] K. Braune. Hochgenaue Standardfunktionen für reelle und komplexe Punkte und Intervalle in beliebigen Gleitpunktrastern [English: High-Accuracy Elementary Functions for Real and Complex Points and Intervals in Arbitrary Floating-Point Systems]. Dissertation, Universität Karlsruhe, Karlsruhe, Germany, 1987. ???? pp.

Brent:1974:AAG

[Bre74] Richard P. Brent. ACM Algorithm 488: a Gaussian pseudorandom number generator [G5]. Communications of the ACM, 17(12):704–706, December 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Brent:1975:FMP

[Bre75a] R. P. Brent. Fast multiple-precision evaluation of elementary functions. Technical Report STAN-CS-75-515, Stanford University, Department of Computer Science, Stanford, CA, USA, August 1975. i + 22 pp. URL https:

//bitsavers.org/pdf/stanford/cs_techReports/STAN-CS-
75_515_Brent_Fast_Multiple-Precision_Evaluation_Of_Elementary_
Functions_Aug75.pdf.

Brent:1975:MZM

[Bre75b] R. P. (Richard P.) Brent. Multiple-precision zero-finding methods and the complexity of elementary function evaluation. Technical report, Department of Computer Science, Carnegie-Mellon University, Pittsburgh, PA, USA, 1975. 26 pp.

Brent:1976:FMP

[Bre76a] Richard P. Brent. Fast multiple-precision evaluation of elementary functions. Journal of the Association for Computing Machinery, 23(2):242–251, April 1976. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Brent:1976:MPZ

[Bre76b] Richard P. Brent. Multiple-precision zero-finding methods and the complexity of elementary function evaluation. In Traub [Tra76], pages 151–176. ISBN 0-12-697560-4. LCCN QA297.S9151 1975.

Brezinski:1977:ACA

[Bre77] Claude Brezinski. Accélération de la convergence en analyse numérique. (French) [Convergence acceleration in numerical analysis]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-08241-7, 3-540-08241-7. 313 pp. LCCN ????

Brezinski:1978:CAS

[Bre78a] C. Brezinski. Convergence acceleration of some sequences by the ϵ-algorithm. Numerische Mathematik, 29(2):173–177, January 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezinski:1978:SCA

[Bre78b] C. Brezinski. Survey on convergence acceleration methods in numerical analysis. The Mathematics Student, 46(1):28–41 (1979), 1978. CODEN MTHSBH. ISSN 0025-5742.

Brezinski:1978:AAC

[Bre78c] Claude Brezinski. Algorithmes d’accélération de la convergence: étude numérique. (French) [Algorithms for convergence acceleration: numerical study]. Éditions Technip, Paris, France, 1978. ISBN 2-7108-0341-0 (paperback). xi + 392 pp. LCCN ????

Brent:1980:UAE

[Bre80a] R. P. Brent. Unrestricted algorithms for elementary and special functions. In Lavington [Lav80], pages 613–619. ISBN 0-444-86034-7. LCCN QA 75.5 I57 1980. URL http://maths-people.

anu.edu.au/~brent/pub/pub052.html.

Brezinski:1980:GEA

[Bre80b] C. Brezinski. A general extrapolation algorithm. Numerische Mathematik, 35(2):175–187, June 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Brezinski:1982:ASG

[Bre82a] C. Brezinski. Algorithm 585: a subroutine for the general interpolation and extrapolation problems. ACM Transactions on Mathematical Software, 8(3):290–301, September 1982. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Brezinski:1982:SNC

[Bre82b] Claude Brezinski. Some new convergence acceleration methods. Mathematics of Computation, 39(159):133–145, July 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brezinski:1983:ECC

[Bre83] Claude Brezinski. Error control in convergence acceleration processes. IMA Journal of Numerical Analysis, 3(1):65–80, 1983. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Brezinski:1985:CAM

[Bre85] Claude Brezinski. Convergence acceleration methods: the past decade. Journal of Computational and Applied Mathematics, 12–13(??):19–36, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Brezinski:1988:NAC

[Bre88] Claude Brezinski. A new approach to convergence acceleration methods. In Nonlinear numerical methods and rational approximation (Wilrijk, 1987), volume 43 of Math. Appl., pages 373– 405. Reidel, Dordrecht, 1988.

Brezinski:1999:EEC

[Bre99] C. Brezinski. Error estimates and convergence acceleration. In Error control and adaptivity in scientific computing (Antalya, 1998), volume 536 of NATO Sci. Ser. C Math. Phys. Sci., pages 87–94. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1999.

Brezinski:2000:CAD

[Bre00] C. Brezinski. Convergence acceleration during the 20th century. Journal of Computational and Applied Mathematics, 122(1–2): 1–21, October 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Numerical analysis 2000, Vol. II: Interpolation and extrapolation.

Brent:2006:FAH

[Bre06] Richard P. Brent. Fast algorithms for high-precision computation of elementary functions. Report ??, Australian National University, Canberra, ACT 0200, Australia, July 12, 2006.

61 pp. URL http://rnc7.loria.fr/brent_invited.pdf;
https://maths-people.anu.edu.au/~brent/pd/RNC7t.pdf.

Brent:2010:UAE

[Bre10] Richard P. Brent. Unrestricted algorithms for elementary and special functions. arxiv.org, ??(??):1–13, April 2010. URL

https://arxiv.org/abs/1004.3621.

Brent:2017:JBP

[Bre17] Richard P. Brent. Jonathan Borwein, pi and the AGM. Talk slides, Australian National University and CARMA, University of Newcastle, Canberra, ACT and Newcastle, NSW, Australia, September 26, 2017. 76 pp. URL https:/

/carma.newcastle.edu.au/meetings/jbcc/abstracts/pdf/
JBCC-Richard_Brent.pdf.

Bremer:2018:ANE

[Bre18] James Bremer. An algorithm for the numerical evaluation of the associated Legendre functions that runs in time independent of degree and order. Journal of Computational Physics, 360(??): 15–38, May 1, 2018. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911830024X.

Bremer:2019:ARN

[Bre19] James Bremer. An algorithm for the rapid numerical evaluation of Bessel functions of real orders and arguments. Advances in Computational Mathematics, 45(1):173–211, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044

(electronic). URL https://link.springer.com/article/10.
1007/s10444-018-9613-9.

Bronstein:1989:SRE

[Bro89a] M. Bronstein. Simplification of real elementary functions. In ACM [ACM89], pages 207–211. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

Bronstein:1989:AIE

[Bro89b] Manuel Bronstein. An algorithm for the integration of elementary functions. Lecture Notes in Computer Science, 378:491–497, 1989. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). EUROCAL ’87 (Leipzig, 1987).

Bronstein:1990:IEF

[Bro90] Manuel Bronstein. Integration of elementary functions. Journal of Symbolic Computation, 9(2):117–173, February 1990. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Brown:2007:PIS

[Bro07] C. W. Brown, editor. Proceedings of the 2007 International Symposium on Symbolic and Algebraic Computation, July 29–August 1, 2007, University of Waterloo, Waterloo, Ontario, Canada. ACM Press, New York, NY 10036, USA, 2007. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Bruguera:2022:LLH

[Bru22] Javier D. Bruguera. Low-latency and high-bandwidth pipelined radix-64 division and square root unit. In IEEE [IEE22], pages 10–17. ISBN 1-66547-827-6, 1-66547-828-4. LCCN ????

Bryc:2002:UAR

[Bry02] W. Bryc. A uniform approximation to the right normal tail integral. Applied Mathematics and Computation, 127(2–3):365– 374, April 15, 2002. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.elsevier.

com/gej-ng/10/9/12/123/27/44/abstract.html;   http:/
/www.sciencedirect.com/science/article/pii/S0096300301000157.

Bowman:1978:ASS

[BS78] K. O. Bowman and L. R. Shenton. Asymptotic series and Stieltjes continued fractions for a gamma function ratio. Journal of Computational and Applied Mathematics, 4(2):105–111, June 1978. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0771050X78900347.

Borjesson:1979:SAE

[BS79] P. Börjesson and C.-E. Sundberg. Simple approximations of the error function Q(x) for communications applications. IEEE Transactions on Communications, 27(3):639–643, March 1979. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Bowman:1983:CFP

[BS83] K. O. Bowman and L. R. Shenton. Continued fractions and the polygamma functions. Journal of Computational and Applied Mathematics, 9(1):29–39, March 1983. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0377042783900262.

Bell:1984:CFN

[BS84] K. L. Bell and N. S. Scott. Coulomb functions (negative energies). Computer Physics Communications, 35(1–3):C–648, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584828088.

Buhring:1998:ACG

[BS98] Wolfgang Bühring and H. M. Srivastava. Analytic continuation of the generalized hypergeometric series near unit argument with emphasis on the zero-balanced series. In Themistocles M. Rassias, editor, Approximation theory and applications, pages 17–35. Hadronic Press, Palm Harbor, FL, USA, 1998. ISBN 1-57485-041-5. LCCN QA297.5 .A685 1998. URL http://catalog.

hathitrust.org/api/volumes/oclc/42786578.html.

Becken:2000:ACG

[BS00] W. Becken and P. Schmelcher. The analytic continuation of the Gaussian hypergeometric function _{2}F_{1}(a, b; c; z) for arbitrary parameters. Journal of Computational and Applied Mathematics, 126(1–2):449–478, December 30, 2000. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042700002673.

Bogolubsky:2006:FEH

[BS06] A. I. Bogolubsky and S. L. Skorokhodov. Fast evaluation of the hypergeometric function _{p}F_{p−1}(a; b; z) at the singular point z = 1 by means of the Hurwitz zeta function ζ(α, s). Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 32(??):145–153, ???? 2006. CODEN PCSODA. ISSN 0361-7688 (print), 1608-3261 (electronic).

Borwein:2011:SVG

[BS11] Jonathan M. Borwein and Armin Straub. Special values of generalized log-sine integrals. In Schost and Emiris [SE11], pages 43–50. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Beals:2013:MFG

[BS13] Richard Beals and Jacek Szmigielski. Meijer G-functions: a gentle introduction. Notices of the American Mathematical Society, 60(7):866–872, August 2013. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Booker:2013:BAB

[BST13] Andrew R. Booker, Andreas Strömbergsson, and Holger Then. Bounds and algorithms for the K-Bessel function of imaginary order. LMS Journal of Computation and Mathematics, 16:78– 108, 2013. CODEN ???? ISSN 1461-1570.

Baratchart:1995:RIE

[BSW95] L. Baratchart, E. B. Saff, and F. Wielonsky. Rational interpolation of the exponential function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(??):1121– 1147, ???? 1995. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bazarov:1982:EEF

[BSY82] M. B. Bazarov, Yu. I. Shokin, and Z. Kh. Yuldashev. On the evaluation of elementary functions in interval analysis (in Russian). Technical report, Applied Mathematics and Mechanics, Tashkent State Univ., Tashkent, USSR, 1982. 26–31 pp.

Bui:1999:DSI

[BT99] H. Bui and S. Tahar. Design and synthesis of an IEEE-754 exponential function. In 1999 IEEE Canadian Conference on Electrical and Computer Engineering, 9–12 May 1999, volume 1, pages 450–455. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. CODEN ???? ISSN ????

Butts:2011:RDR

[BTDS11] J. Adam Butts, Ping Tak Peter Tang, Ron O. Dror, and David E. Shaw. Radix-8 digit-by-rounding: Achieving high-performance reciprocals, square roots, and reciprocal square roots. In Schwarz and Oklobdzija [SO11], pages 149–158. ISBN 0-7695-4318-9, 1-4244-9457-5. LCCN QA76.6. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=5992120.

Bao:2016:SAO

[BTT16] Vo Nguyen Quoc Bao, Luu Pham Tuyen, and Huynh Huu Tue. A survey on approximations of one-dimensional Gaussian Qfunction. REV Journal on Electronics and Communications, 5

(1–2), February 2016. URL http://www.rev-jec.org/index.
php/rev-jec/article/view/92.

Buchholz:1969:CHF

[Buc69] Herbert Buchholz. The confluent hypergeometric function with special emphasis on its applications, volume 15 of Springer tracts in natural philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. xviii + 238 pp.

LCCN QA351 .B813. Translation by H. Lichtblau and K. Wetzel to English from German “Die konfluente hypergeometrische Funktion.”.

Buhring:1987:BUA

[Büh87] Wolfgang Bühring. The behavior at unit argument of the hypergeometric function _{3}F_{2}. SIAM journal on mathematical analysis, 18(5):1227–1234, September 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Buhring:1992:GHF

[Büh92] Wolfgang Bühring. Generalized hypergeometric functions at unit argument. Proceedings of the American Mathematical Society, 114(1):145–153, ???? 1992. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Buhring:2003:PSH

[Büh03] Wolfgang Bühring. Partial sums of hypergeometric functions of unit argument. Proceedings of the American Mathematical Society, 132(2):407–415, ???? 2003. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Bujnowski:1975:EFC

[Buj75] G. Rudnicki Bujnowski. Explicit formulas for Clebsch–Gordan coefficients. Computer Physics Communications, 10(4):245–250, October 1975. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465575900697.

Bulirsch:1965:NCEa

[Bul65a] R. Bulirsch. Numerical calculation of elliptic integrals and elliptic functions. Numerische Mathematik, 7(1):78–90, February 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bulirsch:1965:NCEb

[Bul65b] R. Bulirsch. Numerical calculation of elliptic integrals and elliptic functions. II. Numerische Mathematik, 7(4):353–354, August 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bulirsch:1969:EBT

[Bul69a] R. Bulirsch. An extension of the Bartky-transformation to incomplete elliptic integrals of the third kind. Numerische Mathematik, 13(3):266–284, July 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bulirsch:1969:NCE

[Bul69b] R. Bulirsch. Numerical calculation of elliptic integrals and elliptic functions. III. Numerische Mathematik, 13(4):305–315, August 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Bunck:2009:FAE

[Bun09] Benjamin F. Bunck. A fast algorithm for evaluation of normalized Hermite functions. BIT Numerical Mathematics, 49(2):281–295, June 2009. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=49&issue=2&spage=281.

Burgoyne:1963:AKF

[Bur63] F. D. Burgoyne. Approximations to Kelvin functions (in Technical Notes and Short Papers). Mathematics of Computation, 17 (83):295–298, July 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burgoyne:1964:GTF

[Bur64] F. D. Burgoyne. Generalized trigonometric functions (in Technical Notes and Short Papers). Mathematics of Computation, 18 (86):314–316, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burrell:1974:AAE

[Bur74] Keith H. Burrell. ACM Algorithm 484: Evaluation of the modified Bessel functions K_{0}(z) and K_{1}(z) for complex arguments [S17]. Communications of the ACM, 17(9):524–526, September 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Burr:1982:CCR

[Bur82] S. A. Burr. Computing cube roots when a fast square root is available. Computers and Mathematics with Applications, 8(3):181–183, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0898122182900414.

Buschman:1974:FSR

[Bus74] R. G. Buschman. Finite sum representations for partial derivatives of special functions with respect to parameters. Mathematics of Computation, 28(127):817–824, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Buschman:1976:IHF

[Bus76] R. G. Buschman. Inequalities for hypergeometric functions. Mathematics of Computation, 30(134):303–305, April 1976. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Bannur:1985:VIS

[BV85] J. Bannur and A. Varma. The VLSI implementation of a square root algorithm. In Hwang [Hwa85], pages 159–165. ISBN 0-8186-0632-0 (paperback), 0-8186-8632-4 (hard), 0-8186-4632-2 (microfiche). LCCN QA76.9.C62 S95 1985. URL http:

//www.acsel-lab.com/arithmetic/arith7/papers/ARITH7_

Bannur_Varma.pdf. IEEE catalog number 85CH2146-9. IEEE Computer Society order number 632.

Birge:1989:SUB

[BW89] John R. Birge and Roger J. Wets. Sublinear upper bounds for stochastic programs with recourse. Mathematical Programming, 43(2):131–149, February 1989. CODEN MHPGA4. ISSN 0025-5610.

Bressoud:2000:CCN

[BW00] David M. Bressoud and S. Wagon. A Course in Computational Number Theory. Key College Publishers in cooperation with Springer, New York, NY, USA, 2000. ISBN 1-930190-10-7. xii + 367 pp. LCCN QA241 .B788

2000.  URL http://www.loc.gov/catdir/enhancements/
fy0818/99016037-d.html;   http://www.loc.gov/catdir/
enhancements/fy0818/99016037-t.html.

Beals:2010:SFG

[BW10] Richard Beals and R. (Roderick) Wong. Special functions: a graduate text, volume 126 of Cambridge studies in advanced mathematics. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19797-X. ix + 456 pp. LCCN QA351 .B34

2010; QA351 BEA 2010. URL http://assets.cambridge.
org/97805211/97977/cover/9780521197977.jpg;   http:/
/www.loc.gov/catdir/enhancements/fy1009/2010017815-b.
html;   http://www.loc.gov/catdir/enhancements/fy1009/
2010017815-d.html;   http://www.loc.gov/catdir/enhancements/
fy1009/2010017815-t.html.

Bohlender:1991:SEF

[BWKM91] G. Bohlender, W. Walter, P. Kornerup, and D. W. Matula. Semantics for exact floating point operations. In Kornerup and Matula [KM91], pages 22–26. ISBN 0-8186-9151-4 (case), 0-8186-6151-8 (microfiche), 0-7803-0187-0 (library binding). LCCN QA76.9.C62 S95 1991. IEEE Catalog number 91CH3015-5.

Barnett:2007:HPV

[BY07] Ross Barnett and J. A. Youngman. High-precision values of the gamma function of real argument. In 1st Joint Meeting of the American Mathematical Society and the New Zealand Mathematical Society, Victoria University of Wellington, Wellington, New Zealand, December 12–15, 2007, page ???? American Mathematical Society, Providence, RI, USA, 2007. ISBN

???? LCCN ???? URL http://homepages.mcs.vuw.ac.nz/~
mathmeet/amsnzms2007/abstracts.pdf.

Bernstein:2019:FCT

[BY19] Daniel J. Bernstein and Bo-Yin Yang. Fast constant-time gcd computation and modular inversion. Technical report, International Association for Cryptologic Research, ????, April 13,

2019. 1–59 pp. URL https://eprint.iacr.org/2019/266.
pdf.

Borwein:1992:FEG

[BZ92] J. M. Borwein and I. J. Zucker. Fast evaluation of the gamma function for small rational fractions using complete elliptic integrals of the first kind. IMA Journal of Numerical Analysis, 12 (4):519–526, 1992. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Cherri:2000:PCC

[CA00] A. K. Cherri and M. S. Alam. Parallel computation of complex elementary functions using quaternary signed-digit arithmetic.

Optics and Laser Technology, 32(6):391–399, 2000. CODEN ???? ISSN 0030-3992.

Cadwell:1951:BNI

[Cad51] J. H. Cadwell. The bivariate normal integral. Biometrika, 38 (3/4):475–479, December 1951. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2332596.

Cahill:1954:PCH

[Cah54] W. F. Cahill. Programs for computing the hypergeometric series (in automatic computing machinery; discussions). Mathematical Tables and Other Aids to Computation, 8(45):36–37, January 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Cai:2011:CSB

[Cai11] Liang-Wu Cai. On the computation of spherical Bessel functions of complex arguments. Computer Physics Communications, 182(3):663–668, March 2011. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465510004650.

Campbell:1979:BFR

[Cam79] J. B. Campbell. Bessel functions J_{ν}(x) and Y_{ν}(x) of real order and real argument. Computer Physics Communications, 18(1):133–142, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465579900304.

Campbell:1981:BFR

[Cam81] J. B. Campbell. Bessel functions I_{ν}(z) and K_{ν}(z) of real order and complex argument. Computer Physics Communications, 24(1):97–105, September/October 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0010465581901090.

Campbell:1984:BFZ

[Cam84a] J. B. Campbell. Bessel functions I_{ν}(z) and K_{ν}(z) of real order and complex argument. Computer Physics Communications, 35(1–3):C–747–C–748, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828805.

Campbell:1984:BFRa

[Cam84b] J. B. Campbell. Bessel functions J_{ν}(x) of real order and real argument. Computer Physics Communications, 35(1–3): C–583, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465584827563.

Campbell:1986:NSR

[Cam86] R. A. Campbell. NS32000 square roots. Dr. Dobb’s Journal of Software Tools, 11(3):122–123, 106, March 1986. CODEN DDJOEB. ISSN 1044-789X.

Carlitz:1963:IEF

[Car63] L. Carlitz. The inverse of the error function. Pacific Journal of Mathematics, 13(2):459–470, 1963. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http:

//pjm.math.berkeley.edu/pjm;   http://projecteuclid.
org/DPubS/Repository/1.0/Disseminate?handle=euclid.pjm/
1103035736&view=body&content-type=pdf_1.

Carlson:1969:CBE

[Car69] B. C. Carlson. A connection between elementary functions and higher transcendental functions. SIAM Journal on Applied Mathematics, 17:116–148, 1969. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Carlitz:1970:SRF

[Car70a] L. Carlitz. Some reduction formulas for generalized hypergeometric functions. SIAM journal on mathematical analysis, 1(2): 243–245, May 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1970:HSS

[Car70b] B. C. Carlson. Hidden symmetries of special functions. SIAM Review, 12(3):332–345, July 1970. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://www.

jstor.org/stable/2028552.

Carlson:1972:ACL

[Car72] B. C. Carlson. An algorithm for computing logarithms and arctangents. Mathematics of Computation, 26(118):543–549, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carta:1974:HLR

[Car74] David G. Carta. Help!!: The lost reference: (A modified Newton method for square roots). ACM SIGNUM Newsletter, 9(4): 9, October 1974. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Carta:1975:LOA

[Car75] David G. Carta. Low-order approximations for the normal probability integral and the error function. Mathematics of Computation, 29(131):856–862, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1977:EIF

[Car77a] B. C. Carlson. Elliptic integrals of the first kind. SIAM journal on mathematical analysis, 8(2):231–242, ???? 1977. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Carlson:1977:SFA

[Car77b] Bille Chandler Carlson. Special Functions of Applied Mathematics. Academic Press, New York, NY, USA, 1977. ISBN 0-12-160150-1. xv + 335 pp. LCCN QA351 .C32.

Carlson:1979:CEI

[Car79] B. C. Carlson. Computing elliptic integrals by duplication. Numerische Mathematik, 33(1):1–16, March 1979. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carroll:1982:TST

[Car82] Robert Wayne Carroll. Transmutation, scattering theory, and special functions, volume 87; 69 of North-Holland mathematics studies. North-Holland, Amsterdam, The Netherlands, 1982. ISBN 0-444-86426-1 (paperback). x + 457 pp. LCCN QA1 .N86 no. 87; QA329.

Carlson:1987:TEI

[Car87] B. C. Carlson. A table of elliptic integrals of the second kind. Mathematics of Computation, 49(180):595–606, S13–S17, October 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1988:TEI

[Car88] B. C. Carlson. A table of elliptic integrals of the third kind. Mathematics of Computation, 51(183):267–280, S1–S5, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1989:TEI

[Car89] B. C. Carlson. A table of elliptic integrals: Cubic cases. Mathematics of Computation, 53(187):327–333, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carre:1990:PEF

[Car90] C. Carre. Plethysm of elementary functions. Bayreuth. Math. Schr., 31:1–18, 1990. ISSN 0172-1062.

Carlson:1991:TEI

[Car91] B. C. Carlson. A table of elliptic integrals: One quadratic factor. Mathematics of Computation, 56(193):267–280, January 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1992:TEI

[Car92] B. C. Carlson. A table of elliptic integrals: Two quadratic factors. Mathematics of Computation, 59(199):165–180, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Carlson:1995:NCR

[Car95] B. C. Carlson. Numerical computation of real or complex elliptic integrals. Numerical Algorithms, 10(1–2):13–26, July 1995. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Special functions (Torino, 1993).

Carlson:1999:TSI

[Car99] B. C. Carlson. Toward symbolic integration of elliptic integrals. Journal of Symbolic Computation, 28(6):739–753, December 1999. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-

855X (electronic). URL http://www.idealibrary.com/links/
doi/10.1006/jsco.1999.0336/production;   http://www.
idealibrary.com/links/doi/10.1006/jsco.1999.0336/production/
pdf;   http://www.idealibrary.com/links/doi/10.1006/
jsco.1999.0336/production/ref.

Carlson:2005:JEF

[Car05] B. C. Carlson. Jacobian elliptic functions as inverses of an integral. Journal of Computational and Applied Mathematics, 174(2):355–359, February 15, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042704002201.

Cash:1983:BRKa

[Cas83] J. R. Cash. Block Runge–Kutta methods for the numerical integration of initial value problems in ordinary differential equations. Part I. The nonstiff case. Mathematics of Computation, 40 (161):175–191, January 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cathey:1985:ISR

[Cat85] James Cathey. 68000 integer square root routine in 16BST. Dr. Dobb’s Journal of Software Tools, 10(5):118–??, May 1985. CODEN DDJOEB. ISSN 1044-789X.

Cathey:1986:LEI

[Cat86] J. Cathey. Letter to the editor [integer square root]. Dr. Dobb’s Journal of Software Tools, 11(8):14, 82–85, August 1986. CO-DEN DDJOEB. ISSN 1044-789X.

Causley:2022:GFI

[Cau22] Matthew F. Causley. The gamma function via interpolation. Numerical Algorithms, 90(2):687–707, June 2022. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11075-021-01204-8.

Cawley:2000:FCA

[Caw00] G. C. Cawley. On a fast, compact approximation of the exponential function. Neural Computation, 12(9):2009–2012, September 1, 2000. CODEN NEUCEB. ISSN 0899-7667 (print), 1530-888x (electronic).

Crandall:1987:EFE

[CB87] R. E. Crandall and J. P. Buhler. Elementary function expansions for Madelung constants. Journal of Physics A (Mathematical and General), 20(16):5497–5510, 1987. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Chen:2009:SPA

[CB09] Yunfei Chen and Norman C. Beaulieu. A simple polynomial approximation to the Gaussian Q-function and its application. IEEE Communications Letters, 13(2):124–126, February 2009. CODEN ICLEF6. ISSN 1089-7798 (print), 1558-2558

(electronic). URL http://ieeexplore.ieee.org/document/
4783779/.

Chen:2012:SIM

[CB12] Chao-Ping Chen and Necdet Batir. Some inequalities and monotonicity properties associated with the gamma and psi functions. Applied Mathematics and Computation, 218(17):8217– 8225, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300312001257.

Cuyt:2010:VSF

[CBBV10] Annie Cuyt, Franky Backeljauw, Stefan Becuwe, and Joris Van Deun. Validated special functions software. Lecture Notes in Computer Science, 6327:32–34, 2010. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-15582-6_6.pdf.

Cyvin:1964:AGF

[CC64] S. J. Cyvin and B. N. Cyvin. Algorithm 225: Gamma function with controlled accuracy. Communications of the ACM, 7(5): 295, May 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Clark:1969:SCE

[CC69] N. W. Clark and W. J. Cody. Self-contained exponentiation. In AFIPS FJCC ’69 [AFI69], pages 701–706. LCCN TK7885.A1 J6 1969.

Chen:2017:UTS

[CC17] Chao-Ping Chen and Junesang Choi. Unified treatment of several asymptotic expansions concerning some mathematical constants. Applied Mathematics and Computation, 305(??):348– 363, July 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300317300978.

Cappuccino:1996:DDH

[CCC96] G. Cappuccino, P. Corsonello, and G. Cocorullo. Design and demonstration of high throughput square rooting circuit. Electronics Letters, 32(5):434, ???? 1996. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Cappuccino:1999:HSS

[CCCP99] G. Cappuccino, G. Cocorullo, P. Corsonello, and S. Perri. High speed self-timed pipelined datapath for square rooting. IEE Proceedings on Circuits, Devices and Systems [see also IEE Proceedings G — Circuits, Devices and Systems], 146(1):16–22, February 1999. CODEN ???? ISSN ????

Chang:2011:CTB

[CCM11] Seok-Ho Chang, Pamela C. Cosman, and Laurence B. Milstein. Chernoff-type bounds for the Gaussian error function. IEEE Transactions on Communications, 59(11):2939–2944, November 2011. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Colman:2011:VCC

[CCV11] Michel Colman, Annie Cuyt, and Joris Van Deun. Validated computation of certain hypergeometric functions. ACM Transactions on Mathematical Software, 38(2):11:1–11:20, December 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Chiani:2002:IEB

[CD02] M. Chiani and D. Dardari. Improved exponential bounds and approximation for the Q-function with application to average error probability computation. In Global Telecommunications Conference, 2002. GLOBECOM ’02. IEEE. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910,

USA, 2002. URL http://ieeexplore.ieee.org/document/
1188428/.

Cerone:2008:AIS

[CD08] Pietro Cerone and Sever Silvestru Dragomir, editors. Advances in Inequalities for Special Functions. Advances in mathematical inequalities. Nova Science Publishers, New York, NY, USA, 2008. ISBN 1-60021-919-5 (hardcover), 1-60692-621-7 (e-book). 170 pp. LCCN QA351 .A375 2008. URL http://catalog.

hathitrust.org/api/volumes/oclc/159822357.html.

Corless:2001:RAE

[CDJ^{+}01] Robert M. Corless, James H. Davenport, David J. Jeffrey, Gurjeet Litt, and Stephen M. Watt. Reasoning about the elementary functions of complex analysis. In Artificial intelligence and symbolic computation (Madrid, 2000), volume 1930 of Lecture Notes in Comput. Sci., pages 115–126. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001.

Choong:1971:RA

[CDR71] K. Y. Choong, D. E. Daykin, and C. R. Rathbone. Rational approximations to π. Mathematics of Computation, 25(114):387– 392, April 1971. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL http://www.ams.org/journals/
mcom/1971-25-114/S0025-5718-1971-0300981-0. See errata

[Sha76].

Cocolicchio:2000:ASF

[CDS00] Decio Cocolicchio, G. Dattoli, and H. M. Srivastava, editors. Advanced special functions and applications: proceedings of the workshop: Melfi (PZ), Italy, 9–12 May 1999, volume 1 of Proceedings of the Melfi school on advanced topics in mathematics and physics. Aracne, Roma, Italy, 2000. ISBN 88-7999-265-X. LCCN QA351 .A38 2000.

Chiani:2003:NEB

[CDS03] M. Chiani, D. Dardari, and M. K. Simon. New exponential bounds and approximations for the computation of error probability in fading channels. IEEE Transactions on Wireless Communications, 24(5):840–845, May 2003. CODEN ITWCAX. ISSN 1536-1276 (print), 1558-2248 (electronic). URL http:/

/ieeexplore.ieee.org/document/1210748/.

Chen:2015:AEC

[CE15] Chao-Ping Chen and Neven Elezović. Asymptotic expansions and completely monotonic functions associated with the gamma, psi and polygamma functions. Applied Mathematics and Computation, 269(??):232–241, October 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300315009637.

Cerone:2007:SFA

[Cer07] Pietro Cerone. Special functions: approximations and bounds. Applicable Analysis and Discrete Mathematics, 1(1):72–91, 2007. ISSN 1452-8630 (print), 2406-100X (electronic). URL http://

pefmath.etf.rs/accepted/AADM-Vol1-No1-72-91.pdf.

Cantor:1962:LEF

[CET62] D. Cantor, G. Estrin, and R. Turn. Logarithmic and exponential function evaluation in a variable structure digital computer. IRE Transactions on Electronic Computers, EC-11(2):155–164, April 1962. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=5219348.

Carlson:2000:RTE

[CF00] B. C. Carlson and James FitzSimons. Reduction theorems for elliptic integrands with the square root of two quadratic factors. Journal of Computational and Applied Mathematics, 118(1–2): 71–85, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037704270000282X.

Carlson:1955:RAF

[CG55] Bengt Carlson and Max Goldstein. Rational approximations of functions. Report LA-1943, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, August 1955. iv + 46 pp. URL http://

www.osti.gov/scitech/biblio/4374577-0deJO9/;   http:
//www.osti.gov/scitech/servlets/purl/4374577.

Carlson:1985:AEF

[CG85a] B. C. Carlson and John L. Gustafson. Asymptotic expansion of the first elliptic integral. SIAM journal on mathematical analysis, 16(5):1072–1092, September 1985. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Conover:1985:AHS

[CG85b] B. Conover and D. L. Gustafson. An algorithm for high speed square roots. In IEEE Region 5 ’85 [IEE85], pages 19–21. LCCN TK 7801 N56 1985.

Cao:1989:ABS

[CG89] J.-D. Cao and H. H. Gonska. Approximation by Boolean sums of positive linear operators. II. Gopengauz-type estimates. Journal of Approximation Theory, 57(1):77–89, April 1989. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Carlson:1994:AAS

[CG94] B. C. Carlson and J. L. Gustafson. Asymptotic approximations for symmetric elliptic integrals. SIAM journal on mathematical analysis, 25(2):288–303, March 1994. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/22847.

Ceberio:2002:HRI

[CG02] M. Ceberio and L. Granvilliers. Horner’s rule for interval evaluation revisited. Computing: Archiv fur informatik und numerik, 69(1):51–81, March 2002. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http:

//link.springer.de/link/service/journals/00607/bibs/
2069001/20690051.htm;   http://link.springer.de/link/
service/journals/00607/papers/2069001/20690051.pdf.

Cameron:2024:AHM

[CG24] Thomas R. Cameron and Stef Graillat. Accurate Horner methods in real and complex floating-point arithmetic. BIT Numerical Mathematics, 64(2):??, June 2024. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s10543-024-01017-w.

Cheng:2005:SEEa

[CGKZ05a] Howard Cheng, Barry Gergel, Ethan Kim, and Eugene Zima. Space-efficient evaluation of hypergeometric series. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(2):41–52, June 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Cheng:2005:SEEb

[CGKZ05b] Howard Cheng, Barry Gergel, Ethan Kim, and Eugene Zima. Space-efficient evaluation of hypergeometric series. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 39(3):81–83, 2005. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). ISSAC 2005 poster abstract.

Cody:1967:CAN

[CH67] W. J. Cody and K. E. Hillstrom. Chebyshev approximations for the natural logarithm of the gamma function. Mathematics of Computation, 21(98):198–203, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cochran:1970:NTF

[CH70a] James Alan Cochran and Judith N. Hoffspiegel. Numerical techniques for finding ν-zeros of Hankel functions. Mathematics of Computation, 24(110):413–422, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1970:CAC

[CH70b] W. J. Cody and K. E. Hillstrom. Chebyshev approximations for the Coulomb phase shift. Mathematics of Computation, 24 (111):671–677, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chin:1978:DAD

[CH78] R. C. Y. Chin and G. W. Hedstrom. A dispersion analysis for difference schemes: tables of generalized Airy functions. Mathematics of Computation, 32(144):1163–1170, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1989:FCR

[CH89a] S.-G. Chen and P. Y. Hsieh. Fast computation of the N th root. Computers and Mathematics with Applications, 17(10):1423– 1427, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0898122189900242.

Chen:1989:FCTa

[CH89b] S.-G. Chen and P. Y. Hsieh. Fast computation of the N-th root. Computers and Mathematics with Applications, 17(10):1423– 1427, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0898122189900242.

Cornea-Hasegan:1998:PIC

[CH98] Marius Cornea-Hasegan. Proving the IEEE correctness of iterative floating-point square root, divide, and remainder algorithms. Intel Technology Journal, Q2(Q2):11, 1998. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/
q21998/articles/art_3.htm;   http://developer.intel.
com/technology/itj/q21998/pdf/ieee.pdf.

Char:1980:SCF

[Cha80] Bruce W. Char. On Stieltjes’ continued fraction for the gamma function. Mathematics of Computation, 34(150):547–551, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chambers:1982:UBF

[Cha82] Ll. G. Chambers. An upper bound for the first zero of Bessel functions. Mathematics of Computation, 38(158):589–591, April 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1970:CGI

[Che70] Reuven Chen. On the computation of the generalized integral in glow curve theory. Journal of Computational Physics, 6(2):314– 316, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Chen:1971:BAU

[Che71] Tien Chi Chen. Binary arithmetic unit implementing a multiplicative iteration for the exponential, logarithm, quotient and square root functions. United States Patent 3,631,230, De-

cember 28, 1971. URL http://www.freepatentsonline.com/
3631230.html.

Chen:1972:ACE

[Che72] Tien Chi Chen. Automatic computation of exponentials, logarithms, ratios and square roots. IBM Journal of Research and Development, 16(4):380–388, July 1972. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://

www.research.ibm.com/journal/rd/164/chen.pdf.

Chen:1981:AFD

[Che81] Gang Chen. An attempt to find the derivatives of elementary functions using the algorithmic language BCY. (Chinese). Zhejiang Daxue Xuebao, 3:141–149, 1981.

Chellali:1988:ACN

[Che88] Mustapha Chellali. Accélération de calcul de nombres de Bernoulli. (French) [Bernoulli number calculation acceleration]. Journal of Number Theory, 28(3):347–362, March 1988. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022314X88900479.

Chevillard:2009:FEE

[Che09] S. Chevillard. The functions ERF and ERFC computed with arbitrary precision. Report RRLIP2009-04, HAL, ????, 2009. 32 pp.

Chen:2011:SPF

[Che11] Chao-Ping Chen. Some properties of functions related to the gamma, psi and tetragamma functions. Computers and Mathematics with Applications, 62(9):3389–3395, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122111007267.

Chen:2013:CFE

[Che13a] Chao-Ping Chen. Continued fraction estimates for the psi function. Applied Mathematics and Computation, 219(19):9865– 9871, June 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300313003962.

Chen:2013:UTS

[Che13b] Chao-Ping Chen. Unified treatment of several asymptotic formulas for the gamma function. Numerical Algorithms, 64(2): 311–319, October 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9667-6.

Chen:2015:IAEb

[Che15a] Chao-Ping Chen. Inequalities and asymptotic expansions associated with the Ramanujan and Nemes formulas for the gamma function. Applied Mathematics and Computation, 261(??):337– 350, June 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300315004610.

Chen:2015:ICM

[Che15b] Chao-Ping Chen. Inequalities and completely monotonic functions associated with the ratios of functions resulting from the gamma function. Applied Mathematics and Computation, 259(??):790–799, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300315003148.

Chen:2016:IAEb

[Che16a] Chao-Ping Chen. Inequalities and asymptotic expansions for the psi function and the Euler–Mascheroni constant. Journal of Number Theory, 163(??):596–607, June 2016. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003558.

Chen:2016:IAEc

[Che16b] Chao-Ping Chen. Inequalities and asymptotics for the Euler– Mascheroni constant based on DeTemple’s result. Numerical Algorithms, 73(3):761–774, November 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-016-0116-9.

Chen:2016:MPI

[Che16c] Chao-Ping Chen. Monotonicity properties, inequalities and asymptotic expansions associated with the gamma function. Applied Mathematics and Computation, 283(??):385–396, June 20, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300316301515.

Chen:2016:MAA

[Che16d] Chao-Ping Chen. A more accurate approximation for the gamma function. Journal of Number Theory, 164(??):417–428, July 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X16000068.

Chen:2016:AEG

[Che16e] Chao-Ping Chen. On the asymptotic expansions of the gamma function related to the Nemes, Gosper and Burnside formulas. Applied Mathematics and Computation, 276(??):417–431, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300315016057.

Chen:2011:TSA

[CHG^{+}11] Jianxun Chen, Yongzhong Huang, Shaozhong Guo, Shimiao Chen, and Wei Wang. Test standardization and analyse model of mathematical functions for precision. In 2011 Third International Conference on Measuring Technology and Mechatronics Automation (ICMTMA), volume 3, pages 652–655. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 0-7695-4296-

4. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=5721571.

Cornea-Hasegan:1999:CPO

[CHGM99] M. A. Cornea-Hasegan, R. A. Golliver, and P. Markstein. Correctness proofs outline for Newton–Raphson based floatingpoint divide and square root algorithms. In Koren and Kornerup [KK99], pages 96–105. CODEN ???? ISBN 0-7803-5609-8, 0-7695-0116-8, 0-7695-0118-4. ISSN ???? LCCN

QA76.6 .S887 1999. URL http://euler.ecs.umass.edu/
paper/final/paper-121.pdf;   http://euler.ecs.umass.
edu/paper/final/paper-121.ps. IEEE Computer Society

Order Number PR00116. IEEE Order Plan Catalog Number 99CB36336.

Cornea:2003:DSR

[CHI^{+}03] M. Cornea, J. Harrison, C. Iordache, B. Norin, and S. Story. Division, square root and remainder algorithms for the Intel Itanium architecture. Report, Intel Corporation, Santa Clara, CA, USA, November 2003.

Chin:2020:PPW

[Chi20] Wooyoung Chin. A probabilistic proof of a Wallis-type formula for the gamma function. American Mathematical Monthly, 127 (1):75–79, 2020. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chlebus:2011:RSI

[Chl11] Edward Chlebus. A recursive scheme for improving the original rate of convergence to the Euler–Mascheroni constant. American Mathematical Monthly, 118(3):268–274, March 2011. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.118.03.268.pdf.

Choudhury:2014:SAA

[Cho14] Amit Choudhury. A simple approximation to the area under standard normal curve. Mathematics and Statistics, 2(3):147– 149, ???? 2014. ISSN 2332-2071 (print), 2332-2144 (electronic).

URL http://www.hrpub.org/journals/article_info.php?
aid=1470;   https://www.hrpub.org/download/20140305/
MS7-13401470.pdf.

Christiansen:1962:APC

[Chr62] Søren Christiansen. Algol programming: Contribution no. 3: Calculation of complementary Fresnel integrals. Nordisk tidskrift for informationsbehandling, 2(3):192–194, 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Christiansen:1965:APE

[Chr65] S. Christiansen. Algol programming: Error integral with complex argument. Nordisk tidskrift for informationsbehandling, 5(4):287–293, December 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=5&issue=4&spage=287.

Cody:1971:CAR

[CHT71] W. J. Cody, K. E. Hillstrom, and Henry C. Thatcher, Jr. Chebyshev approximations for the Riemann zeta function. Mathematics of Computation, 25(115):537–547, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chen:1994:ABB

[CIM94] Yang Chen, Mourad E. H. Ismail, and K. A. Muttalib. Asymptotics of basic Bessel functions and q-Laguerre polynomials. Journal of Computational and Applied Mathematics, 54 (3):263–272, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037704279200128V.

Chevillard:2009:CFC

[CJL09] Sylvain Chevillard, Mioara Joldes, and Christoph Lauter. Certified and fast computation of supremum norms of approximation errors. In Bruguera et al. [BCDH09], pages 169–176. ISBN 0-7695-3670-0, 1-4244-4329-6. ISSN 1063-6889. LCCN QA76.6

.S887 2009. URL http://www.ac.usc.es/arith19/.

Chevillard:2010:SED

[CJL10] Sylvain Chevillard, Mioara Jolde, and Christoph Lauter. Sollya: An environment for the development of numerical codes. In Fukuda et al. [FvdHJT10], pages 28–31. ISBN 3-642-15581-2 (paperback), 3-642-15582-0 (e-book). LCCN QA76.95 .I5654

2010. URL http://link.springer.com/book/10.1007/978-
3-642-15582-6.

Corless:1991:NEA

[CJR91] R. M. Corless, D. J. Jeffrey, and H. Rasmussen. Numerical evaluation of Airy functions with complex arguments. Journal of Computational Physics, 93(1):252–253, March 1991. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0021999191900894.

Corless:1992:NEAa

[CJR92a] R. M. Corless, D. J. Jeffrey, and H. Rasmussen. Numerical evaluation of Airy functions with complex arguments. Journal of Computational Physics, 98(2):347, February 1992. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/002199919290150W.

Corless:1992:NEAb

[CJR92b] R. M. Corless, D. J. Jeffrey, and H. Rasmussen. Numerical evaluation of Airy functions with complex arguments. Journal of Computational Physics, 99(1):106–114, March 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0021999192902798.

Conde:1979:Z

[CK79] S. Conde and Shyam L. Kalla. The ν-zeros of J_{−ν}(x). Mathematics of Computation, 33(145):423–426, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cartwright:1988:JTC

[CK88] Donald I. Cartwright and Krzysztof Kucharski. Jackson’s Theorem for compact connect Lie groups. Journal of Approximation Theory, 55(3):352–359, December 1988. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Campbell-Kelly:1989:WCB-3

[CK89a] Martin Campbell-Kelly, editor. The works of Charles Babbage, Vol. 3, The analytical engine and mechanical notation. William Pickering, London, UK, 1989. ISBN 1-85196-503-3, 1-85196-005-8 (set). 253 pp. LCCN ????

Corliss:1989:IIV

[CK89b] George Corliss and Gary Krenz. Indefinite integration with validation. ACM Transactions on Mathematical Software, 15(4): 375–393, December 1989. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/

pubs/citations/journals/toms/1989-15-4/p375-corliss/
; http://www.acm.org/pubs/toc/Abstracts/toms/76915.html.

Chen:1989:EMB

[CKL89] X. R. Chen, P. R. Krishnaiah, and W. W. Liang. Estimation of multivariate binary density using orthogonal functions. Journal of Multivariate Analysis, 31(2):178–186, November 1989. CO-DEN JMVAAI. ISSN 0047-259x (print), 1095-7243 (electronic).

Chakraborty:2007:VSF

[CKT07] Kalyan Chakraborty, Shigeru Kanemitsu, and Haruo Tsukada. Vistas of special functions II. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2007. ISBN 981-270-774-3. xii + 215 pp. LCCN QA351 .K35 2007.

Cheney:1961:TNA

[CL61] E. W. Cheney and H. L. Loeb. Two new algorithms for rational approximation. Numerische Mathematik, 3(1):72–75, December 1961. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Chen:1994:EDU

[CL94a] Sau-Gee Chen and Chieh-Chih Li. Efficient designs of unified 2’s complement division and square root algorithm and architecture. In Proceedings of TENCON ’94. IEEE Region 10’s Ninth Annual International Conference. Theme: ‘Frontiers of Computer Technology’, volume 2, pages 943–947. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. CODEN ???? ISSN ????

Cortadella:1994:HRD

[CL94b] J. Cortadella and T. Lang. High-radix division and square-root with speculation. IEEE Transactions on Computers, 43(8):919– 931, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:1995:UCA

[CL95] San-Gee Chen and Chieh-Chih Li. A unified cellular array for multiplication, division and square root. In IEEE Signal Processing Society Workshop on VLSI Signal Processing, VIII, 1995, pages 533–541. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN ???? ISSN ????

Chu:2000:CPT

[CL00] Wanming Chu and Yamin Li. Cost/performance tradeoff of nselect square root implementations. In ACAC 2000: 5th Australasian Computer Architecture Conference, pages 9–16. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. CODEN ???? ISSN ????

Cardoso:2011:IFP

[CL11] João Ribeiro Cardoso and Ana F. Loureiro. Iteration functions for p th roots of complex numbers. Numerical Algorithms, 57(3):329–356, July 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
1017-1398&volume=57&issue=3&spage=329.

Chen:2016:IAEa

[CL16] Chao-Ping Chen and Long Lin. Inequalities and asymptotic expansions for the gamma function related to Mortici’s formula. Journal of Number Theory, 162(??):578–588, May 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003133.

Clenshaw:1954:PAE

[Cle54] C. W. Clenshaw. Polynomial approximations to elementary functions. Mathematical Tables and Other Aids to Computation, 8(47):143–147, July 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Clenshaw:1955:NSC

[Cle55] C. W. Clenshaw. A note on the summation of Chebyshev series. Mathematical Tables and Other Aids to Computation, 9(51):118– 120, July 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Clemm:1969:ACV

[Cle69] Donald S. Clemm. Algorithm 352: Characteristic values and associated solutions of Mathieu’s differential equation [S22]. Communications of the ACM, 12(7):399–407, July 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Fri72].

Chiccoli:1988:EGE

[CLM88] C. Chiccoli, S. Lorenzutta, and G. Maino. On the evaluation of generalized exponential integrals E_{ν}(x). Journal of Computational Physics, 78(2):278–287, October 1988. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0021999188900502.

Clenshaw:1986:GEL

[CLOT86] C. W. Clenshaw, Daniel W. Lozier, F. W. J. Olver, and P. R. Turner. Generalized exponential and logarithmic functions. Computers and Mathematics with Applications, 12(5–6):1091– 1101, September/December 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Cornea:2020:ISC

[CLT20] Marius Cornea, Weiqiang Liu, and Arnaud Tisserand, editors. 2020 27th IEEE Symposium on Computer Arithmetic: ARITH 2020: proceedings: Portland, Oregon, USA, 7–10 June 2020.

IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2020. ISBN 1-72817-120-2, 1-

72817-121-0. LCCN ???? URL https://ieeexplore.ieee.
org/servlet/opac?punumber=9146973.

Croot:2004:ACC

[CLZ04] Ernie Croot, Ren-Cang Li, and H. J. Hui June Zhu. The abc conjecture and correctly rounded reciprocal square roots. Theoretical Computer Science, 315(2–3):405–417, May 6, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Coleman:1983:CEB

[CM83] J. P. Coleman and A. J. Monaghan. Chebyshev expansions for the Bessel function J_{n}(z) in the complex plane. Mathematics of Computation, 40(161):343–366, January 1983. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ciminiera:1990:HRS

[CM90] L. Ciminiera and P. Montuschi. Higher radix square rooting. IEEE Transactions on Computers, 39(10):1220–1231, October 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-

9956 (electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=59853.

Chen:2013:LIA

[CM13a] Chao-Ping Chen and Cristinel Mortici. Limits and inequalities associated with the Euler–Mascheroni constant. Applied Mathematics and Computation, 219(18):9755–9761, May 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300313003500.

Chevillard:2013:MPE

[CM13b] Sylvain Chevillard and Marc Mezzarobba. Multiple-precision evaluation of the Airy Ai function with reduced cancellation. In IEEE [IEE13], pages 175–182. ISBN 0-7695-4957-8. ISSN 1063-6889. LCCN QA76.9.C62 S95 2013.

Cody:1977:CRF

[CMF77] W. J. Cody, Rose M. Motley, and L. Wayne Fullerton. The computation of real fractional order Bessel functions of the second kind. ACM Transactions on Mathematical Software, 3(3):

232–239, September 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cody:1969:CRA

[CMV69] W. J. Cody, G. Meinardus, and R. S. Varga. Chebyshev rational approximations to e^{−x} on [0, ∞) and applications to heat conduction problems. Journal of Approximation Theory, 2(??): 50–65, ?? 1969. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Clenshaw:1963:ASF

[CMW63] C. W. Clenshaw, G. F. Miller, and M. Woodger. Algorithms for special functions I. Numerische Mathematik, 4:403–419, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Carlson:1981:AAI

[CN81] B. C. Carlson and Elaine M. Notis. Algorithm 577: Algorithms for incomplete elliptic integrals [S21]. ACM Transactions on Mathematical Software, 7(3):398–403, September 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Clenshaw:1980:UAE

[CO80] C. W. Clenshaw and Frank W. J. Olver. An unrestricted algorithm for the exponential function. SIAM Journal on Numerical Analysis, 17(2):310–331, April 1980. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Clenshaw:1986:UAR

[CO86] C. W. Clenshaw and F. W. J. Olver. Unrestricted algorithms for reciprocals and square roots. BIT (Nordisk tidskrift for informationsbehandling), 26(4):475–492, December 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0006-3835&volume=
26&issue=4&spage=475.

Cobb:1969:CAS

[Cob69] S. M. Cobb. Certification of Algorithm 47 [S16]: Associated Legendre functions of the first kind for real or imaginary arguments. Communications of the ACM, 12(11):635–636, November 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cochran:1965:ZHF

[Coc65] J. A. Cochran. The zeros of Hankel functions as functions of their order. Numerische Mathematik, 7(3):238–250, June 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Cody:1964:DPS

[Cod64] William J. Cody, Jr. Double-precision square root for the CDC-3600. Communications of the ACM, 7(12):715–718, December 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cody:1965:CAC

[Cod65a] W. J. Cody, Jr. Chebyshev approximations for the complete elliptic integrals K and E. Mathematics of Computation, 19(89– 92):105–112, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.jstor.org/ stable/2004103. See also [Cod66].

Cody:1965:CPE

[Cod65b] W. J. Cody, Jr. Chebyshev polynomial expansions of complete elliptic integrals. Mathematics of Computation, 19(89–92):249– 259, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL http://www.jstor.org/stable/
2003350.

Cody:1966:CCA

[Cod66] W. J. Cody, Jr. Corrigenda: “Chebyshev approximations for the complete elliptic integrals K and E”. Mathematics of Computation, 20(93):207, January 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Cod65a].

Cody:1967:LEA

[Cod67] William J. Cody, Jr. Letter to the Editor: Another aspect of economical polynomials. Communications of the ACM, 10(9):531, September 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Fik67b].

Cody:1969:RCA

[Cod69] W. J. Cody. Rational Chebyshev approximations for the error function. Mathematics of Computation, 23(107):631–637, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1971:SEF

[Cod71] W. J. Cody. Software for the elementary functions. In Mathematical Software [Ric71], pages 171–186. ISBN 0-12-587250-X. LCCN QA1 .M26. Based on the proceedings of the Mathematical Software Symposium held at Purdue University, Lafayette, Indiana, USA, April 1–3, 1970.

Cody:1975:FPS

[Cod75] W. J. Cody. The FUNPACK package of special function subroutines. ACM Transactions on Mathematical Software, 1(1):13–25, March 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cody:1982:TTP

[Cod82] W. J. Cody. Transportable test procedures for elementary function software. In Mulvey [Mul82], pages 236–247. ISBN 0-387-11495-5. LCCN QA402.5 .E94 1982.

Cody:1983:ASM

[Cod83] W. J. Cody. Algorithm 597: Sequence of modified Bessel functions of the first kind. ACM Transactions on Mathematical Software, 9(2):242–245, June 1983. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cody:1990:PEP

[Cod90] W. J. Cody. Performance evaluation of programs for the error and complementary error functions. ACM Transactions on Mathematical Software, 16(1):29–37, March 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295

(electronic). URL http://www.acm.org/pubs/citations/
journals/toms/1990-16-1/p29-cody/;   http://www.acm.
org/pubs/toc/Abstracts/0098-3500/77628.html.

Cody:1991:CPT

[Cod91] W. J. Cody. CELEFUNT: a portable test package for complex elementary functions. Technical Report ANL-91/1, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, January 1991. iii + 21 pp.

Cody:1993:ACP

[Cod93a] W. J. Cody. Algorithm 714: CELEFUNT: a portable test package for complex elementary functions. ACM Transactions on Mathematical Software, 19(1):1–21, March 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/pubs/citations/journals/toms/
1993-19-1/p1-cody/;   http://www.acm.org/pubs/toc/Abstracts/
toms/151272.html.

Cody:1993:ASP

[Cod93b] W. J. Cody, Jr. Algorithm 715: SPECFUN: a portable FOR-TRAN package of special function routines and test drivers. ACM Transactions on Mathematical Software, 19(1):22–32, March 1993. CODEN ACMSCU. ISSN 0098-3500 (print),

1557-7295 (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/0098-3500/151273.html.

Cohl:2012:TEF

[Coh12] Howard S. Cohl. Table errata to “Formulas and theorems for the special functions of mathematical physics” by W. Magnus, F. Oberhettinger & R. P. Soni (1966). Mathematics of Computation, 81(280):2251, October 2012. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/journals/mcom/2012-81-
280/S0025-5718-2012-02612-3;   http://www.ams.org/journals/
mcom/2012-81-280/S0025-5718-2012-02612-3/S0025-5718-
2012-02612-3.pdf. See [MOS66].

Coleman:1978:RSN

[Col78] John P. Coleman. Remark on “Algorithm 49: Spherical Neumann Function”. ACM Transactions on Mathematical Software, 4(3):295, September 1978. CODEN ACMSCU. ISSN 0098-3500. See [Her61a].

Coleman:1980:FSB

[Col80] J. P. Coleman. A Fortran subroutine for the Bessel function J_{n}(x) of order 0 to 10. Computer Physics Communications, 21(1):109–118, December 1, 1980. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465580900806.

Coleman:1984:FSB

[Col84] J. P. Coleman. A Fortran subroutine for the Bessel function J_{n}(x) of order 0 to 10. Computer Physics Communications, 35(1–3):C–654, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584828143.

Corrington:1961:ACE

[Cor61] Murlan S. Corrington. Applications of the complex exponential integral. Mathematics of Computation, 15(73):1–6, January 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cowgill:1964:LEB

[Cow64] D. Cowgill. Logic equations for a built-in square root method. IEEE Transactions on Electronic Computers, EC-13(2):156–157, April 1964. CODEN IEECA8. ISSN 0367-

7508. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=4038119.

Cowell:1977:PMS

[Cow77] Wayne Cowell, editor. Portability of Numerical Software Workshop, Oak Brook, Illinois, June 21–23, 1976, volume 57 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-08446-0. LCCN QA297 .W65 1976.

Clenshaw:1966:CSB

[CP66] C. W. Clenshaw and Susan M. Picken. Chebyshev series for Bessel functions of fractional order, volume 8 of National Physical Laboratory. Mathematical tables. Her Majesty’s Stationery Office, London, UK, 1966. iii + 53 pp.

Cody:1970:RAC

[CP70] W. J. Cody and Kathleen A. Paciorek. Remark on Algorithm 292 [S22]: Regular Coulomb wave functions. Communications of the ACM, 13(9):573, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cole:1979:EI

[CP79] R. J. Cole and C. Pescatore. Evaluation of the integral ∫_{0}^{∞} t^{n} exp(−t^{2} − x/t) dt. Journal of Computational Physics, 32(2):280–287, August 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179901359.

Carsky:1998:IGF

[CP98] Petr Cársky and Martin Polásek. Incomplete gamma F_{m}(x) functions for real negative and complex arguments. Journal of Computational Physics, 143(1):259–265, June 10, 1998. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0021999198959757.

Corsonello:1999:HPS

[CP99] P. Corsonello and S. Perri. High performance square rooting circuit using hybrid radix-2 adders. Electronics Letters, 35(3): 185–186, February 4, 1999. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Chen:2015:IAEa

[CP15] Chao-Ping Chen and Richard B. Paris. Inequalities, asymptotic expansions and completely monotonic functions related to the gamma function. Applied Mathematics and Computation, 250(??):514–529, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S009630031401515X.

Campos-Pinto:2019:APP

[CPCD19] Martin Campos-Pinto, Frédérique Charles, and Bruno Després. Algorithms for positive polynomial approximation. SIAM Journal on Numerical Analysis, 57(1):148–172, ???? 2019. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Cote:2012:CTL

[CPG12] F. D. Côté, I. N. Psaromiligkos, and W. J. Gross. A Chernofftype lower bound for the Gaussian Q-function. arxiv.org, ??(??):

??, 2012. URL https://arxiv.org/abs/1202.6483.

Cody:1970:CAD

[CPT70] W. J. Cody, Kathleen A. Paciorek, and Henry C. Thacher, Jr. Chebyshev approximations for Dawson’s integral. Mathematics of Computation, 24(109):171–178, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cuyt:2008:HCF

[CPV^{+}08] Annie Cuyt, Vigdis B. Petersen, Brigitte Verdonk, Haakon Waadeland, and William B. Jones. Handbook of Continued Fractions for Special Functions. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-6948-0. xx + 440 pp. LCCN QA295 .H275 2008.

Chaudhry:2004:EHC

[CQSP04] M. Aslam Chaudhry, Asghar Qadir, H. M. Srivastava, and R. B. Paris. Extended hypergeometric and confluent hypergeometric functions. Applied Mathematics and Computation, 159(2):589– 602, December 6, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chiarella:1968:EIR

[CR68] C. Chiarella and A. Reichel. On the evaluation of integrals related to the error function. Mathematics of Computation, 22 (101):137–143, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Chatterjee:2008:CNT

[CR08] S. Chatterjee and D. Roy. A class of new transforms tailored for the hypergeometric series. Computer Physics Communications, 179(8):555–561, October 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Crandall:2006:NFP

[Cra06] Richard E. Crandall. Note on fast polylogarithm computation. Report, Reed College, Portland, OR, USA,

January 2006. 6 pp. URL http://people.reed.edu/
~crandall/papers/Polylog.pdf;   https://web.archive.
org/web/20120916145721/http://people.reed.edu/~crandall/
papers/Polylog.pdf.

Crandall:2012:UAP

[Cra12a] R. E. Crandall. Unified algorithms for polylogarithm, L-series,

and zeta variants, 2012. URL http://www.perfscipress.com/
papers/universalTOC25.pdf;   https://web.archive.org/
web/20130430193005/http://www.perfscipress.com/papers/

universalTOC25.pdf. Published in [Cra12b].

Crandall:2012:ARS

[Cra12b] Richard E. Crandall. Algorithmic Reflections: Selected Works. PSI Press, Portland, OR, USA, first perfectly scientific press paperback edition, 2012. ISBN 1-935638-19-X. 410 pp. LCCN QA958 .C736 2012.

Crenshaw:1991:SRS

[Cre91] J. W. Crenshaw. Square roots are simple? Embedded Systems Programming, 4(11):30–52, November 1991. CODEN EYPRE4. ISSN 1040-3272.

Crenshaw:1998:ISR

[Cre98] Jack W. Crenshaw. Integer square roots. Embedded Systems Programming, 11(2):15–32, February 1998. CODEN

EYPRE4. ISSN 1040-3272. URL http://www.embedded.com/
98/9802fe2.htm.

Crick:1982:IRB

[Cri82] S. E. Crick. Inclusion relations for Bernstein quasi-analytic classes. Journal of Approximation Theory, 34(4):375–379, April 1982. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Critchfield:1989:CEF

[Cri89] Charles L. Critchfield. Computation of elliptic functions. Journal of Mathematical Physics, 30(2):295–297, February 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i2/p295_s1.

Croft:1992:ACA

[Cro92] A. Croft. An application of convergence acceleration techniques to a class of two-point boundary value problems on a semiinfinite domain. Numerical Algorithms, 2(3–4):307–320, September 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Cohen:2000:CAA

[CRZ00] Henri Cohen, Fernando Rodriguez Villegas, and Don Zagier. Convergence acceleration of alternating series. Experimental mathematics, 9(1):3–12, 2000. ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://projecteuclid.org/euclid.
em/1046889587;   http://www.math.u-bordeaux.fr/~cohen/
sumalt2new.ps.

Cruz:1982:ZHF

[CS82] Andrés Cruz and Javier Sesma. Zeros of the Hankel function of real order and of its derivative. Mathematics of Computation, 39 (160):639–645, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cruz:1983:MPR

[CS83] Andrés Cruz and Javier Sesma. Modulus and phase of the reduced logarithmic derivative of the Hankel function. Mathematics of Computation, 41(164):597–605, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Choi:2011:AFT

[CS11] Junesang Choi and H. M. Srivastava. Asymptotic formulas for the triple Gamma function Γ_{3} by means of its integral representation. Applied Mathematics and Computation, 218(6): 2631–2640, November 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300311010289.

Cardoso:2019:CMG

[CS19] João R. Cardoso and Amir Sadeghi. Computation of matrix gamma function. BIT Numerical Mathematics, 59(2):343– 370, June 2019. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://link.springer.

com/article/10.1007/s10543-018-00744-1.

Cody:1973:CAP

[CST73] W. J. Cody, Anthony J. Strecok, and Henry C. Thacher, Jr. Chebyshev approximations for the psi function. Mathematics of Computation, 27(121):123–127, January 1973. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ceretani:2018:AME

[CST18] Andrea N. Ceretani, Natalia N. Salva, and Domingo A. Tarzia. Approximation of the modified error function. Applied Mathematics and Computation, 337(??):591–606, November 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300318304715.

Cody:1967:CRC

[CT67a] W. J. Cody and Henry C. Thacher, Jr. Corrigendum: “Rational Chebyshev approximations for Fermi–Dirac integrals of orders −1/2, 1/2, and 3/2”. Mathematics of Computation, 21(99): 525, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL http://www.jstor.org/stable/

2003289. See [CT67b].

Cody:1967:RCA

[CT67b] W. J. Cody and Henry C. Thacher, Jr. Rational Chebyshev approximations for Fermi–Dirac integrals of orders −1/2, 1/2, and 3/2. Mathematics of Computation, 21(97):30–40, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.jstor.org/stable/2003468. See

also [CT67a].

Cody:1968:RCAa

[CT68] W. J. Cody and H. C. Thacher, Jr. Rational Chebyshev approximations for the exponential integral E_{1}(x). Mathematics of Computation, 22(103):641–649, July 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cody:1969:CAE

[CT69] W. J. Cody and Henry C. Thacher, Jr. Chebyshev approximations for the exponential integral Ei(x). Mathematics of Computation, 23(106):289–303, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cover:1988:DII

[CT88] Thomas M. Cover and Joy A. Thomas. Determinant inequalities via information theory. SIAM Journal on Matrix Analysis and Applications, 9(3):384–392, July 1988. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Chen:2016:SIAa

[CT16] Chao-Ping Chen and Wei-Wei Tong. Sharp inequalities and asymptotic expansions for the gamma function. Journal of Number Theory, 160(??):418–431, March 2016. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0022314X15003200.

Cojean:2022:GML

[CTA22] Terry Cojean, Yu-Hsiang Mike Tsai, and Hartwig Anzt. Ginkgo — a math library designed for platform portability. Parallel Computing, 111(??):??, July 2022. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167819122000096.

Chaudhry:1996:ACF

[CTV96] M. A. Chaudhry, N. M. Temme, and E. J. M. Veling. Asymptotics and closed form of a generalized incomplete gamma function. Journal of Computational and Applied Mathematics, 67(2):371–379, March 29, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0377042795000186.

Corbato:1959:GSB

[CU59] Fernando J. Corbató and Jack L. Uretsky. Generation of spherical Bessel functions in digital computers. Journal of the Association for Computing Machinery, 6(3):366–375, July 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cundiff:1962:AEA

[Cun62a] John L. Cundiff. Algorithm 88: Evaluation of asymptotic expression for the Fresnel sine and cosine integrals. Communications of the ACM, 5(5):280, May 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cundiff:1962:AEFa

[Cun62b] John L. Cundiff. Algorithm 89: Evaluation of the Fresnel sine integral. Communications of the ACM, 5(5):280, May 1962. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cundiff:1962:AEFb

[Cun62c] John L. Cundiff. Algorithm 90: Evaluation of the Fresnel cosine integral. Communications of the ACM, 5(5):281, May 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Curtiss:1964:EIB

[Cur64] C. F. Curtiss. Expansions of integrals of Bessel functions of large order. Journal of Mathematical Physics, 5(4):561–564, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v5/i4/p561_s1.

Cusick:1983:CCL

[Cus83] David Cusick. Computers & calculators: a logarithm algorithm for four-function calculators. Two-Year College Mathematics Journal, 14(4):322–324, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972706;   https://www.jstor.org/stable/3027283.

Cuyt:1994:NNM

[Cuy94] Annie Cuyt, editor. Nonlinear numerical methods and rational approximation II, volume 296 of Mathematics and its applications. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1994. ISBN 0-7923-2967-8. LCCN QA297 .N642 1994. Proceedings of an international conference held at the University of Antwerp, Belgium, Sept. 5–11, 1993.

Coussement:2003:AMO

[CV03] Els Coussement and Walter Van Assche. Asymptotics of multiple orthogonal polynomials associated with the modified Bessel functions of the first kind. Journal of Computational and Applied Mathematics, 153(1–2):141–149, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042702005964.

Combet:1965:CBT

[CVV65] M. Combet, H. Van Zonneveld, and L. Verbeek. Computation of the base two logarithm of binary numbers. IEEE Transactions on Electronic Computers, EC-14(6):863–867, December 1965. CODEN IEECA8. ISSN 0367-7508. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=4038605.

Cuyt:2006:ERM

[CVW06] Annie Cuyt, Brigitte Verdonk, and Haakon Waadeland. Efficient and reliable multiprecision implementation of elementary and special functions. SIAM Journal on Scientific Computing, 28 (4):1437–1462, January 2006. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Cody:1980:SME

[CW80] William J. Cody, Jr. and William Waite. Software Manual for the Elementary Functions. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1980. ISBN 0-13-822064-6. x + 269 pp. LCCN QA331 .C635 1980.

Chen:2018:NEH

[CY18] Ruyun Chen and Gang Yang. Numerical evaluation of highly oscillatory Bessel transforms. Journal of Computational and Applied Mathematics, 342(??):16–24, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042718301894.

Cyvin:1964:AND

[Cyv64] S. J. Cyvin. Algorithm 226: Normal distribution function. Communications of the ACM, 7(5):295, May 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Chaudhry:1994:GIG

[CZ94] M. Aslam Chaudhry and S. M. Zubair. Generalized incomplete gamma functions with applications. Journal of Computational and Applied Mathematics, 55(1):99–123, October 31, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042794901872.

Chaudhry:1995:DGI

[CZ95] M. Aslam Chaudhry and S. M. Zubair. On the decomposition of generalized incomplete gamma functions with applications to Fourier transforms. Journal of Computational and Applied Mathematics, 59(3):253–284, May 30, 1995. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/037704279400026W.

Celledoni:2010:AFF

[CZ10] Elena Celledoni and Antonella Zanna. Algorithm 903: FRB — Fortran routines for the exact computation of free rigid body motions. ACM Transactions on Mathematical Software, 37(2): 23:1–23:24, April 2010. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Darley:1989:FPI

[D^{+}89] H. M. Darley et al. Floating point/integer processor with divide and square root functions, 1989. U.S. Patent No. 4,878,190.

Damnjanovic:1994:EFL

[Dam94] Zlatan Damnjanovic. Elementary functions and loop programs. Notre Dame Journal of Formal Logic, 35(4):496–522, 1994. CO-DEN NDJFAM. ISSN 0029-4527 (print), 1939-0726 (electronic).

Danilcenko:1977:ETC

[Dan77] L. S. Danil’čenko. An efficient technique for the construction of rational approximations of elementary functions. (Russian) optimization of computations (approximation and minimization of functions). Akad. Nauk Ukrain. SSR Inst. Kibernet. Preprint, 18:17–21, 1977.

Danielopoulos:1982:CEP

[Dan82] S. D. Danielopoulos. On the cost of evaluating polynomials and their derivatives. Computing: Archiv fur informatik und numerik, 29(4):373–380, 1982. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Darlington:1970:AAA

[Dar70] Sidney Darlington. Analytical approximations to approximations in the Chebyshev sense. The Bell System Technical Journal, 49(1):1–32, January 1970. CODEN BSTJAN. ISSN

0005-8580. URL http://bstj.bell-labs.com/BSTJ/images/
Vol49/bstj49-1-1.pdf.

Das:1994:ATE

[Das94] Mrinal Kanti Das. Analysis of two elementary functions. International journal of mathematical education in science and technology, 25(1):17–24, 1994. CODEN IJMEBM. ISSN 0020-739X (print), 1464-5211 (electronic).

Davis:1935:EPF

[Dav35] H. T. Davis. An extension to polygamma functions of a theorem of Gauss. Bulletin of the American Mathematical Society, 41 (4):243–248, April 1935. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Davis:1959:LEI

[Dav59] Philip J. Davis. Leonhard Euler’s integral: a historical profile of the gamma function: In memoriam: Milton Abramowitz. American Mathematical Monthly, 66(10):849–869, December 1959.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2309786.

Davis:1964:GFR

[Dav64] Philip J. Davis. Gamma function and related functions. In Abramowitz and Stegun [AS64], pages 253–294. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also avail-

able online at http://www.convertit.com/Go/ConvertIt/

Reference/AMS55.ASP in bitmap image format.

Davis:1981:EFA

[Dav81] Diane F. Davis. Elementary functions on an array processor. In IEEE [IEE81a], pages 170–178. LCCN QA76.9.C62 S95 1981. IEEE Catalog No. 81CH1630-3. Computer Society Order No. 347.

Davenport:1988:ICF

[Dav88] J. H. Davenport. Integration in closed form. In N. M. Stephens and M. P. Thorne, editors, Computers in mathematical research: based on the proceedings of a conference organized by the Institute of Mathematics and its Applications on computers in mathematical research, held at University College, Cardiff in September 1986, volume 14 of Institute of Mathematics and its applications conference series, new series. Clarendon Press, Oxford, UK, 1988. ISBN 0-19-853620-8. LCCN QA11.A1 C618 1986. US$57.50.

DeSchrijver:2012:DPRa

[DAV12a] Steven K. De Schrijver, El-Houssaine Aghezzaf, and Hendrik Vanmaele. Double precision rational approximation algorithm for the inverse standard normal first order loss function. Applied Mathematics and Computation, 219(3):1375–1382, October 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300312007114.

DeSchrijver:2012:DPRb

[DAV12b] Steven K. De Schrijver, El-Houssaine Aghezzaf, and Hendrik Vanmaele. Double precision rational approximation algorithms for the standard normal first and second order loss functions. Applied Mathematics and Computation, 219(4):2320–2330, November 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300312008041.

Dawson:1897:NV

[Daw97] H. G. Dawson. On the numerical value of ∫_{0}^{h} e^{x^{2}} dx. Proceedings of the London Mathematical Society. First Series, s1-29(1): 519–522, November 1897. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (electronic).

Develi:2013:HAA

[DB13] I. Develi and A. Basturk. Highly accurate analytic approximation to the Gaussian Q-function based on the use of nonlinear least squares optimization algorithm. Journal of Optimization Theory and Applications, 159(1):183–191, October 01, 2013. CODEN JOTABN. ISSN 1573-2878. URL https://link.

springer.com/article/10.1007/s10957-012-0217-0.

Drachman:1981:TTH

[DC81] B. Drachman and C. I. Chuang. A table of two hundred zeros of the derivative of the modified Bessel function K_{n}(z) and a graph of their distribution. Journal of Computational and Applied Mathematics, 7(3):167–171, September 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0771050X81900140.

Dattoli:1992:GFM

[DCL^{+}92] G. Dattoli, C. Chiccoli, S. Lorenzutta, G. Maino, M. Richetta, and A. Torre. Generating functions of multivariable generalized Bessel functions and Jacobi-elliptic functions. Journal of Mathematical Physics, 33(1):25–36, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v33/i1/p25_s1.

Davies:1976:IPS

[DD76] M. Davies and B. Dawson. The incrementation parameter in square root iteration. Journal of the Institute of Mathematics and its Applications, 17(2):219–223, 1976. CODEN JMTAA8. ISSN 0020-2932.

Dyer:2007:AEF

[DD07] Stephen Dyer and Justin Dyer. Approximations to error functions. IEEE Instrumentation & Measurement Magazine, 10(6):

45–48, December 2007. URL http://ieeexplore.ieee.org/
abstract/document/4428581/.

Dyer:2008:CCI

[DD08] J. S. Dyer and S. A. Dyer. Corrections to, and comments on, “An improved approximation for the Gaussian Q-Function”. IEEE Communications Letters, 12(4):231, April 2008. CO-DEN ICLEF6. ISSN 1089-7798 (print), 1558-2558 (electronic).

URL http://ieeexplore.ieee.org/document/4489650/. See

[KL07].

deDinechin:2005:TPU

[dDEG05] Florent de Dinechin, Alexey Ershov, and Nicolas Gast. Towards the post-ultimate libm. In IEEE [IEE05], page ?? ISBN ????

LCCN ???? URL http://arith17.polito.it/final/paper-
165.pdf.

deDinechin:2015:HIF

[dDI15] Florent de Dinechin and Matei Istoan. Hardware implementations of fixed-point Atan2. In Muller et al. [MTV15], pages 34–41. ISBN 1-4799-8665-8, 1-4799-8663-1. ISSN 1063-6889. LCCN QA76.9.C62 S95 2015. URL http://ieeexplore.ieee.

org/servlet/opac?punumber=7193754.

deDinechin:2013:FPT

[dDIS13] Florent de Dinechin, Matei Istoan, and Guillaume Sergent. Fixed-point trigonometric functions on FPGAs. ACM SIGARCH Computer Architecture News, 41(5):83–88, December 2013. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

deDinechin:2004:PCR

[dDLM04] Florent de Dinechin, Catherine Loirat, and Jean-Michel Muller. A proven correctly rounded logarithm in double-precision. In Frougny et al. [FBM04], pages 71–85. ISSN 0944-

0488. URL http://www.informatik.uni-trier.de/Reports/
TR-08-2004/rnc6_07_dinechin.pdf. Forschungsbericht Nr.

04-8.

deDinechin:2011:CFP

[dDLM11] Florent de Dinechin, Christoph Lauter, and Guillaume Melquiond. Certifying the floating-point implementation of an elementary function using Gappa. IEEE Transactions on Computers, 60 (2):242–253, February 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

deDinechin:2006:STP

[dDM06] Florent de Dinechin and Sergey Maidanov. Software techniques for perfect elementary functions in floating-point interval arithmetic. In Anonymous [Ano06], page ?? ISBN ???? LCCN ????

deDinechin:2010:FPE

[dDP10] Florent de Dinechin and Bogdan Pasca. Floating-point exponential functions for DSP-enabled FPGAs. In Jinian Bian, Qiang Zhou, and Kang Zhao, editors, Proceedings 2010 International Conference on Field-Programmable Technology, 8–10 December 2010, Beijing, China, pages 110–117. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, December 2010.

deAPMartins:1977:DSB

[de 77] Pedro de A.P.Martins. Determination of spherical Bessel functions of an order larger than the argument. Journal of Computational Physics, 25(2):194–198, October 1977. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0021999177900213.

DeDoelder:1991:SSC

[De 91] P. J. De Doelder. On some series containing ψ(x) − ψ(y >) and (ψ(x) − ψ(y >))^{2} for certain values of x and y. Journal of Computational and Applied Mathematics, 37(1–3):125–141, November 18, 1991. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037704279190112W.

Dederick:1926:QDDc

[Ded26] L. S. Dederick. Questions and discussions: Discussions: a modified method for cube roots and fifth roots. American Mathematical Monthly, 33(9):469–472, November 1926. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/2299613.

Dehling:1989:FLI

[Deh89] Herold Dehling. The functional law of the iterated logarithm for von Mises functionals and multiple Wiener integrals. Journal of Multivariate Analysis, 28(2):177–189, February 1989. CODEN JMVAAI. ISSN 0047-259x (print), 1095-7243 (electronic).

DeLugish:1970:CAA

[DeL70] Bruce Gene DeLugish. A class of algorithms for automatic evaluation of certain elementary functions in a binary computer. Technical Report 399, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, Illinois, 1970. 191 pp.

Delic:1979:CSS

[Del79] G. Delic. Chebyshev series for the spherical Bessel function j_{l}(r). Computer Physics Communications, 18(1):73–86, September 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465579900250.

Delahaye:1981:ACS

[Del81] J.-P. Delahaye. Accélération de la convergence des suites dont le rapport des erreurs est borné. (French) [Convergence acceleration for sequences with bounded error ratios]. Calcolo, 18 (2):1–116, 1981. CODEN CDABAE. ISSN 0008-0624 (print), 1126-5434 (electronic).

Delic:1984:CSS

[Del84] G. Delic. Chebyshev series for the spherical Bessel function j_{l}(r). Computer Physics Communications, 35(1–3):C–577, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584827514.

Demirbas:1989:MSE

[Dem89] K. Demirbas. Multidimensional state estimation using stacks for dynamic systems with interference. Automatica: the journal of IFAC, the International Federation of Automatic Control, 25(4): 617–621, July 1989. CODEN ATCAA9. ISSN 0005-1098 (print), 1873-2836 (electronic).

Derenzo:1977:AHC

[Der77] Stephen E. Derenzo. Approximations for hand calculators using small integer coefficients. Mathematics of Computation, 31(137): 214–222, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Deuflhard:1976:ASC

[Deu76] P. Deuflhard. On algorithms for the summation of certain special functions. Computing: Archiv fur informatik und numerik, 17 (1):37–48, March 1976. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Develi:2012:NAB

[Dev12] I. Develi. A new approximation based on the differential evolution algorithm for the Gaussian Q-function. Int. J. Innov. Comput. Inf. Control, 8(10(B)):7095–7102, ????

2012. URL http://link.springer.com/content/pdf/10.
1007/s10957-012-0217-0.pdf;   http://www.ijicic.org/
ijicic-11-08039.pdf.

Dzjadyk:1978:CRP

[DF78] V. K. Dzjadyk and L. Ī. Fílozof. The convergence rate of Padé approximants for some elementary functions. (Russian). Mat. Sb. (N.S.), 107(149)(3):347–363, 463, 1978. ISSN 0368-8666.

Dhanoa:1984:BPE

[DF84] M. S. Dhanoa and J. France. A BASIC program for the evaluation of the gamma functions. Journal of Applied Statistics, 11(2):225–228, 1984. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

DiDonato:1999:TFC

[DG99] Armido R. DiDonato and Russ Gnoffo. Testing a Fortran 90 compiler using the NSWC Fortran 77 Mathematics Library. Technical Report NSWCDD/TR-98/75, Naval Surface Warfare Center, Dahlgren, VA 22448-5100, USA, February 1999. v + 64

pp. URL https://apps.dtic.mil/sti/pdfs/ADA360604.pdf.

DelPunta:2018:LRC

[DGA18] Jessica A. Del Punta, Gustavo Gasaneo, and Lorenzo U. Ancarani. On the Laguerre representation of Coulomb functions and the relation to orthogonal polynomials. Advances in Quantum Chemistry, 76:79–101, 2018. CODEN AQCHA9. ISSN

0065-3276. URL https://www.sciencedirect.com/science/
article/pii/S0065327617300643.

Delahaye:1980:RNA

[DGB80] J. P. Delahaye and B. Germain-Bonne. Résultats négatifs en accélération de la convergence. (French) [Negative results in convergence acceleration]. Numerische Mathematik, 35(4):443–457, November 1980. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Delahaye:1982:SLC

[DGB82] J. P. Delahaye and B. Germain-Bonne. The set of logarithmically convergent sequences cannot be accelerated. SIAM Journal on Numerical Analysis, 19(4):840–844, August 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dahlquist:1965:CAP

[DGS65] Germund Dahlquist, Sven-Åke Gustafson, and Károly Siklósi. Convergence acceleration from the point of view of linear programming. Nordisk tidskrift for informationsbehandling, 5 (1):1–16, March 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=5&issue=1&spage=1.

Dunster:2018:UAE

[DGS18] T. M. Dunster, A. Gil, and J. Segura. Uniform asymptotic expansions for Laguerre polynomials and related confluent hypergeometric functions. Advances in Computational Mathematics, 44(5):1441–1474, October 2018. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL https://link.

springer.com/article/10.1007/s10444-018-9589-5.

Dunster:2024:CPC

[DGS24] T. M. Dunster, A. Gil, and J. Segura. Computation of parabolic cylinder functions having complex argument. Applied Numerical Mathematics: Transactions of IMACS, 197(??):230–242, March 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0168927423002969.

DiDonato:1959:NFC

[DH59] A. R. DiDonato and A. V. Hershey. New formulas for computing incomplete elliptic integrals of the first and second kind. Journal of the Association for Computing Machinery, 6(4):515– 526, October 1959. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

DiDonato:1976:CIG

[DH76] Armido R. DiDonato and R. K. Hageman. Computation of the incomplete gamma function ratios. Report NSWC/DL TR-3482, Naval Surface Weapons Center, Dahlgren, VA, USA, April

1976. 86 pp. URL https://apps.dtic.mil/sti/citations/
tr/ADA031812.

DiDonato:1977:CPP

[DH77] Armido R. DiDonato and R. K. Hageman. Computation of the percentage points of the chi-square distribution. Report NSWC/DL TR-3569, Naval Surface Weapons Center, Dahlgren,

VA, USA, 1977. 200 pp. URL https://apps.dtic.mil/sti/
pdfs/ADA043997.pdf.

Defour:2004:PSM

[DHL^{+}04] David Defour, Guillaume Hanrot, Vincent Lefèvre, Jean-Michel Muller, Nathalie Revol, and Paul Zimmermann. Proposal for a standardization of mathematical function implementation in floating-point arithmetic. Numerical Algorithms, 37(1–4): 367–375, December 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp009.

kluweronline.com/IPS/content/ext/x/J/5058/I/58/A/30/
abstract.htm; http://perso.ens-lyon.fr/jean-michel.muller/
NumAlg04.pdf;   http://www.loria.fr/~zimmerma/papers/
PropStandFunctions.pdf.

Duprat:1989:SRA

[DHM89] J. Duprat, Y. Herreros, and J.-M. Muller. Some results about online computation of functions. In Ercegovac and Swartzlander, Jr. [ES89], pages 112–118. ISBN 0-8186-8963-3 (case), 0-8186-5963-7 (microfiche). LCCN QA 76.9 C62 S95 1989. IEEE catalog no. 89CH2757-3.

DiMarzio:1986:IPA

[Di 86] F. Di Marzio. An improved procedure for the accurate evaluation of polygamma functions with integer and half-integer argument. Computer Physics Communications, 39(3):343–345, April 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465586900950.

DiDonato:1978:AR

[DiD78] A. R. DiDonato. An approximation for ∫_{x}^{∞} e^{−t^{2}/2}t^{p}dt, x > 0, p real. Mathematics of Computation, 32(141):271–275, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DiDonato:1984:IGF

[DiD84] Armido R. DiDonato. The incomplete gamma function ratios using Temme’s asymptotic expansions. Report NSWC TR 84-79, Naval Surface Weapons Center, Dahlgren, VA, USA, 1984.

DiDonato:1990:SDC

[DiD90] Armido R. DiDonato. Significant digit computation of the elliptical coverage function. Technical Report NAVSWC TR 90-513, Naval Surface Warfare Center, Dahlgren, VA 22448-5000, USA and Silver Spring, MD 20903-5000, USA, September 1990. v +

13 + A-7 + B-9 + 5 pp. URL http://www.dtic.mil/dtic/tr/
fulltext/u2/a230523.pdf.

Dieudonne:1980:SFL

[Die80] Jean Dieudonné, editor. Special functions and linear representations of Lie groups, volume 42 of Regional conference series in mathematics. Published for the Conference Board of the Mathematical Sciences by the American Mathematical Society, Providence, RI, USA, 1980. ISBN 0-8218-1692-6. LCCN ????

Dietrich:1983:VQF

[Die83] D. Dietrich. Verfahren zur Lösung von Quadratwurzeln für Mikrorechnerprozeduren [English: Methods for the Solution of Square Roots for Microprocessor Subroutines]. Elektroniker (Switzerland), 8:EL–1–EL–6, 1983. CODEN ELKRBL. ISSN 0531-9218.

Dijkstra:1977:CFE

[Dij77] D. Dijkstra. A continued fraction expansion for a generalization of Dawson’s integral. Mathematics of Computation, 31 (138):503–510, April 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dingle:1958:AEC

[Din58] R. B. Dingle. Asymptotic expansions and converging factors. III. gamma, psi and polygamma functions, and Fermi–Dirac and Bose–Einstein integrals. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 244(1239): 484–490, April 22, 1958. CODEN PRLAAZ. ISSN 0080-4630.

URL http://www.jstor.org/stable/100264.

Dingle:1973:AET

[Din73] Robert B. Dingle. Asymptotic expansions: their derivation and interpretation. Academic Press, New York, NY, USA, 1973. ISBN 0-12-216550-0. xv + 521 pp. LCCN QA295 .D45.

Divgi:1979:CUB

[Div79] D. R. Divgi. Calculation of univariate and bivariate normal probability functions. Annals of Statistics, 7(4):903–910, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://projecteuclid.org/euclid.aos/
1176344739.

Dunkl:2000:PIW

[DIW00] Charles Dunkl, Mourad Ismail, and Roderick Wong, editors. Proceedings of the international workshop, special functions: Hong Kong, 21–25 June 1999. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2000. ISBN 981-02-4393-6. LCCN QA351 .P76 2000.

DiDonato:1962:MCG

[DJ62a] A. R. DiDonato and M. P. Jamagin. A method for computing the generalized circular error function and the circular coverage function. NWL Report 1768, Naval Surface Weapons Center, Dahlgren, VA 22448, USA, January 23, 1962.

DiDonato:1962:MCC

[DJ62b] A. R. DiDonato and M. P. Jarnagin. A method for computing the circular coverage function. Mathematics of Computation, 16 (79):347–355, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

DiDonato:1967:ECI

[DJ67] A. R. DiDonato and M. P. Jarnagin. The efficient calculation of the incomplete beta-function ratio for half-integer values of the parameters a, b. Mathematics of Computation, 21(100):652–662, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dzjadyk:1977:TPA

[DK77] V. K. Dzjadyk and S. F. Karpenko. Tables of polynomials for the approximate solution of elementary functions. (Russian). Technical Report 28, Akad. Nauk Ukrain. SSR Inst. Mat. Preprint, 1977. 28 pp.

Decker:1982:CAN

[DK82] D. W. Decker and C. T. Kelley. Convergence acceleration for Newton’s method at singular points. SIAM Journal on Numerical Analysis, 19(1):219–229, February 1982. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dvorak:1990:NCI

[DK90] Steven L. Dvorak and Edward F. Kuester. Numerical computation of the incomplete Lipschitz–Hankel integral Je_{0}(a, z). Journal of Computational Physics, 87(2):301–327, April 1990. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/002199919090255Y.

Ditkin:1980:CSF

[DKK80] V. A. Ditkin, K. A. Karpov, and M. K. Kerimov. The computation of special functions. U.S.S.R. Computational Mathematics and Mathematical Physics, 20(5):3–12, 1980. CODEN CMMPA9. ISSN 0041-5553, 0502-9902.

Lugish:1970:CAA

[dL70] B. G. de Lugish. A class of algorithms for automatic evaluation of certain elementary function in a binary computer. Report 399, Department of Computer Science, University of Illinois, June 1970. ???? pp.

Daumas:2009:VRN

[DLM09] M. Daumas, D. Lester, and C. Muoz. Verified real number calculations: a library for interval arithmetic. IEEE Transactions on Computers, 58(2):226–237, February 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

deLamarliere:2023:SFP

[dLMF23] Paul Geneau de Lamarlière, Guillaume Melquiond, and Florian Faissole. Slimmer formal proofs for mathematical libraries. In IEEE [IEE23], pages 32–35. LCCN ????

Saint-Genies:2018:EFT

[dLSG18] Hugues de Lassus Saint-Geniès. Elementary functions: towards automatically generated, efficient, and vectorizable implementations. Ph.D. thesis, Université de Perpignan, Perpignan, France, July 17, 2018. xxxviii + 128 pp. URL https://tel.archives-

ouvertes.fr/tel-01841424/document.

Saint-Genies:2017:ELT

[dLSGDR17] Hugues de Lassus Saint-Geniès, David Defour, and Guillaume Revy. Exact lookup tables for the evaluation of trigonometric and hyperbolic functions. IEEE Transactions on Computers, 66(12):2058–2071, ???? 2017. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.

ieee.org/document/7927421/.

DiDonato:1986:CIG

[DM86] Armido R. DiDonato and Alfred H. Morris, Jr. Computation of the incomplete gamma function ratios and their inverse. ACM Transactions on Mathematical Software, 12(4):377–393, December 1986. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/pubs/

citations/journals/toms/1986-12-4/p377-didonato/.

DiDonato:1987:AFS

[DM87] Armido R. DiDonato and Alfred H. Morris Jr. Algorithm 654: FORTRAN subroutines for computing the incomplete gamma function ratios and their inverse. ACM Transactions on Mathematical Software, 13(3):318–319, September 1987. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/pdf/10.1145/29380.214348.

DiDonato:1988:SDC

[DM88a] Armido I. DiDonato and Alfred H. Morris, Jr. Significant digit computation of the incomplete beta function ratios. Technical Report NSWC TR 88-365, Naval Surface Warfare Center (K33), Dahlgren, VA 22448-5000, USA, November 1988. URL http:/

/www.dtic.mil/dtic/tr/fulltext/u2/a210118.pdf.

Duprat:1988:EPE

[DM88b] J. Duprat and J. M. Muller. Evaluation of polynomials and elementary functions by integrated circuits. Technical Report RR698-I, IMAG, Grenoble, France, January 1988.

Duprat:1988:HPE

[DM88c] Jean Duprat and Jean-Michel Muller. Hardwired polynomial evaluation. Journal of Parallel and Distributed Computing, 5 (3):291–309, June 1988. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Duprat:1991:WND

[DM91] J. Duprat and J.-M. Muller. Writing numbers differently for faster calculation. Technique et Science Informatiques, 10(3): 211–224, ???? 1991. CODEN TTSIDJ. ISSN 0752-4072, 0264-7419.

DiDonato:1992:ASD

[DM92] Armido R. DiDonato and Alfred H. Morris, Jr. Algorithm 708: Significant digit computation of the incomplete beta function ratios. ACM Transactions on Mathematical Software, 18(3):360–373, September 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://doi.

acm.org/10.1145/131766.131776;   http://www.acm.org/
pubs/citations/journals/toms/1992-18-3/p360-didonato/

. See also [BL94].

Dominici:2008:SFO

[DM08] Diego Dominici and Robert S. Maier, editors. Special functions and orthogonal polynomials: AMS Special Session on Special Functions and Orthogonal Polynomials, April 21–22, 2007, Tucson, Arizona, volume 471 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, 2008. ISBN 0-8218-4650-7. LCCN ????

Daumas:1995:MRR

[DMMM95] Marc Daumas, Christophe Mazenc, Xavier Merrheim, and Jean-Michel Muller. Modular range reduction: a new algorithm for fast and accurate computation of the elementary functions. J.UCS: The Journal of Universal Computer Science, 1(3):162– 175 (electronic), 1995. CODEN ???? ISSN 0948-6968.

Das:1995:IFC

[DMS95] D. Das, K. Mukhopadhyaya, and B. P. Sinha. Implementation of four common functions on an LNS co-processor. IEEE Transactions on Computers, 44(1):155–161, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Dominici:2003:NDS

[Dom03] Diego Dominici. Nested derivatives: a simple method for computing series expansions of inverse functions. International Journal of Mathematics and Mathematical Sciences, 58:3699– 3715, 2003. CODEN ???? ISSN 0161-1712 (print), 1687-0425

(electronic). URL http://www2.newpaltz.edu/~dominicd/
NESTED14.pdf.

Dopper:1942:AOV

[Dop42] Herman Pieter Dopper. Asymptotische Ontwikkelingen van de Onvolledige Gammafuncties. (Dutch) [Asymptotic developments of the Incomplete Gamma Functions]. Ph.D. thesis, Rijksuniversiteit Groningen, Groningen, The Netherlands, 1942. ???? pp.

Dorn:1962:GHR

[Dor62] William S. Dorn. Generalizations of Horner’s rule for polynomial evaluation. IBM Journal of Research and Development, 6(2): 239–245, April 1962. CODEN IBMJAE. ISSN 0018-8646 (print),

2151-8556 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5392358.

Doring:1966:CZC

[Dor66] Boro Doring. Complex zeros of cylinder functions. Mathematics of Computation, 20(94):215–222, April 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dorrer:1968:ADS

[Dor68] Egon Dorrer. Algorithm 322: F -distribution [S14]. Communications of the ACM, 11(2):116–117, February 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dubois:1992:CFQ

[DP92] D. Dubois and H. Prade. Calculation with fuzzy quantities. In AEF’92 [AEF92], pages 24–27. ISBN 2-906899-78-X. LCCN ????

Doman:1995:SAP

[DPC95] B. G. S. Doman, C. J. Pursglove, and W. M. Coen. A set of Ada packages for high precision calculations. ACM Transactions on Mathematical Software, 21(4):416–431, December 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Dyn:1989:PEB

[DR89] N. Dyn and A. Ron. Periodic exponential box splines on a three direction mesh. Journal of Approximation Theory, 56(3):287– 296, March 1989. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Dunkl:1994:AHI

[DR94a] Charles F. Dunkl and Donald E. Ramirez. Algorithm 736: Hyperelliptic integrals and the surface measure of ellipsoids. ACM Transactions on Mathematical Software, 20(4):427–435, December 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/pubs/

citations/journals/toms/1994-20-4/p427-dunkl/.

Dunkl:1994:CHI

[DR94b] Charles F. Dunkl and Donald E. Ramirez. Computing hyperelliptic integrals for surface measure of ellipsoids. ACM Transactions on Mathematical Software, 20(4):413–426, December 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295

(electronic). URL http://www.acm.org/pubs/citations/
journals/toms/1994-20-4/p413-dunkl/.

Deleglise:1998:C

[DR98] Marc Deléglise and Joël Rivat. Computing ψ(x). Mathematics of Computation, 67(224):1691–1696, October 1998. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/jourcgi/jour-pbprocess?fn=110&arg1=
S0025-5718-98-00977-6&u=/mcom/1998-67-224/.

Doss:2004:FBI

[DR04] C. C. Doss and R. L. Riley, Jr. FPGA-based implementation of a robust IEEE-754 exponential unit. In FCCM 2004. 12th Annual IEEE Symposium on Field-Programmable Custom Computing Machines, 20–23 April 2004, pages 229–238. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. CODEN ???? ISSN ????

Drezner:1978:CBN

[Dre78] Z. Drezner. Computation of the bivariate normal integral. Mathematics of Computation, 32(141):277–279, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Dritz:1989:RPS

[Dri89] K. W. Dritz. Rationale for the proposed standard for a generic package of elementary functions for Ada. Report ANL-89/2 Rev. 1, Argonne National Laboratory, Mathematics and Computer Science Division, Argonne, IL, USA, October 1989. ???? pp.

Dritz:1991:PSGa

[Dri91a] K. W. Dritz. Proposed standard for a generic package of elementary functions for Ada. ACM SIGADA Ada Letters, 11(7): 9–46, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dritz:1991:PSGb

[Dri91b] K. W. Dritz. Proposed standard for a generic package of primitive functions for Ada. ACM SIGADA Ada Letters, 11(7):66–82, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dritz:1991:RPS

[Dri91c] K. W. Dritz. Rationale for the proposed standard for a generic package of elementary functions for Ada. ACM SIGADA Ada Letters, 11(7):47–65, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Dritz:1991:IPS

[Dri91d] Kenneth W. Dritz. Introduction to the proposed standard for the elementary functions in Ada. ACM SIGADA Ada Letters, 11 (7):3–8, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Driver:1995:NQH

[Dri95] Kathy Driver. Nondiagonal quadratic Hermite–Padé approximation to the exponential function. Journal of Computational and Applied Mathematics, 65(1–3):125–134, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042795001069.

Demsky:1983:MMC

[DSK83] J. Demsky, M. Schlesinger, and R. D. Kent. Micro/mini computer program for calculating the square root of rationals at arbitrary precision. Computer Physics Communications, 29(3): 237–244, May 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465583900048.

Demsky:1984:MMC

[DSK84] J. Demsky, M. Schlesinger, and R. D. Kent. Micro/mini computer program for calculating the square root of rationals at arbitrary precision. Computer Physics Communications, 35(1–3): C–877, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465584829811.

duToit:1993:BFI

[dT93] C. F. du Toit. Bessel functions J_{n}(z) and Y_{n}(z) of integer order and complex argument. Computer Physics Communications, 78(1–2):181–189, December 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0010465593901534.

Deano:2009:MAS

[DT09] Alfredo Deaño and Nico M. Temme. On modified asymptotic series involving confluent hypergeometric functions. Electronic Transactions on Numerical Analysis (ETNA), 35:88– 103, 2009. CODEN ???? ISSN 1068-9613 (print), 1097-4067

(electronic). URL http://etna.mcs.kent.edu/vol.35.2009/
pp88-103.dir/pp88-103.pdf.

Dattoli:1998:GBF

[DTLM98] G. Dattoli, A. Torre, S. Lorenzutta, and G. Maino. Generalized Bessel functions and Kapteyn series. Computers and Mathematics with Applications, 35(8):117–125, April 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122198000509.

Dubrulle:1983:CNM

[Dub83] Augustin A. Dubrulle. Class of numerical methods for the computation of Pythagorean sums. IBM Journal of Research and Development, 27(6):582–589, November 1983. CODEN IB-MJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). See [MM83] and generalization [Jam89].

Dumbgen:2010:BSG

[Düm10] L. Dümbgen. Bounding standard Gaussian tail probabilities. arxiv.org, ??(??):??, December 9, 2010. URL https://arxiv.

org/abs/1012.2063.

Dunham:1987:PMAa

[Dun87a] Charles B. Dunham. Provably monotone approximations. ACM SIGNUM Newsletter, 22(2):6–11, April 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1987:PMAb

[Dun87b] Charles B. Dunham. Provably monotone approximations, II. ACM SIGNUM Newsletter, 22(3):30–31, July 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1988:PMA

[Dun88] Charles B. Dunham. Provably monotone approximations, III. ACM SIGNUM Newsletter, 23(1):10, January 1988. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1990:FPF

[Dun90] C. B. Dunham. Feasibility of “perfect” function evaluation. ACM SIGNUM Newsletter, 25(4):25–26, October 1990. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dunham:1994:PMAa

[Dun94a] Charles B. Dunham. Provably monotone approximations, IV. Technical report TR-417, Department of Computer Science, University of Western Ontario, London, Ontario, Canada,

March 8, 1994. URL http://www.csd.uwo.ca/tech-reports/.

Dunham:1994:PMAb

[Dun94b] Charles B. Dunham. Provably monotone approximations, IV, revised. Technical report TR-422, Department of Computer Science, University of Western Ontario, London, Ontario, Canada,

April 4, 1994. URL http://www.csd.uwo.ca/tech-reports/.

Dutka:1984:EHH

[Dut84] Jacques Dutka. The early history of the hypergeometric function. Archive for History of Exact Sciences, 31(1):15–34, March 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
31&issue=1&spage=15.

Dutka:1986:SRT

[Dut86] Jacques Dutka. On square roots and their representations. Archive for History of Exact Sciences, 36(1):21–39, March 1986. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
36&issue=1&spage=21.

Dunkl:2001:OPS

[DX01] Charles F. Dunkl and Yuan Xu. Orthogonal Polynomials of Several Variables. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-511-56571-2 (e-book), 0-521-80043-9 (hardcover), 1-107-09582-4. xv + 390 pp. LCCN QA404.5 .D86 2001.

Dunkl:2014:OPS

[DX14] Charles F. Dunkl and Yuan Xu. Orthogonal Polynomials of Several Variables, volume 155 of Encyclopedia of mathematics and its applications. Cambridge University Press, Cambridge, UK, second edition, 2014. ISBN 1-107-07189-5, 1-316-05717-8 (e-book). xvii + 420 pp. LCCN QA404.5 .D86

2014. URL http://catdir.loc.gov/catdir/enhancements/
fy1408/2014001846-b.html;   http://catdir.loc.gov/catdir/
enhancements/fy1408/2014001846-d.html;   http://catdir.
loc.gov/catdir/enhancements/fy1408/2014001846-t.html;
 http://digitale-objekte.hbz-nrw.de/storage2/2014/11/
17/file_11/5877666.pdf.

Dzyadyk:1977:OPM

[DZKK77] V. K. Dzjadyk, Z. V. Zarickaja, S. F. Karpenko, and N. F. Kononova. Ob èffektivnom priblizhenii mnogochlenami èlementarnykh funktsiĭ. (Russian) [efficient approximation by polynomials of elementary functions]. Preprint IM-77-21, Akad. Nauk Ukrain. SSR Inst. Mat., Kiev, USSR, 1977. 42 pp.

Eberlein:1989:SAC

[Ebe89] E. Eberlein. Strong approximation of continuous time stochastic processes. Journal of Multivariate Analysis, 31(2):220–235, November 1989. CODEN JMVAAI. ISSN 0047-259x (print), 1095-7243 (electronic).

Epstein:1979:STE

[EC79] H. I. Epstein and B. F. Caviness. A structure theorem for the elementary functions and its application to the identity problem. International Journal of Computer and Information Sciences, 8 (1):9–37, 1979. CODEN IJCIAH. ISSN 0091-7036.

Erricolo:2013:AFS

[EC13] Danilo Erricolo and Giuseppe Carluccio. Algorithm 934: Fortran 90 subroutines to compute Mathieu functions for complex values of the parameter. ACM Transactions on Mathematical Software, 40(1):8:1–8:19, September 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Eckhardt:1976:RAW

[Eck76] Ulrich Eckhardt. A rational approximation to Weierstrass’ ℘- function. Mathematics of Computation, 30(136):818–826, October 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Eckhardt:1977:RAW

[Eck77] Ulrich Eckhardt. A rational approximation to Weierstrass’ elliptic function. II. The lemniscatic case. Computing: Archiv fur informatik und numerik, 18(4):341–349, 1977. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Eckhardt:1980:AWE

[Eck80] Ulrich Eckhardt. Algorithm 549: Weierstrass’ elliptic functions [S21]. ACM Transactions on Mathematical Software, 6(1):112– 120, March 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ewart:2020:PES

[ECSD20] Timothée Ewart, Francesco Cremonesi, Felix Schürmann, and Fabien Delalondre. Polynomial evaluation on superscalar architecture, applied to the elementary function e^{x}. ACM Transactions on Mathematical Software, 46(3):28:1–28:22, September 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-

7295 (electronic). URL https://dl.acm.org/doi/10.1145/
3408893.

El-Gabali:1997:MTA

[EG97] Magdi A. El-Gabali. Multiple-term approximations for Appell’s F_{1} function. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 39(1):135– 148, July 1997. CODEN JAMMDU. ISSN 0334-2700.

URL https://www.cambridge.org/core/journals/anziam-
journal/article/multipleterm-approximations-for-appells-
f1-function/8E218109899D68FD0DC15B6E9D61E8BD.

Egbert:1977:PCAa

[Egb77] W. E. Egbert. Personal calculator algorithms. I. square roots. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 28(9):22–23, May 1977. CODEN HPJOAX. ISSN 0018-1153.

Egger:1989:PAC

[EH89] A. Egger and R. Huotari. The Polya algorithm on convex sets. Journal of Approximation Theory, 56(2):212–216, February 1989. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Erascu:2016:RQE

[EH16] Madalina Erascu and Hoon Hong. Real quantifier elimination for the synthesis of optimal numerical algorithms (case study: Square root computation). Journal of Symbolic Computation, 75(??):110–126, July/August 2016. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0747717115001091.

Ehrling:1961:NCI

[Ehr61] G. Ehrling. On the numerical computation of incomplete elliptic integrals. Nordisk tidskrift for informationsbehandling, 1(1):8–14, March 1961. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=1&issue=1&spage=8.

Ehrenmark:1989:ONF

[Ehr89] Ulf T. Ehrenmark. Overconvergence of the near-field expansion for linearized waves normally incident on a sloping beach. SIAM Journal on Applied Mathematics, 49(3):799–815, June 1989. CO-DEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Ercegovac:1999:IGD

[EIM^{+}99] Miloš D. Ercegovac, Laurent Imbert, David W. Matula, Jean-Michel Muller, and Guoheng Wei. Improving Goldschmidt division, square root, and square root reciprocal. Research Report 99-41, Laboratoire de l’Informatique du Parallélisme, Lyon, France, September 1999. ii + 17 pp. URL https://inria.hal.

science/inria-00072909/file/RR1999-41.pdf.

Ercegovac:2000:IGD

[EIM^{+}00] Milos D. Ercegovac, Laurent Imbert, David W. Matula, Jean-Michel Muller, and Guoheng Wei. Improving Goldschmidt division, square root, and square root reciprocal. IEEE Transactions on Computers, 49(7):759–763, July 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Einarsson:1979:BEN

[Ein79] Bo Einarsson. Bibliography on the evaluation of numerical software. Journal of Computational and Applied Mathematics, 5(2): 145–159, June 1979. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Einmahl:1989:ERK

[Ein89] U. Einmahl. Extensions of results of Komlos, Major, and Tusnady to the multivariate case. Journal of Multivariate Analysis, 28(1):20–68, January 1989. CODEN JMVAAI. ISSN 0047-259x (print), 1095-7243 (electronic).

Eisman:1963:PER

[Eis63] S. H. Eisman. Polynomial evaluation revisited. Communications of the ACM, 6(7):384–385, July 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Eklund:2001:CEF

[Ekl01] Neil Eklund. CORDIC: Elementary function computation using recursive sequences. College Mathematics Journal, 32(5):330– 333, November 2001. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.2001.11921899.

ElAttar:2006:SFO

[El 06] Refaat A. El Attar. Special functions and orthogonal polynomials, volume 3 of Mathematical series. Lulu Press, Morrisville, NC, USA, 2006. ISBN 1-4116-6690-9 (paperback). vi + 302 pp. LCCN QA404.5 .E5 2006; QA351 .E5 2006.

elLozy:1976:RAC

[eL76] Mohamed el Lozy. Remark on “Algorithm 299: Chi-Squared Integral [S15]”. ACM Transactions on Mathematical Software, 2(4):393–395, December 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [HP67, HP85].

elLozy:1979:RAS

[eL79] Mohamed el Lozy. Remark on “Algorithm 395: Student’s t-Distribution” and remark on “Algorithm 396: Student’s Quantiles [S14]”. ACM Transactions on Mathematical Software, 5 (2):238–239, June 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Hil70a, Hil70b, Hil81a, HP85].

Ercegovac:1989:IMC

[EL89a] M. D. Ercegovac and T. Lang. Implementation of module combining multiplication, division, and square root. In IEEE International Symposium on Circuits and Systems, 8–11 May 1989, volume 1, pages 150–153. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. CODEN ???? ISSN ????

Ercegovac:1989:FRD

[EL89b] M. D. Ercegovac and T. Lang. On-the-fly rounding for division and square root. In Ercegovac and Swartzlander, Jr. [ES89], pages 169–173. ISBN 0-8186-8963-3 (case), 0-8186-5963-7 (microfiche). LCCN QA 76.9 C62 S95 1989. URL http:

//www.acsel-lab.com/arithmetic/arith9/papers/ARITH9_

Ercegovac_rounding.pdf. IEEE catalog no. 89CH2757-3.

Ercegovac:1989:RSR

[EL89c] Miloš D. Ercegovac and Tomas Lang. Radix-4 square root without initial PLA. In Ercegovac and Swartzlander, Jr.

[ES89], pages 162–168. ISBN 0-8186-8963-3 (case), 0-8186-5963-7 (microfiche). LCCN QA 76.9 C62 S95 1989. URL http:

//www.acsel-lab.com/arithmetic/arith9/papers/ARITH9_

Ercegovac_radix4.pdf. IEEE catalog no. 89CH2757-3.

Ercegovac:1990:RSR

[EL90] M. D. Ercegovac and T. Lang. Radix-4 square root without initial PLA. IEEE Transactions on Computers, 39(8):1016–1024, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-

9956 (electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=57040.

Ercegovac:1994:DSR

[EL94] Miloš D. (Dragutin) Ercegovac and Tomas Lang. Division and Square Root: Digit-recurrence Algorithms and Implementations. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1994. ISBN 0-7923-9438-0. x + 230 pp. LCCN QA76.9.C62 E73 1994.

Elbert:2001:CZB

[EL01] Árpád Elbert and Andrea Laforgia. A conjecture on the zeros of Bessel functions. Journal of Computational and Applied Mathematics, 133(1–2):683, August 1, 2001. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042700007172.

Elbert:2008:ZCE

[EL08] Árpád Elbert and Andrea Laforgia. The zeros of the complementary error function. Numerical Algorithms, 49(1–4):153–157, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
49&issue=1&spage=153.

Elbert:2001:SRR

[Elb01] Á. Elbert. Some recent results on the zeros of Bessel functions and orthogonal polynomials. Journal of Computational and Applied Mathematics, 133(1–2):65–83, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037704270000635X.

Elezovic:2015:EPF

[Ele15] Neven Elezović. Estimations of psi function and harmonic numbers. Applied Mathematics and Computation, 258(??):192–205, May 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300315001617.

Ellacott:1983:FTE

[Ell83] S. W. Ellacott. On the Faber transform and efficient numerical rational approximation. SIAM Journal on Numerical Analysis, 20(5):989–1000, October 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Ercegovac:2000:RSR

[ELMT00] Milos D. Ercegovac, Tomás Lang, Jean-Michel Muller, and Arnaud Tisserand. Reciprocation, square root, inverse square root, and some elementary functions using small multipliers. IEEE Transactions on Computers, 49(7):628–637, July 2000. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Evans:1986:RIU

[EM86] D. J. Evans and G. M. Megson. Romberg integration using systolic arrays. Parallel Computing, 3(4):289–304, October 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Everitt:1994:GBF

[EM94] W. N. Everitt and C. Markett. On a generalization of Bessel functions satisfying higher-order differential equations. Journal of Computational and Applied Mathematics, 54(3):325–349, October 20, 1994. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0377042794902550.

Ercegovac:2003:DRA

[EM03] M. D. Ercegovac and J.-M. Muller. Digit-recurrence algorithms for division and square root with limited precision primitives. In Conference Record of the Thirty-Seventh Asilomar Conference on Signals, Systems and Computers, 2003, volume 2, pages 1440–1444. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. CODEN ???? ISSN ????

Ercegovac:2004:CSRa

[EM04a] Miloš Ercegovac and Jean-Michel Muller. Complex square root with operand prescaling. Research Report RR2004-42, École Normale Supérieure de Lyon, 69364 Lyon Cedex 07, France, September 2004. 2 + 12 pp. URL http://www.ens-lyon.fr/

LIP/Pub/Rapports/RR/RR2004/RR2004-42.pdf.

Ercegovac:2004:CSRb

[EM04b] Miloš Ercegovac and Jean-Michel Muller. Complex square root with operand prescaling. In Proceedings of the 15th IEEE International Conference on Application-Specific Systems, Architectures and Processors, 2004, pages 52–62. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. CODEN ???? ISBN 0-7695-2226-2. ISSN 1063-6862.

Ercegovac:2007:CSR

[EM07] Miloš D. Ercegovac and Jean-Michel Muller. Complex square root with operand prescaling. Journal of VLSI Signal Processing, 49(1):19–30, October 2007. CODEN JVSPED. ISSN 0922-5773 (print), 1573-109x (electronic).

Emde:1940:TEF

[Emd40] Fritz Emde. Tafeln Elementarer Funktionen (German) [Tables of Elementary Functions]. B. T. Teubner, Leipzig, Germany and Berlin, Germany, 1940. xii + 181 pp. LCCN QA47 .E5. URL

http://lccn.loc.gov/45006177.

Emde:1945:TEF

[Emd45] Fritz Emde. Tafeln Elementarer Funktionen (German) [Tables of Elementary Functions]. Edwards Bros., Ann Arbor, MI, USA, 1945. xii + 181 pp. LCCN ???? Original edition published in 1940.

Emde:1948:TEF

[Emd48] Fritz Emde. Tafeln Elementarer Funktionen (German) [Tables of Elementary Functions]. Teubner, Stuttgart, Leipzig, second edition, 1948. xii + 181 pp. LCCN QA55 .E5 1948.

Emde:1959:TEF

[Emd59] Fritz Emde. Tafeln Elementarer Funktionen (German) [Tables of Elementary Functions]. B. T. Teubner, Leipzig, Germany and Berlin, Germany, third edition, 1959. xii + 181 pp. LCCN QA47

.E5. URL http://lccn.loc.gov/61020314.

Epstein:1982:UAF

[EMR82a] C. Epstein, W. L. Miranker, and T. J. Rivlin. Ultra-arithmetic I: function data types. Mathematics and Computers in Simulation, 24(1):1–18, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Epstein:1982:UAI

[EMR82b] C. Epstein, W. L. Miranker, and T. J. Rivlin. Ultra-arithmetic II: intervals of polynomials. Mathematics and Computers in Simulation, 24(1):19–29, February 1982. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Epperson:1989:UIM

[Epp89] J. F. Epperson. On the use of iteration methods for approximating the natural logarithm. American Mathematical Monthly, 96 (9):831–835, November 1989. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Epstein:1975:AET

[Eps75] Harvey Irwin Epstein. Algorithms for elementary transcendental function arithmetic. PhD thesis, University of Wisconsin, Madison, WI, USA, 1975. 409 pp.

Ercegovac:1972:RES

[Erc72] Milos D. Ercegovac. Radix 16 evaluation of some elementary functions. Technical Report 540, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana, Illinois, 1972. 30 pp.

Ercegovac:1973:REC

[Erc73] M. D. Ercegovac. Radix-16 evaluation of certain elementary functions. IEEE Transactions on Computers, C-22(6):561–566, June 1973. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ercegovac:1975:GMEa

[Erc75] Miloš Dragutin Ercegovac. A General Method for Evaluation of Functions and Computations in a Digital Computer. Ph.D. thesis, Department of Computer Science, University of Illinois at Urbana-Champaign, Urbana-Champaign, IL, USA, July

1975. viii + 109 pp. URL https://search.proquest.com/
pqdtglobal/docview/302756306.

Ercegovac:1977:GHO

[Erc77] Miloš D. Ercegovac. A general hardware-oriented method for evaluation of functions and computations in a digital computer. IEEE Transactions on Computers, C-26(7):667–680, July 1977. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-

9956 (electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1674900.

Ercegovac:1978:LSR

[Erc78] Miloš D. Ercegovac. An on-line square rooting algorithm. In IEEE [IEE78], pages 183–189. ISSN 1063-6889. LCCN QA76.6

.S919a. URL http://www.acsel-lab.com/arithmetic/arith4/
papers/ARITH4_Ercegovac.pdf. IEEE catalog no. 78 CH1412-

6C.

Ercegovac:1989:PSC

[ES89] Miloš D. Ercegovac and Earl E. Swartzlander, Jr., editors. Proceedings: 9th Symposium on Computer Arithmetic: September 6–8, 1989, Santa Monica, California, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-8963-3 (case), 0-8186-5963-7 (microfiche). LCCN QA 76.9 C62 S95 1989. IEEE catalog no. 89CH2757-3.

Elbert:1999:SFZ

[ES99] Árpád Elbert and Panayiotis D. Siafarikas. On the square of the first zero of the Bessel function J_{ν}(z). Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 42(1):56–77, March 1999. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Erocal:2010:SPU

[ES10] Burçin Eröcal and William Stein. The Sage Project: Unifying free mathematical software to create a viable alternative to Magma, Maple, Mathematica and MATLAB. In Fukuda et al. [FvdHJT10], pages 12–27. ISBN 3-642-15581-2 (paperback), 3-642-15582-0 (e-book). LCCN QA76.95 .I5654 2010. URL http:

//link.springer.com/book/10.1007/978-3-642-15582-6.

Escott:1937:QDN

[Esc37] E. B. Escott. Questions, discussions, and notes: Rapid method for extracting a square root. American Mathematical Monthly, 44(10):644–646, December 1937. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2301484.

Eve:1963:SAI

[Eve63] J. Eve. Starting approximations for the iterative calculation of square roots. The Computer Journal, 6(3):274–276, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Eve:1964:EP

[Eve64] J. Eve. The evaluation of polynomials. Numerische Mathematik, 6(1):17–21, December 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Erdelyi:1963:AEC

[EW63] A. Erdélyi and M. Wyman. The asymptotic evaluation of certain integrals. Archive for Rational Mechanics and Analysis, 14(1): 217–260, January 1963. CODEN AVRMAW. ISSN 0003-9527 (print), 1432-0673 (electronic). URL http://link.springer.

com/article/10.1007/BF00250704.

Ellacott:1976:RCA

[EW76] S. Ellacott and Jack Williams. Rational Chebyshev approximation in the complex plane. SIAM Journal on Numerical Analysis, 13(3):310–323, June 1976. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Fabijonas:2002:LMC

[Fab02] Bruce R. Fabijonas. Laplace’s method on a computer algebra system with an application to the real valued modified Bessel functions. Journal of Computational and Applied Mathematics, 146(2):323–342, September 15, 2002. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042702003643.

Fabijonas:2004:AAF

[Fab04] B. R. Fabijonas. Algorithm 838: Airy functions. ACM Transactions on Mathematical Software, 30(4):491–501, December 2004.

CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fandrianto:1989:AHS

[Fan89] Jan Fandrianto. Algorithms for high-speed shared radix 8 division and radix 8 square root. In Ercegovac and Swartzlander, Jr. [ES89], pages 68–75. ISBN 0-8186-8963-3 (case), 0-8186-5963-7 (microfiche). LCCN QA 76.9 C62 S95

1989. URL http://www.acsel-lab.com/arithmetic/arith9/
papers/ARITH9_Fandrianto.pdf. IEEE catalog no. 89CH2757-

3.

Farkas:1969:CAS

[Far69] I. Farkas. Certification of Algorithm 165 [S21]: Complete elliptic integrals. Communications of the ACM, 12(1):38, January 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Farmwald:1981:HBE

[Far81] P. Michael Farmwald. High bandwidth evaluation of elementary functions. In IEEE [IEE81a], pages 139–142. LCCN QA76.9.C62 S95 1981. IEEE Catalog No. 81CH1630-3. Computer Society Order No. 347.

Fateman:2016:CUA

[Fat16] Richard J. Fateman. Comments on unrestricted algorithms for Bessel functions in computer algebra: Arbitrary precision, the backwards recurrence, Taylor series, Hermite interpolation. Report, University of California, Berkeley, Berkeley, CA 947220-1776, USA, June 4, 2016. URL https://people.eecs.

berkeley.edu/~fateman/papers/hermite.pdf.

Ferguson:1991:AMA

[FB91] W. E. Ferguson, Jr. and T. Brightman. Accurate and monotone approximations of some transcendental functions. In Kornerup and Matula [KM91], pages 237–244. ISBN 0-8186-9151-4 (case), 0-8186-6151-8 (microfiche), 0-7803-0187-0 (library binding). LCCN QA76.9.C62 S95 1991.

Frougny:2004:RCR

[FBM04] Christiane Frougny, Vasco Brattka, and Norbert Müller, editors. RNC’6, 6th Conference on Real Numbers and Computers: Nov 15–17, 2004, Dagstuhl, Germany. Universitaät Trier, Fachbereich IV, Mathematik, Informatik, Trier, Germany, 2004. ISSN 0944-0488. URL http://www.

informatik.uni-trier.de/Reports/TR-08-2004;   http:/
/www.informatik.uni-trier.de/Reports/TR-08-2004/rnc6-

complete.pdf. Forschungsbericht Nr. 04-8.

Fisher:1960:PPD

[FC60] Ronald A. Fisher and E. A. Cornish. The percentile points of distributions having known cumulants. Technometrics, 2(2):209– 225, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1266546. See errata [Ano60b].

Fisherkeller:1964:TCE

[FC64] M. A. Fisherkeller and W. J. Cody, Jr. Tables of the complete elliptic integrals K, K′, E, and E′. Technical Memo ANL AMD 71, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1964. 14 pp. See review by John W. Wrench in Mathematics of Computation, 19(89–92), 342, 1965.

Fettis:1973:CZE

[FCC73a] Henry E. Fettis, James C. Caslin, and Kenneth R. Cramer. Complex zeros of the error function and of the complementary error function. Mathematics of Computation, 27(122):401–407, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1973:SPC

[FCC73b] Henry E. Fettis, James C. Caslin, and Kenneth R. Cramer. Saddle points of the complementary error function. Mathematics of Computation, 27(122):409–412, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fdil:1997:SRC

[Fdi97] A. Fdil. Some results of convergence acceleration for a general Θ-type algorithm. Applied Numerical Mathematics: Transactions of IMACS, 23(2):219–240, March 21, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.elsevier.com/cas/tree/store/
apnum/sub/1997/23/2/738.pdf.

Feinsilver:1978:SFP

[Fei78] Philip J. (Philip Joel) Feinsilver. Special functions, probability semigroups, and Hamiltonian flows, volume 696 of Lecture notes in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. ISBN 0-387-09100-9. vi + 112 pp. LCCN QA3 .L28 no. 696; QA273.

FernandezVelicia:1986:HPA

[Fer86] F. J. Fernández Velicia. High-precision analytic approximations for the Fermi–Dirac functions by means of elementary functions. Physical Review A (3), 34(5):4387–4395, 1986. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Ferraro:2007:FAG

[Fer07] Giovanni Ferraro. The foundational aspects of Gauss’s work on the hypergeometric, factorial and digamma functions. Archive for History of Exact Sciences, 61(5):457–518, September 2007. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
61&issue=5&spage=457.

Fettis:1963:AMH

[Fet63] Henry E. Fettis. Algorithm 163: Modified Hankel function. Communications of the ACM, 6(4):161–162, April 1963. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fettis:1965:CEI

[Fet65] Henry E. Fettis. Calculation of elliptic integrals of the third kind by means of Gauss’ transformation. Mathematics of Computation, 19(89):97–104, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

jstor.org/stable/2004102.

Fettis:1967:MCI

[Fet67] Henry E. Fettis. More on the calculation of the integral I_{n}(b) = \frac{2}{π} ∫_{0}^{∞} (\frac{sin x}{x})^{n} cos bx dx (in Technical Notes and Short Papers). Mathematics of Computation, 21(100):727–730, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1974:SAC

[Fet74] Henry E. Fettis. A stable algorithm for computing the inverse error function in the ‘tail-end’ region. Mathematics of Computation, 28(126):585–587, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1976:CR

[Fet76] Henry E. Fettis. Complex roots of sin z = az, cos z = az, and cosh z = az. Mathematics of Computation, 30(135):541–545, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1979:AEU

[Fet79] Henry E. Fettis. An asymptotic expansion for the upper percentage points of the χ^{2}-distribution. Mathematics of Computation, 33(147):1059–1064, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1981:TETb

[Fet81a] Henry E. Fettis. Table errata: A table of the complete elliptic integral of the first kind for complex values of the modulus, Part I [Rep. No. ARL 69-0172, Aerospace Res. Lab., Wright–Patterson Air Force Base, Ohio, 1969; MR 40 #6725] by Fettis and J. C. Caslin. Mathematics of Computation, 36(153):318, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fettis:1981:TEHb

[Fet81b] Henry E. Fettis. Table errata: Handbook of elliptic integrals for engineers and physicists [second edition, Springer, New York, 1971 and MR 43 #3506] by P. F. Byrd and M. D. Friedman. Mathematics of Computation, 36(153):317, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Franklin:1954:CARa

[FF54] J. Franklin and B. Friedman. A convergent asymptotic representation for integrals. Technical report, Division of Electromagnetic Research, Institute of Mathematical Sciences, New York University, New York, NY, USA, 1954. i + 17 pp.

Franklin:1957:CARb

[FF57] Joel Franklin and Bernard Friedman. A convergent asymptotic representation for integrals. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 53:612– 619, 1957. CODEN PCPSA4. ISSN 0008-1981. URL https:/

/www.cambridge.org/core/services/aop-cambridge-core/
content/view/2F167020D3A1F98182882E99549E7751/S0305004100032667a.
pdf/a-convergent-asymptotic-representation-for-integrals.
pdf.

Fessler:1983:HAA

[FFS83] Theodore Fessler, William F. Ford, and David A. Smith. HURRY: An acceleration algorithm for scalar sequences and series. ACM Transactions on Mathematical Software, 9(3):346– 354, September 1983. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fenton:1982:RCM

[FGG82] J. D. Fenton and R. S. Gardiner-Garden. Rapidly-convergent methods for evaluating elliptic integrals and theta and elliptic functions. Journal of the Australian Mathematical Society: Series B, Applied Mathematics, 24(1):47–58, July 1982. CODEN JAMMDU. ISSN 0334-2700. URL https://www.cambridge.

org/core/journals/anziam-journal/article/rapidlyconvergent-
methods-for-evaluating-elliptic-integrals-and-theta-
and-elliptic-functions/2D993C9A7C9EB1D4B61B856E22B45A34.

Fraser:1962:CRA

[FH62] W. Fraser and J. F. Hart. On the computation of rational approximations to continuous functions. Communications of the ACM, 5(7):401–403, July 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fousse:2007:MMP

[FHL^{+}07] Laurent Fousse, Guillaume Hanrot, Vincent Lefèvre, Patrick Pélissier, and Paul Zimmermann. MPFR: a multiple-precision binary floating-point library with correct rounding. ACM Transactions on Mathematical Software, 33(2):1–15, June 2007. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fox:1977:PMS

[FHS77] P. A. Fox, A. D. Hall, and N. L. Schryer. The PORT mathematical subroutine library. Computing Science Technical Report 47, AT&T Bell Laboratories, Murray Hill, NJ, USA, March

22, 1977. ii + 50 pp. URL http://history.siam.org/%5C/
sup/Fox_bell_subroutine.pdf.

Fukushima:1994:NCI

[FI94] Toshio Fukushima and Hideharu Ishizaki. Numerical computation of incomplete elliptic integrals of a general form. Celestial Mechanics and Dynamical Astronomy, 59(3):237–251, July 1994. CODEN CLMCAV. ISSN 0923-2958 (print), 1572-9478 (elec-

tronic). URL http://www.springerlink.com/content/0923-
2958/.

Fields:1965:RAG

[Fie65] Jerry L. Fields. Rational approximations to generalized hypergeometric functions. Mathematics of Computation, 19(92):606– 624, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fike:1966:SAS

[Fik66] C. T. Fike. Starting approximations for square root calculation on IBM System/360. Communications of the ACM, 9(4):297– 299, April 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See letter [Fik67a].

Fike:1967:LER

[Fik67a] C. T. Fike. Letter to the Editor: A rational approximation optimal by Moursund’s criterion. Communications of the ACM, 10 (11):683–684, November 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Mou67, Fik66].

Fike:1967:MEP

[Fik67b] C. T. Fike. Methods of evaluating polynomial approximations in function evaluation routines. Communications of the ACM, 10 (3):175–178, March 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark on efficiency [Cod67].

Fike:1968:CEM

[Fik68] C. T. Fike. Computer Evaluation of Mathematical Functions. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1968. xii + 227 pp. LCCN QA297 .F5.

Filippi:1966:BEE

[Fil66] S. Filippi. Die Berechnung einiger elementarer transzendenter Funktionen mit Hilfe des Richardson-Algorithmus [English: The Computation of Some Elementary Transcendental Functions by Means of the Richardson Algorithm]. Computing: Archiv fur informatik und numerik, 1(2):127–132, June 1966. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Fillebrown:1992:FCB

[Fil92] Sandra Fillebrown. Faster computation of Bernoulli numbers. Journal of Algorithms, 13(3):431–445, September 1992. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/019667749290048H.

Fisher:1973:NEI

[Fis73] N. I. Fisher. A note on the evaluation of the incomplete gamma function. Journal of Statistical Computation and Simulation, 2 (4):325–332, 1973. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Fornberg:1975:CZJ

[FK75] B. Fornberg and K. S. Kolbig. Complex zeros of the Jonquiere or polylogarithm function. Mathematics of Computation, 29 (130):582–599, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Feng:1988:AIN

[FK88] Y. Y. Feng and J. Kozak. An approach to the interpolation of nonuniformly spaced data. Journal of Computational and Applied Mathematics, 23(2):169–178, August 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Fu:2011:ETB

[FK11] Hua Fu and Pooi-Yuen Kam. Exponential-type bounds on the first-order Marcum q-function. In 2011 IEEE Global Telecommunications Conference — GLOBECOM 2011. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, December 2011. URL http://ieeexplore.ieee.

org/document/6133801/.

Fedotov:2019:CWM

[FK19] Alexander Fedotov and Frédéric Klopp. The complex WKB method for difference equations and Airy functions. SIAM journal on mathematical analysis, 51(6):4413–4447, ???? 2019. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Froman:1986:PIF

[FKY86] Per Olof Fröman, Finn Karlsson, and Staffan Yngve. Phaseintegral formulas for Bessel functions and their relation to already existing asymptotic formulas. Journal of Mathematical Physics, 27(11):2738–2747, November 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/
i11/p2738_s1.

Fair:1967:RAI

[FL67] Wyman G. Fair and Yudell L. Luke. Rational approximations to the incomplete elliptic integrals of the first and second kinds. Mathematics of Computation, 21(99):418–422, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Flammer:1957:SWF

[Fla57] Carson Flammer. Spheroidal Wave Functions. A Stanford Research Institute monograph. Stanford University Press, Stanford, CA, USA, 1957. ix + 220 pp. LCCN QA405 .F55.

Fleckner:1968:MCF

[Fle68] Oscar L. Fleckner. A method for the computation of the Fresnel integrals and related functions. Mathematics of Computation, 22 (103):635–640, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Favati:2000:SAC

[FLMR00] P. Favati, G. Lotti, O. Menchi, and F. Romani. Separable asymptotic cost of evaluating elementary functions. Numerical Algorithms, 24(3):255–274, 2000. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Floyd:1961:ACE

[Flo61] Robert W. Floyd. An algorithm for coding efficient arithmetic operations. Communications of the ACM, 4(1):42–51, January 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Fabijonas:2004:CCA

[FLO04] B. R. Fabijonas, D. W. Lozier, and F. W. J. Olver. Computation of complex Airy functions and their zeros using asymptotics and the differential equation. ACM Transactions on Mathematical Software, 30(4):471–490, December 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Flocke:2015:AAE

[Flo15] N. Flocke. Algorithm 954: an accurate and efficient cubic and quartic equation solver for physical applications. ACM Transactions on Mathematical Software, 41(4):30:1–30:24, October 2015. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fabijonas:2003:ACM

[FLR03] B. R. Fabijonas, Daniel W. Lozier, and J. M. Rappoport. Algorithms and codes for the Macdonald function: Recent progress and comparisons. Journal of Computational and Applied Mathematics, 161(1):179–192, ???? 2003. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://math.nist.gov/acmd/Staff/DLozier/
publications/nistir6596.ps.

Ferreira:2006:GHF

[FLS06] C. Ferreira, J. L. López, and E. P. Sinusía. The Gauss hypergeometric function F (a; b; c; z) for large c. Journal of Computational and Applied Mathematics, 197(2):568–577, January 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Fernandez:1982:HCI

[FMC82] F. M. Fernández, A. Mesón, and E. A. Castro. Hypervirial calculation of integrals involving Bessel functions. Journal of Mathematical Physics, 23(2):254–255, February 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v23/
i2/p254_s1.

Ferguson:1975:PFI

[FNC75] Helaman Rolfe Pratt Ferguson, Dale E. Nielsen, and Grant Cook. A partition formula for the integer coefficients of the theta function nome. Mathematics of Computation, 29(131):851–855, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Frappier:1993:QFI

[FO93] Clément Frappier and Patrick Olivier. A quadrature formula involving zeros of Bessel functions. Mathematics of Computation, 60(201):303–316, January 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fabijonas:1999:RAE

[FO99] Bruce R. Fabijonas and F. W. J. Olver. On the reversion of an asymptotic expansion and the zeros of the Airy functions. SIAM Review, 41(4):762–773, December 1999. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL

http://epubs.siam.org/sam-bin/dbq/article/34953.

Fogel:1951:FTE

[Fog51] È. K. Fogel’. A finite theory of elementary functions. I. Logarithmic and exponential functions. (Russian). Latvijas PSR Zinātņu Akad. Vēstis, 5(46):801–813, 1951.

Forrey:1997:CHF

[For97] Robert C. Forrey. Computing the hypergeometric function. Journal of Computational Physics, 137(1):79–100, October 1997. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Fowkes:1993:HEA

[Fow93] R. E. Fowkes. Hardware efficient algorithms for trigonometric functions. IEEE Transactions on Computers, 42(2):235–239, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-

9956 (electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=204796.

Fox:1954:STB

[Fox54] L. Fox. A Short Table for Bessel Functions of Integer Orders and Large Arguments, volume 3 of Royal Society Shorter Mathematical Tables. Cambridge University Press, Cambridge, UK, 1954. 28 pp.

Fox:1960:TWP

[Fox60] L. Fox. Tables of Weber Parabolic Cylinder Functions and Other Functions for Large Arguments, volume 4 of National Physical Laboratory Mathematical Tables. Her Majesty’s Stationery Office, London, UK, 1960. iii + 40 pp. Department of Scientific and Industrial Research.

Fox:1968:CPN

[FP68] L. Fox and I. B. Parker. Chebyshev Polynomials in Numerical Analysis. Oxford mathematical handbooks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1968. ISBN 0-19-859614-6. ix + 205 pp. LCCN QA297 .F65. Reprinted in 1972 with corrections, but same ISBN.

Fullerton:1976:AEM

[FR76] L. W. Fullerton and G. A. Rinker, Jr. Accurate and efficient methods for the evaluation of vacuum-polarization potentials of order Zα and Zα^{2}. Physical Review A, 13(3):1283–1287, March 1976. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Fowler:1998:SRA

[FR98] David Fowler and Eleanor Robson. Square root approximations in old Babylonian mathematics: YBC 7289 in context. Historia Mathematica, 25(4):366–378, November 1998. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086098922091.

Franke:1965:NEE

[Fra65a] Charles H. Franke. Numerical evaluation of the elliptic integral of the third kind (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):494–496, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fraser:1965:SMC

[Fra65b] W. Fraser. A survey of methods for computing minimax and near-minimax polynomial approximations for functions of a single independent variable. Journal of the Association for Computing Machinery, 12(3):295–314, July 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Frankowski:1978:RME

[Fra78] Krzysztof S. Frankowski. A realistic model for error estimates in the evaluation of elementary functions. In IEEE [IEE78], pages 70–74. ISSN 1063-6889. LCCN QA76.6 .S919a. IEEE catalog no. 78 CH1412-6C.

Fransen:1981:ACH

[Fra81] Arne Fransén. Addendum and corrigendum to: “High-precision values of the gamma function and of some related coefficients” [Math. Comp. 34 (1980), no. 150, 553–566, MR 81f:65004] by Fransén and S. Wrigge. Mathematics of Computation, 37 (155):233–235, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [FW80].

Fredette:1981:RES

[Fre81] G. Fredette. 68000 routine extracts square roots. EDN, 26(16): 185–194, August 1981. CODEN EDNSBH. ISSN 0012-7515, 0364-6637.

Freitas:2005:IPF

[Fre05] Pedro Freitas. Integrals of polylogarithmic functions, recurrence relations, and associated Euler sums. Mathematics of Computation, 74(251):1425–1440, July 2005. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

ams.org/mcom/2005-74-251/S0025-5718-05-01747-3/home.
html;   http://www.ams.org/mcom/2005-74-251/S0025-5718-
05-01747-3/S0025-5718-05-01747-3.dvi;   http://www.ams.
org/mcom/2005-74-251/S0025-5718-05-01747-3/S0025-5718-
05-01747-3.pdf;   http://www.ams.org/mcom/2005-74-251/
S0025-5718-05-01747-3/S0025-5718-05-01747-3.ps;   http:
//www.ams.org/mcom/2005-74-251/S0025-5718-05-01747-3/
S0025-5718-05-01747-3.tex.

FreitasDeAbreu:2009:JCU

[Fre09] Giuseppe Thadeu Freitas De Abreu. Jensen–Cotes upper and lower bounds on the Gaussian Q-function and related functions. IEEE Transactions on Communications, 57(11):3328– 3338, November 2009. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic).

Friedland:1967:AAV

[Fri67] Paul Friedland. Algorithm 312: Absolute value and square root of a complex number. Communications of the ACM, 10(10):

665, October 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frisch:1972:RAR

[Fri72] Michael J. Frisch. Remark on “Algorithms 352, 385, 392: Remarks on characteristic values and associated solutions of Mathieu’s differential equation, exponential integral, and systems of hyperbolic P.D.E.”. Communications of the ACM, 15(12):1074– ??, 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Cle69, Pac70, SM70].

Froberg:1955:NTC

[Frö55] Carl-Erik Fröberg. Numerical treatment of Coulomb wave functions. Reviews of Modern Physics, 27(4):399–411, October 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.27.399;   http://rmp.aps.org/
abstract/RMP/v27/i4/p399_1.

Froberg:1961:RCA

[Frö61] Carl-Erik Fröberg. Rational Chebyshev approximations of elementary functions. Nordisk tidskrift for informationsbehandling, 1(4):256–262, December 1961. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0006-3835&volume=1&issue=4&spage=

256. See erratum [Frö63a].

Froberg:1963:ERC

[Frö63a] C.-E. Fröberg. Erratum: “Rational Chebyshev Approximations of Elementary Functions” [BIT 1(4), 1961, p. 261, line 12]. Nordisk tidskrift for informationsbehandling, 3(1):68, March 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0006-3835&volume=
3&issue=1&spage=68. See [Frö61].

Froberg:1963:APC

[Frö63b] Carl-Erik Fröberg. Algol programming: Contribution no. 5: Computation of the Fermi function. Nordisk tidskrift for informationsbehandling, 3(2):141–142, 1963. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Feinsilver:1992:BFR

[FS92] P. Feinsilver and R. Schott. On Bessel functions and rate of convergence of zeros of Lommel polynomials. Mathematics of Computation, 59(199):153–156, July 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Feinsilver:1993:ASO

[FS96] Philip J. (Philip Joel) Feinsilver and René Schott. Algebraic Structures and Operator Calculus, volume 241, 292, 347 of Mathematics and its applications. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1993, 1994, 1996. ISBN 0-7923-2116-2 (v. 1), 0-7923-2921-X (v. 2), 0-7923-3834-0 (v. 3). ????? pp. LCCN QA432 .F45 1993. Volume 1: Representations and probability theory. Volume 2: Special functions and computer science. Volume 3: Representations of Lie groups.

Fukushima:1983:OAA

[Fuk83] M. Fukushima. An outer approximation algorithm for solving general convex programs. Operations Research, 31(1):101–113, February 1983. CODEN OPREAI. ISSN 0030-364X (print), 1526-5463 (electronic).

Fukushima:2009:FCC

[Fuk09a] Toshio Fukushima. Fast computation of complete elliptic integrals and Jacobian elliptic functions. Celestial Mechanics and Dynamical Astronomy, 105(4):305–328, December 2009. CO-DEN CLMCAV. ISSN 0923-2958 (print), 1572-9478 (electronic).

URL http://www.springerlink.com/content/0923-2958/.

Fukushima:2009:FCJ

[Fuk09b] Toshio Fukushima. Fast computation of Jacobian elliptic functions and incomplete elliptic integrals for constant values of elliptic parameter and elliptic characteristic. Celestial Mechanics and Dynamical Astronomy, 105(1–3):245–260, 2009. CODEN CLMCAV. ISSN 0923-2958 (print), 1572-9478 (electronic). URL

http://www.springerlink.com/content/0923-2958/.

Fukushima:2010:FCI

[Fuk10] Toshio Fukushima. Fast computation of incomplete elliptic integral of first kind by half argument transformation. Numerische Mathematik, 116(4):687–719, October 2010. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0029-599X&volume=116&issue=4&spage=687.

Fukushima:2011:PFC

[Fuk11] Toshio Fukushima. Precise and fast computation of the general complete elliptic integral of the second kind. Mathematics of Computation, 80(275):1725–1743, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/journals/mcom/2011-80-
275/S0025-5718-2011-02455-5/home.html;   http://www.
ams.org/journals/mcom/2011-80-275/S0025-5718-2011-02455-
5/S0025-5718-2011-02455-5.pdf.

Fukushima:2012:SES

[Fuk12] Toshio Fukushima. Series expansions of symmetric elliptic integrals. Mathematics of Computation, 81(278): 957–990, 2012. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.ams.org/

journals/mcom/2012-81-278/S0025-5718-2011-02531-7;
http://www.ams.org/journals/mcom/2012-81-278/S0025-5718-
2011-02531-7/S0025-5718-2011-02531-7.pdf.

Fukushima:2013:PFC

[Fuk13a] Toshio Fukushima. Precise and fast computation of Jacobian elliptic functions by conditional duplication. Numerische Mathematik, 123(4):585–605, April 2013. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.

springer.com/article/10.1007/s00211-012-0498-0.

Fukushima:2013:RCD

[Fuk13b] Toshio Fukushima. Recursive computation of derivatives of elliptic functions and of incomplete elliptic integrals. Applied Mathematics and Computation, 221(??):21–31, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300313006152.

Fukushima:2014:ACG

[Fuk14a] Toshio Fukushima. Analytical computation of generalized Fermi–Dirac integrals by truncated Sommerfeld expansions. Applied Mathematics and Computation, 234(??):417–433, May 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300314002926.

Fukushima:2014:CGI

[Fuk14b] Toshio Fukushima. Computation of a general integral of Fermi–Dirac distribution by McDougall–Stoner method. Applied Mathematics and Computation, 238(??):485–510, July 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S009630031400561X.

Fukushima:2015:PFCc

[Fuk15a] Toshio Fukushima. Precise and fast computation of elliptic integrals and functions. In Muller et al. [MTV15], pages 50–57. ISBN 1-4799-8665-8, 1-4799-8663-1. ISSN 1063-6889. LCCN

QA76.9.C62 S95 2015. URL http://ieeexplore.ieee.org/
servlet/opac?punumber=7193754.

Fukushima:2015:PFCb

[Fuk15b] Toshio Fukushima. Precise and fast computation of Fermi– Dirac integral of integer and half integer order by piecewise minimax rational approximation. Applied Mathematics and Computation, 259(??):708–729, May 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0096300315003033.

Fukushima:2015:PFCd

[Fuk15c] Toshio Fukushima. Precise and fast computation of generalized Fermi–Dirac integral by parameter polynomial approximation. Applied Mathematics and Computation, 270(??):802–807, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300315011509.

Fukushima:2015:PFCa

[Fuk15d] Toshio Fukushima. Precise and fast computation of inverse Fermi–Dirac integral of order 1/2 by minimax rational function approximation. Applied Mathematics and Computation, 259(??):698–707, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300315003094.

Fullerton:1972:MIG

[Ful72] W. Fullerton. ACM Algorithm 435: Modified incomplete gamma function. Communications of the ACM, 15(11):993–995, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Sch78b].

Fullerton:1977:PSF

[Ful77] L. W. Fullerton. Portable special function routines. In Cowell [Cow77], pages 452–483. ISBN 0-387-08446-0. LCCN QA297 .W65 1976.

Fullerton:1980:BEM

[Ful80] L. W. Fullerton. A bibliography on the evaluation of mathematical functions. Technical report TM 80-1274-4 and CSTR 86, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 1980.

Fullerton:1981:FUMb

[Ful81a] L. W. Fullerton. FNLIB user’s manual. Technical report CSTR 95, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 1981.

Fullerton:1981:FUMa

[Ful81b] L. W. Fullerton. FNLIB user’s manual explanatory table of contents. Technical report CSTR 92, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 1981.

Fuller:1999:HVH

[Ful99] A. Thomas Fuller. Horner versus Holdred: an episode in the history of root computation. Historia Mathematica, 26(1):29–51, February 1, 1999. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0315086098922145.

Fukuda:2010:MSI

[FvdHJT10] Komei Fukuda, Joris van der Hoeven, Michael Joswig, and Nobuki Takayama, editors. Mathematical software — ICMS 2010: third International Congress on Mathematical Software, Kōbe, Japan, September 13–17, 2010: proceedings, volume 6327 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-15581-2 (paperback), 3-642-15582-0 (e-book).

LCCN QA76.95 .I5654 2010. URL http://link.springer.

com/book/10.1007/978-3-642-15582-6.

Fields:1961:EHF

[FW61] Jerry L. Fields and Jet Wimp. Expansions of hypergeometric functions in hypergeometric functions. Mathematics of Computation, 15(76):390–395, October 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Fransen:1980:HPV

[FW80] Arne Fransén and Staffan Wrigge. High-precision values of the gamma function and of some related coefficients. Mathematics of Computation, 34(150):553–566, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See addendum and corrigendum [Fra81].

Fransen:1984:CMM

[FW84] Arne Fransén and Staffan Wrigge. Calculation of the moments and the moment generating function for the reciprocal gamma distribution. Mathematics of Computation, 42(166):601–616, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Frenzen:1985:NAE

[FW85] C. L. Frenzen and R. Wong. A note on asymptotic evaluation of some Hankel transforms. Mathematics of Computation, 45(172): 537–548, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Feng:2013:TFA

[FW13] Lei Feng and Weiping Wang. Two families of approximations for the gamma function. Numerical Algorithms, 64(3):403– 416, November 2013. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-012-9671-x.

Gabutti:2008:EQG

[GA08] Bruno Gabutti and Giampietro Allasia. Evaluation of q-gamma function and q-analogues by iterative algorithms. Numerical Algorithms, 49(1–4):159–168, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
1017-1398&volume=49&issue=1&spage=159.

Gabutti:1979:HPM

[Gab79] Bruno Gabutti. On high precision methods for computing integrals involving Bessel functions. Mathematics of Computation, 33(147):1049–1057, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gal:1985:CEF

[Gal85] Shmuel Gal. Computing elementary functions: a new approach for achieving high accuracy and good performance. In Miranker and Toupin [MT85], pages 1–16. ISBN 0-387-16798-6. LCCN QA76.95 .A231 1986.

Gal:1986:CEF

[Gal86] Shmuel Gal. Computing elementary functions: a new approach for achieving high accuracy and good performance. In Miranker and Toupin [MT86], pages 1–16. CODEN LNCSD9. ISBN 0-387-16798-6 (USA: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.95 .A231 1986. Symposium sponsored by IBM Deutschland.

Gargantini:1975:PSR

[Gar75] I. Gargantini. Parallel square root iterations. In K. Nickel, editor, Interval Mathematics, volume 29 of Lecture Notes In Computer Science, pages 196–204. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975.

Gargantini:1979:NSS

[Gar79] Irene Gargantini. The numerical stability of simultaneous iterations via square-rooting. Computers and Mathematics with Applications, 5(1):25–31, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/089812218190136X.

Gargantini:1980:PSR

[Gar80] Irene Gargantini. Parallel square-root iterations for multiple roots. Computers and Mathematics with Applications, 6(3):279– 288, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0898122180900358.

Gasper:1981:SFB

[Gas81] George Gasper. Summation formulas for basic hypergeometric series. SIAM journal on mathematical analysis, 12(2):196–200, March 1981. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gautschi:1959:EIL

[Gau59] W. Gautschi. Exponential integral ∫_{1}^{∞} e^{−xt}t^{−n} dt for large values of n. Journal of Research of the National Bureau of Standards (1934), 62(3):123–125, March 1959. ISSN 0091-0635.

Gautschi:1961:RCR

[Gau61] Walter Gautschi. Recursive computation of the repeated integrals of the error function. Mathematics of Computation, 15 (75):227–232, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1963:RCS

[Gau63] Walter Gautschi. Recursive computation of special functions. In ????, editor, The University of Michigan Engineering Summer Conferences, Numerical Analysis, Summer 1963, page ?? ????, ????, 1963.

Gautschi:1964:AAB

[Gau64a] W. Gautschi. ACM Algorithm 236: Bessel functions of the first kind [S17]. Communications of the ACM, 7(8):479–480, August 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Sko75a].

Gautschi:1964:AGF

[Gau64b] Walter Gautschi. Algorithm 221: Gamma functions. Communications of the ACM, 7(3):143, March 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1964:AIB

[Gau64c] Walter Gautschi. Algorithm 222: Incomplete beta functions ratios. Communications of the ACM, 7(3):143, March 1964. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1964:CAI

[Gau64d] Walter Gautschi. Certification of Algorithm 222: Incomplete beta function ratios. Communications of the ACM, 7(4):244, April 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1964:EFF

[Gau64e] Walter Gautschi. Error function and Fresnel integrals. In Abramowitz and Stegun [AS64], pages 295–330. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also avail-

able online at http://www.convertit.com/Go/ConvertIt/

Reference/AMS55.ASP in bitmap image format.

Gautschi:1965:ALF

[Gau65a] W. Gautschi. Algorithm 259: Legendre functions for arguments larger than one [S16]. Communications of the ACM, 8(8):488– 492, August 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Jan77].

Gautschi:1965:CAS

[Gau65b] Walter Gautschi. Certification of Algorithm 236 [S17]: Bessel functions of the first kind. Communications of the ACM, 8(2): 105–106, February 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1966:AD

[Gau66a] Walter Gautschi. Algorithm 282: Derivatives of e^{x}/x, cos(x)/x, and sin(x)/x. Communications of the ACM, 9(4):272, April 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [GK70b].

Gautschi:1966:ARC

[Gau66b] Walter Gautschi. Algorithm 292: Regular Coulomb wave functions. Communications of the ACM, 9(11):793–795, November 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1967:CAT

[Gau67] Walter Gautschi. Computational aspects of three-term recurrence relations. SIAM Review, 9(1):24–82, January 1967. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://link.aip.org/link/?SIR/9/24/1.

Gautschi:1969:ACE

[Gau69a] Walter Gautschi. Algorithm 363: Complex error function [S15]. Communications of the ACM, 12(11):635, November 1969. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Köl72a].

Gautschi:1969:RAS

[Gau69b] Walter Gautschi. Remark on Algorithm 292 [S22]: Regular Coulomb wave functions. Communications of the ACM, 12(5): 280, May 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1973:AAE

[Gau73] Walter Gautschi. ACM Algorithm 471: Exponential integrals [S13]. Communications of the ACM, 16(12):761–763, December 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1975:CMS

[Gau75] W. Gautschi. Computational methods in special functions — a survey. In Askey [Ask75], pages 1–98. ISBN 0-12-064850-4. LCCN QA3 .U45 no. 35 QA351.

Gautschi:1977:ACC

[Gau77a] Walter Gautschi. Anomalous convergence of a continued fraction for ratios of Kummer functions. Mathematics of Computation, 31(140):994–999, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gautschi:1977:ERI

[Gau77b] Walter Gautschi. Evaluation of repeated integrals of the coerror function. ACM Transactions on Mathematical Software, 3(3): 240–252, September 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gautschi:1979:AIG

[Gau79a] W. Gautschi. Algorithm 542: Incomplete gamma functions [S14]. ACM Transactions on Mathematical Software, 5(4):482– 489, December 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gautschi:1979:CPI

[Gau79b] Walter Gautschi. A computational procedure for incomplete gamma functions. ACM Transactions on Mathematical Software, 5(4):466–481, December 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gautschi:1994:MCH

[Gau94] Walter Gautschi, editor. Mathematics of computation, 1943– 1993: a half-century of computational mathematics: Mathematics of Computation 50th Anniversary Symposium, August 9–13, 1993, Vancouver, British Columbia, volume 48 of Proceedings of Symposia in Applied Mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. See also SIAM Review, September 1995, 37(3), p. 483.

Gautschi:1998:IGF

[Gau98] Walter Gautschi. The incomplete gamma functions since Tricomi. In Anonymous [Ano98b], pages 203–237. ISSN 0391-

805X. LCCN QA299.6. URL https://web.archive.org/web/
20070503103643/http://citeseer.ist.psu.edu/gautschi98incomplete.
html.

Gautschi:1999:NRC

[Gau99] Walter Gautschi. A note on the recursive calculation of incomplete gamma functions. ACM Transactions on Mathematical Software, 25(1):101–107, March 1999. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://

doi.acm.org/10.1145/305658.305717;   http://www.acm.
org/pubs/citations/journals/toms/cgi-bin/TOMSbibget?
Gautschi:1999:NRC;   http://www.acm.org:80/pubs/citations/
journals/toms/1999-25-1/p101-gautschi/.

Gautschi:2002:GQA

[Gau02] W. Gautschi. Gauss quadrature approximations to hypergeometric and confluent hypergeometric functions. Journal of Computational and Applied Mathematics, 139(1):173–187, February 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Gautschi:2011:LWF

[Gau11] Walter Gautschi. The Lambert W -functions and some of their integrals: a case study of high-precision computation. Numerical Algorithms, 57(1):27–34, May 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
1017-1398&volume=57&issue=1&spage=27.

Gautschi:2016:AER

[Gau16] Walter Gautschi. Algorithm 957: Evaluation of the repeated integral of the coerror function by half-range Gauss–Hermite quadrature. ACM Transactions on Mathematical Software, 42 (1):9:1–9:10, February 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Galue:2003:FRG

[GAZK03] L. Galué, A. Al-Zamel, and Shyam L. Kalla. Further results on generalized hypergeometric functions. Applied Mathematics and Computation, 136(1):17–25, March 25, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Galant:1968:HAG

[GB68] D. C. Galant and P. F. Byrd. High accuracy gamma function values for some rational arguments (in Technical Notes and Short Papers). Mathematics of Computation, 22(104):885–887, October 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Germain-Bonne:1990:CAN

[GB90] B. Germain-Bonne. Convergence acceleration of numbermachine sequences. Journal of Computational and Applied Mathematics, 32(1–2):83–88, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Extrapolation and rational approximation (Luminy, 1989).

Gal:1991:AEM

[GB91] Shmuel Gal and Boris Bachelis. An accurate elementary mathematical library for the IEEE floating point standard. ACM Transactions on Mathematical Software, 17(1):26–45, March 1991. CODEN ACMSCU. ISSN 0098-3500 (print),

1557-7295 (electronic). URL http://www.acm.org/pubs/toc/
Abstracts/toms/103151.html.

Guralnik:2009:ISV

[GBKK09] Elena Guralnik, Ariel J. Birnbaum, Anatoly Koyfman, and Avi Kaplan. Implementation specific verification of divide and square root instructions. In Bruguera et al. [BCDH09], pages 114–121. ISBN 0-7695-3670-0, 1-4244-4329-6. ISSN 1063-6889. LCCN QA76.6 .S887 2009. URL http://www.ac.usc.

es/arith19/.

Green:2019:DFE

[GBS19] Kevin R. Green, Tanner A. Bohn, and Raymond J. Spiteri. Direct function evaluation versus lookup tables: When to use which? SIAM Journal on Scientific Computing, 41(3):C194– C218, ???? 2019. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Gautschi:1964:EIR

[GC64] Walter Gautschi and William F. Cahill. Exponential integral and related functions. In Abramowitz and Stegun [AS64], pages 227–252. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book

is also available online at http://www.convertit.com/Go/
ConvertIt/Reference/AMS55.ASP in bitmap image format.

Goldstein:1967:CSB

[GCP67] Max Goldstein, C. W. Clenshaw, and Susan M. Picken. Chebyshev series for Bessel functions of fractional order. Mathematics of Computation, 21(99):509–??, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ge:1989:OCL

[Ge89] Renpu Ge. Optimal choice of linear interval extension. Applied Mathematics and Computation, 30(2):165–189, March 1989. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Gebali:2004:EAF

[GEK04] Fayez Gebali and Mohamed Watheq El-Kharashi. ERL: an algorithm for fast evaluation of exponential, reciprocal, and logarithmic functions. In Wahdan [Wah04], pages 269–272. ISBN 0-7803-8575-6. LCCN TK7801 .I5125 2004. IEEE catalog number 04EX893.

Gellman:1951:CCH

[Gel51] Harvey Gellman. The calculation of complex hypergeometric functions with the IBM Type 602-A calculating punch. In Proceedings of a Computation Seminar [IBM Department of Education, Endicot, NY, from December 5 to 9, 1949], pages 161–168. IBM, New York, NY, USA, 1951.

Gillman:1988:ARG

[GF88] E. Gillman and H. R. Fiebig. Accurate recursive generation of spherical Bessel and Neumann functions for a large range of indices. Computers in Physics, 2(1):62–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168296.

Ghanem:1997:SBF

[GF97] Riadh Ben Ghanem and Clément Frappier. Spherical Bessel functions and explicit quadrature formula. Mathematics of Computation, 66(217):289–296, January 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/jourcgi/jour-pbprocess?fn=110&arg1=S0025-
5718-97-00794-1&u=/mcom/1997-66-217/.

Gautschi:2008:LGW

[GG08] Walter Gautschi and Carla Giordano. Luigi Gatteschi’s work on asymptotics of special functions and their zeros. Numerical Algorithms, 49(1–4):11–31, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
1017-1398&volume=49&issue=1&spage=11.

Gautschi:2003:EEI

[GHT03] W. Gautschi, F. E. Harris, and N. M. Temme. Expansions of the exponential integral in incomplete gamma functions. Applied Mathematics Letters, 16(7):1095–1099, October 2003. CODEN AMLEEL. ISSN 0893-9659 (print), 1873-5452 (electronic).

Garvan:2001:SCN

[GI01] Frank (Frank G.) Garvan and Mourad Ismail, editors. Symbolic Computation, Number Theory, Special Functions, Physics, and Combinatorics, volume 4 of Developments in mathematics. Kluwer Academic Publishers, Dordrecht, The Netherlands, 2001. ISBN 1-4020-0101-0. x + 283 pp. LCCN QA295 .S86

Graillat:2023:PCH

[GIJL23] Stef Graillat, Youness Ibrahimy, Clothilde Jeangoudoux, and Christoph Lauter. A parallel compensated Horner scheme for SIMD architecture. In IEEE [IEE23], pages 131–138. LCCN ????

Giles:2016:AAI

[Gil16] Michael B. Giles. Algorithm 955: Approximation of the inverse Poisson cumulative distribution function. ACM Transactions on Mathematical Software, 42(1):7:1–7:22, February 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gimbutas:2020:EAF

[GJL20] Zydrunas Gimbutas, Shidong Jiang, and Li-Shi Luo. Evaluation of Abramowitz functions in the right half of the complex plane. Journal of Computational Physics, 405(??):Article 109169, March 15, 2020. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119308745.

Gautschi:1970:RCC

[GK70a] Walter Gautschi and Bruce J. Klein. Recursive computation of certain derivatives — a study of error propagation. Communications of the ACM, 13(1):7–9, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gautschi:1970:RAD

[GK70b] Walter Gautschi and Bruce J. Klein. Remark on Algorithm 282, Derivatives of e^{x}/x, cos(x)/x, and sin(x)/x. Communications of the ACM, 13(1):53–54, January 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gau66a].

Gersch:1989:SPT

[GK89a] W. Gersch and G. Kitagawa. Smoothness priors transfer function estimation. Automatica: the journal of IFAC, the International Federation of Automatic Control, 25(4):603–608, July 1989. CODEN ATCAA9. ISSN 0005-1098 (print), 1873-2836 (electronic).

Guo:1989:RCS

[GK89b] S.-S. Guo and M. K. Khan. On the rate of convergence of some operators on functions of bounded variation. Journal of Approximation Theory, 58(1):90–101, July 1989. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Galue:2000:MTG

[GKK00] L. Galué, H. G. Khajah, and Shyam L. Kalla. Multiplication theorems for generalized and double-index Bessel functions. Journal of Computational and Applied Mathematics, 118(1–2):143– 150, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042700002855.

Gil:2004:CRZ

[GKS04] Amparo Gil, Wolfram Koepf, and Javier Segura. Computing the real zeros of hypergeometric functions. Numerical Algorithms, 36(2):113–134, June 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp009.kluweronline.com/IPS/content/ext/x/J/5058/
I/54/A/2/abstract.htm.

Giordano:1983:EAZ

[GL83] C. Giordano and A. Laforgia. Elementary approximations for zeros of Bessel functions. Journal of Computational and Applied Mathematics, 9(3):221–228, September 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Giordano:2001:IMP

[GL01] C. Giordano and A. Laforgia. Inequalities and monotonicity properties for the gamma function. Journal of Computational and Applied Mathematics, 133(1–2):387–396, August 1, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042700006592.

Glasser:1966:ESI

[Gla66] M. L. Glasser. Evaluation of some integrals involving the ψfunction (in Technical Notes and Short Papers). Mathematics of Computation, 20(94):332–333, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1971:EII

[Gla71] M. L. Glasser. An elliptic integral identity. Mathematics of Computation, 25(115):533–534, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1974:SDI

[Gla74] M. L. Glasser. Some definite integrals of the product of two Bessel functions of the second kind: (order zero). Mathematics of Computation, 28(126):613–615, April 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1979:NI

[Gla79] M. L. Glasser. A note on the integral ∫_{0}^{∞} t^{2α−1}(1 + t^{2})^{1−α−β}J_{ν}(x\sqrt{1 + t^{2}})dt. Mathematics of Computation, 33(146): 792–793, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Glasser:1981:CBS

[Gla81] M. L. Glasser. A class of Bessel summations. Mathematics of Computation, 37(156):499–501, October 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Graillat:2015:MRE

[GLM15] Stef Graillat, Vincent Lefèvre, and Jean-Michel Muller. On the maximum relative error when computing integer powers by iterated multiplications in floating-point arithmetic. Numerical Algorithms, 70(3):653–667, November 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

link.springer.com/article/10.1007/s11075-015-9967-8.

Glover:1989:THN

[Glo89] Keith Glover. A tutorial on Hankel-norm approximation. In Jan C. Willems, editor, From data to model, pages 26–48. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-51571-2. LCCN QA279 .F76 1989.

Giordano:1998:UTG

[GLP98] C. Giordano, A. Laforgia, and J. Pecarić. Unified treatment of Gautschi–Kershaw type inequalities for the gamma function. Journal of Computational and Applied Mathematics, 99 (1–2):167–175, November 16, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S037704279800154X.

Glaser:2007:FAC

[GLR07] Andreas Glaser, Xiangtao Liu, and Vladimir Rokhlin. A fast algorithm for the calculation of the roots of special functions.

SIAM Journal on Scientific Computing, 29(4):1420–1438, ???? 2007. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Graillat:2015:ECF

[GLT^{+}15] Stef Graillat, Christoph Lauter, Ping Tak Peter Tang, Naoya Yamanaka, and Shin’ichi Oishi. Efficient calculations of faithfully rounded l_{2}-norms of n-vectors. ACM Transactions on Mathematical Software, 41(4):24:1–24:20, October 2015. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Glynn:1989:OSS

[Gly89] P. W. Glynn. Optimization of stochastic systems via simulation. In Edward A. MacNair, Kenneth J. Musselman, and Philip Heidelberger, editors, 1989 Winter Simulation Conference proceedings: December 4–6, 1989, the Capital Hilton Hotel, Washington, DC, pages 90–105. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-911801-58-8. LCCN QA76.9.C65

W56 1989. URL http://ieeexplore.ieee.org/servlet/

opac?punumber=5823. IEEE order no. 89CH2778-9.

Glaeske:1984:LTS

[GM84] H.-J. Glaeske and O. I. Maričev. The Laguerre transform of some elementary functions. Zeitschrift für Analysis und ihre Anwendungen, 3(3):237–244, 1984. ISSN 0232-2064 (print), 1661-4534 (electronic).

Guerrero:1988:HOT

[GM88] Antonio L. Guerrero and Pablo Martin. Higher order twopoint quasi-fractional approximations to the Bessel functions J_{0}(x) and J_{1}(x). Journal of Computational Physics, 77(1):276– 281, July 1988. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999188901684.

Guseinov:2004:EIG

[GM04] I. I. Guseinov and B. A. Mamedov. Evaluation of incomplete gamma functions using downward recursion and analytical relations. Journal of Mathematical Chemistry, 36(4):341–346, August 2004. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-

8897 (electronic). URL http://link.springer.com/article/
10.1023/B:JOMC.0000044521.18885.d3.

Guseinov:2007:UTE

[GM07] I. I. Guseinov and B. A. Mamedov. Unified treatment for the evaluation of generalized complete and incomplete gamma functions. Journal of Computational and Applied Mathematics, 202(2):435–439, May 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042706001506.

Gonzalez-Morales:2013:NII

[GMMIDM13] M. J. González-Morales, R. Mahillo-Isla, and C. Dehesa-Martínez. A new integral identity involving the elliptic integral E(m). Applied Mathematics and Computation, 221(??): 568–570, September 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300313007303.

Gragg:1989:FSE

[GN89] W. Gragg and B. Neta. Fortran subroutines for the evaluation of the confluent hypergeometric functions. Technical Report NPS-MA-89-014, Department of Mathematics, Naval Postgraduate School, Monterey CA 93943, USA, 1989. Software available URL http://math.nps.navy.mil.

Goano:1995:ACC

[Goa95] Michele Goano. Algorithm 745: Computation of the complete and incomplete Fermi–Dirac integral. ACM Transactions on Mathematical Software, 21(3):221–232, September 1995. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/pubs/citations/journals/toms/
1995-21-3/p221-goano/. See remark [Goa97].

Goano:1997:RA7

[Goa97] Michele Goano. Remark on Algorithm 745. ACM Transactions on Mathematical Software, 23(2):295, June 1997. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Goa95].

Godfrey:2001:NCC

[God01] P. Godfrey. A note on the computation of the convergent Lanczos complex gamma approximation. Unpublished Web file.,

2001. URL http://my.fit.edu/~gabdo/gamma.txt.

Godunov:2020:ACC

[God20] A. Godunov. Algorithms for calculating correctly rounded exponential function in double-precision arithmetic. IEEE Transactions on Computers, 69(9):1388–1400, September 2020. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Goerg:2014:ULW

[Goe14] Georg M. Goerg. Usage of the Lambert W function in statistics. Annals of Applied Statistics, 8(4):2567, December 2014. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic).

URL http://projecteuclid.org/euclid.aoas/1419001755.

Gomes:1989:GGL

[Gom89] M. I. Gomes. Generalized Gumbel and likelihood ratio test statistics in the multivariate GEV model. Computational Statistics & Data Analysis, 7(3):259–267, February 1989. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic).

Goncarov:1952:EFC

[Gon52a] V. L. Gončarov. Elementary functions of a complex variable, pages 491–552. Gosudarstv. Izdat. Tehn-Teoret. Lit., Moscow-Leningrad, USSR, 1952.

Goncarov:1952:EFR

[Gon52b] V. L. Gončarov. Elementary functions of a real variable. limits of sequences and functions. The general concept of a function. In Encyclopaedia of elementary mathematics. Book III. Functions and limits (the foundations of analysis), pages 9–296. Gosudarstv. Izdat. Tehn-Teoret. Lit., Moscow-Leningrad, USSR, 1952.

Gonzaga:1989:ASL

[Gon89] Clóvis C. Gonzaga. An algorithm for solving linear programming programs in O(n^{3}L) operations. In Megiddo [Meg89b], pages 1–28. ISBN 0-387-96847-4. LCCN QA402.5 .P7851 1989. Proceedings of the conference held at the Asilomar conference center in Pacific Grove, California, March 1–4, 1987.

Gordon:1982:RAN

[Gor82] H. T. Gordon. Rough approximation numerical algorithms. Dr. Dobb’s Journal of Software Tools, 7(7):54–56, July 1982. CO-DEN DDJOEB. ISSN 1044-789X.

Gil:2023:EGF

[GOST23] Amparo Gil, Andrzej Odrzywołek, Javier Segura, and Nico M. Temme. Evaluation of the generalized Fermi–Dirac integral and its derivatives for moderate/large values of the parameters. Computer Physics Communications, 283(??):Article 108563, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046552200282X.

Gargantini:1964:RCA

[GP64] I. Gargantini and T. Pomentale. Rational Chebyshev approximations to the Bessel function integrals Ki_{s}(x). Communications of the ACM, 7(12):727–730, December 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Grodd:1984:REN

[GP84] Laurence W. Grodd and Charles M. Patton. ROM extends numerical function set of handheld computer. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 35(7):25–36, July 1984. CODEN HPJOAX.

ISSN 0018-1153. URL http://www.hpl.hp.com/hpjournal/
pdfs/IssuePDFs/1984-07.pdf.

Guo:2009:CLC

[GQ09] Senlin Guo and Feng Qi. A class of logarithmically completely monotonic functions associated with the gamma function. Journal of Computational and Applied Mathematics, 224(1):127–132, February 1, 2009. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042708001829.

Giordano:1992:FMC

[GR92] Carla Giordano and Lucia G. Rodonò. Further monotonicity and convexity properties of the zeros of cylinder functions. Journal of Computational and Applied Mathematics, 42(2):245–251, October 12, 1992. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037704279290078C.

Greengard:2019:AEI

[GR19] Philip Greengard and Vladimir Rokhlin. An algorithm for the evaluation of the incomplete gamma function. Advances in Computational Mathematics, 45(1):23–49, February 2019. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044

(electronic). URL https://link.springer.com/article/10.
1007/s10444-018-9604-x.

Gray:1961:BFI

[Gra61] Marion C. Gray. Bessel functions of integral order and complex argument. Communications of the ACM, 4(4):169, April 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gray:1963:AFI

[Gra63a] Malcolm D. Gray. Algorithm 213: Fresnel integrals. Communications of the ACM, 6(10):617, October 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Gra64] and related remark [Gra63b].

Gray:1963:RAE

[Gra63b] Malcolm D. Gray. Remark on algorithms 88, 89, and 90 evaluation of the Fresnel integrals. Communications of the ACM, 6(10):618, October 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gray:1964:CAF

[Gra64] Malcolm Gray. Certification of Algorithm 213: Fresnel integrals. Communications of the ACM, 7(11):661, November 1964. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gra63a, Gra63b].

Gray:2002:ARE

[Gra02] Norman Gray. Automatic reduction of elliptic integrals using Carlson’s relations. Mathematics of Computation, 71 (237):311–318, January 2002. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.ams.org/journal-getitem?pii=S0025-5718-01-01333-
3;   http://www.ams.org/mcom/2002-71-237/S0025-5718-
01-01333-3/S0025-5718-01-01333-3.dvi;   http://www.ams.
org/mcom/2002-71-237/S0025-5718-01-01333-3/S0025-5718-
01-01333-3.pdf;   http://www.ams.org/mcom/2002-71-237/
S0025-5718-01-01333-3/S0025-5718-01-01333-3.ps;   http:
//www.ams.org/mcom/2002-71-237/S0025-5718-01-01333-3/
S0025-5718-01-01333-3.tex.

Gil:2016:ACI

[GRAST16] Amparo Gil, Diego Ruiz-Antolín, Javier Segura, and Nico M. Temme. Algorithm 969: Computation of the incomplete gamma function for negative values of the argument. ACM Transactions on Mathematical Software, 43(3):26:1–26:9, November 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://dl.acm.org/citation.cfm?id=2972951.

Greenhill:1892:AEF

[Gre92] Alfred George Greenhill. The Applications of Elliptic Functions. MacMillan Publishing Company, New York, NY, USA, 1892. xi + 357 pp.

Greenhill:1959:AEF

[Gre59] Alfred George Greenhill. The Applications of Elliptic Functions. Dover, New York, NY, USA, 1959. xi + 357 pp. See [Gre92].

Greaves:1982:AHM

[Gre82] G. Greaves. An algorithm for the Hausdorff moment problem. Numerische Mathematik, 39(2):231–238, August 1982. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Greuel:2016:MSI

[Gre16] Gert-Martin Greuel, editor. Mathematical Software — ICMS 2016: 5th International Conference, Berlin, Germany, July 11– 14, 2016: proceedings, volume 9725 of Lecture Notes in Computer Science/Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-42431-9 (print), 3-319-42432-7 (electronic). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.9.M35. URL http://zbmath.org/?q=an:1342.68017.

Guyot:1996:STD

[GREL96] A. Guyot, M. Renaudin, B. El Hassan, and V. Levering. Self timed division and square-root extraction. In Proceedings of the Ninth International Conference on VLSI Design, 1996, pages 376–381. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. CODEN ???? ISSN ????

Gradshteyn:2007:TIS

[GRJZ07] I. S. Gradshteyn, I. M. Ryzhik, Alan Jeffrey, and Daniel Zwillinger. Table of Integrals, Series and Products. Academic Press, New York, NY, USA, seventh edition, 2007. ISBN 0-12-373637-4 (hardcover). xlv + 1171 pp. LCCN QA55 .G6613 2007.

Gervais:1987:RAF

[GRS87] R. Gervais, Q. I. Rahman, and G. Schmeisser. Representation and approximation of functions via (0, 2)-interpolation. Journal of Approximation Theory, 50(2):89–110, June 1987. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Gautschi:1978:CMB

[GS78] Walter Gautschi and Josef Slavik. On the computation of modified Bessel function ratios. Mathematics of Computation, 32 (143):865–875, July 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gatto:1981:NEM

[GS81] M. A. Gatto and J. B. Seery. Numerical evaluation of the modified Bessel functions I and K. Computers and Mathematics with Applications, 7(3):203–209, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0898122181900808.

Gawronski:1982:ACF

[GS82a] W. Gawronski and U. Stadtmüller. Approximation of continuous functions by generalized Favard operators. Journal of Approximation Theory, 34(4):384–396, April 1982. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Gessel:1982:SEH

[GS82b] Ira Gessel and Dennis Stanton. Strange evaluations of hypergeometric series. SIAM journal on mathematical analysis, 13 (2):295–308, March 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Gawronski:1988:ZLT

[GS88] Wolfgang Gawronski and Ulrich Stadtmüller. On the zeros of Lerch’s transcendental function of real parameters. Journal of Approximation Theory, 53(3):354–364, June 1988. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Gustafson:1991:CAA

[GS91] Sven-Åke Gustafson and Frank Stenger. Convergence acceleration applied to Sinc approximation with application to approximation of |x|^{α}. In Bowers and Lund [BL91], pages 161–171. ISBN 0-8176-3611-0. LCCN TA329 .C644 1991. US$65.00.

Gourdon:1999:NCC

[GS99] Xavier Gourdon and Pascal Sebah. Numbers, constants, and computation. Technical report, ????, Paris, France, 1999. URL

http://numbers.computation.free.fr/Constants/home.html.

World-Wide Web site.

Gil:2012:CRZ

[GS12] Amparo Gil and Javier Segura. Computing the real zeros of cylinder functions and the roots of the equation xC′_{ν}(x) + γC_{ν}(x) = 0. Computers and Mathematics with Applications, 64(1):11–21, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0898122112001460.

Gil:2014:CZA

[GS14a] Amparo Gil and Javier Segura. On the complex zeros of Airy and Bessel functions and those of their derivatives. Analysis and Applications (Singapore), 12(5):537–561, August 2014. ISSN 0219-5305 (print), 1793-6861 (electronic). Special Issue: Dedicated to the Memory of Frank Olver (Part II).

Greuel:2014:SIS

[GS14b] Gert-Martin Greuel and Wolfram Sperber. swMATH — an information service for mathematical software. In Hong and Yap [HY14], pages 691–701. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (e-book). LCCN QA76.9.M35.

Greynat:2014:NAE

[GS14c] David Greynat and Javier Sesma. A new approach to the epsilon expansion of generalized hypergeometric functions. Computer Physics Communications, 185(2):472–478, February 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046551300324X.

Gaudreau:2012:CTP

[GSS12] Philippe Gaudreau, Richard M. Slevinsky, and Hassan Safouhi. Computation of tail probabilities via extrapolation methods and connection with rational and Padé approximants. SIAM Journal on Scientific Computing, 34(1):B65–B85, January 2012. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/doi/abs/10.1137/100803778.

Gil:2002:CSF

[GST02a] A. Gil, J. Segura, and N. M. Temme. Computing special functions by using quadrature rules. Report MAS-R0230, CWI, Amsterdam, The Netherlands, 2002. 11 pp. Published in [GST03b].

Gil:2002:AAB

[GST02b] Amparo Gil, Javier Segura, and Nico M. Temme. Algorithm 819: AIZ, BIZ: two Fortran 77 routines for the computation of complex Airy functions. ACM Transactions on Mathematical Software, 28(3):325–336, September 2002. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2002:AGH

[GST02c] Amparo Gil, Javier Segura, and Nico M. Temme. Algorithm 822: GIZ, HIZ: two Fortran 77 routines for the computation of complex Scorer functions. ACM Transactions on Mathematical Software, 28(4):436–447, December 2002. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2002:EMB

[GST02d] Amparo Gil, Javier Segura, and Nico M. Temme. Evaluation of the modified Bessel function of the third kind of imaginary orders. Journal of Computational Physics, 175(2):398–411, January 20, 2002. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999101968949.

Gil:2003:CMB

[GST03a] Amparo Gil, Javier Segura, and Nico M. Temme. Computation of the modified Bessel function of the third kind of imaginary orders: uniform Airy-type asymptotic expansion. Journal of Computational and Applied Mathematics, 153(1–2):225–234, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042702006088.

Gil:2003:CSF

[GST03b] Amparo Gil, Javier Segura, and Nico M. Temme. Computing special functions by using quadrature rules. Numerical Algorithms, 33(1–4):265–275, August 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/ipsapp007.kluweronline.com/content/getfile/5058/46/
22/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/46/22/fulltext.pdf.

Gil:2004:AMB

[GST04a] Amparo Gil, Javier Segura, and Nico M. Temme. Algorithm 831: Modified Bessel functions of imaginary order and positive argument. ACM Transactions on Mathematical Software, 30 (2):159–164, June 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2004:CSM

[GST04b] Amparo Gil, Javier Segura, and Nico M. Temme. Computing solutions of the modified Bessel differential equation for imaginary orders and positive arguments. ACM Transactions on Mathematical Software, 30(2):145–158, June 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2006:ARP

[GST06a] Amparo Gil, Javier Segura, and Nico M. Temme. Algorithm 850: Real parabolic cylinder functions U(a, x), V (a, x). ACM Transactions on Mathematical Software, 32(1):102–112, March 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2006:CRP

[GST06b] Amparo Gil, Javier Segura, and Nico M. Temme. Computing the real parabolic cylinder functions U(a, x), V (a, x). ACM Transactions on Mathematical Software, 32(1):70–101, March 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2007:NMS

[GST07a] Amparo Gil, Javier Segura, and N. M. Temme. Numerical Methods for Special Functions. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2007. ISBN 0-89871-634-

9. xvi + 415 pp. LCCN QA351 .G455 2007. US$99.00.

Gil:2007:NSS

[GST07b] Amparo Gil, Javier Segura, and Nico M. Temme. Numerically satisfactory solutions of hypergeometric recursions. Mathematics of Computation, 76(259):1449–1468, July 2007. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/mcom/2007-76-259/S0025-5718-
07-01918-7/home.html;   http://www.ams.org/mcom/2007-
76-259/S0025-5718-07-01918-7/S0025-5718-07-01918-7.dvi;
 http://www.ams.org/mcom/2007-76-259/S0025-5718-07-01918-
7/S0025-5718-07-01918-7.pdf;   http://www.ams.org/mcom/
2007-76-259/S0025-5718-07-01918-7/S0025-5718-07-01918-
7.ps.

Gil:2011:APC

[GST11a] Amparo Gil, Javier Segura, and Nico M. Temme. Algorithm 914: Parabolic cylinder function W (a, x) and its derivative. ACM Transactions on Mathematical Software, 38(1):6:1–6:5, November 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2011:FAC

[GST11b] Amparo Gil, Javier Segura, and Nico M. Temme. Fast and accurate computation of the Weber parabolic cylinder function W (a, x). IMA Journal of Numerical Analysis, 31(3):1194–1216, July 2011. CODEN IJNADH. ISSN 0272-4979 (print), 1464-

3642 (electronic). URL http://imajna.oxfordjournals.org/
content/31/3/1194.full.pdf+html.

Gil:2012:EAA

[GST12a] Amparo Gil, Javier Segura, and Nico M. Temme. Efficient and accurate algorithms for the computation and inversion of the incomplete gamma function ratios. SIAM Journal on Scientific Computing, 34(6):A2965–A2981, ???? 2012. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Gil:2012:IAF

[GST12b] Amparo Gil, Javier Segura, and Nico M. Temme. An improved algorithm and a Fortran 90 module for computing the conical function p_{1/2+iτ}^{m} (x). Computer Physics Communications, 183(3):794–799, March 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465511003936.

Gil:2014:ACM

[GST14a] Amparo Gil, Javier Segura, and Nico M. Temme. Algorithm 939: Computation of the Marcum Q-function. ACM Transactions on Mathematical Software, 40(3):20:1–20:21, April 2014. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Gil:2014:RSD

[GST14b] Amparo Gil, Javier Segura, and Nico M. Temme. Recent software developments for special functions in the Santander–Amsterdam project. Science of Computer Programming, 90 (part A)(??):42–54, September 15, 2014. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167642313002888.

Gil:2015:CKF

[GST15a] Amparo Gil, Javier Segura, and Nico M. Temme. Computing the Kummer function U(a, b, z) for small values of the arguments. Applied Mathematics and Computation, 271(??): 532–539, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300315012837.

Gil:2015:GPI

[GST15b] Amparo Gil, Javier Segura, and Nico M. Temme. GammaCHI: a package for the inversion and computation of the gamma and chi-square cumulative distribution functions (central and noncentral). Computer Physics Communications, 191(??):132– 139, June 2015. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465515000107.

Gil:2017:ECL

[GST17] Amparo Gil, Javier Segura, and Nico M. Temme. Efficient computation of Laguerre polynomials. Computer Physics Communications, 210(??):124–131, January 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465516302727.

Gil:2020:NEA

[GST20] A. Gil, J. Segura, and N. M. Temme. Numerical evaluation of Airy-type integrals arising in uniform asymptotic analysis. Journal of Computational and Applied Mathematics, 371(??):Article 112717, June 2020. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S037704272030008X.

Goldstein:1958:BFL

[GT58] M. Goldstein and R. M. Thaler. Bessel functions for large arguments. Mathematical Tables and Other Aids to Computation, 12 (61):18–26, January 1958. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Goldstein:1959:RTC

[GT59] M. Goldstein and R. M. Thaler. Recurrence techniques for the calculation of Bessel functions. Mathematical Tables and Other Aids to Computation, 13(66):102–108, April 1959. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Gunn:1965:AZV

[Gun65a] J. H. Gunn. Algorithm 252 [Z]: Vector coupling or Clebsch– Gordan coefficients. Communications of the ACM, 8(4):217, April 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gunn:1965:ASa

[Gun65b] J. H. Gunn. Algorithm 260: 6-J symbols. Communications of the ACM, 8(8):492, August 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gunn:1965:ASb

[Gun65c] J. H. Gunn. Algorithm 261: 9-J symbols. Communications of the ACM, 8(8):492–493, August 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gunn:1967:ACW

[Gun67] J. H. Gunn. Algorithm 300: Coulomb wave functions. Communications of the ACM, 10(4):244–245, April 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Vos73].

Gupta:1960:OSG

[Gup60] Shanti S. Gupta. Order statistics from the gamma distribution. Technometrics, 2(2):243–262, May 1960. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://

www.jstor.org/stable/1266548. See errata [Ano60b].

Gustafson:1966:CAM

[Gus66] Sven-Åke Gustafson. Convergence acceleration by means of numerical quadrature. Nordisk tidskrift for informationsbehandling, 6(2):117–128, June 1966. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=6&issue=2&spage=117.

Gustafson:1978:ATC

[Gus78a] S.-Å. Gustafson. Algorithm 38. Two computer codes for convergence acceleration. Computing: Archiv fur informatik und numerik, 21(1):87–91, 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gustafson:1978:CAG

[Gus78b] S.-Å. Gustafson. Convergence acceleration on a general class of power series. Computing: Archiv fur informatik und numerik, 21(1):53–69, 1978. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Gustafson:1984:SCC

[Gus84] Sven-Åke Gustafson. On the stability of a class of convergence acceleration methods for power series. BIT (Nordisk tidskrift for informationsbehandling), 24(4):510–519, December 1984. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0006-3835&volume=
24&issue=4&spage=510.

Gustafson:1985:SCA

[Gus85] Sven-Åke Gustafson. Stable convergence acceleration using Laplace transforms. Numerische Mathematik, 47(3):387–394, November 1985. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Gustavson:1986:FEF

[Gus86] F. G. Gustavson. Fast elementary function algorithms for 370 machines. In Miranker and Toupin [MT86], page 17. CO-DEN LNCSD9. ISBN 0-387-16798-6 (USA: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.95 .A231 1986. Symposium sponsored by IBM Deutschland.

Glasser:1971:CFE

[GW71] M. L. Glasser and V. E. Wood. A closed form evaluation of the elliptic integral. Mathematics of Computation, 25(115):535– 536, July 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gray:1991:GMA

[GW91] H. L. Gray and Suojin Wang. A general method for approximating tail probabilities. Journal of the American Statistical Association, 86(413):159–166, March 1991. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

www.jstor.org/stable/2289726.

Ginsberg:1975:AAD

[GZ75] E. S. Ginsberg and Dorothy Zaborowski. ACM Algorithm 490: The dilogarithm function of a real argument [S22]. Communications of the ACM, 18(4):200–202, April 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Mor76].

Gao:2022:TFI

[GZC22] Zhanyuan Gao, Laiping Zhao, and Haonan Chen. A trigonometric function instruction set extension method based on RISC-V. In IEEE, editor, 2022 IEEE/ACIS 22nd International Conference on Computer and Information Science (ICIS), pages 119– 126. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2022.

Hull:1985:PRV

[HA85] T. E. Hull and A. Abrham. Properly rounded variable precision square root. ACM Transactions on Mathematical Software, 11(3):229–237, September 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org/pubs/citations/journals/toms/1985-11-3/p229-
hull/;   http://www.acm.org/pubs/toc/Abstracts/toms/
214413.html.

Hull:1986:VPE

[HA86] T. E. Hull and A. Abrham. Variable precision exponential function. ACM Transactions on Mathematical Software, 12(2):79–91, June 1986. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.

org/pubs/citations/journals/toms/1986-12-2/p79-hull/
;   http://www.acm.org/pubs/toc/Abstracts/toms/6498.
html.

Haimovici:1960:MRE

[Hai60] Corina Haimovici. A method of representing elementary functions in algebras of finite order. (Romanian). An. Şti. Univ. “Al. I. Cuza” Iaşi Secţ I. (N.S.), 6:507–515, 1960.

Hamza:1991:MBD

[HAK91] K. M. Hamza and M. A. H. Abdul-Karim. Microprocessor based direct square root extractor. Modelling, 34(1):45–48, ???? 1991.

Hamaker:1978:MAC

[Ham78] Hugo C. Hamaker. Miscellanea: Approximating the cumulative normal distribution and its inverse. Applied Statistics, 27(1):76– 77, January 1978. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Hartree:1949:NIP

[Har49] D. R. Hartree. Notes on iterative processes. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 45(2):230–236, April 1949. CODEN PCPSA4. ISSN 0008-

1981. URL https://ui.adsabs.harvard.edu/abs/1949PCPS.

..45..230H. Correction: in equation (29) at the bottom of page 233, replace the denominator term 2y_{0} by 2y_{1}, matching the denominator in equation (32) on page 234.

Hart:1966:CAR

[Har66] Roger G. Hart. A close approximation related to the error function (in Technical Notes and Short Papers). Mathematics of Computation, 20(96):600–602, October 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Harris:1977:CAT

[Har77] F. E. Harris. Convergence acceleration technique for lattice sums arising in electronic-structure studies of crystalline solids. Journal of Mathematical Physics, 18(12):2377–2381, December 1977.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hare:1997:CPB

[Har97a] D. E. G. Hare. Computing the principal branch of log-Gamma. Journal of Algorithms, 25(2):221–236, November 1997. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0196677497908816.

Harris:1997:NAC

[Har97b] Frank E. Harris. New approach to calculation of the leaky aquifer function. International Journal of Quantum Chemistry, 63(5):913–916, ???? 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract?ID=42641;   http:
//www3.interscience.wiley.com/cgi-bin/fulltext?ID=42641&
PLACEBO=IE.pdf.

Harris:1998:MAL

[Har98] Frank E. Harris. More about the leaky aquifer function. International Journal of Quantum Chemistry, 70(4–5):623–626, ???? 1998. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X

(electronic). URL http://www3.interscience.wiley.com/
cgi-bin/abstract?ID=75040;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=75040&PLACEBO=IE.pdf.

Harris:2000:SBE

[Har00] Frank E. Harris. Spherical Bessel expansions of sine, cosine, and exponential integrals. Applied Numerical Mathematics: Transactions of IMACS, 34(1):95–98, June 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460

(electronic). URL http://www.elsevier.nl/gej-ng/29/17/
21/62/27/32/abstract.html;   http://www.elsevier.nl/
gej-ng/29/17/21/62/27/32/article.pdf;   http://www.sciencedirect.
com/science/article/pii/S0168927499000318.

Harris:2001:KFL

[Har01] Frank E. Harris. On Kryachko’s formula for the leaky aquifer function. International Journal of Quantum Chemistry, 81(5): 332–334, ???? 2001. CODEN IJQCB2. ISSN 0020-7608 (print),

1097-461X (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/76507286/START;   http://

www3.interscience.wiley.com/cgi-bin/fulltext/76507286/
FILE?TPL=ftx_start;   http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=76507286&PLACEBO=IE.pdf.

Harrison:2003:FVS

[Har03] John Harrison. Formal verification of square root algorithms. Formal Methods in System Design, 22(2):143–153, March 2003. CODEN FMSDE6. ISSN 0925-9856 (print),

1572-8102 (electronic). URL https://dl.acm.org/doi/abs/
10.1023/A:1022973506233.

Harris:2008:IBG

[Har08] Frank E. Harris. Incomplete Bessel, generalized incomplete gamma, or leaky aquifer functions. Journal of Computational and Applied Mathematics, 215(1):260–269, 2008. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042707002014.

Harrison:2009:DTB

[Har09a] John Harrison. Decimal transcendentals via binary. In Bruguera et al. [BCDH09], pages 187–194. ISBN 0-7695-3670-0, 1-4244-4329-6. ISSN 1063-6889. LCCN QA76.6 .S887 2009. URL http:

//www.ac.usc.es/arith19/.

Harrison:2009:FAB

[Har09b] John Harrison. Fast and accurate Bessel function computation. In Bruguera et al. [BCDH09], pages 104–113. ISBN 0-7695-3670-0, 1-4244-4329-6. ISSN 1063-6889. LCCN QA76.6 .S887 2009.

URL http://www.ac.usc.es/arith19/.

Harvey:2014:SAC

[Har14] David Harvey. A subquadratic algorithm for computing the n-th Bernoulli number. Mathematics of Computation, 83 (289):2471–2477, 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.ams.org/

journals/mcom/2014-83-289/S0025-5718-2014-02832-9;
http://www.ams.org/journals/mcom/2014-83-289/S0025-5718-
2014-02832-9/S0025-5718-2014-02832-9.pdf.

Harper:2020:AGH

[Har20] J. F. Harper. Asymptotics of a Gauss hypergeometric function with two large parameters: a new case. The ANZIAM Journal, 62(4):446–452, October 2020. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

URL https://www.cambridge.org/core/journals/anziam-
journal/article/asymptotics-of-a-gauss-hypergeometric-
function-with-two-large-parameters-a-new-case/32B40986E7DB85F500FC9024F846E527.

Hastings:1955:ADC

[Has55] Cecil Hastings, Jr. Approximations for Digital Computers. The Rand series. Princeton University Press, Princeton, NJ, USA, 1955. ISBN 0-691-07914-5. viii + 201 pp. LCCN QA76 .H37. Assisted by Jeanne T. Hayward and James P. Wong, Jr.

Hastings:1966:RCB

[Has66] C. W. Hastings. R66-78 computation of the base two logarithm of binary number. IEEE Transactions on Electronic Computers, EC-15(6):956–957, December 1966. CODEN IEECA8. ISSN

0367-7508. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4038956.

Hashemian:1990:SRA

[Has90] R. Hashemian. Square rooting algorithms for integer and floating-point numbers. IEEE Transactions on Computers, C-39 (8):1025–1029, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hassibi:2000:ESR

[Has00a] B. Hassibi. An efficient square-root algorithm for BLAST. In Proceedings. 2000 IEEE International Conference on Acoustics, Speech, and Signal Processing: ICASSP ’00, 5–9 June 2000, volume 2, pages II737–II740. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. CODEN ???? ISSN ????

Hassibi:2000:FSR

[Has00b] B. Hassibi. A fast square-root implementation for BLAST. In Conference Record of the Thirty-Fourth Asilomar Conference on Signals, Systems and Computers, 2000, volume 2, pages 1255– 1259. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. CODEN ???? ISSN ????

Hassenpflug:2002:EAS

[Has02] W. C. Hassenpflug. Error analysis in the series evaluation of the exponential type integral e^{z}E_{1}(z). Computers and Mathematics with Applications, 43(1–2):207–266, January 2002. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122101002838.

Hautot:1988:CAC

[Hau88] A. Hautot. Convergence acceleration of continued fractions of Poincaré type. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4):309–322, June 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Haavie:1979:GNT

[Håv79] Tore Håvie. Generalized Neville type extrapolation schemes. BIT (Nordisk tidskrift for informationsbehandling), 19(2): 204–213, June 1979. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=19&issue=2&spage=204.

Haavie:1981:RUT

[Håv81] Tore Håvie. Remarks on a unified theory for classical and generalized interpolation and extrapolation. BIT (Nordisk tidskrift for informationsbehandling), 21(4):465–474, December 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0006-3835&volume=
21&issue=4&spage=465.

Hawkes:1982:ANT

[Haw82] Alan G. Hawkes. Approximating the normal tail. Journal of the Royal Statistical Society. Series D (The Statistician), 31(3):231– 236, September 1982. CODEN ???? ISSN 0039-0526 (print),

1467-9884 (electronic). URL http://www.jstor.org/stable/
2987989.

Haynes:1956:EIE

[Hay56] John G. Haynes. Evaluation of incomplete elliptic integrals by Gaussian integration. In ????, editor, ACM’56: Proceedings of the 1956 11th ACM national meeting, pages 56–59. ACM Press, New York, NY 10036, USA, 1956. URL https://dl.acm.org/

ft_gateway.cfm?id=808948.

Hayashi:1999:SRR

[Hay99] Takao Hayashi. A set of rules for the root-extraction prescribed by the sixteenth-century Indian mathematicians, Nīlakan.t.ha Somastuvan and Śaṅkara Vāriyar. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(2):135–153, November 1999. CODEN HISCDU. ISSN 0285-4821.

Headley:1975:DZG

[HB75] V. B. Headley and V. K. Barwell. On the distribution of the zeros of generalized Airy functions. Mathematics of Computation, 29 (131):863–877, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Henner:2009:MMP

[HBF09] Victor Henner, Tatyana Belozerova, and Kyle Forinash. Mathematical Methods in Physics: Partial Differential Equations, Fourier Series, and Special Functions. A. K. Peters, Ltd., Wellesley, MA, USA, 2009. ISBN 1-56881-335-X. xviii + 841 pp. LCCN QC20 .H487 2009.

Holmgren:2000:CAL

[HBS00] Sverker Holmgren, Henrik Brandén, and Erik Sterner. Convergence acceleration for the linearized Navier–Stokes equations using semicirculant approximations. SIAM Journal on Scientific Computing, 21(4):1524–1550, July 2000. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http://

epubs.siam.org/sam-bin/dbq/article/31798.

Han:2009:ICS

[HCK09] Dong-Guk Han, Dooho Choi, and Howon Kim. Improved computation of square roots in specific finite fields. IEEE Transactions on Computers, 58(2):188–196, February 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4663058.

Hart:1968:CAa

[HCL^{+}68a] John F. Hart, E. W. Cheney, Charles L. Lawson, Hans J. Maehly, Charles K. Mesztenyi, John R. Rice, Henry G. Thatcher, Jr., and Christoph Witzgall. Computer Approximations. Robert E. Krieger Publishing Company, Huntington, NY, USA, 1968. ISBN 0-88275-642-7. x + 343 pp. LCCN QA 297 C64 1978. Reprinted 1978 with corrections.

Hart:1968:CAb

[HCL^{+}68b] John F. Hart, E. W. Cheney, Charles L. Lawson, Hans J. Maehly, Charles K. Mesztenyi, John R. Rice, Henry G. Thatcher, Jr., and Christoph Witzgall. Computer Approximations. The SIAM series in applied mathematics. Wiley, New York, NY, USA, 1968. ISBN 0-471-35630-1. x + 343 pp. LCCN QA297 .C64.

Hill:1973:AAN

[HD73] G. W. Hill and A. W. Davis. ACM Algorithm 442: Normal deviate [S14]. Communications of the ACM, 16(1):51–52, January 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Higham:2015:PCA

[HDG^{+}15] Nicholas J. Higham, Mark R. Dennis, Paul Glendinning, Paul A. Martin, Fadil Santosa, and Jared Tanner, editors. The Princeton Companion to Applied Mathematics. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-15039-7 (hardcover). 994 (est.) pp. LCCN QA155 .P75 2015.

Heatley:1965:ETT

[Hea65] A. H. Heatley. An extension of the table of the Toronto function (in Technical Notes and Short Papers). Mathematics of Computation, 19(89):118–123, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Heilmann:1988:SSM

[Hei88] Margareta Heilmann. L_{p}-saturation of some modified Bernstein operators. Journal of Approximation Theory, 54(3):260–273, September 1988. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Heinrich:1996:AAF

[Hei96] Peter Heinrich. Algorithm alley: a fast integer square root. Dr. Dobb’s Journal of Software Tools, 21(4):113–114, 130, April 1996. CODEN DDJOEB. ISSN 1044-789X.

Henrici:1977:ART

[Hen77] Peter Henrici. Analytische Rechenverfahren für den Taschenrechner HP-25. (German) [Analytical Calculations for the HP-25 Calculator]. Oldenbourg, München, West Germany, 1977. ISBN 0-471-02938-6. 280 pp. LCCN ????

Herndon:1961:ASN

[Her61a] J. R. Herndon. Algorithm 49: Spherical Neumann function. Communications of the ACM, 4(4):179, April 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Col78].

Herndon:1961:AEC

[Her61b] John R. Herndon. Algorithm 46: Exponential of a complex number. Communications of the ACM, 4(4):178, April 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herndon:1961:AGF

[Her61c] John R. Herndon. Algorithm 54: Gamma function for range 1 to 2. Communications of the ACM, 4(4):180, April 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herndon:1961:ACEa

[Her61d] John R. Herndon. Algorithm 55: Complete elliptic integral of the first kind. Communications of the ACM, 4(4):180, April 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herndon:1961:ACEb

[Her61e] John R. Herndon. Algorithm 56: Complete elliptic integral of the second kind. Communications of the ACM, 4(4):180–181, April 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herndon:1961:ABF

[Her61f] John R. Herndon. Algorithm 57: Ber or Bei function. Communications of the ACM, 4(4):181, April 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Herman:1969:NHE

[Her69] G. T. Herman. A new hierarchy of elementary functions. Proceedings of the American Mathematical Society, 20:557–562, 1969. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Herman:1971:EDH

[Her71] G. T. Herman. The equivalence of different hierarchies of elementary functions. Z. Math. Logik Grundlagen Math., 17:219– 224, 1971.

Hermann:1982:SAG

[Her82] Robert Hermann. Some algebraic, geometric, and systemtheoretic properties of the Special Functions of mathematical physics. Journal of Mathematical Physics, 23(7):1282–1294, July 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v23/i7/p1282_s1.

Hobson:1995:EMR

[HF95] R. F. Hobson and M. W. Fraser. An efficient maximumredundancy radix-8 SRT division and square-root method. IEEE Journal of Solid-State Circuits, 30(1):29–38, January 1995. CO-DEN IJSCBC. ISSN 0018-9200 (print), 1558-173X (electronic).

Harris:2009:MIB

[HF09] Frank E. Harris and J. G. Fripiat. Methods for incomplete Bessel function evaluation. International Journal of Quantum Chemistry, 109(8):1728–1740, February 2009. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hyogo:1999:LVF

[HFS99] A. Hyogo, Y. Fukutomi, and K. Sekine. Low voltage fourquadrant analog multiplier using square-root circuit based on CMOS pair. In Proceedings of the 1999 IEEE International Symposium on Circuits and Systems: ISCAS ’99, 2 June 1999, volume 2, pages 274–277. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. CODEN ???? ISSN ????

Hull:1994:ICE

[HFT94] T. E. Hull, Thomas F. Fairgrieve, and Ping Tak Peter Tang. Implementing complex elementary functions using exception handling. ACM Transactions on Mathematical Software, 20(2): 215–244, June 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/

pubs/citations/journals/toms/1994-20-2/p215-hull/. See

correction [Ano94], and improved analysis, tightened bounds, and exhibition of worst cases for complex square roots [JM17].

Haynsworth:1964:BEP

[HG64] Emilie V. Haynsworth and Karl Goldberg. Bernoulli and Euler polynomials, Riemann Zeta function. In Abramowitz and Stegun [AS64], pages 803–820. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://www.convertit.com/

Go/ConvertIt/Reference/AMS55.ASP in bitmap image format.

Hanson:2018:RAM

[HH18] Richard J. Hanson and Tim Hopkins. Remark on Algorithm 539: A Modern Fortran Reference Implementation for Carefully Computing the Euclidean Norm. ACM Transactions on Mathematical Software, 44(3):24:1–24:23, April 2018. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/citation.cfm?id=3134441.

Hasan:2000:FPI

[HHR00] M. A. Hasan, A. A. Hasan, and S. Rahman. Fixed point iterations for computing square roots and the matrix sign function of complex matrices. In Proceedings of the 39th IEEE Conference on Decision and Control, volume 5, pages 4253–4258. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. CODEN ???? ISSN ????

Hemker:1973:SDL

[HHv^{+}73] P. W. Hemker, W. Hoffmann, S. P. N. van Kampen, H. L. Oudshoorn, and D. T. Winter. Single- and double-length computation of elementary functions. Technical Report NW 7, Mathematical Centre, Amsterdam, 1973.

Higgins:1977:CBP

[Hig77] John Rowland Higgins. Completeness and basis properties of sets of special functions. Cambridge University Press, Cambridge, UK, 1977. ISBN 0-521-21376-2 (hardcover), 0-521-60488-5 (paperback). x + 134 pp. LCCN ????

Higginbotham:1993:ISR

[Hig93] T. F. Higginbotham. The integer square root of N via a binary search. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 25(4):41–45, December 1993. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Higham:2018:UN

[Hig18] Nicholas J. Higham. The unwinding number. SIAM News, 51 (8):??, October 2018. ISSN 0036-1437. URL https://sinews.

siam.org/Details-Page/the-unwinding-number.

Hill:1970:AASa

[Hil70a] G. W. Hill. ACM Algorithm 395: Student’s t-distribution. Communications of the ACM, 13(10):617–619, October 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [eL79, Hil81a].

Hill:1970:AASb

[Hil70b] G. W. Hill. ACM Algorithm 396: Student’s t-quantiles. Communications of the ACM, 13(10):619–620, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Hil81a, Hil81b, eL79].

Hill:1973:AAS

[Hil73a] G. W. Hill. ACM Algorithm 465: Student’s t frequency [S14]. Communications of the ACM, 16(11):690, November 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hill:1973:SAA

[Hil73b] I. D. Hill. Statistical algorithms: Algorithm AS 66: The normal integral. Applied Statistics, 22(3):424–427, September 1973. CO-DEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

URL http://lib.stat.cmu.edu/apstat/66.

Hill:1977:AIB

[Hil77] G. W. Hill. Algorithm 518: Incomplete Bessel function I_{0}. The von Mises distribution [S14]. ACM Transactions on Mathematical Software, 3(3):279–284, September 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Hill:1981:RSD

[Hil81a] G. W. Hill. Remark on “Algorithm 395: Student’s t-Distribution”. ACM Transactions on Mathematical Software, 7(2):247–249, June 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Hil70a, Hil70b, eL79].

Hill:1981:RSQ

[Hil81b] G. W. Hill. Remark on “Algorithm 396: Student’s t-Quantiles”. ACM Transactions on Mathematical Software, 7(2):250–251, June 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Hil70b].

Hinden:1977:PAR

[Hin77] Harvey J. Hinden. Phi again: a relationship between the golden ratio and the limit of a ratio of modified Bessel functions. Fibonacci Quarterly, 15(2):112, 152, April 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-2/hinden-a.pdf.

Hitchcock:1957:PAB

[Hit57] A. J. M. Hitchcock. Polynomial approximations to Bessel functions of order zero and one and to related functions. Mathematical Tables and Other Aids to Computation, 11(58):86–88, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853

(electronic). URL http://www.jstor.org/stable/2002156.

Hitotumatu:1974:NMC

[Hit74] Sin Hitotumatu. A new method for the computation of square root, exponential and logarithmic functions through hyperbolic CORDIC. Revue d’Analyse Numérique et de la Théorie de l’Approximation, 3(2):173–180, 1974. ISSN 1010-3376 (print),

2457-8118 (electronic). URL https://ictp.acad.ro/jnaat/
journal/article/view/1974-vol3-no2-art7.

Hitotumatu:1975:SRU

[Hit75] Sin Hitotumatu. Some remarks on the unified treatment of elementary functions by microprogramming. In Miller [Mil75a], pages 51–56 (vol. 2). ISBN 0-12-496950-X. LCCN QA297 .R69 1973.

Hill:1967:ANCa

[HJ67a] I. D. Hill and S. A. Joyce. Algorithm 304: Normal curve integral. Communications of the ACM, 10(6):374–375, June 1967. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [HJ67b, Ber68].

Hill:1967:RAS

[HJ67b] I. D. Hill and S. A. Joyce. Remarks on Algorithm 123 [S15]: Real error function, ERF(x); Algorithm 180 [S15]: Error function — large X; Algorithm 181 [S15]: Complementary error function — large X; Algorithm 209 [S15]: Gauss; Algorithm 226 [S15]: Normal distribution function; Algorithm 272 [S15]: Procedure for the normal distribution functions; Algorithm 304 [S15]: Normal curve integral. Communications of the ACM, 10 (6):377–378, June 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Cyv64, Mac65, HJ67a].

Hickey:1996:FSP

[HJ96] Timothy J. Hickey and Qun Ju. Fast, sound, and precise narrowing of the exponential function. Technical report, Computer Science Department, Brandeis University, Waltham, MA, USA

02254, March 1996. URL http://www.cs.brandeis.edu/~tim/
Papers/eiianuia.ps.gz.

Hubrecht:2023:TCRa

[HJZ23a] Tom Hubrecht, Claude-Pierre Jeannerod, and Paul Zimmermann. Towards a correctly-rounded and fast power function in binary64 arithmetic. Report, DI-ENS — Département d’informatique — ENS Paris, Paris, France, July 12, 2023. URL

https://inria.hal.science/hal-04159652v1/.

Hubrecht:2023:TCRb

[HJZ23b] Tom Hubrecht, Claude-Pierre Jeannerod, and Paul Zimmermann. Towards a correctly-rounded and fast power function in binary64 arithmetic. In IEEE [IEE23], pages 111–118.

LCCN ???? URL https://arith2023.arithsymposium.org/
slides/S6_PaulZimmermannS6P1.pdf;   https://inria.hal.
science/hal-04159652v1/file/pow.pdf;   https://inria.
hal.science/hal-04326201.

Hubrecht:2024:TCR

[HJZ^{+}24] Tom Hubrecht, Claude-Pierre Jeannerod, Paul Zimmermann, Laurence Rideau, and Laurent Théry. Towards a correctlyrounded and fast power function in binary64 arithmetic. Report, DI-ENS — Département d’informatique — ENS Paris, Paris,

France, February 8, 2024. URL https://inria.hal.science/
hal-04159652v2/.

Hu:1993:BRS

[HKA93] Chen-Yi Hu, R. Baker Kearfott, and Abdulhamid Awad. On bounding the range of some elementary functions in FORTRAN 77. Interval Computations = Interval’nye vychisleniia, 1993(3): 29–39, 1993. ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Harrison:1999:CTF

[HKST99] John Harrison, Ted Kubaska, Shane Story, and Ping Tak Peter Tang. The computation of transcendental functions on the IA-64 architecture. Intel Technology Journal, (Q4):7, November 22,

1999. URL http://developer.intel.com/technology/itj/
q41999/articles/art_5.htm;   http://developer.intel.
com/technology/itj/q41999/pdf/transendental.pdf.

Harumi:1960:VTN

[HKW60] Kasaburô Harumi, Shigetoshi Katsura, and John W. Wrench, Jr. Values of \frac{2}{π} ∫_{0}^{∞} (\frac{sin t}{t})^{n}dt (in Technical Notes and Short Papers). Mathematics of Computation, 14(72):379, October 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Huang:2013:NNE

[HL13] Zhi-Wei Huang and Jueping Liu. NumExp: Numerical epsilon expansion of hypergeometric functions. Computer Physics Communications, 184(8):1973–1980, August 2013. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465513001136.

Horyachyy:2019:SEF

[HMO19] Oleh Horyachyy, Leonid Moroz, and Viktor Otenko. Simple effective fast inverse square root algorithm with two magic constants. International Journal of Computing, 18 (4):461–470, December 2019. ISSN 1727-6209 (print), 2312-

5381 (electronic). URL https://www.computingonline.net/
computing/article/view/1616;   https://www.researchgate.
net/publication/349173096_SIMPLE_EFFECTIVE_FAST_INVERSE_
SQUARE_ROOT_ALGORITHM_WITH_TWO_MAGIC_CONSTANTS.

Hobson:1931:TSE

[Hob31] Ernest William Hobson. The Theory of Spherical and Ellipsoidal Harmonics. Cambridge University Press, Cambridge, UK, 1931. xi + 500 pp. LCCN QA406 .H7.

Hobson:1955:TSE

[Hob55] Ernest William Hobson. The Theory of Spherical and Ellipsoidal Harmonics. Chelsea Pub. Co., New York, NY, USA, 1955. 500 pp. LCCN QA406 .H7 1955.

Hochstadt:1961:SFM

[Hoc61] Harry Hochstadt. Special functions of mathematical physics. Holt, Rinehart, and Winston, New York, NY, USA, 1961. ISBN 0-236-73011-8. viii + 81 pp. LCCN QA351 H65 1961.

Hochstrasser:1964:OP

[Hoc64] Urs W. Hochstrasser. Orthogonal polynomials. In Abramowitz and Stegun [AS64], pages 771–802. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://

www.convertit.com/Go/ConvertIt/Reference/AMS55.ASP in

bitmap image format.

Hochstadt:1971:FMP

[Hoc71] Harry Hochstadt. The Functions of Mathematical Physics, volume XXIII of Pure and applied mathematics: a series of texts and monographs. Wiley-Interscience, New York, NY, USA, 1971. ISBN 0-471-40170-6 (hardcover). xi + 322 pp. LCCN QA351 .H68.

Hochstadt:1986:FMP

[Hoc86] Harry Hochstadt. The Functions of Mathematical Physics. Dover, New York, NY, USA, 1986. ISBN 0-486-65214-9 (paperback), 0-486-16878-6 (e-book). xi + 322 pp. LCCN QA351 .H68 1986.

Holzwarth:1969:VBB

[Hol69] A. Holzwarth. Ein Verfahren zur Bestimmung bester Tschebyscheff-Approximationen der Quadratwurzelfunktion. (German) A Method for Determination of Best Chebyshev Approximations to the Square Root Function. Computing: Archiv fur informatik und numerik, 4(2):168–177, June 1969. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Holmgren:1970:RAN

[Hol70] Bo Holmgren. Remark on Algorithm 304, Normal Curve Integral. Communications of the ACM, 13(10):624, October 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Holton:1990:IJE

[Hol90] P. G. W. Holton. An introduction to the Jacobian elliptic functions with some applications. Thesis (m.sc.), University of Newcastle upon Tyne, Newcastle upon Tyne, UK, 1990. 117 pp.

Homeier:1994:NCA

[Hom94] Herbert H. H. Homeier. Nonlinear convergence acceleration for orthogonal series. In Ralf Gruber and Marco Tomassini, editors, Proceedings of the 6th Joint EPS-APS International Conference on Physics Computing, Physics Computing ’94, pages 47–50. European Physical Society, Boite Postale 69, CH-1213 Petit-Lancy, Geneva, Switzerland, Lugano, 1994. ISBN 2-

88270-011-3. URL http://www.chemie.uni-regensburg.de/
pub/preprint/preprint.html#TCNA942.

Homeier:1996:CAI

[Hom96a] H. H. H. Homeier. On convergence acceleration for the iterative solution of the inverse Dyson equation. Journal of Molecular Structure. Theochem, 368:81–91, 1996. CO-DEN THEODJ. ISSN 0166-1280 (print), 1872-7999 (elec-

tronic). URL http://www.chemie.uni-regensburg.de/pub/
preprint/preprint.html#TCQM954. Proceedings of the 2nd

Electronic Computational Chemistry Conference.

Homeier:1996:KMP

[Hom96b] H. H. H. Homeier. Zur Konvergenzverbesserung der Møller– Plesset Störungsreihe (English: On convergence acceleration of the Møller–Plesset perturbation series). Technical Report Homeier:1996:KMP, Institut für Physikalische und Theoretische Chemie, Universität Regensburg, D-93040 Regens-

burg, 1996. URL http://www.chemie.uni-regensburg.de/
pub/preprint/preprint.html#TCQM962. Poster CP 14.77,

Frühjahrstagung des Arbeitskreises Festkörperphysik bei der DPG, Regensburg 1996. Abstract: Verhandl. DPG (VI) 31, 2165-2166 (1996).

Homeier:1998:CAM

[Hom98a] H. H. H. Homeier. On convergence acceleration of multipolar and orthogonal expansions. Internet Journal of Chemistry, 1(Article 28):????, 1998. CODEN IJCHFJ. ISSN 1099-8292. URL http:

//www.ijc.com/articles/1998v1/28/. Proceedings of the 4^{th}

Electronic Computational Chemistry Conference.

Homeier:1998:AHC

[Hom98b] Herbert H. H. Homeier. An asymptotically hierarchy-consistent, iterative sequence transformation for convergence acceleration of Fourier series. Numerical Algorithms, 18(1):1–30, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://ipsapp007.

kluweronline.com/content/getfile/5058/13/2/abstract.
htm; http://ipsapp007.kluweronline.com/content/getfile/
5058/13/2/fulltext.pdf;   http://www.chemie.uni-regensburg.
de/pub/preprint/preprint.html#TCNA972.

Homeier:1999:CAL

[Hom99] H. H. H. Homeier. Convergence acceleration of logarithmically convergent series avoiding summation. Applied Mathematics Letters, 12(3):29–32, 1999. CODEN AMLEEL. ISSN 0893-9659 (print), 1873-5452 (electronic).

Honey:1971:CCD

[Hon71] D. W. Honey. Correspondence: Calculation of a double-length square root from a double length number using single precision techniques. The Computer Journal, 14(4):443, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140443.sgm.abs.html;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/443.tif.

Horner:1819:XNM

[Hor19] William George Horner. XXI. A new method of solving numerical equations of all orders, by continuous approximation. Philosophical Transactions of the Royal Society of London, 109:308– 335, 1819. CODEN PTRSAV. ISSN 0370-2316 (print), 2053-

9207 (electronic). URL https://royalsocietypublishing.
org/doi/pdf/10.1098/rstl.1819.0023.

Horsley:2017:BPF

[Hor17] David E. Horsley. Bessel phase functions: calculation and application. Numerische Mathematik, 136(3):679–702, July 2017. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hough:1981:API

[Hou81] D. Hough. Application of the proposed IEEE 754 standard for floating-point arithmetic. Computer, 14(3):70–74, 1981. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Howard:2022:AAA

[How22] Roy M. Howard. Arbitrarily accurate analytical approximations for the error function. Mathematical and Computational Applications, 27(1):14:1–14:44, February 2022. CODEN ???? ISSN

2297-8747. URL https://www.mdpi.com/2297-8747/27/1/14.

Hansen:1962:SRV

[HP62] Eldon R. Hansen and Merrell L. Patrick. Some relations and values for the generalized Riemann zeta function. Mathematics of Computation, 16(79):265–274, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Hill:1967:ACS

[HP67] I. D. Hill and M. C. Pike. Algorithm 299: Chi-squared integral. Communications of the ACM, 10(4):243–244, April 1967. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [HP85, eL76].

Hill:1985:RCS

[HP85] I. D. Hill and M. C. Pike. Remark on “Algorithm 299: Chi-Squared Integral”. ACM Transactions on Mathematical Software, 11(2):185, June 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [HP67, eL76, eL79].

Han:1993:CAS

[HP93] Weimin Han and Florian A. Potra. Convergence acceleration for some rootfinding methods. In Albrecht et al. [AAS93], pages 67–78. CODEN COSPDM. ISBN 0-387-82451-0 (New York), 3-211-82451-0 (Vienna). ISSN 0344-8029. LCCN QA297 .V27 1993. Dedicated to Ulrich Kulisch on the occasion of his 60th birthday.

Hunter:1972:NEC

[HR72] D. B. Hunter and T. Regan. A note on the evaluation of the complementary error function. Mathematics of Computation, 26 (118):539–541, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Horwitz:1988:TPF

[HR88] Alan L. Horwitz and Lee A. Rubel. Totally positive functions and totally bounded functions on [−1, 1]. Journal of Approximation Theory, 52(2):204–216, February 1988. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Hernandez:2005:ACN

[HR05] M. A. Hernández and N. Romero. Accelerated convergence in Newton’s method for approximating square roots. Journal of Computational and Applied Mathematics, 177(1):225–229, May 1, 2005. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042704004315.

Hernandez:2007:MPO

[HR07] M. A. Hernández and N. Romero. Methods with prefixed order for approximating square roots with global and general convergence. Applied Mathematics and Computation, 194(2):346– 353, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Hasson:1983:C

[HS83] M. Hasson and O. Shisha. On the condition ∑_{n=1}^{∞} n^{p−1}E_{n}^{∗}(f) < ∞. Journal of Approximation Theory, 39(4):389–398, December 1983. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Hrycak:2020:ELP

[HS20] Tomasz Hrycak and Sebastian Schmutzhard. Evaluation of Legendre polynomials by a three-term recurrence in floating-point arithmetic. IMA Journal of Numerical Analysis, 40(1):587–605, January 2020. CODEN IJNADH. ISSN 0272-4979 (print), 1464-

3642 (electronic). URL http://academic.oup.com/imajna/
article/40/1/587/5162990.

Harris:2024:UDS

[HSE^{+}24] David Harris, James Stine, Milo Ercegovac, Alberto Nannarelli, Katherine Parry, and Cedar Turek. Unified digit selection for radix-4 recurrence division and square root. IEEE Transactions on Computers, 73(1):292–300, January 2024. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Hornik:1989:MFN

[HSW89] K. Hornik, M. Stinchcombe, and H. White. Multilayer feedforward networks are universal approximators. Neural Networks, 2 (5):359–366, 1989. CODEN NNETEB. ISSN 0893-6080 (print), 1879-2782 (electronic).

Hwang:1973:RRS

[HT73] W. G. Hwang and John Todd. A recurrence relation for the square root. Journal of Approximation Theory, 9:299–306, 1973. CODEN JAXTAZ. ISSN 0021-9045,1096-0430.

Hale:2013:FAC

[HT13] Nicholas Hale and Alex Townsend. Fast and accurate computation of Gauss–Legendre and Gauss–Jacobi quadrature nodes and weights. SIAM Journal on Scientific Computing, 35(2): A652–A674, ???? 2013. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Hahn:1994:UDF

[HTHR94] H. Hahn, D. Timmermann, B. J. Hosticka, and B. Rix. A unified and division-free CORDIC argument reduction method with unlimited convergence domain including inverse hyperbolic functions. IEEE Transactions on Computers, 43(11):1339–1344, November 1994. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=324568.

Hull:1978:DFP

[Hul78] T. E. Hull. Desirable floating-point arithmetic and elementary functions for numerical computation. In IEEE [IEE78], pages 63–69. ISSN 1063-6889. LCCN QA76.6 .S919a. IEEE catalog no. 78 CH1412-6C.

Hummer:1964:EDF

[Hum64] David G. Hummer. Expansions of Dawson’s function in a series of Chebyshev polynomials (in Technical Notes and Short Papers). Mathematics of Computation, 18(86):317–319, April 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Humblet:1985:BFE

[Hum85] J. Humblet. Bessel function expansions of Coulomb wave functions. Journal of Mathematical Physics, 26(4):656–659, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v26/i4/p656_s1.

Holzapfel:2007:AGA

[HUY07] Rolf-Peter Holzapfel, A. Muhammed Uludag, and Masaaki Yoshida, editors. Arithmetic and geometry around hypergeometric functions: lecture notes of a CIMPA Summer School held at Galatasaray University, Istanbul, Turkey, June 13–25, 2005, volume 235 of Progress in mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2007. ISBN 3-7643-8283-X. LCCN

QA245 .S86 2005.  URL http://www.loc.gov/catdir/
enhancements/fy0913/2006939568-d.html;   http://www.
loc.gov/catdir/enhancements/fy0913/2006939568-t.html.

Hofsommer:1963:NCE

[HvdR63] D. J. Hofsommer and R. van de Riet. On the numerical calculation of elliptic integrals of the first and second kind and the elliptic functions of Jacobi. Numerische Mathematik, 5:291–302, December 1963. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Hwang:1985:PSC

[Hwa85] Kai Hwang, editor. Proceedings: 7th Symposium on Computer Arithmetic, June 4–6, 1985, University of Illinois, Urbana, Illinois. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ISBN 0-8186-0632-0 (paperback), 0-8186-8632-4 (hard), 0-8186-4632-2 (microfiche). LCCN QA76.9.C62 S95 1985. IEEE catalog number 85CH2146-9. IEEE Computer Society order number 632.

Hao:2022:DVP

[HXX22] Jiangwei Hao, Jinchen Xu, and YuanYuan Xia. Design of variable precision transcendental function automatic generator. The Journal of Supercomputing, 78(2):2196–2218, February 2022. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484

(electronic). URL https://link.springer.com/article/10.
1007/s11227-021-03937-8.

Hong:2014:MSI

[HY14] Hoon Hong and Chee Yap, editors. Mathematical Software — ICMS 2014: 4th International Conference, Seoul, South Korea, August 5–9, 2014, Proceedings, volume 8592. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (e-book). LCCN QA76.9.M35.

Ismail:2020:ESF

[IA20] Mourad H. Ismail and Walter Van Assche, editors. Encyclopedia of Special Functions: the Askey–Bateman Project. Volume 1, Univariate Orthogonal Polynomials. Cambridge University Press, Cambridge, UK, 2020. ISBN 0-511-97915-0 (e-book), 0-521-19742-2 (hardcover), 1-108-76433-9 (e-book). xiv + 388 pp. LCCN QA351 .E63 2020.

Iacono:2021:BEF

[Iac21] Roberto Iacono. Bounding the error function. Computing in Science and Engineering, 23(4):65–68, 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Ibbetson:1963:AG

[Ibb63] D. Ibbetson. Algorithm 209: Gauss. Communications of the ACM, 6(10):616, October 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

IBM:2005:MAS

[IBM05] IBM Corporation. Mathematical Acceleration Subsystem for Linux. World Wide Web document, 2005. URL http://www-

306.ibm.com/software/awdtools/mass/linux/mass-linux.
html.

IEEE:1978:PSC

[IEE78] IEEE, editor. Proceedings of the Symposium on Computer Arithmetic (4th: 1978: Santa Monica, CA). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1978. ISSN 1063-6889. LCCN QA76.6 .S919a. IEEE catalog no. 78 CH1412-6C.

IEEE:1981:PIS

[IEE81a] IEEE, editor. Proceedings: 5th Symposium on Computer Arithmetic, May 18-19, 1981, University of Michigan, Ann Arbor, Michigan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. LCCN QA76.9.C62 S95 1981. IEEE Catalog No. 81CH1630-3. Computer Society Order No. 347.

IEEE:1981:PSC

[IEE81b] Proceedings: 5th Symposium on Computer Arithmetic: May 18– 19, 1981, University of Michigan, Ann Arbor, Michigan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. LCCN QA 76.6 S985t 1981. IEEE catalog number 81CH1630-C.

IEEE:1985:ERC

[IEE85] 1985 IEEE Region 5 Conference, March 13–15, 1985, Holiday Inn Civic Center, Lubbock, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. LCCN TK 7801 N56 1985.

IEE:1989:EEC

[IEE89a] IEE, editor. ECCTD 89: European Conference on Circuit Theory and Design, 5–8 September 1989: venue, University of Sussex, Brighton, United Kingdom. IEE, London, UK, 1989. ISBN 0-85296-383-1. LCCN ???? Conference publication no. 308.

IEEE:1989:ASF

[IEE89b] IEEE, editor. 30th annual Symposium on Foundations of Computer Science, October 30–November 1, 1989, Research Triangle Park, North Carolina. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. CODEN ASFPDV. ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

IEEE:1989:PII

[IEE89c] Proceedings: 1989 IEEE International Conference on Computer Design: VLSI in Computer and Processors, ICCD ’89, Hyatt Regency Cambridge, Cambridge, Massachusetts, October 2– 4, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-1971-6 (paper), 0-8186-5971-8 (microfiche), 0-8186-8971-4 (case). LCCN TK 7888.4 I23 1989. IEEE catalog number 89CH2794-6.

IEEE:1991:PSA

[IEE91] IEEE, editor. Proceedings, Supercomputing ’91: Albuquerque, New Mexico, November 18–22, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

IEEE:1999:PIF

[IEE99] IEEE, editor. Proceedings of the IEEE Forum on Research and Technology Advances in Digital Libraries, May 19–21, 1999, Baltimore, Maryland. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7695-0219-9. LCCN ZA4080 .F67 1999.

IEEE:2005:PIS

[IEE05] IEEE, editor. Proceedings of the 17th IEEE Symposium on Computer Arithmetic, ARITH-17, June 27–29, 2005, Cape Cod, Massachusetts, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN ???? LCCN ????

IEEE:2008:ICA

[IEE08] IEEE, editor. 2008 International Conference on Application-Specific Systems, Architectures and Processors: Leuven, Belgium, 2–4 July 2008. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2008. ISBN 1-4244-1897-6 (paperback), 1-4244-1898-4. LCCN

????  URL http://ieeexplore.ieee.org/servlet/opac?
punumber=4569858;   http://www.gbv.de/dms/tib-ub-hannover/

631855815.pdf. IEEE catalog number CFP08063-PRT.

IEEE:2013:PIS

[IEE13] IEEE, editor. Proceedings of the 21st IEEE Symposium on Computer Arithmetic, Austin, Texas, USA, 8–10 April 2013. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 0-7695-4957-8. ISSN 1063-6889. LCCN QA76.9.C62 S95 2013.

IEEE:2021:ISC

[IEE21] IEEE, editor. 2021 IEEE 28th Symposium on Computer Arithmetic: ARITH 2021: virtual conference, 14–16 June 2021: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2021. ISBN 1-66542-293-9 (print), 1-66544-648-X (e-book). LCCN ????

IEEE:2022:ISC

[IEE22] IEEE, editor. 2022 IEEE 29th Symposium on Computer Arithmetic: ARITH 2022: virtual conference, 12–14 September 2022: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2022. ISBN 1-66547-827-6, 1-66547-828-4. LCCN ????

IEEE:2023:PIS

[IEE23] IEEE, editor. Proceedings: 2023 IEEE 30th Symposium on Computer Arithmetic: ARITH 2023, 4–6 September 2023 Portland, United States. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2023. LCCN ????

Ismail:2005:TAS

[IK05] Mourad E. H. Ismail and Erik Koelink, editors. Theory and Applications of Special Functions: a Volume Dedicated to Mizan Rahman, volume 13 of Developments in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-24231-7 (hardcover), 0-387-24233-3 (e-book). x + 491 pp.

LCCN QA351 .T44 2005. URL http://www.loc.gov/catdir/
enhancements/fy0663/2005275626-d.html;   http://www.
loc.gov/catdir/toc/fy0605/2005275626.html.

Ikebe:1975:ZRC

[Ike75] Yasuhiko Ikebe. The zeros of regular Coulomb wave functions and of their derivatives. Mathematics of Computation, 29 (131):878–887, July 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ikebe:1976:CZB

[Ike76] Y. Ikebe. Computing zeros of Bessel and regular Coulomb wave functions and of their derivatives by matrix theoretic approach — practical accuracy criteria. SIAM Review, 18(4):810, ???? 1976. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Ikebe:1991:CZO

[IKF91] Yasuhiko Ikebe, Yasushi Kikuchi, and Issei Fujishiro. Computing zeros and orders of Bessel functions. Journal of Computational and Applied Mathematics, 38(1–3):169–184, December 23, 1991. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037704279190169K.

Ifantis:1991:PZS

[IKK91] E. K. Ifantis, C. G. Kokologiannaki, and C. B. Kouris. On the positive zeros of the second derivative of Bessel functions. Journal of Computational and Applied Mathematics, 34(1):21–31, February 10, 1991. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0377042791901449.

Iske:2007:AAP

[IL07] Armin Iske and Jeremy Levesley, editors. Algorithms for Approximation: Proceedings of the 5th International Conference, Chester, July 2005. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-46551-0, 3-540-33283-9. LCCN QA221 .A44 2007.

Iordache:1999:ARS

[IM99a] Cristina Iordache and David W. Matula. Analysis of reciprocal and square root reciprocal instructions in the AMD K6-2 implementation of 3DNow! Electronic Notes in Theoretical Computer Science, 24:34–62, 1999. CODEN ???? ISSN 1571-0661.

Iordache:1999:IPR

[IM99b] Cristina Iordache and David W. Matula. On infinitely precise rounding for division, square root, reciprocal and square root reciprocal. In Koren and Kornerup [KK99], pages 233–240. CO-DEN ???? ISBN 0-7803-5609-8, 0-7695-0116-8, 0-7695-0118-4.

ISSN ???? LCCN QA76.6 .S887 1999. URL http://euler.ecs.

umass.edu/paper/final/paper-164.pdf;   http://euler.
ecs.umass.edu/paper/final/paper-164.ps;   http://www.
acsel-lab.com/arithmetic/arith14/papers/ARITH14_Iordache.

pdf. IEEE Computer Society Order Number PR00116. IEEE Order Plan Catalog Number 99CB36336.

Intel:2000:DSR

[Int00] Intel. Divide, square root, and remainder algorithms for the Itanium architecture. Intel Software Development Products, July

2000. URL https://studylib.net/doc/7921762/divide--
square-root-and-remainder-algorithms-for-the-ia-64.

Intel:2003:DSR

[Int03a] Intel. Divide, square root, and remainder algorithms for the Itanium architecture. Intel Software Development Products, De-

cember 18, 2003. URL http://www.intel.com/cd/software/
products/asmo-na/eng/219863.htm.

Intel:2003:NID

[Int03b] Intel. Non-IEEE division, square root, reciprocal, and reciprocal square root algorithms for the Intel Itanium architecture. Intel Software Development Products, December 18,

2003. URL http://www.intel.com/cd/software/products/
asmo-na/eng/219864.htm.

Iserles:1981:ARA

[IP81] A. Iserles and M. J. D. Powell. On the A-acceptability of rational approximations that interpolate the exponential function. IMA Journal of Numerical Analysis, 1(3):241–251, July 1981. CO-DEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Iserles:1987:SAN

[IP87] A. Iserles and M. J. D. Powell, editors. The State of the Art in Numerical Analysis: Proceedings of the Joint IMA/SIAM Conference on the State of the Art in Numerical Analysis held at the University of Birmingham, 14–18 April 1986. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1987. ISBN 0-19-853614-3. LCCN QA297 .S781 1987. UK£55.00, US$77.50.

Istoan:2017:FFP

[IP17] M. Istoan and B. Pasca. Flexible fixed-point function generation for FPGAs. In Burgess et al. [BBdD17], pages 123–130. ISBN 1-5386-1966-0 (print), 1-5386-1965-2, 1-5386-1964-4. ISSN 1063-6889. LCCN QA76.9.C62 S95 2017. URL http://ieeexplore.

ieee.org/servlet/opac?punumber=8019911.

Isukapalli:2008:ATA

[IR08] Yogananda Isukapalli and Bhaskar D. Rao. An analytically tractable approximation for the Gaussian Q-function. IEEE Communications Letters, 12(9):669–671, September 2008. CO-DEN ICLEF6. ISSN 1089-7798 (print), 1558-2558 (electronic).

Ifantis:1988:BFP

[IS88] E. K. Ifantis and P. D. Siafarikas. Bounds for the first positive zero of a mixed Bessel function. Journal of Computational and Applied Mathematics, 21(2):245–249, February 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042788902737.

Ifantis:1990:DIP

[IS90] E. K. Ifantis and P. D. Siafarikas. Differential inequalities for the positive zeros of Bessel functions. Journal of Computational and Applied Mathematics, 30(2):139–143, May 28, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037704279090022R.

Ifantis:1992:DIP

[IS92] E. K. Ifantis and P. D. Siafarikas. A differential inequality for the positive zeros of Bessel functions. Journal of Computational and Applied Mathematics, 44(1):115–120, December 9, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042792900553.

Iserles:1991:CDC

[Ise91] A. Iserles. Complex dynamics of convergence acceleration. IMA Journal of Numerical Analysis, 11(2):205–240, 1991. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Iserles:1994:CAD

[Ise94] A. Iserles. Convergence acceleration as a dynamical system. Applied Numerical Mathematics: Transactions of IMACS, 15(2):101–121, September 13, 1994. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

URL http://www.elsevier.com/cgi-bin/cas/tree/store/
apnum/cas_sub/browse/browse.cgi?year=1994&volume=15&

issue=2&aid=496. Innovative methods in numerical analysis (Bressanone, 1992).

Ifantis:1987:UBF

[ISK87] E. K. Ifantis, P. D. Siafarikas, and C. B. Kouris. Upper bounds for the first zeros of Bessel functions. Journal of Computational and Applied Mathematics, 17(3):355–358, March 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042787901117.

Ismail:1977:IRC

[Ism77] Mourad E. H. Ismail. Integral representations and complete monotonicity of various quotients of Bessel functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 29(??):1198–1207, ???? 1977. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

ISO:2010:IIIa

[ISO10] ISO. ISO/IEC 29124:2010: Information technology — Programming languages, their environments and system software interfaces — Extensions to the C++ Library to support mathematical special functions. Technical report. International Organization for Standardization, Geneva, Switzerland, 2010.

LCCN ???? URL http://www.iso.org/iso/iso_catalogue/
catalogue_tc/catalogue_detail.htm?csnumber=50511.

Ito:1995:EIA

[ITY95] M. Ito, N. Takagi, and S. Yajima. Efficient initial approximation and fast converging methods for division and square root. In Knowles and McAllister [KM95], pages 2–9. CODEN ???? ISBN 0-8186-7089-4 (paperback), 0-8186-7089-4 (case), 0-8186-7149-1 (microfiche), 0-8186-7089-4 (softbound), 0-7803-2949-X (casebound). ISSN ???? LCCN QA 76.9 C62 S95 1995.

Ito:1996:SRI

[ITY96] Masayuki Ito, Naofumi Takagi, and Shuzo Yajima. Square rooting by iterative multiply-additions. Information Processing Letters, 60(5):267–269, December 8, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ito:1997:EIA

[ITY97] M. Ito, N. Takagi, and S. Yajima. Efficient initial approximation for multiplicative division and square root by a multiplication with operand modification. IEEE Transactions on Computers, 46(4):495–498, April 1997. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=588066.

Jablonski:1994:NES

[Jab94] Aleksander Jablonski. Numerical evaluation of spherical Bessel functions of the first kind. Journal of Computational Physics, 111(2):256–259, April 1994. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999184710606.

Jablonski:2012:EAC

[Jab12] A. Jablonski. An effective algorithm for calculating the Chandrasekhar function. Computer Physics Communications, 183(8):1773–1782, August 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465512000847.

Jablonski:2013:IAC

[Jab13] A. Jablonski. Improved algorithm for calculating the Chandrasekhar function. Computer Physics Communications, 184 (2):440–442, February 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046551200286X.

Jablonski:2020:IAC

[Jab20] A. Jablonski. Improved algorithm for calculating high accuracy values of the Chandrasekhar function. Computer Physics Communications, 251(??):Article 107237, June 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465520300709.

James:1981:LTS

[Jam81] D. G. James. Linear transformations of the second elementary function. Linear and Multilinear Algebra, 10(4):347–349, 1981.

CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Jamieson:1989:RCI

[Jam89] M. J. Jamieson. Rapidly converging iterative formulae for finding square roots and their computational efficiencies. The Computer Journal, 32(1):93–94, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_32/Issue_
01/tiff/93.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/94.tif. This work generalizes

the Pythagorean sums in [Dub83, MM83].

Jamieson:1999:NRF

[Jam99] M. J. Jamieson. Notes: On rational function approximations to square roots. American Mathematical Monthly, 106(1):50–52, January 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jameson:2016:IGF

[Jam16] G. J. O. Jameson. The incomplete gamma functions. The Mathematical Gazette, 100(548):298–306, July 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
9373A31AD28D793AB5431E35EA5C2AF6.

Jansen:1977:RLF

[Jan77] J. K. M. Jansen. Remark on “Algorithm 259: Legendre Functions for Arguments Larger than One”. ACM Transactions on Mathematical Software, 3(2):204–250, June 1977. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Gau65a].

Jang:2011:CTS

[Jan11a] Won Mee Jang. Corrections to “A Simple Upper Bound of the Gaussian Q-Function with Closed-form Error Bound”. IEEE Communications Letters, 15(12):1274, December 2011. CODEN ICLEF6. ISSN 1089-7798 (print), 1558-2558 (electronic). URL

http://ieeexplore.ieee.org/document/6065242/.

Jang:2011:SUB

[Jan11b] Won Mee Jang. A simple upper bound of the Gaussian Qfunction with closed-form error bound. IEEE Communications Letters, 15(2):157–159, February 2011. CODEN ICLEF6. ISSN 1089-7798 (print), 1558-2558 (electronic). URL http://

ieeexplore.ieee.org/document/5692888/.

Jargstorff:2012:AEF

[Jar12] Frank Jargstorff. Approximating the erfinv function. In mei Hwu [mH12], chapter 10, page ?? ISBN 0-12-385963-8 (hardback). LCCN T385 .G6875 2012.

Jeffrey:2008:HMF

[JD08] Alan Jeffrey and Hui-Hui Dai. Handbook of Mathematical Formulas and Integrals. Elsevier Academic Press, Amsterdam, The Netherlands, fourth edition, 2008. ISBN 0-12-374288-9 (paperback), 0-08-055684-1 (e-book). xlv + 541 pp. LCCN QA47 .J38

2008. URL https://shop.elsevier.com/books/handbook-
of-mathematical-formulas-and-integrals/jeffrey/978-0-
12-374288-9.

Jefferson:1962:RAI

[Jef62] David K. Jefferson. Remark on Algorithm 73: Incomplete elliptic integrals. Communications of the ACM, 5(10):514, October 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jeffrey:1995:HMF

[Jef95] Alan Jeffrey. Handbook of Mathematical Formulas and Integrals. Academic Press, New York, NY, USA, 1995. ISBN 0-08-052301-3, 0-12-382580-6 (e-book), 1-4832-9514-1 (e-book). xxiv + 410

pp. LCCN QA47 .J38 1995. URL http://www.sciencedirect.
com/science/book/9780123825803.

Jeffrey:2000:HMF

[Jef00] Alan Jeffrey. Handbook of Mathematical Formulas and Integrals. Academic Press, New York, NY, USA, second edition, 2000. ISBN 0-12-382251-3. xxvi + 433 pp. LCCN QA47 .J38 2000.

Jeffrey:2004:HMF

[Jef04] Alan Jeffrey. Handbook of Mathematical Formulas and Integrals. Elsevier Academic Press, Amsterdam, The Netherlands, third edition, 2004. ISBN 0-12-382256-4. xxvi + 453 pp. LCCN QA47 .J38 2004.

Jeffrey:2017:BSI

[Jef17] David J. Jeffrey. Branch structure and implementation of Lambert W. Mathematics in Computer Science, 11(3–4):341–350, December 2017. CODEN ???? ISSN 1661-8270 (print), 1661-8289 (electronic).

Johansson:2016:FAE

[JF16] H. T. Johansson and C. Forssén. Fast and accurate evaluation of Wigner 3j, 6j, and 9j symbols using prime factorization and multiword integer arithmetic. SIAM Journal on Scientific Computing, 38(1):A376–A384, ???? 2016. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Jiang:2013:AFE

[JGB13] Hao Jiang, Stef Graillat, and Roberto Barrio. Accurate and fast evaluation of elementary symmetric functions. In IEEE [IEE13], pages 183–190. ISBN 0-7695-4957-8. ISSN 1063-6889. LCCN QA76.9.C62 S95 2013.

Jeffrey:1996:UBL

[JHC96] D. J. Jeffrey, D. E. G. Hare, and Robert M. Corless. Unwinding the branches of the Lambert W function. The Mathematical Scientist, 21(1):1–7, June 1996. ISSN 0312-3685 (print), 1475-

6080 (electronic). URL http://www.appliedprobability.
org/data/files/TMS%20articles/21_1_1.pdf.

James:1965:GSR

[JJ65] Wendy James and P. Jarratt. The generation of square roots on a computer with rapid multiplication compared with division (in Technical Notes and Short Papers). Mathematics of Computation, 19(91):497–500, July 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Jacobsen:1986:FRC

[JJW86] Lisa Jacobsen, William B. Jones, and Haakon Waadeland. Further results on the computation of incomplete gamma functions. Lecture Notes in Mathematics, 1199:67–89, 1986. CODEN LN-MAA2. ISBN 3-540-16768-4 (print), 3-540-38817-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/BFb0075936/.

Jiang:1992:CCM

[JKD92] Thomas J. Jiang, Joseph B. Kadane, and James M. Dickey. Computation of Carlson’s multiple hypergeometric function R for Bayesian applications. Journal of Computational and Graphical Statistics, 1(3):231–251, September 1992. CODEN ???? ISSN 1061-8600 (print), 1537-2715 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/10618600.1992.
10474583.

Johnson:1992:UDD

[JKK92] Norman Lloyd Johnson, Samuel Kotz, and Adrienne W. Kemp. Univariate Discrete Distributions. Wiley series in probability and mathematical statistics. Applied probability and statistics. Wiley, New York, NY, USA, second edition, 1992. ISBN 0-471-54897-9 (hardcover). xx + 565 pp. LCCN QA273.6

.J64 1992. URL http://www.loc.gov/catdir/description/
wiley031/92011685.html;   http://www.loc.gov/catdir/
enhancements/fy0607/92011685-b.html;   http://www.loc.
gov/catdir/toc/onix01/92011685.html.

Jeannerod:2011:CFP

[JKMR11] Claude-Pierre Jeannerod, Hervé Knochel, Christophe Monat, and Guillaume Revy. Computing floating-point square roots via bivariate polynomial evaluation. IEEE Transactions on Computers, 60(2):214–227, February 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Jain:1994:SRR

[JL94] V. K. Jain and Lei Lin. Square-root, reciprocal, sine/cosine, arctangent cell for signal and image processing. In IEEE International Conference on Acoustics, Speech, and Signal Processing: ICASSP-94, 19–22 April 1994, volume 2, pages II/521–II/524. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. CODEN ???? ISSN ????

Jentschura:2012:NCB

[JL12] U. D. Jentschura and E. Lötstedt. Numerical calculation of Bessel, Hankel and Airy functions. Computer Physics Communications, 183(3):506–519, March 2012. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465511003729.

Jeannerod:2017:REC

[JM17] Claude-Pierre Jeannerod and Jean-Michel Muller. On the relative error of computing complex square roots in floating-point arithmetic. In Michael B. Matthews, editor, 2017 51st Asilomar Conference on Signals, Systems, and Computers. October 29– November 1, 2017. Pacific Grove, California, pages 737–740. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2017. ISBN 1-5386-1823-0.

Johansson:2018:FRA

[JM18] Fredrik Johansson and Marc Mezzarobba. Fast and rigorous arbitrary-precision computation of Gauss–Legendre quadrature nodes and weights. SIAM Journal on Scientific Computing, 40 (6):C726–C747, ???? 2018. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Johnson:1979:RAF

[JMMW79] Donald B. Johnson, Webb Miller, Brian Minnihan, and Celia Wrathall. Reducibility among floating-point graphs. Journal of the Association for Computing Machinery, 26(4):739–760, October 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jukic:1998:DTN

[JMS98] D. Jukić, T. Marosšević, and R. Scitovski. Discrete total l_{p}-norm approximation problem for the exponential function. Applied Mathematics and Computation, 94(2–3):137–143, August 15, 1998. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300397100686.

Johnson:1983:MGC

[Joh83] Gary M. Johnson. Multiple-grid convergence acceleration of viscous and inviscid flow computations. Applied Mathematics and Computation, 13(3–4):375–398, November 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Johnson:1987:AES

[Joh87a] Kenneth C. Johnson. Algorithm 650: Efficient square root implementation on the 68000. ACM Transactions on Mathematical Software, 13(2):138–151, June 1987. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [Joh87b].

Johnson:1987:CES

[Joh87b] Kenneth C. Johnson. Corrigendum: “Algorithm 650: efficient square root implementation on the 68000” [ACM Trans. Math. Software 13 (1987), no. 2, 138–151]. ACM Transactions on Mathematical Software, 13(3):320, September 1987. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Joh87a].

Johnsson:1988:DPP

[Joh88] S. L. Johnsson. Data parallel programming and basic linear algebra subroutines. In J. R. Rice, editor, Mathematical aspects of scientific software, volume 14 of The IMS volumes in mathematics and its applications, pages 183–196. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96706-0. LCCN QA76.76.D47 M366 1988.

Johnson:1989:IMA

[Joh89] K. R. Johnson. An iterative method for approximating square roots. Mathematics Magazine, 62(4):253–259, October 1989. CODEN MAMGA8. ISSN 0025-570X.

Johansson:2011:CRE

[Joh11] Bo Göran Johansson. Cube root extraction in medieval mathematics. Historia Mathematica, 38(3):338–367, August 2011. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086010000546.

Johnson:2012:FPF

[Joh12] Steven G. Johnson. Faddeeva package, a free/open-source C++ software to compute the various error functions of arbitrary complex arguments. Web site, 2012. URL http://ab-initio.mit.

edu/wiki/index.php/Faddeeva_Package.

Johansson:2014:EIE

[Joh14a] Fredrik Johansson. Efficient implementation of elementary functions in the medium-precision range. arxiv.org, ??(??):??, Octo-

ber 27, 2014. URL https://arxiv.org/abs/1410.7176.

Johansson:2014:FRC

[Joh14b] Fredrik Johansson. Fast and Rigorous Computation of Special Functions to High Precision. Ph.D. thesis, Johannes Kepler University, Linz, Austria, March 24, 2014. ix + 109 pp. URL http:

//permalink.obvsg.at/AC10776210.

Johansson:2015:EIE

[Joh15a] Fredrik Johansson. Efficient implementation of elementary functions in the medium-precision range. In Muller et al. [MTV15], pages 83–89. ISBN 1-4799-8665-8, 1-4799-8663-1. ISSN 1063-6889. LCCN QA76.9.C62 S95 2015. URL http://ieeexplore.

ieee.org/servlet/opac?punumber=7193754.

Johansson:2015:RHP

[Joh15b] Fredrik Johansson. Rigorous high-precision computation of the Hurwitz zeta function and its derivatives. Numerical Algorithms, 69(2):253–270, June 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9893-1.

Johansson:2016:CHF

[Joh16] Fredrik Johansson. Computing hypergeometric functions rigorously. arxiv.org, ??(??):2–29, June 22, 2016. URL http://

arxiv.org/abs/1606.06977.

Johansson:2019:CHF

[Joh19] Fredrik Johansson. Computing hypergeometric functions rigorously. ACM Transactions on Mathematical Software, 45(3): 30:1–30:26, August 2019. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://dl.acm.org/

citation.cfm?id=3328732.

Johansson:2020:CLW

[Joh20a] Fredrik Johansson. Computing the Lambert W function in arbitrary-precision complex interval arithmetic. Numerical Algorithms, 83(1):221–242, January 2020. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Johansson:2020:FSL

[Joh20b] Fredrik Johansson. FunGrim: A symbolic library for special functions. In Bigatti et al. [BCD^{+}20], pages 315–323.

Johnson:2020:EAHa

[Joh20c] Jeff Johnson. Efficient, arbitrarily high precision hardware logarithmic arithmetic for linear algebra. arxiv.org, ??(??):1–8, May

14, 2020. URL https://arxiv.org/pdf/2004.09313.pdf.

Johnson:2020:EAHb

[Joh20d] Jeff Johnson. Efficient, arbitrarily high precision hardware logarithmic arithmetic for linear algebra. In Cornea et al. [CLT20], pages 25–32. ISBN 1-72817-120-2, 1-72817-121-0. ISSN 2576-

2265. LCCN ???? URL https://ieeexplore.ieee.org/
servlet/opac?punumber=9146973.

Johansson:2021:APC

[Joh21] Fredrik Johansson. Arbitrary-precision computation of the gamma function. arXiv.org, pages 1–51, September 17, 2021.

URL https://arxiv.org/abs/2109.08392.

Jones:1970:GHF

[Jon70] Alan L. Jones. The generalized hypergeometric function. ACM SIGPLAN Notices, 5(3):26–27, March 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:2006:PCF

[Jon06] D. S. Jones. Parabolic cylinder functions of large order. Journal of Computational and Applied Mathematics, 190(1–2):453– 469, June 1, 2006. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042705002463.

Jeffreys:1956:MMP

[JS56] Harold Jeffreys and Bertha Swirles Jeffreys. Methods of Mathematical Physics. Cambridge University Press, Cambridge, UK, third edition, 1956. 714 pp. LCCN QA401 .J4 1956.

URL https://en.wikipedia.org/wiki/Bertha_Swirles;
https://en.wikipedia.org/wiki/Harold_Jeffreys.

Jeffries:1989:GFA

[JS89] John S. Jeffries and Donald R. Smith. A Green function approach for a singularly perturbed vector boundary-value problem. Advances in Applied Mathematics, 10(1):1–50, March 1989. CODEN ???? ISSN 0196-8858 (print), 1090-2074 (electronic).

Jeffreys:1999:MMP

[JS99] Harold Jeffreys and Bertha Swirles Jeffreys. Methods of Mathematical Physics. Cambridge University Press, Cambridge, UK, third edition, 1999. ISBN 0-521-66402-0 (paperback). viii + 718 pp. LCCN QA401 .J4 1999. URL https://en.wikipedia.

org/wiki/Bertha_Swirles;   https://en.wikipedia.org/
wiki/Harold_Jeffreys. Reprint of [JS56].

Jaime:2011:HSA

[JSH^{+}11] F. J. Jaime, M. A. Sánchez, J. Hormigo, J. Villalba, and E. L. Zapata. High-speed algorithms and architectures for range reduction computation. IEEE Transactions on Very Large Scale Integration (VLSI) Systems, 19(3):512–516, ???? 2011. CODEN IEVSE9. ISSN 1063-8210 (print), 1557-9999 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5308221.

Jones:1985:CIG

[JT85] William B. Jones and W. J. Thron. On the computation of incomplete gamma functions in the complex domain. Journal of Computational and Applied Mathematics, 12–13(??):401– 417, May 1985. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0377042785900342.

Jurimae:1966:KFT

[Jür66] E. Jürimäe. Kompleksmuutuja funktsioonide teooria. I: Elementaarsed funktsioonid. (Estonian) [Theory of functions of a complex variable. I: Elementary functions]. Tartu Riiklik Ülikool, Tartu, Estonia, 1966. 131 pp.

Jacobsen:1988:CAL

[JW88] Lisa Jacobsen and Haakon Waadeland. Convergence acceleration of limit periodic continued fractions under asymptotic side conditions. Numerische Mathematik, 53(3):285–298, July 1988. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kuki:1971:FEP

[KA71] H. Kuki and J. Ascoly. FORTRAN extended-precision library. IBM Systems Journal, 10(1):39–61, 1971. CODEN IBMSA7. ISSN 0018-8670.

Kahan:1980:SPI

[Kah80] W. Kahan. Software \sqrt{x} for the proposed IEEE Floating-Point Standard. Manuscript, August 25, 1980.

Kahan:1982:BCC

[Kah82] W. Kahan. Branch cuts for complex elementary functions. Technical Report PAM-105, Center for Pure and Applied Mathematics, University of California, Berkeley, Berkeley, CA, USA, 1982.

Kahan:1987:BCC

[Kah87] W. Kahan. Branch cuts for complex elementary functions or much ado about nothing’s sign bit. In Iserles and Powell [IP87], pages 165–211. ISBN 0-19-853614-3. LCCN QA297 .S781 1987. UK£55.00, US$77.50.

Kahan:1996:TCR

[Kah96] W. Kahan. A test for correctly rounded SQRT. Lecture notes.,

1996. URL http://www.cs.berkeley.edu/~wkahan/SQRTest.
ps.

Kahan:1999:SRD

[Kah99] W. Kahan. Square root without division. World-Wide Web document, February 23, 1999. URL http://www.cs.berkeley.

edu/~wkahan/ieee754status/reciprt.pdf.

Kahan:2004:LTC

[Kah04] W. Kahan. A logarithm too clever by half. World-Wide Web document, August 9, 2004. URL http://www.cs.berkeley.

edu/~wkahan/LOG10HAF.TXT.

Kahan:2018:TD

[Kah18] William Kahan. The tanpi dilemma. Web document., April 16,

2018. URL http://754r.ucbtest.org/background/tanpi.
txt;   https://www.math.utah.edu/pub/tex/bib/elefunt.
bib.

Kaishev:1989:SSC

[Kai89] A. I. Kaishev. A sharpened scheme for constructing an a posteriori interval extension of an elementary function. (Russian). Voprosy Kibernet. (Moscow), 149:14–18, 1989. ISBN 0134-6388.

Kalbfell:1934:QDN

[Kal34] D. C. Kalbfell. Questions, discussions and notes: On a method for calculating square roots. American Mathematical Monthly, 41(8):504–506, October 1934. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.

org/stable/2300417.

Kalmykov:2004:SEH

[Kal04] M. Y. Kalmykov. Series and ϵ-expansion of the hypergeometric functions. Nuclear Physics B, Proceedings Supplements, 135(??): 280–284, ???? 2004. CODEN NPBSE7. ISSN 0920-5632 (print), 1873-3832 (electronic).

Kantabutra:1996:HCE

[Kan96] V. Kantabutra. On hardware for computing exponential and trigonometric functions. IEEE Transactions on Computers, 45(3):328–339, March 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=485571.

Karp:1984:ELS

[Kar84] A. H. Karp. Exponential and logarithm by sequential squaring. IEEE Transactions on Computers, C-33(5):462–464, May 1984. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1676464.

Karatsuba:2001:ARE

[Kar01] Ekatherina A. Karatsuba. On the asymptotic representation of the Euler gamma function by Ramanujan. Journal of Computational and Applied Mathematics, 135(2):225–240, October 15, 2001. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042700005860.

Kasperkovitz:1980:AAM

[Kas80] P. Kasperkovitz. Asymptotic approximations for modified Bessel functions. Journal of Mathematical Physics, 21(1):6–13, January 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v21/i1/p6_s1.

Katholi:1978:CVP

[Kat78] Charles R. Katholi. On the computation of values of the psi function from rapidly converging power series expansions. Journal of Statistical Computation and Simulation, 8(1):25–42, 1978. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Kazangapov:1965:REF

[Kaz65] A. N. Kazangapov. Representation of elementary function in the system of residual classes. (Russian). Izv. Akad. Nauk Kazah. SSR Ser. Fiz.-Mat. Nauk, 3:79–84, 1965.

Knuth:1967:CTE

[KB67] Donald E. Knuth and Thomas J. Buckholtz. Computation of tangent, Euler, and Bernoulli numbers. Mathematics of Computation, 21(100):663–688, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kushner:1986:ECC

[KB86] Ed Kushner and Rick Broussard. Efficient computation of the cylindrical Bessel functions of complex argument. Computer Physics Communications, 42(3):345–349, November 1986. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0010465586900044.

Kak:1989:BAS

[KB89] S. C. Kak and A. O. Barbir. The Brahmagupta algorithm for square rooting. In Proceedings of the Twenty-First Southeastern Symposium on System Theory, 26–28 March 1989, pages 456– 459. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. CODEN ???? ISSN ????

Kiernan:1990:FAE

[KB90] J. M. Kiernan and T. B. Blachowiak. Fast, accurate elementary functions for the Cray Y-MP computer. In ???? [???90], pages 243–252. ISBN ???? LCCN ????

Kerridge:1976:YAS

[KC76] D. F. Kerridge and G. W. Cook. Yet another series for the normal integral. Biometrika, 63(2):401–403, August 1976. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL

http://www.jstor.org/stable/2335636.

Knessl:2011:EAF

[KC11] Charles Knessl and Mark W. Coffey. An effective asymptotic formula for the Stieltjes constants. Mathematics of Computation, 80(273):379–386, January 2011. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.ams.

org/journals/mcom/2011-80-273/S0025-5718-2010-02390-
7/home.html;   http://www.ams.org/journals/mcom/2011-
80-273/S0025-5718-2010-02390-7/S0025-5718-2010-02390-
7.pdf.

Kong:2006:IGA

[KCYL06] Fanyu Kong, Zhun Cai, Jia Yu, and Daxing Li. Improved generalized Atkin algorithm for computing square roots in finite fields. Information Processing Letters, 98(1):1–5, April 15, 2006. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kearfott:1992:IPF

[KDDH92] Baker Kearfott, Milind Dawande, Kaisheng Du, and Chen-Yi Hu. INTLIB: a portable FORTRAN 77 elementary function library. Interval Computations = Interval’nye vychisleniia, 3 (5):96–105, 1992. ISSN 0135-4868. Interval ’92 (Moscow, 1992).

Kearfott:1994:AIP

[KDDH94] R. B. Kearfott, M. Dawande, K. Du, and C. Hu. Algorithm 737: INTLIB: a portable Fortran-77 elementary function library. ACM Transactions on Mathematical Software, 20(4):447–459, December 1994. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Keener:1982:CLB

[Kee82] L. L. Keener. Characterizing local best SAIN approximations. Journal of Approximation Theory, 36(1):55–63, September 1982. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Kennelly:1914:TCH

[Ken14] Arthur E. (Arthur Edwin) Kennelly. Tables of Complex Hyperbolic and Circular Functions. Harvard University Press, Cambridge, MA, USA, 1914. iii + 212 pp. LCCN QA342 .K45.

Kennelly:1921:TCH

[Ken21] Arthur Edwin Kennelly. Tables of Complex Hyperbolic and Circular Functions. Harvard University Press, Cambridge, MA, USA, 1921. iii + 240 pp. LCCN QA342 .K45 1921.

Kershaw:1983:SEW

[Ker83] D. Kershaw. Some extensions of W. Gautschi’s inequalities for the gamma function. Mathematics of Computation, 41(164): 607–611, October 1983. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kyurkchiev:2004:FCN

[KI04] N. Kyurkchiev and A. Iliev. Failure of convergence of the Newton–Weierstrass iterative method for simultaneous root finding of generalized polynomials. Computers and Mathematics with Applications, 47(2–3):441–446, January/February 2004. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122104900363.

Kutsuna:2016:ARM

[KI16] Takuro Kutsuna and Yoshinao Ishii. Abstraction and refinement of mathematical functions toward SMT-based test-case generation. International Journal on Software Tools for Technology Transfer (STTT), 18(1):109–120, February 2016. CODEN ???? ISSN 1433-2779 (print), 1433-2787 (electronic). URL http://

link.springer.com/article/10.1007/s10009-015-0389-7.

Kim:1972:AEH

[Kim72] Shoon K. Kim. The asymptotic expansion of a hypergeometric function _{2}F_{2}(1, α; ρ_{1}, ρ_{2}; z). Mathematics of Computation, 26 (120):963, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

King:1921:SNF

[Kin21] Louis Vessot King. On some new formulae for the numerical calculation of the mutual induction of coaxial circles. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 100(702):60–66, October 4, 1921. ISSN 0950-1207 (print), 2053-9150 (electronic). URL

http://www.jstor.org/stable/93861. This is the first known

publication of the AGM method, discovered by the author in 1913, for computing Jacobian elliptic functions. See also [Kin24, Kin07].

King:1924:DNC

[Kin24] Louis Vessot King. On the Direct Numerical Calculation of Elliptic Functions and Integrals. Cambridge University Press, Cambridge, UK, 1924. viii + 42 pp. LCCN QA343.

King:1965:LED

[Kin65] R. King. Letter to the Editor: On the double-precision square root routine. Communications of the ACM, 8(4):202, April 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

King:2007:DNC

[Kin07] Louis Vessot King. On the Direct Numerical Calculation of Elliptic Functions and Integrals. Mellon Press, 2007. ISBN

1-4067-4226-0. 56 pp. URL https://www.google.com/books/
edition/On_the_Direct_Numerical_Calculation_of_E/CMM4AAAAMAAJ.

Dedicated to the memory of James Harkness, Peter Redpath Professor of Pure Mathematics, McGill University 1903–1923.

Kirby:1988:ELA

[Kir88] James C. Kirby. An efficient logarithm algorithm for calculators. College Mathematics Journal, 19(3):257–260, May 1988. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/07468342.1988.11973125.

Kalantari:1996:HOI

[KK96] B. Kalantari and I. Kalantari. High order iterative methods for approximating square roots. BIT Numerical Mathematics, 36(2):395–399, June 1996. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

URL http://www.mai.liu.se/BIT/contents/bit36.html;
http://www.springerlink.com/openurl.asp?genre=article&
issn=0006-3835&volume=36&issue=2&spage=395.

Kiranon:1998:SRV

[KK98] W. Kiranon and N. Kumprasert. Square-rooting and vector summation circuits using current conveyors. IEE Proceedings on Circuits, Devices and Systems [see also IEE Proceedings G - Circuits, Devices and Systems], 145(2):139, April 1998. CODEN ???? ISSN ????

Koren:1999:ISC

[KK99] Israel Koren and Peter Kornerup, editors. 14th IEEE Symposium on Computer Arithmetic: proceedings: April 14–16, 1999, Adelaide, Australia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7803-5609-8, 0-7695-0116-8, 0-7695-0118-4. ISSN 1063-6889. LCCN QA76.6 .S887 1999. URL http://

computer.org/conferen/home/arith/;   http://www.ecs.
umass.edu/ece/arith14/program.html. IEEE Computer So-

ciety Order Number PR00116. IEEE Order Plan Catalog Number 99CB36336.

Korotkov:2020:IRE

[KK20] N. E. (Nikolaýi Efimovich) Korotkov and Alexander N. Korotkov. Integrals Related to the Error Function. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2020. ISBN 0-367-40820-1 (hardcover), 0-367-80923-0 (e-book), 1-000-03307-4 (e-book), 1-000-03308-2 (Mobipocket e-book), 1-000-03309-0 (e-Pub). xx + 227 pp. LCCN QA308 .K67 2020.

Kilpatrick:1967:CIP

[KKI67] J. E. Kilpatrick, Shigetoshi Katsura, and Yuji Inoue. Calculations of integrals of products of Bessel functions. Mathematics of Computation, 21(99):407–412, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Karagiannidis:2007:IAG

[KL07] George Karagiannidis and Athanasios Lioumpas. An improved approximation for the Gaussian Q-function. IEEE Communications Letters, 11(8):644–646, August 2007. CODEN ICLEF6. ISSN 1089-7798 (print), 1558-2558 (electronic). See corrections and comments [DD08].

Kupriianova:2014:MMF

[KL14] Olga Kupriianova and Christoph Lauter. Metalibm: A mathematical functions code generator. In Hong and Yap [HY14], pages 713–717. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (e-book). LCCN QA76.9.M35.

Kornerup:1991:PIS

[KM91] Peter Kornerup and David W. Matula, editors. Proceedings: 10th IEEE Symposium on Computer Arithmetic: June 26–28, 1991, Grenoble, France. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-9151-4 (case), 0-8186-6151-8 (microfiche), 0-7803-0187-0 (library binding). LCCN QA76.9.C62 S95 1991.

Karp:1993:HPD

[KM93] A. H. Karp and P. Markstein. High precision division and square root. Technical Report HPL-93-42, Hewlett–Packard Lab., Palo Alto, CA, USA, June 1993. 20 pp.

Knowles:1995:PSC

[KM95] Simon Knowles and William H. McAllister, editors. Proceedings of the 12th Symposium on Computer Arithmetic, July 19– 21, 1995, Bath, England. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7089-4 (paperback), 0-8186-7089-4 (case), 0-8186-7149-1 (microfiche), 0-8186-7089-4 (softbound), 0-7803-2949-X (casebound). LCCN QA 76.9 C62 S95 1995.

Karp:1997:HPD

[KM97] Alan H. Karp and Peter Markstein. High-precision division and square root. ACM Transactions on Mathematical Software, 23 (4):561–589, December 1997. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.

org/pubs/articles/journals/toms/forthcoming/a0-karp/
a0-karp.ps;   http://www.acm.org/pubs/citations/journals/
toms/1997-23-4/p561-karp/.

Kornerup:2007:PIS

[KM07] Peter Kornerup and Jean-Michel Muller, editors. Proceedings of the 18th IEEE Symposium on Computer Arithmetic, June 25– 27, 2007, Montpellier, France. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2007. ISBN 0-7695-2854-6. ISSN 1063-6889. LCCN QA76.9.C62.

URL http://www.lirmm.fr/arith18/.

Karp:1994:FPA

[KMB94] Alan H. Karp, Peter Markstein, and Dennis Brzezinski. Floating point arithmetic unit using modified Newton–Raphson technique for division and square root. US Patent 5,341,321, Au-

gust 23, 1994. URL http://patft.uspto.gov/netahtml/PTO/
search-bool.html;   https://patents.google.com/patent/

US5341321A. Patent filed 5 May 1993, granted to Hewlett-Packard Company on 23 August 1994. Patent expired 5-May-2013. See criticism in [Zim05].

Kojima:1989:PDI

[KMY89] M. Kojima, S. Mizuno, and A. Yoshise. A primal-dual interior point algorithm for linear programming. In Megiddo [Meg89b], pages 29–47. ISBN 0-387-96847-4. LCCN QA402.5 .P7851 1989. Proceedings of the conference held at the Asilomar conference center in Pacific Grove, California, March 1–4, 1987.

Kwan:1995:CII

[KNS95] H. Kwan, R. L. Nelson, Jr., and E. E. Swartzlander, Jr. Cascaded implementation of an iterative inverse-square-root algorithm, with overflow lookahead. In Knowles and McAllister [KM95], pages 115–122. CODEN ???? ISBN 0-8186-7089-4 (paperback), 0-8186-7089-4 (case), 0-8186-7149-1 (microfiche), 0-8186-7089-4 (softbound), 0-7803-2949-X (casebound). ISSN ???? LCCN QA 76.9 C62 S95 1995.

Knuth:1962:ECP

[Knu62a] Donald E. Knuth. Euler’s constant to 1271 places. Mathematics of Computation, 16(79):275–281, July 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Knuth:1962:EPC

[Knu62b] Donald E. Knuth. Evaluation of polynomials by computer. Communications of the ACM, 5(12):595–599, December 1962. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See letter [Knu63].

Knuth:1963:LEE

[Knu63] Donald E. Knuth. Letter to the Editor: Evaluation of polynomials by computer. Communications of the ACM, 6(2):51, February 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Knu62b].

Knuth:1998:EP

[Knu98a] Donald E. Knuth. Evaluation of polynomials, chapter 4, pages 485–524. Volume 2 of The Art of Computer Programming [Knu98b], third edition, 1998. ISBN 0-201-89684-2. LCCN QA76.6 .K64 1997. US$52.75.

Knuth:1998:SA

[Knu98b] Donald E. Knuth. Seminumerical Algorithms, volume 2 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, third edition, 1998. ISBN 0-201-89684-2. xiii + 762 pp. LCCN QA76.6 .K64 1997. US$52.75.

Khajah:1994:UHP

[KO94] H. G. Khajah and E. L. Ortiz. Ultra-high precision computations. Computers and Mathematics with Applications, 27(7): 41–57, April 1994. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0898122194901481.

Kodama:2007:RA

[Kod07] Masao Kodama. Remark on Algorithm 644. ACM Transactions on Mathematical Software, 33(4):28:1–28:3, August 2007. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Amo86, Amo90, Amo95].

Kodama:2008:ASP

[Kod08] Masao Kodama. Algorithm 877: a subroutine package for cylindrical functions of complex order and nonnegative argument. ACM Transactions on Mathematical Software, 34(4):22:1–22:21, July 2008. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kodama:2011:AMC

[Kod11] Masao Kodama. Algorithm 912: a module for calculating cylindrical functions of complex order and complex argument. ACM Transactions on Mathematical Software, 37(4):47:1–47:25, February 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Koelink:2016:AST

[Koe16] Erik Koelink. Applications of spectral theory to special functions. ArXiv e-prints, ??:1–63, December 2016. URL https://

arxiv.org/abs/1612.07035.

Kogbetliantz:1957:CEN

[Kog57] E. G. Kogbetliantz. Computation of e^{N} for −∞ < N < +∞ using an electronic computer. IBM Journal of Research and Development, 1(2):110–115, April 1957. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Kogbetliantz:1958:CANb

[Kog58a] E. G. Kogbetliantz. Computation of arcsin N for 0 < N < 1 using an electronic computer. IBM Journal of Research and Development, 2(3):218–222, July 1958. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Kogbetliantz:1958:CANa

[Kog58b] E. G. Kogbetliantz. Computation of arctan N for −∞ < N < +∞ using an electronic computer. IBM Journal of Research and Development, 2(1):43–53, January 1958. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Kogbetliantz:1959:CSC

[Kog59] E. G. Kogbetliantz. Computation of sin N, cos N, and M th root of N using an electronic computer. IBM Journal of Research and Development, 3(2):147–152, April 1959. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Kogbetliantz:1960:GEF

[Kog60] E. G. Kogbetliantz. Generation of elementary functions. In Ralston and Wilf [RW77], pages 7–35. ISBN 0-471-70690-6. LCCN QA39 .R26. Three volumes.

Kogbetliantz:1966:GEF

[Kog66] E. G. Kogbetliantz. Generation of elementary functions. In Ralston and Wilf [RW77], pages 7–35. ISBN 0-471-70690-6. LCCN QA39 .R26. Three volumes.

Kogan:1989:GBF

[Kog89] B. J. Kogan. General background of functional memory algorithms. Transactions of the Society for Computer Simulation, 5 (4):285–317, October 1989. CODEN TSCSEV. ISSN 0740-6797.

Kolbig:1968:ADS

[Köl68] K. S. Kölbig. Algorithm 327: Dilogarithm [S22]. Communications of the ACM, 11(4):270–271, April 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kolbig:1969:CASa

[Köl69a] K. S. Kölbig. Certification of Algorithm 292 [S22]: Regular Coulomb wave functions and of remark on Algorithm 292 [S22]: Regular Coulomb wave functions. Communications of the ACM, 12(5):278–279, May 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kolbig:1969:CASb

[Köl69b] K. S. Kölbig. Certification of Algorithm 300 [S22]: Coulomb wave functions. Communications of the ACM, 12(5):279–280, May 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kolbig:1969:RAS

[Köl69c] K. S. Kölbig. Remark on Algorithm 300 [S22]: Coulomb wave functions. Communications of the ACM, 12(12):692, December 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kolbig:1970:CZI

[Kol70] K. S. Kolbig. Complex zeros of an incomplete Riemann zeta function and of the incomplete gamma function. Mathematics of Computation, 24(111):679–696, July 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1972:CAC

[Köl72a] K. S. Kölbig. Certification of “Algorithm 363: Complex error function”. Communications of the ACM, 15(6):465–466, June 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gau69a].

Kolbig:1972:ZIG

[Kol72b] K. S. Kolbig. On the zeros of the incomplete gamma function. Mathematics of Computation, 26(119):751–755, July 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kolbig:1972:PCL

[Köl72c] K. S. Kölbig. Programs for computing the logarithm of the gamma function, and the digamma function, for complex argument. Computer Physics Communications, 4(2):221–226, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Kolbig:1972:RCC

[Köl72d] K. S. Kölbig. Remarks on the computation of Coulomb wavefunctions. Computer Physics Communications, 4(2):214–220, November 1972. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Kolbig:1984:PCL

[Köl84] K. S. Kölbig. Programs for computing the logarithm of the gamma function, and the digamma function, for complex argument. Computer Physics Communications, 35(1–3):C–152, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584824042.

Kolbig:1987:BRC

[Köl87] K. S. Kölbig. Book review: Calculation of Special Functions, the Gamma Function, the Exponential Integrals and Error-Like Functions (C. G. van der Laan and N. M. Temme). SIAM Review, 29(4):660–661, ???? 1987. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Kolbig:1990:BRC

[Köl90] K. S. Kölbig. Book review: Computation of Functions on Electronic Computers — Handbook (in Russian). Naukova Dumka, Kiev, 194, 599pp, by B. A. Popov, G. S. Tesler. Mathematics of Computation, 55(191):395–397, July 1990. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

www.jstor.org/stable/2008818.

Kolbig:1996:PF

[Köl96] K. S. Kölbig. The polygamma function ψ^{(k)}(x) for x = 1/4 and x = 3/4. Journal of Computational and Applied Mathematics, 75(1):43–46, November 12, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042796000556.

Kolbig:1997:TEH

[Köl97] K. S. Kölbig. Table errata: Handbook of elliptic integrals for engineers and scientists [Second edition, Springer, New York, 1971, MR 43 #3506] by P. F. Byrd and M. D. Friedman. Mathematics of Computation, 66(220):1767, October 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Kononova:1976:CEF

[Kon76] N. F. Kononova. The computation of elementary functions by means of polynomial approximations by the method of V. K. Dzjadik. (Russian). Vyčisl. Prikl. Mat. (Kiev), 29:27–39, 1976. ISSN 0321-4117.

Koopman:1991:PST

[Koo91] Philip J. Koopman, Jr., editor. The proceedings of the second and third annual workshops for the ACM Special Interest Group on Forth: SIGForth ’90, February 16–18, 1990, Dallas, Texas … SIGForth ’91, March 7–9, 1991, San Antonio, Texas. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-462-7. LCCN QA 76.73 F24 S53 1991. ACM order number 817911.

Koppelaar:1974:CRA

[Kop74] Henk Koppelaar. Certification and remark on “Algorithm 191: Hypergeometric”. Communications of the ACM, 17(10):589– 590, October 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Rel63b].

Korenev:2002:BFT

[Kor02] B. G. (Boris Grigorevich) Korenev. Bessel Functions and Their Applications. Analytical methods and special functions. Taylor and Francis, Boca Raton, FL, USA, 2002. ISBN 0-415-28130-X (hardcover). ix + 276 pp. LCCN QA408 .K67 2002.

Kornerup:2005:DSS

[Kor05] Peter Kornerup. Digit selection for SRT division and square root. IEEE Transactions on Computers, 54(3):294–303, March 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956

(electronic). URL http://csdl.computer.org/comp/trans/
tc/2005/03/t0294abs.htm;   http://csdl.computer.org/
dl/trans/tc/2005/03/t0294.htm;   http://csdl.computer.
org/dl/trans/tc/2005/03/t0294.pdf;   http://ieeexplore.
ieee.org/iel5/12/30205/01388194.pdf?isnumber=30205&prod=
JNL&arnumber=1388194&arSt=+294&ared=+303&arAuthor=Kornerup%
2C+P.;   http://ieeexplore.ieee.org/xpls/abs_all.jsp?
isnumber=30205&arnumber=1388194&count=13&index=5;   http:
//ieeexplore.ieee.org/xpls/references.jsp?arnumber=1388194.

Kormanyos:2011:APC

[Kor11] Christopher Kormanyos. Algorithm 910: a portable C++ multiple-precision system for special-function calculations. ACM Transactions on Mathematical Software, 37(4):45:1–45:27, February 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

King:1969:LEN

[KP69] Richard F. King and David L. Phillips. The logarithmic error and Newton’s method for the square root. Communications of the ACM, 12(2):87–88, February 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kioustelidis:1975:PLA

[KP75] J. B. Kioustelidis and J. K. Petrou. A piecewise linear approximation of log_{2} x with equal maximum errors in all intervals. IEEE Transactions on Computers, C-24(9):858–861, September 1975. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-

9956 (electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1672923.

Kuhlmann:1998:FLP

[KP98] M. Kuhlmann and K. K. Parhi. Fast low-power shared division and square-root architecture. In Proceedings of the 1998 International Conference on Computer Design: VLSI in Computers and Processors. ICCD ’98, pages 128–135. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. CODEN ???? ISSN ????

Kzaz:2003:CAG

[KP03] M. Kzaz and M. Prévost. Convergence acceleration of Gauss– Chebyshev quadrature formulae. Numerical Algorithms, 34 (2–4):379–391, December 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/ipsapp007.kluweronline.com/content/getfile/5058/48/
6/abstract.htm;   http://ipsapp007.kluweronline.com/
content/getfile/5058/48/6/fulltext.pdf.

Kiani:2008:AND

[KPPS08] M. Kiani, J. Panaretos, S. Psarakis, and M. Saleem. Approximations to the normal distribution function and an extended table for the mean range of the normal variables. J. Iran. Stat. Soc., 7(1):57–72, ???? 2008.

Kusterer:1979:SEP

[KR79] Roland Kusterer and Manfred Reimer. Stable evaluation of polynomials in time log n. Mathematics of Computation, 33(147): 1019–1031, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

S:1986:CEF

[KR86] A. S. Kuz’menko and K. I. Rogozin. Calculation of elementary functions in a number system with arbitrary basis on the basis of order-differential transformations. (Russian). Prace Nauk. Inst. Cybernet. Tech. Politech. Wrocław. Ser. Konfer., 74(31):259– 262, 1986.

Kravchuk:1985:ACE

[Kra85a] V. R. Kravchuk. Approximation of certain elementary functions by rational functions of order (n, 2). (Russian). Akad. Nauk Ukrain. SSR Inst. Mat. Preprint, 18:7–40, 1985.

Kravchuk:1985:EAE

[Kra85b] V. R. Kravchuk. Effective approximation of elementary functions by rational polynomials of order (n, 1). (Russian). Ukrainskii matematicheskii zhurnal, 37(2):175–180, 270, 1985. CODEN UMZHAA. ISSN 0041-6053.

Kramer:1987:ISR

[Krä87] W. Krämer. Inverse Standardfunktionen für reelle und komplexe Intervallargumente mit a priori Fehlerabschätzungen für beliebige Datenformate [English: Inverse Elementary Functions for Real and Complex Interval Arguments with A-Priori Error Estimates for Arbitrary Data Formats]. Dissertation, Universität Karlsruhe, Karlsruhe, Germany, 1987. ???? pp.

Kramer:1988:ISF

[Krä88] W. Krämer. Inverse standard functions for real and complex point and interval arguments with dynamic accuracy. Computing. Supplementum, pages 185–212, 1988. CODEN COSPDM. ISSN 0344-8029. Computer Arithmetic and Scientific Computation.

Kraaikamp:1989:SEP

[Kra89] Cor Kraaikamp. Statistic and ergodic properties of Minkowski’s diagonal continued fraction. Theoretical Computer Science, 65 (2):197–212, June 28, 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Kramer:1993:MPC

[Krä93] Walter Krämer. Multiple-precision computations with result verification. In Mathematics in Science and Engineering: Scientific Computing with Automatic Result Verification, volume 189, pages 325–356. Elsevier BV, Amsterdam, The Netherlands, 1993.

Krattenthaler:1995:HHM

[Kra95] C. Krattenthaler. HYP and HYPQ: Mathematica packages for the manipulation of binomial sums and hypergeometric series, respectively q-binomial sums and basic hypergeometric series. Journal of Symbolic Computation, 20(5–6):737–744, November– December 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Symbolic computation in combinatorics Δ_{1} (Ithaca, NY, 1993).

Kramer:1998:PWC

[Kra98] W. Kramer. A priori worst case error bounds for floating-point computations. IEEE Transactions on Computers, 47(7):750– 756, July 1998. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=709374. See [Tan92].

Krasikov:2014:ABA

[Kra14] Ilia Krasikov. Approximations for the Bessel and Airy functions with an explicit error term. LMS Journal of Computation and Mathematics, 17(1):209–225, 2014. CODEN ???? ISSN 1461-1570.

Kreider:1989:WSA

[Kre89] K. L. Kreider. A wave splitting approach to time dependent inverse scattering for the stratified cylinder. SIAM Journal on Applied Mathematics, 49(3):932–943, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Kilbas:2002:ARH

[KRT02] Anatoly A. Kilbas, Luis Rodríguez, and Juan J. Trujillo. Asymptotic representations for hypergeometric-Bessel type function and fractional integrals. Journal of Computational and Applied Mathematics, 149(2):469–487, December 15, 2002. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042702005629.

Krukier:1999:CAT

[Kru99] L. A. Krukier. Convergence acceleration of triangular iterative methods based on the skew-symmetric part of the matrix. Applied Numerical Mathematics: Transactions of IMACS, 30(2–3):281–290, June 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-bin/cas/tree/store/apnum/cas_
sub/browse/browse.cgi?year=1999&volume=30&issue=2-3&
aid=981.

Kravanja:1998:ZMS

[KRVZ98] P. Kravanja, O. Ragos, M. N. Vrahatis, and F. A. Zafiropoulos. ZEBEC: a mathematical software package for computing simple zeros of Bessel functions of real order and complex argument. Computer Physics Communications, 113(2–3):220–238, October 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000642.

Kowalski:1988:ASP

[KS88] Marek Kowalski and Waldemar Sielski. Approximation of smooth periodic functions in several variables. Journal of complexity, 4(4):356–372, December 1988. CODEN JOCOEH. ISSN 0885-064X (print), 1090-2708 (electronic).

Kilbas:2004:TTA

[KS04] A. A. (Anatolii Aleksandrovich) Kilbas and Megumi Saigo. Htransforms: theory and applications, volume 9 of Analytical methods and special functions. Chapman and Hall/CRC, Boca Raton, FL, USA, 2004. ISBN 0-203-48737-0, 0-415-29916-0, 1-58488-116-X. xii + 389 pp. LCCN ????

Koornwinder:2020:ESF

[KS20] T. H. Koornwinder and Jasper V. Stokman, editors. Encyclopedia of Special Functions: the Askey–Bateman Project. Volume 2. Multivariable Special Functions. Cambridge University Press, Cambridge, UK, 2020. ISBN 0-511-77716-7 (e-book), 1-107-00373-3 (hardcover). xii + 427 pp. LCCN QA351 .E63 2021.

Kuijlaars:2007:TIH

[KSVW07] A. B. J. Kuijlaars, H. Stahl, W. Van Assche, and F. Wielonsky. Type II Hermite–Padé approximation to the exponential function. Journal of Computational and Applied Mathematics, 207(2):227–244, October 15, 2007. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042706005978.

Kreyszig:1959:RUE

[KT59] Erwin Kreyszig and John Todd. The radius of univalence of the error function. Numerische Mathematik, 1:78–89, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Kilbas:2000:CFI

[KT00] A. A. Kilbas and J. J. Trujillo. Computation of fractional integrals via functions of hypergeometric and Bessel type. Journal of Computational and Applied Mathematics, 118(1–2):223– 239, June 1, 2000. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042700002910.

Kuipers:1952:PSE

[Kui52] L. Kuipers. Properties of some elementary functions. Nederl. Akad. Wetensch. Proc. Ser. A. = Indagationes Math., 55(14): 388–393, 1952.

Kuki:1966:CAE

[Kuk66] H. Kuki. Comments on the ANL evaluation of the OS/360 FOR-TRAN math function library. ???? SSD 169, C4773, SHARE Secretary Distribution, ????, 1966. 47–53 pp.

Kuki:1971:MFS

[Kuk71] H. Kuki. Mathematical function subprograms for basic system libraries—objectives, constraints, and trade-off. In Mathematical Software [Ric71], pages 187–199. ISBN 0-12-587250-X. LCCN QA1 .M26. Based on the proceedings of the Mathematical Software Symposium held at Purdue University, Lafayette, Indiana, USA, April 1–3, 1970.

Kuki:1972:AAC

[Kuk72a] Hirondo Kuki. ACM Algorithm 421: Complex gamma function with error control [S14]. Communications of the ACM, 15 (4):271–272, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuki:1972:CGF

[Kuk72b] Hirondo Kuki. Complex gamma function with error control [S14]. Communications of the ACM, 15(4):262–267, April 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kuliamin:2007:STI

[Kul07] V. V. Kuliamin. Standardization and testing of implementations of mathematical functions in floating point numbers. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 33(3):154–173, 2007. CODEN PC-SODA. ISSN 0361-7688 (print), 1608-3261 (electronic).

Kunz:1981:QZ

[Kun81] W. Kunz. Quadratwurzel mit dem μP Z80 [English: Square Roots with the Z80 Microprocessor]. Elektronik, 7:109–110, 1981. CODEN EKRKAR. ISSN 0013-5658.

Kuznetsov:2015:CTT

[Kuz15] A. Kuznetsov. Computing the truncated theta function via Mordell integral. Mathematics of Computation, 84(296): 2911–2926, 2015. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.ams.org/

journals/mcom/2015-84-296/S0025-5718-2015-02953-6;
http://www.ams.org/journals/mcom/2015-84-296/S0025-5718-
2015-02953-6/S0025-5718-2015-02953-6.pdf.

Koelink:2003:OPS

[KV03] Erik Koelink and Walter Van Assche, editors. Orthogonal polynomials and special functions: Leuven 2002, volume 1817 of Lecture notes in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-40375-2. ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN 33 33-06 33C 68W.

Kang:2021:NEE

[KW21] Hongchao Kang and Hong Wang. Numerical evaluation and error analysis of many different oscillatory Bessel transforms via confluent hypergeometric function. Applied Numerical Mathematics: Transactions of IMACS, 160(??):23–41, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0168927420302932.

Kalmykov:2007:AOEa

[KWY07a] M. Y. Kalmykov, B. F. L. Ward, and Y. Yost. All order ϵexpansion of Gauss hypergeometric functions with integer and half/integer values of parameters. Journal of High Energy Physics, 2007(02):040–??, ???? 2007. CODEN JHEPAB. ISSN 1126-6708.

Kalmykov:2007:AOEb

[KWY07b] M. Y. Kalmykov, B. F. L. Ward, and Y. Yost. On the all-order ϵ-expansion of generalized hypergeometric functions with integer values of parameters. Journal of High Energy Physics, 2007(11): 009, ???? 2007. CODEN JHEPAB. ISSN 1126-6708.

Kyriakopoulos:1974:GFH

[Kyr74] E. Kyriakopoulos. Generating functions of the hypergeometric functions. Journal of Mathematical Physics, 15(6):753–759, June 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v15/i6/p753_s1.

Koren:1990:EEF

[KZ90] I. Koren and O. Zinaty. Evaluating elementary functions in a numerical coprocessor based on rational approximations. IEEE Transactions on Computers, C-39(8):1030–1037, August 1990. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Kzaz:1992:CAS

[Kza92] M. Kzaz. Convergence acceleration of some Gaussian quadrature formulas for analytic functions. Applied Numerical Mathematics: Transactions of IMACS, 10(6):481–496, November 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kzaz:1999:CAG

[Kza99] M. Kzaz. Convergence acceleration of the Gauss–Laguerre quadrature formula. Applied Numerical Mathematics: Transactions of IMACS, 29(2):201–220, February 1, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.elsevier.com/cas/tree/store/
apnum/sub/1999/29/2/940.pdf.

L:1961:BRF

[L.61] Y. L. L. Book review: L. Fox, Tables of Weber Parabolic Cylinder Functions and Other Functions for Large Arguments, National Physical Laboratory Mathematical Tables Volume 4, Her Majesty’s Stationery Office, London, 1960, iii + 40 p., 28 cm. (Paperback). Mathematics of Computation, 15(75): 310–311, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.jstor.org/

stable/2002928;   https://www.ams.org/journals/mcom/
1961-15-075/S0025-5718-61-99215-8/S0025-5718-61-99215-
8.pdf.

Lo:1985:GPB

[LA85] Hao-Yung Lo and Y. Aoki. Generation of a precise binary logarithm with difference grouping programmable logic array. IEEE Transactions on Computers, C-34(8):681–691, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1676614.

Lang:2001:CRR

[LA01] Tomás Lang and Elisardo Antelo. Correctly rounded reciprocal square-root by digit recurrence and radix-4 implementation. In Burgess and Ciminiera [BC01], pages 83–93. CODEN ???? ISBN 0-7695-1150-3; 0-7695-1152-X. ISSN ???? LCCN QA76.9.C62 S95 2001. US$145. IEEE order no. PR01150.

Lang:2003:RRS

[LA03] Tomás Lang and Elisardo Antelo. Radix-4 reciprocal square-root and its combination with division and square root. IEEE Transactions on Computers, 52(9):1100–1114, September 2003. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

LaFara:1954:MCI

[LaF54] Robert L. LaFara. A method for calculating inverse trigonometric functions. Mathematical Tables and Other Aids to Computation, 8(47):132–139, July 1954. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Laforgia:1984:FIG

[Laf84] Andrea Laforgia. Further inequalities for the gamma function. Mathematics of Computation, 42(166):597–600, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Laforgia:1986:IBF

[Laf86] Andrea Laforgia. Inequalities for Bessel functions. Journal of Computational and Applied Mathematics, 15(1):75–81, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042786902396.

Laforgia:1991:BMB

[Laf91] Andrea Laforgia. Bounds for modified Bessel functions. Journal of Computational and Applied Mathematics, 34(3):263–267, April 26, 1991. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037704279190087Z.

LakshmanYN:1996:IPI

[Lak96] Lakshman Y.N., editor. ISSAC ’96: Proceedings of the 1996 International Symposium on Symbolic and Algebraic Computation, July 24–26, 1996, Zurich, Switzerland. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-796-0. LCCN QA 76.95 I59 1996.

Lancaster:1942:MME

[Lan42] Otis E. Lancaster. Machine method for the extraction of cube root. Journal of the American Statistical Association, 37(217):

112–115, March 1942. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/

stable/2279437.

Langer:1959:NAP

[Lan59] R. E. Langer, editor. On numerical approximation. Proceedings of a Symposium, Madison, April 21–23, 1958. The University of Wisconsin Press, Madison, WI, USA, 1959. x + 462 pp. LCCN QA3 .U45 no. 1. Publication no. 1 of the Mathematics Research Center, U.S. Army, the University of Wisconsin.

Lang:1960:ECC

[Lan60] H. A. Lang. On the evaluation of certain complex elliptic integrals (in Technical Notes and Short Papers). Mathematics of Computation, 14(70):195–199, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lanczos:1964:PAG

[Lan64] Cornelius Lanczos. A precision approximation of the gamma function. Journal of the Society for Industrial and Applied Mathematics: Series B, Numerical Analysis, 1(1):86–96, ???? 1964. ISSN 0887-459X (print), 1095-7170 (electronic). URL http://

www.jstor.org/stable/2949767.

Langebartel:1980:FER

[Lan80] R. G. Langebartel. Fourier expansions of rational fractions of elliptic integrals and Jacobian elliptic functions. SIAM journal on mathematical analysis, 11(3):506–513, May 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Larssen:1966:CAC

[Lar66] Gerhard Meidell Larssen. Certification of Algorithm 56: Complete elliptic integral of the second kind. Communications of the ACM, 9(1):12, January 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lardner:1969:RBB

[Lar69] Thomas J. Lardner. Relations between_{0}F_{3} and Bessel functions. SIAM Review, 11(1):69–72, ???? 1969. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lassey:1982:CCI

[Las82] Keith R. Lassey. On the computation of certain integrals containing the modified Bessel function I_{0}(ξ). Mathematics of Computation, 39(160):625–637, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lynch:1995:KTF

[LASC95] T. Lynch, A. Ahmed, M. Schulte, and T. Callaway. The K5 transcendental functions. In Knowles and McAllister [KM95], pages 163–171. ISBN 0-8186-7089-4 (paperback), 0-8186-7089-4 (case), 0-8186-7149-1 (microfiche), 0-8186-7089-4 (softbound), 0-7803-2949-X (casebound). LCCN QA 76.9 C62 S95 1995. URL

http://mesa.ece.wisc.edu/publications/cp_1995-04.pdf.

Laurenzi:1973:DWF

[Lau73] Bernard J. Laurenzi. Derivatives of Whittaker functions W_{k,1/2} and M_{k,1/2} with respect to order K. Mathematics of Computation, 27(121):129–132, January 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lau:1995:NLC

[Lau95] H. T. (Hang Tong) Lau. A Numerical Library in C for Scientists and Engineers. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1995. ISBN 0-8493-7376-X. xvii + 795 pp. LCCN QA76.73.C15 L38 1995. URL http://www.loc.

gov/catdir/enhancements/fy0744/94037928-d.html.

Lau:2004:NLJ

[Lau04] H. T. (Hang Tong) Lau. A Numerical Library in Java for Scientists and Engineers. Chapman and Hall/CRC, Boca Raton, FL, USA, 2004. ISBN 1-58488-430-4. xxiii + 1063 pp. LCCN

QA76.73.J38 L363 2004. URL http://www.loc.gov/catdir/
enhancements/fy0646/2003055149-d.html.

Lavington:1980:IPP

[Lav80] Simon Hugh Lavington, editor. Information Processing 80: Proceedings of IFIP Congress 80, Tokyo, Japan, October 6– 9, 1980, Melbourne, Australia, October 14–17, 1980. Elsevier North-Holland, Inc., New York, NY, USA, 1980. ISBN 0-444-86034-7. LCCN QA 75.5 I57 1980.

Lavoie:1986:SEG

[Lav86] J. L. Lavoie. Some evaluations for the generalized hypergeometric series. Mathematics of Computation, 46(173):215–218, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lazard:1985:PFD

[Laz85] Daniel Lazard. Primitives des fonctions élémentaires (d’après Risch et Davenport). (French) [primitives of elementary functions (following Risch and Davenport)] Seminar Bourbaki, Vol. 1983/84, No. 121-122. Astérisque, 121–122:295–308, 1985.

Levrie:1990:CAN

[LB90] Paul Levrie and Adhemar Bultheel. Convergence acceleration for the numerical solution of second-order linear recurrence relations. SIAM Journal on Numerical Analysis, 27(1):166–177, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Loskot:2009:PPA

[LB09] P. Loskot and N. C. Beaulieu. Prony and polynomial approximations for evaluation of the average probability of error over slowfading channels. IEEE Transactions on Vehicular Technology, 58(3):1269–1280, March 2009. CODEN ITUTAB. ISSN 0018-9545 (print), 1939-9359 (electronic). URL http://ieeexplore.

ieee.org/document/4529094/.

Lopez-Benitez:2011:VAA

[LBC11] Miguel Löpez-Benitez and Fernando Casadevall. Versatile, accurate, and analytically tractable approximation for the Gaussian Q-function. IEEE Transactions on Communications, 59(4):917– 922, April 2011. CODEN IECMBT. ISSN 0090-6778 (print),

1558-0857 (electronic). URL http://ieeexplore.ieee.org/
document/5706433/.

LeMaire:2016:CFP

[LBdDM16] Julien Le Maire, Nicolas Brunie, Florent de Dinechin, and Jean-Michel Muller. Computing floating-point logarithms with fixedpoint operations. In Montuschi et al. [MSH^{+}16], pages 156– 163. ISBN 1-5090-1615-5. ISSN 1063-6889. LCCN QA76.9.C62

S95 2016. URL http://ieeexplore.ieee.org/servlet/opac?
punumber=7562813.

Luke:1961:EHF

[LC61] Yudell L. Luke and Richard L. Coleman. Expansion of hypergeometric functions in series of other hypergeometric functions. Mathematics of Computation, 15(75):233–237, July 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lo:1987:HGA

[LC87] Hao-Yung Lo and Jau-Ling Chen. A hardwired generalized algorithm for generating the logarithm base-k by iteration. IEEE Transactions on Computers, C-36(11):1363–1367, November 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-

9956 (electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=5009477.

Li:1996:NNR

[LC96] Yamin Li and Wanming Chu. A new non-restoring square root algorithm and its VLSI implementations. In Proceedings of the IEEE International Conference on Computer Design: VLSI in Computers and Processors: ICCD ’96, pages 538–544. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. CODEN ???? ISSN ????

Li:1997:ISP

[LC97a] Yamin Li and Wanming Chu. Implementation of single precision floating point square root on FPGAs. In Proceedings of the 5th Annual IEEE Symposium on FPGAs for Custom Computing Machines, 16–18 April 1997, pages 226–232. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. CODEN ???? ISSN ????

Li:1997:PAI

[LC97b] Yamin Li and Wanming Chu. Parallel-array implementations of a non-restoring square root algorithm. In Proceedings of the 1997 IEEE International Conference on Computer Design: VLSI in Computers and Processors: ICCD ’97, pages 690–695. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. CODEN ???? ISSN ????

Ljusternik:1963:MVF

[LČJ63] L. A. Ljusternik, O. A. Červonenkis, and A. R. Janpolśkiĭ. Matematicheskiĭanaliz. Vychislenie èlementarnykh funktsiĭ.

[Mathematical analysis. Computation of the elementary functions]. Gosudarstv. Izdat. Fiz-Mat. Lit., Moscow, USSR, 1963. 247 pp.

Lyusternik:1965:HCE

[LCY65] L. A. Lyusternik, O. A. Chervonenkis, and A. R. Yanpolśkii. Handbook for Computing Elementary Functions, volume 76 of International series of monographs on pure and applied mathematics. Pergamon Press, New York, NY, USA, 1965. xiii + 251 pp. LCCN QA221.L513. Translated from the Russian by G. J. Tee. Translation edited by K. L. Stewart.

Lu:1989:VMI

[LD89] P. Y. Lu and K. Dawallu. A VLSI module for IEEE floatingpoint multiplication/division/square root. In IEEE ICCD ’89 [IEE89c], pages 366–368. ISBN 0-8186-1971-6 (paper), 0-8186-5971-8 (microfiche), 0-8186-8971-4 (case). LCCN TK 7888.4 I23 1989. IEEE catalog number 89CH2794-6.

Li:1993:CAF

[LDP93] Y. Li, X. Dong, and S. Pan. Computation of auxiliary functions in STO molecular integrals up to arbitrary accuracy. I. Evaluation of incomplete gamma function E_{n} (X) by forward recursion. International Journal of Quantum Chemistry, 45(1):3–??, 1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Louie:1993:DRS

[LE93] M. E. Louie and M. D. Ercegovac. A digit-recurrence square root implementation for field programmable gate arrays. In Proceedings of the IEEE Workshop on FPGAs for Custom Computing Machines, 5–7 April 1993, pages 178–183. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. CODEN ???? ISSN ????

Louie:1995:VPS

[LE95] Marianne E. Louie and Miloš D. Ercegovac. A variable-precision square root implementation for field programmable gate arrays. The Journal of Supercomputing, 9(3):315–336, September 1995. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0920-8542&volume=9&issue=3&spage=
315; http://www.wkap.nl/oasis.htm/95692.

Lebedev:1965:SFT

[Leb65] N. N. (Nikolaĭi Nikolaevich) Lebedev. Special Functions and Their Applications. Selected Russian publications in the mathematical sciences. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1965. xii + 308 pp. LCCN QA351 .L3613. Revised English edition translated and edited by Richard A. Silverman.

Lebedev:1972:SFT

[Leb72] N. N. (Nikolai Nikolaevich) Lebedev. Special functions and their applications. Dover, New York, NY, USA, 1972. ISBN 0-486-60624-4 (paperback). xii + 308 pp. LCCN QA351 .L3613 1972.

URL http://www.loc.gov/catdir/description/dover031/

72086228.html. Translated to English and edited by Richard A. Silverman.

Lee:1992:LCF

[Lee92] Chu-In Charles Lee. On Laplace continued fraction for the normal integral. Annals of the Institute of Statistical Mathematics, 44(1):107–120, March 1992. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (electronic). URL http://link.springer.

com/article/10.1007/BF00048673.

Lee:1997:PRF

[Lee97] M. Howard Lee. Polylogarithms and Riemann’s ζ function. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 56(4):3909–3912, October 1997. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic).

Lefevre:2005:NRD

[Lef05] Vincent Lefèvre. New results on the distance between a segment and Z^{2}. Application to the exact rounding. In IEEE [IEE05], page ?? ISBN ???? LCCN ???? URL http://arith17.

polito.it/final/paper-147.pdf.

Lehmer:1944:NCB

[Leh44] Derrick Henry Lehmer. Note on the computation of the Bessel function I_{n}(X). Mathematical Tables and Other Aids to Computation, 1(5):133–135, April 1944. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Lehmer:1956:RRZ

[Leh56] D. H. Lehmer. On the roots of the Riemann zeta-function. Acta Mathematica, 95(1):291–298, December 1956. CODEN ACMAA8. ISSN 0001-5962 (print), 1871-2509 (electronic).

Lehman:1970:DZR

[Leh70] R. S. Lehman. On the distribution of zeros of the Riemann zetafunction. Proceedings of the London Mathematical Society. First Series, 3(20):303–320, ???? 1970. ISSN 0024-6115 (print), 1460-244X (electronic). This paper corrects several errors in [Tur53]. See also [Tru11].

Lehmer:1971:CCM

[Leh71] D. H. Lehmer. On the compounding of certain means. Journal of Mathematical Analysis and Applications, 36(1):183–200, October 1971. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic).

Lehnhoff:1983:NPT

[Leh83a] H.-G Lehnhoff. A new proof of Teljakowskii’s theorem. Journal of Approximation Theory, 38(2):177–181, June 1983. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Lehnhoff:1983:SPF

[Leh83b] H.-G Lehnhoff. A simple proof of a A. F. Timan’s theorem. Journal of Approximation Theory, 38(2):172–176, June 1983. CO-DEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Lehoucq:1996:CEU

[Leh96] R. B. Lehoucq. The computation of elementary unitary matrices. ACM Transactions on Mathematical Software, 22(4):393–400, December 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lembarki:1988:CAL

[Lem88] Alami Lembarki. Convergence acceleration of limit k-periodic continued fractions. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4):337–349, June 1988. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lentz:1976:GBF

[Len76] William J. Lentz. Generating Bessel functions in Mie scattering calculations using continued fractions. Applied Optics, 15(3): 668–671, March 1976. CODEN APOPAI. ISSN 0003-6935.

Lentz:1990:CFC

[Len90] William J. Lentz. Continued fraction calculation of spherical Bessel functions. Computers in Physics, 4(4):403–407, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168382.

Lether:1996:RAF

[Let96] Frank G. Lether. Rational approximation formulas for computing the positive zeros of J_{0}(x). Journal of Computational and Applied Mathematics, 67(1):167–172, February 20, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042795002197.

Lether:1997:CNM

[Let97] Frank G. Lether. Constrained near-minimax rational approximations to Dawson’s integral. Applied Mathematics and Computation, 88(2–3):267–274, December 30, 1997. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S009630039600330X.

Lether:2001:VPA

[Let01] F. G. Lether. Variable precision algorithm for the numerical computation of the Fermi–Dirac function F_{j}(x) of order j = −3/2. Journal of Scientific Computing, 16(1): 69–79, March 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/content/pdf/10.1023/A%3A1011150530703;   http://
www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=16&issue=1&spage=69-79.

Levin:1973:DNL

[Lev73] D. Levin. Development of non-linear transformations for improving convergence of sequences. International Journal of Computer Mathematics, 3(1–4):371–388, ???? 1973. CODEN IJCMAT. ISSN 0020-7160.

Levrie:1991:CFC

[Lev91a] Paul Levrie. G-continued fractions and convergence acceleration in the solution of third-order linear recurrence relations of Poincaré type. Applied Numerical Mathematics: Transactions of IMACS, 8(3):225–242, October 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Levrie:1991:CAF

[Lev91b] Paul Levrie. Convergence acceleration for n-fractions. Applied Numerical Mathematics: Transactions of IMACS, 7(6):481–492, June 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lewis:1944:SCH

[Lew44] Charles J. Lewis. A survey of the confluent hypergeometric function. ????, Washington, DC, USA, 1944. 155 pp. LCCN QA351 .L53.

Lewin:1958:DAF

[Lew58] Leonard Lewin. Dilogarithms and Associated Functions. Macdonald, London, UK, 1958. 353 pp. LCCN QA351 .L5.

Lewis:1969:FII

[Lew69] Richard L. Lewis. On finite integrals involving trigonometric, Bessel, and Legendre functions. Mathematics of Computation, 23(106):259–273, April 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewis:1975:CPF

[Lew75] John Gregg Lewis. Certification of “Algorithm 349: Polygamma functions with arbitrary precision”. ACM Transactions on Mathematical Software, 1(4):380–382, December 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [TS69].

Lewin:1981:PAF

[Lew81] Leonard Lewin. Polylogarithms and Associated Functions. North-Holland, Amsterdam, The Netherlands, 1981. ISBN 0-444-00550-1. xvii + 359 pp. LCCN QA342 .L47.

Lewanowicz:1985:RRH

[Lew85] Stanisław Lewanowicz. Recurrence relations for hypergeometric functions of unit argument. Mathematics of Computation, 45 (172):521–535, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewanowicz:1987:CRR

[Lew87] Stanisław Lewanowicz. Corrigendum: “Recurrence relations for hypergeometric functions of unit argument” [Math. Comp. 45 (1985), no. 172, 521–535, MR 86m:33004]. Mathematics of Computation, 48(178):853, April 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lewin:1991:SPP

[Lew91] Leonard Lewin, editor. Structural Properties of Polylogarithms, volume 37 of Mathematical surveys and monographs. American Mathematical Society, Providence, RI, USA, 1991. ISBN 0-8218-1634-9, 1-4704-1264-0 (e-book). ISSN 0076-5376. xviii + 412 pp.

Lewanowicz:1994:SAS

[Lew94] Stanisław Lewanowicz. A simple approach to the summation of certain slowly convergent series. Mathematics of Computation, 63(208):741–745, October 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Liu:1992:QBS

[LF92] K. J. R. Liu and E. Frantzeskakis. Qrd-based square root free and division free algorithms and architectures. In Workshop on VLSI Signal Processing, V, 1992, pages 459–468. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. CODEN ???? ISSN ????

Lebedev:1960:GMT

[LFB60] A. V. (Aleksandr Vasil’evich) Lebedev, R. M. (Rimma Maksimova) Fedorova, and Nina Mikhollovna Burunova. A Guide to Mathematical Tables. Pergamon Press, New York, NY, USA, 1960. xlvi + 586 pp. LCCN Z6654.T3 L42. English edition prepared by D. G. Fry from the Russian original.

Lu:2014:GAF

[LFM14] Dawei Lu, Jinghai Feng, and Congxu Ma. A general asymptotic formula of the gamma function based on the Burnside’s formula. Journal of Number Theory, 145(??):317–328, December 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14002224.

Lotsch:1964:AFI

[LG64] Helmut Lotsch and Malcolm Gray. Algorithm 244: Fresnel integrals [S20]. Communications of the ACM, 7(11):660–661, November 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Leung:1979:AFE

[LG79] K. V. Leung and S. S. Ghaderpanah. An application of the finite element approximation method to find the complex zeros of the modified Bessel function K_{n}(z). Mathematics of Computation, 33(148):1299–1306, October 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lee:2005:OHF

[LGML05] Dong-U. Lee, Altaf Abdul Gaffar, Oskar Mencer, and Wayne Luk. Optimizing hardware function evaluation. IEEE Transactions on Computers, 54(12):1520–1531, December 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lide:2001:CEM

[Lid01] David R. Lide, editor. A Century of Excellence in Measurements, Standards, and Technology, volume 958 of NIST special publication. National Institute for Standards and Technology, Gaithersburg, MD, USA, 2001. ix + 386 pp.

Lide:2002:CEM

[Lid02] David R. Lide, editor. A Century of Excellence in Measurements, Standards, and Technology. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2002. ISBN 0-8493-1247-7. ix + 386 pp. LCCN QC100.U6

.C46 2002. URL http://www.loc.gov/catdir/toc/fy031/

2002283707.html. Republication of [Lid01].

Linskii:1957:CEF

[Lin57] V. S. Linskiĭ. Calculation of elementary functions on automatic digital machines. (Russian). Vyčisl. Mat., 2:90–119, 1957.

Linz:1972:MCB

[Lin72] Peter Linz. A method for computing Bessel function integrals. Mathematics of Computation, 26(118):509–513, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ling:1978:EWZ

[Lin78] Chin-Bing Ling. Evaluation of Weierstrass zeta functions. SIAM Review, 20(1):183, ???? 1978. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lindstrom:1979:MSM

[Lin79a] F. T. Lindstrom. A modified 3-spline method for evaluating the Euler digamma function. Technometrics, 21(3):307–311, August 1979. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1267752.

Ling:1979:EWZ

[Lin79b] C. B. Ling. Evaluation of Weierstrass zeta functions. SIAM Review, 21(1):146–147, ???? 1979. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Lin:1989:ANT

[Lin89] Jinn Tyan Lin. Approximating the normal tail probability and its inverse for use on a pocket calculator. Applied Statistics, 38 (1):69–70, 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Lin:1990:MSL

[Lin90] Jinn Tyan Lin. Miscellanea: a simpler logistic approximation to the normal tail probability and its inverse. Applied Statistics, 39 (2):255–257, 1990. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Linhart:2009:ACL

[Lin09] Jean Marie Linhart. Algorithm 885: Computing the logarithm of the normal distribution. ACM Transactions on Mathematical Software, 35(3):20:1–20:10, October 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lipovetsky:2021:BRI

[Lip21] Stan Lipovetsky. Book review: Integrals Related to the Error Function, by Nikolai E. Korotkov and Alexander N. Korotkov. Boca Raton, FL: Chapman and Hall/CRC Press, Taylor & Francis Group, 2020, 228 pp., $140.00 (hardback), $46.36 (eBook), ISBN: 978-0-367-40820-6 (hardback). Technometrics, 62(4):560, 2021. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liron:1971:ISR

[Lir71] N. Liron. Infinite sums of roots for a class of transcendental equations and Bessel functions of order one-half. Mathematics of Computation, 25(116):769–781, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Little:1983:CCS

[Lit83] F. Little. Convex combination surfaces. In Robert E. Barnhill and Wolfgang Boehm, editors, Surfaces in computer aided geometric design: proceedings of a conference held at Mathematisches Forschungsinstitut Oberwolfach, F.R.G., April 25–30, 1982, organized by Wolfgang Boehm and Josef Hoschek, pages 99–109. North-Holland, Amsterdam, The Netherlands, 1983. ISBN 0-444-86550-0. LCCN T385 .S827 1982.

Littlestone:1989:MBL

[Lit89] N. Littlestone. Mistake bounds and logarithmic linear-threshold learning algorithms. Ph.D thesis, University of California at Santa Cruz, Santa Cruz, CA, USA, 1989. ???? pp.

Litvinov:1993:ACR

[Lit93] Grigori L. Litvinov. Approximate construction of rational approximations and the effect of error autocorrection. Technical report 8, Institute of Mathematics, University of Oslo, Oslo, Norway, May 1993. See [Lit94].

Litvinov:1994:ACR

[Lit94] Grigori L. Litvinov. Approximate construction of rational approximations and the effect of error autocorrection. Russian Journal of Mathematical Physics, 1(3):313–352, ???? 1994. CO-DEN RJMPEL. ISSN 1061-9208 (print), 1555-6638 (electronic).

URL http://arxiv.org/abs/math/0101042.

Liu:1987:BEF

[Liu87] Z. A. Liu. Berkeley elementary function test suite. M.S. thesis, Computer Science Division, Department of Electrical Engineering and Computer Science, Univerity of California at Berkeley, Berkeley, CA, USA, December 1987.

Liu:1993:DSC

[Liu93] Hui Min Liu. Determination of several classes of elementary functions by functional inequalities. (Chinese). Hunan Jiaoyu Xueyuan Xuebao (Ziran Kexue), 11(2):30–35, 12, 1993. ISSN 1001-6074.

Liu:1995:SRV

[Liu95] S.-I. Liu. Square-rooting and vector summation circuits using current conveyors. IEE Proceedings on Circuits, Devices and Systems [see also IEE Proceedings G - Circuits, Devices and Systems], 142(4):223–226, August 1995. CODEN ???? ISSN ????

Low:2013:MET

[LJ13] Joshua Yung Lih Low and Ching Chuen Jong. A memoryefficient tables-and-additions method for accurate computation of elementary functions. IEEE Transactions on Computers, 62 (5):858–872, May 2013. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Linz:1973:NCI

[LK73] Peter Linz and T. E. Kropp. A note on the computation of integrals involving products of trigonometric and Bessel functions. Mathematics of Computation, 27(124):871–872, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Langdon:2021:GID

[LK21] William B. Langdon and Oliver Krauss. Genetic improvement of data for maths functions. ACM Transactions on Evolutionary Learning and Optimization (TELO), 1(2):7:1–7:30, June 2021. CODEN ???? ISSN 2688-299X (print), 2688-3007 (electronic).

URL https://dl.acm.org/doi/10.1145/3461016.

Li:2010:NRB

[LKF10] Rong Li, Pooi Yuen Kam, and Hua Fu. New representations and bounds for the generalized Marcum Q-function via a geometric approach, and an application. IEEE Transactions on Communications, 58(1):157–169, January 2010. CODEN IECMBT. ISSN 0090-6778 (print), 1558-0857 (electronic). URL http://

ieeexplore.ieee.org/document/5397910/.

Lee:1993:DAE

[LKH93] Joong-Eon Lee, Oh-Young Kwon, and Tack-Don Han. Design of an area efficient unit for floating-point division and square root. Journal of the Korea Information Science Society = Chongbo Kwahakhoe nonmunji, 20(7):1060–1071, July 1993. CODEN HJKHDC. ISSN 0258-9125.

Lemire:2019:FRD

[LKK19] Daniel Lemire, Owen Kaser, and Nathan Kurz. Faster remainder by direct computation: Applications to compilers and software libraries. Software—Practice and Experience, 49(6):953– 970, June 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lee:1999:STS

[LKKH99] Young-Sang Lee, Jun-Woo Kang, Lee-Sup Kim, and Seung-Ho Hwang. Self-timed shared division and square-root implementation using full redundant signed digit numbers. In 6th International Conference on VLSI and CAD: ICVC ’99, pages 541–544. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. CODEN ???? ISSN ????

Li:2002:SWF

[LKL02] Le-Wei Li, Xiao-Kang Kang, and Mook-Seng Leong. Spheroidal Wave Functions in Electromagnetic Theory. Wiley series in microwave and optical engineering. Wiley, New York, NY, USA, 2002. ISBN 0-471-03170-4 (hardcover). xiii + 295 pp.

LCCN QC670 .L49 2002. URL http://www.loc.gov/catdir/
bios/wiley043/2001045399.html;   http://www.loc.gov/
catdir/description/wiley036/2001045399.html;   http:
//www.loc.gov/catdir/toc/onix07/2001045399.html.

Ling:1972:EM

[LL72] Chih-Bing Ling and Jung Lin. On evaluation of moments of K_{ν}(t)/I_{ν}(t). Mathematics of Computation, 26(118):529–537, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lauter:2009:ERB

[LL09] C. Q. Lauter and V. Lefevre. An efficient rounding boundary test for pow(x, y) in double precision. IEEE Transactions on Computers, 58(2):197–207, February 2009. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Li:2014:ICH

[LLK14] Dingfang Li, Ping Liu, and Jisheng Kou. An improvement of Chebyshev–Halley methods free from second derivative. Applied Mathematics and Computation, 235(??):221–225, May 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300314003312.

Lang:1992:HRS

[LM92] T. Lang and P. Montuschi. Higher radix square root with prescaling. IEEE Transactions on Computers, 41(8):996–1009, August 1992. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Laforgia:1993:AMR

[LM93] Andrea Laforgia and Maria Luisa Mathis. Additional monotonicity results for the zeros of Bessel functions. Journal of Computational and Applied Mathematics, 47(1):135–139, June 28, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037704279390095S.

Lang:1995:VHR

[LM95] T. Lang and P. Montuschi. Very-high radix combined division and square root with prescaling and selection by rounding. In Knowles and McAllister [KM95], pages 124–131. CODEN ???? ISBN 0-8186-7089-4 (paperback), 0-8186-7089-4 (case), 0-8186-7149-1 (microfiche), 0-8186-7089-4 (softbound), 0-7803-2949-X (casebound). ISSN ???? LCCN QA 76.9 C62 S95 1995.

Lang:1999:VHR

[LM99] T. Lang and P. Montuschi. Very high radix square root with prescaling and rounding and a combined division/square root unit. IEEE Transactions on Computers, 48(8):827–841, August 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lefevre:2000:CRF

[LM00] V. D. Lefevre and J.-M. Muller. Correctly rounded functions for better arithmetic. In Conference Record of the Thirty-Fourth Asilomar Conference on Signals, Systems and Computers, 2000, volume 2, pages 875–878. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. CODEN ???? ISSN ????

Lefevre:2003:WCC

[LM03] Vincent Lefèvre and Jean-Michel Muller. Worst cases for correct rounding for the elementary functions in double precision. Technical report, INRIA, Projet Spaces, LORIA, Campus Scientifique, B.P. 239, 54506 Vandoeuvre-lès-Nancy Cedex,

France, August 14, 2003. URL http://perso.ens-lyon.fr/
jean-michel.muller/TMDworstcases.pdf.

Lefevre:2007:SNP

[LM07] Vincent Lefévre and Jean-Michel Muller. Some notes on the possible under/overflow of the most common elementary functions. Report, LIP, École Normale Supérieure de Lyon, Lyon,

France, 2007. 7 pp. URL http://prunel.ccsd.cnrs.fr/ensl-
00149414.

Lorch:2008:MSR

[LM08] Lee Lorch and Martin E. Muldoon. Monotonic sequences related to zeros of Bessel functions. Numerical Algorithms, 49 (1–4):221–233, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
1017-1398&volume=49&issue=1&spage=221.

Lauter:2015:SAF

[LM15a] Christoph Lauter and Marc Mezzarobba. Semi-automatic floating-point implementation of special functions. In Muller et al. [MTV15], pages 58–65. ISBN 1-4799-8665-8, 1-4799-8663-1. ISSN 1063-6889. LCCN QA76.9.C62 S95 2015. URL http:/

/ieeexplore.ieee.org/servlet/opac?punumber=7193754.

Lu:2015:SNQ

[LM15b] Dawei Lu and Congxu Ma. Some new quicker continued fraction approximations for the gamma function related to the Nemes’ formula. Numerical Algorithms, 70(4):825–833, December 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11075-015-9975-8.

Li:2001:LLF

[LMOT01] Ren-Cang Li, Peter Markstein, Jon P. Okada, and James W. Thomas. The libm library and floating-point arithmetic for HP-UX on Itanium. Technical report, Hewlett–Packard Company, Palo Alto, CA, USA, April 2001. ?? pp. URL http:

//h21007.www2.hp.com/dspp/ddl/ddl_Download_File_TRX/
1,1249,942,00.pdf;   http://h21007.www2.hp.com/dspp/
tech/tech_TechDocumentDetailPage_IDX/1,1701,981,00.html.

Lozier:1973:BCP

[LMS73] D. W. Lozier, L. C. Maximon, and W. L. Sadowski. A bit comparison program for algorithm testing. The Computer Journal, 16(2):111–117, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/160111.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_02/tiff/111.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/112.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/113.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_16/Issue_02/tiff/114.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_02/tiff/115.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_02/tiff/116.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/117.tif.

Lozier:1999:DDM

[LMS99] Daniel W. Lozier, B. R. Miller, and B. V. Saunders. Design of a digital mathematical library for science, technology and education. In IEEE [IEE99], pages 118–128. ISBN 0-7695-0219-

9. LCCN ZA4080 .F67 1999. URL http://dlmf.nist.gov/

about/publications/nistir6297.ps.gz.

Lynch:1992:HSD

[LMT^{+}92] T. Lynch, S. McIntyre, K. Tseng, S. Shaw, and T. Hurson. High speed divider with square root capability, 1992. U.S. Patent No. 5,128,891.

Lang:1997:ISC

[LMT97] Tomas Lang, Jean-Michel Muller, and Naofumi Takagi, editors. 13th IEEE Symposium on Computer Arithmetic: proceedings, July 6–9, 1997, Asilomar, California, USA, volume 13 of Symposium on Computer Arithmetic. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7846-1, 0-8186-7847-X, 0-8186-7848-8. ISSN 1063-6889. LCCN QA76.9.C62 S95 1997. IEEE Computer Society order number PR07846. IEEE Order Plan catalog number 97CB36091.

Lefevre:1998:TCR

[LMT98] V. Lefèvre, J.-M. Muller, and A. Tisserand. Toward correctly rounded transcendentals. IEEE Transactions on Computers, 47 (11):1235–1243, November 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Lim:2021:HPC

[LN21] Jay P. Lim and Santosh Nagarakatte. High performance correctly rounded math libraries for 32-bit floating point representations. In Proceedings of the 42nd ACM SIGPLAN International Conference on Programming Language Design and Implementation. ACM Press, New York, NY 10036, USA, June

2021. URL https://github.com/rutgers-apl/rlibm-32.

Lim:2022:OPA

[LN22] Jay P. Lim and Santosh Nagarakatte. One polynomial approximation to produce correctly rounded results of an elementary function for multiple representations and rounding modes. Proceedings of the ACM on Programming Languages (PACMPL), 6(POPL):3:1–3:28, January 2022. CODEN ???? ISSN 2475-

1421 (electronic). URL https://dl.acm.org/doi/10.1145/
3498664.

Lozier:1993:ABF

[LO93] Daniel W. Lozier and F. W. J. Olver. Airy and Bessel functions by parallel integration of ODEs. In Sincovec et al. [SKL93], pages 530–538. ISBN 0-89871-315-3. LCCN QA76.58 .S55 1993 v.1-2. Two volumes.

Lozier:1994:NESa

[LO94a] D. W. Lozier and F. W. J. Olver. Numerical evaluation of special functions. Report NISTIR 5383, Computing and Applied Mathematics Laboratory, U. S. Department of Commerce, Washington, DC, USA, March 1994. 47 pp. URL https://math.nist.

gov/~DLozier/publications/nistir5383.pdf.

Lozier:1994:NESb

[LO94b] D. W. Lozier and F. W. J. Olver. Numerical evaluation of special functions. In Gautschi [Gau94], pages 79–125. ISBN 0-8218-0291-7, 0-8218-0353-0 (pt. 1), 0-8218-0354-9 (pt. 2). ISSN 0160-7634. LCCN QA1 .A56 v.48 1994; QA297.M385 1993. URL

http://math.nist.gov/mcsd/Reports/2001/nesf/. See also

SIAM Review, September 1995, 37(3), p. 483.

Leeser:1995:VSR

[LO95] M. Leeser and J. O’Leary. Verification of a subtractive radix-2 square root algorithm and implementation. In Proceedings of the IEEE International Conference on Computer Design: VLSI in Computers and Processors, ICCD ’95, pages 526–531. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN ???? ISSN ????

Loenko:2001:CEF

[Loe01] M. Yu. Loenko. Computation of elementary functions with guaranteed accuracy. Programmirovanie, 2:68–80, 2001. CODEN PROGD3. ISSN 0132-3474, 0361-7688.

Longman:1959:STT

[Lon59] I. M. Longman. A short table of ∫_{x}^{∞} J_{0}(t)t^{−n}dt and ∫_{x}^{∞} J_{1}(t)t^{−n}dt (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 13(68):306–311, October 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Longman:1981:DCA

[Lon81] I. M. Longman. Difficulties of convergence acceleration. In Padé approximation and its applications, Amsterdam 1980 (Amsterdam, 1980), volume 888 of Lecture Notes in Mathematics, pages 273–289. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981.

Lopez:1998:SSC

[Lóp98] JoséL. López. Several series containing gamma and polygamma functions. Journal of Computational and Applied Mathematics, 90(1):15–23, April 6, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0377042798000077.

Lopez:2000:AES

[Lóp00] José L. López. Asymptotic expansions of symmetric standard elliptic integrals. SIAM journal on mathematical analysis, 31(4):

754–775, 2000. CODEN SJMAAH. ISSN 0036-1410 (print),

1095-7154 (electronic). URL http://epubs.siam.org/sam-
bin/dbq/article/35117.

Lopez:2018:CEB

[Lóp18] José L. López. Convergent expansions of the Bessel functions in terms of elementary functions. Advances in Computational Mathematics, 44(1):277–294, February 2018. CO-DEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic).

URL http://link.springer.com/article/10.1007/s10444-
017-9543-y.

Lorentz:1989:NA

[Lor89] G. G. Lorentz. Notes on approximation. Journal of Approximation Theory, 56(3):360–365, March 1989. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Lovelace:1843:SAE

[Lov43] Ada Augusta Lovelace. Sketch of the Analytical Engine. Scientific Memoirs, 3(??):666–731, ???? 1843. Reprinted in [Lov89].

Lovelace:1953:AET

[Lov53] Ada Augusta Lovelace. Appendix 1: Extracts from Taylor’s Scientific Memoirs, Vol. III. In Bowden [Bow53a], pages 341–408.

LCCN QA76.5 .B66. URL https://archive.org/details/

FasterThanThought. Reprint of [Lov43]. Pages 400–408 describe the computation of the Bernoulli numbers.

Lovelace:1989:SAE

[Lov89] Augusta Ada Lovelace. Sketch of the Analytical Engine (1843). In Campbell-Kelly [CK89a], pages 89–170. ISBN 1-85196-503-3, 1-85196-005-8 (set). LCCN ????

Lowan:1949:CLN

[Low49] Arnold N. Lowan. The Computation Laboratory of the National Bureau of Standards. Scripta Mathematica, 15(??):33–63, ???? 1949. ISSN 0036-9713.

Lowan:1964:SWF

[Low64] Arnold N. Lowan. Spheroidal wave functions. In Abramowitz and Stegun [AS64], pages 751–770. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://

www.convertit.com/Go/ConvertIt/Reference/AMS55.ASP in

bitmap image format.

Lozier:1994:SNS

[Loz94] Daniel W. Lozier. Software needs in special functions. Technical Report NISTIR 5490, National Institute for Standards and Technology, Gaithersburg, MD, USA, August 1994.

16 pp. URL http://math.nist.gov/acmd/Staff/DLozier/
publications/nistir5490.ps. Published in [Loz96b].

Lozier:1996:PST

[Loz96a] Daniel W. Lozier. A proposed software test service for special functions. Technical Report NISTIR 5916, National Institute for Standards and Technology, Gaithersburg, MD, USA,

October 1996. 11 pp. URL http://math.nist.gov/acmd/
Staff/DLozier/publications/nistir5916.ps. Published in

[Loz97a].

Lozier:1996:SNS

[Loz96b] Daniel W. Lozier. Software needs in special functions. Journal of Computational and Applied Mathematics, 66(??):345–358, ???? 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Lozier:1997:PST

[Loz97a] Daniel W. Lozier. A proposed software test service for special functions. In Boisvert [Boi97], pages 167–178. ISBN 0-412-80530-8. LCCN QA297 .I35 1996.

Lozier:1997:TRN

[Loz97b] Daniel W. Lozier. Toward a revised NBS handbook of mathematical functions. Technical Report NISTIR 6072, National Institute for Standards and Technology, Gaithersburg, MD, USA,

September 1997. 8 pp. URL http://math.nist.gov/acmd/
Staff/DLozier/publications/nistir6072.ps.gz.

Lozier:2000:DPN

[Loz00] Daniel W. Lozier. The DLMF Project: a new initiative in classical special functions. In Dunkl et al. [DIW00], pages 207–220. ISBN 981-02-4393-6. LCCN QA351 .P76 2000.

Lozier:2003:NDL

[Loz03] Daniel W. Lozier. NIST Digital Library of Mathematical Functions. Annals of Mathematics and Artificial Intelligence, 38(1– 3):105–119, May 2003. CODEN AMAIEC. ISSN 1012-2443 (print), 1573-7470 (electronic). URL http://math.nist.gov/

acmd/Staff/DLozier/publications/Linz01.ps.

Landau:1961:PSW

[LP61] H. J. Landau and H. O. Pollak. Prolate spheroidal wave functions, Fourier analysis and uncertainty. II. The Bell System Technical Journal, 40(1):65–84, January 1961. CODEN BST-

JAN. ISSN 0005-8580. URL http://bstj.bell-labs.com/
BSTJ/images/Vol40/bstj40-1-65.pdf.

Landau:1962:PSW

[LP62] H. J. Landau and H. O. Pollak. Prolate spheroidal wave functions, Fourier analysis and uncertainty. III. The dimension of the space of essentially time- and band-limited signals. The Bell System Technical Journal, 41(4):1295–1336, July 1962. CODEN BSTJAN. ISSN 0005-8580. URL http://bstj.bell-labs.

com/BSTJ/images/Vol41/bstj41-4-1295.pdf.

Lutskii:1983:VFT

[LP83] G. M. Lutskiĭ and O. I. Penchev. Vychislenie èlementarnykh funktsiĭ metodom tsifra za tsifroĭ v izbytochnykh sistemakh schisleniya. (Russian) [calculation of elementary functions by the digit-by-digit method in redundant number systems]. Preprint 83-22, Akad. Nauk Ukrain. SSR, Inst. Kibernet., Kiev, USSR, 1983. 30 pp.

Levrie:1988:CAM

[LP88] Paul Levrie and Robert Piessens. Convergence acceleration for Miller’s algorithm. In Nonlinear numerical methods and rational approximation (Wilrijk, 1987), volume 43 of Math. Appl., pages 349–370. Reidel, Dordrecht, The Netherlands, 1988.

Lewanowicz:1995:AMC

[LP95] Stanisław Lewanowicz and Stefan Paszkowski. An analytic method for convergence acceleration of certain hypergeometric series. Mathematics of Computation, 64(210):691–713, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Langhammer:2016:SPN

[LP16] Martin Langhammer and Bogdan Pasca. Single precision natural logarithm architecture for hard floating-point and DSPenabled FPGAs. In Montuschi et al. [MSH^{+}16], pages 164– 171. ISBN 1-5090-1615-5. ISSN 1063-6889. LCCN QA76.9.C62

S95 2016. URL http://ieeexplore.ieee.org/servlet/opac?
punumber=7562813.

Langhammer:2017:FPT

[LP17a] M. Langhammer and B. Pasca. Floating point tangent implementation for FPGAs. In Burgess et al. [BBdD17], pages 64– 65. ISBN 1-5386-1966-0 (print), 1-5386-1965-2, 1-5386-1964-4. ISSN 1063-6889. LCCN QA76.9.C62 S95 2017. URL http://

ieeexplore.ieee.org/servlet/opac?punumber=8019911.

Langhammer:2017:SPL

[LP17b] Martin Langhammer and Bogdan Pasca. Single precision logarithm and exponential architectures for hard floating-point enabled FPGAs. IEEE Transactions on Computers, 66(12):2031– 2043, ???? 2017. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
document/7927449/.

Latham:1974:CPC

[LR74] W. P. Latham and Rogers W. Redding. On the calculation of the parabolic cylinder functions. Journal of Computational Physics, 16(1):66–75, September 1974. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0021999174901041.

Laforgia:1988:MRI

[LS88] Andrea Laforgia and Silvana Sismondi. Monotonicity results and inequalities for the gamma and error functions. Journal of Computational and Applied Mathematics, 23(1):25–33, July 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042788903287.

Lucas:1995:EII

[LS95] S. K. Lucas and H. A. Stone. Evaluating infinite integrals involving Bessel functions of arbitrary order. Journal of Computational and Applied Mathematics, 64(3):217–231, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/B6TYH-4002HHC-J/2/54f1e67d9bea3e951acc3c39556ab452.

Lu:2014:GAG

[LSM14] Dawei Lu, Lixin Song, and Congxu Ma. A generated approximation of the gamma function related to Windschitl’s formula. Journal of Number Theory, 140(??):215–225, July 2014. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000687.

Lu:2015:SNA

[LSM15] Dawei Lu, Lixin Song, and Congxu Ma. Some new asymptotic approximations of the gamma function based on Nemes’ formula, Ramanujan’s formula and Burnside’s formula. Applied Mathematics and Computation, 253(??):1–7, February 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300314017317.

Lu:2016:QCF

[LSM16] Dawei Lu, Lixin Song, and Congxu Ma. A quicker continued fraction approximation of the gamma function related to the Windschitl’s formula. Numerical Algorithms, 72(4):865–874, August 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://link.springer.com/article/
10.1007/s11075-015-0070-y.

Lu:2015:NSC

[LSY15] Dawei Lu, Lixin Song, and Yang Yu. New sequences with continued fraction towards Euler’s constant. Applied Mathematics and Computation, 259(??):12–20, May 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300315001745.

Lefevre:2008:WCE

[LSZ08] Vincent Lefèvre, Damien Stehlé, and Paul Zimmermann. Worst cases for the exponential function in the IEEE 754r decimal64 format. Lecture Notes in Computer Science, 5045:114–126, 2008.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-540-85521-7_7.pdf.

Lucas:1971:AAC

[LT71] C. W. Lucas, Jr. and C. W. Terrill. ACM Algorithm 404: Complex gamma function [S14]. Communications of the ACM, 14 (1):48–49, January 1971. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lopez:2013:NSE

[LT13] José L. López and Nico M. Temme. New series expansions of the Gauss hypergeometric function. Advances in Computational Mathematics, 39(2):349–365, August 2013. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-9044 (electronic). URL http://

link.springer.com/article/10.1007/s10444-012-9283-y.

Lorch:1996:BPU

[LU96] Lee Lorch and Riccardo Uberti. “Best possible” upper bounds for the first positive zeros of Bessel functions — the finite part. Journal of Computational and Applied Mathematics, 75(2):249–258, November 28, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042796000659.

Lu:2014:SNI

[Lu14] Dawei Lu. Some new improved classes of convergence towards Euler’s constant. Applied Mathematics and Computation, 243 (??):24–32, September 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S009630031400798X.

Ludwig:1963:AIB

[Lud63] Oliver G. Ludwig. Algorithm 179: Incomplete beta ratio. Communications of the ACM, 6(6):314, June 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [PSB76, BB74].

Luke:1956:RFAa

[Luk56a] Yudell L. Luke. On rational function approximations to the exponential function with application to the practical solution of linear differential difference equations with constant coefficients.

In ????, editor, ACM’56: Proceedings of the 1956 11th ACM national meeting, pages 13–16. ACM Press, New York, NY 10036,

USA, 1956. URL https://dl.acm.org/ft_gateway.cfm?id=

808937. See errata and addenda [Luk56b].

Luke:1956:RFAb

[Luk56b] Yudell L. Luke. On rational function approximations to the exponential function with application to the practical solution of linear differential difference equations with constant coefficients: Errata and addenda. In ????, editor, ACM’56: Proceedings of the 1956 11th ACM national meeting, pages 177–178. ACM Press, New York, NY 10036, USA, 1956. URL https://

dl.acm.org/ft_gateway.cfm?id=808979. See [Luk56a].

Luke:1957:C

[Luk57] Yudell L. Luke. On the computation of log Z and arc tan Z. Mathematical Tables and Other Aids to Computation, 11(57):16– 18, January 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Luke:1959:ECH

[Luk59] Yudell L. Luke. Expansion of the confluent hypergeometric function in series of Bessel functions. Mathematical Tables and Other Aids to Computation, 13(68):261–271, October 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Luke:1964:IBF

[Luk64] Yudell L. Luke. Integrals of Bessel functions. In Abramowitz and Stegun [AS64], pages 479–494. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://

www.convertit.com/Go/ConvertIt/Reference/AMS55.ASP in

bitmap image format.

Luke:1968:AEI

[Luk68] Yudell L. Luke. Approximations for elliptic integrals. Mathematics of Computation, 22(103):627–634, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1969:SFTa

[Luk69a] Yudell L. Luke. The Special Functions and Their Approximations, volume I of Mathematics in Science and Engineering, Volume 53-I, Editor: Richard Bellman. Academic Press, New York, NY, USA, 1969. ISBN 0-12-459901-X. xx + 349 pp. LCCN QA351 .L94 1969.

Luke:1969:SFTb

[Luk69b] Yudell L. Luke. The Special Functions and Their Approximations, volume II of Mathematics in Science and Engineering, Volume 53-II, Editor: Richard Bellman. Academic Press, New York, NY, USA, 1969. ISBN 0-12-459902-8. xx + 485 pp.

LCCN QA351 .L797. URL http://www.sciencedirect.com/
science/book/9780124599024.

Luke:1970:FAE

[Luk70] Yudell L. Luke. Further approximations for elliptic integrals. Mathematics of Computation, 24(109):191–198, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1971:MTBa

[Luk71a] Yudell L. Luke. Miniaturized tables of Bessel functions. Mathematics of Computation, 25(114):323–330, April 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1971:MTBb

[Luk71b] Yudell L. Luke. Miniaturized tables of Bessel functions, II. Mathematics of Computation, 25(116):789–795, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1972:MTB

[Luk72] Yudell L. Luke. Miniaturized tables of Bessel functions. III. Mathematics of Computation, 26(117):237–240, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Luke:1975:SRU

[Luk75a] Y. L. Luke. Some remarks on uniform asymptotic expansions for Bessel functions. Computers and Mathematics with Applications, 1(3–4):285–290, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0898122175900279.

Luke:1975:MFT

[Luk75b] Yudell L. Luke. Mathematical Functions and Their Approximations. Academic Press, New York, NY, USA, 1975. ISBN 0-12-459950-8, 1-4832-6245-6 (e-book). xvii + 568 pp. LCCN QA55 .L96 1975. URL https://shop.

elsevier.com/books/mathematical-functions-and-their-
approximations/luke/978-0-12-459950-5.

Luke:1976:CER

[Luk76] Yudell L. Luke. Chebyshev expansions and rational approximations. Journal of Computational and Applied Mathematics, 2(2): 85–93, June 1976. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0771050X76900139.

Luke:1977:ACM

[Luk77] Yudell L. Luke. Algorithms for the Computation of Mathematical Functions. Academic Press, New York, NY, USA, 1977. ISBN 0-12-459940-0. xiii + 284 pp. LCCN QA351 .L7961.

Luk:1999:PSA

[Luk99] Franklin T. Luk, editor. Proceedings of SPIE: Advanced signal processing algorithms, architectures, and implementations IX: 19–21 July, 1999, Denver, Colorado, volume 3807. Society of Photo-optical Instrumentation Engineers (SPIE), Bellingham, WA, USA, 1999. ISBN 0-8194-3293-8. LCCN TK5102.5 .A3325 1999; TK5102.5 .A3173 1999eb; TK5102.9 .A37 1999; TK5102.5; TS510 .S63.

Luther:1995:HAT

[Lut95] Wolfram Luther. Highly accurate tables for elementary functions. BIT Numerical Mathematics, 35(3):352–360, September 1995. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print),

1572-9125 (electronic). URL http://www.mai.liu.se/BIT/
contents/bit35.html;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0006-3835&volume=35&issue=3&spage=
352.

Ledoux:2005:CME

[LVV05] V. Ledoux, M. Van Daele, and G. Vanden Berghe. CP methods and the evaluation of negative energy Coulomb Whittaker functions. Journal of Computational and Applied Mathematics, 183(1):168–176, November 1, 2005. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042705000233.

Lee-Whiting:1963:EFC

[LW63a] G. E. Lee-Whiting. Erratum: “Formulas for computing incomplete elliptic integrals of the first and second kinds”. Journal of the Association for Computing Machinery, 10:412, 1963. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See [LW63b].

Lee-Whiting:1963:FCI

[LW63b] G. E. Lee-Whiting. Formulas for computing incomplete elliptic integrals of the first and second kinds. Journal of the Association for Computing Machinery, 10(2):126–130, April 1963. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). See also [LW63a].

Ling:1982:EIH

[LW82] Chih Bing Ling and Ming Jing Wu. Evaluation of integrals of Howland type involving a Bessel function. Mathematics of Computation, 38(157):215–222, January 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lo:1989:RBP

[LW89] Shaw-Hwa Lo and Jane-Ling Wang. Representations for the bivariate product limit estimators and the bootstrap versions. Journal of Multivariate Analysis, 28(2):211–226, February 1989. CODEN JMVAAI. ISSN 0047-259x (print), 1095-7243 (electronic).

Lether:1995:MAZ

[LW95] F. G. Lether and P. R. Wenston. Minimax approximations to the zeros of P_{n}(x) and Gauss–Legendre quadrature. Journal of Computational and Applied Mathematics, 59(2):245–252, May 19, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042794000305.

Lang:1996:BPU

[LW96] T. Lang and R. Wong. “Best possible” upper bounds for the first two positive zeros of the Bessel function J_{v}(x): the infinite case. Journal of Computational and Applied Mathematics, 71(2):311– 329, July 27, 1996. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0377042795002200.

Lu:2014:NAE

[LW14] Dawei Lu and Xiaoguang Wang. A new asymptotic expansion and some inequalities for the gamma function. Journal of Number Theory, 140(??):314–323, July 2014. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0022314X14000705.

Lyons:1991:FMF

[Lyo91] Ken Lyons. A fast method for finding an integer square root. In Koopman, Jr. [Koo91], pages 27–30. ISBN 0-89791-462-7. LCCN

QA 76.73 F24 S53 1991. URL http://www.acm.org:80/pubs/
citations/proceedings/plan/259965/p27-lyons/. ACM or-

der number 817911.

MacLaren:1965:APN

[Mac65] M. D. MacLaren. Algorithm 272: Procedure for the normal distribution functions [S15]. Communications of the ACM, 8(12): 789–790, December 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remarks [HJ67b, Mac68].

MacLaren:1968:RAP

[Mac68] M. D. MacLaren. Remark on Algorithm 272: Procedure for the normal distribution functions. Communications of the ACM, 11 (7):498, July 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Mac65].

MacKinnon:1972:AEH

[Mac72] Robert F. MacKinnon. The asymptotic expansions of Hankel transforms and related integrals. Mathematics of Computation, 26(118):515–527, April 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Macleod:1989:SAA

[Mac89] Allan J. Macleod. Statistical algorithms: Algorithm AS 245: a robust and reliable algorithm for the logarithm of the gamma function. Applied Statistics, 38(2):397–402, June 1989. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic). URL

http://lib.stat.cmu.edu/apstat/245.

MacLeod:1994:CIA

[Mac94a] Allan J. MacLeod. Computation of inhomogeneous Airy functions. Journal of Computational and Applied Mathematics, 53(1):109–116, July 29, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0377042794901961.

MacLeod:1994:TBT

[Mac94b] Allan J. MacLeod. Table-based tests for Bessel function software. Advances in Computational Mathematics, 2(2):251–260, March 1994. CODEN ACMHEX. ISSN 1019-7168 (print), 1572-

9044 (electronic). URL http://link.springer.com/article/
10.1007/BF02521111.

Macleod:1996:AMS

[Mac96] Allan J. Macleod. Algorithm 757: MISCFUN, a software package to compute uncommon special functions. ACM Transactions on Mathematical Software, 22(3):288–301, September 1996. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://doi.acm.org/10.1145/232826.232846;
 http://www.acm.org/pubs/citations/journals/toms/1996-
22-3/p288-macleod/.

MacLeod:1997:AEE

[Mac97] Allan J. MacLeod. Accurate and efficient evaluation of the Bose–Einstein functions g_{3/2} and g_{5/2}. Computers in Physics, 11(4):385–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168609.

Maehly:1960:ACD

[Mae60] Hans J. Maehly. Approximations for the Control Data 1604. Institute for Advanced Study, Princeton, NJ, USA, January 15,

1960. ii + 44 pp. URL http://www.bitsavers.org/pdf/cdc/
1604/Approximations_For_The_Control_Data_1604_Mar60.
pdf.

Magnus:1994:ASA

[Mag94] Alphonse P. Magnus. Asymptotics and super asymptotics of best rational approximation error norms for the exponential function (the ‘1/9’ problem) by the Carathéodory–Fejér method. In Cuyt [Cuy94], pages 173–185. ISBN 0-7923-2967-8. LCCN QA297 .N642 1994. Proceedings of an international conference held at the University of Antwerp, Belgium, Sept. 5–11, 1993.

Mahler:1930:NUG

[Mah30] K. Mahler. Über die Nullstellen der unvollstaendigen Gammafunktionen. (German) [On the zeros of the incomplete gammafunctions]. Rendiconti del Circolo matematico di Palermo, 54(??):1–41, ???? 1930. CODEN RCMMAR. ISSN 0009-725X (print), 1973-4409 (electronic). URL https://carma.

newcastle.edu.au/mahler/collected.html;   https://carma.
newcastle.edu.au/mahler/docs/002.pdf.

Majithia:1972:CAE

[Maj72] J. C. Majithia. Cellular array for extraction of squares and square roots of binary numbers. IEEE Transactions on Computers, C-21(9):1023–1024, September 1972. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5009084.

Majerski:1985:SRA

[Maj85] S. Majerski. Square-rooting algorithms for high-speed digital circuits. IEEE Transactions on Computers, C-34(8):724–733, August 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-

9956 (electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1676618.

Maklovic:1965:IIC

[Mak65] S. T. Maklovič. Investigation of integrals containing Bessel and elementary functions. (Russian). Kišinev. Gos. Univ. Učen. Zap., 82:75–81, 1965.

Maksymiv:1980:APT

[Mak80] E. M. Maksymiv. Approximate properties of Thiele’s formula in a class of elementary functions. (Russian). In Differentsialnye Uravneniya i ikh Prilozhen, volume 141, pages 55–56. Vestnik L’vov. Politekhn. Inst., L’vov, USSR, 1980.

Mansfield:2023:MSR

[Man23] Daniel F. Mansfield. Mesopotamian square root approximation by a sequence of rectangles. British Journal for the History of Mathematics, 38(3):175–188, 2023. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/26375451.2023.
2215652.

Markman:1965:RZF

[Mar65] B. Markman. The Riemann zeta function. Nordisk tidskrift for informationsbehandling, 5(2):138–141, 1965.

Marino:1972:NAA

[Mar72] D. Marino. New algorithms for the approximate evaluation in hardware of binary logarithms and elementary functions. IEEE Transactions on Computers, C-21(12):1416–1421, December 1972. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Marsaglia:1977:SMG

[Mar77] George Marsaglia. The squeeze method for generating gamma variates. Computers and Mathematics with Applications, 3(4): 321–325, 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Markov:1981:ICE

[Mar81] S. M. Markov. On the interval computation of elementary functions. Comptes rendus de l’Académie bulgare des sciences, 34 (3):319–322, 1981. CODEN DBANAD. ISSN 0366-8681.

Marron:1985:FEP

[Mar85] J. S. Marron, editor. Function estimates: proceedings of a conference held July 28–August 3, 1985, volume 59 of Contemporary mathematics (American Mathematical Society). American Mathematical Society, Providence, RI, USA, 1985. ISBN 0-8218-5062-8. ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN QA276.8 .C651 1985. Held at Humboldt State University, Arcata, California.

Marsaglia:1986:CIG

[Mar86] John C. W. Marsaglia. C249. The incomplete gamma function and Ramanujan’s rational approximation to e^{x}. Journal of Statistical Computation and Simulation, 24(2):163–168, 1986. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Marsaglia:1988:CES

[Mar88] John Christopher Winston Marsaglia. Computer Evaluation of the special functions of probability and statistics. Ph.D. dissertation, Department of Computer Science, Washington State University, Pullman, WA, USA, August 1988. vii + 79 pp.

Markstein:1990:CEF

[Mar90] P. W. Markstein. Computation of elementary functions on the IBM RISC System/6000 processor. IBM Journal of Research and Development, 34(1):111–119, January 1990. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Maroni:1995:IRB

[Mar95] P. Maroni. An integral representation for the Bessel form. Journal of Computational and Applied Mathematics, 57(1– 2):251–260, February 20, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0377042793E0249L.

Markstein:2000:IEF

[Mar00] Peter Markstein. IA-64 and elementary functions: speed and precision. Hewlett–Packard professional books. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-018348-2. xix + 298 pp. LCCN QA76.9.A73 M365 2000. URL http://

www.markstein.org/.

Markstein:2003:FQP

[Mar03] Peter Markstein. A fast quad precision elementary function library for Itanium. In Anonymous [Ano03], pages 5–12. ISBN ???? LCCN ????

Markstein:2004:SDS

[Mar04a] Peter Markstein. Software division and square root using Goldschmidt’s algorithms. In Frougny et al. [FBM04], pages 146–157.

ISSN 0944-0488. URL http://www.informatik.uni-trier.
de/Reports/TR-08-2004/rnc6_12_markstein.pdf. Forschungs-

bericht Nr. 04-8.

Marsaglia:2004:END

[Mar04b] George Marsaglia. Evaluating the normal distribution. Journal of Statistical Software, 11(4):1–7, ???? 2004. CODEN JSSOBK.

ISSN 1548-7660. URL http://www.jstatsoft.org/counter.
php?id=100&url=v11/i04/cphi.pdf&ct=1.

Markstein:2005:FSM

[Mar05] Peter Markstein. A fast-start method for computing the inverse tangent. In IEEE [IEE05], page ?? ISBN ???? LCCN ????

URL http://arith17.polito.it/final/paper-112.pdf.

Masser:1975:EFT

[Mas75] D. W. Masser. Elliptic Functions and Transcendence, volume 437 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNMAA2. ISBN 3-540-07136-9 (print), 3-540-37410-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 112 (est.) pp. LCCN QA3 .L28 no. 437. URL http://

link.springer.com/book/10.1007/BFb0069432;   http://
www.springerlink.com/content/978-3-540-37410-7.

Mason:1983:CBF

[Mas83] Janet P. Mason. Cylindrical Bessel functions for a large range of complex arguments. Computer Physics Communications, 30(1):1–11, July/August 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583901169.

Mathis:1987:EFP

[Mat87] Robert F. Mathis. Elementary functions package for Ada. In ACM [ACM87], pages 95–100. ISBN 0-89791-243-8. LCCN QA 76.73 A35 U85 1987.

Matos:1990:CAM

[Mat90a] Ana C. Matos. A convergence acceleration method based on a good estimation of the absolute value of the error. IMA Journal of Numerical Analysis, 10(2):243–251, 1990. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Matula:1990:HPD

[Mat90b] D. Matula. Highly parallel divide and square root algorithms for a new generation floating point processor. In Ullrich [Ull90], page ?? ISBN ???? LCCN QA76.9.C62 C664 1990.

Matos:1992:CAP

[Mat92] Ana C. Matos. Convergence and acceleration properties for the vector ϵ-algorithm. Numerical Algorithms, 3(1–4):313–319, December 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). Extrapolation and rational approximation (Puerto de la Cruz, 1992).

Mathai:1993:HGS

[Mat93] A. M. Mathai. A handbook of generalized special functions for statistical and physical sciences. Oxford science publications. Clarendon Press, Oxford, UK, 1993. ISBN 0-19-853595-3. xi + 235 pp. LCCN QA351 .M35 1993. URL http://www.loc.gov/

catdir/enhancements/fy0635/92036065-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0635/92036065-t.html.

Mathar:2004:NRI

[Mat04] Richard J. Mathar. Numerical representations of the incomplete gamma function of complex-valued argument. Numerical Algorithms, 36(3):247–264, July 2004. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

ipsapp009.kluweronline.com/IPS/content/ext/x/J/5058/
I/57/A/4/abstract.htm.

Maximon:1991:EIP

[Max91] Leonard C. Maximon. On the evaluation of the integral over the product of two spherical Bessel functions. Journal of Mathematical Physics, 32(3):642–648, March 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v32/i3/p642_s1.

Mason:1987:AAB

[MC87] J. C. Mason and M. G. Cox, editors. Algorithms for approximation: based on the proceedings of the IMA Conference on Algorithms for the Approximation of Functions and Data, held at the Royal Military College of Science, Shrivenham, July 1985, volume 10 of The Institute of Mathematics and Its Applications conference series, new series. Clarendon Press, Oxford, UK, 1987. ISBN 0-19-853612-7. LCCN QA221 .A5361 1987; QA221 .I47 1985. US$90.

Montuschi:1989:EIH

[MC89] Paolo Montuschi and Luigi Cinimera. On the efficient implementation of higher radix square root algorithms. In Ercegovac and Swartzlander, Jr. [ES89], pages 154–161. ISBN 0-8186-8963-3 (case), 0-8186-5963-7 (microfiche). LCCN QA 76.9 C62 S95

1989. URL http://www.acsel-lab.com/arithmetic/arith9/
papers/ARITH9_Montuschi.pdf. IEEE catalog no. 89CH2757-

3.

Mason:1990:AAI

[MC90] J. C. Mason and M. G. Cox, editors. Algorithms for approximation II: based on the proceedings of the Second International Conference on Algorithms for Approximation, held at Royal Military College of Science, Shrivenham, July 1988. Chapman and Hall, Ltd., London, UK, 1990. ISBN 0-412-34580-3. LCCN QA221 .I54 1988.

Montuschi:1991:SRD

[MC91] Paolo Montuschi and Luigi Ciminiera. Simple radix 2 division and square root with skipping of some addition steps. In Kornerup and Matula [KM91], pages 202–209. ISBN 0-8186-9151-4 (case), 0-8186-6151-8 (microfiche), 0-7803-0187-0 (library binding). LCCN QA76.9.C62 S95 1991. URL http://www.acsel-

lab.com/arithmetic/arith10/papers/ARITH10_Montuschi.
pdf.

Montuschi:1993:RIT

[MC93] P. Montuschi and L. Ciminiera. Reducing iteration time when result digit is zero for radix 2 SRT division and square root with redundant remainders. IEEE Transactions on Computers, 42 (2):239–246, February 1993. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See remark [MC95].

Montuschi:1995:RRI

[MC95] P. Montuschi and L. Ciminiera. A remark on “Reducing iteration time when result digit is zero for radix-2 SRT division and square root with redundant remainders”. IEEE Transactions on Computers, 44(1):144–146, January 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). See [MC93].

Moore:2004:PSW

[MC04] Ian C. Moore and Michael Cada. Prolate spheroidal wave functions, an introduction to the Slepian series and its properties. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 16(3):208–230, May 2004. ISSN 1063-5203 (print), 1096-603x (electronic).

McCabe:1974:CFE

[McC74] J. H. McCabe. A continued fraction expansion, with a truncation error estimate, for Dawson’s integral. Mathematics of Computation, 28(127):811–816, July 1974. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCarthy:1977:OAE

[McC77] D. P. McCarthy. The optimal algorithm to evaluate x^{n} using elementary multiplication methods. Mathematics of Computation, 31(137):251–256, January 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McCullagh:1981:RCS

[McC81] Peter McCullagh. A rapidly convergent series for computing ψ(z) and its derivatives. Mathematics of Computation, 36(153): 247–248, January 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.jstor.org/

stable/2007741.

McCabe:1983:ASC

[McC83] J. H. McCabe. On an asymptotic series and corresponding continued fraction for a gamma function ratio. Journal of Computational and Applied Mathematics, 9(2):125–130, June 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042783900353.

McCurley:1984:EE

[McC84] Kevin S. McCurley. Explicit estimates for θ(x; 3, l) and ψ(x; 3, l). Mathematics of Computation, 42(165):287–296, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

McConnell:1990:LEP

[McC90] C. R. McConnell. Letter to the Editor: Pocket computer approximation for areas under the standard normal curve. The American Statistician, 44(1):63, February 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL

http://www.jstor.org/stable/2684963.

McCluskey:2002:MLF

[McC02] Glen McCluskey. Math library functions in C9X. ;login: the USENIX Association newsletter, 27(2):8–13, April 2002. CO-DEN LOGNEM. ISSN 1044-6397. URL http://www.usenix.

org/publications/login/2002-04/pdfs/mccluskey.pdf.

McLachlan:1934:BFE

[McL34] N. W. (Norman William) McLachlan. Bessel Functions for Engineers. The Oxford engineering science series. Clarendon Press, Oxford, UK, 1934. xi + 1 + 192 pp. LCCN QA408 .M3.

McLachlan:1955:BFE

[McL55] N. W. (Norman William) McLachlan. Bessel Functions for Engineers. The Oxford engineering science series. Clarendon Press, Oxford, UK, second edition, 1955. xii + 239 pp. LCCN QA408 .M3 1955.

Mora:1990:EFI

[MCM90] Gerardo Mora, Edwin Castro, and Ioan Muntean. Elementary functions. I. (Spanish). In Mathematics in Costa Rica, Vol. 1 (Spanish) (San José, 1990), pages 76–86. Univ. Costa Rica, San José, Costa Rica, 1990.

McNolty:1973:SPD

[McN73] Frank McNolty. Some probability density functions and their characteristic functions. Mathematics of Computation, 27(123): 495–504, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Munoz-Coreas:2018:CQO

[MCT18] Edgard Muñoz-Coreas and Himanshu Thapliyal. T-count and qubit optimized quantum circuit design of the non-restoring square root algorithm. ACM Journal on Emerging Technologies in Computing Systems (JETC), 14(3):36:1–36:15, October 2018. CODEN ???? ISSN 1550-4832.

Martin-Dorel:2016:PTB

[MDM16] Érik Martin-Dorel and Guillaume Melquiond. Proving tight bounds on univariate expressions with elementary functions in Coq. Journal of Automated Reasoning, 57(3):187–217, October 2016. CODEN JAREEW. ISSN 0168-7433 (print), 1573-0670

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10817-015-9350-4.

Mechel:1966:CMB

[Mec66] Fr. Mechel. Calculation of the modified Bessel functions of the second kind with complex argument. Mathematics of Computation, 20(95):407–412, July 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Mechel:1968:IRT

[Mec68] Fr. Mechel. Improvement in recurrence techniques for the computation of Bessel functions of integral order (in Technical Notes and Short Papers). Mathematics of Computation, 22(101):202– 205, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Megiddo:1989:POS

[Meg89a] N. Megiddo. Pathways to the optimal set in linear programming. In Progress in Mathematical Programming: Interior-Point and Related Methods [Meg89b], pages 131–158. ISBN 0-387-96847-4. LCCN QA402.5 .P7851 1989. Proceedings of the conference held at the Asilomar conference center in Pacific Grove, California, March 1–4, 1987.

Megiddo:1989:PMP

[Meg89b] N. Megiddo, editor. Progress in Mathematical Programming: Interior-Point and Related Methods. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-96847-4. LCCN QA402.5 .P7851 1989. Proceedings of the conference held at the Asilomar conference center in Pacific Grove, California, March 1–4, 1987.

Meinardus:1964:AFI

[Mei64] Gunter Meinardus. Approximation von Funktionen und ihre numerische Behandlung. (German) [Approximation of functions and their numerical treatment], volume 4 of Springer tracts in natural philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1964. ISBN 3-540-03219-3, 3-642-85646-2, 3-642-85647-0 (print). viii + 180 pp. LCCN QA320 .M4.

Meinardus:1967:AFT

[Mei67] Günter Meinardus. Approximation of functions: Theory and numerical methods, volume 13 of Springer Tracts in Natural Philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 0-387-03985-6, 3-540-03985-6. ISSN 0081-3877. viii + 198 pp. LCCN QA221 .M3813.

URL http://catalogue.bnf.fr/ark:/12148/cb37378349d.

Meister:1983:MYF

[Mei83] B. Meister. On Murphy’s yield formula. IBM Journal of Research and Development, 27(6):545–548, November 1983. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Menezes:2006:PAS

[Men06] Ronaldo Menezes, editor. Proceedings of the 44th annual Southeast Regional Conference 2006: Melbourne, Florida, March 10– 12, 2006. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-315-8 (print). LCCN QA75.5 A184 2006 E.

Merner:1962:AAC

[Mer62] J. N. Merner. ACM Algorithm 149: Complete elliptic integral. Communications of the ACM, 5(12):605, December 1962. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Sko78].

Merrheim:1994:CEF

[Mer94] X. Merrheim. The computation of elementary functions in radix 2^{p}. Computing: Archiv fur informatik und numerik, 53(3–4): 219–232, 1994. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). International Symposium on Scientific Computing, Computer Arithmetic and Validated Numerics (Vienna, 1993).

Merkle:2005:GRG

[Mer05] M. Merkle. Gurland’s ratio for the gamma function. Computers and Mathematics with Applications, 49(2–3):389–406, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122105000416.

Metze:1965:MSR

[Met65] Gernot Metze. Minimal square rooting. IEEE Transactions on Electronic Computers, EC-14(2):181–185, April 1965. CODEN IEECA8. ISSN 0367-7508. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=4038397.

Meyer:1963:CAI

[Mey63] Noelle A. Meyer. Certification of Algorithm 73: Incomplete elliptic integrals. Communications of the ACM, 6(2):69, February 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meyer:2001:JEF

[Mey01] Kenneth R. Meyer. Jacobi elliptic functions from a dynamical systems point of view. American Mathematical Monthly, 108(8): 729–737, October 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2695616.

Martin:1985:FAB

[MG85] Pablo Martín and Antonio L. Guerrero. Fractional approximations to the Bessel function J_{0}(x). Journal of Mathematical Physics, 26(4):705–707, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i4/p705_s1.

Martin:1989:TPQ

[MG89] Pablo Martin and Antonio Luis Guerrero. Two-point quasifractional approximations to the Bessel function J_{ν}(x) of fractional order. Journal of Computational Physics, 85(2):487–492, December 1989. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999189901617.

Morita:1972:CAG

[MH72] Tohru Morita and Tsuyoshi Horiguchi. Convergence of the arithmetic-geometric mean procedure for the complex variables and the calculation of the complete elliptic integrals with complex modulus. Numerische Mathematik, 20(5):425–430, October 1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0029-599X&volume=
20&issue=5&spage=425.

Mason:2003:CP

[MH03] J. C. Mason and D. C. Handscomb. Chebyshev Polynomials. Chapman and Hall/CRC, Boca Raton, FL, USA, 2003. ISBN 0-8493-0355-9. xiii + 341 pp. LCCN QA404.5 .M37 2003.

Mathai:2008:SFA

[MH08] A. M. Mathai and H. J. Haubold. Special Functions for Applied Scientists. Springer Science+Business Media, New York, NY, USA, 2008. ISBN 0-387-75893-3. xxv + 464 pp. LCCN QA351 .M37X 2008.

Hwu:2012:GCG

[mH12] Wen mei Hwu, editor. GPU computing gems. Applications of GPU computing series. Morgan Kaufmann, Boston, MA, jade edition, 2012. ISBN 0-12-385963-8 (hardback). xvi + 541 + 16 pp. LCCN T385 .G6875 2012.

Munshi:2016:OCS

[MHSK16] Aaftab Munshi, Lee Howes, Bartosz Sochacki, and Khronos OpenCL Working Group. The OpenCL C specification version: 2.0 document revision: 33. Web document., April

13, 2016. URL https://www.khronos.org/registry/OpenCL/
specs/opencl-2.0-openclc.pdf.

Matsubara:1997:LPZ

[MI97] G. Matsubara and N. Ide. A low power zero-overhead self-timed division and square root unit combining a single-rail static circuit with a dual-rail dynamic circuit. In Proceedings of the Third International Symposium on Advanced Research in Asynchronous Circuits and Systems, 7–10 April 1997, pages 198–209. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. CODEN ???? ISSN ????

Midy:1975:ICG

[Mid75] P. Midy. An improved calculation of the general elliptic integral of the second kind in the neighbourhood of x = 0. Numerische Mathematik, 25(1):99–101, March 1975. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miller:1957:NGS

[Mil57] J. C. P. Miller. Note on the general solution of the confluent hypergeometric equation. Mathematical Tables and Other Aids to Computation, 11(58):97–99, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic). URL http://

www.jstor.org/stable/2002156.

Miller:1958:LNN

[Mil58] J. C. P. Miller. Lecture notes on numerical analysis. Report, Cambridge University, Cambridge, England, 1958. URL https:

//www.youtube.com/watch?v=81YrRSeReKo.

Miller:1964:PCF

[Mil64] J. C. P. Miller. Parabolic cylinder functions. In Abramowitz and Stegun [AS64], pages 685–720. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://

www.convertit.com/Go/ConvertIt/Reference/AMS55.ASP in

bitmap image format.

Miller:1965:ASF

[Mil65] G. F. Miller. Algorithms for special functions II. Numerische Mathematik, 7:194–196, 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Miller:1968:LTS

[Mil68] Willard Miller. Lie theory and special functions, volume 43 of Mathematics in science and engineering. Academic Press, New York, NY, USA, 1968. ISBN 0-12-497450-3. xv + 338 pp. LCCN QA387 .M55 1968eb.

Miller:1975:TNA

[Mil75a] John J. H. Miller, editor. Topics in numerical analysis: proceedings of the Royal Irish Academy Conference on Numerical Analysis, 1972, 1974, 1976. Academic Press, New York, NY, USA, 1975. ISBN 0-12-496950-X. LCCN QA297 .R69 1973.

Miller:1975:CCN

[Mil75b] Webb Miller. Computational complexity and numerical stability. SIAM Journal on Computing, 4(2):97–107, June 1975. CODEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Milgram:1985:GIE

[Mil85] M. S. Milgram. The generalized integro-exponential function. Mathematics of Computation, 44(170):443–458, April 1985. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miler:1989:EEM

[Mil89] T. H. Miler. Error evaluation of microcomputer intrinsic functions. Ph.D thesis, University of Idaho, Moscow, ID, 1989. ???? pp.

Miller:2004:AMF

[Mil04] Alan Miller. Alan Miller’s Fortran software. Web site, February

4, 2004. URL https://jblevins.org/mirror/amiller/. From

the Web site: All code written by Alan Miller is released into the public domain.

Mimachi:1988:PRI

[Mim88] Katsuhisa Mimachi. A proof of Ramanujan’s identity by use of loop integrals. SIAM journal on mathematical analysis, 19(6): 1490–1493, November 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Minton:1970:GHF

[Min70] Barry M. Minton. Generalized hypergeometric function of unit argument. Journal of Mathematical Physics, 11(4):1375– 1376, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i4/p1375_s1.

Mitchell:1949:TFA

[Mit49] K. Mitchell. Tables of the function ∫_{0}^{z} − log |1 − y|/y dy with an account of some properties of this and related functions. Philosophical Magazine, 40(302):351–368, 1949. CODEN PHMAA4.

ISSN 0031-8086. URL https://www.tandfonline.com/doi/
abs/10.1080/14786444908561256.

Mitchell:1992:VFA

[Mit92] H. B. Mitchell. Very fast accurate square-root algorithm for use with gradient edge operators. Electronics Letters, 28(10):922– 923, May 7, 1992. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Matsubara:1995:NBS

[MIT^{+}95] G. Matsubara, N. Ide, H. Tago, S. Suzuki, and N. Goto. 30ns 55-b shared radix 2 division and square root using a selftimed circuit. In Knowles and McAllister [KM95], pages 98–105. CODEN ???? ISBN 0-8186-7089-4 (paperback), 0-8186-7089-4 (case), 0-8186-7149-1 (microfiche), 0-8186-7089-4 (softbound), 0-7803-2949-X (casebound). ISSN ???? LCCN QA 76.9 C62 S95 1995.

Miyajima:2019:VCM

[Miy19] Shinya Miyajima. Verified computation for the matrix Lambert W function. Applied Mathematics and Computation, 362(??):Article 124555, December 1, 2019. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0096300319305387.

Mastin:2013:LQB

[MJ13] Andrew Mastin and Patrick Jaillet. Log-quadratic bounds for the Gaussian Q-function. arxiv.org, ??(??):??, April 9, 2013.

URL http://arxiv.org/abs/1304.2488.

Majithia:1971:CAN

[MK71] J. C. Majithia and R. Kitai. A cellular array for the nonrestoring extraction of square roots. IEEE Transactions on Computers, C-20(12):1617–1618, December 1971. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1671784.

Mlodzki:1983:PPC

[MKS83] J. Mlodzki, J. Kuszkowski, and M. Suffczynski. A Pascal program for calculating the reduced Coulomb Green’s functions and their partial waves. Computer Physics Communications, 29(4): 341–350, June 1983. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465583900139.

Mikami:1992:NDO

[MKY92] N. Mikami, M. Kobayashi, and Y. Yokoyama. A new DSPoriented algorithm for calculation of the square root using a nonlinear digital filter. IEEE Transactions on Signal Processing, 40(7):1663–1669, July 1992. CODEN ITPRED. ISSN 1053-587X (print), 1941-0476 (electronic).

Mikaitis:2018:AFP

[MLS^{+}18] Mantas Mikaitis, David R. Lester, Delong Shang, Steve Furber, Gengting Liu, Jim Garside, Stefan Scholze, Sebastian Höppner, and Andreas Dixius. Approximate fixed-point elementary function accelerator for the SpiNNaker-2 neuromorphic chip. In Tenca and Takagi [TT18], pages 37–44. ISBN 1-5386-2612-8 (USB), 1-5386-2665-9. ISSN 2576-2265. LCCN QA76.9.C62. IEEE catalog number CFP18121-USB.

Moler:1983:RSR

[MM83] Cleve B. Moler and Donald Morrison. Replacing square roots by Pythagorean sums. IBM Journal of Research and Development, 27(6):577–581, November 1983. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http:

//www.research.ibm.com/journal/rd/276/ibmrd2706P.pdf.

See [Dub83] and generalization [Jam89].

Milone:1988:EDF

[MM88] L. A. Milone and A. A. E. Milone. Evaluation of Dawson’s function. Astrophysics and Space Science, 147(1):189–191, 1988. CO-DEN APSSBE. ISSN 0004-640X (print), 1572-946X (electronic).

Magnus:1990:CFL

[MM90a] Arne Magnus and John McCabe. On a continued fraction for log_{e}^{2}(1 + x). Journal of Computational and Applied Mathematics, 30(1):81–86, April 10, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037704279090007M.

Montuschi:1990:SSR

[MM90b] P. Montuschi and P. M. Mezzalama. Survey of square rooting algorithms. IEE Proceedings. Computers and Digital Techniques, 137(1):31–40, January 1990. CODEN ICDTEA. ISSN 1350-2387 (print), 1359-7027 (electronic).

McQuillan:1991:VAM

[MM91a] S. E. McQuillan and J. V. McCanny. A VLSI architecture for multiplication, division and square root. In 1991 International Conference on Acoustics, Speech, and Signal Processing: ICASSP-91, 14–17 April 1991, volume 2, pages 1205–1208. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. CODEN ???? ISSN ????

Montuschi:1991:OAE

[MM91b] P. Montuschi and M. Mezzalama. Optimal absolute error starting values for Newton–Raphson calculation of square root. Computing: Archiv fur informatik und numerik, 46(1):67–86, March 1991. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

McQuillan:1992:VMH

[MM92] S. E. McQuillan and J. V. McCanny. VLSI module for highperformance multiply, square root and divide. IEE Proceedings. Computers and Digital Techniques, 139(6):505–510, November 1992. CODEN ICDTEA. ISSN 1350-2387 (print), 1359-7027 (electronic).

Miller:1995:RCF

[MM95] A. R. Miller and I. S. Moskowitz. Reduction of a class of Fox– Wright psi functions for certain rational parameters. Computers and Mathematics with Applications, 30(11):73–82, December 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/089812219500165U.

Miller:1998:CGI

[MM98] Allen R. Miller and Ira S. Moskowitz. On certain generalized incomplete gamma functions. Journal of Computational and Applied Mathematics, 91(2):179–190, May 4, 1998. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042798000314.

MacNair:1989:WSC

[MMH89] Edward A. MacNair, Kenneth J. Musselman, and Philip Heidelberger, editors. 1989 Winter Simulation Conference proceedings: December 4–6, 1989, the Capital Hilton Hotel, Washington, DC. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-911801-58-8. LCCN QA76.9.C65 W56 1989. URL http://ieeexplore.

ieee.org/servlet/opac?punumber=5823. IEEE order number

89CH2778-9.

McQuillan:1993:NAV

[MMH93] S. E. McQuillan, J. V. McCanny, and R. Hamill. New algorithms and VLSI architectures for SRT division and square root. In Swartzlander, Jr. et al. [SIJ93], pages 80–86. CODEN ???? ISBN 0-7803-1401-8 (softbound), 0-8186-3862-1 (casebound), 0-8186-3861-3 (microfiche). ISSN ???? LCCN QA 76.9

C62 S95 1993. URL http://www.acsel-lab.com/arithmetic/
arith11/papers/ARITH11_McQuillan.pdf. IEEE Transactions

on Computers 43(8), 1994.

Mazenc:1993:CFU

[MMM93] Christophe Mazenc, Xavier Merrheim, and Jean-Michel Muller. Computing functions cos^{−1} and sin^{−1} using Cordic. IEEE Transactions on Computers, 42(1):118–122, January 1993. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=192222.

Markstein:1991:WFF

[MMNP91] V. Markstein, P. Markstein, T. Nguyen, and S. Poole. Wide format floating-point math libraries. In IEEE [IEE91], pages 130–138. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Maino:1981:CPC

[MMV81] G. Maino, E. Menapace, and A. Ventura. Computation of parabolic cylinder functions by means of a Tricomi expansion. Journal of Computational Physics, 40(2):294–304, April 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181902114.

McQuillan:1991:HPV

[MMW91] S. E. McQuillan, J. V. McCanny, and R. F. Woods. High performance VLSI architecture for division and square root. Electronics Letters, 27(1):19–21, January 3, 1991. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Mohankumar:2016:VAN

[MN16] N. Mohankumar and A. Natarajan. On the very accurate numerical evaluation of the Generalized Fermi–Dirac Integrals. Computer Physics Communications, 207(??):193–201, October 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465516301667.

McCurdy:1983:ACD

[MNP83] A. McCurdy, K. C. Ng, and Beresford N. Parlett. Accurate computation of divided differences of the exponential function. Report PAM-160, Center for Pure and Applied Mathematics, University of California, Berkeley, Berkeley, CA, USA, June 1983.

Magnus:1949:FTS

[MO49] Wilhelm Magnus and Fritz Oberhettinger. Formulas and theorems for the special functions of mathematical physics. Chelsea Pub. Co., New York, NY, USA, 1949. 172 pp. LCCN QA41 M19e. Translated from the German by John Wermer.

Myland:2018:JEF

[MO18] Jan C. Myland and Keith B. Oldham. Jacobian elliptic functions describe the properties of the diffuse charge region in narrow electrochemical cells. Journal of Mathematical Chemistry, 56 (4):1184–1205, April 2018. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (electronic). URL https://link.springer.

com/article/10.1007/s10910-017-0847-4.

Moler:1983:MSV

[Mol83] C. Moler. Mathematical software for vector computers. In ????, editor, SIAM Conference on Parallel Processing for Scientific Computing, Norfolk, VA, November 10–11, 1983, page ?? Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1983.

Monk:1983:SFE

[Mon83] P. B. Monk. Some finite element methods for the approximation of the biharmonic equation. Ph.D thesis, Rutgers University, The State University of New Jersey, New Brunswick, NJ, USA, 1983. 242 pp.

Moody:1967:ADF

[Moo67] William T. Moody. Approximations for the psi (digamma) function (in technical notes and short notices). Mathematics of Computation, 21(97):112, January 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Moon:1981:AFC

[Moo81] Wooil Moon. Airy function with complex arguments. Computer Physics Communications, 22(4):411–417, May 1981. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0010465581901387.

Moon:1984:AFC

[Moo84] Wooil Moon. Airy function with complex arguments. Computer Physics Communications, 35(1–3):C–692, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465584828428.

Morelock:1964:AAE

[Mor64] J. C. Morelock. ACM algorithm 229: Elementary functions by continued fractions. Communications of the ACM, 7(5):296, May 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morris:1975:LRS

[Mor75] Robert Morris. A library of reference standard mathematical subroutines. Technical Memorandum 1074 (TM 75-1271-6), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1, 1975. ?? pp.

Morris:1976:RDF

[Mor76] Robert Morris. Remark on “Algorithm 490: The dilogarithm function of a real argument [S22]”. ACM Transactions on Mathematical Software, 2(1):112, March 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [GZ75].

Morita:1978:CCE

[Mor78] T. Morita. Calculation of the complete elliptic integrals with complex modulus. Numerische Mathematik, 29(2):233–236, January 1978. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Morris:1979:DFR

[Mor79] Robert Morris. The dilogarithm function of a real argument. Mathematics of Computation, 33(146):778–787, April 1979. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Moran:1980:CND

[Mor80] P. A. P. Moran. Calculation of the normal distribution function. Biometrika, 67(3):675–676, December 1980. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://

www.jstor.org/stable/2335138.

Morris:1981:NDL

[Mor81] Alfred H. Morris, Jr. NSWC/DL library of mathematics subroutines. Report NSWC/TR-79-338, Naval Surface Warfare Center, Dahlgren, VA 22448-5000, USA; Silver Spring, MD 20903-5000,

USA, 1981. 235 pp. URL https://ntrl.ntis.gov/NTRL/
dashboard/searchResults/titleDetail/ADA108106.xhtml.

See also later editions [Mor90, Mor93].

Morris:1984:NLM

[Mor84] A. H. Morris. NSWC library of mathematics subroutines. Report NSWC TR 84-143, Naval Surface Weapons Center, Dahlgren, VA, USA, 1984.

Morris:1990:NLM

[Mor90] Alfred H. Morris, Jr. NSWC library of mathematics subroutines. Report NSWC TR 90-21, Naval Surface Warfare Center, Dahlgren, VA 22448-5000, USA; Silver Spring, MD 20903-5000, USA, January 1990. xii + 492 + 9 pp.

URL https://apps.dtic.mil/sti/citations/ADA476840;
https://apps.dtic.mil/sti/pdfs/ADA476840.pdf;   https:
//people.math.sc.edu/Burkardt/f_src/nswc/nswc.f90;
https://people.math.sc.edu/Burkardt/f_src/nswc/nswc.

html. See also later edition [Mor93].

Morris:1993:NLM

[Mor93] Alfred H. Morris, Jr. NSWC library of mathematics subroutines. Report NSWCDD/TR-92/425, Naval Surface Warfare Center, Dahlgren, VA 22448-5000, USA; Silver Spring, MD 20903-5000, USA, January 1993. 464 pp. URL https://ntrl.ntis.gov/

NTRL/dashboard/searchResults/titleDetail/ADA261511.xhtml.

See also earlier edition [Mor90].

Morita:1999:CEI

[Mor99] T. Morita. Calculation of the elliptic integrals of the first and second kinds with complex modulus. Numerische Mathematik, 82(4):677–688, June 1999. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://link.springer-

ny.com/link/service/journals/00211/bibs/9082004/90820677.
htm;   http://link.springer-ny.com/link/service/journals/
00211/papers/9082004/90820677.pdf.

Mortici:2011:IAF

[Mor11] Cristinel Mortici. Improved asymptotic formulas for the gamma function. Computers and Mathematics with Applications, 61(11):3364–3369, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0898122111003373.

Mortici:2014:SBG

[Mor14] Cristinel Mortici. Sharp bounds for gamma function in terms of x^{x−1}. Applied Mathematics and Computation, 249(??): 278–285, December 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300314013939.

Magnus:1966:FTS

[MOS66] Wilhelm Magnus, Fritz Oberhettinger, and Raj Pal Soni. Formulas and theorems for the special functions of mathematical physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1966. viii + 508 pp. LCCN QA1 G88 v. 52, 1966. See errata [Coh12, Szm13].

Moses:1969:ICS

[Mos69] Joel Moses. The integration of a class of special functions with the Risch algorithm. AI Memo (180) MAC-M-421, Artificial Intelligence Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1969. 13 + 1 pp.

Moses:1972:TGT

[Mos72] Joel Moses. Toward a general theory of special functions. Communications of the ACM, 15(7):550–554, July 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Twenty-fifth anniversary of the Association for Computing Machinery.

Moshier:1989:MPM

[Mos89] Stephen L. B. Moshier. Methods and Programs for Mathematical Functions. Ellis Horwood, New York, NY, USA, 1989. ISBN 0-7458-0289-3. vii + 415 pp. LCCN QA331 .M84 1989.

US£48.00. URL http://www.moshier.net/;   http://www.
netlib.org/cephes.

Motzkin:1955:EP

[Mot55a] T. S. Motzkin. Evaluation of polynomials. Bulletin of the American Mathematical Society, 61(2):163, March 1955. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Motzkin:1955:ERF

[Mot55b] T. S. Motzkin. Evaluation of rational functions. Bulletin of the American Mathematical Society, 61(2):163, March 1955. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Moursund:1967:OSV

[Mou67] David G. Moursund. Optimal starting values for Newton– Raphson calculation of \sqrt{x}. Communications of the ACM, 10 (7):430–432, July 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See letter [Fik67a].

Maurone:1979:ABD

[MP79] Philip A. Maurone and Alain J. Phares. On the asymptotic behavior of the derivatives of Airy functions. Journal of Mathematical Physics, 20(11):2191, November 1979. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v20/
i11/p2191_s1.

Martin:1992:TPQa

[MPG92a] Pablo Martin, Ricardo Pérez, and Antonio L. Guerrero. Twopoint quasi-fractional approximations to the Airy function Ai(x). Journal of Computational Physics, 98(2):349, February 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290165U.

Martin:1992:TPQb

[MPG92b] Pablo Martín, Ricardo Pérez, and Antonio L. Guerrero. Twopoint quasi-fractional approximations to the Airy function Ai(x). Journal of Computational Physics, 99(2):337–340, April 1992. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199919290212H.

Medhurst:1965:EI

[MR65] R. G. Medhurst and J. H. Roberts. Evaluation of the integral

I_{n}(b) = \frac{2}{π} ∫_{0}^{∞} (\frac{sin x}{x})^{n} cos(bx)dx. Mathematics of Computation,

19(89):113–117, April 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Matta:1971:UCE

[MR71] F. Matta and A. Reichel. Uniform computation of the error function and other related functions. Mathematics of Computation, 25(114):339–344, April 1971. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/journals/mcom/1971-25-
114/S0025-5718-1971-0295538-4.

Melquiond:2019:FVS

[MRH19] Guillaume Melquiond and Raphael Rieu-Helft. Formal verification of a state-of-the-art integer square root. In Takagi et al. [TBL19], pages 183–186. ISBN 1-72813-366-1. ISSN 1063-6889.

Matic:2017:PBA

[MRS17] Ivan Matić, Radoš Radoičić, and Dan Stefanica. Pólyabased approximation for the ATM-forward implied volatility. International Journal of Financial Engineering, 4(2–3):1–15, June/September 2017. ISSN 2424-7863. URL http://www.

worldscientific.com/doi/abs/10.1142/S2424786317500323.

Matic:2018:SPB

[MRS18] Ivan Matić, Radoš Radoičić, and Dan Stefanica. A sharp Pólyabased approximation to the normal cumulative distribution function. Applied Mathematics and Computation, 322(??):111– 122, April 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S009630031730718X.

MacRobert:1967:SHE

[MS67] Thomas Murray MacRobert and Ian Naismith Sneddon. Spherical harmonics; an elementary treatise on harmonic functions, with applications, volume 98 of International series of monographs in pure and applied mathematics. Pergamon Press, New York, NY, USA, third edition, 1967. xviii + 349 pp. LCCN QA406 .M3 1967.

Mavromatis:1968:IFP

[MS68] H. A. Mavromatis and K. Schilcher. Inverse functions of the products of two Bessel functions and applications to potential scattering. Journal of Mathematical Physics, 9(10):1627–1632, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i10/p1627_s1.

Mori:1982:ARS

[MS82] S. Mori and C. Y. Suen. Automatic recognition of symbols and architecture of the recognition unit. In Ching Y. Suen and Renato De Mori, editors, Computer analysis and perception: vol. I, Visual signals, pages 17–40. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1982. ISBN 0-8493-6305-5 (vol. 1), 0-8493-6306-3 (vol. 2). LCCN TA1650 .C65 1982.

Moricz:1987:ACF

[MS87] Ferenc Moricz and Xianliang Shi. Approximation to continuous functions by Cesaro means of double Fourier series and conjugate series. Journal of Approximation Theory, 49(4):346–377, April 1987. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Mudge:1994:PTS

[MS94] Trevor N. Mudge and Bruce D. Shriver, editors. Proceedings of the Twenty-Seventh Hawaii International Conference on System Sciences Vol. I: Architecture. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5050-8 (paper), 0-8186-5051-6 (microfiche). LCCN ???? First of five volumes. IEEE Catalog No. 94TH0607-2.

Meunier:2003:EAG

[MS03] Ludovic Meunier and Bruno Salvy. ESF: an automatically generated encyclopedia of special functions. In Hoon Hong, editor, Proceedings of the 2003 International Symposium on Symbolic and Algebraic Computation: Philadelphia, PA, USA, August 3–6, 2003. ACM Press, New York, NY 10036, USA, August 2003. ISBN 1-58113-641-2 (paperback). LCCN QA76.5 S98 2003; QA76.95.I59 2003.

Mortici:2015:PAG

[MS15] Cristinel Mortici and Hari M. Srivastava. A product approximation of the gamma function. Numerical Algorithms, 69 (3):595–610, July 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://link.springer.

com/article/10.1007/s11075-014-9915-z.

Maignan:2016:FGL

[MS16] Aude Maignan and Tony C. Scott. Fleshing out the generalized Lambert W function. ACM Communications in Computer Algebra, 50(2):45–60, June 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Montuschi:2016:ISC

[MSH^{+}16] Paolo Montuschi, Michael Schulte, Javier Hormigo, Stuart Oberman, and Nathalie Revol, editors. 2016 IEEE 23nd Symposium on Computer Arithmetic (ARITH 2016), Santa Clara, California, USA, 10–13 July 2016. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2016. ISBN 1-5090-1615-5. ISSN 1063-6889. LCCN QA76.9.C62

S95 2016. URL http://ieeexplore.ieee.org/servlet/opac?
punumber=7562813.

Mansour:1989:CAS

[MST89] Y. Mansour, B. Schieber, and P. Tiwari. The complexity of approximating the square root. In IEEE [IEE89b], pages 325–330. CODEN ???? ISBN 0-8186-1982-1 (casebound), 0-8186-5982-3 (microfiche). ISSN ???? LCCN QA 76 S979 1989; TK7885.A1 S92 1989. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog no. 89CH2808-4. Computer Society order no. 1982.

Milne-Thomson:1964:EI

[MT64a] L. M. Milne-Thomson. Elliptic integrals. In Abramowitz and Stegun [AS64], pages 587–626. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://

www.convertit.com/Go/ConvertIt/Reference/AMS55.ASP in

bitmap image format.

Milne-Thomson:1964:JEF

[MT64b] L. M. Milne-Thomson. Jacobian elliptic functions and theta functions. In Abramowitz and Stegun [AS64], pages 567–586.

LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also avail-

able online at http://www.convertit.com/Go/ConvertIt/

Reference/AMS55.ASP in bitmap image format.

Manos:1972:CCA

[MT72] Paul Manos and L. Richard Turner. Constrained Chebyshev approximations to some elementary functions suitable for evaluation with floating-point arithmetic. NASA Technical Note TN D-6698, NASA, Washington, DC, USA, March 1972. iii + 68

pp. URL http://ntrs.nasa.gov/archive/nasa/casi.ntrs.
nasa.gov/19720010958_1972010958.pdf.

Miranker:1985:ASC

[MT85] Willard L. Miranker and Richard A. Toupin, editors. Accurate Scientific Computations: Symposium, Bad Neuenahr, FRG, March 12–14, 1985: Proceedings, volume 235 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-16798-6. LCCN QA76.95 .A231 1986.

Miranker:1986:ASC

[MT86] Willard L. Miranker and Richard A. Toupin, editors. Accurate scientific computations: symposium, Bad Neuenahr, FRG, March 12–14, 1985: proceedings, volume 235 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. CODEN LNCSD9. ISBN 0-387-16798-6 (USA: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.95 .A231 1986. Symposium sponsored by IBM Deutschland.

McCormick:1982:EFM

[MTP82] S. F. McCormick, G. D. Taylor, and D. V. Pryor. Evaluation of functions on microcomputers: ln(x). Computers and Mathematics with Applications, 8(5):389–392, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0898122182900323.

Muller:2015:ISC

[MTV15] Jean-Michel Muller, Arnaud Tisserand, and Julio Villalba, editors. 2015 IEEE 22nd Symposium on Computer Arithmetic (ARITH 2015) Lyon, France, 22–24 June 2015. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2015. ISBN 1-4799-8665-8, 1-4799-8663-1. ISSN 1063-6889. LCCN QA76.9.C62 S95 2015. URL http:/

/ieeexplore.ieee.org/servlet/opac?punumber=7193754.

Mulvey:1982:EMP

[Mul82] J. M. Mulvey, editor. Evaluating Mathematical Programming Techniques: Proceedings of a Conference Held at the National Bureau of Standards, Boulder, Colorado, January 5–6, 1981, volume 199 of Lecture Notes in Economics and Mathematical Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-11495-5. LCCN QA402.5 .E94 1982.

Muller:1985:DBC

[Mul85a] Jean-Michel Muller. Discrete basis and computation of elementary functions. IEEE Transactions on Computers, C-34 (9):857–862, September 1985. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=1676643.

Muller:1985:RNR

[Mul85b] Jean-Michel Muller. Représentation des nombres réels et calcul des fonctions élémentaires. (French) [representation of real numbers and calculation of elementary functions]. In Seminar on number theory, 1984–1985 (Talence, 1984/1985), volume 12, page 22. Univ. Bordeaux I, Talence, France, 1985.

Muller:1986:MDC

[Mul86] Jean-Michel Muller. Une méthodologie du calcul hardware des fonctions élémentaires. (French) [A methodology for the hardware computation of elementary functions]. Mathematical modelling and numerical analysis = Modelisation mathématique et analyse numérique: M^{2}AN, 20(4):667–695, 1986. CODEN RM-MAEV. ISSN 0764-583X (print), 1290-3841 (electronic).

Muller:1990:HCA

[Mül90] Volker Müller. Hochgenaue CORDIC-Algorithmen für reelle Standardfunktionen mittels dynamischer Defektberechnung [English: High-accuracy CORDIC Algorithms for Real Elementary Functions by Means of Dynamic Error Computation].

Diplomarbeit, Institut für angewandte Mathematik, Universität Karlsruhe, Karlsruhe, Germany, December 1990. ???? pp.

Muller:1993:NAC

[Mül93] Jürgen Müller. On numerical analytic continuation and convergence acceleration by summability methods. Analysis, 13(3): 279–291, 1993. ISSN 0174-4747.

Muldoon:1995:EZB

[Mul95] Martin E. Muldoon. Electrostatics and zeros of Bessel functions. Journal of Computational and Applied Mathematics, 65 (1–3):299–308, December 29, 1995. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0377042795001182.

Muller:1997:EFA

[Mul97] Jean-Michel Muller. Elementary Functions: Algorithms and Implementation. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1997. ISBN 0-8176-3990-X. xv + 204 pp. LCCN QA331.M866 1997. US$59.95. URL http://www.

birkhauser.com/cgi-win/ISBN/0-8176-3990-X;   http://
www.ens-lyon.fr/~jmmuller/book_functions.html.

Muller:1999:CAT

[Mül99] J. Müller. Convergence acceleration of Taylor sections by convolution. Constructive Approximation, 15(4):523–536, 1999. ISSN 0176-4276 (print), 1432-0940 (electronic).

Muller:2001:CCH

[Mul01] Keith E. Muller. Computing the confluent hypergeometric function, M(a, b, x). Numerische Mathematik, 90(1):179–196, November 2001. CODEN NUMMA7. ISSN 0029-599X (print),

0945-3245 (electronic). URL http://link.springer-ny.com/
link/service/journals/00211/bibs/1090001/10900179.htm;
 http://link.springer-ny.com/link/service/journals/00211/
papers/1090001/10900179.pdf.

Muller:2004:CSR

[Mül04] Siguna Müller. On the computation of square roots in finite fields. Designs, Codes, and Cryptography, 31(3):301–312, March 2004. CODEN DCCREC. ISSN 0925-1022 (print), 1573-

7586 (electronic). URL http://ipsapp008.kluweronline.
com/IPS/content/ext/x/J/4630/I/61/A/8/abstract.htm.

Muller:2006:GFA

[Mul06] Jean-Michel Muller. Generating function approximations at compile time. In Michael B. Matthews, editor, 2006 Fortieth Asilomar Conference on Signals, Systems and Computers. October 29–November 1, 2006. Pacific Grove, California, pages 328– 331. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 1-4244-0785-0.

Muller:2020:EFA

[Mul20] Jean-Michel Muller. Elementary functions and approximate computing. Proceedings of the IEEE, 108(12):2136–2149, December 2020. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Muller:2024:SNC

[Mul24] Jean-Michel Muller. Some notes on correct rounding of functions. Attachment to STDS-754 mailing list, September 18, 2024.

Muroi:1988:ECR

[Mur88] Kazuo Muroi. Extraction of cube roots in Babylonian mathematics. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 31(3):181–188, October 1988. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Muroi:1999:ESR

[Mur99] Kazuo Muroi. Extraction of square roots in Babylonian mathematics. Historia Scientiarum. Second Series. International Journal of the History of Science Society of Japan, 9(2):127– 133, November 1999. CODEN HISCDU. ISSN 0285-4821.

Musielak:1987:AEG

[Mus87] J. Musielak. Approximation of elements of a generalized Orlicz sequence space by nonlinear, singular kernels. Journal of Approximation Theory, 50(4):366–372, August 1987. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Morozov:1998:NWR

[MV98] D. Kh. Morozov and V. V. Voitsekhovich. A new wide-range approximation of modified Bessel functions in terms of elementary functions. Rev. Mexicana Fís., 44(3):231–234, 1998. CODEN RMXFAT. ISSN 0035-001X.

Muench:1983:LAF

[MW83] Donald L. Muench and Gerald Wildenberg. A logarithm algorithm for a five-function calculator. Two-Year College Mathematics Journal, 14(4):324–326, September 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.1080/00494925.1983.
11972707.

Moroz:2016:FCI

[MWH^{+}16] Leonid V. Moroz, Cezary J. Walczyk, Andriy Hrynchyshyn, Vijay Holimath, and Jan L. Cieśliński. Fast calculation of inverse square root with the use of magic constant — analytical approach. arXiv.org, ??(??):1–23, March 14, 2016. URL https:/

/arxiv.org/pdf/1603.04483.pdf.

Moroz:2018:FCI

[MWH^{+}18] Leonid V. Moroz, Cezary J. Walczyk, Andriy Hrynchyshyn, Vijay Holimath, and Jan L. Cieśliński. Fast calculation of inverse square root with the use of magic constant — analytical approach. Applied Mathematics and Computation, 316 (??):245–255, January 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300317305763.

Midy:1991:CSE

[MY91] P. Midy and Y. Yakovlev. Computing some elementary functions of a complex variable. Mathematics and Computers in Simulation, 33(1):33–49, 1991. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Marsaglia:1994:REI

[MZM94] George Marsaglia, Arif Zaman, and John C. W. Marsaglia. Rapid evaluation of the inverse of the normal distribution function. Statistics & Probability Letters, 19(4):259–266, March 15, 1994. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Nadarajah:2015:CGH

[Nad15] Saralees Nadarajah. On the computation of Gauss hypergeometric functions. The American Statistician, 69(2):146–148, 2015. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Nagashima:1968:EFN

[Nag68] Takashi Nagashima. On elementary functions of natural numbers. Hitotsubashi J. Arts Sci., 9:50–58, 1968. ISSN 0073-2788.

Nagel:2001:EHF

[Nag01] Bengt Nagel. An expansion of the hypergeometric function in Bessel functions. Journal of Mathematical Physics, 42(12):5910– 5914, December 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nagel:2004:CEG

[Nag04] Bengt Nagel. Confluence expansions of the generalized hypergeometric function. Journal of Mathematical Physics, 45(1):495– 508, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i1/p495_s1.

Nannarelli:2011:RCD

[Nan11] Alberto Nannarelli. Radix-16 combined division and square root unit. In Schwarz and Oklobdzija [SO11], pages 169– 176. ISBN 0-7695-4318-9, 1-4244-9457-5. LCCN QA76.6.

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=&
arnumber=5992122.

Nasell:1974:IMB

[Nas74] Ingemar Nasell. Inequalities for modified Bessel functions. Mathematics of Computation, 28(125):253–256, January 1974. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nave:1983:ITF

[Nav83] Rafi Nave. Implementation of transcendental functions on a numerics processor. Microprocessing and Microprogramming, 11: 221–225, 1983. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

Nellis:1966:REE

[NC66] W. J. Nellis and B. C. Carlson. Reduction and evaluation of elliptic integrals. Mathematics of Computation, 20(94):223–231, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ng:1970:CDF

[ND70] E. W. Ng and C. J. Devine. On the computation of Debye functions of integer orders. Mathematics of Computation, 24 (110):405–407, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nemes:2019:AEI

[ND19] Gergő Nemes and Adri B. Olde Daalhuis. Asymptotic expansions for the incomplete gamma function in the transition regions. Mathematics of Computation, 88(318):1805– 1827, April 2019. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL https://www.ams.org/

journals/mcom/2019-88-318/S0025-5718-2018-03391-9;
https://www.ams.org/journals/mcom/2019-88-318/S0025-
5718-2018-03391-9/S0025-5718-2018-03391-9.pdf;   https:
//www.ams.org/mathscinet/search/authors.html?authorName=
Nemes%2C%20Gergo; https://www.ams.org/mathscinet/search/
authors.html?mrauthid=293428.

Ng:1969:CPE

[NDT69] Edward W. Ng, C. J. Devine, and R. F. Tooper. Chebyshev polynomial expansion of Bose–Einstein functions of orders 1 to 10. Mathematics of Computation, 23(107):639–643, July 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Neher:2007:CSF

[Neh07] Markus Neher. Complex standard functions and their implementation in the CoStLy library. ACM Transactions on Mathematical Software, 33(1):2:1–2:27, March 2007. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Nemeth:1965:CEF

[Ném65] G. Németh. Chebyshev expansions for Fresnel integrals. Numerische Mathematik, 7(4):310–312, August 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Nemes:2013:EBE

[Nem13] Gergő Nemes. Error bounds and exponential improvement for Hermite’s asymptotic expansion for the gamma function. Applicable Analysis and Discrete Mathematics, 7(1):161–179, April 2013. ISSN 1452-8630 (print), 2406-100X (electronic). URL

https://doi.org/10.2298/AADM130124002N.

Nesbet:1984:ARI

[Nes84] R. K. Nesbet. Algorithms for regular and irregular Coulomb and Bessel functions. Computer Physics Communications, 32(4): 341–347, July 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465584900511.

Newman:1984:ABS

[New84] J. N. Newman. Approximations for the Bessel and Struve functions. Mathematics of Computation, 43(168):551–556, October 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Newman:1963:ICS

[NF63] J. N. Newman and W. Frank. An integral containing the square of a Bessel function. Mathematics of Computation, 17(81):64– 70, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ng:1968:DSS

[Ng68] Edward W. Ng. On the direct summation of series involving higher transcendental functions. Journal of Computational Physics, 3(2):334–338, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999168900296.

Ng:1970:CAE

[Ng70] E. N. Ng. Certification of Algorithm 385, Exponential Integral Ei(x). Communications of the ACM, 13(7):448–449, July 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ng:1975:CCM

[Ng75] Edward W. Ng. A comparison of computational methods and algorithms for the complex gamma function. ACM Transactions on Mathematical Software, 1(1):56–70, March 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ng:1977:CAL

[Ng77] E. W. Ng. Computations and applications of linear hypergeometric transformations. Computers and Mathematics with Applications, 3(1):65–70, 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Ng:1984:DAA

[Ng84] K. C. Ng. Contributions to the computation of the matrix exponential. Report PAM-212, Center for Pure and Applied Mathematics, University of California, Berkeley, Berkeley, CA, USA, February 1984. Based on the author’s Ph.D. thesis.

Nguyen:1998:MLS

[Ngu98] Phong Nguyen. A Montgomery-like square root for the number field sieve. Lecture Notes in Computer Science, 1423:151– ??, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1423/14230151.htm;   http:/
/link.springer-ny.com/link/service/series/0558/papers/
1423/14230151.pdf.

Ninomiya:1970:BRS

[Nin70] Ichizo Ninomiya. Best rational starting approximations and improved Newton iteration for the square root. Mathematics of Computation, 24(110):391–404, April 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Nishimoto:1984:TFD

[Nis84] Katsuyuki Nishimoto. Tables of fractional differintegrations of elementary functions. J. College Engrg. Nihon Univ. Ser. B, 25: 41–46, 1984. ISSN 0285-6182.

Nishioka:1994:EFB

[Nis94] Keiji Nishioka. Elementary functions based on elliptic curves. Tokyo journal of mathematics, 17(2):439–446, 1994. ISSN 0387-3870.

Nam:2008:PAE

[NKY08] Byeong-Gyu Nam, Hyejung Kim, and Hoi-Jun Yoo. Power and area-efficient unified computation of vector and elementary functions for handheld 3D graphics systems. IEEE Transactions on Computers, 57(4):490–504, April 2008. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4432232.

Narayanaswami:1994:AE

[NL94] Chandrasekhar Narayanaswami and William Luken. Approximating x^{n} efficiently. Information Processing Letters, 50(4): 205–210, May 25, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nannarelli:1999:LPR

[NL99] A. Nannarelli and T. Lang. Low-power radix-4 combined division and square root. In (ICCD ’99) International Conference on Computer Design, pages 236–242. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. CODEN ???? ISSN ????

Norlund:1950:HF

[Nør50] Niels Erik Nørlund. Hypergeometric functions. Mat. Tidsskr. B., 1950(??):18–21, 1950.

Norlund:1955:HF

[Nør55] Niels Erik Nørlund. Hypergeometric functions. Acta Mathematica, 94(??):289–349, ???? 1955. CODEN ACMAA8. ISSN 0001-5962 (print), 1871-2509 (electronic).

Norton:1989:PCA

[Nor89] Robert M. Norton. Pocket-calculator approximation for areas under the standard normal curve. The American Statistician, 43(1):24–26, February 1989. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2685163.

Nowak:2006:MCA

[Now06] Rafal Nowak. A method of convergence acceleration of some continued fractions. Numerical Algorithms, 41(3):297–317, March 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
41&issue=3&spage=297.

Navas-Palencia:2018:FAA

[NP18a] Guillermo Navas-Palencia. Fast and accurate algorithm for the generalized exponential integral E_{ν}(x) for positive real order. Numerical Algorithms, 77(2):603–630, February 2018.

CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s11075-017-0331-z.

Navas-Palencia:2018:HPC

[NP18b] Guillermo Navas-Palencia. High-precision computation of the confluent hypergeometric functions via Franklin–Friedman expansion. Advances in Computational Mathematics, 44(3):841– 859, June 2018. CODEN ACMHEX. ISSN 1019-7168 (print),

1572-9044 (electronic). URL https://link.springer.com/
article/10.1007/s10444-017-9565-5.

Navas-Palencia:2016:CCH

[NPA16] Guillermo Navas-Palencia and Argimiro Arratia. On the computation of confluent hypergeometric functions for large imaginary part of parameters b and z. In Greuel [Gre16], pages 241–248. ISBN 3-319-42431-9 (print), 3-319-42432-7 (electronic). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.9.M35.

URL http://zbmath.org/?q=an:1342.68017.

Natalini:2015:BPM

[NR15] Pierpaolo Natalini and Paolo Emilio Ricci. Bell polynomials and modified Bessel functions of half-integral order. Applied Mathematics and Computation, 268(??):270–274, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300315008504.

Nagayama:2009:CGB

[NS09] S. Nagayama and T. Sasao. Complexities of graph-based representations for elementary functions. IEEE Transactions on Computers, 58(1):106–119, January 2009. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4599569.

Neta:2013:FHL

[NS13] Beny Neta and Melvin Scott. On a family of Halley-like methods to find simple roots of nonlinear equations. Applied Mathematics and Computation, 219(15):7940–7944, April 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300313001574.

Nandagopal:2010:NEF

[NSR10] Mohankumar Nandagopal, Soubhadra Sen, and Ajay Rawat. A note on the error function. Computing in Science and Engineering, 12(4):84–88, July/August 2010. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). See improvements [Iac21].

Naterop:2020:HRN

[NSU20] L. Naterop, A. Signer, and Y. Ulrich. handyG — rapid numerical evaluation of generalised polylogarithms in Fortran. Computer Physics Communications, 253(??):Article 107165, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465520300230.

Noble:1984:CPE

[NT84] C. J. Noble and I. J. Thompson. COULN, a program for evaluating negative energy Coulomb functions. Computer Physics Communications, 33(4):413–419, October 1984. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0010465584901462.

Nikiforov:1988:SFM

[NU88] Arnol’d F. Nikiforov and Vasilij B. Uvarov. Special functions of mathematical physics: a unified introduction with applications. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1988. ISBN 3-7643-3183-6, 0-8176-3183-6. xviii + 427 pp. LCCN QC20.7.F87 N692e. URL http://www.

gbv.de/dms/hbz/toc/ht002696178. Translated from the Rus-

sian by Ralph P. Boas.

OBrien:1980:SBF

[O’B80] D. M. O’Brien. Spherical Bessel functions of large order. Journal of Computational Physics, 36(1):128–132, June 1980. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999180901771.

Oberhettinger:1964:HF

[Obe64] Frtiz Oberhettinger. Hypergeometric functions. In Abramowitz and Stegun [AS64], pages 555–566. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://

www.convertit.com/Go/ConvertIt/Reference/AMS55.ASP in

bitmap image format.

Oberman:1999:FPD

[Obe99] S. F. Oberman. Floating point division and square root algorithms and implementation in the AMD-K7^{T M} microprocessor. In Koren and Kornerup [KK99], pages 106–115. CODEN ???? ISBN 0-7803-5609-8, 0-7695-0116-8, 0-7695-0118-4. ISSN ???? LCCN QA76.6 .S887 1999. URL http://euler.ecs.

umass.edu/paper/final/paper-139.pdf;   http://euler.
ecs.umass.edu/paper/final/paper-139.ps. IEEE Computer

Society Order Number PR00116. IEEE Order Plan Catalog Number 99CB36336.

Oklobdzija:1982:LSR

[OE82] V. G. Oklobdzija and M. D. Ercegovac. An on-line square root algorithm. IEEE Transactions on Computers, C-31(1):70–75, January 1982. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-

9956 (electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=1675887.

Ozcag:2016:RPI

[ÖE16] Emin Özca.ğ and İnci Ege. Remarks on polygamma and incomplete gamma type functions. Journal of Number Theory, 169(??):369–387, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16301378.

Ollin:1982:CFE

[OG82] H. Z. Ollin and I. Gerst. Classes of functions with explicit best uniform approximations. Journal of Approximation Theory, 34 (3):264–276, March 1982. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Ohta:1994:INP

[OGWY94] Shigemi Ohta, Eiichi Goto, Weng Fai Wong, and Nobuaki Yoshida. Improvement and new proposal on fast evaluation of elementary functions. (Japanese). Transactions of the Information Processing Society of Japan, 35(5):926–933, May 1994. CODEN JSGRD5. ISSN 0387-5806.

Ortiz:2004:SPI

[OJ04] I. Ortiz and M. Jimenez. Scalable pipeline insertion in floatingpoint division and square root units. In MWSCAS ’04. The 2004 47th Midwest Symposium on Circuits and Systems. 25–28 July 2004, volume 2, pages II–225–II–228. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. CODEN ???? ISSN ????

Oh:2022:EPA

[OJL22] Hyun Woo Oh, Won Sik Jeong, and Seung Eun Lee. Evaluation of posit arithmetic on machine learning based on approximate exponential functions. In IEEE, editor, 2022 19th International SoC Design Conference (ISOCC), pages 358–359. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2022.

Schulten:1979:AEC

[oKSAG79] Z. (or K. ??) Schulten, D. G. M. Anderson, and R. G. Gordon. An algorithm for the evaluation of the complex Airy functions. Journal of Computational Physics, 31(1):60–75, April 1979. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Olver:2010:NHM

[OLBC10] Frank W. J. Olver, Daniel W. Lozier, Ronald F. Boisvert, and Charles W. Clark, editors. NIST Handbook of Mathematical Functions. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19225-0. xv + 951 pp. LCCN QA331 .N57

2010. US$99.00. URL http://dlmf.nist.gov/;   http://
www.cambridge.org/9780521140638.

Oleksy:1996:CAM

[Ole96] Cz. Oleksy. A convergence acceleration method of Fourier series. Computer Physics Communications, 96(1):17–26, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

OLeary:1995:NRI

[OLHA95] J. O’Leary, M. Leeser, J. Hickey, and M. Aagaard. Non-restoring integer square root: a case study in design by principled optimization. Lecture Notes in Computer Science, 901:52–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oliver:1975:SME

[Oli75] J. Oliver. Stable methods for evaluating the points cos(iπ/n). Journal of the Institute of Mathematics and its Applications, 16 (2):247–257, October 1975. CODEN JMTAA8. ISSN 0020-2932.

Olver:1964:BFI

[Olv64] F. W. J. Olver. Bessel functions of integer order. In Abramowitz and Stegun [AS64], pages 355–434. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://

www.convertit.com/Go/ConvertIt/Reference/AMS55.ASP in

bitmap image format.

Olver:1967:BSS

[Olv67a] F. W. J. Olver. Bounds for the solutions of second-order linear difference equations [Anger–Weber and Struve functions]. Journal of Research of the National Bureau of Standards (1934), 71B (4):161–166, October 1967. ISSN 0091-0635.

Olver:1967:NSS

[Olv67b] F. W. J. Olver. Numerical solution of second-order linear difference equations. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 71B(2–3):111–129, April 1967. CODEN JNBBAU. ISSN 0022-

4340. URL https://nvlpubs.nist.gov/nistpubs/jres/71B/
jresv71Bn2-3p111_A1b.pdf.

Olver:1974:ASF

[Olv74] F. W. J. Olver. Asymptotics and Special Functions. Academic Press, New York, NY, USA, 1974. ISBN 0-12-525850-X. xvi + 572 pp. LCCN QA351 .O481 1974.

Olver:1980:UAG

[Olv80] F. W. J. Olver. Unrestricted algorithms for generating elementary functions. In Alefeld et al. [AGA^{+}80], pages 131–140. ISBN 0-387-81566-X. ISSN 0344-8029. LCCN QA297 .F84.

Olver:1988:EBL

[Olv88] F. W. J. Olver. Error bounds for linear recurrence relations. Mathematics of Computation, 50(182):481–499, April 1988. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Olver:1991:UEIb

[Olv91] F. W. J. Olver. Uniform, exponentially improved, asymptotic expansions for the confluent hypergeometric function and other integral transforms. SIAM journal on mathematical analysis, 22(5):1475–1489, September 1991. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Olver:1994:GEI

[Olv94] F. W. J. Olver. The generalized exponential integral. In Zahar [Zah94], pages 497–510. ISBN 0-8176-3753-2. LCCN QA221 .A634 1994.

Osborn:1968:IBF

[OM68] David Osborn and Richard Madey. The incomplete beta function and its ratio to the complete beta function. Mathematics of Computation, 22(101):159–162, January 1968. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Oldham:2009:AF

[OMS09] Keith B. Oldham, Jan Myland, and Jerome Spanier, editors. An Atlas of Functions: With Equator, the Atlas Function Calculator. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2009. ISBN 0-387-48807-3 (softcover), 0-387-48806-5 (hardcover). xi + 748 pp. LCCN QA331 .S685 2009. US$129.95.

Olver:1972:NBR

[OS72] F. W. J. Olver and D. J. Sookne. Note on backward recurrence algorithms. Mathematics of Computation, 26(120):941–947, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

OSullivan:2016:ZD

[O’S16] Cormac O’Sullivan. Zeros of the dilogarithm. Mathematics of Computation, 85(302):2967–2993, November 2016. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/journals/mcom/2016-85-
302/S0025-5718-2016-03065-3;   http://www.ams.org/journals/
mcom/2016-85-302/S0025-5718-2016-03065-3/S0025-5718-
2016-03065-3.pdf;   http://www.ams.org/mathscinet/search/
author.html?mrauthid=658848.

Osada:1990:CAM

[Osa90] Naoki Osada. A convergence acceleration method for some logarithmically convergent sequences. SIAM Journal on Numerical Analysis, 27(1):178–189, February 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Osada:1994:CAM

[Osa94] Naoki Osada. Convergence acceleration methods. Sūrikaisekikenkyūsho Kōkyūroku, 880(??):28–43, ???? 1994. The state of the art of scientific computing and its prospects (Japanese) (Kyoto, 1993).

Osipov:2013:PSW

[Osi13] Andrei Osipov. Prolate Spheroidal Wave Functions of Order Zero: Mathematical Tools for Bandlimited Approximation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-8258-5.

???? pp. LCCN ???? URL http://www.loc.gov/catdir/
enhancements/fy1315/2013945079-b.html;   http://www.
loc.gov/catdir/enhancements/fy1315/2013945079-d.html;
 http://www.loc.gov/catdir/enhancements/fy1315/2013945079-
t.html.

Opps:2009:RFA

[OSS09] Sheldon B. Opps, Nasser Saad, and H. M. Srivastava. Recursion formulas for Appell’s hypergeometric function F_{2} with some applications to radiation field problems. Applied Mathematics and Computation, 207(2):545–558, January 15, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ostrowski:1937:KAN

[Ost37] A. M. Ostrowski. Über die Konvergenz und die Abründungsfestigkeit des Newtonschen Verfahren. (German) [On the convergence and rounding strength of Newton’s method]. Rec. Math., 2(??):1073– 1098, ???? 1937.

Ostrowski:1938:NMA

[Ost38] A. M. Ostrowski. On Newton’s method of approximation. British Association for the Advancement of Science, ??(??):392– ??, ???? 1938.

Ostrowski:1954:TPA

[Ost54] A. M. Ostrowski. On two problems in abstract algebra connected with Horner’s rule. In Birkhoff [Bir54], pages 40–48. ISBN 1-4832-6356-8. LCCN QA3 .S853.

Okabe:1991:LDC

[OTY91] Y. Okabe, N. Takagi, and S. Yaima. Log-depth circuits for elementary functions using residue number system. Electronics and communications in Japan, 74(8):31–37, 1991. CODEN ECO-JAL. ISSN 0424-8368. Translated from Denshi Joho Tsushin Gakkai Ronbunshi, vol. 21-DI, no. 9, September 1990, pp. 723-728.

Ovtchinnikov:2003:CEGb

[Ovt03] E. Ovtchinnikov. Convergence estimates for the generalized Davidson method for symmetric eigenvalue problems II: The subspace acceleration. SIAM Journal on Numerical Analysis, 41(1):272–286, February 2003. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://epubs.siam.

org/sam-bin/dbq/article/41176.

OBrien:1968:CAC

[OW68] William M. O’Brien and Joan Wood. Certification of Algorithm 299 [S15]: Chi-squared integral. Communications of the ACM, 11(4):271, April 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ogburn:2014:FDC

[OWS^{+}14] Daniel X. Ogburn, Colin L. Waters, Murray D. Sciffer, Jeff A. Hogan, and Paul C. Abbott. A finite difference construction of the spheroidal wave functions. Computer Physics Communications, 185(1):244–253, January 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465513002610.

OGrady:1991:HOA

[OY91] E. Pearse O’Grady and Baek-Kyu Young. A hardware-oriented algorithm for floating-point function generation. IEEE Transactions on Computers, 40(2):237–241, February 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Ozban:2006:NMA

[Özb06] Ahmet Yaşar Özban. New methods for approximating square roots. Applied Mathematics and Computation, 175(1):532–540, April 1, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Piessens:1986:ATP

[PA86] Robert Piessens and Shafique Ahmed. Approximation for the turning points of Bessel functions. Journal of Computational Physics, 64(1):253–257, May 1986. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199918690029X.

Paciorek:1970:AEI

[Pac70] K. A. Paciorek. Algorithm 385: Exponential integral Ei(x). Communications of the ACM, 13(7):446–447, July 1970. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Red70, Fri72].

Page:1977:MAC

[Pag77] E. Page. Miscellanea: Approximations to the cumulative normal function and its inverse for use on a pocket calculator. Applied Statistics, 26(1):75–76, 1977. CODEN APSTAG. ISSN 0035-9254 (print), 1467-9876 (electronic).

Pairman:1919:TDT

[Pai19] Eleanor Pairman. Tables of the Digamma and Trigamma Functions, volume I of Tracts for computers. Cambridge University Press, Cambridge, UK, 1919. 9 + 11 pp. LCCN QA47 .T7 no.1.

Pakes:2018:LFN

[Pak18] Anthony G. Pakes. The Lambert W function, Nuttall’s integral, and the Lambert law. Statistics & Probability Letters, 139(??): 53–60, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218301354.

Palumbo:1998:GSI

[Pal98] Biagio Palumbo. A generalization of some inequalities for the gamma function. Journal of Computational and Applied Mathematics, 88(2):255–268, March 2, 1998. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042797001878.

Pan:1959:CSC

[Pan59a] V. Ya. Pan. Certain schemes for the calculation of values of polynomials with real coefficients. Problemy Kibernetiki, 5(??): 17–29, ???? 1959. Also available as English translation JPRS No. 11045.

Pan:1959:SCP

[Pan59b] V. Ya. Pan. Schemes for the computation of polynomials with real coefficients. Doklady Akademii nauk SSSR, 127(2):266–269, ???? 1959. CODEN DANKAS. ISSN 0002-3264.

Parnes:1972:CZM

[Par72] R. Parnes. Complex zeros of the modified Bessel function K_{n}(Z). Mathematics of Computation, 26(120):949–953, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Parhami:1999:ALT

[Par99] B. Parhami. Analysis of the lookup table size for square-rooting. In Conference Record of the Thirty-Third Asilomar Conference on Signals, Systems, and Computers, 1999, volume 2, pages 1327–1330. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. CODEN ???? ISSN ????

Paris:2002:EBU

[Par02a] R. B. Paris. Error bounds for the uniform asymptotic expansion of the incomplete gamma function. Journal of Computational and Applied Mathematics, 147(1):215–231, October 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037704270200434X.

Paris:2002:UAE

[Par02b] R. B. Paris. A uniform asymptotic expansion for the incomplete gamma function. Journal of Computational and Applied Mathematics, 148(2):323–339, November 2002. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). See related work [Par16].

Paris:2003:AEG

[Par03] R. B. Paris. The asymptotic expansion of a generalised incomplete gamma function. Journal of Computational and Applied Mathematics, 151(2):297–306, February 15, 2003. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042702008099.

Parks:2006:UTS

[Par06] Michael Parks. Unifying tests for square root. In Anonymous [Ano06], page ?? ISBN ???? LCCN ????

Paris:2009:HPE

[Par09] R. B. Paris. High-precision evaluation of the Bessel functions via Hadamard series. Journal of Computational and Applied Mathematics, 224(1):84–100, February 1, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042708001799.

Paris:2016:UAE

[Par16] R. B. Paris. A uniform asymptotic expansion for the incomplete gamma functions revisited. arxiv.org, ??(??):1–9, November 2,

2016. URL https://arxiv.org/abs/1611.00548.

Paszkowski:1992:CAC

[Pas92] Stefan Paszkowski. Convergence acceleration of continued fractions of Poincaré’s type 1. Numerical Algorithms, 2(2):155–170, ???? 1992. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Paszkowski:1995:QHS

[Pas95] S. Paszkowski. Quasipower and hypergeometric series— construction and evaluation. Numerical Algorithms, 10(3–4): 337–361, October 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Paszkowski:2003:CAS

[Pas03] Stefan Paszkowski. Convergence acceleration of some continued fractions. Numerical Algorithms, 32(2–4):193–247, April 2003. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.
com/content/getfile/5058/45/5/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/45/
5/fulltext.pdf.

Paszkowski:2008:CAO

[Pas08] Stefan Paszkowski. Convergence acceleration of orthogonal series. Numerical Algorithms, 47(1):35–62, January 2008. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=47&issue=1&spage=35.

Paszkowski:2010:UMC

[Pas10] Stefan Paszkowski. Untypical methods of convergence acceleration. Numerical Algorithms, 54(??):??, ???? 2010. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=1017-1398&volume=0&issue=0&spage=
??.

Paszkowski:2011:UMC

[Pas11] Stefan Paszkowski. Untypical methods of convergence acceleration. Numerical Algorithms, 56(2):185–209, ???? 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Patterson:2018:SCS

[Pat18] T. N. L. Patterson. Sines, cosines, square roots, and binary bits. American Mathematical Monthly, 125(8):750–754, 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Pawellek:2011:GJE

[Paw11] Michael Pawellek. On a generalization of Jacobi’s elliptic functions and the double sine–Gordon kink chain. Journal of Mathematical Physics, 52(11):113701, November 2011. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v52/
i11/p113701_s1.

Piessens:1982:ABF

[PB82] R. Piessens and Maria Branders. Approximation for Bessel functions and their application in the computation of Hankel transforms. Computers and Mathematics with Applications, 8(4):305– 311, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0898122182900128.

Piessens:1983:MCC

[PB83] Robert Piessens and Maria Branders. Modified Clenshaw– Curtis method for the computation of Bessel function integrals. BIT (Nordisk tidskrift for informationsbehandling), 23 (3):370–381, September 1983. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=23&issue=3&spage=370.

Pineiro:2002:HSD

[PB02] J. A. Piñeiro and J. D. Bruguera. High-speed double precision computation of reciprocal, division, square root, and inverse square root. IEEE Transactions on Computers, 51(12): 1377–1388, December 2002. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Piso:2008:NRA

[PB08] D. Piso and J. D. Bruguera. A new rounding algorithm for variable latency division and square root implementations. In Luca Fanucci, editor, Proceedings: 11th Euromicro Symposium on Digital Systems Design: Architectures, Methods and Tools (DSD 2008), Parma, Italy, September 3–5, 2008, pages 760– 767. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2008. ISBN 0-7695-3277-2.

Pineiro:2003:LHR

[PBE03] J.-A. Pineiro, J. D. Bruguera, and M. D. Ercegovac. On-line high-radix exponential with selection by rounding. In ISCAS ’03. Proceedings of the 2003 International Symposium on Circuits and Systems. 25–28 May 2003, volume 4, pages IV–121– IV–124. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. CODEN ???? ISSN ????

Pineiro:2008:RDD

[PBLM08] J.-A. Pineiro, J. D. Bruguera, F. Lamberti, and P. Montuschi. A radix-2 digit-by-digit architecture for cube root. IEEE Transactions on Computers, 57(4):562–566, April 2008.

CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4407683.

Prudnikov:1986:ISE

[PBM86a] Anatolij P. Prudnikov, Jurij A. Bryčkov, and Oleg I. Maričev. Integrals and series. Elementary functions, volume 1. Gordon and Breach Science Publishers, New York, NY, USA, 1986. ISBN 2-88124-089-5. 798 pp. LCCN QA308.P7813 1986.

Prudnikov:1986:ISS

[PBM86b] Anatolij P. Prudnikov, Jurij A. Bryčkov, and Oleg I. Maričev. Integrals and series. Special functions, volume 2. Gordon and Breach Science Publishers, New York, NY, USA, 1986. ISBN 2-88124-090-9. 750 pp. LCCN QA308.P7813 1986.

Prudnikov:1992:IS

[PBM92] Anatolij P. Prudnikov, Jurij A. Bryčkov, and Oleg I. Maričev. Integrals and series. More special functions, volume 3. Gordon and Breach Science Publishers, New York, NY, USA, 1992. ISBN 2-88124-097-6. xx + 618 pp. LCCN QA308 P68 1986.

Perger:1993:NEG

[PBN93] Warren F. Perger, Atul Bhalla, and Mark Nardin. A numerical evaluator for the generalized hypergeometric series. Computer Physics Communications, 77(2):249–254, October 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Petkovic:1993:SII

[PC93] Miodrag S. Petković and Carsten Carstensen. Some improved inclusion methods for polynomial roots with Weierstrass’ corrections. Computers and Mathematics with Applications, 25(3):59– 67, February 1993. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/089812219390143J.

Polyanin:2011:CHM

[PC11] A. D. (Andrei Dmitrievich) Polyanin and A. I. Chernoutisan, editors. A Concise Handbook of Mathematics, Physics, and Engineering Sciences. Chapman and Hall/CRC, Boca Raton, FL, USA, 2011. ISBN 0-429-13137-2, 1-282-90264-4, 1-4398-0639-X (hardcover), 1-4398-0640-3 (PDF). xxviii + 1097 pp. LCCN QA40 .C65 2010.

Panteliou:1996:DII

[PDK96] S. D. Panteliou, A. D. Dimarogonas, and I. N. Katz. Direct and inverse interpolation for Jacobian elliptic functions, zeta function of Jacobi and complete elliptic integrals of the second kind. Computers and Mathematics with Applications, 32(8):51– 57, October 1996. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0898122196001666.

Pearson:2009:CHF

[Pea09] J. Pearson. Computation of hypergeometric functions. Master’s thesis, Oxford University, Oxford, UK, 2009.

Pineiro:2002:HRL

[PEB02] J.-A. Pineiro, M. D. Ercegovac, and J. D. Bruguera. High-radix logarithm with selection by rounding. In The IEEE International Conference on Application-Specific Systems, Architectures and Processors, 2002. Proceedings. 17–19 July 2002, pages 101– 110. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. CODEN ???? ISSN ????

Pineiro:2004:AAL

[PEB04] J. A. Piñeiro, M. D. Ercegovac, and J. D. Bruguera. Algorithm and architecture for logarithm, exponential and powering computation. IEEE Transactions on Computers, 53(9):1085– 1096, 2004. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-

9956 (electronic). URL http://www.ac.usc.es/arquivos/
articulos/2004/gac2004-j05.ps.

Pedersen:1980:HBM

[Ped80] P. W. Pedersen. Hvordan beregner man kvadratroden? [English: How do you calculate the square root?]. Elektronik (Denmark), ??(4):18–21, April 1980.

Pedersen:2003:DGF

[Ped03] Henrik L. Pedersen. The double gamma function and related Pick functions. Journal of Computational and Applied Mathematics, 153(1–2):361–369, April 1, 2003. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042702006040.

Peuizulaev:1963:AEF

[Peĭ63] Š. I. Peĭzulaev. An approximation by elementary functions. (Russian). Ž. Vyčisl. Mat. i Mat. Fiz., 3:769–770, 1963.

Peng:1981:AES

[Pen81] Hong Peng. Algorithms for extracting square roots and cube roots. In IEEE CA5 ’81 [IEE81b], pages 121–126. LCCN

QA 76.6 S985t 1981. URL http://www.acsel-lab.com/
arithmetic/arith5/papers/ARITH5_Peng.pdf. IEEE catalog

number 81CH1630-C.

Peralta:1985:TRN

[Per85a] Rene Caupolican Peralta. Three results in number theory and cryptography: a new algorithm to compute square roots modulo a prime number; On the bit complexity of the discrete logarithm; a framework for the study of cryptoprotocols. Thesis (Ph.D.), Department of Computer Science, University of California, Berkeley, Berkeley, CA, USA, December 1985. 52 pp.

Pereira:1985:ECF

[Per85b] N. Costa Pereira. Estimates for the Chebyshev function ψ(x) − θ(x). Mathematics of Computation, 44(169):211–221, January 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [Per87].

Pereira:1987:CEC

[Per87] N. Costa Pereira. Corrigendum: “Estimates for the Chebyshev function ψ(x) − θ(x)” [Math. Comp. 44 (1985), no. 169, 211– 221, MR 86k:11005]. Mathematics of Computation, 48(177):447, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See [Per85b].

Punta:2018:CFL

[PGA18] Jessica A. Del Punta, Gustavo Gasaneo, and Lorenzo U. Ancarani. On the Laguerre representation of Coulomb functions and the relation to orthogonal polynomials. Advances in Quantum Chemistry, 76:79–101, 2018. CODEN AQCHA9. ISSN

0065-3276. URL https://www.sciencedirect.com/science/
article/pii/S0065327617300643.

Pike:1966:ALG

[PH66] M. C. Pike and I. D. Hill. Algorithm 291: Logarithm of gamma function. Communications of the ACM, 9(9):684, September 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pike:1967:RAI

[PH67] M. C. Pike and I. D. Hill. Remark on Algorithm 179: Incomplete Beta ratio. Communications of the ACM, 10(6):375–376, June 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Palmore:1990:CAC

[PH90] J. Palmore and C. Herring. Computer arithmetic, chaos and fractals. Physica D, 42(1–3):99–110, June 1990. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic). Ninth Annual International Conference of the Center for Nonlinear Studies on Self-Organizing, Collective and Cooperative Phenomena in Natural and Artificial Networks.

Pinchon:2008:NEL

[PHH08] Didier Pinchon, Philip E. Hoggan, and Frank E. Harris. A new expansion of the leaky aquifer function. International Journal of Quantum Chemistry, 108(15):3042–3046, ???? 2008. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Philip:1960:FI

[Phi60] J. R. Philip. The function inverfcθ. Australian Journal of Physics, 13(1):13–20, March 1960. CODEN AUJPAS. ISSN 0004-9506 (print), 1446-5582 (electronic).

Phillips:1979:FAC

[Phi79] G. Phillips. A fast approximation to the complementary error function for use in fitting gamma-ray peaks. Nuclear Instruments and Methods, 164(??):561–563, September 1979. CODEN NUIMAL. ISSN 0029-554x (print), 1878-3759 (electronic). URL

http://adsabs.harvard.edu/abs/1979NucIM.164..561P.

Piparo:2016:CPT

[PI16] D. Piparo and V. Innocente. The CptnHook Profiler — a tool to investigate usage patterns of mathematical functions. Journal of Physics: Conference Series, 762:012038:1–012038:, October 2016. CODEN ???? ISSN 1742-6588 (print), 1742-6596 (electronic).

Piessens:1982:ACB

[Pie82] R. Piessens. Automatic computation of Bessel function integrals. Computer Physics Communications, 25(3):289–295, March 1982. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465582900248.

Piessens:1984:ACB

[Pie84a] R. Piessens. Automatic computation of Bessel function integrals. Computer Physics Communications, 35(1–3):C–791, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S001046558482915X.

Piessens:1984:CSA

[Pie84b] R. Piessens. Chebyshev series approximations for the zeros of the Bessel functions. Journal of Computational Physics, 53(1):188– 192, January 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184900603.

Piessens:1984:SEF

[Pie84c] R. Piessens. A series expansion for the first positive zero of the Bessel functions. Mathematics of Computation, 42(165):195– 197, January 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Piessens:1984:CBF

[Pie84d] Robert Piessens. The computation of Bessel functions on a small computer. Computers and Mathematics with Applications, 10(2):161–166, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0898122184900452.

Piparo:2014:SHE

[PIH14] Danilo Piparo, Vincenzo Innocente, and Thomas Hauth. Speeding up HEP experiment software with a library of fast and auto-vectorisable mathematical functions. Journal of Physics: Conference Series, 513(5):052027, June 2014. ISSN 1742-6588 (print), 1742-6596 (electronic).

Pitteway:1967:RAA

[Pit67] M. L. V. Pitteway. Remark on Algorithm 301: Airy function. Communications of the ACM, 10(7):453, July 1967. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pisarskii:1974:QCD

[PKP74] A. V. Pisarskiĭ, A. F. Kurgaev, and A. V. Palagin. On the question of the convergence of the “digit by digit” methods of computing the elementary functions. Kibernetika (Kiev), 4:147– 149, 1974. CODEN KBRNA5. ISSN 0023-1274.

Paliouras:2000:FPP

[PKS00] V. Paliouras, K. Karagianni, and T. Stouraitis. A floating-point processor for fast and accurate sine/cosine evaluation. IEEE Transactions on Circuits and Systems. 2, Analog and Digital Signal Processing, 47(5):441–451, May 2000. CODEN ICSPE5. ISSN 1057-7130 (print), 1558-125X (electronic).

Parhi:2019:CAF

[PL19] Keshab K. Parhi and Yin Liu. Computing arithmetic functions using stochastic logic by series expansion. IEEE Transactions on Emerging Topics in Computing, 7(1):44–59, January/March 2019. ISSN 2168-6750 (print), 2376-4562 (electronic).

Plofker:1996:ESM

[Plo96] Kim Plofker. An example of the secant method of iterative approximation in a Fifteenth-Century Sanskrit text. Historia Mathematica, 23(3):246–256, August 1996. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0315086096900269.

Plofker:2001:EIT

[Plo01] Kim Plofker. The “Error” in the Indian “Taylor series approximation” to the sine. Historia Mathematica, 28(4):283–295, November 2001. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0315086001923316.

Pradhan:2023:ETB

[PLT23] Chetana Pradhan, Martin Letras, and Jürgen Teich. Efficient table-based function approximation on FPGAs using interval splitting and BRAM instantiation. ACM Transactions on Embedded Computing Systems, 22(4):73:1–73:??, July 2023. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL

https://dl.acm.org/doi/10.1145/3580737.

Parlett:1985:DAA

[PN85] Beresford N. Parlett and K. C. Ng. Development of an accurate algorithm for exp(Bt). Technical Report PAM-294, Center for Pure and Applied Mathematics, University of California, Berkeley, Berkeley, CA, USA, August 1985.

Plagianakos:2001:LCP

[PNV01] V. P. Plagianakos, N. K. Nousis, and M. N. Vrahatis. Locating and computing in parallel all the simple roots of special functions using PVM. Journal of Computational and Applied Mathematics, 133(1–2):545–554, August 1, 2001. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042700006750.

Polya:1949:RCP

[Pól49] G. Pólya. Remarks on computing the probability integral in one and two dimensions. In J. Neyman, editor, Proceedings of the First Berkeley Symposium on Mathematical Statistics and Probability, pages 63–78. University of California Press, Berkeley, CA, USA, 1949.

Polyak:1988:SOM

[Pol88] R. A. Polyak. Smooth optimization methods for minimax problems. SIAM Journal on Control and Optimization, 26(6):1274– 1286, November 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pomeranz:1974:AAE

[Pom74] John Pomeranz. ACM Algorithm 487: Exact cumulative distribution of the Kolmogorov–Smirnov statistic for small samples. Communications of the ACM, 17(12):703–704, December 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Pom76].

Pomeranz:1976:REC

[Pom76] J. Pomeranz. Remark on “Algorithm 487: Exact Cumulative Distribution of the Kolmogorov–Smirnov Statistic for Small Samples [S14]”. ACM Transactions on Mathematical Software, 2(1):111, March 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Pom74].

Pineiro:2005:HSF

[POMB05] Jose-Alejandro Piñeiro, Stuart F. Oberman, Jean-Michel Muller, and Javier D. Bruguera. High-speed function approximation using a minimax quadratic interpolator. IEEE Transactions on Computers, 54(3):304–318, March 2005. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.computer.org/comp/trans/tc/
2005/03/t0304abs.htm;   http://csdl.computer.org/dl/
trans/tc/2005/03/t0304.htm;   http://csdl.computer.org/
dl/trans/tc/2005/03/t0304.pdf;   http://ieeexplore.ieee.
org/iel5/12/30205/01388195.pdf?isnumber=30205&prod=JNL&
arnumber=1388195&arSt=+304&ared=+318&arAuthor=Pineiro%
2C+J.-A.%3B+Oberman%2C+S.F.%3B+Muller%2C+J.-M.%3B+Bruguera%
2C+J.D.;   http://ieeexplore.ieee.org/xpls/abs_all.jsp?
isnumber=30205&arnumber=1388195&count=13&index=6;   http:
//ieeexplore.ieee.org/xpls/references.jsp?arnumber=1388195.

Pearson:2017:NMC

[POP17] John W. Pearson, Sheehan Olver, and Mason A. Porter. Numerical methods for the computation of the confluent and Gauss hypergeometric functions. Numerical Algorithms, 74(3):821–866, March 2017. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://link.springer.com/article/
10.1007/s11075-016-0173-0.

Pornin:2020:OBG

[Por20] Thomas Pornin. Optimized binary GCD for modular inversion. Report ??, International Association for Cryptologic Research, ????, August 23, 2020. 16 pp. URL https://eprint.iacr.

org/2020/972.pdf.

Potter:2014:APC

[Pot14] Ronald W. Potter. Arbitrary Precision Calculation of Selected Higher Functions. Lulu, ????, 2014. ISBN 1-312-59943-X. ???? pp. LCCN QA76.9.A43 P56 2014.

Powell:1988:RBF

[Pow88] J. D. Powell. Radial basis function approximations to polynomials. In D. F. (David Francis) Griffiths and G. A. Watson, editors, Numerical analysis 1987, volume 170 of Pitman research notes in mathematics series, pages 223–241. Longman, Inc., New York, NY, USA, 1988. ISBN 0-582-02157-X. LCCN QA297.N828 1988. US$54.95.

Poelke:2012:DCC

[PP12] Konstantin Poelke and Konrad Polthier. Domain coloring of complex functions: An implementation-oriented introduction. IEEE Computer Graphics and Applications, 32(5):90–97, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Patel:1982:HND

[PR82] Jagdish K. Patel and Campbell B. Read. Handbook of the Normal Distribution, volume 40 of Statistics, textbooks and monographs. Marcel Dekker, Inc., New York, NY, USA, 1982. ISBN 0-8247-1541-1. ix + 337 pp. LCCN QA273.6 .P373.

Patel:1996:HND

[PR96] Jagdish K. Patel and Campbell B. Read. Handbook of the Normal Distribution, volume 150 of Statistics, textbooks and monographs. Marcel Dekker, Inc., New York, NY, USA, second revised and expanded edition, 1996. ISBN 0-8247-9342-0 (hardcover). ix + 431 pp. LCCN QA273.6 .P373 1996. URL http://www.

loc.gov/catdir/enhancements/fy0647/95049404-d.html.

Petkovic:2004:GCS

[PR04] M. S. Petković and L. Rancić. On the guaranteed convergence of the square-root iteration method. Journal of Computational and Applied Mathematics, 170(1):169–179, September 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042704000184.

Preston:1955:ACS

[Pre55] F. S. Preston. An analog computer for the solution of tangents. IRE Transactions on Electronic Computers, EC-4 (3):101–106, ???? 1955. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=5407908.

Preston:1978:NAT

[Pre78] F. S. Preston. A new algorithm for the tangent. IEEE Transactions on Computers, C-27(2):167, February 1978. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=1675052.

Prevost:2010:RVZ

[Pré10] Marc Prévost. Recurrence for values of the zeta function. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1382– 1394, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Price:1966:NAR

[Pri66] James F. Price. Numerical analysis and related literature for scientific computer users. Mathematical Note 456 (D1-82-0517), Mathematics Research Laboratory, Boeing Scientific Research Laboratories, Seattle, WA, USA, March 1966. ix + 191 pp. URL

http://www.dtic.mil/dtic/tr/fulltext/u2/632244.pdf.

Prince:1975:AAF

[Pri75] P. J. Prince. Algorithm 498: Airy functions using Chebyshev series approximations. ACM Transactions on Mathematical Software, 1(4):372–379, December 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [RS81].

Prosser:1983:NCS

[Pro83] C. J. Prosser. A note on computing the square root of an integer. The Computer Journal, 26(2):187–188, May 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_02/tiff/187.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
02/tiff/188.tif.

Proinov:1987:NIA

[Pro87] Petko D. Proinov. Numerical integration and approximation of differentiable functions, II. Journal of Approximation Theory, 50 (4):373–393, August 1987. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Proinov:1988:ISF

[Pro88] Petko D. Proinov. Integration of smooth functions and ϕdiscrepancy. Journal of Approximation Theory, 52(3):284–292, March 1988. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Pexton:1977:RTTa

[PS77] Robert L. Pexton and Arno D. Steiger. Roots of two transcendental equations involving spherical Bessel functions. Mathematics of Computation, 31(139):752–753, July 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pexton:1978:RTT

[PS78] Robert L. Pexton and Arno D. Steiger. Roots of two transcendental equations as functions of a continuous real parameter. Mathematics of Computation, 32(142):511–518, April 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Pexton:1979:DRT

[PS79] Robert L. Pexton and Arno D. Steiger. Degenerate roots of three transcendental equations involving spherical Bessel functions. Mathematics of Computation, 33(147):1041–1048, July 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Petkovic:1986:SIS

[PS86] M. S. Petković and L. V. Stefanović. On some improvements of square root iteration for polynomial complex zeros. Journal of Computational and Applied Mathematics, 15(1):13–25, May 1986. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042786902359.

Povolotskii:1993:OFR

[PS93] A. I. Povolotskiĭ and G. A. Sviridyuk. Odnomernyĭ matematicheskiĭ analiz èlementarnykh funktsiĭ. (Russian) [One-dimensional mathematical analysis of elementary functions] Nepreryvnye funktsii. Differentsiruemye funktsii. Integriruemye funktsii. [Continuous functions. Differentiable functions. Integrable functions]. Chelyabinsk. Gos. Univ., Chelyabinsk, USSR, 1993. ISBN 5-230-17764-0. 92 pp.

Perram:2005:EFW

[PS05] John W. Perram and Edgar R. Smith. Elliptic functions of the worst kind: Non-linear quantisation of the classical spherical pendulum. Advances in Quantum Chemistry, 48:111–125, 2005. CODEN AQCHA9. ISSN 0065-3276. URL http://www.

sciencedirect.com/science/article/pii/S0065327605480089.

Pegoraro:2011:ECV

[PS11] Vincent Pegoraro and Philipp Slusallek. On the evaluation of the complex-valued exponential integral. Journal of Graphics, GPU, and Game Tools, 15(3):183–198, 2011. CODEN ???? ISSN 2151-237X.

Pike:1976:RIB

[PSB76] Malcolm C. Pike, Jennie SooHoo, and N. E. Bosten. Remark on “Algorithm 179: Incomplete Beta Ratio [S14]”. ACM Transactions on Mathematical Software, 2(2):207–208, June 1976. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [Lud63, BB74].

Petropoulou:2003:CZB

[PSS03] Eugenia N. Petropoulou, Panayiotis D. Siafarikas, and Ioannis D. Stabolas. On the common zeros of Bessel functions. Journal of Computational and Applied Mathematics, 153(1–2):387– 393, April 1, 2003. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042702006416.

Popov:1980:PFD

[PT80] B. A. Popov and G. S. Tesler. Priblizhenie funktsii dlya tekhnicheskikh prilozhenii. (Russian) [Approximation of functions for technical applications]. Naukova Dumka, Kiev, USSR, 1980. 351 pp. Nauka i tekhnicheskii progress. [Science and technical progress].

Press:1990:EI

[PT90] William H. Press and Saul A. Teukolsky. Elliptic integrals. Computers in Physics, 4(1):92–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.1063/1.4822893.

Press:1991:BFF

[PT91a] William H. Press and Saul A. Teukolsky. Bessel functions of fractional order. Computers in Physics, 5(2):244–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.
4822982.

Press:1991:MBF

[PT91b] William H. Press and Saul A. Teukolsky. Modified Bessel functions of fractional order. Computers in Physics, 5(3):330–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822991.

Press:2007:NRA

[PTVF07] William H. Press, Saul A. Teukolsky, William T. Vetterling, and Brian P. Flannery. Numerical Recipes — The Art of Scientific Computing. Cambridge University Press, Cambridge, UK, third edition, 2007. ISBN 0-521-88068-8 (hardcover), 0-521-88407-1 (with source code CD ROM), 0-521-70685-8 (source code CD ROM). xxi + 1235 pp. LCCN QA297 .N866 2007. URL http:

//www.cambridge.org/numericalrecipes.

Pugh:2004:ALG

[Pug04] Glendon Ralpha Pugh. An Analysis of the Lanczos Gamma Approximation. Ph.d. thesis, Department of Mathematics, University of British Columbia, Vancouver, BC, Canada, ???? 2004. viii + 154 pp.

Phong:2012:EAG

[PUHM12] Dao Ngoc Phong, Nguyen Quang Uy, Nguyen Xuan Hoai, and R. I. (Bob) McKay. Evolving approximations for the Gaussian Q-function by genetic programming with semantic based crossover. In ????, editor, Proceedings of the 2012 IEEE Congress on Evolutionary Computation, June 10–15, 2012 — Brisbane, QLD, Australia, pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910,

USA, 2012. URL http://ieeexplore.ieee.org/document/
6256588/.

Puoskari:1988:MCB

[Puo88] M. Puoskari. A method for computing Bessel function integrals. Journal of Computational Physics, 75(2):334–344, April 1988. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999188901167.

Paul:1976:SEF

[PW76] George Paul and M. Wayne Wilson. Should the elementary function library be incorporated into computer instruction sets? ACM Transactions on Mathematical Software, 2(2):132–142, June 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Poppe:1990:AEC

[PW90a] G. P. M. Poppe and C. M. J. Wijers. Algorithm 680: Evaluation of the complex error function. ACM Transactions on Mathematical Software, 16(1):47, March 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/pubs/citations/journals/toms/
1990-16-1/p47-poppe/. See remark [Zag19].

Poppe:1990:MEC

[PW90b] G. P. M. Poppe and C. M. J. Wijers. More efficient computation of the complex error function. ACM Transactions on Mathematical Software, 16(1):38–46, March 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/pubs/citations/journals/toms/
1990-16-1/p38-poppe/.

Paris:1992:EIA

[PW92] R. B. Paris and A. D. Wood. Exponentially-improved asymptotics for the gamma function. Journal of Computational and Applied Mathematics, 41(1–2):135–143, August 20, 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037704279290243Q.

Prabhu:1995:MRD

[PZ95] J. A. Prabhu and G. B. Zyner. 167 MHz radix-8 divide and square root using overlapped radix-2 stages. In Knowles and McAllister [KM95], pages 155–162. CODEN ???? ISBN 0-8186-7089-4 (paperback), 0-8186-7089-4 (case), 0-8186-7149-1 (microfiche), 0-8186-7089-4 (softbound), 0-7803-2949-X (casebound). ISSN ???? LCCN QA 76.9 C62 S95 1995.

Qi:2010:CMS

[QGG10] Feng Qi, Senlin Guo, and Bai-Ni Guo. Complete monotonicity of some functions involving polygamma functions. Journal of Computational and Applied Mathematics, 233(9):2149–2160, March 1, 2010. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042709006682.

Qi:2014:IRC

[Qi14] Feng Qi. Integral representations and complete monotonicity related to the remainder of Burnside’s formula for the gamma function. Journal of Computational and Applied Mathematics, 268(??):155–167, October 1, 2014. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042714001356.

Qi:2019:CMD

[QL19] Feng Qi and Ai-Qi Liu. Completely monotonic degrees for a difference between the logarithmic and psi functions. Journal of Computational and Applied Mathematics, 361(??):366–371, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S0377042719302298.

Qi:2015:SIT

[QM15] Feng Qi and Cristinel Mortici. Some inequalities for the trigamma function in terms of the digamma function. Applied Mathematics and Computation, 271(??):502–511, November 15, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300315012758.

Quan:2018:ACA

[QN18] Le Phuong Quan and Thái Anh Nhan. Applying computer algebra systems in approximating the trigonometric functions. Mathematical and Computational Applications, 23(3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:/

/www.mdpi.com/2297-8747/23/3/37.

Quan:2018:CMM

[Qua18] Le Phuong Quan. A computational method with MAPLE for a piecewise polynomial approximation to the trigonometric functions. Mathematical and Computational Applications, 23(4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:/

/www.mdpi.com/2297-8747/23/4/63.

Qiu:1996:SEC

[QV96] S.-L. Qiu and M. K. Vamanamurthy. Sharp estimates for complete elliptic integrals. SIAM journal on mathematical analysis, 27(3):823–834, May 1996. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Qiu:1998:SIG

[QVV98] S.-L. Qiu, M. K. Vamanamurthy, and M. Vuorinen. Some inequalities for the growth of elliptic integrals. SIAM journal on mathematical analysis, 29(5):1224–1237, September 1998. CO-DEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

URL http://epubs.siam.org/sam-bin/dbq/article/31049.

Qian:2006:HMP

[QW06] Jianbo Qian and Cao An Wang. How much precision is needed to compare two sums of square roots of integers? Information Processing Letters, 100(5):194–198, December 16, 2006. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Qi:2018:DME

[QXG18] Hongyuan Qi, Jinchen Xu, and Shaozhong Guo. Detection of the maximum error of mathematical functions. The Journal of Supercomputing, 74(11):6275–6290, November 2018. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Rack:1982:GIV

[Rac82] H.-J Rack. A generalization of an inequality of V. Markov to multivariate polynomials. Journal of Approximation Theory, 35(1):94–97, May 1982. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Raff:1970:CGF

[Raf70] Morton S. Raff. On calculating the gamma function of nonintegral arguments. The American Statistician, 24(2):22–24, April 1970. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2681926.

Rahavachary:2001:LSS

[Rah01] Saty Rahavachary. Letters: Setting the sqrt() record straight. Dr. Dobb’s Journal of Software Tools, 26(4):12, April 2001. CO-DEN DDJOEB. ISSN 1044-789X. URL http://www.ddj.com/.

Rainville:1960:SF

[Rai60] Earl David Rainville. Special Functions. MacMillan Publishing Company, New York, NY, USA, 1960. ISBN 0-8284-0258-2. xii + 365 pp. LCCN QA331 R15; QA351 .R3 1971. Reprinted in 1971.

Randazzo:1977:DFE

[Ran77] D. J. Randazzo. Data fits with exponential functions. In Numerical analysis (Proceedings of the Colloquium, Lausanne, 1976), volume 37 of Internat. Ser. Numer. Math., pages 77– 94. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1977. ISBN 3-7643-0939-3.

Rappoport:1994:TMC

[Rap94] Juri M. Rappoport. The Tau-Method and the computation of the Bessel functions of the complex order. In Brown et al. [BCEP94], pages 353–355. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Rappoport:2001:CVP

[Rap01] J. M. Rappoport. Canonical vector polynomials for the computation of complex order Bessel functions with the tau method. Computers and Mathematics with Applications, 41 (3–4):399–406, February 2001. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0898122100002820.

Rappoport:2014:MSM

[Rap14] Juri Rappoport. Mathematical software for modified Bessel functions. In Hong and Yap [HY14], pages 325–332. ISBN 3-662-44198-5 (paperback), 3-662-44199-3 (e-book). LCCN QA76.9.M35.

Rudnicki-Bujnowski:1975:EFC

[RB75] Georges Rudnicki-Bujnowski. Explicit formulas for Clebsch– Gordan coefficients. Computer Physics Communications, 10 (4):245–250, October 1975. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465575900697.

Ratis:1993:CCH

[RdC93] Yu. L. Ratis and P. Fernández de Córdoba. A code to calculate (high order) Bessel functions based on the continued fractions method. Computer Physics Communications, 76(3):381– 388, August 1993. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/001046559390062H.

Rusev:1998:TMS

[RDK98] Petur Rusev, I. Dimovski, and Virginia Kiryakova, editors. Transformation methods and special functions, Varna ’96: second international workshop: proceedings. Institute of Mathematics and Informatics, Bulgarian Academy of Sciences, Sofia, Bulgaria, 1998. ISBN 954-8986-05-1. LCCN ????

Rabi:2001:OCA

[RdL01] J. A. Rabi and M. J. S. de Lemos. Optimization of convergence acceleration in multigrid numerical solutions of conductiveconvective problems. Applied Mathematics and Computation, 124(2):215–226, October 25, 2001. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

elsevier.com/gej-ng/10/9/12/113/31/31/abstract.html.

Redish:1970:RAE

[Red70] K. A. Redish. Remark on Algorithm 385, Exponential Integral Ei(x). Communications of the ACM, 13(12):750, December 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Pac70].

Reemtsen:1990:MFR

[Ree90] Rembert Reemtsen. Modifications of the first Remez algorithm. SIAM Journal on Numerical Analysis, 27(2):507–518, April 1990. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Reichel:1969:IPV

[Rei69] Alex Reichel. The integral of the n th power of the Voigt function. Mathematics of Computation, 23(107):645–649, July 1969.

CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Reichel:1986:PAU

[Rei86] L. Reichel. On polynomial approximation in the uniform norm by the discrete least squares method. BIT (Nordisk tidskrift for informationsbehandling), 26(3):350–368, January 1986. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Relph:1963:AAH

[Rel63a] A. P. Relph. ACM Algorithm 191: Hypergeometric. Communications of the ACM, 6(7):388–389, July 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Kop74].

Relph:1963:AH

[Rel63b] A. P. Relph. Algorithm 191: Hypergeometric. Communications of the ACM, 6(7):388, July 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Kop74].

Relph:1963:ACH

[Rel63c] A. P. Relph. Algorithm 192: Confluent hypergeometric. Communications of the ACM, 6(7):388, July 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Revfeim:1990:LEM

[Rev90] K. J. A. Revfeim. Letter to the Editor: More approximations for the cumulative and inverse normal distribution. The American Statistician, 44(1):63, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.

jstor.org/stable/2684963.

Revy:2016:ADF

[Rev16] Guillaume Revy. Automated design of floating-point logarithm functions on integer processors. In Montuschi et al. [MSH^{+}16], pages 172–180. ISBN 1-5090-1615-5. ISSN 1063-6889. LCCN

QA76.9.C62 S95 2016. URL http://ieeexplore.ieee.org/
servlet/opac?punumber=7562813.

Ramamoorthy:1972:SPI

[RGK72] C. V. Ramamoorthy, James R. Goodman, and K. H. Kim. Some properties of iterative square-rooting methods using high-speed multiplication. IEEE Transactions on Computers, C-21(8):837– 847, August 1972. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5009039.

Rodriguez-Henriquez:2008:LCB

[RHMLL08] F. Rodriguez-Henriquez, G. Morales-Luna, and J. Lopez. Lowcomplexity bit-parallel square root computation over GF(2^{m}) for all trinomials. IEEE Transactions on Computers, 57(4):472– 480, April 2008. CODEN ITCOB4. ISSN 0018-9340 (print),

1557-9956 (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=4358282.

Rice:1965:CPR

[Ric65] John R. Rice. On the conditioning of polynomial and rational forms. Numerische Mathematik, 7(5):426–435, October 1965. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Richardson:1968:SUP

[Ric68] Daniel Richardson. Some undecidable problems involving elementary functions of a real variable. Journal of Symbolic Logic, 33(4):514–520, December 1968. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic).

Rice:1971:MS

[Ric71] John R. Rice. Mathematical Software. Academic Press, New York, NY, USA, 1971. ISBN 0-12-587250-X. xvii + 515 pp. LCCN QA1 .M26. Based on the proceedings of the Mathematical Software Symposium held at Purdue University, Lafayette, Indiana, USA, April 1–3, 1970.

Rice:1973:EEI

[Ric73] S. O. Rice. Efficient evaluation of integrals of analytic functions by the trapezoidal rule. The Bell System Technical Journal, 52(5):707–722, May–June 1973. CODEN BSTJAN. ISSN

0005-8580. URL http://bstj.bell-labs.com/BSTJ/images/
Vol52/bstj52-5-707.pdf.

Richardson:1988:NMT

[Ric88] Daniel Richardson. Nonstandard models of the theory of elementary functions of a real variable. Zeitschrift für mathematische Logik und Grundlagen der Mathematik, 34(4):355–372, 1988. CODEN ZMLGAQ. ISBN 0044-3050. ISSN 0044-3050.

Richards:1992:HFD

[Ric92] Donald St P. Richards, editor. Hypergeometric functions on domains of positivity, Jack polynomials, and applications: proceedings of an AMS Special Session held March 22–23, 1991 in Tampa, Florida, volume 138 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, 1992. ISBN 0-8218-5159-4. ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN QA353.H9 H97 1992.

Risch:1979:APE

[Ris79] Robert H. Risch. Algebraic properties of the elementary functions of analysis. American Journal of Mathematics, 101(4): 743–759, 1979. CODEN AJMAAN. ISSN 0002-9327 (print), 1080-6377 (electronic).

Ritt:1925:EFT

[Rit25] J. F. Ritt. Elementary functions and their inverses. Transactions of the American Mathematical Society, 27(1):68–90, 1925. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-

6850 (electronic). URL http://www.ams.org/journals/tran/
1925-027-01/S0002-9947-1925-1501299-9/.

Rix:1982:UQA

[Rix82] P. Rix. Universeller Quadratwurzel-Algorithmus [English: Universal Square Root Algorithms]. Elektronik, 23:81–82, 1982. CO-DEN EKRKAR. ISSN 0013-5658.

Raveendran:2020:NPF

[RJM^{+}20] Aneesh Raveendran, Sandra Jean, J. Mervin, D. Vivian, and David Selvakumar. A novel parametrized fused division and square-root POSIT arithmetic architecture. In IEEE, editor, 2020 33rd International Conference on VLSI Design and 2020 19th International Conference on Embedded Systems (VLSID), Bengaluru, India, 4–8 January 2020, pages 207–212. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, January 2020. ISBN 1-72815-701-3. ISSN 1063-9667 (print), 2380-6923 (electronic).

Ratnanather:2014:ATI

[RKZ^{+}14] J. Tilak Ratnanather, Jung H. Kim, Sirong Zhang, Anthony M. J. Davis, and Stephen K. Lucas. Algorithm 935: IIPBF, a MATLAB toolbox for infinite integral of products of two Bessel functions. ACM Transactions on Mathematical Software, 40(2): 14:1–14:12, February 2014. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Redding:1976:CPC

[RL76] R. W. Redding and W. P. Latham. On the calculation of the parabolic cylinder functions. II. The function V (a, x). Journal of Computational Physics, 20(2):256–258, February 1976. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999176900711.

Rengarajan:1980:MFI

[RL80] S. R. Rengarajan and J. E. Lewis. Mathieu functions of integral order and real arguments. IEEE Transactions on Microwave Theory and Techniques, 28(3):276–277, March 1980. CODEN IETMAB. ISSN 0018-9480 (print), 1557-9670 (electronic).

Ren:2007:CFA

[RM07] C. Ren and A. R. MacKenzie. Closed-form approximations to the error and complementary error functions and their applications in atmospheric science. Atmospheric Science Letters, 8(3):70–73, ???? 2007. ISSN 1530-261X. URL http://

onlinelibrary.wiley.com/doi/10.1002/asl.154/full.

Rao:1996:RTS

[RN96] V. M. Rao and B. Nowrouzian. Rounding techniques for signed binary arithmetic. In Canadian Conference on Electrical and Computer Engineering. 26–29 May 1996, volume 1, pages 294– 297. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. CODEN ???? ISSN ????

Robinson:1955:EAC

[Rob55] A. S. Robinson. An electronic analog computing technique for the solution of trigonometric problems. IRE Transactions on Electronic Computers, EC-4(3):95–101, ???? 1955. CODEN IRELAO. ISSN 0367-9950. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=5407907.

Robertson:1969:CNC

[Rob69] G. H. Robertson. Computation of the noncentral chi-square distribution. The Bell System Technical Journal, 48(1):201–207, January 1969. CODEN BSTJAN. ISSN 0005-8580. URL http:

//bstj.bell-labs.com/BSTJ/images/Vol48/bstj48-1-201.
pdf.

Rockey:1982:DMS

[Roc82] S. A. Rockey. Discrete methods of state approximation, parameter identification and optimal control for hereditary systems. Ph.D thesis, Brown University, Providence, RI, 1982. 208 pp.

Rolfe:1987:FIS

[Rol87] Timothy J. Rolfe. On a fast integer square root algorithm. ACM SIGNUM Newsletter, 22(4):6–11, October 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Ronning:1986:CTF

[Ron86] Gerd Ronning. On the curvature of the trigamma function. Journal of Computational and Applied Mathematics, 15(3):397– 399, July 1986. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0377042786902311.

Rothmaier:1970:BQN

[Rot70a] B. Rothmaier. Die Berechnung der Quadratwurzel nebst Schranken auf Dualmaschinen [English: The Computation of the Square Root together with [Interval] Bounds on Binary Machines]. Interner Bericht Nr. 70/17, Institut für Informatik, Universität Karlsruhe, 1970. ?? pp.

Rothmaier:1970:DSB

[Rot70b] B. Rothmaier. Dokumentation der Standardfunktionen des Betriebssystems Hydra X8 [English: Documentation of the elementary functions of the operating system Hydra X8]. Interner Bericht Nr. 70/8, Institut für Informatik, Universität Karlsruhe, Karlsruhe, Germany, 1970. ???? pp.

Rothmaier:1971:BEF

[Rot71] B. Rothmaier. Die Berechnung der elementaren Funktionen mit beliebiger Genauigkeit [English: The Computation of Elementary Functions with Arbitrary Accuracy]. Dissertation, Universität Karlsruhe, Karlsruhe, Germany, 1971. ???? pp.

Razaz:1981:RAF

[RS81] M. Razaz and J. L. Schonfelder. Remark on “Algorithm 498: Airy functions using Chebyshev series approximations”. ACM Transactions on Mathematical Software, 7(3):404–405, September 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Pri75].

Rivolo:1998:CDR

[RS98] M. T. Rivolo and A. Simi. Il calcolo delle radici quadrate e cubiche in Italia da Fibonacci a Bombelli. (Italian) [The calculation of square and cube roots in Italy from Fibonacci to Bombelli]. Archive for History of Exact Sciences, 52(2):161–193, February 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=
52&issue=2&spage=161.

Rader:1961:CAC

[RT61] P. J. Rader and Henry C. Thacher, Jr. Certification of Algorithm 14 [not 13]: Complex exponential integral. Communications of the ACM, 4(2):105, February 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rhee:1989:MII

[RT89] W. T. Rhee and M. Talagrand. Martingale inequalities, interpolation and NP-complete problems. Mathematics of Operations Research, 14(1):91–96, February 1989. CODEN MOREDQ. ISSN 0364-765x (print), 1526-5471 (electronic).

Rump:2001:RPS

[Rum01] Siegfried M. Rump. Rigorous and portable standard functions. BIT Numerical Mathematics, 41(3):540–562, June 2001. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0006-3835&volume=
41&issue=3&spage=540.

Ruscheweyh:1988:EST

[Rus88] Stephan Ruscheweyh. Extension of Szegő’s theorem on the sections of univalent functions. SIAM journal on mathematical analysis, 19(6):1442–1449, November 1988. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Russinoff:1998:MCP

[Rus98] David M. Russinoff. A mechanically checked proof of IEEE compliance of the floating point multiplication, division and square root algorithms of the AMD-K7 processor. LMS Journal of Computation and Mathematics, 1:148–200, 1998. CODEN

???? ISSN 1461-1570. URL http://www.lms.ac.uk/jcm/1/

lms1998-001/. Appendices A and B available to subscribers electronically (http://www.lms.ac.uk/jcm/1/lms98001/appendixa/ and http://www.lms.ac.uk/jcm/1/lms98001/appendix-b/).

Russinoff:1999:MCP

[Rus99] David M. Russinoff. A mechanically checked proof of correctness of the AMD K5 floating point square root microcode. Formal Methods in System Design, 14(1):75–125, January 1999. CODEN FMSDE6. ISSN 0925-9856 (print), 1572-8102 (elec-

tronic). URL http://www.wkap.nl/jrnltoc.htm/0925-9856;
 http://www.wkap.nl/oasis.htm/194808. Special issue on

arithmetic circuits.

Russinoff:2013:CFV

[Rus13] David M. Russinoff. Computation and formal verification of SRT quotient and square root digit selection tables. IEEE Transactions on Computers, 62(5):900–913, May 2013. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Rutishauser:1951:BAK

[Rut51] Heinz Rutishauser. Bemerkungen zur Arbeit von K. Emden “Eine Lösung für ∫ e^{b(x+a cos x)} dx”. (German) [Remarks on the work by K. Emden, “A solution for ∫ e^{b(x+a cosx)} dx”]. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 2(4):292–293, July 1951. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Rutishauser:1963:BQG

[Rut63] H. Rutishauser. Betrachtungen zur Quadratwurzeliteration. (German) [Considerations on square root iteration]. Monatshefte für Mathematik, 67:452–464, 1963. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic).

Ruymgaart:1989:SPB

[Ruy89] F. H. Ruymgaart. Some properties of bivariate empirical hazard processes under random censoring. Journal of Multivariate Analysis, 28(2):271–281, February 1989. CODEN JMVAAI. ISSN 0047-259x (print), 1095-7243 (electronic).

Rabin:1972:FEP

[RW72] Michael O. Rabin and Shmuel Winograd. Fast evaluation of polynomials by rational preparation. Communications on Pure and Applied Mathematics (New York), 25(4):433–458, July 1972. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Ralston:1960:MMD

[RW77] Anthony Ralston and Herbert S. Wilf, editors. Mathematical methods for digital computers. Wiley, New York, NY, USA, 1960–1977. ISBN 0-471-70690-6. various pp. LCCN QA39 .R26. Three volumes.

Raade:2004:MHS

[RW04] Lennart Råde and Bertil Westergren. Mathematics Handbook for Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fifth edition, 2004. ISBN 3-540-21141-1 (hardcover). 562 pp.

LCCN QA41 .R34 2004. URL http://www.loc.gov/catdir/
enhancements/fy0818/2006286513-d.html;   http://www.
loc.gov/catdir/toc/fy0704/2006286513.html.

Rokhlin:2007:AFC

[RX07] Vladimir Rokhlin and Hong Xiao. Approximate formulae for certain prolate spheroidal wave functions valid for large values of both order and band-limit. Applied and Computational Harmonic Analysis. Time-Frequency and Time-Scale Analysis, Wavelets, Numerical Algorithms, and Applications, 22(1):105– 123, January 2007. ISSN 1063-5203 (print), 1096-603x (electronic).

Rybicki:1989:DIS

[Ryb89] George B. Rybicki. Dawson’s integral and the sampling theorem. Computers in Physics, 3(2):85–87, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822832.

Rzadkowski:2012:SEE

[Rza̧12] Grzegorz Rza̧dkowski. On some expansions for the Euler gamma function and the Riemann zeta function. Journal of Computational and Applied Mathematics, 236(15):3710–3719, September 2012. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042711004663.

Stegun:1957:GBF

[SA57] Irene A. Stegun and Milton Abramowitz. Generation of Bessel functions on high speed computers. Mathematical Tables and Other Aids to Computation, 11(60):255–257, October 1957. CO-DEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Simon:2000:DCF

[SA00] Marvin Kenneth Simon and Mohamed-Slim Alouini. Digital Communication over Fading Channels: a Unified Approach to Performance Analysis. Wiley series in telecommunications and signal processing. Wiley, New York, NY, USA, 2000. ISBN 0-471-31779-9. xix + 544 pp. LCCN TK5103.7 .S523

2000. URL http://www.loc.gov/catdir/bios/wiley043/
99056352.html;   http://www.loc.gov/catdir/description/
wiley035/99056352.html;   http://www.loc.gov/catdir/
toc/onix06/99056352.html.

Simon:2005:DCF

[SA05] Marvin Kenneth Simon and Mohamed-Slim Alouini. Digital Communication over Fading Channels. Wiley series in telecommunications and signal processing. Wiley-Interscience, New York, NY, USA, second edition, 2005. ISBN 0-471-64953-8 (hardcover). xxxiv + 900 pp. LCCN TK5103.7 .S523 2005.

URL http://www.loc.gov/catdir/description/wiley042/
2004042040.html;   http://www.loc.gov/catdir/enhancements/
fy0617/2004042040-b.html;   http://www.loc.gov/catdir/
toc/wiley041/2004042040.html.

Saan:1991:VFP

[Saa91] T. Saan. Vychislenie èlementarnykh funktsiĭs pomoshch’yu drobno-ratsional’nykh priblizheniĭ. (Russian) [Calculation of elementary functions by means of rational approximations]. Èston. Nauchno-Proizvod. Ob”ed. Vychisl. Tekhn. Inform., Tartu, 1991. 139 pp.

Sablonniere:2008:BSH

[Sab08] Paul Sablonnière. B-splines and Hermite–Padé approximants to the exponential function. Journal of Computational and Applied Mathematics, 219(2):509–517, October 1, 2008. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037704270700252X.

Safouhi:2010:BSC

[Saf10] Hassan Safouhi. Bessel, sine and cosine functions and extrapolation methods for computing molecular multi-center integrals. Numerical Algorithms, 54(1):141–167, May 2010. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=1017-1398&volume=54&issue=1&spage=141.

Saigo:1989:FID

[Sai89] Megumi Saigo. Fractional integrals and derivatives associated with elementary functions and Bessel functions. In Srivastava and Owa [SO89], pages 283–306. ISBN 0-470-21630-1, 0-7458-0701-1. LCCN QA331 .U55 1989. UK£39.95.

Salzer:1951:FCE

[Sal51a] H. E. Salzer. Formulas for calculating the error function of a complex variable. Mathematical Tables and Other Aids to Computation, 5(34):67–70, April 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1951:RTT

[Sal51b] H. E. Salzer. Radix table for trigonometric functions and their inverses to high accuracy. Mathematical Tables and Other Aids to Computation, 5(33):9–11, January 1951. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Salzer:1953:RTO

[Sal53] Herbert E. Salzer. Radix table for obtaining hyperbolic and inverse hyperbolic functions to many places. Journal of Mathematics and Physics (MIT), 32(1–4):197–202, April 1953. CODEN JMPHA9. ISSN 0097-1421. URL https://onlinelibrary.

wiley.com/doi/epdf/10.1002/sapm1953321197.

Salzer:1955:CZE

[Sal55] Herbert E. Salzer. Complex zeros of the error function. Journal of The Franklin Institute, 260(3):209–211, September 1955. CO-DEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

Salamin:1976:CUA

[Sal76a] Eugene Salamin. Computation of π using arithmetic-geometric mean. Mathematics of Computation, 30(135):565–570, July 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Salimov:1976:OCE

[Sal76b] F. I. Salimov. The organization of calculations of elementary functions into tables. (Russian). In Probabilistic methods and cybernetics, volume 12–13, pages 77–90. Kazan University, Kazan, USSR, 1976.

Salzer:1983:NDG

[Sal83] Herbert E. Salzer. Note on the Dočev–Grosswald asymptotic series for generalized Bessel polynomials. Journal of Computational and Applied Mathematics, 9(2):131–135, June 1983. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0377042783900365. See errata [Ano84].

Sala:1989:TJA

[Sal89] Kenneth L. Sala. Transformations of the Jacobian amplitude function and its calculation via the arithmetic-geometric mean. SIAM journal on mathematical analysis, 20(6):1514– 1528, November 1989. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Salwin:1992:UPE

[Sal92] Arthur E. Salwin. Using the proposed elementary functions standard to build a strongly typed trig package. ACM SIGADA Ada Letters, 12(5):59–63, September/October 1992. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Samko:2002:HIT

[Sam02] S. G. (Stefan Grigorevich) Samko. Hypersingular Integrals and Their Applications, volume 5 of Analytical methods and special functions. Taylor and Francis, Boca Raton, FL, USA, 2002. ISBN 0-415-27268-8. xvii + 359 pp. LCCN QA403.5 .S26 2002.

Sarafyan:1959:NMC

[Sar59] Diran Sarafyan. A new method of computation of square roots without using division. Communications of the ACM, 2(11):

23–24, November 1959. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See comments [Tra60].

Sarafyan:1960:DCS

[Sar60] Diran Sarafyan. Divisionless computation of square roots through continued squaring. Communications of the ACM, 3 (5):319–321, May 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Saunders:1992:EFS

[Sau92] L. R. Saunders. An exact formula for the symmetrical incomplete beta function where the parameter is an integer or halfinteger. Australian Journal of Statistics, 34(2):261–264, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Saurer:1993:BSF

[Sau93] Josef Saurer. Bases of special functions and their domains of convergence, volume 73 of Mathematical research. Akademie Verlag GmbH, Berlin, Germany, 1993. ISBN 3-05-501613-0. ISSN 0138-3019. 158 pp. LCCN QA351 .S28 1993.

Sawada:2002:FVD

[Saw02] J. Sawada. Formal verification of divide and square root algorithms using series calculation. In Borrione [Bor02], pages 31–49. ISBN ???? LCCN ???? URL http://www.cs.utexas.

edu/users/moore/acl2/workshop-2002/.

Shenton:1971:CFP

[SB71] L. R. Shenton and K. O. Bowman. Continued fractions for the psi function and its derivatives. SIAM Journal on Applied Mathematics, 20(4):547–554, June 1971. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic). URL http://www.

jstor.org/stable/2099856.

Shipman:1971:HSE

[SC71] L. L. Shipman and R. E. Christoffersen. High speed evaluation of F_{0}(x). Computer Physics Communications, 2(4):201–206, May/June 1971. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465571900531.

Steinhardt:1981:ASF

[SC81] Paul J. Steinhardt and P. Chaudhari. Airy stress function for atomic models. Journal of Computational Physics, 42(2):266– 276, August 1981. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999181902448.

Siafarikas:2001:PFI

[SC01] Panayiotis D. Siafarikas and Theodore Seio Chihara, editors. Proceedings of the Fifth International Symposium on Orthogonal Polynomials, Special Functions and their Applications: Patras, Greece, 20–24 September 1999, volume 133(1/2) of Journal of computational and applied mathematics. Elsevier, Amsterdam, The Netherlands, 2001. ISSN 0377-0427 (print), 1879-1778 (electronic). LCCN QA76 J86 v. 133, no. 1/2.

Scarton:1971:DPF

[Sca71] Henry A. Scarton. Double precision FORTRAN subroutines to compute both ordinary and modified Bessel functions of the first kind and of integer order with arbitrary complex argument: J_{n}(x + jy) and I_{n}(x + jy). Journal of Computational Physics, 8(2):295–299, October 1971. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0021999171900106.

Schmidt:1964:AEC

[Sch64] Paul W. Schmidt. Asymptotic expansion of certain integrals containing the Bessel function J_{0}(x). Journal of Mathematical Physics, 5(8):1183–1184, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i8/p1183_s1.

Schmidt:1968:AEK

[Sch68] Jochen W. Schmidt. Asymptotische Einschließung bei konvergenzbeschleunigenden Verfahren. II. (German) [Asymptotic enclosure with convergence acceleration method. II]. Numerische Mathematik, 11(1):53–56, January 1968. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Schmid:1973:BLVa

[Sch73] H. Schmid. BCD logic V: BCD square root. Electronic Design, 21(17):62–69, August 1973. CODEN ELODAW. ISSN 0013-4872 (print), 1944-9550 (electronic).

Schett:1976:PTS

[Sch76a] Alois Schett. Properties of the Taylor series expansion coefficients of the Jacobian elliptic functions. Mathematics of Computation, 30(133):143–147, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schonfelder:1976:PSF

[Sch76b] J. L. Schonfelder. The production of special function routines for a multi-machine library. Software—Practice and Experience, 6 (1):71–82, January/March 1976. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schett:1977:RFT

[Sch77a] Alois Schett. Recurrence formula of the Taylor series expansion coefficients of the Jacobian elliptic functions. Mathematics of Computation, 31(140):1003–1005, October 1977. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schonfelder:1977:PTS

[Sch77b] J. L. Schonfelder. The production and testing of special function software in the NAG library. In Cowell [Cow77], pages 425–451. ISBN 0-387-08446-0. LCCN QA297 .W65 1976.

Schmidt:1978:EI

[Sch78a] Paul W. Schmidt. Evaluation of the integral ∫_{0}^{∞} ^{\frac{t^{2α−1}J_{ν}(x\sqrt{1+t^{2}})(1+t^{2})^{α+β−1}}{d}} t. Mathematics of Computation, 32(141):265–269, January 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schoene:1978:RMI

[Sch78b] Andrew Y. Schoene. Remark on “Algorithm 435: Modified incomplete gamma function [S14]”. ACM Transactions on Mathematical Software, 4(3):296–304, September 1978. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Ful72].

Schonfelder:1978:CEE

[Sch78c] J. L. Schonfelder. Chebyshev expansions for the error and related functions. Mathematics of Computation, 32(144):1232– 1240, October 1978. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schell:1980:AEU

[Sch80a] Hans-Joachim Schell. Asymptotische Entwicklungen für die unvollständige Gammafunktion. (German) [Asymptotic developments for the incomplete gamma function]. Wissenschaftliche Zeitschrift der Technischen Hochschule Karl-Marx-Stadt, 22(5): 477 485, ???? 1980.

Schonfelder:1980:VHA

[Sch80b] J. L. Schonfelder. Very high accuracy Chebyshev expansions for the basic trigonometric functions. Mathematics of Computation, 34(149):237–244, January 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schmidt:1984:TAI

[Sch84] J. W. Schmidt. Two-sided approximations of inverses, square roots and Cholesky factors. Comput. Math., Banach Center Publ., 13:483–497, 1984.

Schoof:1985:ECF

[Sch85] René Schoof. Elliptic curves over finite fields and the computation of square roots modp. Mathematics of Computation, 44 (170):483–494, April 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schappacher:1988:EIG

[Sch88a] Norbert Schappacher. Elliptic integrals and the gamma function. Lecture Notes in Mathematics, 1301:117–127, 1988. CODEN LNMAA2. ISBN 3-540-18915-7 (print), 3-540-38842-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082098/.

Schwarz:1988:CLI

[Sch88b] Jerry Schwarz. A C++ library for infinite precision floating point. In USENIX Association [USE88], pages 271–281.

Schwarz:1993:HRAa

[Sch93a] E. Schwarz. High-radix algorithms for high-order arithmetic operations. Technical Report CSL-TR-93-559, Computer Systems Laboratory, Stanford University, Stanford, CA, USA, January 1993.

Schwarz:1993:HRAb

[Sch93b] Eric Mark Schwarz. High-radix algorithms for high-order arithmetic operations. Thesis (Ph.D.), Department of Electrical Engineering, Stanford University, Stanford, CA, USA, April 1993. 224 pp.

Schwarz:1995:RQC

[Sch95] E. M. Schwarz. Rounding for quadratically converging algorithms for division and square root. In Singh [Sin95], pages 600–603. CODEN ???? ISBN 0-8186-7370-2. ISSN ???? LCCN TK7801 .A83 1995. Two volumes.

Schatzman:1996:ADF

[Sch96] James C. Schatzman. Accuracy of the discrete Fourier transform and the fast Fourier transform. SIAM Journal on Scientific Computing, 17(5):1150–1166, September 1996. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL

http://epubs.siam.org/sam-bin/dbq/article/24702.

Schraudolph:1999:FCA

[Sch99] N. N. Schraudolph. A fast, compact approximation of the exponential function. Neural Computation, 11(4):853–862, May 1, 1999. CODEN NEUCEB. ISSN 0899-7667 (print), 1530-888x

(electronic). URL https://www.mitpressjournals.org/doi/
abs/10.1162/089976699300016467.

Schwalm:2015:EFE

[Sch15] William A. Schwalm. Elliptic Functions and Elliptic Integrals. IOP concise physics. Morgan and Claypool Publishers and IOP Publishing, San Rafael, CA, USA and Bristol, UK, 2015. ISBN 1-68174-166-0 (print), 1-68174-230-6 (e-book), 1-68174-102-4 (mobi). ISSN 2054-7307. 67 pp. LCCN QA343 .S355 2015. URL

http://iopscience.iop.org/book/978-1-6817-4230-4.

Schmidt:2016:ZSG

[Sch16] Maxie D. Schmidt. Zeta series generating function transformations related to generalized Stirling numbers and partial sums of the Hurwitz zeta function. arxiv.org, November 2016. URL

http://adsabs.harvard.edu/abs/2016arXiv161100957S.

Sreedharan:1985:ASS

[SD85] J. Sreedharan and A. Dhurkadas. 8086 algorithm solves square roots. EDN, 30(7):272, April 1985. CODEN EDNSBH. ISSN 0012-7515, 0364-6637.

Segura:1997:CEM

[SdCR97] J. Segura, P. Fernández de Córdoba, and Yu. L. Ratis. A code to evaluate modified Bessel functions based on the continued fraction method. Computer Physics Communications, 105(2–3):263–272, October 1, 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465597000696.

Strollo:2011:EFH

[SDP11] Antonio Giuseppe Maria Strollo, Davide De Caro, and Nicola Petra. Elementary functions hardware implementation using constrained piecewise-polynomial approximations. IEEE Transactions on Computers, 60:418–432, 2011. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Schost:2011:IPI

[SE11] Éric Schost and Ioannis Z. Emiris, editors. ISSAC 2011: Proceedings of the 2011 International Symposium on Symbolic and Algebraic Computation, June 7–11, 2011, San Jose, CA, USA. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-0675-6. LCCN QA76.95 .I59 2011.

Soranzo:2014:VSE

[SE14] Alessandro Soranzo and Emanuela Epure. Very simply explicitly invertible approximations of normal cumulative and normal quantile function. Applied Mathematical Sciences (Ruse), 8:4323–4341, 2014. ISSN 1312-885X (print), 1314-7552 (elec-

tronic). URL http://m-hikari.com/ams/ams-2014/ams-85-
88-2014/epureAMS85-88-2014.pdf.

Seaton:1984:CFA

[Sea84] M. J. Seaton. Coulomb functions analytic in the energy. Computer Physics Communications, 35(1–3):C–771, ???? 1984.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465584828994.

Sedogbo:1990:CAS

[Sed90] Guy Antoine Sedogbo. Convergence acceleration of some logarithmic sequences. Journal of Computational and Applied Mathematics, 32(1–2):253–260, November 26, 1990. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Extrapolation and rational approximation (Luminy, 1989).

Segura:2008:IZC

[Seg08] Javier Segura. Interlacing of the zeros of contiguous hypergeometric functions. Numerical Algorithms, 49(1–4):387–407, December 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=1017-1398&volume=
49&issue=1&spage=387.

Seidel:1999:HSR

[Sei99] P.-M. Seidel. High-speed redundant reciprocal approximation. Integration, the VLSI journal, 28(1):1–12, September 1999. CO-DEN IVJODL. ISSN 0167-9260 (print), 1872-7522 (electronic).

Sellers:1993:CDC

[Sel93] H. Sellers. The C^{2}-DIIS convergence acceleration algorithm. International Journal of Quantum Chemistry, 45(1):31–??, 1993. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sen:1967:TSF

[Sen67] Bibhutibhusan Sen. A treatise on special functions, for scientists and engineers. Allied Publishers, Bombay, India, 1967. 164 pp. LCCN QA351 .S45.

Sherry:1959:CGF

[SF59] M. E. Sherry and S. Fulda. Calculation of gamma functions to high accuracy (in Technical Notes and Short Papers). Mathematical Tables and Other Aids to Computation, 13(68):314–315, October 1959. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Smith:1979:ALL

[SF79] David A. Smith and William F. Ford. Acceleration of linear and logarithmic convergence. SIAM Journal on Numerical Analysis, 16(2):223–240, April 1979. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Stoyanov:1987:AE

[SF87] Basil J. Stoyanov and Richard A. Farrell. On the asymptotic evaluation of ∫_{0}^{π/2} J_{0}^{2}(γ sin x)dx. Mathematics of Computation, 49(179):275–279, July 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schwarz:1993:HSA

[SF93a] E. M. Schwarz and M. J. Flynn. Hardware starting approximation for the square root operation. In Swartzlander, Jr. et al. [SIJ93], pages 103–111. CODEN ???? ISBN 0-7803-1401-8 (softbound), 0-8186-3862-1 (casebound), 0-8186-3861-3 (microfiche). ISSN ???? LCCN QA 76.9 C62 S95 1993. IEEE Transactions on Computers 43(8), 1994.

Schwarz:1993:UFM

[SF93b] Eric Mark Schwarz and M. J. (Michael J.) Flynn. Using a floating-point multiplier’s internals for high-radix division and square root. Technical report CSL-TR-93-554, Computer Systems Laboratory, Stanford University, Stanford, CA, USA, 1993. iv + 45 pp.

Schwarz:1996:HSA

[SF96] Eric M. Schwarz and Michael J. Flynn. Hardware starting approximation method and its application to the square root operation. IEEE Transactions on Computers, 45(12):1356–1369, December 1996. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Sofotasios:2010:NEM

[SF10] Paschalis C. Sofotasios and Steven Freear. Novel expressions for the Marcum and one dimensional Q-functions. In 2010 7th International Symposium on Wireless Communication Systems. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 2010. ISBN 1-4244-6315-7 (print), 1-4244-6317-3 (e-book), 1-4244-6316-5 (CD-ROM). ISSN 2154-0217 (print), 2154-0225 (electronic). URL

http://ieeexplore.ieee.org/document/5624374/.

Sayed:2016:WCR

[SF16] Wafaa S. Sayed and Hossam A. H. Fahmy. What are the correct results for the special values of the operands of the power operation? ACM Transactions on Mathematical Software, 42 (2):14:1–14:17, June 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Strecok:1972:HPE

[SG72] A. J. Strecok and J. A. Gregory. High precision evaluation of the irregular Coulomb wave functions. Mathematics of Computation, 26(120):955–961 + s1–s10, October 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schulten:1976:REC

[SG76] K. Schulten and R. G. Gordon. Recursive evaluation of 3j and 6j coefficients. Computer Physics Communications, 11(2):269–278, March/May 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Schulte:1992:AHD

[Sg92] Michael Joseph Schulte and Function generation. Algorithms and hardware designs for parallel elementary function generation. Thesis (M.S. in engin.), University of Texas at Austin, Austin, TX, USA, 1992. ix + 73 pp.

Segura:1998:PCF

[SG98] J. Segura and A. Gil. Parabolic cylinder functions of integer and half-integer orders for nonnegative arguments. Computer Physics Communications, 115(1):69–86, December 1, 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000976.

Segura:1999:EGT

[SG99] J. Segura and A. Gil. ELF and GNOME: Two tiny codes to evaluate the real zeros of the Bessel functions of the first kind for real orders. Computer Physics Communications, 117(3):250–262, March 11, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465598001933.

Sawada:2002:MVS

[SG02a] Jun Sawada and Ruben Gamboa. Mechanical verification of a square root algorithm using Taylor’s theorem. Lecture Notes in Computer Science, 2517:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.de/link/service/series/0558/bibs/2517/
25170274.htm; http://link.springer.de/link/service/series/
0558/papers/2517/25170274.pdf.

Sebah:2002:IGF

[SG02b] Pascal Sebah and Xavier Gourdon. Introduction to the gamma function. World-Wide Web document, February

4, 2002.  URL http://numbers.computation.free.fr/
Constants/constants.html;   http://numbers.computation.
free.fr/Constants/Miscellaneous/gammaFunction.ps.

Schulten:1981:NAE

[SGA81] Z. Schulten, R. G. Gordon, and D. G. M. Anderson. A numerical algorithm for the evaluation of Weber parabolic cylinder functions U(a, x), V (a, x), and W (a, ±x). Journal of Computational Physics, 42(2):213–237, August 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0021999181902412.

Samet:1972:CDL

[SH72] P. A. Samet and D. W. Honey. Calculation of a double-length square root from double-length number using single precision techniques. The Computer Journal, 15(2):116, May 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/2/
116.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/116.tif.

Shanks:1976:TER

[Sha76] D. Shanks. Table errata: “Regular continued fractions for π and γ”, (Math. Comp. 25 (1971), 403); “Rational approximations to π” (ibid. 25 (1971), 387–392) by K. Y. Choong, D. E. Daykin and C. R. Rathbone. Mathematics of Computation, 30(134): 381, 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL http://www.ams.org/journals/mcom/
1976-30-134/S0025-5718-1976-0386215-4.

Shah:1985:SAA

[Sha85] Arvind K. Shah. A simpler approximation for areas under the standard normal curve. The American Statistician, 39(1):80, February 1985. CODEN ASTAAJ. ISSN 0003-1305 (print),

1537-2731 (electronic). URL http://www.jstor.org/stable/
2683918.

Shenton:1954:INI

[She54] L. R. Shenton. Inequalities for the normal integral including a new continued fraction. Biometrika, 41(1/2):177–189, June 1954. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2333015.

Sheorey:1974:CEW

[She74] V. B. Sheorey. Chebyshev expansions for wavefunctions. Computer Physics Communications, 7(1):1–12, January 1974. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Sheorey:1976:DCE

[She76] V. B. Sheorey. Double Chebyshev expansions for wave functions. Computer Physics Communications, 12(2):125–134, November 1976. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465576900618.

Shishkov:1993:RDB

[Shi93] Dimit”r Shishkov. Reduction of domains of basic elementary functions to arbitrary small intervals. Annuaire Univ. Sofia Fac. Math. Inform., 87(1–2):3–32 (1999), 1993. ISSN 0205-0808.

Sholander:1960:AEE

[Sho60] Marlow Sholander. Analytical expressions and elementary functions. American Mathematical Monthly, 67(3):213–214, March 1960. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/2309678.

Shore:1986:AID

[Sho86] Haim Shore. An approximation for the inverse distribution function of a combination of random variables, with an application to operating theatres. Journal of Statistical Computation and Simulation, 23(3):157–181, 1986. CODEN JSCSAJ. ISSN 0094-9655 (print), 1563-5163 (electronic).

Shore:2002:RMM

[Sho02] Haim Shore. Response modeling methodology (RMM)-exploring the properties of the implied error distribution. Communications in Statistics: Theory and Methods, 31(12):2225–2249, 2002. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shore:2005:ARB

[Sho05] Haim Shore. Accurate RMM-based approximations for the CDF of the normal distribution. Communications in Statistics: Theory and Methods, 34(3):507–513, 2005. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shterenlikht:2019:QIF

[Sht19] A. Shterenlikht. On quality of implementation of Fortran 2008 complex intrinsic functions on branch cuts. ACM Transactions on Mathematical Software, 45(1):11:1–11:9, March 2019. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL https://dl.acm.org/citation.cfm?id=3301318.

Sidhu:1995:EIF

[Sid95] Satinder S. Sidhu. Elliptic integrals and functions. Computers in Physics, 9(3):268–276, May/June 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Sidi:2003:PEM

[Sid03] Avram Sidi. Practical Extrapolation Methods: Theory and Applications, volume 10 of Cambridge monographs on applied and computational mathematics. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-66159-5 (hardcover), 0-511-54681-5 (e-book), 0-511-06018-1 (Adobe Reader), 0-511-06649-X (ebook). xxii + 519 pp. LCCN QA281 .S555 2003.

Sidi:2006:CTC

[Sid06] Avram Sidi. A challenging test for convergence accelerators: Summation of a series with a special sign pattern. App. Math. E-Notes, 6(??):225–234, ???? 2006. CODEN ???? ISSN ????

Siemieniuch:1976:PCR

[Sie76] J. L. Siemieniuch. Properties of certain rational approximations to e^{−z}. BIT (Nordisk tidskrift for informationsbehandling), 16(2):172–191, June 1976. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=16&issue=2&spage=172.

Swartzlander:1993:SCA

[SIJ93] Earl Swartzlander, Jr., Mary Jane Irwin, and Graham Jullien, editors. Proceedings: 11th Symposium on Computer Arithmetic, June 29–July 2, 1993, Windsor, Ontario. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1401-8 (softbound), 0-8186-3862-1 (casebound), 0-8186-3861-3 (microfiche). ISSN 0018-9340 (print), 1557-9956 (electronic). LCCN QA 76.9 C62 S95 1993. IEEE Transactions on Computers 43(8), 1994.

Simauti:1964:AFS

[Sim64] Takakazu Simauti. Approximation formulas for some elementary functions. Comment. Math. Univ. St. Paul., 12:23–35, 1964. CODEN COMAAC. ISSN 0010-258X.

Singh:1995:CRT

[Sin95] Avtar Singh, editor. Conference record of the Twenty-Ninth Asilomar Conference on Signals, Systems & Computers: October 30–November 1, 1995 Pacific Grove, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7370-2. LCCN TK7801 .A83 1995. Two volumes.

Sinclair:1996:ORS

[Sin96] R. Sinclair. Optimization of reciprocals and square roots on the i860 microprocessor. International Journal of High Speed Computing, 8(1):57–64, 1996. CODEN IHSCEZ. ISSN 0129-0533.

Srivastava:2001:SAZ

[SJ01] H. M. Srivastava and Choi Junesang. Series Associated with the Zeta and Related Functions. Kluwer Academic Publishers, Dordrecht, The Netherlands, 2001. ISBN 0-7923-7054-6. ix + 388 pp. LCCN QA351 .S74 2001. URL http://

www.loc.gov/catdir/enhancements/fy0822/2001035764-d.
html;   http://www.loc.gov/catdir/enhancements/fy0822/
2001035764-t.html.

Srivastava:2012:ZZF

[SJ12] H. M. Srivastava and Choi Junesang. Zeta and q-Zeta functions and associated series and integrals. Elsevier Insights. Elsevier, Amsterdam, The Netherlands, 2012. ISBN 0-12-385218-8. xvi + 657 pp. LCCN QA351 .S745 2012. URL http://www.

sciencedirect.com/science/book/9780123852182.

Sarkar:1971:EPP

[SK71] B. P. Sarkar and E. V. Krishnamurthy. Economic pseudodivision processes for obtaining square root, logarithm, and arctan. IEEE Transactions on Computers, C-20(12):1589–1593, December 1971. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Schreier:2008:OIR

[SK08] Franz Schreier and Dieter Kohlert. Optimized implementations of rational approximations — a case study on the Voigt and complex error function. Computer Physics Communications, 179(7):457–465, October 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465508001495.

Shi:2011:AEA

[SK11] Qinghua Shi and Y. Karasawa. An accurate and efficient approximation to the Gaussian Q-function and its applications in performance analysis in Nakagami-m fading. IEEE Communications Letters, 15(5):479–481, May 2011. CODEN ICLEF6. ISSN 1089-7798 (print), 1558-2558 (electronic). URL http://

ieeexplore.ieee.org/document/5740503/.

Skeel:1986:CVS

[Ske86a] Robert D. Skeel. Construction of variable-stepsize multistep formulas. Mathematics of Computation, 47(176):503–510, S45–S52, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Skeel:1986:SCV

[Ske86b] Robert D. Skeel. Supplement to construction of variable-stepsize multistep formulas. Mathematics of Computation, 47(176):S45– S52, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sincovec:1993:PSS

[SKL93] Richard F. Sincovec, David E. Keyes, and Michael R. Leuze, editors. Proceedings of the Sixth SIAM Conference on Parallel Processing for Scientific Computing, held March 22–24, 1993, in Norfolk, VA, USA. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993. ISBN 0-89871-315-3. LCCN QA76.58 .S55 1993 v.1-2. Two volumes.

Skovgaard:1975:RBF

[Sko75a] Ove Skovgaard. Remark on “Algorithm 236: Bessel Functions of the First Kind [S17]”. ACM Transactions on Mathematical Software, 1(3):282–284, September 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Gau64a].

Skovgaard:1975:RAJ

[Sko75b] Ove Skovgaard. Remark on “Algorithm 332: Jacobi polynomials”. Communications of the ACM, 18(2):116–117, 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Wit68].

Skovgaard:1978:RCE

[Sko78] Ove Skovgaard. Remark on “Algorithm 149: Complete Elliptic Integral [S21]”. ACM Transactions on Mathematical Software, 4 (1):95, March 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Mer62].

Skorokhodov:2004:STP

[Sko04] S. L. Skorokhodov. Symbolic transformations in the problem of analytic continuation of the hypergeometric function _{p}F_{p−1}(z) in the neighborhood of the point z = 1 in the logarithmic case. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 30(3):150–156, ???? 2004. CODEN PCSODA. ISSN 0361-7688 (print), 1608-3261 (electronic).

Skorokhodov:2005:MCG

[Sko05] S. L. Skorokhodov. A method for computing generalized hypergeometric function _{p}F_{p−1}(a_{1}, …, a_{p}; b_{1}, …, b_{p−1}; 1) in terms of the Riemann zeta function. Computational Mathematics and Mathematical Physics, 45(4):550–562, ???? 2005. CODEN ???? ISSN 0965-5425 (print), 1555-6662 (electronic).

Stern:1963:CSR

[SL63] T. E. Stern and R. M. Lerner. A circuit for the square root of the sum of the squares. Proceedings of the IEEE, 51(4):593– 596, April 1963. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Shepherd:1981:CA

[SL81] M. M. Shepherd and J. G. Laframboise. Chebyshev approximation of (1 + 2x) exp(x^{2})(x) in 0 ≤ x < ∞. Mathematics of Computation, 36(153):249–253, January 1981. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Sidi:1983:ZSP

[SL83] Avram Sidi and Doron S. Lubinsky. On the zeros of some polynomials that arise in numerical quadrature and convergence acceleration. SIAM Journal on Numerical Analysis, 20(2):400– 405, April 1983. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Soderquist:1995:APC

[SL95] Peter Soderquist and Miriam Leeser. An area/performance comparison of subtractive and multiplicative divide/square root implementations. In Knowles and McAllister [KM95], pages 132–139. ISBN 0-8186-7089-4 (paperback), 0-8186-7089-4 (case), 0-8186-7149-1 (microfiche), 0-8186-7089-4 (softbound), 0-7803-2949-X (casebound). LCCN QA 76.9 C62

S95 1995. URL http://www.acsel-lab.com/arithmetic/
arith12/papers/ARITH12_Soderquist.pdf.

Soderquist:1997:DSR

[SL97] Peter Soderquist and Miriam Leeser. Division and square root: Choosing the right implementation: Exploring the major design choices for microprocessor implementations of floating-point division and square root. IEEE Micro, 17(4):56–66, July/August 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143

(electronic). URL http://pascal.computer.org/mi/books/
mi1997/pdf/m4056.pdf.

Slavjanov:2007:SFU

[SL07] Sergej J. Slavjanov and Wolfgang Lay. Special Functions: a Unified Theory Based on Singularities. Oxford mathematical monographs; Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-850573-6. xvi + 293 pp. LCCN ????

Slater:1960:CHF

[Sla60] Lucy Joan Slater. Confluent hypergeometric functions. Cambridge University Press, Cambridge, UK, 1960. 247 pp. LCCN QA351 .S56.

Slater:1964:CHF

[Sla64] Lucy Joan Slater. Confluent hvpergeometric functions. In Abramowitz and Stegun [AS64], pages 503–536. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also avail-

able online at http://www.convertit.com/Go/ConvertIt/

Reference/AMS55.ASP in bitmap image format.

Slater:1966:GHF

[Sla66] Lucy Joan Slater. Generalized Hypergeometric Functions. Cambridge University Press, Cambridge, UK, 1966. xiii + 273 pp. LCCN QA351 .S565.

Stange:2016:CAM

[SLD16] J. Stange, N. Loginova, and T. Dickhaus. Computing and approximating multivariate chi-square probabilities. Journal of Statistical Computation and Simulation, 86(6):1233–1247, 2016. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (elec-

tronic), 1563-5163. URL http://www.tandfonline.com/doi/
abs/10.1080/00949655.2015.1058798.

Slepian:1964:PSW

[Sle64] David Slepian. Prolate spheroidal wave functions, Fourier analysis and uncertainty — IV: Extensions to many dimensions; generalized prolate spheroidal functions. The Bell System Technical Journal, 43(6):3009–3057, November 1964. CODEN BST-

JAN. ISSN 0005-8580. URL http://bstj.bell-labs.com/
BSTJ/images/Vol43/bstj43-6-3009.pdf.

Slepian:1965:SAE

[Sle65] David Slepian. Some asymptotic expansions for prolate spheroidal wave functions. J. Math. and Physics XLIV(2), ??(??):99–140, June 1965. URL http://www.ams.org/

mathscinet/search/publications.html?pg1=IID&s1=189661.

Slepian:1978:PSW

[Sle78] D. Slepian. Prolate spheroidal wave functions, Fourier analysis, and uncertainty — V: The discrete case. The Bell System Technical Journal, 57(5):1371–1430, May–June 1978. CODEN BST-

JAN. ISSN 0005-8580. URL http://bstj.bell-labs.com/
BSTJ/images/Vol57/bstj57-5-1371.pdf.

Shi:2006:NAS

[SLH06] Xiquan Shi, Fengshan Liu, and Minghan Hu. A new asymptotic series for the gamma function. Journal of Computational and Applied Mathematics, 195(1–2):134–154, October 15, 2006. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042705004802.

Stehle:2005:SWC

[SLZ05] Damien Stehlé, Vincent Lefèvre, and Paul Zimmermann. Searching worst cases of a one-variable function using lattice reduction. IEEE Transactions on Computers, 54(3):340–346, March 2005. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://csdl.computer.org/comp/trans/tc/
2005/03/t0340abs.htm;   http://csdl.computer.org/dl/
trans/tc/2005/03/t0340.htm;   http://csdl.computer.org/
dl/trans/tc/2005/03/t0340.pdf;   http://ieeexplore.ieee.
org/iel5/12/30205/01388198.pdf?isnumber=30205&prod=JNL&
arnumber=1388198&arSt=+340&ared=+346&arAuthor=Stehle%
2C+D.%3B+Lefevre%2C+V.%3B+Zimmermann%2C+P.;   http:/
/ieeexplore.ieee.org/xpls/abs_all.jsp?isnumber=30205&
arnumber=1388198&count=13&index=9;   http://ieeexplore.
ieee.org/xpls/references.jsp?arnumber=1388198.

Smith:1970:ASH

[SM70] Robert R. Smith and Dennis McCall. Algorithm 392: Systems of hyperbolic P.D.E. Communications of the ACM, 13(9):567–570, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Fri72].

Schindler:1977:CCS

[SM77] Susan Schindler and R. Mirman. The Clebsch–Gordan coefficients of S_{n}. Journal of Mathematical Physics, 18(8):1697–1704, August 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schindler:1978:GCG

[SM78] Susan Schindler and R. Mirman. Generation of the Clebsch-Gordan coefficients for S_{n}. Computer Physics Communications, 15(1–2):131–145, September 1978. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Stratton:1941:ECS

[SMCH41] Julius Adams Stratton, Philip M. (Philip McCord) Morse, Lan Jen Chu, and Reina Albagli Hutner. Elliptic cylinder and spheroidal wave functions, including tables of separation constants and coefficients. Wiley, New York, NY, USA, 1941. xii + 127 pp. LCCN QC174.2 .S78.

Skaf:1994:LHI

[SMG94] Ali Skaf, Jean-Michel Muller, and Alain Guyot. On-line hardware implementation for complex exponential and logarithm. In Anonymous, editor, ESSCIRC ’94: Twentieth European Solid-State Circuits Conference. Ulm, Germany. September 20–22, 1994, pages 252–255. Éditions Frontières, B. P. 33. 91192 Gifsur-Yvette Cedex, France, 1994. ISBN 2-86332-160-9.

Smith:1989:EMP

[Smi89] David M. Smith. Efficient multiple-precision evaluation of elementary functions. Mathematics of Computation, 52(185):131– 134, January 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Smith:1991:AFP

[Smi91] David M. Smith. Algorithm 693: A FORTRAN package for floating-point multiple-precision arithmetic. ACM Transactions on Mathematical Software, 17(2):273–283, June 1991. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
toms/108585.html.

Smith:1995:CFA

[Smi95] Roger Alan Smith. A continued-fraction analysis of trigonometric argument reduction. IEEE Transactions on Computers, 44 (11):1348–1351, November 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Smith:2001:AFS

[Smi01] David M. Smith. Algorithm 814: Fortran 90 software for floating-point multiple precision arithmetic, gamma and related functions. ACM Transactions on Mathematical Software, 27(4): 377–387, December 2001. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Smith:2011:AMP

[Smi11] David M. Smith. Algorithm 911: Multiple-precision exponential integral and related functions. ACM Transactions on Mathematical Software, 37(4):46:1–46:16, February 2011. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Sneddon:1956:SFM

[Sne56] Ian Naismith Sneddon. Special Functions of Mathematical Physics and Chemistry, volume 19 of University mathematical texts. Oliver and Boyd, Edinburgh, UK, third edition, 1956. viii + 164 pp. LCCN QA1 U588 v. 19.

Sneddon:1961:SFM

[Sne61] Ian Naismith Sneddon. Special Functions of Mathematical Physics and Chemistry. University mathematical texts. Oliver and Boyd, Edinburgh, UK, ssecond edition, 1961. 184 pp. LCCN QA331 .S65 1961.

Sneddon:1963:SFM

[Sne63] Ian Naismith Sneddon. Spezielle Funktionen der mathematischen Physik und Chemie. Mathematische Formelsammlung II. (German) [Special functions of mathematical physics and chemistry. Mathematical formula collection II], volume 54 of B. I.- Hochschultaschenbücher. Bibliographisches Institut, Mannheim, Germany, 1963. 166 pp. LCCN QA351 .S6415.

Sneddon:1980:SFM

[Sne80] Ian Naismith Sneddon. Special Functions of Mathematical Physics and Chemistry. Longman mathematical texts. Longman, London, UK, third edition, 1980. ISBN 0-582-44396-2 (paperback). ix + 182 pp. LCCN QA351 .S64 1980.

Snyder:1993:AFI

[Sny93] W. Van Snyder. Algorithm 723: Fresnel integrals. ACM Transactions on Mathematical Software, 19(4):452–456, December 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295

(electronic). URL http://www.acm.org/pubs/citations/
journals/toms/1993-19-4/p452-van_snyder/. See remarks

[Sny96, Sny21].

Snyder:1996:RAF

[Sny96] W. Van Snyder. Remark on Algorithm 723: Fresnel integrals. ACM Transactions on Mathematical Software, 22(4):498–500, December 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [Sny93].

Snyder:2021:CRA

[Sny21] W. Van Snyder. Corrigendum: Remark on Algorithm 723: Fresnel Integrals. ACM Transactions on Mathematical Software, 47(4):37:1, December 2021. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://dl.acm.

org/doi/10.1145/3452336. See [Sny93].

Spanier:1987:AF

[SO87] Jerome Spanier and Keith B. Oldham. An Atlas of Functions. Hemisphere Publishing Corp., Washington, DC, USA, 1987. ISBN 0-89116-573-8, 3-540-17395-1. ix + 700 pp. LCCN QA331 .S685 1987. See also the second edition [OMS09].

Srivastava:1989:UFF

[SO89] H. M. Srivastava and Shigeyoshi Owa, editors. Univalent functions, fractional calculus, and their applications (Kōriyama, 1988). Ellis Horwood, New York, NY, USA, 1989. ISBN 0-470-21630-1, 0-7458-0701-1. 404 pp. LCCN QA331 .U55 1989. UK£39.95.

Schwarz:2011:PIS

[SO11] Eric Schwarz and Vojin G. Oklobdzija, editors. Proceedings of the 20th IEEE Symposium on Computer Arithmetic, July 25– 27, 2011, Tübingen, Germany. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 0-7695-4318-9, 1-4244-9457-5. LCCN QA76.6.

Sobczyk:1988:SMA

[Sob88] Kazimierz Sobczyk. Stochastic modelling and analysis of fatigue. In W. Schiehlen and W. Wedig, editors, Analysis and estimation of stochastic mechanical systems, volume 303 of Courses and lectures, pages 269–313. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-82058-2. LCCN TA350.3.I5 no.303; TJ173 .A531 1988.

Sollfrey:1969:IFP

[Sol69] William Sollfrey. Inverse functions of the products of two Bessel functions. Journal of Mathematical Physics, 10(8):1429–1430, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i8/p1429_s1.

Smith:1981:ERA

[SOL81] J. M. Smith, F. W. J. Olver, and D. W. Lozier. Extended-range arithmetic and normalized Legendre polynomials. ACM Transactions on Mathematical Software, 7(1):93–105, March 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://doi.acm.org/10.1145/355934.355940.

Sommer:1982:EPL

[Som82] M. Sommer. Existence of pointwise-Lipschitz-continuous selections of the metric projection for a class of Z-spaces. Journal of Approximation Theory, 34(2):115–130, February 1982. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Sookne:1973:BFC

[Soo73a] D. J. Sookne. Bessel functions I and J of complex argument and integer order. Journal of Research of the National Bureau of Standards (1934), 77B(3–4):111–114, July 1973. ISSN 0091-0635.

Sookne:1973:BFR

[Soo73b] D. J. Sookne. Bessel functions of real argument and integer order. Journal of Research of the National Bureau of Standards (1934), 77A(3–4):125–132, July 1973. ISSN 0091-0635.

Sookne:1973:CABb

[Soo73c] D. J. Sookne. Certification of an algorithm for Bessel functions of complex argument. Journal of Research of the National Bureau of Standards (1934), 77B(3):133–136, July 1973. ISSN 0091-0635.

Sookne:1973:CABa

[Soo73d] D. J. Sookne. Certification of an algorithm for Bessel functions of real argument. Journal of Research of the National Bureau of Standards (1934), 77B(3–4):115–124, July 1973. ISSN 0091-0635.

Sorenson:1994:TFG

[Sor94] J. Sorenson. Two fast GCD algorithms. Journal of Algorithms, 16(1):110–144, January 1994. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677484710066.

Southard:1957:ATW

[Sou57] Thomas H. Southard. Approximation and table of the Weierstrass ℘ function in the equianharmonic case for real argument. Mathematical Tables and Other Aids to Computation, 11(58):99– 100, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print),

2326-4853 (electronic). URL http://www.jstor.org/stable/
2002156.

Southard:1964:WER

[Sou64] Thomas H. Southard. Weierstrass elliptic and related functions. In Abramowitz and Stegun [AS64], pages 627–684. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also avail-

able online at http://www.convertit.com/Go/ConvertIt/

Reference/AMS55.ASP in bitmap image format.

Slepian:1961:PSW

[SP61] D. Slepian and H. O. Pollak. Prolate spheroidal wave functions, Fourier analysis and uncertainty. I. The Bell System Technical Journal, 40(1):43–63, January 1961. CODEN BSTJAN. ISSN

0005-8580. URL http://bstj.bell-labs.com/BSTJ/images/
Vol40/bstj40-1-43.pdf.

Shibata:2020:SPV

[SP20] Naoki Shibata and Francesco Petrogalli. SLEEF: A portable vectorized library of C standard mathematical functions. IEEE Transactions on Parallel and Distributed Systems, 31(6):1316– 1327, June 2020. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Spellucci:1971:DPA

[Spe72] P. Spellucci. Double precision approximations to the elementary functions using Jacobi-fractions. Numerische Mathematik, 18:127–143, 1971/1972. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Spielberg:1961:ECF

[Spi61] Kurt Spielberg. Efficient continued fraction approximations to elementary functions. Mathematics of Computation, 15(76):409– 417, October 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spira:1971:CGF

[Spi71] Robert Spira. Calculation of the gamma function by Stirling’s formula. Mathematics of Computation, 25(114):317–322, April 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spijker:1985:SRS

[Spi85] M. N. Spijker. Stepsize restrictions for stability of one-step methods in the numerical solution of initial value problems. Mathematics of Computation, 45(172):377–392, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Spouge:1994:CGD

[Spo94] John L. Spouge. Computation of the gamma, digamma, and trigamma functions. SIAM Journal on Numerical Analysis, 31 (3):931–944, June 1994. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Sprague:2000:PAH

[Spr00] Ralph H. Sprague, editor. Proceedings of the 33rd Annual Hawaii International Conference on System Sciences: abstracts and CD-ROM of full papers: January 4–7, 2000, Maui, Hawaii. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7695-0493-0, 0-7695-0494-9, 0-7695-0495-7. LCCN TA168 .H37 2000; TA168 .H37 2000xeb; TA168.

Squire:1970:RAI

[Squ70] William Squire. A rational approximation to an integral appearing in glow curve theory. Journal of Computational Physics, 6 (1):152–253, August 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Squire:1991:PSG

[Squ91a] J. S. Squire. Proposed standard for a generic package of complex elementary functions (Ada). ACM SIGADA Ada Letters, 11(7): 140–165, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Squire:1991:TVG

[Squ91b] J. S. Squire. Towards validation of generic elementary functions and other standard Ada numerics packages. ACM SIGADA Ada Letters, 11(7):217–243, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Squire:1991:ANS

[Squ91c] Jon S. Squire. Ada numerics standardization and testing. ACM SIGADA Ada Letters, 11(7):1–286, 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Squire:1991:RPS

[Squ91d] Jon S. Squire. Rationale for the proposed standard for a generic package of complex elementary functions (Ada). ACM SIGADA Ada Letters, 11(7):166–179, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Scherberg:1953:ACP

[SR53] Max G. Scherberg and John F. Riordan. Analogue calculation of polynomial and trigonometric expansions (in Other Aids to Computation). Mathematical Tables and Other Aids to Computation, 7(41):61–65, January 1953. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Steinberg:1981:LSE

[SR81] D. Steinberg and M. Rodeh. A layout for the shuffle-exchange network with O(N^{2}/ log^{3/2N}) area. IEEE Transactions on Computers, C-30(12):977–982, December 1981. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Shukla:2014:LLH

[SR14] R. Shukla and K. C. Ray. Low latency hybrid CORDIC algorithm. IEEE Transactions on Computers, 63(12):3066–3078, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Stefanica:2016:SAA

[SR16] Dan Stefanica and Radoš Radoičić. A sharp approximation for ATM-forward option prices and implied volatilities. International Journal of Financial Engineering, 3(1):1650002, March 2016. ISSN 2424-7863.

Srivastava:1970:CRI

[Sri70] H. M. Srivastava. Certain results involving generalized hypergeometric functions. SIAM journal on mathematical analysis, 1 (1):75–81, February 1970. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Srinivasan:2007:GFE

[Sri07] Gopala Krishna Srinivasan. The gamma function: An eclectic tour. American Mathematical Monthly, 114(4):297–315, April 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/27642193.

Slepian:1965:EAP

[SS65] David Slepian and Estelle Sonnenblick. Eigenvalues associated with prolate spheroidal wave functions of zero order. The Bell System Technical Journal, 44(8):1745–1759, October 1965. CO-DEN BSTJAN. ISSN 0005-8580. URL http://bstj.bell-

labs.com/BSTJ/images/Vol44/bstj44-8-1745.pdf.

Spain:1970:FMP

[SS70] Barry Spain and M. G. (Michael Gambier) Smith. Functions of Mathematical Physics. The New university mathematics series. Van Nostrand Reinhold Co., New York, NY, USA, 1970. ISBN 0-442-07871-4, 0-442-07876-5 (hardcover). xi + 208 pp.

LCCN QA351 .S69. URL http://books.google.com/books?
id=kYgZAQAAIAAJ.

Smith:1987:BAM

[SS87] P. W. Smith and J. J. Swetits. Best approximation by monotone functions. Journal of Approximation Theory, 49(4):398– 403, April 1987. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Schulte:1993:ERC

[SS93a] M. Schulte and E. Swartzlander. Exact rounding of certain elementary functions. In Swartzlander, Jr. et al. [SIJ93], pages 138–145. ISBN 0-7803-1401-8 (softbound), 0-8186-3862-1 (casebound), 0-8186-3861-3 (microfiche). ISSN 0018-9340 (print), 1557-9956 (electronic). LCCN QA 76.9 C62 S95

1993. URL http://mesa.ece.wisc.edu/publications/cp_

1993-01.pdf. IEEE Transactions on Computers 43(8), 1994.

Schulte:1993:PHD

[SS93b] Michael J. Schulte and Earl E. Swartzlander, Jr. Parallel hardware designs for correctly rounded elementary functions. Interval Computations = Interval’nye vychisleniia, 4:65–88, 1993. ISSN 0135-4868. Proceedings of the International Conference on Numerical Analysis with Automatic Result Verification (Lafayette, LA, 1993).

Schulte:1994:HDE

[SS94] M. J. Schulte and E. E. Swartzlander, Jr. Hardware design for exactly rounded elementary functions. IEEE Transactions on Computers, 43(8):964–973, August 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Schulte:1997:SBT

[SS97a] M. Schulte and J. Stine. Symmetric bipartite tables for accurate function approximation. In Lang et al. [LMT97], pages 175–183. ISBN 0-8186-7846-1, 0-8186-7847-X, 0-8186-7848-8. ISSN 1063-6889. LCCN QA76.9.C62 S95 1997. URL http:

//mesa.ece.wisc.edu/publications/cp_1997-01.pdf. IEEE

Computer Society order number PR07846. IEEE Order Plan catalog number 97CB36091.

Schulte:1997:AFA

[SS97b] M. J. Schulte and James E. Stine. Accurate function approximations by symmetric table lookup and addition. In Thiele et al. [T^{+}97], pages 144–153. ISBN 0-8186-7959-X, 0-8186-7960-3, 0-8186-7958-1. LCCN TK7874.6 .I57 1997eb; TK7874.6 .I57

1997; TK7874.6 .I58 1997. URL http://mesa.ece.wisc.edu/
publications/cp_1997-02.pdf.

Sidi:1998:UBC

[SS98] Avram Sidi and Yair Shapira. Upper bounds for convergence rates of acceleration methods with initial iterations. Numerical Algorithms, 18(2):113–132, September 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-

9265 (electronic). URL http://ipsapp007.kluweronline.

com/content/getfile/5058/14/1/abstract.htm;   http:/
/ipsapp007.kluweronline.com/content/getfile/5058/14/
1/fulltext.pdf.

Schulte:1999:AEF

[SS99a] M. Schulte and J. Stine. Approximating elementary functions with symmetric bipartite tables. IEEE Transactions on Computers, 48(8):842–847, 1999. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://mesa.ece.wisc.

edu/publications/cp_1999-10.pdf.

Stine:1999:STA

[SS99b] E. Stine and M. J. Schulte. The symmetric table addition method for accurate function approximation. Journal of VLSI Signal Processing, 21(2):167–177, June 1999. CODEN JVSPED. ISSN 0922-5773 (print), 1573-109x (electronic). URL http://

mesa.ece.wisc.edu/publications/cp_1999-11.pdf.

Slevinsky:2010:RAT

[SS10] Richard M. Slevinsky and Hassan Safouhi. A recursive algorithm for the G transformation and accurate computation of incomplete Bessel functions. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1411–1417, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Slevinsky:2023:RAE

[SS23] Richard M. Slevinsky and Hassan Safouhi. A recursive algorithm for an efficient and accurate computation of incomplete Bessel functions. Numerical Algorithms, 92(1):973–983, January 2023. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/10.
1007/s11075-022-01438-0.

Schneider:2018:NFP

[SSG^{+}18] Barry I. Schneider, Javier Segura, Amparo Gil, Xiaoxu Guan, and Klaus Bartschat. A new Fortran 90 program to compute regular and irregular associated Legendre functions (new version announcement). Computer Physics Communications, 225(??):192–193, April 2018. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465517304186.

Shaw:1974:NME

[ST74] Mary Shaw and J. F. Traub. On the number of multiplications for the evaluation of a polynomial and some of its derivatives. Journal of the Association for Computing Machinery, 21(1):161– 167, January 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Story:1999:NAI

[ST99] S. Story and P. T. P. Tang. New algorithms for improved transcendental functions on IA-64. In Koren and Kornerup [KK99], pages 4–11. ISBN 0-7803-5609-8, 0-7695-0116-8, 0-7695-0118-4. ISSN 1063-6889. LCCN QA76.6 .S887

1999.  URL http://euler.ecs.umass.edu/paper/final/
paper-118.pdf;   http://euler.ecs.umass.edu/paper/final/

paper-118.ps. IEEE Computer Society Order Number PR00116. IEEE Order Plan Catalog Number 99CB36336.

Schmelzer:2007:CGF

[ST07] Thomas Schmelzer and Lloyd N. Trefethen. Computing the gamma function using contour integrals and rational approximations. SIAM Journal on Numerical Analysis, 45(2):558–571, ???? 2007. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Staunton:2017:PP

[Sta17] Mike Staunton. Power to Pólya. Wilmott Magazine, 90:36–37,

July 2017. URL http://onlinelibrary.wiley.com/doi/10.
1002/wilm.10605/full.

Stegun:1964:LF

[Ste64a] Irene A. Stegun. Legendre functions. In Abramowitz and Stegun [AS64], pages 331–354. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://www.convertit.com/

Go/ConvertIt/Reference/AMS55.ASP in bitmap image format.

Stegun:1964:MF

[Ste64b] Irene A. Stegun. Miscellaneous functions. In Abramowitz and Stegun [AS64], pages 997–1010. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available online at http://

www.convertit.com/Go/ConvertIt/Reference/AMS55.ASP in

bitmap image format.

Stearns:1989:SFD

[Ste89] C. C. Stearns. Subtractive floating-point division and square root for VLSI DSP. In IEE [IEE89a], pages 405–409. ISBN 0-85296-383-1. LCCN ???? Conference publication no. 308.

Steele:2008:FPSb

[Ste08] Guy L. Steele Jr. Floating point square root provider with embedded status information. US Patent 7430576, September 30,

2008. URL http://www.patentstorm.us/patents/7430576/
fulltext.html.

Stiefel:1953:ITF

[Sti53] E. Stiefel. Zur Interpolation von tabellierten Funktionen durch Exponentialsummen und zur Berechnung von Eigenwerten aus den Schwarzschen Konstanten. (German) [On interpolation of tabulated functions by exponential sums and on the calculation of eigenvalues from the Schwarz’s constants]. Zeitschrift für Angewandte Mathematik und Mechanik, 33:260–262, 1953. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Stiefel:1959:NMT

[Sti59] Eduard L. Stiefel. Numerical methods of Tchebycheff approximation. In Langer [Lan59], pages 217–232. LCCN QA3 .U45 no. 1. Publication no. 1 of the Mathematics Research Center, U.S. Army, the University of Wisconsin.

Stillinger:1988:CPS

[Sti88] Frank H. Stillinger. Collective phenomena in statistical mechanics and the geometry of potential energy hypersurfaces. In Donald G. Truhlar, editor, Mathematical frontiers in computational chemical physics, volume 15 of The IMA volumes in mathematics and its applications, pages 157–173. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96782-6. LCCN QD455.3.M3 M38 1988. US$36.80. Proceedings of the Workshop on Atomic and Molecular Structure and Dynamics, held June 15–July 24, 1987, at the Institute for Mathematics and Its Applications, University of Minnesota.

Stoner:1941:FEF

[Sto41] Paul Matthew Stoner. Fitting the exponential function and the Gompertz function by the method of least squares. Journal of the American Statistical Association, 36(216):515–518, December 1941. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2278959.

Stratton:1956:SWF

[Str56] Julius Adams Stratton. Spheroidal Wave Functions, Including Tables of Separation Constants and Coefficients. Wiley, New York, NY, USA, 1956. xiii + 613 pp. LCCN QA405 .S8.

Strachey:1959:TSR

[Str59] C. Strachey. On taking the square root of a complex number. The Computer Journal, 2(2):89, July 1959. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/020089.sgm.abs.html;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
02/tiff/89.tif.

Strecok:1968:CIE

[Str68] Anthony J. Strecok. On the calculation of the inverse of the error function. Mathematics of Computation, 22(101):144–158, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL http://www.jstor.org/stable/
2004772.

Suetin:1999:OPT

[Sue99] P. K. (Pavel Kondratevich) Suetin. Orthogonal Polynomials in Two Variables, volume 3 of Analytical methods and special functions. Gordon and Breach Science Publishers, Amsterdam, The Netherlands, 1999. ISBN 90-5699-167-1. xx + 348 pp. LCCN QA404.5 .S8813 1999. Translated from the Russian by E. V. Pankratiev.

Sugai:1958:ERR

[Sug58] Iwao Sugai. Extraction of roots by repeated subtractions for digital computers. Communications of the ACM, 1(12):6–8, December 1958. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sundblad:1971:AFT

[Sun71] Y. Sundblad. The Ackermann function. A theoretical, computational, and formula manipulative study. BIT (Nordisk tidskrift for informationsbehandling), 11(1):107–119, March 1971. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0006-3835&volume=
11&issue=1&spage=107.

Sun:1988:SAF

[Sun88] Xiehua Sun. On the simultaneous approximation of functions and their derivatives by the Szasz–Mirakyan operator. Journal of Approximation Theory, 55(3):279–288, December 1988. CO-DEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Sun:2015:LEG

[Sun15] Qiming Sun. Libcint: an efficient general integral library for Gaussian basis functions. Journal of Computational Chemistry, 36(22):1664–1671, August 15, 2015. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Stanley:1950:TRG

[SW50] John Pearson Stanley and Maurice V. Wilkes. Table of the reciprocal of the gamma function for complex argument. Report, Computation Centre, University of Toronto, Toronto, ON, Canada, 1950. ???? pp.

Stephens:1988:SCR

[SW88] A. J. Stephens and H. C. Williams. Some computational results on a problem concerning powerful numbers. Mathematics of Computation, 50(182):619–632, April 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Schulte:1999:ESO

[SW99a] M. J. Schulte and K. E. Wires. Efficient second order approximations for reciprocals and square roots. In Luk [Luk99], pages 10–18. ISBN 0-8194-3293-8. LCCN TK5102.5 .A3325 1999; TK5102.5 .A3173 1999eb; TK5102.9 .A37 1999; TK5102.5;

TS510 .S63. URL http://mesa.ece.wisc.edu/publications/
cp_1999-05.pdf.

Schulte:1999:HSI

[SW99b] Michael J. Schulte and Kent E. Wires. High-speed inverse square roots. In Koren and Kornerup [KK99], pages 124–131. CODEN ???? ISBN 0-7803-5609-8, 0-7695-0116-8, 0-7695-0118-4. ISSN ???? LCCN QA76.6 .S887 1999. URL http://euler.ecs.

umass.edu/paper/final/paper-109.pdf;   http://euler.
ecs.umass.edu/paper/final/paper-109.ps. IEEE Computer

Society Order Number PR00116. IEEE Order Plan Catalog Number 99CB36336.

Swarztrauber:1965:LED

[Swa65] P. N. Swarztrauber. Letter to the Editor: On the doubleprecision square root routine. Communications of the ACM, 8 (4):202, April 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Swartzlander:1990:CA

[Swa90] Earl E. Swartzlander. Computer arithmetic. IEEE Computer Society Press tutorial. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-8931-5 (v. 1), 0-8186-5931-9 (v. 1 microfiche). various pp. LCCN QA76.6.C633 1990. Two volumes.

Sweeney:1963:CEC

[Swe63] Dura W. Sweeney. On the computation of Euler’s constant. Mathematics of Computation, 17(82):170–178, April 1963. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Stegun:1970:ACM

[SZ70] I. A. Stegun and R. Zucker. Automatic computing methods for special functions. part 1. Error, probability, and related functions. Journal of Research of the National Bureau of Standards (1934), 74B(3):211–224, July 1970. ISSN 0091-0635.

Stegun:1974:ACM

[SZ74] I. A. Stegun and R. Zucker. Automatic computing methods for special functions. Part II. The exponential integral E_{n}(x). Journal of Research of the National Bureau of Standards (1934), 78B(4):199–216, October 1974. ISSN 0091-0635.

Stegun:1976:ACM

[SZ76] I. A. Stegun and R. Zucker. Automatic computing methods for special functions. Part III. The sine, cosine, exponential integrals, and related functions. Journal of Research of the National Bureau of Standards (1934), 80B(2):291–311, April 1976. ISSN 0091-0635.

Sharma:1977:GMA

[SZ77] R. R. Sharma and Bahman Zohuri. A general method for an accurate evaluation of exponential integrals E_{1}(x), x > 0. Journal of Computational Physics, 25(2):199–204, October 1977. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999177900225.

Stehle:2005:GAT

[SZ05] Damien Stehlé and Paul Zimmermann. Gal’s accurate tables method revisited. In IEEE [IEE05], page ?? ISBN ???? LCCN

???? URL http://arith17.polito.it/final/paper-152.
pdf.

Szmytkowski:2013:EBT

[Szm13] Radosław Szmytkowski. Erratum to Formulas and Theorems for the Special Functions of Mathematical Physics by W. Magnus, F. Oberhettinger, R. P. Soni. Mathematics of Computation, 82 (283):1709–1710, ???? 2013. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.ams.

org/journals/mcom/2013-82-283/S0025-5718-2013-02671-
3;   http://www.ams.org/journals/mcom/2013-82-283/S0025-
5718-2013-02671-3/S0025-5718-2013-02671-3.pdf.

Srivastava:2011:ADZ

[SZW11] H. M. Srivastava, Jian-Rong Zhou, and Zhi-Gang Wang. Asymptotic distributions of the zeros of certain classes of hypergeometric functions and polynomials. Mathematics of Computation, 80(275):1769–1784, July 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/journals/mcom/2011-80-
275/S0025-5718-2011-02409-9/home.html;   http://www.
ams.org/journals/mcom/2011-80-275/S0025-5718-2011-02409-
9/S0025-5718-2011-02409-9.pdf.

T:1957:CHS

[T.57] C. B. T. Comment: T. H. Southard, Approximation and table of the Weierstrass ℘ function in the equianharmonic case for real argument. [MTAC, this issue, p. 99–100]. Mathematical Tables and Other Aids to Computation, 11(58):110, April 1957. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853

(electronic). URL http://www.jstor.org/stable/2002157.

Thiele:1997:IIC

[T^{+}97] Lothar Thiele et al., editors. IEEE International Conference on Application-Specific Systems, Architectures and Processors: proceedings, July 14–16, 1997, Zürich, Switzerland. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7959-X, 0-8186-7960-3, 0-8186-7958-1. LCCN TK7874.6 .I57 1997eb; TK7874.6 .I57 1997; TK7874.6 .I58 1997.

Takenaga:1966:EIG

[Tak66] Roy Takenaga. On the evaluation of the incomplete gamma function (in Technical Notes and Short Papers). Mathematics of Computation, 20(96):606–610, October 1966. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Takahashi:2000:IMP

[Tak00] D. Takahashi. Implementation of multiple-precision parallel division and square root on distributed-memory parallel computers. In Proceedings of the 2000 International Workshops on Parallel Processing, pages 229–235. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. CODEN ???? ISSN ????

Takagi:2001:HAC

[Tak01] N. Takagi. A hardware algorithm for computing reciprocal square root. In Burgess and Ciminiera [BC01], pages 94–100. CODEN ???? ISBN 0-7695-1150-3; 0-7695-1152-X. ISSN ???? LCCN QA76.9.C62 S95 2001. US$145. IEEE order no. PR01150.

Takekawa:2022:FPC

[Tak22] Takashi Takekawa. Fast parallel calculation of modified Bessel function of the second kind and its derivatives. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN

2352-7110. URL http://www.sciencedirect.com/science/
article/pii/S2352711021001655.

Talman:1968:SFG

[Tal68] James D. Talman. Special Functions: a Group Theoretic Approach Based on Lectures by Eugene P. Wigner. The mathematical physics monograph series. W. A. Benjamin, Inc., New York, NY, USA, 1968. xii + 260 pp. LCCN ???? With an introduction by Eugene P. Wigner.

Talman:1983:LSC

[Tal83] James D. Talman. LSFBTR: a subroutine for calculating spherical Bessel transforms. Computer Physics Communications, 30(1):93–99, July/August 1983. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465583901261.

Talman:1984:LSC

[Tal84] James D. Talman. LSFBTR: a subroutine for calculating spherical Bessel transforms. Computer Physics Communications, 35(1–3):C–903, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465584830027.

Talman:2009:NSC

[Tal09] J. D. Talman. NumSBT: a subroutine for calculating spherical Bessel transforms numerically. Computer Physics Communications, 180(2):332–338, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508003329.

Tanese:1987:PGA

[Tan87] Reiko Tanese. Parallel genetic algorithms for a hypercube. In John J. Grefenstette, editor, Genetic algorithms and their applications: proceedings of the second International Conference on Genetic Algorithms: July 28–31, 1987 at the Massachusetts Institute of Technology, Cambridge, MA, pages 177–183. Lawrence Erlbaum Associates, Inc., Hillsdale, NJ, USA, 1987. ISBN 0-8058-0158-8, 0-8058-0159-6 (paperback). LCCN Q334 .I5561 1987. US$39.95.

Tang:1988:PIG

[Tan88] Ping Tak Peter Tang. Portable implementation of a generic exponential function. Technical report ANL-88-3, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1988.

Tang:1989:TDI

[Tan89a] Ping Tak Peter Tang. Table-driven implementation of the exponential function in IEEE floating-point arithmetic. ACM Transactions on Mathematical Software, 15(2):144– 157, June 1989. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://doi.acm.

org/10.1145/63522.214389;   http://www.acm.org/pubs/
citations/journals/toms/1989-15-2/p144-tang/.

Tang:1989:TCA

[Tan89b] Ping Tak Peter Tang. Testing computer arithmetic by elementary number theory. Technical report, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, IL, USA, August 1989. ???? pp.

Tang:1990:AET

[Tan90a] Ping Tak Peter Tang. Accurate and efficient testing of the exponential and logarithm functions. ACM Transactions on Mathematical Software, 16(3):185–200, September 1990. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://doi.acm.org/10.1145/79505.79506;   http:/
/www.acm.org/pubs/citations/journals/toms/1990-16-3/
p185-tang/.

Tang:1990:FAL

[Tan90b] Ping Tak Peter Tang. A fast algorithm for linear complex Chebyshev approximation. In Mason and Cox [MC90], pages 265–274. ISBN 0-412-34580-3. LCCN QA221 .I54 1988.

Tang:1990:PGE

[Tan90c] Ping Tak Peter Tang. A portable generic elementary function package in Ada and an accurate test suite. Technical report ANL-90/35, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, November 1990. iii +

35 pp. URL http://www.osti.gov/bridge/servlets/purl/
6310184-4n5sOR/6310184.PDF.

Tang:1990:SSI

[Tan90d] Ping Tak Peter Tang. Some software implementations of the functions sine and cosine. Technical report ANL-90/3, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, April 1990. URL http://www-fp.mcs.anl.

gov/division/publications/abstracts/abstracts90.htm.

Tang:1990:TDI

[Tan90e] Ping Tak Peter Tang. Table-driven implementation of the logarithm function in IEEE floating-point arithmetic. ACM Transactions on Mathematical Software, 16(4):378–400, December 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-

7295 (electronic). URL http://doi.acm.org/10.1145/98267.
98294;   http://www.acm.org/pubs/citations/journals/
toms/1990-16-4/p378-tang/.

Tang:1991:PGE

[Tan91a] Ping Tak Peter Tang. A portable generic elementary function package in Ada and an accurate test suite. ACM SIGADA Ada Letters, 11(7):181–216, Fall 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Tang:1991:TLA

[Tan91b] Ping Tak Peter Tang. Table-lookup algorithms for elementary functions and their error analysis. In Kornerup and Matula [KM91], pages 232–236. ISBN 0-8186-9151-4 (case), 0-8186-6151-8 (microfiche), 0-7803-0187-0 (library binding). LCCN QA76.9.C62 S95 1991.

Tang:1992:TDI

[Tan92] Ping Tak Peter Tang. Table-driven implementation of the Expm1 function in IEEE floating-point arithmetic. ACM Transactions on Mathematical Software, 18(2):211–222, June 1992. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org/pubs/citations/journals/toms/

1992-18-2/p211-tang/. See independent analysis and accuracy confirmation of this algorithm in [Kra98].

Taylor:1981:CHD

[Tay81] George S. Taylor. Compatible hardware for division and square root. In IEEE CA5 ’81 [IEE81b], pages 127–134. LCCN QA 76.6

S985t 1981. URL http://www.acsel-lab.com/arithmetic/
arith5/papers/ARITH5_Taylor.pdf. IEEE catalog number

81CH1630-C.

Takagi:1991:RCM

[TAY91] N. Takagi, T. Asada, and S. Yajima. Redundant CORDIC methods with a constant scale factor for sine and cosine computation. IEEE Transactions on Computers, 40(9):989–995, September 1991. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-

9956 (electronic). URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=83660.

Thompson:1986:CBF

[TB86] I. J. Thompson and A. R. Barnett. Coulomb and Bessel functions of complex arguments and order. Journal of Computational Physics, 64(2):490–509, June 1986. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/002199918690046X.

Thompson:1987:MBF

[TB87] I. J. Thompson and A. R. Barnett. Modified Bessel functions I_{ν}(z) and K_{ν}(z) of real order and complex argument, to selected accuracy. Computer Physics Communications, 47(2–3):245–257, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0010465587901111. See erratum [Tho04].

Tang:2011:TCT

[TBDS11] Ping Tak Peter Tang, J. Adam Butts, Ron O. Dror, and David E. Shaw. Tight certification techniques for digit-by-rounding algorithms with application to a new 1/\sqrt{x} design. In Schwarz and Oklobdzija [SO11], pages 159–168. ISBN 0-7695-4318-9, 1-4244-

9457-5. LCCN QA76.6. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=5992121.

Takagi:2019:ISC

[TBL19] Naofumi Takagi, Sylvie Boldo, and Martin Langhammer, editors. 2019 IEEE 26th Symposium on Computer Arithmetic ARITH-26 (2019), Kyoto, Japan, 10–12 June 2019. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 2019. ISBN 1-72813-366-1. ISSN 1063-6889.

Thorsley:2001:AEH

[TC01] Michael D. Thorsley and Marita C. Chidichimo. An asymptotic expansion for the hypergeometric function _{2}F_{1}(a, b; c; x). Journal of Mathematical Physics, 42(4):1921–1930, April 2001. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thakkar:2006:PDP

[TE06] Anuja J. Thakkar and Abdel Ejnioui. Pipelining of double precision floating point division and square root operations. In Menezes [Men06], pages 488–493. ISBN 1-59593-315-8 (print). LCCN QA75.5 A184 2006 E.

Temme:1975:NEM

[Tem75a] N. M. Temme. On the numerical evaluation of the modified Bessel function of the third kind. Journal of Computational Physics, 19(3):324–337, November 1975. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Temme:1975:UAE

[Tem75b] N. M. Temme. Uniform asymptotic expansions of the incomplete gamma functions and the incomplete beta function. Mathematics of Computation, 29(132):1109–1114, October 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Temme:1976:NEO

[Tem76] Nico M. Temme. On the numerical evaluation of the ordinary Bessel function of the second kind. Journal of Computational Physics, 21(3):343–350, July 1976. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0021999176900322.

Temme:1978:UAE

[Tem78] N. M. Temme. Uniform asymptotic expansions of confluent hypergeometric functions. Journal of the Institute of Mathematics and its Applications, 22(2):215–223, 1978. CODEN JMTAA8. ISSN 0020-2932.

Temme:1979:AAP

[Tem79a] N. M. Temme. An algorithm with ALGOL 60 program for the computation of the zeros of ordinary Bessel functions and those of their derivatives. Journal of Computational Physics, 32(2): 270–279, August 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Temme:1979:AEI

[Tem79b] N. M. Temme. The asymptotic expansion of the incomplete gamma functions. SIAM journal on mathematical analysis, 10 (4):757–766, July 1979. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Temme:1981:ECH

[Tem81] N. M. Temme. On the expansion of confluent hypergeometric functions in terms of Bessel functions. Journal of Computational and Applied Mathematics, 7(1):27–32, March 1981. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0771050X81900048.

Temme:1983:NCC

[Tem83a] N. M. Temme. The numerical computation of the confluent hypergeometric function U(a, b, z). Numerische Mathematik, 41(1):63–82, April 1983. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Temme:1983:TTR

[Tem83b] N. M. Temme. Traces to Tricomi in recent work on special functions and asymptotics of integrals. Report TW 239/83, Stichting mathematisch Centrum, Amsterdam, The Netherlands, 1983. 15 pp.

Temme:1983:UAE

[Tem83c] Nico M. Temme. Uniform asymptotic expansions of Laplace integrals. Analysis, 3(1–4):221–249, 1983. ISSN 0174-4747 (print), 2196-6753 (electronic).

Temme:1985:LTI

[Tem85] N. M. Temme. Laplace type integrals: transformation to standard form and uniform asymptotic expansions. Quarterly of Applied Mathematics, 43(1):103–123, 1985. CODEN QAMAAY. ISSN 0033-569x (print), 1552-4485 (electronic).

Temme:1986:DIC

[Tem86] N. M. Temme. A double integral containing the modified Bessel function: asymptotics and computation. Mathematics of Computation, 47(176):683–691, October 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Temme:1987:CIG

[Tem87] N. M. Temme. On the computation of the incomplete gamma functions for large values of the parameters. In Mason and Cox [MC87], pages 479–489. ISBN 0-19-853612-7. LCCN QA221 .A5361 1987; QA221 .I47 1985. US$90.

Temme:1992:AII

[Tem92] N. M. Temme. Asymptotic inversion of incomplete gamma functions. Mathematics of Computation, 58(198):755–764, April 1992. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Temme:1994:CAI

[Tem94a] N. M. Temme. Computational aspects of incomplete gamma functions with large complex parameters. In Zahar [Zah94], pages 551–562. ISBN 0-8176-3753-2. LCCN QA221 .A634 1994.

Temme:1994:SAI

[Tem94b] N. M. Temme. A set of algorithms for the incomplete gamma functions. Probability in the Engineering and Informational Sciences, 8(2):291–307, April 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.cambridge.

org/core/product/F5677268895A0805A0BDF31E4B20A106.

Temme:1996:SFI

[Tem96a] N. M. Temme. Special Functions: an Introduction to the Classical Functions of Mathematical Physics. Wiley, New York, NY, USA, 1996. ISBN 0-471-11313-1. xii + 374 pp. LCCN QC20.7.F87 T46 1996.

Temme:1996:UAI

[Tem96b] N. M. Temme. Uniform asymptotics for the incomplete gamma functions starting from negative values of the parameters. Methods and Applications of Analysis, 3(3):335–344, 1996. ISSN 1073-2772 (print), 1945-0001 (electronic).

Temme:2000:NAA

[Tem00] Nico M. Temme. Numerical and asymptotic aspects of parabolic cylinder functions. Journal of Computational and Applied Mathematics, 121(1–2):221–246, September 1, 2000. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042700003472.

Temme:2007:NAS

[Tem07] Nico M. Temme. Numerical aspects of special functions. Acta Numerica, 16:379–478, 2007. CODEN ANUMFU. ISBN 0-521-87743-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Temme:2015:AMI

[Tem15] Nico M. Temme. Asymptotic Methods for Integrals, volume 6 of Series in Analysis. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2015. ISBN 981-4612-15-4 (hardcover), 981-4612-16-2 (e-book). xxii + 605 pp.

Terras:1979:DIG

[Ter79] Riho Terras. The determination of incomplete gamma functions through analytic integration. Journal of Computational Physics, 31(1):146–151, April 1979. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999179900664.

Terras:1981:ASI

[Ter81a]  Riho Terras.  Algorithms for some integrals of Bessel

functions and multivariate Gaussian integrals. Journal of Computational Physics, 41(1):192–199, May 1981. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0021999181900875.

Terras:1981:MAI

[Ter81b] Riho Terras. A Miller algorithm for an incomplete Bessel function. Journal of Computational Physics, 39(1):233–240, January 1981. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999181901479.

Tesler:1969:AEF

[Tes69] G. S. Tesler. The approximation of elementary functions by means of polynomials of degree zero and one. (Russian). In Mathematical provisioning of electronic digital computers and effective organization of the computing process (Proc. Sem., Kiev, 1969) (Russian), No. 2, pages 75–88. Akad. Nauk Ukrain. SSR, Kiev, USSR, 1969.

Thomas:1989:SNL

[TGC89] Marlin A. Thomas, Gary W. Gemmill, and John R. Crigler. STATLIB: NSWC library of statistical programs and subroutines. Technical Report NSWC TR 89-97, Naval Surface Warfare Center, Dahlgren, VA 22448-5000, USA and Silver Spring, MD 20903-5000, USA, August 1989. viii + 280 pp. URL http:

//www.dtic.mil/dtic/tr/fulltext/u2/a221538.pdf.

Thacher:1961:ISR

[Tha61] Henry C. Thacher, Jr. Iterated square root expansions for the inverse cosine and inverse hyperbolic cosine. Mathematics of Computation, 15(76):399–403, October 1961. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thacher:1962:CAB

[Tha62a] Henry C. Thacher, Jr. Certification of Algorithm 57: ber or bei function. Communications of the ACM, 5(8):438, August 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thacher:1962:CAH

[Tha62b] Henry C. Thacher, Jr. Certification of Algorithms 191 and 192, Hypergeometric and Confluent Hypergeometric Functions. Communications of the ACM, 7(4):244, April 1962. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thacher:1963:ACEa

[Tha63a] Henry C. Thacher, Jr. Algorithm 165: Complete elliptic integrals. Communications of the ACM, 6(4):163–164, April 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thacher:1963:CACb

[Tha63b] Henry C. Thacher, Jr. Certification of Algorithm 149: Complete elliptic integral. Communications of the ACM, 6(4):166–167, April 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thacher:1963:CACa

[Tha63c] Henry C. Thacher, Jr. Certification of Algorithm 55: Complete elliptic integral of the first kind. Communications of the ACM, 6(4):166–167, April 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Thacher:1979:NBR

[Tha79] Henry C. Thacher, Jr. New backward recurrences for Bessel functions. Mathematics of Computation, 33(146):744–764, April 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

MPFR:2021:MLA

[The21] The MPFR Team. The MPFR library: Algorithms and proofs. Report, ????, ????, November 5, 2021. 69 pp. URL https://

www.mpfr.org/algorithms.pdf.

Thompson:1965:AEI

[Tho65] G. T. Thompson. The asymptotic expansion of the integrals psi and chi in terms of Tchebycheff polynomials (in Technical Notes and Short Papers). Mathematics of Computation, 19(92): 661–663, October 1965. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thompson:1966:ESI

[Tho66] Rory Thompson. Evaluation of I_{n}(b) = 2π^{−1} ∫_{0}^{∞} (\frac{sinx}{x})^{n} cos(bx)dx and of similar integrals (in Technical Notes and Short Papers). Mathematics of Computation, 20(94):330–332, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thompson:1987:IEF

[Tho87] Peter Thompson. Implementing an elementary function library. ACM SIGNUM Newsletter, 22(2):2–5, April 1987. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Thompson:1993:CCQ

[Tho93] William J. Thompson. Cutting corners: Quick square roots and trig functions. Computers in Physics, 7(1):18–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.
4823136.

Thompson:1997:ACMb

[Tho97a] William J. (William Jackson) Thompson. Atlas for Computing Mathematical Functions: an Illustrated Guide for Practitioners with Programs in Fortran 90 and Mathematica. Wiley, New York, NY, USA, 1997. ISBN 0-471-18171-4 (cloth). xiv + 888 pp. LCCN QA331.T386 1997. Includes CD-ROM.

Thompson:1997:ACMa

[Tho97b] William J. (William Jackson) Thompson. Atlas for Computing Mathematical Functions: an Illustrated Guidebook for Practitioners: with Programs in C and Mathematica. Wiley, New York, NY, USA, 1997. ISBN 0-471-00260-7 (cloth). xiv + 903 pp. LCCN QA331.T385 1997.

Thompson:2004:EBB

[Tho04] I. J. Thompson. Erratum to Modified Bessel functions I nu(z) and K_{ν}(z) of real order and complex argument [Comput. Phys. Commun. 47 (1987) 245–257]. Computer Physics Communications, 159(3):243–244, June 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465504001067. See [TB87].

Thompson:2013:AIG

[Tho13] Ian Thompson. Algorithm 926: Incomplete Gamma functions with negative arguments. ACM Transactions on Mathematical Software, 39(2):14:1–14:9, February 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Timan:1987:DFP

[Tim87] A. F. Timan. Distribution of fractional parts and approximation of functions with singularities by Bernstein polynomials. Journal of Approximation Theory, 50(2):167–174, June 1987. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Tkachev:2003:EFI

[Tka03] Vladimir G. Tkachev. Elliptic functions: Introduction course. Web lecture notes., November 25, 2003. URL http://users.

mai.liu.se/vlatk48/teaching/lect2-agm.pdf.

Tihanyi:2017:CEL

[TKK17] Norbert Tihanyi, Attila Kovács, and József Kovács. Computing extremely large values of the Riemann zeta function. Journal of Grid Computing, 15(4):527–534, December 2017. CODEN ???? ISSN 1570-7873 (print), 1572-9184 (electronic). URL https://

link.springer.com/article/10.1007/s10723-017-9416-0.

Tooper:1968:SCP

[TM68] Robert F. Tooper and John Mark. Simplified calculation of Ei(x) for positive arguments, and a short table of Shi(x) (in Technical Notes and Short Papers). Mathematics of Computation, 22 (102):448–449, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Tocher:1963:AS

[Toc63] K. D. Tocher. The Art of Simulation. Van Nostrand, Princeton, NJ, USA, 1963. viii + 184 pp. LCCN TA177 .T6.

Todd:1954:MWN

[Tod54] John Todd. Motivation for working on numerical analysis. In ????, editor, Transactions of 2nd Symposium on Applied Mathematics, 29 April 1954, University of Chicago, page ???? American Mathematical Society, Providence, RI, USA, 1954.

Todd:1955:MWN

[Tod55] John Todd. Motivation for working in numerical analysis. Communications on Pure and Applied Mathematics (New York), 8 (1):97–116, February 1955. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Todd:1990:WMP

[Tod90] John Todd. The Weierstrass mean. I. the periods of ℘(z|e_{1}, e_{2}, e_{3}). Numerische Mathematik, 57(8):737–746, August 1990. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Tolke:1943:PFE

[Töl43] Friedrich Tölke. Praktische Funktionenlehre. 1. Elementare und elementare transzendente Funktionen, Unterstufe. (German) [Practical functional theory. 1. Elementary and elementary transcendental functions, lower stage]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1943. viii + 261 pp.

Tolke:1950:PFE

[Töl50] Friedrich Tölke. Praktische Funktionenlehre. 1. Elementare und elementare transzendente Funktionen. (German) [Practical functional theory. 1. Elementary and elementary transcendental functions]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1950. xi + 440 pp.

Tolke:1966:PFT

[Töl66]  Friedrich Tölke.  Praktische Funktionenlehre. 2. Theta-

Funktionen und spezielle Weierstrasssche Funktionen. (German) [Practical functional theory. 2. Theta functions and special Weierstrass functions]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1966. vii + 248 pp. LCCN ????

Tolke:1967:PFJ

[Töl67a] Friedrich Tölke. Praktische Funktionenlehre. 3. Jacobische elliptische Funktionen, Legendresche elliptische Normalintegrale und spezielle Weierstrasssche Zeta- und Sigma-Funktionen. (German) [Practical functional theory. 3. Jacobian elliptic functions, Legendre elliptical normal Integrals and special Weierstrass zetaand sigma functions]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 3-642-50264-4, 3-662-36379-8, 3-662-35550-7 (print), 3-662-36379-8 (e-book).

viii + 180 pp. LCCN ???? URL http://www.springerlink.
com/content/978-3-662-36379-9.

Tolke:1967:PFE

[Töl67b] Friedrich Tölke. Praktische Funktionenlehre. 4. Elliptische Integralgruppen und Jacobische elliptische Funktionen im Komplexen. (German) [Practical functional theory. 4. Elliptical integral groups and Jacobian elliptic functions in the complex plane]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 3-662-36381-X, 3-662-35552-3 (print), 3-662-36381-X (e-book). viii + 191 pp. LCCN

???? URL http://www.springerlink.com/content/978-3-
662-36381-2.

Tolke:1968:PFA

[Töl68] Friedrich Tölke. Praktische Funktionenlehre. 5. Allgemeine Weierstrasssche Funktionen und Ableitungen nach dem Parameter: Integrale der Theta-Funktionen und Bilinear-Entwicklungen. (German) [Practical functional theory. 5. General information on Weierstrass functions and derivatives according to the parameters: integrals of theta functions and bilinear developments]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1968. ISBN 3-662-11121-7, 3-662-11120-9. viii + 158 pp. LCCN ????

Tolke:1969:PFT

[Töl69] Friedrich Tölke. Praktische Funktionenlehre. 6. Tafeln aus dem Gebiet der Theta-Funktionen und der elliptischen Funtionen. (German) [Practical functional theory. 6. Tables from the field of theta functions and elliptic functions]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1969. lxxxii + 449 (vol. 1) pp. LCCN ???? Two volumes.

Tolke:1970:PFT

[Töl70] Friedrich Tölke. Praktische Funktionenlehre. 6. Tafeln aus dem Gebiet der Theta-Funktionen und der elliptischen Funtionen. (German) [Practical functional theory. 6. Tables from the field of theta functions and elliptic functions]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. ISBN 3-662-13079-3 (print), 3-662-13078-5. 452–1047 (vol. 2) pp. LCCN ???? Two volumes.

Tommiska:2000:AEI

[Tom00] M. T. Tommiska. Area-efficient implementation of a fast square root algorithm. In Proceedings of the 2000 Third IEEE International Caracas Conference on Devices, Circuits and Systems, 15–17 March 2000, pages S18/1–S18/4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. CODEN ???? ISSN ????

Tornaria:2002:SRM

[Tor02] Gonzalo Tornaría. Square roots modulo p. Lecture Notes in Computer Science, 2286:430–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer-ny.com/link/service/series/0558/bibs/
2286/22860430.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/2286/22860430.pdf.

Tamura:1984:CFC

[TR84] Taro Tamura and Frank Rybicki. Coulomb functions for complex energies. Computer Physics Communications, 35(1–3):C–5, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465584822766.

Traub:1960:CNM

[Tra60] J. F. Traub. Comments on a recent paper [“A New Method of Computation of Square Roots Without Using Division”]. Communications of the ACM, 3(2):86, February 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Sar59].

Traub:1976:ACC

[Tra76] J. F. (Joseph Frederick) Traub, editor. Analytic computational complexity: Proceedings of the Symposium on Analytic Computational Complexity, held by the Computer Science Department, Carnegie-Mellon University, Pittsburgh, Pennsylvania, on April 7–8, 1975. Academic Press, New York, NY, USA, 1976. ISBN 0-12-697560-4. LCCN QA297.S9151 1975.

Tretjakov:1980:PSE

[Tre80] V. A. Tret’jakov. On the properties of some elementary functions that are defined on the algebra of bicomplex numbers. (Russian). In Mathematical analysis and the theory of functions (Russian), pages 99–106. Moskov. Oblast. Ped. Inst., Moscow, USSR, 1980.

Timmermann:1994:CFV

[TRHH94] D. Timmermann, B. Rix, H. Hahn, and B. J. Hosticka. A CMOS floating-point vector-arithmetic unit. IEEE Journal of Solid-State Circuits, 29(5):634–639, May 1994. CODEN IJSCBC. ISSN 0018-9200 (print), 1558-173X (electronic).

Tricomi:1950:AEU

[Tri50] F. G. Tricomi. Asymptotische Eigenschaften der unvollständigen Gammafunktion. (German) [Asymptotic properties of the incomplete gamma function]. Mathematische Zeitschrift, 53(2): 136–148, April 1950. CODEN MAZEAX. ISSN 0025-5874 (print), 1432-1823 (electronic). URL http://link.springer.

com/article/10.1007/BF01162409.

Tricomi:1966:RUS

[Tri66] F. G. Tricomi. Lectures on the use of special functions by calculations with electronic computers. Lecture Series 47, The Institute for Fluid Dynamics and Applied Mathematics, University of Maryland, College Park, College Park, MD, USA, 1966.

Trojan:1984:LBF

[Tro84] George M. Trojan. Lower bounds and fast algorithms for sequence acceleration. Journal of the Association for Computing Machinery, 31(2):329–335, April 1984. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Trudgian:2011:ITM

[Tru11] Timothy Trudgian. Improvements to Turing’s method. Mathematics of Computation, 80(276):2259–2279, October 2011. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (elec-

tronic). URL http://www.ams.org/journals/mcom/2011-80-
276/S0025-5718-2011-02470-1/home.html;   http://www.
ams.org/journals/mcom/2011-80-276/S0025-5718-2011-02470-

1/S0025-5718-2011-02470-1.pdf; http://www.ams.org/mathscinet-
getitem?mr=2813359. See [Tur53, Leh70].

TadeudeMedeiros:1969:APF

[TS69] Adilson Tadeu de Medeiros and Georges Schwachheim. Algorithm 349: Polygamma functions with arbitrary precision [S14]. Communications of the ACM, 12(4):213–214, April 1969. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See certification [Lew75].

Tugov:1976:MCA

[TS76] I. I. Tugov and Yu. L. Shitkov. A method of calculating the Appell functions F_{a}(α; β, β′; γ; x, y). U.S.S.R. Computational Mathematics and Mathematical Physics, 16(6):1587–1590, 1976. CODEN CMMPA9. ISSN 0041-5553, 0502-9902.

Tenca:2018:PIS

[TT18] Alexandre Tenca and Naofumi Takagi, editors. Proceedings of the 25th International Symposium on Computer Arithmetic, 25– 27 June 2018 Amherst, MA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 2018. ISBN 1-5386-2612-8 (USB), 1-5386-2665-9. ISSN 2576-2265. LCCN QA76.9.C62. IEEE catalog number CFP18121-USB.

Takemasa:1979:CFC

[TTW79] T. Takemasa, T. Tamura, and H. H. Wolter. Coulomb functions with complex angular momenta. Computer Physics Communications, 17(4):351–355, July/August 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Takemasa:1984:CFC

[TTW84] T. Takemasa, T. Tamura, and H. H. Wolter. Coulomb functions with complex angular momenta. Computer Physics Communications, 35(1–3):C–562, ???? 1984. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046558482737X.

Tukey:1948:NSR

[Tuk48] John W. Tukey. A note on the square-root iteration. SRG Memorandum report 10, Princeton University, Princeton, NJ, USA, 1948. 18 pp.

Turing:1943:MCZ

[Tur43] A. M. Turing. A method for the calculation of the zeta-function. Proceedings of the London Mathematical Society. Second Series, 48:180–197, 1943. ISSN 0024-6115 (print), 1460-244X (elec-

tronic). URL http://turing.ecs.soton.ac.uk/browse.php/
B/17.

Turing:1953:SCR

[Tur53] A. M. Turing. Some calculations of the Riemann zeta-function. Proceedings of the London Mathematical Society. Third Series, 3(3):99–117, 1953. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244X (electronic). URL http://turing.ecs.

soton.ac.uk/browse.php/B/21. See corrections and improve-

ments [Leh70], and [Tru11]. The latter comments: “Turing’s Method has become the standard technique used in modern verification of the Riemann hypothesis.” See also [Leh56].

Turner:1969:DSC

[Tur69] L. R. Turner. Difficulty in sin /cos routine. ACM SIGNUM Newsletter, 4(3):13, 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Turunov:1972:EFD

[Tur72] M. Turunov. Elementary functions of a discrete real and complex argument. (Russian). Taškent. Gos. Univ. Naučn. Trudy, 418 Voprosy Mat.:263–271, 386, 1972.

Turner:1994:SRM

[Tur94] Stephen M. Turner. Square roots mod p. American Mathematical Monthly, 101(5):443–449, May 1994. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Temme:2009:AER

[TV09]  Nico M. Temme and Vladimir Varlamov.  Asymptotic

expansions for Riesz fractional derivatives of Airy functions and applications. Journal of Computational and Applied Mathematics, 232(2):201–215, October 15, 2009. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042709003410.

Tchoumatchenko:2000:FBS

[TVG00] V. Tchoumatchenko, T. Vassileva, and P. Gurov. A FPGA based square-root coprocessor. In Proceedings of the 22nd EUROMI-CRO Conference EUROMICRO 96. ’Beyond 2000: Hardware and Software Design Strategies’, pages 520–525. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. CODEN ???? ISSN ????

Tretter:1980:ASA

[TW80] Marietta J. Tretter and G. W. Walster. Analytic subtraction applied to the incomplete gamma and beta functions. SIAM Journal on Scientific and Statistical Computing, 1(3):321–326, September 1980. CODEN SIJCD4. ISSN 0196-5204.

Ubhaya:1989:LAN

[Ubh89] V. A. Ubhaya. L_{p} approximation from nonconvex subsets of special classes of functions. Journal of Approximation Theory, 57(2):223–238, May 1989. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Uzer:2005:CAS

[UE05] A. Uzer and T. Ege. On the convergence acceleration of slowly convergent sums involving oscillating terms. Computing: Archiv fur informatik und numerik, 75(4):311–318, August 2005. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0010-485X&volume=
75&issue=4&spage=311.

Ullrich:1990:CCA

[Ull90] Christian Ullrich, editor. Contributions to Computer Arithmetic and Self-Validating Numerical Methods. (Proceedings of SCAN 89, held in Basel, Oct. 2–6, 1989), volume 7 of IMACS annals on computing and applied mathematics. J. C. Baltzer AG, Scientific Publishing Company, Basel, Switzerland, 1990. ISBN ???? LCCN QA76.9.C62 C664 1990.

USNBSCL:1949:TCH

[Uni49] United States National Bureau of Standards Computation Laboratory . Tables of the confluent hypergeometric function F (n/2, 1/2, x) and related functions, volume 3 of Applied mathematics series. U.S. Government Printing Office, Washington, DC, USA, 1949. xxii + 73 pp. LCCN QA3 .U5 no. 3.

URL http://openlibrary.org/works/OL1218358W/Tables_
of_the_confluent_hypergeometric_function_F%28n_2_1_2_
x%29_and_related_functions.

USENIX:1988:UPC

[USE88] USENIX Association, editor. USENIX proceedings: C++ Conference, Denver, CO, October 17–21, 1988. USENIX, San Francisco, CA, USA, 1988.

VanBuren:1975:TAS

[V^{+}75] A. L. (Arnie Lee) Van Buren et al. Tables of Angular Spheroidal Wave Functions. Naval Research Laboratory, Washington, DC, USA, 1975. ???? pp. LCCN QC174.26.W3 T27.

Vaidya:1989:LWB

[Vai89] P. M. Vaidya. A locally well-behaved potential function and a simple Newton-type method for finding the center of a polytype. In Megiddo [Meg89b], pages 79–90. ISBN 0-387-96847-4. LCCN QA402.5 .P7851 1989. Proceedings of the conference held at the Asilomar conference center in Pacific Grove, California, March 1–4, 1987.

VandeVel:1969:SEM

[Van69] H. Van de Vel. On the series expansion method for computing incomplete elliptic integrals of the first and second kinds. Mathematics of Computation, 23(105):61–69, January 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

VanHalen:1989:AAA

[Van89] P. Van Halen. Accurate analytical approximations for error function and its integral. Electronics Letters, 25(9):561–563, April 27, 1989. CODEN ELLEAK. ISSN 0013-5194 (print),

1350-911X (electronic). URL http://ieeexplore.ieee.org/
document/19780/.

Varchenko:1995:MHF

[Var95] A. N. (Aleksandr Nikolaevich) Varchenko. Multidimensional Hypergeometric Functions and Representation Theory of Lie Algebras and Quantum Groups, volume 21 of Advanced series in mathematical physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1995. ISBN 981-02-1880-X. ix + 371 pp. LCCN QA353.H9 V37 1995.

Vavasis:1989:CFP

[Vav89] S. A. Vavasis. Complexity of fixed point computations. Ph.D thesis, Stanford University, Stanford, CA, USA, 1989. ???? pp.

Vogelius:1981:DRM

[VB81] M. Vogelius and I. Babuska. On a dimensional reduction method. II. Some approximation-theoretic results. Mathematics of Computation, 37(155):47–68, July 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

VanDeun:2004:IAO

[VB04] J. Van Deun and A. Bultheel. An interpolation algorithm for orthogonal rational functions. Journal of Computational and Applied Mathematics, 164–165(??):749–762, March 2004. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Proceedings of the 10th International Congress on Computational and Applied Mathematics University of Leuven, Belgium, 22–26 July 2002. Edited by M. J. Goovaerts, S. Vandewalle, and L. Wuytack.

VanDeun:2006:ACI

[VC06a] Joris Van Deun and Ronald Cools. Algorithm 858: Computing infinite range integrals of an arbitrary product of Bessel functions. ACM Transactions on Mathematical Software, 32(4): 580–596, December 2006. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

VanDeun:2006:SRI

[VC06b] Joris Van Deun and Ronald Cools. A stable recurrence for the incomplete gamma function with imaginary second argument. Numerische Mathematik, 104(4):445–456, October 2006. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Verdonk:2001:PRIa

[VCV01] Brigitte Verdonk, Annie Cuyt, and Dennis Verschaeren. A precision- and range-independent tool for testing floatingpoint arithmetic I: Basic operations, square root, and remainder. ACM Transactions on Mathematical Software, 27(1): 92–118, March 2001. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://doi.acm.org/10.

1145/382043.382404;   http://www.win.ua.ac.be/~cant/
ieeecc754.html.

Lou:1988:ETR

[vdD88] Lou van den Dries. On the elementary theory of restricted elementary functions. Journal of Symbolic Logic, 53(3):796–808, 1988. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic).

vanderLaan:1984:CSF

[vdLT84] C. G. van der Laan and N. M. Temme. Calculation of special functions: the gamma function, the exponential integrals and error-like functions, volume 10 of CWI tract / Centrum voor Wiskunde en Informatica. Centre for Mathematics and Computer Science, Amsterdam, The Netherlands, 1984. ISBN 90-6196-277-3. iv + 231 pp. LCCN QA1 M4591 no. 10.

vandeRiet:1963:CAI

[vdR63] R. P. van de Riet. Certification of Algorithm 73: Incomplete elliptic integrals. Communications of the ACM, 6(4):167, April 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Vassiliadis:2005:IIC

[VDR05] Stamatis Vassiliadis, Nikitas J. Dimopoulos, and Sanjay Vishnu Rajopadhye, editors. 16th IEEE International Conference on Application-Specific Systems, Architectures, and Processors: ASAP 2005: 23–25 July 2005, Samos, Greece. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2407-9. LCCN TK7874.6 .I58 2005.

Veberic:2012:LFA

[Veb12] Darko Veberič. Lambert W function for applications in physics. Computer Physics Communications, 183(12):2622–2628, December 2012. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465512002366.

Vedder:1993:IAN

[Ved93] John D. Vedder. An invertible approximation to the normal distribution function. Computational Statistics & Data Analysis, 16(1):119–123, June 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http://www.

sciencedirect.com/science/article/pii/016794739390248R.

Veling:2011:GIG

[Vel11] E. J. M. Veling. The generalized incomplete gamma function as sum over modified Bessel functions of the first kind. Journal of Computational and Applied Mathematics, 235(14):4107– 4116, May 15, 2011. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042711001245.

Vepstas:2008:EAA

[Vep08] Linas Vepstas. An efficient algorithm for accelerating the convergence of oscillatory series, useful for computing the polylogarithm and Hurwitz zeta functions. Numerical Algorithms, 47(3):211–252, March 2008. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
1017-1398&volume=47&issue=3&spage=211.

Verma:1967:NSG

[Ver67] Arun Verma. A note on the summation of the generalised hypergeometric functions (in Technical Notes and Short Papers). Mathematics of Computation, 21(98):232–236, April 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Verma:1977:CSF

[Ver77] Arun Verma. Certain summation formulae for basic hypergeometric series. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 20(??):369–376, ???? 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Vial:1989:APP

[Via89] J.-P. Vial. Approximate projections in a projective method for the linear feasibility problem. In Megiddo [Meg89b], pages 65– 78. ISBN 0-387-96847-4. LCCN QA402.5 .P7851 1989. Proceedings of the conference held at the Asilomar conference center in Pacific Grove, California, March 1–4, 1987.

Vilenkin:1988:SFT

[Vil88] N. Ja. (Naum Jakovlevich) Vilenkin. Special functions and the theory of group representations, volume 22 of Translations of mathematical monographs. American Mathematical Society, Providence, RI, USA, 1988. ISBN 0-8218-1572-5. x + 613 pp. LCCN QA3 .A5 v.22 1988. Reprint of 1968 edition.

Vilenkin:1995:RLG

[VK95] N. Ja. (Naum Jakovlevich) Vilenkin and A. U. (Anatolii Ulsianovich) Klimyk. Representation of Lie Groups and Special Functions: Recent Advances, volume 316 of Mathematics and its applications. Kluwer Academic Publishers, Dordrecht, The Netherlands, 1995. ISBN 0-7923-3210-5. xvi + 497 pp. LCCN QA176 .V55 1995. URL http://www.loc.gov/

catdir/enhancements/fy0823/95108075-d.html;   http:/
/www.loc.gov/catdir/enhancements/fy0823/95108075-t.html.

Vladimirov:2002:MTG

[Vla02] V. S. (Vasilii Sergeevich) Vladimirov. Methods of the Theory of Generalized Functions, volume 6 of Analytical methods and special functions. Taylor and Francis, Boca Raton, FL, USA, 2002. ISBN 0-415-27356-0. xiv + 311 pp. LCCN QC20.7.T45 V53 2002.

Vazquez-Leal:2012:HAS

[VLCSFN^{+}12] Hector Vazquez-Leal, Roberto Castaneda-Sheissa, Uriel Filobello-Nino, Arturo Sarmiento-Reyes, and Jesus Sanchez Orea. High accurate simple approximation of normal distribution integral. Mathematical Problems in Engineering, 2012:1–22, 2012. ISSN 1024-123X (print), 1563-5147 (electronic). URL https://www.

hindawi.com/journals/mpe/2012/124029/.

Visentin:1987:FAE

[VM87] Kley Visentin and Pablo Martín. Fractional approximation to elliptic functions. Journal of Mathematical Physics, 28(2):330– 333, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i2/p330_s1.

Volkova:2019:SAI

[VM19] Anastasia Volkova and Jean-Michel Muller. Semi-automatic implementation of the complementary error function. In Takagi et al. [TBL19], pages 167–174. ISBN 1-72813-366-1. ISSN 1063-6889.

Volder:1959:CCT

[Vol59a] Jack Volder. The CORDIC computing technique. Proceedings of the Western Joint Computer Conference, pages 257–261, 1959.

Volder:1959:CTC

[Vol59b] Jack E. Volder. The CORDIC trigonometric computing technique. IRE Transactions on Electronic Computers, EC-8(3): 330–334, September 1959. CODEN IRELAO. ISSN 0367-

9950. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=5222693.

Volz:1983:CAA

[Völ83] H. Völz. CORDIC und ähnliche Algorithmen der elementaren Funktionen mit besonderer Eignung für Mikrorechner [English: CORDIC and Similar Algorithms for Elementary Functions with Particular Aptitude for Microcomputers]. Nachrichtentechnik Elektronik, 33(12):506–510, ???? 1983. CODEN NTELAP. ISSN 0323-4657.

vonGudenberg:1980:EAR

[von80] J. Wolff von Gudenberg. Einbettung allgemeiner Rechnerarithmetik in Pascal mittels eines Operatorkonzepts und Implementierung der Standardfunktionen mit optimaler Genauigkeit [English: Embedding a General Computer Arithmetic in Pascal by Means of an Operator Concept and the Implementation of Elementary Functions with Optimal Accuracy]. Dissertation, Universität Karlsruhe, Karlsruhe, Germany, 1980. ???? pp.

vonGudenberg:1984:BMG

[von84] J. Wolff von Gudenberg. Berechnung maximal genauer Standardfunktionen mit einfacher Mantissenlänge [English: Computation of Maximally Accurate Elementary Functions Using Simple Mantissa Length]. Elektronische Rechenanlagen, 26(5): 230–238, October 1984. CODEN ELRAA4. ISSN 0013-5720.

Vos:1973:RAC

[Vos73] H. Vos. Remark on “Algorithm 300: Coulomb wave functions”. Communications of the ACM, 16(5):308–309, May 1973. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Gun67].

Vyridis:1999:ICA

[VPK99] D. G. Vyridis, S. D. Panteliou, and I. N. Katz. An inverse convergence approach for arguments of Jacobian elliptic functions. Computers and Mathematics with Applications, 37(2):21– 26, January 1999. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122198002508.

vonRosen:1989:MLE

[vR89] D. von Rosen. Maximum likelihood estimators in multivariate linear normal models. Journal of Multivariate Analysis, 31(2): 187–200, November 1989. CODEN JMVAAI. ISSN 0047-259x (print), 1095-7243 (electronic).

vanRijckevorsek:1988:CCA

[vRdL88] Jan L. A. van Rijckevorsek and Jan de Leeus, editors. Component and correspondence analysis: dimension reduction by functional approximation. Wiley series in probability and mathematical statistics. Wiley, New York, NY, USA, 1988. ISBN 0-471-91847-4. xiii + 146 pp. LCCN QA278.5 .C6571 1988. US$40.

Vrahatis:1995:RPP

[VRS^{+}95] M. N. Vrahatis, O. Ragos, T. Skiniotis, F. A. Zafiropoulos, and T. N. Grapsa. RFSFNS: a portable package for the numerical determination of the number and the calculation of roots of Bessel functions. Computer Physics Communications, 92(2–3):252–266, December 1995. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0010465595001159. See erratum [VRS^{+}99].

Vrahatis:1999:ESP

[VRS^{+}99] M. N. Vrahatis, O. Ragos, T. Skiniotis, F. A. Zafiropoulos, and T. N. Grapsa. Erratum to: RFSFNS: a portable package for the numerical determination of the number and the calculation of roots of Bessel functions [Comput. Phys. Commun. 92 (1995) 252–266]. Computer Physics Communications, 117(3):290, March 11, 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046559800109X. See [VRS^{+}95].

Vallee:2004:AFA

[VS04] Olivier Vallée and Manuel Soares. Airy Functions and Applications to Physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2004. ISBN 1-86094-478-7 (hardcover). x + 194 pp. LCCN QA351 .V35 2004.

Vallee:2010:AFA

[VS10] Olivier Vallée and Manuel Soares. Airy Functions and Applications to Physics. Imperial College Press, London WC26 9HE, UK, second edition, 2010. ISBN 1-84816-548-X, 1-84816-550-1. x + 202 pp. LCCN QA351.

VanDeun:2011:RIC

[VT11] Joris Van Deun and Lloyd N. Trefethen. A robust implementation of the Carathéodory-Fejér method for rational approximation. BIT Numerical Mathematics, 51(??):??, ???? 2011. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125

(electronic). URL http://www.springerlink.com/content/
ag2514840142707r/.

Vazquez:2009:CDT

[VVA09] Álvaro Vázquez, Julio Villalba, and Elisardo Antelo. Computation of decimal transcendental functions using the CORDIC algorithm. In Bruguera et al. [BCDH09], pages 179–186. ISBN 0-7695-3670-0, 1-4244-4329-6. ISSN 1063-6889. LCCN QA76.6

.S887 2009. URL http://www.ac.usc.es/arith19/.

Wachspress:2000:EEF

[Wac00] E. L. Wachspress. Evaluating elliptic functions and their inverses. Computers and Mathematics with Applications, 39 (3–4):131–136, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0898122199003399.

Wadey:1958:TSR

[Wad58] W. G. Wadey. Two square-root approximations. Communications of the ACM, 1(11):13–14, November 1958. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wahdan:2004:IHE

[Wah04] Abdel-Moniem Wahdan, editor. ICEEC’04: 2004 International Conference on Electrical, Electronic and Computer Engineering: proceedings: 5–7 September, 2004, Cairo, Egypt. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7803-8575-6. LCCN TK7801 .I5125 2004. IEEE catalog number 04EX893.

Walther:1971:UAE

[Wal71] J. S. Walther. A unified algorithm for elementary functions. In AFIPS SJCC ’71 [AFI71], pages 379–385. LCCN ????

Waldecker:1980:NSR

[Wal80] D. E. Waldecker. Nonrestoring square root with simplified answer generation. IBM Technical Disclosure Bulletin, 22(11): 4807–4808, April 1980. CODEN IBMTAA. ISSN 0018-8689.

Walmsley:1984:EEM

[Wal84] John L. Walmsley. On the efficient evaluation of modified Bessel functions of zeroth and first orders for arguments of the form x exp(iπ/4). Journal of Computational Physics, 56(2):349–355, November 1984. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999184901001.

Walz:1996:AE

[Wal96] Guido Walz. Asymptotics and Extrapolation, volume 88 of Mathematical research. Akademie-Verlag, Berlin, Germany, 1996. ISBN 3-05-501732-3. 330 pp. LCCN QA281 .W349 1996.

Wall:2000:ATC

[Wal00] H. S. Wall. Analytic Theory of Continued Fractions. American Mathematical Society, Providence, RI, USA, 2000. ISBN 0-8218-2106-7. xiii + 433 pp. LCCN QA295 .W28 2000. This is a reprint of the definitive, and widely cited, treatise first published in 1948.

Wang:1974:UEZ

[Wan74] Paul S. Wang. The undecidability of the existence of zeros of real elementary functions. Journal of the Association for Computing Machinery, 21(4):586–589, October 1974. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Wang:1978:EPF

[Wan78] J. Y. Wang. The evaluation of periodic functions with large input arguments. ACM SIGNUM Newsletter, 13(4):7–8, December 1978. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Wang:1980:ISM

[Wan80] J. Y. Wang. On the improvement of some mathematical subroutines in the IBM S/360 FORTRAN IV libraries. In SHARE-54, Anaheim, CA, March 3, 1980, volume 1, pages 75–77. ????, ????, 1980.

Wang:1982:AME

[Wan82] J.-L Wang. Asymptotically minimax estimators for distributions with increasing failure rate. Ph.D thesis, University of California, Berkeley, Berkeley, CA, USA, 1982. 42 pp.

Wang:2016:UAA

[Wan16] Weiping Wang. Unified approaches to the approximations of the gamma function. Journal of Number Theory, 163(??):570– 595, June 2016. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X16000470.

Ward:1960:CCE

[War60a] Morgan Ward. The calculation of the complete elliptic integral of the third kind. American Mathematical Monthly, 67(3):205– 213, March 1960. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/
2309677.

Warmus:1960:TEF

[War60b] Mieczysław Warmus. Tables of elementary functions. Pergamon Press, New York, NY, USA, 1960. vii + 567 pp. LCCN QA55 .W3 1960. Państwowe Wydawnictwo Naukowe, Warsaw. Separately bound table of proportional parts, 30 pp.

Warner:1975:PDG

[War75] D. D. Warner. A partial derivative generator. Computing Science Technical Report 28, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1975.

Ware:2008:RIE

[War08] Willis H. Ware. RAND and the information evolution: a history in essays and vignettes. Rand Corporation, Santa Monica, CA, 2008. ISBN 0-8330-4513-X, 0-8330-4816-3, 1-282-45123-5. xxvi + 201 pp. LCCN QA76.27. URL http://

www.jstor.org/stable/10.7249/cp537rc;   https://www.

rand.org/content/dam/rand/pubs/corporate_pubs/2008/RAND_
CP537.pdf.

Wasilkowski:1989:AIV

[Was89] G. W. Wasilkowski. On adaptive information with varying cardinality for linear problems with elliptically contoured measures. Journal of complexity, 5(3):363–368, September 1989. CODEN JOCOEH. ISSN 0885-064X (print), 1090-2708 (electronic).

Watson:1966:TTB

[Wat66] G. N. Watson. A Treatise on the Theory of Bessel Functions. Cambridge University Press, Cambridge, UK, second edition, 1966. ISBN 0-521-09382-1. vi + 804 pp. LCCN QA 408 W33t 1966.

Watson:1990:NMC

[Wat90] G. Alistair Watson. Numerical methods for Chebyshev approximation of complex-valued functions. In Mason and Cox [MC90], pages 246–264. ISBN 0-412-34580-3. LCCN QA221 .I54 1988.

Watson:1995:TTB

[Wat95] G. N. (George Neville) Watson. A Treatise on the Theory of Bessel Functions. Cambridge mathematical library. Cambridge University Press, Cambridge, UK, second edition, 1995. ISBN 0-521-48391-3 (paperback), 0-521-06743-X (hardcover). vi + 804 pp. LCCN QA408 .W2 1995. URL http://www.loc.

gov/catdir/description/cam028/96139881.html;   http:
//www.loc.gov/catdir/toc/cam029/96139881.html.

Wang:1978:SMR

[WB78] J. Y. Wang and J. Boyer. A studv of the mathematical routines in the IBM System/360 FORTRAN IV and FORTRAN IV (Mod II) libraries. Report AMD-TM 304, Applied Mathematics Division, Argonne National Laboratory, Argonne, IL, USA, January 1978. Cited in [ACG^{+}86, Reference 8] in elefunt.bib and fparith.bib.

Wills:1982:RCA

[WBR82] C. A. Wills, J. M. Blair, and P. L. Ragde. Rational Chebyshev approximations for the Bessel functions J_{0}(x), J_{1}(x), Y_{0}(x), Y_{1}(x). Mathematics of Computation, 39(160):617–623, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Werner:1961:CAG

[WC61] Helmut Werner and Robert Collinge. Chebyshev approximations to the gamma function (in Technical Notes and Short Papers). Mathematics of Computation, 15(74):195–197, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Weniger:1990:RAM

[WC90] Ernst Joachim Weniger and Jiři Cížek. Rational approximations for the modified Bessel function of the second kind. Computer Physics Communications, 59(3):471–493, July 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/001046559090089J.

Wang:2016:AFG

[WCS16] Miao-Kun Wang, Yu-Ming Chu, and Ying-Qing Song. Asymptotical formulas for Gaussian and generalized hypergeometric functions. Applied Mathematics and Computation, 276(??):44– 60, March 5, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300315015908.

Weingarten:1961:TGC

[WD61] Harry Weingarten and A. R. Di Donato. A table of generalized circular error. Mathematics of Computation, 15(74):169–173, April 1961. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wang:2004:CHP

[WdZ04] Ren-Hong Wang and Cheng de Zheng. Cubic Hermite–Padé approximation to the exponential function. Journal of Computational and Applied Mathematics, 163(1):259–268, February 1, 2004. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0377042703008124.

Wei:1998:NFS

[Wei98] Liqiang Wei. New formula for 9–j symbols and their direct calculation. Computers in Physics, 12(6):632–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.
168745.

Wejntrob:1983:ASR

[Wej83] Leon Wejntrob. Approximation of square roots. Two-Year College Mathematics Journal, 14(5):427–431, November 1983. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/00494925.1983.11972733.

Wengert:1964:SAD

[Wen64] R. E. Wengert. A simple automatic derivative evaluation program. Communications of the ACM, 7(8):463–464, 1964. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Weniger:1989:NST

[Wen89] Ernst Joachim Weniger. Nonlinear sequence transformations for the acceleration of convergence and the summation of divergent series. Computer Physics Reports, 10(5–6):189–371, December 1989. CODEN CPHREF. ISSN 0167-7977 (print), 1878-1004 (electronic). Available as math.NA/0306302.

Weniger:1996:CWF

[Wen96] Ernst Joachim Weniger. Computation of the Whittaker function of the second kind by summing its divergent asymptotic series with the help of nonlinear sequence transformations. Computers in Physics, 10(5):496–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

aip.scitation.org/doi/10.1063/1.168579.

Weniger:2001:IID

[Wen01] Ernst Joachim Weniger. Irregular input data in convergence acceleration and summation processes: General considerations and some special Gaussian hypergeometric series as model problems. Computer Physics Communications, 133(2–3):202–228, January 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

Weniger:2007:AAT

[Wen07] Ernst Joachim Weniger. Asymptotic approximations to truncation errors of series representations for special functions. In Iske and Levesley [IL07], pages 331–348. ISBN 3-540-46551-0, 3-540-33283-9. LCCN QA221 .A44 2007.

Weniger:2009:SHF

[Wen09] Ernst Joachim Weniger. The strange history of B functions or how theoretical chemists and mathematicians do (not) interact. International Journal of Quantum Chemistry, 109(8): 1728–1740, February 2009. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weniger:2010:SDP

[Wen10] E. J. Weniger. Summation of divergent power series by means of factorial series. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1429–1441, ???? 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Presented at Approximation and Extrapolation of Convergent and Divergent Sequences and Series (CIRM, Luminy — France, 2009).

West:2005:BAC

[Wes05] G. West. Better approximations to cumulative normal functions. Wilmott Magazine, ??(??):70–76, ???? 2005. ISSN 1540-6962 (print), 1541-8286 (electronic).

Wang:2014:CFA

[WEX14] Dong Wang, Miloš D. Ercegovac, and Yang Xiao. Complex function approximation using two-dimensional interpolation. IEEE Transactions on Computers, 63(12):2948–2960, December 2014. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Wang:1989:SF

[WG89] Z. X. Wang and D. R. Guo. Special Functions. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1989. ISBN 9971-5-0659-9. xiii + 422 pp. LCCN QA331 .W296 1989.

Wong:1991:FHA

[WG91] W. F. Wong and E. Goto. Fast hardware-based algorithms for elementary function computations. In Anonymous [Ano91], pages 56–65. ISBN 4-87378-284-8. LCCN QA76.88.I 1991.

Wong:1992:DSR

[WG92] W. F. Wong and E. Goto. Division and square-rooting using a split multiplier. Electronics Letters, 28(18):1758–1759, August 27, 1992. CODEN ELLEAK. ISSN 0013-5194 (print), 1350-911X (electronic).

Wong:1994:FEE

[WG94a] W. F. Wong and E. Goto. Fast evaluation of the elementary functions in double precision. In Mudge and Shriver [MS94], pages 349–358. ISBN 0-8186-5050-8 (paper), 0-8186-5051-6 (microfiche). LCCN ???? First of five volumes. IEEE Catalog No. 94TH0607-2.

Wong:1994:FHB

[WG94b] W. F. Wong and E. Goto. Fast hardware-based algorithms for elementary function computations using rectangular multipliers. IEEE Transactions on Computers, 43(3):278–294, March 1994. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=272429.

Wong:1995:FEE

[WG95] W. F. Wong and E. Goto. Fast evaluation of the elementary functions in single precision. IEEE Transactions on Computers, 44(3):453–457, March 1995. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Whittlesey:1963:IGF

[Whi63] John R. B. Whittlesey. Incomplete gamma functions for evaluating Erlang process probabilities. Mathematics of Computation, 17(81):11–17, January 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Whitley:1982:MBI

[Whi82] R. Whitley. Markov and Bernstein’s inequalities, and compact and strictly singular operators. Journal of Approximation Theory, 34(3):277–285, March 1982. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Wieder:1999:ANH

[Wie99] Thomas Wieder. Algorithm 794: Numerical Hankel transform by the Fortran program HANKEL. ACM Transactions on Mathematical Software, 25(2):240–250, June 1999. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL ftp:

//netlib.bell-labs.com/netlib/toms/794.gz;   http://
phase.etl.go.jp/netlib/toms/794;   http://www.acm.org/
pubs/citations/journals/toms/1999-25-2/p240-wieder/;

 http://www.acm.org/pubs/citations/journals/toms/cgi-
bin/TOMSbibget?Wieder:1999:ANH; http://www.hensa.ac.
uk/netlib/toms/794.gz;   http://www.netlib.no/netlib/
toms/794; http://www.netlib.org/toms/794.

Wigner:1967:BRW

[Wig67] Eugene P. Wigner. Book review: Wilhelm Magnus, Fritz Oberhettinger, and R. P. Soni, Formulas and Theorems for the Special Functions of Mathematical Physics. Physics Today, 20 (12):81, December 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v20/i12/p81/s1.

Wilcox:1968:ZTN

[Wil68] Peter H. Wilcox. The zeros of P_{ν}^{1}(cos θ) and \frac{∂}{∂θ} P_{μ}^{1}(cos θ) (in Technical Notes and Short Papers). Mathematics of Computation, 22(101):205–208, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.

jstor.org/stable/2004783.

Wilson:1970:OSA

[Wil70] M. Wayne Wilson. Optimal starting approximations for generating square root for slow or no divide. Communications of the ACM, 13(9):559–560, September 1970. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wills:1971:URR

[Wil71] John G. Wills. On the use of recursion relations in the numerical evaluation of spherical Bessel functions and Coulomb functions. Journal of Computational Physics, 8(1):162–166, August 1971. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199917190043X.

Williams:1996:TMF

[Wil96] K. B. Williams. Testing math functions: When requirements are tight, we must carefully examine all potential sources of error. Make sure your math library isn’t the weak link in the chain. C/C++ Users Journal, 14(12):49–54, 58–65, December 1996. CODEN CCUJEX. ISSN 1075-2838. Describes a package that extends the Cody-Waite-Plauger work on the ELEFUNT package for the testing of the elementary functions, including the inverse hyperbolic functions, cube root, and Bessel functions of the first and second kinds. The C++ package implements 192bit extended precision versions of all of the functions, so that accurate results are available for comparison with the normal double-precision results.

Willis:2012:AGH

[Wil12] Joshua L. Willis. Acceleration of generalized hypergeometric functions through precise remainder asymptotics. Numerical Algorithms, 59(??):??, ???? 2012. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:/

/arxiv.org/abs/1102.3003;   http://www.springerlink.
com/content/k413064448600815/.

Wimp:1962:PEB

[Wim62] Jet Wimp. Polynomial expansions of Bessel functions and some associated functions. Mathematics of Computation, 16(80):446– 458, October 1962. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). See corrigendum [Wim72].

Wimp:1968:RFH

[Wim68] Jet Wimp. Recursion formulae of hypergeometric functions. Mathematics of Computation, 22(102):363–373, April 1968. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wimp:1972:CPE

[Wim72] Jet Wimp. Corrigendum: Polynomial expansions of Bessel functions and some associated functions (Math. Comp. 16 (1962), 446–458). Mathematics of Computation, 26(117): 309, January 1972. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.ams.org/

journals/mcom/1972-26-117/S0025-5718-1972-0400659-X;
 http://www.ams.org/journals/mcom/1972-26-117/S0025-
5718-1972-0400659-X/S0025-5718-1972-0400659-X.pdf; https:
//www.ams.org/mathscinet-getitem?mr=400659;   https:
//www.ams.org/mathscinet/search/authors.html?authorName=
Wimp%2C%20Jet. See [Wim62].

Wimp:1974:CTP

[Wim74] J. Wimp. On the computation of Tricomi’s Psi function. Computing: Archiv fur informatik und numerik, 13(3–4):195–203, 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Wimp:1981:STT

[Wim81] Jet Wimp. Sequence Transformations and Their Applications, volume 154 of Mathematics in Science and Engineering. Academic Press, New York, NY, USA, 1981. ISBN 0-08-095662-9 (e-book), 0-12-757940-0. xix + 257 pp. LCCN QA292 .W54.

URL http://public.eblib.com/EBLPublic/PublicView.do?
ptiID=453177.

Wimp:1982:CMS

[Wim82] Jet (Jesse Jet) Wimp. Computational methods and special functions. D.Sc. thesis, University of Edinburgh, Edinburgh, UK, 1982.

Winograd:1970:NMN

[Win70] Shmuel Winograd. On the number of multiplications necessary to compute certain functions. Communications on Pure and Applied Mathematics (New York), 23(2):165–179, March 1970. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Wise:1948:IBF

[Wis48] M. E. Wise. The incomplete beta function and the incomplete gamma function: An acknowledgment. Journal of the Royal Statistical Society. Series B (Methodological), 10(2):264, ???? 1948. CODEN JSTBAJ. ISSN 0035-9246. URL http://www.

jstor.org/stable/2983781.

Witte:1968:AAJ

[Wit68] B. F. W. Witte. ACM Algorithm 332: Jacobi polynomials. Communications of the ACM, 11(6):436–437, June 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See remark [Sko75b].

Wood:1966:DBI

[WKG66] Van E. Wood, R. P. Kenan, and M. L. Glasser. Doppler broadening integrals (in Technical Notes and Short Papers). Mathematics of Computation, 20(96):610–611, October 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wu:2011:NEL

[WLK11] Mingwei Wu, Xuzheng Lin, and Pooi-Yuen Kam. New exponential lower bounds on the Gaussian Q-function via Jensen’s inequality. In 2011 IEEE 73rd Vehicular Technology Conference (VTC Spring). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 2011.

URL http://ieeexplore.ieee.org/document/5956392/.

Wang:2014:FPT

[WMBE14] Dong Wang, Jean-Michel Muller, Nicolas Brisebarre, and Milo D. Ercegovac. (M, p, k)-Friendly points: a table-based method to evaluate trigonometric function. IEEE Transactions on Circuits and Systems II: Express Briefs, 61(9):711–715, 2014. ISSN 1549-7747 (print), 1558-3791 (electronic).

Walczyk:2018:IAF

[WMC18] Cezary J. Walczyk, Leonid V. Moroz, and Jan L. Cieśliński. Improving the accuracy of the fast inverse square root algorithm. arXiv.org, ??(??):1–21, February 17, 2018. URL https:

//arxiv.org/abs/1802.06302.

Walczyk:2021:IAF

[WMC21] Cezary J. Walczyk, Leonid V. Moroz, and Jan L. Cieśliński. Improving the accuracy of the fast inverse square root by modifying Newton–Raphson corrections. Entropy, 23(1):86:1–86:20, January 2021. CODEN ENTRFG. ISSN 1099-4300.

Woods:1992:HPD

[WMDM92] R. F. Woods, S. E. McQuillan, J. Dowling, and J. V. McCanny. High performance DSP ASIC for multiply, divide and square root. In Proceedings of Fifth Annual IEEE International ASIC Conference and Exhibit, 1992, pages 209–213. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. CODEN ???? ISSN ????

Wang:2003:TDF

[WN03] Xiaojun Wang and B. E. Nelson. Tradeoffs of designing floatingpoint division and square root on Virtex FPGAs. In FCCM 2003: 11th Annual IEEE Symposium on Field-Programmable Custom Computing Machines, 9–11 April 2003, pages 195–203. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. CODEN ???? ISSN ????

Wozny:2009:MSS

[WN09] Paweł Woźny and Rafał Nowak. Method of summation of some slowly convergent series. Applied Mathematics and Computation, 215(4):1622–1645, ???? 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Watanabe:1994:MSP

[WNO94] T. (Tsutomu) Watanabe, Makoto Natori, and Tsutomu Oguni. Mathematical Software for the P.C. and Work Stations: a Collection of Fortran 77 Programs. North-Holland, Amsterdam, The Netherlands, June 1994. ISBN 0-444-82000-0. xiv + 387 pp. LCCN QA 76.73 F25 F6813 1994. US$178.50. URL http://

www.cbooks.com/sqlnut/SP/search/gtsumt?source=&isbn=

0444820000. Translation of: FORTRAN 77 ni yoru suchi keisan sofutowea.

Weiss:2015:ROS

[WO15] Alexander K. H. Weiss and Christian Ochsenfeld. A rigorous and optimized strategy for the evaluation of the Boys function kernel in molecular electronic structure theory. Journal of Computational Chemistry, 36(18):1390–1398, July 5, 2015. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wong:1995:EHS

[WOG95] W. F. Wong, Yoshio Oyanagi, and Eiichi Goto. Evaluation of the Hitachi S-3800 supercomputer using six benchmarks. The International Journal of Supercomputer Applications and High Performance Computing, 9(1):58–70, Spring 1995. CODEN IJSAE9. ISSN 0890-2720.

Wojcicki:1961:ABF

[Woj61] Maria E. Wojcicki. Algorithm 44: Bessel functions computed recursively. Communications of the ACM, 4(4):177–178, April 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wong:1973:AEL

[Won73a] R. Wong. An asymptotic expansion of W_{k,m}(z) with large variable and parameters. Mathematics of Computation, 27(122): 429–436, April 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wong:1973:UAE

[Won73b] R. Wong. On uniform asymptotic expansion of definite integrals. Journal of Approximation Theory, 7(1):76–86, January 1973. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021904573900555.

Wong:1988:AE

[Won88] R. Wong. Asymptotic expansion of ∫_{0}^{π/2} J_{ν}^{2}(λ cos θ)dθ. Mathematics of Computation, 50(181):229–234, January 1988. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wood:1967:CEI

[Woo67] Van E. Wood. Chebyshev expansions for integrals of the error function (in Technical Notes and Short Papers). Mathematics of Computation, 21(99):494–496, July 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wood:1992:CP

[Woo92] David C. Wood. The computation of polylogarithms. Report, University of Kent, Canterbury, Kent CT2 7NZ, UK, 1992. URL

https://www.cs.kent.ac.uk/pubs/1992/110/content.pdf.

Wozny:2010:EAS

[Woź10] Paweł Woźny. Efficient algorithm for summation of some slowly convergent series. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1442–1453, ???? 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Presented at Approximation and Extrapolation of Convergent and Divergent Sequences and Series (CIRM, Luminy - France, 2009).

Werner:1963:AFI

[WR63] H. Werner and G. Raymann. An approximation to the Fermi integral F_{1/2}(x). Mathematics of Computation, 17(82):193– 194, April 1963. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL http://www.jstor.org/stable/
2003641.

Wong:1971:SEI

[WR71] R. Wong and E. Rosenbloom. Series expansions of W_{k,m}(z) involving parabolic cylinder functions. Mathematics of Computation, 25(116):783–787, October 1971. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Waissi:1996:SAS

[WR96] Gary R. Waissi and Donald F. Rossin. A sigmoid approximation of the standard normal integral. Applied Mathematics and Computation, 77(1):91–95, June 1996. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0096300395001905.

Wrench:1968:CTS

[Wre68] John W. Wrench, Jr. Concerning two series for the gamma function. Mathematics of Computation, 22(103):617–626, July 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrench:1973:ECT

[Wre73] John W. Wrench. Erratum: Concerning Two Series for the Gamma Function. Mathematics of Computation, 27(123):681– 682, July 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrigge:1973:EII

[Wri73] H. S. Wrigge. An elliptic integral identity. Mathematics of Computation, 27(124):839–840, October 1973. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wrigge:1984:NMG

[Wri84] Staffan Wrigge. A note on the moment generating function for the reciprocal gamma distribution. Mathematics of Computation, 42(166):617–621, April 1984. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wawrzynczyk:1984:GRS

[WS84] Antoni Wawrzyńczyk and Aleksander Strasburger. Group representations and special functions, volume 8 of Mathematics and its applications. East European series. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1984. ISBN 90-277-1269-7. xvi + 688 pp. LCCN QA171 .W3513 1984; QA1 M4281 v. 8.

Walters:2005:EFA

[WS05a] George Walters and Michael Schulte. Efficient function approximation using truncated multipliers and squarers. In IEEE [IEE05], page ?? ISBN ???? LCCN ???? URL http://

arith17.polito.it/final/paper-190.pdf.

Wang:2005:DFPa

[WS05b] L.-K. Wang and M. J. Schulte. Decimal floating-point square root using Newton–Raphson iteration. In Vassiliadis et al. [VDR05], pages 309–315. ISBN 0-7695-2407-9. LCCN TK7874.6

.I58 2005. URL http://mesa.ece.wisc.edu/publications/
cp_2005-05.pdf.

Wires:2006:RRS

[WS06] Kent E. Wires and Michael J. Schulte. Reciprocal and reciprocal square root units with operand modification and multiplication. Journal of VLSI Signal Processing, 42(3):257–272, March 2006. CODEN JVSPED. ISSN 0922-5773 (print), 1573-

109x (electronic). URL http://springerlink.metapress.
com/content/t6027p6713727606/fulltext.pdf.

Wells:1990:LE

[WTA^{+}90] Martin T. Wells, Ram C. Tiwari, I. Arizono, H. Ohta, James W. Mergerson, Gabriella M. Belli, Christopher Cox, Murray A. Jorgensen, Jacques Benichou, Mitchell H. Gail, Warren F. Kuhfeld, Brian Dawkins, Walter B. Studdiford, Colin Goodall, W. D. Kaigh, Stephen W. Looney, Robert Kinnison, James A. Gibbons, Joel R. Levin, Ronald C. Serlin, K. J. A. Revfeim, Charles R. McConnell, Robert M. Norton, R. W. Farebrother, I. J. Good, Stanley Lebergott, and Vedula N. Murty. Letters to the editor. The American Statistician, 44(1):56–65, February 1990. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731

(electronic). URL http://www.jstor.org/stable/2684963.

Weber:2005:MIG

[WTM05] Kenneth Weber, Vilmar Trevisan, and Luiz Felipe Martins. A modular integer GCD algorithm. Journal of Algorithms, 54(2):152–167, February 2005. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0196677404001075.

Wakhare:2020:TCJ

[WV20] Tanay Wakhare and Christophe Vignat. Taylor coefficients of the Jacobi θ_{3}(q) function. Journal of Number Theory, 216(??):280–306, November 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20300858.

Wilkes:1951:PPE

[WWG51] Maurice V. Wilkes, David J. Wheeler, and Stanley Gill. The Preparation of Programs for an Electronic Digital Computer. Addison-Wesley, Reading, MA, USA, 1951. 167 pp. LCCN QA76.5 .W55 1951. See also second edition [WWG57], and reprint [WWG82].

Wilkes:1957:PPE

[WWG57] Maurice V. Wilkes, David J. Wheeler, and Stanley Gill. The Preparation of Programs for an Electronic Digital Computer. Addison-Wesley, Reading, MA, USA, second edition, 1957. xiv + 238 pp. LCCN QA76.5 .W52 1957. URL https://b-ok.org/ book/3668116/b363ff. See also first edition [WWG51].

Wilkes:1982:PPE

[WWG82] M. V. (Maurice Vincent) Wilkes, David J. Wheeler, and Stanley Gill. The Preparation of Programs for an Electronic Digital Computer: with Special Reference to the EDSAC and the Use of a Library of Subroutines, volume 1 of Charles Babbage Institute reprint series for the history of computing. Tomash Publishers, Los Angeles, CA, USA, 1982. ISBN 0-262-23118-2 (MIT Press 1984), 0-938228-03-X. xxxi + 167 pp. LCCN QA76.6 .W545 1982. With a new introduction by Martin Campbell-Kelly.

Wynn:1960:RAF

[Wyn60] Peter Wynn. The rational approximation of functions which are formally defined by a power series expansion. Mathematics of Computation, 14(70):147–186, April 1960. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wynn:1962:AAP

[Wyn62] P. Wynn. An arsenal of ALGOL procedures for complex arithmetic. Nordisk tidskrift for informationsbehandling, 2 (4):232–255, December 1962. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://

www.springerlink.com/openurl.asp?genre=article&issn=
0006-3835&volume=2&issue=4&spage=232.

Wynn:1972:CAM

[Wyn72] Peter Wynn. Convergence acceleration by a method of intercalation. Computing: Archiv fur informatik und numerik, 9(4): 267–273, 1972. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Xu:2014:SII

[XC14] Ai-Min Xu and Zhong-Di Cen. Some identities involving exponential functions and Stirling numbers and applications. Journal of Computational and Applied Mathematics, 260(??):201– 207, April 2014. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042713005323.

Xu:2015:CFC

[XC15] Ai-Min Xu and Zhong-Di Cen. Closed formulas for computing higher-order derivatives of functions involving exponential functions. Applied Mathematics and Computation, 270(??):136–141, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300315011066.

Xu:2017:AEP

[XC17] Aimin Xu and Zhongdi Cen. Asymptotic expansions for the psi function and the Euler–Mascheroni constant. Journal of Number Theory, 180(??):360–372, November 2017. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302007.

Xue:2018:RCL

[XD18] Changfeng Xue and Shaozhong Deng. Recursive computation of logarithmic derivatives, ratios, and products of spheroidal harmonics and modified Bessel functions and applications. Journal of Scientific Computing, 75(1):128–156, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.
1007/s10915-017-0527-3;   https://link.springer.com/
content/pdf/10.1007/s10915-017-0527-3.pdf.

Xu:2016:AEG

[XHT16] Aimin Xu, Yongcai Hu, and Peipei Tang. Asymptotic expansions for the gamma function. Journal of Number Theory, 169(??):134–143, December 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X16301366.

Xu:1994:VPC

[XL94] Guo Liang Xu and Jia Kai Li. Variable precision computation of elementary functions. (Chinese). Journal on Numerical Methods and Computer Applications, 15(3):161–171, 1994. ISSN 1000-3266.

Xiao:2020:PAH

[XLW^{+}20] Feibao Xiao, Feng Liang, Bin Wu, Junzhe Liang, Shuting Cheng, and Guohe Zhang. Posit arithmetic hardware implementations with the minimum cost divider and square root. Electronics, 9 (10):1622:1–1622:16, October 2020. ISSN 2079-9292.

Xu:1983:HPG

[XLX^{+}83] Xian Yu Xu, Jia Kai Li, Gui Jing Xiong, Guo Liang Xu, and Chun Qing Lu. High-precision generation of elementary functions. (Chinese). Appl. Math. Math. Comput., 6:24–32, 1983.

Yoshida:2024:CIB

[YA24] Toshio Yoshida and Yoshinori Adachi. Computation of incomplete beta function ratios I_{x}(a, b) with Deuflhard’s algorithm. Numerical Algorithms, 97(1):373–390, September 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/10.
1007/s11075-023-01707-6.

Yarbrough:1967:PCC

[Yar67] Lynn Yarbrough. Precision calculations of e and π constants. Communications of the ACM, 10(9):537, September 1967. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Yerukala:2011:ACS

[YBR11] R. Yerukala, N. K. Boiroju, and M. K. Reddy. An approximation to the CDF of standard normal distribution. International Journal of Mathematical Archive, 2(7):1077–1079, ????

2011. ISSN 2229-5046. URL http://ijma.info/index.php/
ijma/article/view/393.

Yang:2015:AFG

[YC15] Zhen-Hang Yang and Yu-Ming Chu. Asymptotic formulas for gamma function with applications. Applied Mathematics and Computation, 270(??):665–680, November 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300315011431.

Yang:2015:SBP

[YCZ15] Zhen-Hang Yang, Yu-Ming Chu, and Xiao-Hui Zhang. Sharp bounds for psi function. Applied Mathematics and Computation, 268(??):1055–1063, October 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300315009248.

Ye:2017:SDP

[Ye17] Liangjie Ye. A symbolic decision procedure for relations arising among Taylor coefficients of classical Jacobi theta functions. Journal of Symbolic Computation, 82(??):134–163, September/ October 2017. CODEN JSYCEH. ISSN 0747-7171 (print),

1095-855X (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0747717117300135.

Yeyios:1992:TSA

[Yey92] A. K. Yeyios. On two sequences of algorithms for approximating square roots. Journal of Computational and Applied Mathematics, 40(1):63–72, June 1992. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Yortsos:1989:LSI

[YH89] Y. C. Yortsos and F. J. Hickernell. Linear stability of immiscible displacement in porous media. SIAM Journal on Applied Mathematics, 49(3):730–748, June 1989. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Yousif:1997:BFF

[YM97] Hashim A. Yousif and Richard Melka. Bessel function of the first kind with complex argument. Computer Physics Communications, 106(3):199–206, November 1997. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0010465597000878.

Yousif:2003:CBF

[YM03] Hashim A. Yousif and Richard Melka. Computing Bessel functions of the second kind in extreme parameter regimes. Computer Physics Communications, 151(1):25–34, March 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006975.

Yohe:1970:RBC

[Yoh70] J. M. Yohe. Rigorous bounds on computed approximations to square roots and cube roots. MRC Technical Summary 1088, University of Wisconsin, Madison, 1970.

Yohe:1973:IBS

[Yoh73] J. M. Yohe. Interval bounds for square roots and cube roots. Computing: Archiv fur informatik und numerik, 11(1):51–57, March 1973. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Yong:1970:GBA

[Yon70] Lam Lay Yong. The geometrical basis of the Ancient Chinese square-root method. Isis, 61(1):92–102, Spring 1970. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/229151.

Yoshida:1997:HFM

[Yos97] Masaaki Yoshida. Hypergeometric Functions, My Love: Modular Interpretations of Configuration Spaces, volume E 32 of Aspects of mathematics. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1997. ISBN 3-528-06925-2. ISSN 0179-2156. xvi + 292 pp. LCCN QA353.H9 Y67 1997.

Ypma:1995:HDN

[Ypm95] Tjalling J. Ypma. Historical development of the Newton– Raphson method. SIAM Review, 37(4):531–551, December 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://epubs.siam.org/23425.htm;   http:
//link.aip.org/link/?SIR/37/531/1.

Yun:2009:ACN

[Yun09] Beong In Yun. Approximation to the cumulative normal distribution using hyperbolic tangent based functions. Journal of the Korean Mathematical Society, 46(6):1267–1276, 2009. ISSN 0304-9914.

Yun:2014:AHA

[Yun14] Beong In Yun. An ad hoc approximation to the Gauss error function and a correction method. Applied Mathematical Sciences (Ruse), 8:4261–4273, 2014. ISSN 1312-885X (print), 1314-7552

(electronic). URL http://www.m-hikari.com/ams/ams-2014/
ams-85-88-2014/yunAMS85-88-2014.pdf.

Zaghloul:2011:ACF

[ZA11] Mofreh R. Zaghloul and Ahmed N. Ali. Algorithm 916: Computing the Faddeyeva and Voigt functions. ACM Transactions on Mathematical Software, 38(2):15:1–15:22, December 2011. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [Zag16].

Zafar:2014:HAD

[ZA14] Saad Zafar and Raviteja Adapa. Hardware architecture design and mapping of “Fast Inverse Square Root” algorithm. In 2014 International Conference on Advances in Electrical Engineering (ICAEE), pages 1–4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, January 2014.

Zaghloul:2024:CFI

[ZA24] Mofreh R. Zaghloul and Leen Alrawas. Calculation of Fresnel integrals of real and complex arguments up to 28 significant digits. Numerical Algorithms, 96(2):489–506, June 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/10.
1007/s11075-023-01654-2.

Zaghloul:2016:RAC

[Zag16] Mofreh R. Zaghloul. Remark on “Algorithm 916: Computing the Faddeyeva and Voigt Functions”: Efficiency improvements and Fortran translation. ACM Transactions on Mathematical Software, 42(3):26:1–26:9, May 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [ZA11].

Zaghloul:2017:ASE

[Zag17] Mofreh R. Zaghloul. Algorithm 985: Simple, efficient, and relatively accurate approximation for the evaluation of the Faddeyeva function. ACM Transactions on Mathematical Software, 44(2):22:1–22:9, September 2017. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http:/

/dl.acm.org/citation.cfm?id=3119904.

Zaghloul:2019:RO

[Zag19] Mofreh R. Zaghloul. Remark on ‘Algorithm 680: Evaluation of the Complex Error Function’: Cause and Remedy for the Loss of Accuracy Near the Real Axis. ACM Transactions on Mathematical Software, 45(2):24:1–24:3, April 2019. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/citation.cfm?id=3309681.

Zaghloul:2024:EMP

[Zag24a] Mofreh R. Zaghloul. Efficient multiple-precision computation of the scaled complementary error function and the Dawson integral. Numerical Algorithms, 95(3):1291–1308, March 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265

(electronic). URL https://link.springer.com/article/10.
1007/s11075-023-01608-8.

Zaghloul:2024:ENA

[Zag24b] Mofreh R. Zaghloul. Efficient numerical algorithms for multiprecision and multi-accuracy calculation of the error functions and Dawson integral with complex arguments. Numerical Algorithms, 95(??):1–19, ???? 2024. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Zahar:1994:ACF

[Zah94] R. V. M. (Ramsay Vincent Michael) Zahar, editor. Approximation and computation: a festschrift in honor of Walter Gautschi: proceedings of the Purdue conference, December 2–5, 1993, volume 119 of International series of numerical mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 0-8176-3753-2. LCCN QA221 .A634 1994.

Zaitsev:1993:IMM

[Zai93] A. V. Zaitsev. Implementation of Miller’s method for evaluation of Bessel functions of first kind. Journal of Soviet Mathematics, 63(5):558–560, February 1993. CODEN JSOMAR. ISSN 0090-4104 (print), 2376-5798 (electronic).

Zaker:1969:CCE

[Zak69] T. A. Zaker. Calculation of the complementary error function of complex argument. Journal of Computational Physics, 4(3): 427–430, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Zalik:1989:NIE

[Zal89] R. A. Zalik. A new inequality for entire functions. Journal of Approximation Theory, 58(3):281–283, September 1989. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Zhang:1995:TMAa

[ZB95] J. Zhang and J. A. Belward. Tau-method approximations for the Bessel function Y_{0}(z). Computers and Mathematics with Applications, 30(7):5–14, October 1995. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/089812219500120N.

Zhang:1997:CSA

[ZB97] Jun Zhang and John A. Belward. Chebyshev series approximations for the Bessel function Y_{n}(z) of complex argument. Applied Mathematics and Computation, 88(2–3):275–286, December 30, 1997. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300396003359.

Zill:1970:SEI

[ZC70] D. G. Zill and B. C. Carlson. Symmetric elliptic integrals of the third kind. Mathematics of Computation, 24(109):199–214, January 1970. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

zuCastell:2005:ILO

[zCFF05] Wolfgang zu Castell, Frank Filbir, and Brigitte Forster, editors. Inzell lectures on orthogonal polynomials, Advances in the theory of special functions and orthogonal polynomials. Nova Science, Hauppauge, NY, USA, 2005. ISBN 1-59454-108-6. LCCN QA404.5 .I595 2005. Lectures from the Inzell Summer School on Orthogonal Polynomials, Harmonic Analysis, Approximation, and Applications, held in Inzell, Germany, September 17–21,

Zhang:2015:EAR

[ZCL15] Jianfeng Zhang, Paul Chow, and Hengzhu Liu. An enhanced adaptive recoding rotation CORDIC. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 9(1):4:1–4:??, November 2015. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Zeilberger:1991:MPP

[Zei91] Doron Zeilberger. A Maple program for proving hypergeometric identities. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(3):4–13, July 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Zeng:2004:AMM

[Zen04] Zhonggang Zeng. Algorithm 835: MultRoot—a Matlab package for computing polynomial roots and multiplicities. ACM Transactions on Mathematical Software, 30(2):218–236, June 2004. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Zurawski:1987:DHS

[ZG87] J. H. P. Zurawski and J. B. Gosling. Design of a highspeed square root multiply and divide unit. IEEE Transactions on Computers, C-36(1):13–23, January 1987. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5009445.

Zhang:1995:TMAb

[Zha95] J. Zhang. Tau-method approximations for the Bessel function Y_{1}(z). Computers and Mathematics with Applications, 30(7):15– 19, October 1995. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/089812219500121E.

Zhang:1996:NTM

[Zha96a] J. Zhang. A note on the tau-method approximations for the Bessel functions Y_{0}(z) and Y_{1}(z). Computers and Mathematics with Applications, 31(9):63–70, May 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/0898122196000430.

Zhang:1996:SNC

[Zha96b] Jun Zhang. Symbolic and numerical computation on Bessel functions of complex argument and large magnitude. Journal of Computational and Applied Mathematics, 75(1):99–118, November 12, 1996. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042796000635.

Zhu:2010:JTI

[Zhu10] Ling Zhu. Jordan type inequalities involving the Bessel and modified Bessel functions. Computers and Mathematics with Applications, 59(2):724–736, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0898122109007196.

Ziemer:1989:WDF

[Zie89] William P. Ziemer. Weakly differentiable functions: Sobolev spaces and functions of bounded variation, volume 120 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-97017-7. xvi + 308 pp. LCCN QA323 .Z53 1989.

Zimmermann:1999:KSR

[Zim99] Paul Zimmermann. Karatsuba square root. Research Report 3805, LORIA/INRIA Lorraine, Bâtiment A, Technopôle de Nancy-Brabois, 615 rue du jardin botanique, F-54602 Villers-lès-Nancy Cedex, France, 1999. 8 pp. URL ftp://ftp.inria.fr/

INRIA/publication/publi-pdf/RR/RR-3805.pdf;   ftp://
ftp.inria.fr/INRIA/publication/publi-ps-gz/RR/RR-3805.
ps.gz;   http://www.inria.fr/rrrt/rr-3805.html.

Zimmermann:2000:PGF

[Zim00] Paul Zimmermann. A proof of GMP fast division and square root implementations. Technical report, LORIA/INRIA Lorraine, Bâtiment A, Technopôle de Nancy-Brabois, 615 rue du jardin botanique, F-54602 Villers-lès-Nancy Cedex, France, September 2000. 14 pp. URL http://www.loria.fr/~

zimmerma/papers/proof-div-sqrt.ps.gz.

Zimmermann:2005:XXX

[Zim05] Paul Zimmermann. 5,341,321. Technical report, LORIA/INRIA Lorraine, Bâtiment A, Technopôle de Nancy-Brabois, 615 rue du jardin botanique, F-54602 Villers-lès-Nancy Cedex, France,

June 8, 2005. 2 pp. URL http://www.loria.fr/~zimmerma/
papers/5341321.ps.gz.

Ziv:1991:FEE

[Ziv91] Abraham Ziv. Fast evaluation of elementary mathematical functions with correctly rounded last bit. ACM Transactions on Mathematical Software, 17(3):410–423, September 1991. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
toms/116813.html.

Ziv:1999:SUR

[Ziv99] Abraham Ziv. Sharp ULP rounding error bound for the hypotenuse function. Mathematics of Computation, 68(227):1143– 1148, July 1999. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL http://www.ams.org/jourcgi/
jour-pbprocess?fn=110&arg1=S0025-5718-99-01103-5&u=/
mcom/1999-68-227/.

Zhang:1996:CSF

[ZJ96] Shanjie Zhang and Jianming Jin. Computation of Special Functions. Wiley, New York, NY, USA, 1996. ISBN 0-471-11963-6. xxvi + 717 pp. LCCN QA351.C45 1996. US$94.00.

Zhong:2013:AKF

[ZLA13] Min Zhong, R. J. Loy, and R. S. Anderssen. Approximating the Kohlrausch function by sums of exponentials. The ANZIAM Journal, 54(4):306–323, April 2013. CO-DEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/approximating-the-kohlrausch-
function-by-sums-of-exponentials/1F2BD299466198D202D9D2355E34116F.

Zhu:2004:ISR

[ZLC04] Hufei Zhu, Zhongding Lei, and F. P. S. Chin. An improved square-root algorithm for BLAST. IEEE Signal Processing Letters, 11(9):772–775, September 2004. CODEN ISPLEM. ISSN 1070-9908 (print), 1558-2361 (electronic).

Ziv:2001:APM

[ZOHR01] Abraham Ziv, Moshe Olshansky, Ealan Henis, and Anna Reitman. Accurate portable mathematical library (IBM APMath-Lib). World-Wide Web document, December 20, 2001. URL

ftp://www-126.ibm.com/pub/mathlib/mathlib12.20.2001.
tar.gz;   http://oss.software.ibm.com/mathlib/.

Zelen:1964:PF

[ZS64] Marvin Zelen and Norman C. Severo. Probability functions. In Abramowitz and Stegun [AS64], pages 925–996. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also avail-

able online at http://www.convertit.com/Go/ConvertIt/

Reference/AMS55.ASP in bitmap image format.

Zhu:2008:SNR

[ZS08] Ling Zhu and Jinju Sun. Six new Redheffer-type inequalities for circular and hyperbolic functions. Computers and Mathematics with Applications, 56(2):522–529, July 2008. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0898122108000813.

Zaritskaya:1986:ACE

[ZSA86] Z. V. Zaritskaya, A. I. Shvaĭ, and P. Ē. Antonyuk. Approximation of certain elementary functions in the metric L. (Russian). Vestnik L’vov. Politekhn. Inst., 202:38–40, 1986.

Zucker:1964:ETF

[Zuc64] Ruth Zucker. Elementary transcendental functions. Logarithmic, exponential, circular and hyperbolic functions. In Abramowitz and Stegun [AS64], pages 65–226. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also avail-

able online at http://www.convertit.com/Go/ConvertIt/

Reference/AMS55.ASP in bitmap image format.

Zwick:1987:BAC

[Zwi87] D. Zwick. Best approximation by convex functions. American Mathematical Monthly, 94(6):528–534, July 1987. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zwillinger:2003:CSM

[Zwi03] Daniel Zwillinger, editor. CRC Standard Mathematical Tables and Formulae. Chapman and Hall/CRC, Boca Raton, FL, USA, 31st edition, 2003. ISBN 1-58488-291-3 (hardcover), 1-4200-3534-7 (e-book). xiv + 910 pp. LCCN QA47 .M315 2003.

Zheng:2001:ARE

[ZXBZH01] Liang Zheng, Shen Xu-Bang, and Peng Zuo-Hui. The application of redundant encoding in iterative implementation of division and square root. In Proceedings of the 4th International Conference on ASIC, pages 603–606. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. CODEN ???? ISSN ????

Zahle:1996:FDW

[ZZ96] M. Zähle and H. Ziezold. Fractional derivatives of Weierstrasstype functions. Journal of Computational and Applied Mathematics, 76(1–2):265–275, December 17, 1996. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0377042796001100.

Zhou:2011:ADZ

[ZZ11] Jian-Rong Zhou and Yu-Qiu Zhao. Asymptotic distributions of the zeros of certain classes of Gauss hypergeometric polynomials. Applied Mathematics and Computation, 218(3):1153–1159, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300311007892. Special Issue in

Honour of Hari M. Srivastava on his 70th birth anniversary.